A Complete Bibliography of Internet RFC Documents

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 585 1640, +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org, beebe@ieee.org (Internet)
                    WWW URL: http://www.math.utah.edu/~beebe/

16 March 2017 Version 3.29

Title word cross-reference

#1 [65, 2300, 2424]. #10 [2301]. #33 [47]. #36 [38, 39]. #5 [296]. #7 [2302].

$500 [420].

/Internet [2149]. /ISO [1128, 1138, 1317]. /Multi [1455].

0 [396]. 0/9 [509].

1 [1, 291, 832, 1805, 2244]. 1.0 [1932, 2111]. 1.1 [2055]. 10 [10, 798]. 100 [100]. 1000 [990]. 1001 [991]. 1002 [992]. 10021 [1128, 1138, 1317]. 1003 [993]. 1004 [994]. 1005 [995]. 1006 [996]. 1007 [997]. 1008 [998]. 1009 [999]. 101 [101]. 1010 [1000]. 1011 [1001]. 1012 [1002]. 1013 [1003]. 1014 [1004]. 1015 [1005]. 1016 [1006]. 1017 [1007]. 1018 [1008]. 1019 [1009]. 102 [102]. 1020 [1010]. 1021 [1011]. 1022 [1012]. 1023 [1013]. 1024 [1014]. 1025 [1015]. 1026 [1016]. 1027 [1017]. 1028 [1018]. 1029 [1019]. 103 [103]. 1030 [1020]. 1031 [1021]. 1032 [1022]. 1033 [1023]. 1034 [1024]. 1035 [1025]. 1036 [1026]. 1037 [1027]. 1038 [1028]. 1039 [1029]. 104 [104]. 1040 [1030]. 1041 [1031]. 1042 [1032]. 1043 [1033]. 1044 [1034]. 1045 [1035]. 1046 [1036]. 1047 [1037]. 1048 [1038]. 1049 [1039]. 105 [105]. 1050 [1040]. 1051 [1041]. 1052 [1042]. 1053 [1043]. 1054 [1044]. 1055 [1045]. 1056 [1046]. 1057 [1047]. 1058 [1048]. 1059 [1049]. 106 [106]. 1060 [1050]. 1061 [1051]. 1062 [1052]. 1063 [1053]. 1064 [1054]. 10646 [1805, 2031]. 1065 [1055]. 1066 [1056]. 1067 [1057]. 1068 [1058]. 1069 [1059]. 107 [107]. 1070 [1060]. 1071 [1061]. 1072 [1062]. 1073 [1063]. 1074 [1064]. 1075 [1065]. 1076 [1066]. 1077 [1067]. 1078 [1068]. 1079 [1069]. 108 [108]. 1080 [1070]. 1081 [1071]. 1082 [1072]. 1083 [1073]. 1084 [1074]. 1085 [1075]. 1086 [1076]. 1087 [1077]. 1088 [1078]. 1089 [1079]. 109 [109]. 1090 [1080]. 1091 [1081]. 1092 [1082]. 1093 [1083]. 1094 [1084]. 1095 [1085]. 1096 [1086]. 1097 [1087]. 1098 [1088]. 1099 [1089]. 11 [11, 1395, 2149]. 110 [110]. 1100 [1090]. 1101 [1091]. 1102 [1092]. 1103 [1093]. 1104 [1094]. 1105 [1095]. 1106 [1096]. 1107 [1097]. 1108 [1098]. 1109 [1099]. 111 [111]. 1110 [1100]. 1111 [1101]. 1112 [1102]. 1113 [1103]. 1114 [1104]. 1115 [1105]. 1116 [1106]. 1117 [1107]. 1118 [1108]. 1119 [1109]. 112 [112]. 1120 [1110]. 1121 [1111]. 1122 [1112]. 1123 [1113]. 1124 [1114]. 1125 [1115]. 1126 [1116]. 1127 [1117]. 1128 [1118]. 1129 [1119]. 113 [113]. 1130 [1120]. 1131 [1121]. 1132 [1122]. 1133 [1123]. 1134 [1124]. 1135 [1125]. 1136 [1126]. 1137 [1127]. 1138 [1128]. 1139 [1129]. 114 [114]. 1140 [1130]. 1141 [1131]. 1142 [1132]. 1143 [1133]. 1144 [1134]. 1145 [1135]. 1146 [1136]. 1147 [1137]. 1148 [1138]. 1149 [1139]. 115 [115]. 1150 [1140]. 1151 [1141]. 1152 [1142]. 1153 [1143]. 1154 [1144]. 1155 [1145]. 1156 [1146]. 1157 [1147]. 1158 [1148]. 1159 [1149]. 116 [116]. 1160 [1150]. 1161 [1151]. 1162 [1152]. 1163 [1153]. 1164 [1154]. 1165 [1155]. 1166 [1156]. 1167 [1157]. 1168 [1158]. 1169 [1159]. 117 [117]. 1170 [1160]. 1171 [1161]. 1172 [1162]. 1173 [1163]. 1174 [1164]. 1175 [1165]. 1176 [1166]. 1177 [1167]. 1178 [1168]. 1179 [1169]. 118 [118]. 1180 [1170]. 1181 [1171]. 1182 [1172]. 1183 [1173]. 1184 [1174]. 1185 [1175]. 1186 [1176]. 1187 [1177]. 1188 [1178]. 1189 [1179]. 119 [119]. 1190 [1180]. 1191 [1181]. 1192 [1182]. 1193 [1183]. 1194 [1184]. 1195 [1185]. 1196 [1186]. 1197 [1187]. 1198 [1188]. 1199 [1189]. 12 [12, 1699]. 12-Days [1869]. 120 [120]. 1200 [1190]. 1201 [1191]. 1202 [1192]. 1203 [1193]. 1204 [1194]. 1205 [1195]. 1206 [1196]. 1207 [1197]. 1208 [1198]. 1209 [1199]. 121 [121]. 1210 [1200]. 1211 [1201]. 1212 [1202]. 1213 [1203]. 1214 [1204]. 1215 [1205]. 1216 [1206]. 1217 [1207]. 1218 [1208]. 1219 [1209]. 122 [122]. 1220 [1210]. 1221 [1211]. 1222 [1212]. 1223 [1213]. 1224 [1214]. 1225 [1215]. 1226 [1216]. 1227 [1217]. 1228 [1218]. 1229 [1219]. 123 [123]. 1230 [1220]. 1231 [1221]. 1232 [1222]. 1233 [1223]. 1234 [1224]. 1235 [1225]. 1236 [1226]. 1237 [1227]. 1238 [1228]. 1239 [1229]. 124 [124]. 1240 [1230]. 1241 [1231]. 1242 [1232]. 1243 [1233]. 1244 [1234]. 1245 [1235]. 1246 [1236]. 1247 [1237]. 1248 [1238]. 1249 [1239]. 125 [125]. 1250 [1240]. 1251 [1241].

1252 [1242]. 1253 [1243]. 1254 [1244]. 1255 [1245]. 1256 [1246]. 1257 [1247]. 1258 [1248]. 1259 [1249]. 126 [126]. 1260 [1250]. 1261 [1251]. 1262 [1252]. 1263 [1253]. 1264 [1254]. 1265 [1255]. 1266 [1256]. 1267 [1257]. 1268 [1258]. 1269 [1259]. 127 [127]. 1270 [1260]. 1271 [1261]. 1272 [1262]. 1273 [1263]. 1274 [1264]. 1275 [1265]. 1276 [1266]. 1277 [1267]. 1278 [1268]. 1279 [1269]. 128 [128]. 128-bit [1741]. 1280 [1270]. 1281 [1271]. 1282 [1272]. 1283 [1273]. 1284 [1274]. 1285 [1275]. 1286 [1276]. 1287 [1277]. 1288 [1278]. 1289 [1279]. 129 [129]. 1290 [1280]. 1291 [1281]. 1292 [1282]. 1293 [1283]. 1294 [1284]. 1295 [1285]. 1296 [1286]. 1297 [1287]. 1298 [1288]. 1299 [1289]. 13 [13]. 130 [130]. 1300 [1290]. 1301 [1291]. 1302 [1292]. 1303 [1293]. 1304 [1294]. 1305 [1295]. 1306 [1296]. 1307 [1297]. 1308 [1298]. 1309 [1299]. 131 [131]. 1310 [1300]. 1311 [1301]. 1312 [1302]. 1313 [1303]. 1314 [1304]. 1315 [1305]. 1316 [1306]. 1317 [1307]. 1318 [1308]. 1319 [1309]. 132 [132]. 1320 [1310]. 1321 [1311]. 1322 [1312]. 1323 [1313]. 1324 [1314]. 1325 [1315]. 1326 [1316]. 1327 [1317]. 1328 [1318]. 1329 [1319]. 133 [133]. 1330 [1320]. 1331 [1321]. 1332 [1322]. 1333 [1323]. 1334 [1324]. 1335 [1325]. 1336 [1326]. 1337 [1327]. 1338 [1328]. 1339 [1329]. 134 [134]. 1340 [1330]. 1341 [1331]. 1342 [1332]. 1343 [1333]. 1344 [1334]. 1345 [1335]. 1346 [1336]. 1347 [1337]. 1348 [1338]. 1349 [1339]. 135 [135]. 1350 [1340]. 1351 [1341]. 1352 [1342]. 1353 [1343]. 1354 [1344]. 1355 [1345]. 1356 [1346]. 1357 [1347]. 1358 [1348]. 1359 [1349]. 136 [136]. 1360 [1350]. 1361 [1351]. 1362 [1352]. 1363 [1353]. 1364 [1354]. 1365 [1355]. 1366 [1356]. 1367 [1357]. 1368 [1358]. 1369 [1359]. 137 [137]. 1370 [1360]. 1371 [1361]. 1372 [1362]. 1373 [1363]. 1374 [1364]. 1375 [1365]. 1376 [1366]. 1377 [1367]. 1378 [1368]. 1379 [1369]. 138 [138]. 1380 [1370]. 1381 [1371]. 1382 [1372]. 1383 [1373]. 1384 [1374]. 1385 [1375]. 1386 [1376]. 1387 [1377]. 1388 [1378]. 1389 [1379]. 139 [139]. 1390 [1380]. 1391 [1381]. 1392 [1382]. 1393 [1383]. 1394 [1384]. 1395 [1385]. 1396 [1386]. 1397 [1387]. 1398 [1388]. 1399 [1389]. 14 [14]. 140 [140]. 1400 [1390]. 1401 [1391]. 1402 [1392]. 1403 [1393]. 1404 [1394]. 1405 [1395]. 1406 [1396]. 1407 [1397]. 1408 [1398]. 1409 [1399]. 141 [141]. 1410 [1400]. 1411 [1401]. 1412 [1402]. 1413 [1403]. 1414 [1404]. 1415 [1405]. 1416 [1406]. 1417 [1407]. 1418 [1408]. 1419 [1409]. 142 [142]. 1420 [1410]. 1421 [1411]. 1422 [1412]. 1423 [1413]. 1424 [1414]. 1425 [1415]. 1426 [1416]. 1427 [1417]. 1428 [1418]. 1429 [1419]. 143 [143]. 1430 [1420]. 1431 [1421]. 1432 [1422]. 1433 [1423]. 1434 [1424]. 1435 [1425]. 1436 [1426]. 1437 [1427]. 1438 [1428]. 1439 [1429]. 144 [144]. 1440 [1430]. 1441 [1431]. 1442 [1432]. 1443 [1433]. 1444 [1434]. 1445 [1435]. 1446 [1436]. 1447 [1437]. 1448 [1438]. 1449 [1439]. 145 [145]. 1450 [1440]. 1451 [1441]. 1452 [1442]. 1453 [1443]. 1454 [1444]. 1455 [1445]. 1456 [1446]. 1457 [1447]. 1458 [1448]. 1459 [1449]. 146 [146]. 1460 [1450]. 1461 [1451]. 1462 [1452]. 1463 [1453]. 1464 [1454]. 1465 [1455]. 1466 [1456]. 1467 [1457]. 1468 [1458]. 1469 [1459]. 147 [147]. 1470 [1460]. 1471 [1461]. 1472 [1462]. 1473 [1463]. 1474 [1464]. 1475 [1465]. 1476 [1466]. 1477 [1467]. 1478 [1468]. 1479 [1469]. 148 [148]. 1480 [1470]. 1481 [1471]. 1482 [1472]. 1483 [1473]. 1484 [1474]. 1485 [1475]. 1486 [1476]. 1487 [1477]. 1488 [1478]. 1489 [1479]. 149 [149]. 1490 [1480]. 1491 [1481]. 1492 [1482]. 1493 [1483]. 1494 [1484]. 1495 [1485]. 1496 [1486]. 1497 [1487]. 1498 [1488]. 1499 [1489]. 15 [15, 542, 835]. 150 [150]. 1500 [1490]. 1501 [1491]. 1502 [1492]. 1503 [1493]. 1504 [1494]. 1505 [1495]. 1506 [1496]. 1507 [1497]. 1508 [1498]. 1509 [1499]. 151 [151]. 1510 [1500]. 1511 [1501]. 1512 [1502]. 1513 [1503]. 1514 [1504]. 1515 [1505]. 1516 [1506]. 1517 [1507]. 1518 [1508]. 1519 [1509]. 152 [152]. 1520 [1510]. 1521 [1511]. 1522 [1512]. 1523 [1513]. 1524 [1514]. 1525 [1515]. 1526 [1516]. 1527 [1517]. 1528 [1518]. 1529 [1519]. 153 [153]. 1530 [1520]. 1531 [1521]. 1532 [1522]. 1533 [1523]. 1534 [1524]. 1535 [1525]. 1536 [1526]. 1537 [1527]. 1538 [1528]. 1539 [1529]. 154 [154]. 1540 [1530]. 1541 [1531]. 1542 [1532]. 1543 [1533]. 1544 [1534]. 1545 [1535]. 1546 [1536]. 1547 [1537]. 1548 [1538]. 1549 [1539]. 155 [155]. 1550 [1540]. 1551 [1541]. 1552 [1542]. 1553 [1543]. 1554 [1544]. 1555 [1545]. 1556 [1546]. 1557 [1547]. 1558 [1548]. 1559 [1549]. 156 [156]. 1560 [1550]. 1561 [1551]. 1562 [1552]. 1563 [1553]. 1564 [1554]. 1565 [1555]. 1566 [1556]. 1567 [1557]. 1568 [1558]. 1569 [1559]. 157 [157]. 1570 [1560]. 1571 [1561]. 1572 [1562]. 1573 [1563]. 1574 [1564]. 1575 [1565]. 1576 [1566]. 1577 [1567]. 1578 [1568]. 1579 [1569]. 158 [158]. 1580 [1570]. 1581 [1571]. 1582 [1572]. 1583 [1573]. 1584 [1574]. 1585 [1575]. 1586 [1576]. 1587 [1577]. 1588 [1578]. 1589 [1579]. 159 [159]. 1590 [1580]. 1591 [1581]. 1592 [1582]. 1593 [1583]. 1594 [1584]. 1595 [1585]. 1596 [1586]. 1597 [1587]. 1598 [1588]. 1599 [1589]. 16 [16]. 160 [160]. 1600 [1590]. 1601 [1591]. 1602 [1592]. 1603 [1593]. 1604 [1594]. 1605 [1595]. 1606 [1596]. 1607 [1597]. 1608 [1598]. 1609 [1599]. 161 [161]. 1610 [1600]. 1611 [1601]. 1612 [1602]. 1613 [1603]. 1614 [1604]. 1615 [1605]. 1616 [1606]. 1617 [1607]. 1618 [1608]. 1619 [1609]. 162 [162]. 1620 [1610]. 1621 [1611]. 1622 [1612]. 1623 [1613]. 1624 [1614]. 1625 [1615]. 1626 [1616]. 1627 [1617]. 1628 [1618]. 1629 [1619]. 163 [163]. 1630 [1620]. 1631 [1621]. 1632 [1622]. 1633 [1623]. 1634 [1624]. 1635 [1625]. 1636 [1626]. 1637 [1627]. 1638 [1628]. 1639 [1629]. 164 [164]. 1640 [1630]. 1641 [1631]. 1642 [1632]. 1643 [1633]. 1644 [1634]. 1645 [1635]. 1646 [1636]. 1647 [1637]. 1648 [1638]. 1649 [1639]. 165 [165]. 1650 [1640]. 1651 [1641]. 1652 [1642]. 1653 [1643]. 1654 [1644]. 1655 [1645]. 1656 [1646]. 1657 [1647]. 1658 [1648]. 1659 [1649]. 166 [166]. 1660 [1650]. 1661 [1651]. 1662 [1652]. 1663 [1653]. 1664 [1654]. 1665 [1655]. 1666 [1656]. 1667 [1657]. 1668 [1658]. 1669 [1659]. 167 [167]. 1670 [1660]. 1671 [1661]. 1672 [1662]. 1673 [1663]. 1674 [1664]. 1675 [1665]. 1676 [1666]. 1677 [1667]. 1678 [1668]. 1679 [1669]. 168 [168]. 1680 [1670]. 1681 [1671]. 1682 [1672]. 1683 [1673]. 1684 [1674]. 1685 [1675]. 1686 [1676]. 1687 [1677]. 1688 [1678]. 1689 [1679]. 169 [169]. 1690 [1680]. 1691 [1681]. 1692 [1682]. 1693 [1683]. 1694 [1684]. 1695 [1685]. 1696 [1686]. 1697 [1687]. 1698 [1688]. 1699 [1689]. 17 [17, 531]. 170 [170]. 1700 [1690]. 1701 [1691]. 1702 [1692]. 1703 [1693]. 1704 [1694]. 1705 [1695]. 1706 [1696]. 1707 [1697]. 1708 [1698].

1709 [1699]. 171 [171]. 1710 [1700]. 1711 [1701]. 1712 [1702]. 1713 [1703]. 1714 [1704]. 1715 [1705]. 1716 [1706]. 1717 [1707]. 1718 [1708]. 1719 [1709]. 172 [172]. 1720 [1710]. 1721 [1711]. 1722 [1712]. 1723 [1713]. 1724 [1714]. 1725 [1715]. 1726 [1716]. 1727 [1717]. 1728 [1718]. 1729 [1719]. 173 [173]. 1730 [1720]. 1731 [1721]. 1732 [1722]. 1733 [1723]. 1734 [1724]. 1735 [1725]. 1736 [1726]. 1737 [1727]. 1738 [1728]. 1739 [1729]. 174 [174]. 1740 [1730]. 1741 [1731]. 1742 [1732]. 1743 [1733]. 1744 [1734]. 1745 [1735]. 1746 [1736]. 1747 [1737]. 1748 [1738]. 1749 [1739]. 175 [175]. 1750 [1740]. 1751 [1741]. 1752 [1742]. 1753 [1743]. 1754 [1744]. 1755 [1745]. 1756 [1746]. 1757 [1747]. 1758 [1748]. 1759 [1749]. 176 [176]. 1760 [1750]. 1761 [1751]. 1762 [1752]. 1763 [1753]. 1764 [1754]. 1765 [1755]. 1766 [1756]. 1767 [1757]. 1768 [1758]. 1769 [1759]. 177 [177]. 1770 [1760]. 1771 [1761]. 1772 [1762]. 1773 [1763]. 1774 [1764]. 1775 [1765]. 1776 [1766]. 1777 [1767]. 1778 [1768]. 1779 [1769]. 178 [178]. 1780 [1770]. 1781 [1771]. 1782 [1772]. 1783 [1773]. 1784 [1774]. 1785 [1775]. 1786 [1776]. 1787 [1777]. 1788 [1778]. 1789 [1779]. 179 [179]. 1790 [1780]. 1791 [1781]. 1792 [1782]. 1793 [1783]. 1794 [1784]. 1795 [1785]. 1796 [1786]. 1797 [1787]. 1798 [1788]. 1799 [1789]. 18 [18, 834, 2502]. 180 [180]. 1800 [1790]. 1801 [1791]. 1802 [1792]. 1803 [1793]. 1804 [1794]. 1805 [1795]. 1806 [1796]. 1807 [1797]. 1808 [1798]. 1809 [1799]. 181 [181]. 1810 [1800]. 1811 [1801]. 1812 [1802]. 1813 [1803]. 1814 [1804]. 1815 [1805]. 1816 [1806]. 1817 [1807]. 1818 [1808]. 1819 [1809]. 182 [182]. 1820 [1810]. 1821 [1811]. 1822 [1812]. 1822L [792, 841, 868]. 1823 [1813]. 1824 [1814]. 1825 [1815]. 1826 [1816]. 1827 [1817]. 1828 [1818]. 1829 [1819]. 183 [183]. 1830 [1820]. 1831 [1821]. 1832 [1822]. 1833 [1823]. 1834 [1824]. 1835 [1825]. 1836 [1826]. 1837 [1827]. 1838 [1828]. 184 [184]. 1840 [1882]. 1841 [1829]. 1842 [1830]. 1843 [1831]. 1844 [1832]. 1845 [1833]. 1846 [1834]. 1847 [1835]. 1848 [1836]. 185 [185]. 1850 [1837]. 1851 [1838]. 1852 [1839]. 1853 [1840]. 1854 [1841]. 1855 [1842]. 1856 [1843]. 1857 [1844]. 1858 [1845]. 1859 [1846]. 186 [186]. 1860 [1847]. 1861 [1848]. 1862 [1849]. 1863 [1850]. 1864 [1851]. 1865 [1852]. 1866 [1853]. 1867 [1854]. 1868 [1855]. 1869 [1856]. 187 [187]. 1870 [1857]. 1871 [1858]. 1872 [1859]. 1873 [1860]. 1874 [1861]. 1875 [1862]. 1876 [1863]. 1877 [1864]. 1878 [1865]. 1879 [1866]. 188 [188]. 1880 [1867]. 1881 [1868]. 1882 [1869]. 1883 [1870]. 1884 [1871]. 1885 [1872]. 1886 [1873]. 1887 [1874]. 1888 [1875]. 1889 [1876]. 189 [189]. 1890 [1877]. 1891 [1878]. 1892 [1879]. 1893 [1880]. 1894 [1881]. 1895 [1882]. 1896 [1883]. 1897 [1884]. 1898 [1885]. 1899 [1886]. 19 [19]. 190 [190]. 1900 [1887]. 1901 [1888]. 1902 [1889]. 1903 [1890]. 1904 [1891]. 1905 [1892]. 1906 [1893]. 1907 [1894]. 1908 [1895]. 1909 [1896]. 191 [191]. 1910 [1897]. 1911 [1898]. 1912 [1899]. 1913 [1900]. 1914 [1901]. 1915 [1902]. 1916 [1903]. 1917 [1904]. 1918 [1905]. 1919 [1906]. 192 [192]. 1920 [1907]. 1921 [1908]. 1922 [1909]. 1923 [1910]. 1924 [1911]. 1925 [1912]. 1926 [1913]. 1927 [1914]. 1928 [1915]. 1929 [1916]. 193 [193]. 1930 [1917]. 1931 [1918]. 1932 [1919]. 1933 [1920]. 1934 [1921]. 1935 [1922]. 1936 [1923]. 1937 [1924]. 1938 [1925]. 1939 [1926]. 194 [194]. 1940 [1927]. 1941 [1928]. 1942 [1929]. 1943 [1930]. 1944 [1931]. 1945 [1932]. 1946 [1933]. 1947 [1934]. 1948 [1935]. 1949 [1936]. 195 [195]. 1950 [1937]. 1951 [1938]. 1952 [1939]. 1953 [1940]. 1954 [1941]. 1955 [1942]. 1956 [1943]. 1957 [1944]. 1958 [1945]. 1959 [1946]. 196 [196]. 1960 [1947]. 1961 [1948]. 1962 [1949]. 1963 [1950]. 1964 [1951]. 1965 [1952]. 1966 [1953]. 1967 [1954]. 1968 [1955]. 1969 [1956]. 197 [197]. 1970 [1957]. 1971 [101, 108, 290, 296, 1958]. 1972 [1959]. 1973 [539, 542, 591, 1960]. 1974 [1961]. 1975 [1962]. 1976 [1963]. 1977 [1964]. 1978 [1965]. 1979 [798, 1966]. 198 [198]. 1980 [1967]. 1981 [1968]. 1982 [1969]. 1983 [834, 835, 836, 1970]. 1984 [1971]. 1984/1988 [1395]. 1985 [1972]. 1986 [1973]. 1987 [1974]. 1988 [1615, 1975]. 1989 [1976]. 199 [199]. 1990 [1200, 1977]. 1991 [1978]. 1992 [1979]. 1993 [1980]. 1994 [1626, 1849, 1981]. 1995 [1814, 1982]. 1995/4 [1814]. 1996 [1983, 2117]. 1997 [1984]. 1998 [1985, 2428]. 1999 [1986].

2 [2, 1491, 1544, 2069]. 20 [20]. 200 [200]. 2000 [1987]. 2001 [1988]. 2002 [1989]. 2003 [1990]. 2004 [1991]. 2005 [1992]. 2006 [1993]. 2007 [1994]. 2008 [1995]. 2009 [1996]. 201 [201]. 2010 [1997]. 2011 [1998]. 2012 [1999]. 2013 [2000]. 2014 [2001]. 2015 [2002]. 2016 [2003]. 2017 [2004]. 2018 [2005]. 2019 [2006]. 202 [202]. 2020 [2007]. 2021 [2008]. 2022 [2009]. 2023 [2010]. 2024 [2011]. 2025 [2012]. 2026 [2013]. 2027 [2014]. 2028 [2015]. 2029 [2016]. 203 [203]. 2030 [2017]. 2031 [2018]. 2032 [2019]. 2033 [2020]. 2034 [2021]. 2035 [2022]. 2036 [2023]. 2037 [2024]. 2038 [2025]. 2039 [2026]. 204 [204]. 2040 [2027]. 2041 [2028]. 2042 [2029]. 2043 [2030]. 2044 [2031]. 2045 [2032]. 2046 [2033]. 2047 [2034]. 2048 [2035]. 2049 [2036]. 205 [205]. 2050 [2037]. 2051 [2038]. 2052 [2039]. 2053 [2040]. 2054 [2041]. 2055 [2042]. 2056 [2043]. 2057 [2044]. 2058 [2045]. 2059 [2046]. 206 [206]. 2060 [2047]. 2061 [2048]. 2062 [2049]. 2063 [2050]. 2064 [2051]. 2065 [2052]. 2066 [2053]. 2067 [2054]. 2068 [2055]. 2069 [2056]. 207 [207]. 2070 [2057]. 2071 [2058]. 2072 [2059]. 2073 [2060]. 2074 [2061]. 2075 [2062]. 2076 [2063]. 2077 [2064]. 2078 [2065]. 2079 [2066]. 208 [208]. 2080 [2067]. 2081 [2068]. 2082 [2069]. 2083 [2070]. 2084 [2071]. 2085 [2072]. 2086 [2073]. 2087 [2074]. 2088 [2075]. 2089 [2076]. 209 [209]. 2090 [2077]. 2091 [2078]. 2092 [2079]. 2093 [2080]. 2094 [2081]. 2095 [2082]. 2096 [2083]. 2097 [2084]. 2098 [2085]. 2099 [2086]. 21 [21]. 210 [210]. 2100 [2087]. 2101 [2088]. 2102 [2089]. 2103 [2090]. 2104 [2091]. 2105 [2092]. 2106 [2093]. 2107 [2094]. 2108 [2095]. 2109 [2096]. 211 [211]. 2110 [2097]. 2111 [2098]. 2112 [2099]. 2113 [2100]. 2114 [2101]. 2115 [2102]. 2116 [2103]. 2117 [2104]. 2118 [2105]. 2119 [2106]. 212 [212]. 2120 [2107]. 2121 [2108]. 2122 [2109]. 2123 [2110]. 2124 [2111]. 2125 [2112]. 2126 [2113]. 2127 [2114]. 2128 [2115]. 2129 [2116]. 213 [213]. 2130 [2117]. 2131 [2118]. 2132 [2119]. 2133 [2120]. 2134 [2121]. 2135 [2122]. 2136 [2123]. 2137 [2124]. 2138 [2125]. 2139 [2126]. 214 [214]. 2140 [2127]. 2141 [2128]. 2142 [2129]. 2143 [2130]. 2144 [2131]. 2145 [2132]. 2146 [2133]. 2147 [2134]. 2148 [2135].

2149 [2136]. 215 [215]. 2150 [2137]. 2151 [2138]. 2152 [2139]. 2153 [2140]. 2154 [2141]. 2155 [2142]. 2156 [2143]. 2157 [2144]. 2158 [2145]. 2159 [2146]. 216 [216]. 2160 [2147]. 2161 [2148]. 2162 [2149]. 2163 [2150]. 2164 [2151]. 2165 [2152]. 2166 [2153]. 2167 [2154]. 2168 [2155]. 2169 [2156]. 217 [217]. 2170 [2157]. 2171 [2158]. 2172 [2159]. 2173 [2160]. 2174 [2161]. 2175 [2162]. 2176 [2163]. 2177 [2164]. 2178 [2165]. 2179 [2166]. 218 [218]. 2180 [2167]. 2181 [2168]. 2182 [2169]. 2183 [2170]. 2184 [2171]. 2185 [2172]. 2186 [2173]. 2187 [2174]. 2188 [2175]. 2189 [2176]. 219 [219]. 2190 [2177]. 2191 [2178]. 2192 [2179]. 2193 [2180]. 2194 [2181]. 2195 [2182]. 2196 [2183]. 2197 [2184]. 2198 [2185]. 2199 [2186]. 21ST [1597]. 22 [22, 836]. 220 [220]. 2200 [2187]. 2201 [2188]. 2202 [2189]. 2203 [2190]. 2204 [2191]. 2205 [2192]. 2206 [2193]. 2207 [2194]. 2208 [2195]. 2209 [2196]. 221 [221]. 2210 [2197]. 2211 [2198]. 2212 [2199]. 2213 [2200]. 2214 [2201]. 2215 [2202]. 2216 [2203]. 2217 [2204]. 2218 [2205]. 2219 [2206]. 222 [222]. 2220 [2207]. 2221 [2208]. 2222 [2209]. 2223 [2210]. 2224 [2211]. 2225 [2212]. 2226 [2213]. 2227 [2214]. 2228 [2215]. 2229 [2216]. 223 [223]. 2230 [2217]. 2231 [2218]. 2232 [2219]. 2233 [2220]. 2234 [2221]. 2235 [2222]. 2236 [2223]. 2237 [2224]. 2238 [2225]. 2239 [2226]. 224 [224]. 2240 [2227]. 2241 [2228]. 2242 [2229]. 2243 [2230]. 2244 [2231]. 2245 [2232]. 2246 [2233]. 2247 [2234]. 2248 [2235]. 2249 [2236]. 225 [225]. 2250 [2237]. 2251 [2238]. 2252 [2239]. 2253 [2240]. 2254 [2241]. 2255 [2242]. 2256 [2243]. 2257 [2244]. 2258 [2245]. 2259 [2246]. 226 [226]. 2260 [2247]. 2261 [2248]. 2262 [2249]. 2263 [2250]. 2264 [2251]. 2265 [2252]. 2266 [2253]. 2267 [2254]. 2268 [2255]. 2269 [2256]. 227 [227]. 2270 [2257]. 2271 [2258]. 2272 [2259]. 2273 [2260]. 2274 [2261]. 2275 [2262]. 2276 [2263]. 2277 [2264]. 2278 [2265]. 2279 [2266]. 228 [228]. 2280 [2267]. 2281 [2268]. 2282 [2269]. 2283 [2270]. 2284 [2271]. 2285 [2272]. 2286 [2273]. 2287 [2274]. 2288 [2275]. 2289 [2276]. 229 [229]. 2290 [2277]. 2291 [2278]. 2292 [2279]. 2293 [2280]. 2294 [2281]. 2295 [2282]. 2296 [2283]. 2297 [2284]. 2298 [2285]. 2299 [2286]. 23 [23]. 230 [230]. 2300 [2287]. 2301 [2288]. 2302 [2289]. 2303 [2290]. 2304 [2291]. 2305 [2292]. 2306 [2293]. 2307 [2294]. 2308 [2295]. 2309 [2296]. 231 [231]. 2310 [2297]. 2311 [2298]. 2312 [2299]. 2313 [2300]. 2314 [2301]. 2315 [2302]. 2316 [2303]. 2317 [2304]. 2318 [2305]. 2319 [2306]. 232 [232, 1307, 1649]. 2320 [2307]. 2321 [2308]. 2322 [2309]. 2323 [2310]. 2324 [2311]. 2325 [2312]. 2326 [2313]. 2327 [2314]. 2328 [2315]. 2329 [2316]. 233 [233]. 2330 [2317]. 2331 [2318]. 2332 [2319]. 2333 [2320]. 2334 [2321]. 2335 [2322]. 2336 [2323]. 2337 [2324]. 2338 [2325]. 2339 [2326]. 234 [234]. 2340 [2327]. 2341 [2328]. 2342 [2329]. 2343 [2330]. 2344 [2331]. 2345 [2332]. 2346 [2333]. 2347 [2334]. 2348 [2335]. 2349 [2336]. 235 [235]. 2350 [2337]. 2351 [2338]. 2352 [2339]. 2353 [2340]. 2354 [2341]. 2355 [2342]. 2356 [2343]. 2357 [2344]. 2358 [2345]. 2359 [2346]. 236 [236]. 2360 [2347]. 2361 [2348]. 2362 [2349]. 2363 [2350]. 2364 [2351]. 2365 [2352]. 2366 [2353]. 2367 [2354]. 2368 [2355]. 2369 [2356]. 237 [237]. 2370 [2357]. 2371 [2358]. 2372 [2359]. 2373 [2360]. 2374 [2361]. 2375 [2362]. 2376 [2363]. 2377 [2364]. 2378 [2365]. 2379 [2366]. 238 [238]. 2380 [2367]. 2381 [2368]. 2382 [2369]. 2383 [2370]. 2384 [2371]. 2385 [2372]. 2386 [2373]. 2387 [2374]. 2388 [2375]. 2389 [2376]. 239 [239]. 2390 [2377]. 2391 [2378]. 2392 [2379]. 2393 [2380]. 2394 [2381]. 2395 [2382]. 2396 [2383]. 2397 [2384]. 2398 [2385]. 2399 [2386]. 24 [24]. 240 [240]. 2400 [2387]. 2401 [2388]. 2402 [2389]. 2403 [2390]. 2404 [2391]. 2405 [2392]. 2406 [2393]. 2407 [2394]. 2408 [2395]. 2409 [2396]. 241 [241]. 2410 [2397]. 2411 [2398]. 2412 [2399]. 2413 [2400]. 2414 [2401]. 2415 [2402]. 2416 [2403]. 2417 [2404]. 2418 [2405]. 2419 [2406]. 242 [242]. 2420 [2407]. 2421 [2408]. 2422 [2409]. 2423 [2410]. 2424 [2411]. 2425 [2412]. 2426 [2413]. 2427 [2414]. 2428 [2415]. 2429 [2416]. 243 [243]. 2430 [2417]. 2431 [2418]. 2432 [2419]. 2433 [2420]. 2434 [2421]. 2435 [2422]. 2436 [2423]. 2437 [2424]. 2438 [2425]. 2439 [2426]. 244 [244]. 2440 [2427]. 2441 [2428]. 2442 [2429]. 2443 [2430]. 2444 [2431]. 2445 [2432]. 2446 [2433]. 2447 [2434]. 2448 [2435]. 2449 [2436]. 245 [245]. 2450 [2437]. 2451 [2438]. 2452 [2439]. 2453 [2440]. 2454 [2441]. 2455 [2442]. 2456 [2443]. 2457 [2444]. 2458 [2445]. 2459 [2446]. 246 [246]. 2460 [2447]. 2461 [2448]. 2462 [2449]. 2463 [2450]. 2464 [2451]. 2465 [2452]. 2466 [2453]. 2467 [2454]. 2468 [2455]. 2469 [2456]. 247 [247]. 2470 [2457]. 2471 [2458]. 2472 [2459]. 2473 [2460]. 2474 [2461]. 2475 [2462]. 2476 [2463]. 2477 [2464]. 2478 [2465]. 2479 [2466]. 248 [248]. 2480 [2467]. 2481 [2468]. 2482 [2469]. 2483 [2470]. 2484 [2471]. 2485 [2472]. 2486 [2473]. 2487 [2474]. 2488 [2475]. 2489 [2476]. 249 [249]. 2490 [2477]. 2491 [2478]. 2492 [2479]. 2493 [2480]. 2494 [2481]. 2495 [2482]. 2496 [2483]. 2497 [2484]. 2498 [2485]. 25 [25]. 250 [250]. 2502 [2486]. 2503 [2487]. 2504 [2488]. 2505 [2489]. 2507 [2490]. 2508 [2491]. 2509 [2492]. 251 [251]. 2513 [2493]. 2514 [2494]. 2515 [2495]. 2517 [2496]. 2518 [2497]. 2519 [2498]. 252 [252]. 2520 [2499]. 253 [253]. 254 [254]. 255 [255]. 256 [256]. 257 [257]. 26 [26]. 263 [258]. 264 [259]. 265 [260]. 266 [261]. 267 [262]. 268 [263]. 269 [264]. 27 [27]. 270 [265]. 271 [266]. 272 [267]. 273 [268]. 274 [269]. 275 [270]. 276 [271]. 277 [272]. 278 [273]. 279 [274]. 28 [28]. 280 [275]. 281 [276]. 282 [277]. 283 [278]. 284 [279]. 285 [280]. 286 [281]. 287 [282]. 288 [283]. 289 [284]. 29 [29, 2117]. 290 [285]. 291 [286]. 292 [287]. 293 [288]. 294 [289]. 295 [290]. 296 [291]. 297 [292]. 298 [293]. 299 [294].

3 [3, 1257]. 3.0/3 [2404]. 3.0/3.1 [2009, 2353]. 30 [30]. 30-Second [405]. 300 [295]. 301 [296]. 302 [297]. 303 [298]. 304 [299]. 305 [300]. 306 [301]. 307 [302]. 308 [303]. 309 [304]. 31 [31]. 310 [305]. 311 [306]. 312 [307]. 313 [308]. 314 [309]. 315 [310]. 316 [311]. 317 [312]. 318 [313]. 319 [314]. 32 [32]. 320 [315]. 321 [316]. 322 [317]. 323 [318]. 324 [319]. 325 [320]. 326 [321]. 327 [322]. 3270 [1031]. 328 [323]. 329 [324]. 33 [33]. 330 [325]. 331 [326]. 332 [327]. 333 [328]. 334 [329]. 335 [330]. 336 [331]. 337 [332]. 338 [333]. 339 [334]. 34 [34]. 340 [335]. 341 [336]. 342 [337]. 343 [338]. 344 [339]. 345 [340]. 346 [341]. 347 [342]. 348 [343]. 349 [344]. 3492 [2500]. 35 [35]. 350 [345]. 351 [346]. 352 [347]. 353 [348]. 354 [349]. 355 [350]. 356 [351]. 357 [352]. 358 [353]. 359 [354]. 36 [36]. 360 [109, 198, 330, 355]. 360/67 [109, 198]. 360/91 [340]. 361 [356]. 362 [357]. 363 [358]. 364 [359]. 365 [360]. 366 [361]. 367 [362]. 368 [363]. 369 [364]. 37 [37]. 370 [365]. 371 [366]. 3718 [2501]. 372 [367]. 373 [368]. 374 [369]. 375 [370]. 376 [371]. 377 [372]. 378 [373]. 379 [374]. 38 [38]. 380 [375]. 381 [376]. 382 [377]. 383 [378]. 384 [379]. 385 [380]. 386 [381]. 387 [382]. 388 [383]. 389 [384]. 39 [39]. 390 [385]. 391 [386]. 392 [387]. 393 [388]. 394 [389]. 395 [390]. 396 [391]. 397 [392]. 398 [393]. 399 [394]. 3DESE [2407].

4 [4, 1644, 1646, 1647, 1761, 1763, 1764, 1814, 2270]. 4.2 [901]. 40 [40]. 400 [395]. 401 [396]. 402 [397]. 403 [398]. 404 [399]. 4042 [2502]. 405 [400]. 406 [401]. 407 [402]. 408 [403]. 409 [404]. 41 [41]. 410 [405]. 411 [406]. 412 [407]. 413 [408]. 414 [409]. 415 [410]. 416 [411]. 417 [412]. 418 [413]. 419 [414]. 42 [42]. 420 [415]. 421 [416]. 422 [417]. 423 [418]. 424 [419]. 425 [420]. 426 [421]. 427 [422]. 429 [423]. 43 [43]. 430 [424]. 431 [425]. 432 [426]. 433 [427]. 434 [428]. 435 [429]. 436 [430]. 437 [431]. 438 [432]. 439 [433]. 44 [44]. 440 [434]. 441 [435]. 442 [436]. 443 [437]. 444 [438]. 445 [439]. 446 [440]. 447 [441]. 448 [442]. 449 [443]. 45 [45]. 450 [444, 793]. 450/500 [793]. 451 [445]. 452 [446]. 453 [447]. 454 [448]. 455 [449]. 456 [450]. 457 [451]. 458 [452]. 459 [453]. 46 [46]. 460 [454]. 461 [455]. 462 [456]. 463 [457]. 464 [458]. 465 [459]. 466 [460]. 467 [461]. 468 [462]. 469 [463]. 47 [47]. 470 [464]. 471 [465]. 472 [466]. 473 [467]. 474 [468]. 475 [469]. 476 [470]. 477 [471]. 478 [472]. 479 [473]. 48 [48]. 48.bit [816]. 480 [474]. 481 [475]. 482 [476]. 483 [477]. 484 [478]. 485 [479]. 486 [480]. 487 [481]. 488 [482]. 489 [483]. 49 [49]. 490 [484]. 491 [485]. 492 [486]. 493 [487]. 494 [488]. 495 [489]. 496 [490]. 497 [491]. 498 [492]. 499 [493]. 4rev1 [2047].

5 [5]. 5/16 [164]. 5/1971 [164]. 50 [50]. 500 [494, 793]. 501 [495]. 502 [496]. 503 [497]. 504 [498]. 505 [499]. 506 [500]. 507 [501]. 508 [502]. 509 [503]. 51 [51]. 510 [504]. 511 [505]. 512 [506]. 513 [507]. 514 [508]. 515 [509]. 516 [510]. 517 [511]. 518 [512]. 519 [513]. 52 [52]. 520 [514]. 521 [515]. 522 [516]. 523 [517]. 524 [518]. 525 [519]. 526 [520]. 527 [521]. 528 [522]. 529 [523]. 53 [53]. 530 [524]. 531 [525]. 532 [526]. 533 [527]. 534 [528]. 535 [529]. 536 [530]. 537 [531]. 538 [532]. 539 [533]. 54 [54]. 540 [534]. 542 [535]. 543 [536]. 544 [537]. 545 [538]. 546 [539]. 547 [540]. 548 [541]. 549 [542]. 55 [55]. 550 [543]. 551 [544]. 552 [545]. 553 [546]. 554 [547]. 555 [548]. 556 [549]. 557 [550]. 558 [551]. 559 [552]. 56 [56]. 560 [553]. 561 [554]. 562 [555]. 563 [556]. 564 [557]. 565 [558]. 566 [559]. 567 [560]. 568 [561]. 569 [562]. 57 [57]. 570 [563]. 571 [564]. 572 [565]. 573 [566]. 574 [567]. 575 [568]. 576 [569]. 577 [570]. 578 [571]. 579 [572]. 58 [58]. 580 [573]. 581 [574]. 582 [575]. 583 [576]. 584 [577]. 585 [578]. 586 [579]. 587 [580]. 588 [581]. 589 [582]. 59 [59]. 590 [583]. 591 [584]. 592 [585]. 593 [586]. 594 [587]. 595 [588]. 596 [589]. 597 [590]. 598 [591]. 599 [592].

6 [6, 1872, 2452, 2453]. 60 [60]. 600 [593]. 600-699 [689]. 601 [594]. 602 [595]. 603 [596]. 604 [597]. 605 [598]. 606 [599]. 607 [600]. 608 [601]. 609 [602]. 61 [61]. 610 [603]. 611 [604]. 612 [605]. 613 [606]. 614 [607]. 615 [608]. 616 [609]. 617 [610]. 618 [611]. 619 [612]. 62 [62]. 620 [613]. 621 [614]. 622 [615]. 623 [616]. 624 [617]. 625 [618]. 626 [619]. 627 [620]. 628 [621]. 629 [622]. 63 [63]. 630 [623]. 631 [624]. 632 [625]. 633 [626]. 634 [627]. 635 [628]. 636 [629]. 637 [630]. 638 [631]. 639 [632]. 64 [64]. 640 [633]. 641 [634]. 642 [635]. 643 [636]. 644 [637]. 645 [638]. 646 [639]. 647 [640]. 648 [641]. 65 [65]. 651 [642]. 652 [643]. 653 [644]. 654 [645]. 655 [646]. 656 [647]. 657 [648]. 658 [649]. 659 [650]. 66 [66]. 660 [651]. 661 [652]. 662 [653]. 663 [654]. 664 [655]. 666 [656]. 667 [657]. 668 [658]. 669 [659]. 67 [67, 109, 198, 460]. 670 [660]. 671 [661]. 672 [662]. 673 [663]. 674 [664]. 675 [665]. 676 [666]. 677 [667]. 678 [668]. 679 [669]. 68 [68]. 680 [670]. 681 [671]. 682 [672]. 683 [673]. 684 [674]. 685 [675]. 686 [676]. 687 [677]. 688 [678]. 689 [679]. 69 [69]. 690 [680]. 691 [681]. 692 [682]. 693 [683]. 694 [684]. 695 [685]. 696 [686]. 697 [687]. 698 [688]. 699 [689].

7 [7, 1632, 2139]. 70 [70]. 700 [690]. 700-799 [790]. 701 [691]. 702 [692]. 703 [693]. 704 [694]. 705 [695]. 706 [696]. 707 [697]. 708 [698]. 709 [699]. 71 [71]. 710 [700]. 711 [701]. 712 [702]. 713 [703]. 714 [704]. 715 [705]. 716 [706]. 717 [707]. 718 [708]. 719 [709]. 72 [72]. 720 [710]. 721 [711]. 722 [712]. 723 [713]. 724 [714]. 725 [715]. 726 [716]. 727 [717]. 728 [718]. 729 [719]. 73 [73]. 730 [720]. 731 [721]. 732 [722]. 733 [723]. 734 [724]. 735 [725]. 736 [726]. 737 [727]. 738 [728]. 739 [729]. 74 [74]. 740 [730]. 741 [731]. 742 [732]. 743 [733]. 744 [734]. 745 [735]. 746 [736]. 747 [737]. 748 [738]. 749 [739]. 75 [75]. 750 [740]. 751 [741]. 752 [742]. 753 [743]. 754 [744]. 755 [745]. 756 [746]. 757 [747]. 758 [748]. 759 [749]. 76 [76]. 760 [750]. 761 [751]. 762 [752]. 763 [753]. 764 [754]. 765 [755]. 766 [756]. 767 [757]. 768 [758]. 769 [759]. 77 [77]. 770 [760]. 771 [761]. 772 [762]. 773 [763]. 774 [764]. 775 [765]. 776 [766]. 777 [767]. 778 [768]. 779 [769]. 78 [78]. 780 [770]. 781 [771]. 782 [772]. 783 [773]. 784 [774]. 785 [775]. 786 [776]. 787 [777]. 788 [778]. 789 [779]. 79 [79]. 790 [780]. 791 [781]. 792 [782]. 793 [783]. 794 [784]. 795 [785]. 7958 [265]. 796 [786]. 797 [787]. 798 [788]. 799 [789].

8 [8, 833, 2031, 2240, 2266]. 80 [80]. 800 [790]. 800-899 [889]. 801 [791]. 802 [792]. 802.12 [2007]. 803 [793]. 804 [794]. 805 [795]. 806 [796]. 807 [797]. 808 [798]. 809 [799]. 81 [81]. 810 [800]. 811 [801]. 812 [802]. 813 [803]. 814 [804]. 815 [805]. 816 [806]. 817 [807]. 818 [808]. 819 [809]. 82 [82]. 820 [810]. 821 [811]. 822 [812, 977, 1016, 1128, 1138, 1317, 1484, 1485, 1828, 2143]. 822/MIME [2143, 2144]. 823 [813]. 824 [814]. 825 [815]. 826 [816]. 827 [817]. 828 [818]. 829 [819]. 83 [83, 832, 833]. 830 [820]. 831 [821]. 832 [822]. 833 [823]. 834 [824]. 835 [825]. 836 [826]. 837 [827]. 838 [828]. 839 [829]. 84 [84, 1486].

840 [830]. 841 [831]. 842 [832]. 843 [833]. 844 [834]. 845 [835]. 846 [836]. 847 [837]. 848 [838]. 849 [839]. 85 [85]. 850 [840]. 851 [841]. 852 [842]. 853 [843]. 854 [844]. 855 [845]. 856 [846]. 857 [847]. 858 [848]. 859 [849]. 86 [86]. 860 [850]. 861 [851]. 862 [852]. 863 [853]. 864 [854]. 865 [855]. 866 [856]. 867 [857]. 868 [858]. 869 [859]. 87 [87]. 870 [860]. 871 [861]. 872 [862]. 873 [863]. 874 [864]. 875 [865]. 876 [866]. 877 [867]. 878 [868]. 879 [869]. 88 [88, 1486]. 880 [870]. 881 [871]. 882 [872]. 883 [873]. 884 [874]. 885 [875]. 886 [876]. 887 [877]. 888 [878]. 889 [879]. 89 [89]. 890 [880]. 891 [881]. 892 [882]. 893 [883]. 894 [884]. 895 [885]. 896 [886]. 897 [887]. 898 [888]. 899 [889]. 8bit [1416, 1418, 1642]. 8bit-MIMEtransport [1416, 1642]. 8bit-SMTP [1418]. 8bit-SMTP/MIME [1418].

9 [9, 2502]. 90 [90]. 900 [890]. 900-999 [989]. 901 [891]. 902 [892]. 903 [893]. 904 [894]. 905 [895]. 906 [896]. 907 [897]. 908 [898]. 909 [899]. 91 [91]. 910 [900]. 911 [901]. 912 [902]. 913 [903]. 914 [904]. 915 [905]. 916 [906]. 917 [907]. 918 [908]. 919 [909]. 92 [92]. 920 [910]. 921 [911]. 922 [912]. 923 [913]. 924 [914]. 925 [915]. 926 [916]. 927 [917]. 928 [918]. 929 [919]. 93 [93]. 930 [920]. 931 [921]. 932 [922]. 933 [923]. 934 [924]. 935 [925]. 936 [926]. 937 [927]. 938 [928]. 939 [929]. 94 [94]. 940 [930]. 941 [931]. 942 [932]. 943 [933]. 944 [934]. 945 [935]. 946 [936]. 947 [937]. 948 [938]. 949 [939]. 95 [95]. 950 [940]. 951 [941]. 952 [942]. 953 [943]. 954 [944]. 955 [945]. 956 [946]. 957 [947]. 958 [948]. 959 [949]. 96 [96, 2390, 2391]. 960 [950]. 961 [951]. 962 [952]. 963 [953]. 964 [954]. 965 [955]. 966 [956]. 967 [957]. 968 [958]. 969 [959]. 97 [97]. 970 [960]. 971 [961]. 972 [962]. 973 [963]. 974 [964]. 975 [965]. 976 [966]. 977 [967]. 978 [968]. 979 [969]. 98 [98]. 980 [970]. 981 [971]. 982 [972]. 983 [973]. 984 [974]. 985 [975]. 986 [976]. 987 [977]. 988 [978]. 989 [979]. 99 [99]. 990 [980]. 991 [981]. 992 [982]. 993 [983]. 994 [984]. 995 [985]. 996 [986]. 997 [987]. 998 [988]. 999 [989].

AAL5 [1616, 2351]. AARNet [1552]. ABNF [2221]. Academic [1606]. ACAP [2231]. ACCESS [1720, 2047, 96, 514, 529, 792, 841, 868, 897, 995, 1054, 1166, 1193, 1211, 1466, 1477, 1482, 1604, 1767, 1788, 1860, 2004, 2044, 2049, 2056, 2093, 2101, 2158, 2162, 2231, 2238, 2239, 2240, 2252, 2262, 2473, 2478, 2501, 110, 195, 216, 499, 545, 821, 1027]. Access-Type [2004]. Accessed [2167]. Accounting [1262, 1336, 1662, 2046, 2126, 2493, 136]. Accounts [345, 512]. Accuracy [1345]. Achieve [1771, 1474]. Achieving [203]. Acknowledgement [803]. Acknowledgment [2005]. ACL [2073]. Acquisition [1918]. Across [1510]. Act [1111]. Active [856]. Activities [1110, 1150, 1252]. Activity [316, 113]. Ad [1099]. Adaptation [1473]. add [2468]. Addendum [931, 1016, 1858]. Addition [584, 276]. Additional [1378, 1713, 1914, 650]. ADDR [2304]. Address [208, 399, 710, 786, 816, 893, 1268, 1283, 1319, 1325, 1328, 1355, 1356, 1357, 1456, 1471, 1508, 1509, 1535, 1587, 1621, 1629, 1654, 1705, 1725, 1734, 1766, 1826, 1868, 1874, 1884, 1905, 1918, 1958, 1995, 2023, 2060, 2088, 2150, 2254, 2281, 2361, 2362, 2377, 2378, 2449, 2458, 583, 627, 630, 915, 972, 1019, 1226, 2151, 2290, 2291]. Addresses [1267, 1337, 1365, 1620, 1671, 1804, 1828, 1864, 1911, 2456, 504, 744, 804, 816, 976, 1059]. Addressing [1337, 1370, 1743, 1871, 1942, 1996, 2162, 2360, 720, 747, 922, 926, 931]. Administration [1343, 1493]. Administrative [1115, 1126, 1341, 1435, 1519, 1879, 1896, 136]. Administratively [2352]. administrators [1022, 1023]. Admission [2111]. ADPCM [2409]. Advanced [1481, 1528, 2279]. Advancement [2425]. Advancing [1212]. Advertisement [1387]. Advice [1425]. affecting [2108]. again [517]. Against [1935]. Agenda [140]. Agent [1333, 1514, 1810, 1832, 2076, 2244, 2308]. Agents [1293, 414, 2003]. AgentX [2244]. Aggregatable [2361]. Aggregate [2097]. Aggregation [1328, 1472, 1509, 2498]. Agreement [1780, 2326]. AH [2390, 2391]. Ahead [588, 589, 848]. AHIP [995]. AHIP-E [995]. aids [376]. Airline [2338]. AIW [1785]. Alert [2100]. alerts [1214]. Algorithm [1176, 1309, 1310, 1311, 2131, 2255, 2283, 2392, 2397, 971, 1036]. Algorithms [946, 1413, 1493, 1988, 2027, 2438, 805, 1105]. Aliases [2206]. ALLOCATION [2037, 59, 68, 1227, 1336, 1508, 1587, 1619, 1868, 1874, 1884, 1905, 1995, 2112, 2227, 2458]. alone [676]. Alpha [1003]. alternate [1135, 1136]. alternative [1850, 1953]. AM [1303, 2040]. Ames [126, 413]. among [2324]. Analysis [1235, 1255, 1377, 1571, 1575, 1672, 1711, 1764, 1775, 2079]. ANL [544]. Announcement [173, 286, 304, 369, 430, 468, 531, 567, 188, 447, 448, 455, 498]. Announcing [580, 650]. Anonymous [1625, 2232]. Answerback [1384]. Answers [1167, 1196, 1197, 1315, 1568, 1584]. Anti [2489]. Anti-Spam [2489]. Anticipate [1349]. Anycasting [1536]. Anyway [1922, 2058]. API [1499, 1948, 1951, 2012, 2279, 2354, 2465, 2466]. APPC [2038]. Appeal [1904]. Appendix [706]. AppleTalk [1368, 1233, 1409, 1494, 1732]. Applicability [1360, 1507, 1712, 1910, 1992, 2026, 2068, 2195, 2320]. Application [1154, 1258, 1498, 1645, 1762, 1813, 1882, 1985, 2065, 2157, 2174, 2207, 2344, 2466, 1113, 2231]. Application/CALS [1882]. Application/CALS-1840 [1882]. Application/MARC [2207]. Applications [1688, 1876, 2250, 2260, 2274, 2364, 2500, 809, 945, 1247]. APPN [1583, 2142, 2153, 2340, 2442, 2443]. APPN/HPR [2340]. Approach [1312, 2294, 1019]. Approaches [2089, 2090]. April [503, 1003]. Arbitrarily [1831]. Arbitrary [368, 1454]. ARC [153, 411, 537, 614]. ARC-NIC [153]. Architectural [1945, 2263]. Architecture [1277, 1348, 1468, 1508, 1591, 1610, 1611, 1623, 1626, 1681, 1697, 1743, 1811, 1815, 1825, 1871, 1874, 1900, 1979, 2050, 2081, 2085, 2092, 2188, 2248, 2258, 2303, 2360, 2365, 2388, 2417, 2462, 1083, 1212]. Architectures [2136, 865]. Archiving [734]. ARCNET [1041, 1191, 2484]. Area [1459]. AREQUIPA [2157]. Arithmetic [1969]. Armenia [2040]. ARP [1017, 1041, 1364, 1380, 1423, 1567, 1855, 2212, 2307]. ARPA [8, 32, 46, 155, 168, 211, 295, 298, 299, 311, 324, 351, 358, 372, 377, 379, 384, 397, 562, 640, 714, 723, 812, 891, 892, 914, 934, 951, 981]. ARPA-Internet [891, 914, 934, 951, 981]. ARPA-like [640]. ARPANET [254, 297, 303, 346, 359, 364, 418, 502, 505, 512, 577, 578, 593, 612, 628, 653, 704, 747, 792, 841, 842, 861, 868, 995]. ARPANET-Type [704]. ARPAWOCKY [521]. arrives [973]. Articles [2121]. Artificial [152]. Arts [2137]. Ascend [1921, 2094]. ASCII [20, 183, 333, 563, 620, 688, 1332, 1512, 1830, 1831, 2034]. Asked [1167, 1568, 1852, 1928, 1196, 1197, 1315, 1584]. Aspects [2172]. Assassination [1327]. Assessment [628]. Assigned [312, 597, 707, 729, 740, 745, 748, 752, 760, 766, 780, 810, 860, 890, 913, 933, 950, 980, 1000, 1050, 2159, 621, 1330, 1690]. Assignment [1164, 1328, 1509, 1516, 1705, 2437, 610, 1209]. Assignments [179, 2362]. Assistance [1192]. Associating [1914]. Association [2395]. Asynchronous [906]. asynchronously [1214]. AT&T [2175, 2435]. AT&T/Neda [2175]. ATCP [1368]. ATM [2404, 1473, 1567, 1616, 1670, 1685, 1744, 1745, 1811, 1913, 1919, 1933, 1941, 2009, 2085, 2136, 2157, 2212, 2213, 2256, 2307, 2318, 2324, 2353, 2366, 2367, 2368, 2369, 2370, 2479, 2494, 2495]. ATMP [2094]. Attachment [1505, 2226]. Attachments [306]. Attacks [1935, 2254]. Attendance [108]. Attendees [1381, 1529, 1708]. attention [178]. Attribute [1478, 1768, 1984, 1985, 2029, 2066, 2116, 2239]. Attributes [1454]. Audio [1877, 2185]. Augmentation [34, 191]. Augmented [2221]. August [386, 559, 691]. Authenticated [1814]. Authentication [902, 921, 1324, 1399, 1406, 1411, 1497, 1500, 1501, 1694, 1721, 1816, 1818, 1839, 1916, 1948, 1981, 2045, 2056, 2069, 2072, 2091, 2125, 2209, 2271, 2389, 2472, 637, 979, 994, 1030, 1103, 1724, 1401, 1402]. Authoring [2278, 2497]. AUTHorize [2082, 2182]. Authors [1533, 2210, 1101]. Autoconfiguration [1958, 2449]. Automatic [1918]. Automating [1493]. Autonomous [965, 1917, 1952, 1116]. Availability [488, 303]. Available [1282, 1622, 490]. averages [539]. AVI [2348]. avian [1139]. Avoidance [1988, 2296]. AX.25 [1216].

b [1558]. B6700 [695]. backbone [1064, 1082]. Background [1058, 1262]. Backup [821]. Backward [1375]. BACP [2112]. bake [1015]. Balancing [1784]. band [711]. Bandwidth [560, 2112, 410, 561, 1067]. Banyan [1753]. BAP [2112]. Base [1056, 1146, 1148, 1152, 1203, 1238, 1242, 1243, 1275, 1305, 1440, 1441, 1502, 1618, 1686, 1687, 1732, 1747, 1837, 1894, 1998, 1999, 2000, 2008, 2102, 2114, 2115, 2193, 2200, 2201, 2452, 2453, 1204, 1233, 1261, 2439, 2441]. Based [831, 967, 1373, 1412, 1472, 1494, 1814, 1830, 1996, 2060, 2176, 2188, 2434, 2480, 230, 308, 697, 746, 796, 1055, 1056, 1064, 1075, 1082, 1094, 1104, 1145, 1146, 1148, 1203, 1293, 1674, 1854, 1888, 1897, 2009, 2136, 2353, 2373, 2404, 2251, 2252, 2261, 2262]. Basic [1688, 2120]. Basis [2227]. Batch [2429].

BBN [47, 265, 296, 657, 706, 798]. BCP [1628]. Be [411, 1765, 538, 1617]. bed [1249]. Before [1869, 958]. Behaviour [2088]. Belated [63]. Benchmarking [1232, 1931, 2272, 2419]. Berkeley [901]. Best [149, 1808]. Between [1524, 2326, 563, 977, 1016, 1076, 1123, 1127, 1128, 1138, 1317, 1391, 1395, 1442, 1484, 1485, 1496, 1780, 1828, 1895, 2143, 2144, 2149, 2423]. Beverage [2312]. BFTP [1058]. BGP [1095, 1153, 1255, 1256, 1257, 1354, 1393, 1644, 1646, 1647, 1761, 1763, 1764, 1850, 1952, 1953, 1984, 1985, 2029, 2270, 2372, 2426]. BGP-3 [1257]. BGP-4 [1644, 1646, 1647, 1761, 1763, 1764, 2270]. BGP/IDRP [1850]. BGP2 [1387]. BGP3 [1387]. BGP4 [1735]. BGP4/IDRP [1735]. Bi [1546, 2076]. Bi-directional [1546]. bi-lingual [2076]. Bibliographic [1347, 1797, 2275]. Bibliography [1002, 1453, 243, 285, 1165]. Big [1535, 1629, 1096, 1100]. Bigger [1337]. Bill [1285]. Binary [31, 846, 1820]. Binding [1488, 1823]. bindings [1499]. BinHex [1731]. Biographies [1241, 1326]. bis [2406]. Bit [2162, 1741]. Bitmap [787]. Block [2127]. blocking [842]. Blocksize [1773, 2335]. BNF [2221]. Board [1110, 1150, 1348, 1591]. Bob [6, 367]. Bodies [1331, 1484, 1485, 1511, 2032, 2144]. Body [1860, 2004, 2145, 2146, 2148]. BOF [1744]. Books [1422]. BOOTP [1038, 1074, 1385, 1487, 1523, 1524, 2119]. Bootstrap [896, 941, 1522, 1532]. Bootstring [2500]. Border [1644, 1761, 2444, 1095, 1153, 1154, 1257, 1258, 1259, 1387, 1645, 1647, 1762]. Boredom [1428]. bound [19]. Boundaries [1510]. Box [196, 221, 273]. Bridge [1464, 1076]. Bridges [1276, 1483, 1515]. Bridging [1628, 1210]. Brief [1407, 1748, 34, 160]. Broadcast [2213, 2478]. Broadcasting [909, 912, 937]. Brussels [1200]. BSD [1248, 1272, 1672, 1964]. BT.656 [2418]. Bud [1792]. Buffers [292, 2403]. Building [1249, 1292, 1930, 2496]. Bulk [1177, 959, 988]. Bundle [2481]. Bundled [2330]. Bus [1220]. BusyTime [2433]. BVCP [1753]. By-Laws [2122]. By-Request [1296]. Byte [176, 719, 725]. Bytes [128].

C [1499]. C-bindings [1499]. c=US [1208, 1245]. Cable [1676]. Cabletron [2111]. Cache [2173, 2174, 2321]. Caching [2295]. Calendaring [2432]. Call [468, 624, 1040, 1047, 1821, 233, 664]. Called [1482]. calling [674]. CALS-1840 [1882]. Campus [384, 418]. Cancellation [477]. Canonical [2456]. Capability [702]. Capture [1751]. Card [1885, 490]. care [595]. carriage [643]. carriage-return [643]. carriers [1139]. Carrying [1378, 1713, 2147]. Case [71]. Cases [2189, 2273]. CAST [2131]. CAST-128 [2131]. Catalog [1282, 1460, 1622, 2103, 1137]. Catalog-96 [2103]. Catalogue [1994]. Categorization [100]. CATNIP [1697]. Caution [2456]. CBC [1819, 2027, 2392, 2438]. CBC-Mode [2438]. CBI [316]. CBT [2176, 2188]. CC [526]. CCA [415]. CCITT [794]. CCN [90, 340, 374, 484, 492, 582]. CCP [1949]. CCSO [2365]. CEASE [68]. CellB [2016]. Cellular [1664]. Center [34, 90, 115, 126, 191, 223, 351, 1345, 2137, 231, 96]. CENTURY [1597].

Certificate [1412, 2299, 2446, 1104]. Certificate-Based [1412, 1104]. Certification [198, 1414, 2301]. Chairman [111, 130]. Challenge [1981, 2082, 2182]. Challenge/Response [2082, 2182]. Challenges [2090]. Change [67, 69, 307, 326, 464, 540, 627, 630, 1164, 213, 256, 266, 338, 444, 450, 461, 583, 586, 685, 694]. Changes [72, 323, 335, 388, 389, 399, 1263, 1983, 217, 604, 651, 677, 680, 682, 686, 963]. Changing [218]. Channels [2475]. CHAP [1981, 2420]. Character [368, 423, 854, 1306, 1335, 1458, 1479, 1492, 1545, 1547, 1648, 1805, 1830, 1909, 1934, 2117, 2171, 2218, 2224, 2264, 2306, 205]. Characterization [2202]. Characters [1830, 1831]. Characters-Based [1830]. Charset [2265, 2053]. Charter [577, 1348, 1591]. Charting [1599]. Chat [1449]. Checking [1329]. Checklist [1810, 1832]. Checkout [193]. Checkpoint [1833, 214]. Checkpoint/Restart [1833]. Checksum [1614, 1923, 1061, 1131, 1135, 1136]. checksumming [522]. chimney [595]. Chinese [1830, 1831, 1909]. Choosing [1168, 1361, 1716]. Christmas [1869]. CIDR [1457, 1507, 1508, 1509, 1510, 1807]. Cipher [2392, 2438]. CIPX [1543]. Circuit [1296, 1963]. Circuit-Switched [1296]. Circuit-Terminating [1963]. Circuits [1571, 1572, 1783, 2078]. Cisco [2092, 2268, 2328, 1603]. Clarification [228]. Clarifications [1522, 1532, 2168]. Class [1787, 1846, 1866, 2066, 2023]. Classes [1365, 482]. Classful [1807]. Classical [1567, 1906, 2212, 2307, 2323]. Classifications [1701]. Classless [1507, 1509, 2304]. Cleaning [107, 102]. Client [1183, 1843, 1967, 2041, 2101]. Client-Server [1843]. Client-Side [1967]. CLNP [1516, 1551, 1565, 1758]. CLNS [1213, 1228]. Clock [768, 34, 947]. clocks [946]. Closed [1785, 2153, 2340]. Cluster [2108]. CMIP [1179]. CMOT [1085, 1179]. Code [1446, 1834, 623]. Codec [2348]. Codes [1384, 1880, 2021, 183, 633]. Codified [1617]. Coexistence [1442, 1895]. Coffee [2311]. Coherent [1225]. Collaboration [2423, 1200]. Collection [2493, 662]. Com [2204]. Come [405]. Coming [45]. Command [446, 1841, 2184, 500, 687, 939, 1724, 2164]. Commands [68, 145, 2356, 765]. Comment [483, 1443]. Commentary [674]. Comments [38, 39, 40, 44, 47, 50, 65, 68, 127, 141, 148, 151, 175, 176, 224, 238, 363, 380, 388, 424, 507, 529, 552, 556, 575, 600, 607, 616, 617, 680, 682, 686, 689, 708, 763, 790, 889, 989, 990, 1002, 1089, 1189, 1663, 1789, 1886, 1986, 2086, 2186, 2286, 2386, 117, 129, 409, 457, 815, 1008, 1101, 1101]. Commercial [1158]. Commercialization [1182]. Committee [173, 102, 107]. Committees [2014, 2269]. Common [562, 1361, 1394, 1501, 1526, 1527, 1697, 1844, 1899, 2063, 2129, 2205, 1085, 1179]. Commonly [1167, 1315, 1568, 1584, 1196, 1197]. Commotion [1207]. Communicating [1796, 2170]. Communication [41, 61, 62, 435, 1667, 32, 203, 624, 955, 982, 1112, 1183]. Communications [157, 366, 1260, 1688, 190, 604, 818]. Communities [1984]. Community [1320, 1606, 1639, 1888, 1904, 1985]. Community-based [1888]. Compact [1911]. Company [2332]. Comparison [1444]. COMPATIBILITY [1722, 2048, 1375]. Compilation [182]. compiler [194]. compiler/interpreter [194]. Components [2023]. Compressed [1937, 1938, 2022, 2422]. Compressing [1134, 1543, 2491]. Compression [1949, 1954, 1961, 1962, 1963, 1964, 1965, 1980, 2105, 2380, 2381, 2382, 2490, 2492, 462]. Computation [1614]. Computational [1009]. Computer [31, 61, 62, 169, 174, 285, 308, 366, 795, 831, 1314, 2130, 2337, 32, 144, 146, 494, 796, 798, 1168]. computer-to-computer [32]. computers [32, 195, 241, 254, 640, 904, 914, 974, 983, 1046]. computers-data [195]. Computing [384, 418, 702, 1061, 1007]. Concept [187]. Concepts [872, 991, 1369, 603, 1024]. conceptualization [191]. Concerns [1665]. Concise [1202]. condition [161, 619]. Conditional [1975]. Confederations [1952, 965]. Conference [366]. Conferences [1877]. Conferencing [1314, 1443]. Configuration [1333, 1514, 1521, 1527, 1531, 1899, 2118, 2231, 2277, 2471, 1162]. Confirmation [2014, 2269]. Conformance [1434, 1891, 2036]. Conformant [2150]. Congestion [886, 1244, 1988, 2296, 2468, 19]. Conjunction [1975, 985]. Connected [1164]. Connecting [32, 241, 1349, 640, 904, 914]. Connection [76, 93, 145, 165, 197, 1782, 1788, 1902, 2493, 461, 483]. Connection-less [1788]. Connection-Oriented [2493]. Connectionless [777, 976, 984, 1059, 1152, 1228, 916, 1230]. Connectionless-Mode [976, 1059, 984]. connections [176, 679]. Connectivity [2247, 2485, 606]. consider [192, 440]. Considerations [341, 1263, 1660, 1661, 1678, 1845, 2071, 2421]. Considered [1253, 1316, 1617]. Console [306, 110]. Consortium [1207, 2501]. Consulting [416]. Content [1039, 1427, 1513, 1534, 1553, 1731, 1796, 1851, 1859, 1860, 1879, 1882, 1883, 2064, 2098, 2099, 2170, 2207, 2282, 2374, 2379, 2411, 2412, 1486]. Content-Disposition [1796, 2170]. Content-ID [1860, 2098, 2379]. Content-MD5 [1534, 1851]. Content-Type [2412, 1039, 1513, 1553, 1859, 1882, 1883, 2099, 2207, 2374]. Content-Types [1427, 1486]. contents [756, 764]. Context [2107]. Continuations [2171, 2218]. Control [18, 23, 59, 68, 86, 145, 210, 351, 665, 711, 751, 767, 782, 783, 954, 1244, 1297, 1322, 1336, 1362, 1366, 1367, 1368, 1461, 1463, 1464, 1482, 1542, 1628, 1737, 1752, 1753, 1754, 1829, 1846, 1864, 1872, 1877, 1902, 1949, 1955, 1999, 2030, 2084, 2112, 2115, 2127, 2204, 2252, 2262, 2311, 2439, 2450, 22, 436, 443, 779, 886, 1013, 1066, 1070, 2044]. Controlled [553, 574, 716, 2198, 2368, 1247]. Controlled-Load [2198, 2368]. Controlling [2493]. Convention [1205, 1293, 1638, 1741, 2156, 2339, 809]. Conventions [24, 27, 30, 110, 175, 1433, 1446, 1890, 2452, 2480, 2494, 3, 10, 167]. Conversation [6, 367]. Conversations [49]. Conversion [396]. converting [816]. Coordinates [396]. Coordination [249, 1455]. Copy [315]. Core [2188, 2356, 2400, 2432, 2176]. Cornell [1681]. Corporate [1668, 1677]. Correction [265, 400, 1525]. Corrections [574]. Correspondence [1391].

COSINE [1264]. costs [387, 420]. could [1006]. Country [560, 561]. course [271]. covering [931]. creation [1917]. Credit [1885]. Criteria [1254, 1659, 1716, 1997, 2036, 2344, 2464]. Critical [1067]. Critique [864]. critiques [548]. CRL [2446]. Crocker [49]. CRONUS [814]. Cross [560, 561, 675]. Cryptographic [1814, 1971, 2302]. Cryptography [2424]. CSCR [1207]. css [2305]. CTS [2027]. Current [436, 443, 1566, 1808]. Cut [97]. CWD [687]. CyberCash [1885]. Cyrillic [1479].

Dacom [793]. daemon [1169]. Damping [2426]. Dangerous [1316]. DARPA [813]. DASS [1497]. Data [42, 80, 86, 125, 138, 144, 157, 163, 166, 173, 188, 194, 195, 227, 242, 259, 286, 289, 299, 304, 305, 311, 322, 431, 480, 566, 608, 638, 703, 721, 722, 818, 898, 1004, 1033, 1141, 1424, 1527, 1737, 1785, 1795, 1822, 1831, 1852, 1924, 1937, 1938, 1941, 1950, 1963, 2011, 2093, 2101, 2185, 2194, 2466, 83, 146, 199, 243, 251, 285, 303, 387, 462, 494, 502, 558, 624, 662, 777, 788, 793, 867, 929, 932, 959, 961, 988, 2375, 2384, 171]. data-tablet [199]. Data/Software [1795]. Database [1472, 1687, 1755]. Databases [1345, 667]. datacomputer [219, 558]. Datagram [758, 2000, 2441, 746, 805]. datagram-based [746]. Datagrams [909, 1178, 1321, 1913, 867, 884, 885, 912, 938, 1032, 1041, 1045, 1078, 1093, 1124, 1139, 1161, 1199]. datalanguage [509, 603]. Date [450]. Day [855, 420]. Days [1869]. Daytime [857]. DCCS [702]. DCE [1963]. DCN [881]. DCNET [768]. DCP [1954]. DDN [1123, 1226]. DDT [515]. Deadlock [202]. Deamon [356]. Debugger [899]. Debugging [636, 1460, 675, 1137, 1703]. DEC [190]. December [437, 605, 591]. Decision [1924]. Declaration [1417, 1643, 1857]. DECnet [1279, 1395, 1549, 1366, 1752]. Decode [5]. Decoding [788]. Dedicated [2257]. Default [1387, 1616]. Defeating [2254]. Defending [1935]. Defense [929, 588, 932, 1098]. Define [1736]. Defining [2476, 993, 1205]. Definition [147, 2004, 2066, 2411, 2461, 931]. Definitions [1173, 1222, 1223, 1259, 1274, 1276, 1294, 1306, 1307, 1308, 1343, 1358, 1388, 1396, 1397, 1461, 1462, 1463, 1464, 1483, 1505, 1506, 1515, 1585, 1586, 1594, 1613, 1633, 1640, 1647, 1648, 1649, 1650, 1655, 1656, 1684, 1685, 1737, 1993, 2007, 2011, 2038, 2095, 2142, 2219, 2225, 2226, 2239, 2253, 2274, 2307, 2312, 2345, 2353, 2404, 2442, 2443, 2444, 2481, 2482, 2483, 2494, 2495, 775, 1014, 1202]. Deflate [1966, 1938, 2381]. degradations [625]. DEL [5]. Delay [405, 879, 1062, 1006]. Delegation [1581, 2217]. Deliberations [1370]. delivered [2428]. Delivery [1760, 1878, 1881, 536, 747]. Demand [59, 1571, 1572, 1783, 1927, 2078]. demo [415]. Demonstration [366]. Denial [2254]. Department [929, 1098, 932]. dependent [474]. Deployed [1525]. Deploying [1420]. Deployment [1320, 1457, 1792, 2135, 2195]. DES-CBC [1819, 2392]. Describing [1293, 1331, 1511, 2248, 2258]. Description [2255, 2314, 177, 206]. descriptions [195]. Descriptive [242]. DESE [1956, 2406]. DESE-bis [2406]. Design [1263, 546, 585, 603]. Designers [573]. Desirability [398]. Destination [1760]. Destinations [1488]. Detailed [992]. details [253]. detection [506, 510, 528, 654].

Determination [2399]. development [1042]. Device [177, 396]. Devices [1306, 1307, 1308, 1358, 1460, 1506, 1648, 1649, 1650, 1931, 2095, 2253, 2272, 2312, 1137, 1232]. dhcp [2309, 1523, 1524, 2119, 2228, 2472, 2476]. diagrams [12]. Dial [2045, 2115, 2125]. Dictionary [2216]. Dies [2504]. Differentiated [2417, 2462, 2461]. Digest [1143, 1176, 1309, 1310, 1311, 2056]. Digital [520, 1681, 2141]. Dilemma [1517]. Directed [1423, 1760, 2044]. Direction [1709]. directional [1546]. Directories [2496, 614]. Directory [765, 1192, 1298, 1299, 1320, 1374, 1420, 1474, 1477, 1552, 1557, 1598, 1599, 1607, 1767, 1771, 1788, 1791, 1792, 1793, 1794, 1826, 1827, 1828, 1930, 2228, 2238, 2239, 2240, 2280, 2281, 2364, 2412, 2413, 1097, 2151, 1265, 1266, 1285]. Directory-Enabled [2364]. DIS [984]. DISA [1391]. Discard [343, 853]. Discovery [1246, 1425, 1957, 1968, 2400, 2448, 1053, 1181]. discuss [447]. Discussion [139, 709, 1314, 87]. Dispatching [2249, 2259]. Display [86, 177, 184, 205, 291, 1086]. Disposition [1796, 2170, 2285, 643, 645, 646, 648, 649]. Distance [1065]. Distant [258, 540, 584]. Distinguished [1475, 1769, 2234, 2240]. Distribute [1419, 1654, 2150]. Distributed [498, 698, 702, 712, 820, 994, 1218, 1266, 1497, 1582, 2278, 2322, 2430, 2486, 2497, 974, 983, 1046, 1426, 1723]. Distributed-protocol [994]. Distributing [1164]. Distribution [69, 95, 1225, 1936, 52, 185, 839]. divers [1020]. DIXIE [1239]. DLSw [1785, 2153, 1785]. DLUR [2219]. DMS [571]. DNCP [1366, 1752]. DNS [2039, 1091, 1173, 1338, 1373, 1391, 1525, 1526, 1527, 1601, 1602, 1627, 1654, 1696, 1702, 1703, 1784, 1873, 1899, 1982, 1983, 2039, 2123, 2150, 2168, 2169, 2206, 2217, 2295, 2496]. do [1006, 1247]. DOCTOR [433]. Document [158, 1646, 1681, 1919, 2153, 2340, 2398, 65, 137, 448, 970, 1426]. Documentation [3, 10, 24, 27, 30, 407, 118, 209, 537]. Documents [1407, 1748, 1914, 2097, 115, 664, 757]. DoD [750, 751, 800, 918, 935, 942, 1029]. DODIIS [1033]. Domain [809, 871, 872, 873, 887, 910, 911, 963, 972, 1022, 1023, 1024, 1025, 1115, 1312, 1376, 1454, 1468, 1469, 1470, 1507, 1509, 1581, 1778, 1801, 1806, 1863, 1943, 2040, 2052, 2123, 2124, 2133, 2227, 2229, 2332, 2339, 2394, 2498, 2500, 964, 1021, 1132, 2155]. Domains [1126, 1269, 1384, 1639, 2234, 789]. Down [541, 615]. downgrading [1318, 1486]. DP [895]. DPI [1218]. Draft [275, 546, 1028, 794, 975]. Drip [2312]. Drip-Type [2312]. DS [2461, 291, 1421, 1474, 1475, 1554]. DS-1 [291]. DS0 [2481]. DS1 [1222, 1396, 2482]. DS2 [2482]. DS3 [1223, 1397, 2483]. DS3/E3 [1397, 2483]. DSA [1554]. DSL [630]. DSP [972]. DTEs [1305, 2102]. DTP [238]. DUA [1421]. Dual [1319, 1185]. DUAs [1363]. Dublin [2400]. due [619]. Duplicate [1037, 667]. Duration [2411]. During [411]. Dynamic [1521, 1531, 1918, 2118, 2123, 2124]. Dynamically [1297].

E-mail [1701, 1810, 1832, 2097]. E-mailing [1347]. E.I.S.S. [1814]. E.I.S.S.-Report [1814]. E1 [1396, 2482]. E2 [2482]. E3 [1397, 2483]. EAP [2271]. EBCDIC [183, 333]. EBCDIC/ASCII [333]. Echo [342, 847, 852, 1129, 1565, 2062]. Echoing [352, 553, 716, 574]. ECN [2468]. economics [1206]. ECP [1955]. EDI [1757, 1852]. Editor [562]. Education [1157]. Efficiency [1705, 807]. Efficient [2502, 2175]. EGP [817, 901, 1082]. EIP [1375]. Electric [1663]. ELECTRONIC [1723, 1411, 1412, 1413, 1414, 1934, 979, 1030, 1103, 1104, 1105, 1852]. Element [2198, 2203]. Elements [698, 2202]. Eliminate [67]. employ [2254]. emption [784]. Enabled [2364]. ENCAPS [1942]. Encapsulating [1817, 2130, 2393]. Encapsulation [1231, 1316, 1473, 1691, 1692, 1730, 1757, 1913, 1990, 1991, 2097, 924, 1216]. encapsulations [883]. encipherment [979, 1030, 1103]. Encode [5]. Encoded [1731, 2171, 2218]. Encoding [1144, 1267, 1446, 1458, 1495, 1545, 1547, 1702, 1830, 1909, 1934, 2016, 2224, 2418, 1091, 1127, 1268, 2500]. encounters [433]. Encrypted [1835]. Encryption [1411, 1902, 1955, 1956, 2131, 2255, 2300, 2397, 2406, 2407]. End [919, 969, 985, 1152, 1228, 640, 695, 875]. End-to-End [969]. Engineering [1254, 1381, 1428, 1529, 1660, 1680, 1708, 1975, 2417]. Enhanced [1494, 1848, 1880, 1973, 2021, 2143, 995]. Enhancement [1411, 1412, 1413, 1414, 979, 1030, 1103, 1104, 1105]. Enhancements [1637, 2153, 2342]. Enhancing [2475]. enough [676]. enriched [1513, 1553, 1883]. entering [544]. Enterprise [505, 1903]. Entity [435, 1011, 1012, 2024]. Entries [2433, 1285]. Entry [88, 105, 302, 320, 355, 363, 402, 721, 722, 1033]. Environment [1398, 1455, 1510, 1561, 1562, 1719, 2486]. Environments [1009, 1551, 1185]. EOF [13]. Epilogue [37]. equations [993]. Equipment [1963, 249]. Equivalences [1484]. Error [71, 132, 2021, 2435, 79, 124, 623]. Errors [1526, 1527, 1899]. ESP [2393, 1817, 1819, 1838, 2390, 2391, 2392, 2438]. ESRO [2175]. Essential [1564]. Establishing [269]. etc [37]. ETFTP [1973]. Ethernet [1079, 1359, 816, 884, 885, 1274, 1388, 1613, 1633, 1640, 1959, 2345, 2451]. Ethernet-Like [1388, 1274, 1613, 1633, 1640, 2345]. Ethernets [1019]. Ethics [1077]. Europe [1606]. European [821]. Evaluating [2344, 2464]. Evaluation [364, 513]. evaluators [538]. Events [2433]. Evolution [1563]. example [519, 779]. Examples [2036]. Exchange [985, 1304, 1542, 1814, 1978, 2217, 2396]. Exchanging [1510, 1831]. Executive [929, 1298, 465]. Exercising [297]. Exhaustion [1325]. Existing [2275]. Expectations [1736, 2337]. Experience [264, 708, 1236, 1256, 1359, 1425, 1575, 1646, 1763, 1903]. experienced [1197]. Experiences [1296, 2110, 2496, 382]. Experiment [96, 1373, 1476, 1787, 1866, 225, 328, 435, 543, 690]. Experimental [563, 971, 1180, 1263, 1913, 885, 1229]. experimentation [1060]. Experiments [947, 1973, 879]. experts [818]. Explaining [1159]. Explicit [1846, 2392, 2468]. Exponential [1814]. Exposition [154]. Expressing [1863]. Expression [1620]. Extended [851, 1072, 1492, 2230, 2444, 688, 1375]. Extending [1369, 1783]. Extensibility [2244]. Extensible [720, 1881, 2271, 2285]. Extension [1355, 1371, 1372, 1379, 1416, 1417, 1427, 1544, 1642, 1643, 1683, 1714, 1772, 1829, 1833, 1841, 1855, 1857, 1878, 1967, 1972, 2021, 2056, 2082, 2160, 2161, 2182, 2184, 2334, 2436, 2474, 673, 727, 733, 736, 956, 1175, 1451, 1846, 2073, 2074, 2346, 2036]. Extensions [1219, 1253, 1279, 1313, 1385, 1415, 1487, 1503, 1512, 1522, 1523, 1532, 1549, 1560, 1571, 1572, 1574, 1601, 1602, 1610, 1634, 1636, 1641, 1758, 1820, 1856, 1864, 1873, 1918, 2034, 2052, 2078, 2085, 2119, 2120, 2140, 2171, 2194, 2201, 2215, 2218, 2270, 2415, 2420, 2497, 737, 978, 1038, 1044, 1062, 1074, 1102, 1210, 1266, 1331, 1511, 1512, 2032, 2033, 2034, 2035, 2064]. Extensive [2178]. Exterior [817, 878, 880, 894]. External [1822, 2004, 1004]. External-Body [1860, 2004].

F [706, 2293]. F.Y.I. [1140]. facilitate [585]. Facilities [150, 126, 263, 872, 1024]. facility [118, 218, 464, 526, 567, 662]. Facsimile [2292, 2293, 759, 788, 793, 799]. factors [192]. famine [525]. FANP [2116]. Fast [1988]. Fault [806, 982, 1019]. FAX [2146, 2291, 2288]. FDDI [1093, 1178, 1275, 1319, 1380, 1502, 2006, 2454]. Feast [525]. Feature [2376, 842]. Features [406]. February [101, 476, 832, 833, 834, 835, 836, 1626, 2117, 108, 669]. Federal [796]. Feinroth [544]. Few [611]. Field [1495, 1534, 1799, 1851, 2170, 2297, 2461, 720, 1039, 1144]. Fields [2356]. File [114, 122, 133, 141, 172, 180, 264, 276, 289, 304, 305, 322, 349, 380, 404, 409, 424, 448, 514, 529, 535, 600, 607, 617, 755, 903, 949, 968, 1058, 1084, 1225, 1304, 1430, 1527, 1751, 1854, 1973, 2191, 2288, 2289, 2293, 2376, 250, 413, 481, 495, 499, 566, 620, 653, 668, 759, 775, 1027, 1939, 260]. Files [1730, 1731, 1831, 442, 673, 787]. Filtering [1845, 2254]. Filters [1548, 1947, 2241]. Final [984]. Finger [1184, 1186, 1278, 732]. finite [679]. Firewall [1569, 2343]. Firewall-Friendly [1569]. First [97]. Five [2036]. Fixed [59, 1781]. Fixes [1526]. Flap [2426]. Flow [59, 210, 1353, 1362, 1799, 1846, 1940, 1941, 2050, 2051, 2110, 2111, 2116, 12, 436, 443, 1070]. flow-control [436, 443]. Flows [2194]. FOOBAR [1535, 1629]. Force [1254, 1007, 1381, 1428, 1529, 1708, 1975]. Form [347, 1854, 346]. Form-based [1854]. form-data [2375]. formal [894]. Format [86, 125, 787, 955, 1304, 1331, 1333, 1347, 1455, 1511, 1514, 1632, 1751, 1797, 1831, 1881, 1927, 1937, 1938, 1946, 2016, 2019, 2022, 2025, 2032, 2060, 2139, 2177, 2237, 2242, 2285, 2288, 2289, 2293, 2330, 2361, 2416, 2418, 2422, 2427, 20, 714, 723, 757, 759, 796, 812, 831, 966, 1143, 1939, 2031, 2266, 2290, 2291]. Formation [318]. Formats [80, 1978, 2502, 22, 668]. formfeed [646]. Forms [31, 2375]. Forthcoming [40]. Fortran [119]. Forum [1744]. Forwarding [1344, 1924, 1927, 2083, 2328]. Four [2035, 2403]. Fourth [1647]. FP [918]. Fragment [1845]. Frame [1284, 1305, 1480, 1576, 1586, 1594, 1960, 2102, 2414]. Frames [2084, 1216].

Framework [1375, 1447, 1919, 2317, 2369, 2373, 2498, 458, 477, 697, 1431, 1442, 1895, 1975]. Frameworks [2248, 2258]. Framing [1539, 1652]. Free [481, 485, 495]. Frequency [1667]. Frequently [1852, 1928]. Friendly [1474, 1569, 1771]. Front [695, 919, 640]. Front-end [695]. FTAM [1405]. FTP [238, 407, 409, 432, 442, 452, 457, 462, 469, 472, 473, 474, 500, 514, 526, 544, 564, 586, 623, 633, 681, 687, 727, 733, 741, 765, 939, 1405, 1535, 1569, 1625, 1629, 2215, 2415]. FTP-FTAM [1405]. FTPSRV [673]. full [1127, 1850, 1953]. Function [1565, 1129]. Functional [8, 1287, 1634, 1726, 1727, 2192, 928, 969, 1116]. Functions [2129]. FUNI [2350]. Further [603, 409]. Future [1277]. FYI [1137, 1165, 1167, 1188, 1196, 1197, 1315, 1452, 1453, 1460, 1568, 1584]. FZA [1980].

Gandalf [1980]. Gateway [817, 878, 880, 888, 894, 1018, 1064, 1095, 1153, 1154, 1244, 1257, 1258, 1259, 1387, 1405, 1644, 1645, 1647, 1761, 1762, 813, 901]. Gatewaying [1496]. Gateways [865, 1334, 2467, 975, 999, 1017]. General [1520, 1974, 2202, 2284, 2452]. generated [1214]. Generation [1540, 1716, 1742]. Generator [423, 854, 962]. Generic [1498, 1499, 1691, 1692, 2065, 2383, 2460, 2466, 1219]. generic-interface [1219]. Geographical [1702]. Getting [64]. Gigabit [1206]. Given [1517]. GKMP [2080, 2081]. Glitch [107, 102]. Global [1263, 2150, 2361]. Glossary [1198, 1382, 1970]. Go [588, 589, 848, 1639]. Go-Ahead [588, 589]. GO-MHS [1639]. Goals [1116, 1263]. Going [541]. Gold [1280, 1392]. Good [1804, 2002]. Gopher [1426]. GORDO [11]. GOSIP [1159]. Government [1801, 1806, 2133]. GPS [1996]. GPS-Based [1996]. Grant [2504, 1812, 1975]. Graphic [331, 178, 184]. graphical [177, 955]. Graphics [86, 94, 125, 126, 134, 174, 191, 192, 246, 253, 263, 277, 287, 309, 346, 382, 391, 487, 542, 2070, 186, 199, 225, 232, 280, 466, 546, 736]. GRE [1691]. Greek [1934]. Group [85, 87, 164, 207, 234, 245, 309, 311, 318, 391, 577, 1099, 1491, 1563, 1593, 1669, 1744, 1758, 2001, 2080, 2081, 2220, 2223, 2405, 2452, 2453, 514, 578, 888, 1142, 101, 542, 1386, 1630, 1680, 2472]. Groups [1679, 956, 982]. Growth [1286]. GSS [2466, 1948, 1951, 2012, 2465]. GSS-API [1948, 1951, 2012, 2465]. Guaranteed [2199, 2201, 2368]. Guide [1381, 1529, 1570, 1708, 2059, 2347, 100, 269, 990, 998, 1022, 1023, 1108]. Guidelines [972, 976, 1059, 1227, 1252, 1271, 1356, 1357, 1374, 1456, 1552, 1576, 1593, 1607, 1619, 1699, 1842, 1917, 2001, 2195, 2310, 2366, 2405, 2421, 2037]. GVB [68]. GZIP [1939].

H [918, 1705]. H-FP [918]. H.261 [2019]. H.263 [2177]. Handbook [1234, 2183, 2488, 756, 764]. handbooks [185]. Handling [1320, 1546, 2299, 115, 178]. Handshake [1981]. HAP [1211]. Hard [315].

Hardware [1307, 1308, 1649, 1650, 1923, 2312, 816]. Harmful [1253, 1617]. Harvard [493]. Hashing [2091]. Haskins [627]. Hazards [1327]. HDLC [1539, 1652]. HDLC-like [1652]. Header [1495, 1512, 1534, 1612, 1796, 1816, 1851, 2034, 2170, 2297, 2356, 2389, 2411, 2490, 2492, 876, 1039, 1144]. Headers [554, 1332, 1543, 2063, 2461, 2491, 1134]. Heated [2312]. Hebrew [1545]. heffalumps [865]. held [798, 2117]. Helminthiasis [1125]. HEMP [1012]. HEMS [1011, 1013, 1014, 1066]. Hewlett [1975]. hierarchy [1826, 2281]. High [2504, 25, 697, 1011, 1012, 1313, 1067, 1142, 1175]. High-level [697, 1011, 1012]. high-speed [1175]. High-Tech [2504]. HIPPI [1364, 2054]. Historic [1910, 89]. Historical [1596]. History [2480]. Hit [2214]. Hit-Metering [2214]. Hitchhikers [1108]. HMAC [2072, 2091, 2189, 2273, 2390, 2391]. HMAC-MD5 [2072, 2189]. HMAC-MD5-96 [2390]. HMAC-RIPEMD128 [2273]. HMAC-RIPEMD160 [2273]. HMAC-SHA-1 [2189]. HMAC-SHA-1-96 [2391]. Hobbe [2222]. Hoc [1099]. Hold [2066]. home [1985]. Homed [2257, 2247]. Hop [2319]. hope [284]. horizontal [644, 645]. Host [1, 2, 7, 9, 11, 12, 17, 22, 33, 65, 67, 102, 106, 109, 136, 142, 209, 239, 258, 262, 275, 288, 300, 306, 307, 310, 312, 314, 321, 325, 327, 337, 339, 356, 357, 361, 365, 371, 389, 399, 446, 461, 474, 540, 584, 587, 590, 596, 599, 601, 604, 651, 677, 680, 682, 685, 686, 694, 704, 711, 792, 841, 859, 868, 897, 919, 956, 978, 995, 1044, 1102, 1117, 1211, 1504, 1521, 1531, 1536, 1671, 1672, 1758, 2062, 2118, 112, 226, 229, 233, 236, 237, 247, 252, 261, 268, 283, 284, 293, 301, 303, 348, 362, 387, 460, 550, 613, 616, 640, 657, 742, 744, 800, 839, 942, 1006, 1163, 18, 107]. Host-dependent [474]. Host-Front [919]. Host-Host [11, 33, 102, 142, 312, 461, 685, 704, 22, 18, 107]. Host-IMP [7, 17, 587]. Host-to-Host [711]. Host/Host [65]. Host/IMP [67, 209, 677, 680, 686, 694]. Hostname [943]. Hostnames [801, 618, 620]. Hosts [282, 405, 541, 1516, 1920, 2087, 15, 110, 255, 1112, 1113]. Hot [2268]. HPN [1669]. HPR [2225, 2340]. HSRP [2268]. HTCPCP [2311]. HTCPCP/1.0 [2311]. HTML [1854, 1929, 1967, 2097]. HTTP [1932, 2055, 2056, 2096, 2132, 2156, 2214, 2282, 2283, 2497]. HTTP/1.0 [1932]. HTTP/1.1 [2055]. Human [1741]. Human-Readable [1741]. Humanities [2137]. hung [595]. Hyper [2311]. HYPERchannel [1034, 1213]. Hypertext [2057, 1853, 1932, 2055]. HZ [1831].

IAB [1042, 1073, 1090, 1120, 1130, 1164, 1190, 1240, 1241, 1270, 1326, 1348, 1350, 1361, 1391, 1400, 1471, 1591, 1626, 1849, 2014, 2117, 2269, 2303]. IANA [2503, 1904, 2265, 2421, 2455]. IBGP [1953]. IBM [110, 1812]. iCalendar [2432, 2433, 2434]. ICCC [367, 415]. ICMP [834, 1246, 1778]. ICMPv6 [2450, 1872, 2453]. ICP [123, 143, 151, 161, 202, 215, 393, 2173, 2174]. ID [527, 1860, 2098, 2379]. ideas [66]. Identification [1403, 1404, 1756, 2310, 917, 1055, 1145]. Identifier [1164, 2473]. Identifiers [1413, 1429, 1516, 2061, 2155, 2275, 1105, 1620, 2066, 2383]. IDENTITIES [2494]. Identity [1814]. Identity-Based [1814]. idents [379]. IDLE [2164].

IDNA [2500]. IDP [1754]. IDPR [1467]. IDRP [1735, 1850]. IDUP [2466]. IDUP-GSS-API [2466]. IEEE [938, 1032, 1220, 1221, 1358, 1505, 1506, 1733, 1738, 1739, 2007, 2095, 2226, 2253]. IEPG [1680]. IESG [1241, 1326, 1361, 1370, 1425, 2014, 2269]. IETF [1381, 1529, 1593, 1708, 2015, 2018, 2197, 2264, 2310, 2326, 2344, 2405, 2423, 2425]. IETF-ISOC [2018]. IFIP [818]. IGP [1361]. II [472, 1104, 1148, 1180, 1203, 1412, 1563, 1732]. III [109, 1105, 1413]. IKE [2396]. ILL [2487]. Illinoi [466]. Illinois [101, 156, 593]. Illusion [863]. Image [1967, 2145, 2289, 2293, 520]. image/tiff [2289]. Images [1304]. IMAP [2082, 2179, 2182]. IMAP/POP [2082, 2182]. IMAP2 [1722]. IMAP2BIS [1722, 2048]. IMAP4 [1721, 1722, 1723, 2048, 2073, 2074, 2075, 2164, 2167, 2180, 2208, 2329, 2346]. iMIP [2434]. IMLAC [190]. IMP [7, 12, 17, 18, 41, 67, 209, 213, 215, 218, 266, 296, 307, 326, 330, 338, 354, 369, 389, 401, 428, 439, 441, 470, 515, 522, 541, 587, 604, 615, 619, 626, 631, 651, 677, 680, 682, 686, 694]. IMP-Host [12, 389]. IMP-IMP [18, 41]. IMP-to-Host [307]. IMP/360 [330]. IMP/Host [604, 651, 677, 682, 686, 694]. IMP/TIP [428, 439, 441, 470, 615, 626, 631]. implement [1017]. Implementation [11, 103, 552, 708, 873, 998, 1005, 1263, 1295, 1359, 1507, 1526, 1646, 1672, 2366, 2367, 55, 166, 191, 194, 206, 215, 586, 635, 678, 774, 807, 880, 887, 911, 1025, 1033, 1049, 1109]. Implementations [1282, 1622, 1944, 2103, 2181, 2445, 866]. Implementer [2153, 2340]. Implementers [573]. Implementing [1923, 382, 1133]. Implementors [2385]. Implications [1334, 1995]. improve [604]. improved [376, 839]. Improvement [210, 653]. improvements [629]. IMPs [390]. IMPSYS [256, 436, 443]. IN-ADDR [2304]. Inc. [1780, 2326]. Inches [1695]. Incident [2337]. Incorporation [2121]. Increased [2402]. Increasing [2401]. Incremental [1131, 1614, 1982]. Independent [2104, 2349, 2433, 177, 1795, 2466]. Index [1900, 591]. Indicate [2106]. Industry [2504, 1664, 1676]. INETPhone [1779]. infinite [960]. INFN [1666]. Information [81, 96, 115, 223, 281, 294, 347, 796, 1048, 1056, 1085, 1146, 1148, 1152, 1179, 1184, 1186, 1203, 1204, 1233, 1238, 1242, 1243, 1261, 1275, 1278, 1292, 1305, 1333, 1345, 1378, 1385, 1432, 1440, 1441, 1446, 1487, 1502, 1510, 1514, 1598, 1604, 1618, 1679, 1686, 1687, 1713, 1717, 1719, 1732, 1747, 1792, 1796, 1824, 1826, 1837, 1849, 1863, 1889, 1894, 1903, 1998, 1999, 2000, 2008, 2102, 2114, 2115, 2170, 2193, 2200, 2201, 2229, 2281, 2294, 2359, 2412, 2439, 2441, 2452, 2453, 2493, 263, 346, 466, 525, 652, 684, 746, 970, 1038, 1055, 1074, 1145, 1165, 1187]. Infrastructure [1292, 1849, 1896, 2446, 1200]. Ingress [2254]. Initial [165, 197, 2401, 2402, 206, 1162, 93, 145]. Initiated/Unsolicited [1430]. Input [71, 331, 1660, 1669, 563]. Institute [1663]. Institutions [1349]. instruction [308]. Instructions [1101, 1533, 2210]. Integer [340]. Integrated [1287, 1623, 1717, 2197, 2200, 2201, 2202, 2369]. Integration [494, 1811]. Integrity [1795]. Intelligence [152]. Inter [435, 1312, 1468, 1469, 1507, 1509, 2445, 2498, 1115, 1116]. inter-autonomous [1116]. Inter-Domain [1312, 1468, 1469, 1507, 1509, 2498]. Inter-Entity [435]. Inter-Networking [2445]. Interact [1901]. Interaction [1354, 1393, 1735, 432, 472]. Interactions [712]. Interactive [96, 1054, 1166, 1193]. interagency [1005]. Interchange [51, 968, 20, 840, 966, 1026, 1446, 1852]. Interconnect [1284, 1346, 1451, 1480, 1931, 2414]. Interconnected [1460, 1137]. interconnecting [1114]. Interconnection [1029, 1232]. Interdependence [2127]. Interest [340, 577, 578, 888]. Interface [330, 1218, 1222, 1223, 1274, 1294, 1388, 1396, 1397, 1498, 1582, 1585, 1613, 1633, 1640, 1813, 2065, 2120, 2130, 2345, 2481, 2482, 2483, 7, 12, 209, 258, 404, 587, 651, 735, 776, 928, 1195, 1219, 2466]. Interfaces [1563, 1684, 2007, 2220]. Interfacing [593]. Interim [189, 706]. Interior [1064]. Intermail [1158]. Intermediate [985, 1152, 1228, 1471]. Internal [1287]. International [624, 2333, 818, 366]. Internationalization [2057, 2471]. Internationalized [2500]. Internet [768, 892, 953, 1254, 1315, 1322, 1331, 1375, 1386, 1426, 1452, 1511, 1512, 1568, 1584, 1626, 1700, 1849, 1864, 2032, 2033, 2034, 2035, 2036, 2064, 2143, 2358, 2359, 1720, 1770, 1790, 1867, 2037, 2047, 2187, 2287, 2387, 2504, 665, 743, 749, 750, 756, 764, 767, 771, 781, 782, 789, 800, 809, 812, 813, 820, 879, 891, 904, 907, 909, 912, 914, 926, 928, 930, 934, 937, 940, 942, 951, 956, 975, 976, 979, 981, 987, 999, 1001, 1005, 1007, 1010, 1030, 1034, 1039, 1042, 1052, 1059, 1060, 1061, 1092, 1097, 1098, 1103, 1104, 1105, 1107, 1108, 1110, 1112, 1113, 1118, 1119, 1125, 1126, 1131, 1142, 1144, 1150, 1154, 1156, 1163, 1164, 1167, 1179, 1180, 1182, 1196, 1197, 1204, 1216, 1227, 1231, 1241, 1252, 1254, 1258, 1262, 1263, 1264, 1265, 1266, 1271, 1277, 1286, 1300, 1303, 1304, 1325, 1326, 1331, 1332, 1334, 1337, 1339, 1348, 1349, 1359, 1361, 1381, 1382, 1384, 1390, 1411, 1412, 1413, 1414, 1418, 1420, 1422, 1428, 1431, 1442, 1447, 1449, 1453, 1457, 1458, 1465]. Internet [1466, 1490, 1495, 1496, 1511, 1529, 1530, 1537, 1545, 1547, 1550, 1564, 1590, 1591, 1600, 1610, 1614, 1619, 1623, 1630, 1645, 1654, 1680, 1694, 1697, 1708, 1710, 1717, 1719, 1726, 1729, 1734, 1762, 1765, 1777, 1779, 1780, 1792, 1796, 1801, 1806, 1809, 1815, 1830, 1852, 1870, 1872, 1895, 1898, 1904, 1907, 1909, 1922, 1923, 1942, 1945, 1970, 1971, 1975, 1987, 1995, 1998, 2013, 2023, 2032, 2044, 2049, 2063, 2121, 2122, 2133, 2135, 2137, 2138, 2149, 2150, 2170, 2173, 2174, 2205, 2222, 2223, 2224, 2245, 2288, 2290, 2291, 2292, 2296, 2326, 2347, 2364, 2373, 2388, 2394, 2395, 2396, 2408, 2432, 2445, 2446, 2447, 2450, 2486]. Internet-IP [976, 1059]. Internet-Standard [1975, 1431, 1442, 1895]. Internets [1460, 1587, 1905, 1055, 1056, 1075, 1137, 1145, 1146, 1148, 1203]. Internetwork [1542, 2308]. Internetworking [1453, 1165, 1699]. internetworks [886]. Interoperability [1561, 1673, 2433, 2434]. Interoperation [1524, 2368]. Interpretation [2132, 2394]. interpreter [194]. Interprocess [61, 62]. Interrupt [103]. Interval [1774, 2336]. Intervals [2480]. Intra [1132, 2324]. Intra-domain [1132]. Intra-LIS [2324]. Introduced [1006]. Introducing [1453, 1680, 1792]. Introduction [918, 1140, 1298, 1301, 1431, 1674, 1888]. Introductory [1453]. Inverse [1283, 2377]. Invitation [157]. Involved [2015]. Involving [399]. IP [2037, 2452, 2453, 771, 805, 867, 884, 885, 886, 938, 976, 978, 1015, 1028, 1032, 1036, 1041, 1044, 1045, 1053, 1059, 1075, 1078, 1085, 1093, 1102, 1134, 1137, 1139, 1170, 1178, 1185, 1191, 1199, 1224, 1226, 1263, 1344, 1355, 1356, 1357, 1361, 1364, 1365, 1373, 1380, 1383, 1444, 1456, 1459, 1460, 1463, 1508, 1528, 1540, 1567, 1596, 1598, 1616, 1621, 1706, 1716, 1729, 1735, 1742, 1744, 1745, 1776, 1802, 1811, 1816, 1817, 1818, 1839, 1840, 1845, 1871, 1873, 1904, 1906, 1913, 1919, 1957, 1968, 1989, 1990, 1991, 1992, 1993, 2010, 2054, 2062, 2072, 2083, 2100, 2130, 2138, 2157, 2212, 2213, 2229, 2254, 2307, 2309, 2317, 2318, 2323, 2324, 2331, 2338, 2340, 2343, 2352, 2360, 2378, 2380, 2381, 2382, 2389, 2393, 2394, 2398, 2419, 2439, 2441, 2448, 2459, 2468, 2477, 2490, 2491, 2492]. IP-ATM [1811]. IP-based [1055, 1056, 1145, 1146, 1148, 1203]. IP-OSPF [1735]. IP/TCP [886]. IP/UDP/RTP [2491]. IPC [150]. IPComp [2380]. IPCP [1322, 2277]. IPNG [1942, 1540, 1657, 1658, 1659, 1660, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1668, 1669, 1670, 1672, 1673, 1676, 1677, 1678, 1695, 1709, 1716, 1743]. IPOA [2307]. IPOA-MIB [2307]. IPPM [2485]. IPSEC [2194, 2397]. Ipsilon [1941, 1974, 2284, 1940]. IPv4 [1692, 1760, 1847, 1865, 1940, 1941, 2017, 2088, 2163, 2277, 2461]. IPv6 [2447, 2448, 2450, 1799, 1868, 1870, 1872, 1874, 1875, 1884, 1911, 1920, 1957, 1958, 1959, 2006, 2017, 2060, 2067, 2120, 2134, 2172, 2279, 2415, 2449, 2451, 2454, 2457, 2458, 2460, 2461, 2484, 2361, 2362, 2478, 2479]. IPX [1122, 1224, 1288, 1352, 1410, 1541, 1543, 1624, 1781, 1782]. IPXCP [1542]. IPXWAN [1352, 1541, 1624]. IRSG [1241, 1326]. IRTF [2001]. IS-IS [1132, 1185]. ISAKMP [2395, 2394]. ISDN [1346, 1608, 2114]. ISI [399, 774, 775, 776]. ISO [1128, 1138, 1317, 882, 895, 972, 973, 976, 985, 996, 997, 998, 1059, 1075, 1076, 1129, 1152, 1228, 1544, 1551, 1565, 1805, 1846, 2031, 2113, 2266, 2487]. ISO-10646 [1805]. ISO-10646-J-1 [1805]. ISO-2022-JP [1544]. ISO-2022-JP-2 [1544]. ISO-TP0 [1076]. ISOC [2018, 2423]. ISOC/IETF [2423]. Isochronous [1247]. isolation [806]. Issue [1471]. Issued [14, 26, 92, 159, 201, 220, 244, 248, 257, 267, 270, 272, 274, 279, 332, 336, 353, 370, 375, 378, 392, 419, 422, 438, 459, 475, 478, 496, 501, 511, 530, 534, 547, 551, 557, 565, 568, 576, 598, 632, 634, 639, 641, 655, 658, 660, 663, 666, 672, 683, 699, 700, 701, 705]. Issues [993, 1345, 1561, 2108, 146, 429, 1067, 1114]. Iteration [391]. iTIP [2433]. ITOT [2113]. ITU [2416, 2423]. ITU-T [2416, 2423]. IV [1279, 1549, 1366, 1414, 1752, 2392]. IX [1465].

J [1805]. January [364, 449, 798]. JANUS [735]. Japanese [1458, 2224]. jitter [1247]. jitter-controlled [1247]. Job [88, 105, 278, 302, 320, 355, 363, 402, 471]. Jon [2504, 2428]. Journal [473, 622, 2433, 536, 621]. JP [1544, 1544]. JPEG [2022, 2422].

JPEG-compressed [2022, 2422]. July [373, 542, 549, 693, 1200, 531, 539]. Jumbograms [2134]. June [532]. junk [696]. Just [1418]. Just-Send-8 [1418].

K-12 [1699]. Kahn [6, 367]. kbit [2409]. kbit/s [2409]. Kerberos [1401, 1500, 1951]. Kernel [1579]. Key [1412, 1414, 1814, 1936, 2012, 2080, 2081, 2106, 2217, 2354, 2395, 2396, 2399, 2446, 1104, 1160, 1750]. Key-Exchange [1814]. Keyed [1818, 1839, 2091]. Keyed-Hashing [2091]. Keys [103, 1741]. known [317]. KOI8 [2306]. Korean [1547]. KRFC [1303].

L2F [2328]. Lab [627]. Label [1447, 1799]. Labelled [1941]. Laboratory [109]. Labs [198]. Laid [149]. LAN [862, 915, 1019, 1829, 2272]. Language [5, 51, 83, 242, 1446, 2057, 2469, 195, 1013, 1066, 1853, 2267]. Language-machine [83]. Languages [1756, 2171, 2218, 2264]. LAPB [1371]. Large [1668, 1677, 1820, 2486, 1201]. Latest [609]. Laws [2122]. Layer [1367, 1372, 1473, 1688, 2328, 2456, 925, 931, 1060, 2209]. Layers [1267, 1443, 1112]. LBL [544]. LCP [1560, 2471]. LDAP [1548, 1813, 1946, 1947, 2151, 2234, 2241, 2242, 2294]. LDAP/X.500 [2234]. LDAPv3 [2243]. Leaving [676]. Left [1695]. Legal [2227, 2339]. Length [13, 1847, 1865]. less [1788]. Letter [360, 381, 544]. letters [233]. Level [56, 109, 287, 331, 382, 445, 656, 1263, 1281, 2274, 66, 88, 697, 1011, 1012]. Levels [2106]. Liaison [384, 418]. liaisons [414]. Library [281, 1681]. licenses [1160]. Light [2111]. Light-weight [2111]. Lightweight [1477, 1767, 1788, 2238, 2239, 2240]. Like [1388, 640, 1274, 1307, 1308, 1613, 1633, 1640, 1649, 1650, 1652, 2345]. Limitations [2486]. Limiting [2214]. Lincoln [109, 198]. Line [74, 216, 315, 345, 563, 1169, 121, 537, 599, 601, 616, 618, 635]. linefeed [649]. Linemode [1106, 1174]. lines [1045]. lingual [2076]. Link [25, 104, 179, 312, 412, 1297, 1323, 1424, 1445, 1461, 1737, 1785, 1924, 1976, 2011, 2093, 2101, 2456, 597, 925]. Link-Layer [2456]. linking [569]. Links [18, 1321, 1941, 1973, 2491, 352, 1036, 1124, 1134, 1161]. LIS [2324]. List [69, 84, 170, 851, 2356, 52, 108, 160, 182, 200, 284, 427, 497]. listings [1285]. Lists [324, 397, 155, 168, 211, 295, 298, 358, 1201]. Listserv [1419]. literals [2075]. little [838]. LL [446, 460]. LL-67 [460]. LLC1 [1213]. Load [1784, 2198, 2368, 2378, 539]. Loader [899, 186]. loading [896]. Local [814, 1459, 1924, 2020, 269, 881]. local-network [881]. Local/Remote [1924]. Locating [537]. Location [877, 1702, 1863, 2152, 769, 936, 1086, 2039, 1795]. Location-Independent [1795]. Locators [1728, 1798, 2043, 2098, 2379, 1726]. lockup [619]. Logger [56, 79, 98, 460]. logger/server [460]. Logical [58, 426]. Login [394, 425, 2208]. Logout [394, 425, 717]. London [581]. Long [1792, 1062]. long-delay [1062]. Look [305]. Lookup [1824]. lose [738]. Lost [510, 528, 654, 506]. Low [2491, 1134]. Low-Speed [2491, 1134]. Lower [1267]. LSA [2357]. LSNAT [2378]. LUname [1636]. LZS [1954, 1961, 2382]. LZS-DCP [1954].

M.I.T. [16, 32]. MAC [1319]. Machine [575, 83, 679]. Macintosh [1730]. MacMIME [1730]. macro [719, 725]. MACSYMA [571]. Magnalink [1962]. MAIL [1723, 196, 221, 273, 452, 518, 554, 570, 710, 761, 762, 770, 774, 775, 776, 778, 811, 964, 1054, 1158, 1166, 1193, 1329, 1331, 1333, 1334, 1411, 1412, 1413, 1414, 1418, 1496, 1511, 1512, 1514, 1556, 1632, 1654, 1879, 1880, 1898, 1934, 2020, 2032, 2033, 2034, 2035, 2036, 2064, 2139, 2236, 2290, 2291, 2292, 2356, 2408, 447, 450, 463, 469, 492, 533, 548, 567, 623, 637, 696, 741, 744, 763, 795, 797, 798, 900, 905, 966, 974, 979, 983, 1030, 1046, 1103, 1104, 1105, 1395, 2097, 2149, 1395, 2149, 1701, 1810, 1832]. Mail-11 [1395, 2149]. Mail-Safe [1632, 2139]. Mailbox [224, 2129, 2167, 2180, 504]. mailboxes [753]. Mailing [324, 397, 155, 168, 211, 295, 298, 358, 1201, 1347]. mailto [2355]. Maintaining [1375]. Maintenance [667, 434, 439, 626, 631, 1201]. make [508]. make-work [508]. Making [2333]. Managed [1259, 1274, 1276, 1294, 1306, 1307, 1308, 1343, 1358, 1388, 1396, 1397, 1461, 1462, 1463, 1464, 1483, 1505, 1506, 1515, 1585, 1586, 1594, 1613, 1633, 1640, 1647, 1648, 1649, 1650, 1655, 1656, 1684, 1685, 1737, 1993, 2007, 2011, 2038, 2095, 2142, 2219, 2225, 2226, 2253, 2274, 2307, 2312, 2345, 2353, 2404, 2442, 2443, 2444, 2481, 2482, 2483, 2493, 2495, 1222, 1223]. Management [173, 286, 294, 311, 1011, 1012, 1056, 1057, 1085, 1088, 1099, 1146, 1147, 1148, 1152, 1179, 1203, 1204, 1218, 1233, 1238, 1242, 1243, 1275, 1305, 1356, 1357, 1412, 1431, 1432, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1441, 1442, 1456, 1460, 1502, 1582, 1618, 1639, 1685, 1686, 1687, 1732, 1734, 1747, 1837, 1868, 1889, 1890, 1891, 1892, 1893, 1894, 1895, 1940, 1974, 1975, 1998, 1999, 2000, 2008, 2026, 2080, 2081, 2094, 2096, 2102, 2114, 2115, 2193, 2200, 2201, 2223, 2248, 2249, 2251, 2252, 2258, 2259, 2261, 2262, 2284, 2296, 2309, 2354, 2395, 2439, 2441, 2494, 2495, 188, 299, 494, 772, 1042, 1055, 1104, 1137, 1145, 1204, 1261, 2452, 2453]. Manager [1441]. Manager-to-Manager [1441]. Managers [1493, 1163]. Managing [2107, 1214]. manuals [185]. Many [1671]. map [426]. MAPOS [2158, 2159, 2160, 2161, 2162, 2163]. Mapping [333, 977, 1016, 1127, 1128, 1138, 1317, 1395, 1485, 1654, 2076, 2143, 2144, 2149, 2338, 183, 563, 1226, 1828, 2151, 2150]. Mappings [1439, 1893, 785, 786]. Maps [1967, 609]. MARC [2207]. March [364, 491, 2117, 296]. Mark [850]. Market [1659]. Marking [67, 64, 923]. Markup [1853, 2057]. MARS [734, 2108, 2136, 2256, 2430]. MARS-based [2136]. Materials [1994]. Mathematical [377]. Mathematics [1009]. MATHLAB [519, 571]. Matter [1780, 2326]. MAUs [1505, 2226]. MaXIM [2149]. MaXIM-11 [2149]. maximum [869]. Maxwell [466]. May [116, 131, 140, 329, 516]. MCGAM [2150]. MD2 [1309]. MD4 [1176, 1310]. MD5 [2390, 1311, 1534, 1800, 1818, 1851, 2069, 2072, 2189, 2372]. Mean [612]. Means [1863]. Measured [1118]. Measurement [318, 387, 1252, 1263, 2050, 2051, 2110, 566]. measurements [550]. Measuring [2485]. Mechanism [142, 1333, 1514, 1951, 1983, 2012, 2096, 2232, 2431, 2436, 2465, 53, 2376]. Mechanisms [1331, 1511, 1721, 1920, 2475]. Media [1352, 1541, 1543, 1580, 1610, 1624, 1861, 2033, 2305, 2341, 2363, 2429, 757]. Medium [1427, 2226, 1505]. Meeting [35, 37, 43, 63, 75, 82, 99, 173, 286, 309, 391, 447, 1578, 21, 77, 85, 87, 101, 108, 116, 131, 134, 140, 164, 188, 207, 212, 232, 234, 245, 246, 253, 277, 319, 448, 450, 455, 463, 468, 477, 520, 531, 542, 578, 624, 795, 797, 798, 888, 900]. Meets [1852, 519]. Members [1241, 1326]. Memo [514, 1207]. Memorandum [450]. Memory [68, 428, 441, 470]. Mesh [1901, 1850, 1953]. MESSAGE [1720, 2047, 13, 31, 42, 58, 292, 328, 527, 670, 703, 734, 743, 749, 767, 782, 979, 1030, 1035, 1149, 1176, 1302, 1309, 1310, 1311, 1320, 1331, 1332, 1411, 1417, 1484, 1485, 1511, 1512, 1643, 1766, 1857, 1860, 1872, 1881, 1972, 1978, 2032, 2034, 2049, 2063, 2091, 2098, 2144, 2196, 2249, 2259, 2285, 2298, 2302, 2356, 2379, 2410, 2427, 2434, 2450, 2463, 22, 506, 510, 528, 541, 619, 654, 747, 796, 831, 876, 924, 1087, 1103, 1143, 1194]. Message-Based [2434]. Message-Digest [1309, 1310, 1311]. Message-ID [527, 2098, 2379]. Message/External [1860]. Message/External-Body [1860]. Messages [23, 1246, 1458, 1495, 1545, 1547, 1778, 1796, 1820, 1830, 1879, 1909, 1934, 2170, 2224, 582, 625, 714, 723, 812, 840, 1026, 1037, 1039, 1144, 1486]. Messaging [2338]. Meta [2356]. Meta-Syntax [2356]. Metadata [2400]. Meter [2051]. Metering [2214]. Method [1948, 1959, 1133]. Methodology [1931]. methods [938, 991]. Metrics [1421, 1554, 2317, 2485]. Meyer [50]. MHS [1455, 1639, 1791]. MHTML [2097]. MIB [1148, 1202, 1203, 1217, 1219, 1220, 1221, 1228, 1279, 1344, 1359, 1371, 1372, 1379, 1404, 1437, 1451, 1503, 1504, 1549, 1555, 1556, 1557, 1563, 1583, 1601, 1602, 1686, 1687, 1714, 1733, 1738, 1739, 1749, 1782, 2024, 2051, 2061, 2083, 2220, 2235, 2236, 2307, 2425, 2480]. MIB-II [1148, 1203, 1563]. MIBs [1229, 2026]. Microsoft [2105, 2420]. Microsystems [1780, 2326]. Mid [1281]. Mid-Level [1281]. Migrating [1605]. MIL [1943]. Military [954, 997, 953]. MILNET [1021]. MIME [1331, 1334, 1418, 1427, 1486, 1511, 1512, 1513, 1546, 1553, 1631, 1730, 1731, 1757, 1810, 1820, 1832, 1835, 1836, 1859, 1883, 1914, 2002, 2004, 2032, 2033, 2034, 2035, 2036, 2097, 2099, 2144, 2146, 2147, 2148, 2171, 2218, 2289, 2298, 2299, 2374, 2409, 2410, 2411, 2412, 2413, 2467, 2487, 2143]. MIMEtransport [1416, 1642]. Minded [122, 404]. minicomputer [230]. minicomputer-based [230]. minicomputers [624]. Minimal [1877, 1991, 2290, 2291]. Minor [718]. Minute [2480]. Minutes [164, 542]. MIT-DMS [571]. MIT-MATHLAB [519, 571]. MITRE [316]. MIX [467, 479, 488]. MIXAL [467, 479, 488]. Mixed [340, 1831]. MIXER [2143, 2150, 2151]. MLC [241]. MLFL [741]. MLTNET [334]. Mnemonics [1335, 226, 239]. Mobile [2028, 2277, 2331, 2343, 2499]. Mobile-IPv4 [2277]. Mobility [1678, 1989, 1992, 1993, 2090]. Mode [976, 1059, 1346, 2104, 2292, 2324, 2349, 916, 984, 2438]. Model [1341, 1394, 1435, 1579, 1843, 1844, 1897, 2064, 2256, 2477, 772, 861, 1126, 2251, 2252, 2261, 2262]. Modeling [1657]. MODELS [1723, 1094]. Modem [1686]. Modernization [1390]. Modes [96, 1413, 184, 1105]. Modification [312]. Modifications [181, 555]. modifying [569]. Modularity [807]. Modules [2480]. moments [89]. monitor [613]. Monitoring [859, 1018, 1261, 1323, 1460, 1503, 1555, 1556, 1557, 1747, 1976, 2008, 2061, 2235, 2236, 1013, 1066, 1137]. Moratorium [72]. MOSPF [1575]. move [602]. MP [1707, 1921, 1977]. MPEG [2330]. MPEG1 [2025, 2237]. MPEG1/MPEG2 [2025, 2237]. MPEG2 [2025, 2237]. MPP [1194]. MPPC [2105]. MPSX [340]. MSDTP [703]. MSDTP-Message [703]. MTAs [2489]. MTP [776]. MTP-NIMAIL [776]. MTU [1053, 1181, 1425, 1616, 1781, 1968]. Muddle [571]. muddling [495]. Multi [662, 915, 937, 1019, 1321, 1455, 1760, 1777, 1985, 2167, 2247, 465, 757, 1124, 1161, 334]. Multi-Accessed [2167]. Multi-Destination [1760]. Multi-home [1985]. Multi-homed [2247]. Multi-LAN [915, 1019]. multi-media [757]. Multi-network [937]. Multi-protocol [1321, 1124, 1161]. Multi-provider [1777, 2247]. Multi-site [662, 465]. Multicast [1065, 1291, 1448, 1459, 1574, 1936, 2009, 2077, 2089, 2104, 2136, 2176, 2188, 2344, 2349, 2352, 2353, 2362, 2404, 2419, 2477, 2486, 956, 2324]. Multicast-Sparse [2104, 2349]. Multicasting [1758, 978, 1044, 1102, 2136]. Multics [653, 406, 444, 583]. Multilingual [1544]. Multilink [1707, 1921, 1977]. Multimedia [797, 900, 1333, 1514, 1604, 1810, 1832, 1187]. Multipart [1835, 1859, 1879, 2099, 2374, 2375]. Multipart/Encrypted [1835]. multipart/form [2375]. multipart/form-data [2375]. Multipart/Related [1859, 2099, 2374]. Multipart/Report [1879]. Multipart/Signed [1835]. Multiparts [1835, 2467]. Multiple [23, 2158, 2162, 2478, 971]. multiple-path [971]. Multiplexer [1068]. Multiplexing [1682]. Multiprotocol [1284, 1346, 1451, 1473, 1480, 1673, 1744, 2270, 2414, 1550]. Multipurpose [1331, 1511, 1512, 2032, 2033, 2034, 2035, 2036, 2064]. munging [876]. must [192]. Mutual [1316]. MUX [1217]. my [420].

NADF [1407, 1748]. NAK [1096]. Name [804, 1454, 1581, 1778, 1863, 1864, 1997, 2052, 2123, 2124, 2155, 2227, 2229, 2332, 76, 129, 616, 746, 789, 820, 887, 911, 1021, 1168, 732, 2263]. NAME/FINGER [732]. named [939]. Names [275, 1475, 1620, 1675, 1727, 1769, 1801, 1806, 2129, 2133, 2234, 2240, 2275, 2332, 2339, 2500, 229, 236, 237, 247, 268, 284, 599, 601, 871, 872, 873, 1024, 1025, 1091]. Nameserver [2365]. Namespace [2329]. Naming [1208, 1245, 1374, 1474, 1488, 1552, 1607, 1771, 2087, 2107, 2364, 500, 747, 809]. NARP [1725]. NASA [413]. NAT [1621]. national [1249, 1157]. Native [1933]. NATs [2415]. NAUs [1655, 1656]. NBFCP [2084]. NBMA [1725, 2256, 2319, 2478]. NCACHE [2295]. NCC [296]. NCP [55, 60, 78, 215, 383, 420, 454, 543, 611, 679, 763, 791]. NCP/TCP [763, 791]. Near [1611]. Near-term [1611]. Necessary [2470]. Neda [2175]. Needs [1887, 456]. Negative [2295]. Negotiation [1775, 2282, 2465, 1133, 2376]. Neighbor [1957, 2448]. net [544, 744]. NETBANK [403]. NetBIOS [991, 992, 1078, 2084]. NETBLT [959, 988, 1020]. NETBUGGER3 [162]. NETCRT [205]. NETED [562]. Netiquette [1842]. NeTraMet [2110]. NETRJS [88, 189, 320, 582, 592, 730]. NETRJT [278]. NetWare [2229]. NetWare/IP [2229]. Network [4, 8, 15, 21, 35, 37, 38, 46, 51, 61, 62, 63, 72, 75, 77, 82, 85, 86, 87, 90, 94, 96, 99, 105, 106, 112, 113, 115, 119, 120, 121, 122, 125, 134, 155, 164, 168, 173, 178, 186, 187, 192, 193, 207, 211, 214, 223, 234, 243, 245, 246, 252, 253, 261, 262, 280, 281, 282, 283, 288, 293, 295, 298, 301, 309, 310, 311, 314, 318, 320, 321, 324, 325, 327, 329, 333, 337, 339, 348, 351, 357, 358, 361, 362, 365, 371, 372, 377, 379, 382, 384, 391, 397, 406, 414, 416, 418, 426, 453, 463, 488, 508, 536, 554, 560, 562, 606, 608, 622, 636, 638, 671, 704, 707, 710, 714, 731, 814, 905, 948, 967, 976, 984, 1007, 1049, 1057, 1059, 1088, 1099, 1109, 1147, 1152, 1155, 1200, 1203, 1213, 1218, 1228, 1261, 1267, 1292, 1295, 1314, 1336, 1345, 1351, 1367, 1431, 1432, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1442, 1443, 1453, 1455, 1460, 1463, 1464, 1488, 1500, 1503]. Network [1517, 1555, 1558, 1570, 1582, 1604, 1617, 1620, 1621, 1635, 1698, 1747, 1759, 1811, 1824, 1829, 1848, 1889, 1890, 1891, 1892, 1893, 1894, 1895, 1918, 1931, 1975, 1994, 2008, 2017, 2028, 2058, 2061, 2070, 2166, 2198, 2202, 2203, 2206, 2235, 2249, 2251, 2252, 2254, 2259, 2261, 2262, 2294, 2327, 2378, 2473, 20, 32, 74, 110, 199, 212, 225, 232, 255, 269, 299, 302, 376, 387, 398, 434, 440, 447, 469, 482, 493, 494, 504, 522, 525, 550, 558, 561, 604, 609, 627, 630, 637, 674, 675, 697, 715, 723, 779, 816, 818, 881, 916, 931, 937, 946, 947, 1042, 1056, 1060, 1091, 1137, 1146, 1148, 1163, 1206, 1232, 1249, 96, 101, 542, 1034, 1084, 1118, 1119, 1157]. network-based [697]. Network/440 [187]. Networked [1679]. Networking [316, 1912, 2445, 89, 1067, 1198]. Networks [1093, 1178, 1280, 1281, 1296, 1319, 1380, 1392, 1459, 1576, 1599, 1668, 1781, 1904, 1959, 2009, 2213, 2256, 2340, 2353, 2404, 2451, 2454, 2457, 2479, 2484, 2493, 144, 146, 169, 285, 640, 867, 884, 885, 929, 932, 938, 1020, 1032, 1041, 1078, 1114, 1122, 1142, 1191, 1224, 1247, 1692, 2478]. New-Protocol [659, 669, 691, 692, 693]. News [967, 464]. Next [1444, 1465, 1540, 1716, 1742, 2319, 87]. NFILE [1027]. NFS [1084, 1803, 2211, 2326]. NGG [531]. NGP [396]. NGP-0 [396]. NGWG [468]. NHCs [2499]. NHRP [2319, 2320, 2322, 2323, 2499]. NIC [109, 66, 95, 153, 185, 237, 239, 265, 271, 340, 466, 473, 505, 543, 602, 614, 746, 1251]. NIC/NLS [602]. NICNAME [802, 944]. NICNAME/WHOIS [802, 944]. night [958]. NIMAIL [776]. Nimrod [1743, 1979, 2089, 2090]. NLA [2437]. NLS [482, 602]. NMG [318]. No [25, 265]. NOC [1287]. Node [1583, 2160, 2444, 581]. nodes [19]. noise [66]. Nomenclator [2245, 2246]. Nominating [2014, 2269]. Non [1267, 1332, 1512, 1846, 2034, 2178, 2478, 2075, 2256]. Non-ASCII [1332, 1512, 2034]. non-ATM [2256]. Non-Broadcast [2478]. Non-OSI [1267]. non-synchronizing [2075]. Non-Unicast [2178]. Non-use [1846].

Nonstandard [1045]. Nortel [2327]. Note [61, 70, 523, 573, 610, 661, 621]. notebook [346, 440, 458, 477]. Notes [36, 82, 101, 367, 1140, 1189, 1301, 1359, 34, 46, 194, 322, 689, 790, 795, 797, 888, 889, 900, 989]. Notification [326, 1983, 2116, 213, 256, 338, 2468]. Notifications [1878, 1881, 2285, 266]. NOTIFY [1983]. Novell [1541, 1624, 1352, 2228]. November [417, 594, 659]. Novice [1453]. NRC [929, 935]. NREN [1249]. NSAP [1338, 972, 1227, 1619, 1627, 1696]. NSAPs [1875]. NSFNET [1064, 1082, 1083, 1123, 1212, 1472]. NSSA [1577]. NTP [948, 1155, 1698]. NULL [2397]. Number [170, 179, 1935, 1969, 200, 427, 497, 610, 936]. Numbers [25, 300, 312, 344, 707, 890, 1089, 1789, 1886, 1986, 2086, 2132, 2159, 2186, 2286, 2386, 317, 527, 597, 621, 729, 740, 745, 748, 752, 760, 766, 780, 810, 860, 913, 933, 950, 980, 987, 1000, 1010, 1050, 1052, 1107, 1156, 1209, 2309, 1330, 1690]. NVP [731]. NVT [563]. NWG [38, 39, 44, 47, 57, 73, 84, 95, 100, 108, 116, 131, 135, 140, 212, 350]. NWG/RFC [38, 39, 44, 47, 57, 73, 84, 95, 135, 350]. NWG/RFCs [100]. NYU [544].

O [1675, 1826, 2281]. O/R [1675, 1826, 2281]. OAKLEY [2399]. Object [1836, 2066, 2432, 2494]. OBJECT-IDENTITIES [2494]. Objects [1259, 1274, 1276, 1294, 1306, 1307, 1308, 1343, 1358, 1388, 1396, 1397, 1461, 1462, 1463, 1464, 1483, 1505, 1506, 1515, 1585, 1586, 1594, 1613, 1620, 1633, 1640, 1647, 1648, 1649, 1650, 1655, 1656, 1684, 1685, 1737, 1757, 1993, 2007, 2011, 2038, 2095, 2142, 2219, 2225, 2226, 2253, 2274, 2307, 2312, 2345, 2353, 2404, 2442, 2443, 2444, 2481, 2482, 2483, 2493, 2495, 1222, 1223]. Observations [1734, 1944, 2023, 611, 963]. Obsolete [2049]. Octet [1688]. October [290, 408, 579, 1849, 2428]. ODA [1187, 2148]. ODETTE [2191]. off [1015]. offerings [1072]. Office [908, 927, 1071, 1072, 1215, 1450, 1715, 1926, 1944]. OFFICIAL [1240, 1270, 1770, 1790, 1867, 2187, 2287, 2387, 53, 54, 123, 312, 379, 685, 830, 870, 891, 914, 934, 951, 981, 1001, 1350, 1400, 1490, 1530, 1590, 1600, 1710, 1907, 1987, 117, 143, 151, 165, 284, 714, 1073, 1090, 1120, 1130, 1190]. OLS [217, 519]. On-Line [74, 216, 345, 618, 121, 537, 599, 601, 616]. ONC [1780, 1823]. One [681, 1511, 1750, 1925, 2032, 2276, 2431, 1111]. One-Time [1750, 1925, 2276]. One-Time-Password [2431]. Only [2403, 287]. onto [2076]. Opaque [2357]. Open [1029, 1249, 2472]. OpenPGP [2427]. Operation [1267, 1271, 1518, 1519, 1520, 1535, 1559, 1629, 1693, 2014, 2169, 2269, 230, 376, 517]. Operational [1394, 1639, 1844, 1899, 1997]. Operations [1155, 1266, 1438, 1638, 1892, 1023, 2175]. operators [121]. Opstat [1843]. Optimization [157]. Option [574, 845, 847, 848, 849, 850, 851, 1174, 1383, 1577, 1760, 1772, 1773, 1775, 2053, 2077, 2100, 2204, 2277, 2334, 2335, 2357, 2372, 2471, 2472, 553, 556, 642, 643, 644, 645, 646, 647, 648, 688, 716, 717, 719, 721, 722, 725, 726, 738, 739, 769, 771, 874, 875, 917, 920, 923, 936, 1028, 1031, 1033, 1043, 1063, 1069, 1070, 1081, 1086, 1087, 1100, 1106, 1133, 1362, 1398, 1399, 1406, 1561, 1562]. Options [851, 1098, 1523, 1774, 2005, 2119, 2228, 2336, 2341, 2476, 580, 650, 1053, 1096, 1135, 1136, 1162]. oral [1163]. Order [1683, 970]. Ordering [2456]. Organization [1386, 1630]. Organizations [2015, 379]. Oriented [2493, 76, 765]. OS/2 [1491]. OSI [1060, 1132, 1151, 1155, 1185, 1204, 1213, 1227, 1230, 1267, 1273, 1320, 1367, 1408, 1421, 1474, 1475, 1554, 1564, 1619, 1688, 1771, 1875, 2017]. OSI-DS [1421, 1474, 1475, 1554]. OSINLCP [1367]. OSPF [1121, 1235, 1236, 1237, 1238, 1242, 1243, 1354, 1360, 1393, 1573, 1574, 1576, 1577, 1735, 1755, 1783, 1837, 2141, 2165, 2315, 2316, 2357]. Other [1661, 66, 1091]. OTP [2230]. Out-of-band [711]. Out-of-net [744]. Output [102, 105, 107, 739, 923, 643, 644, 645, 646, 647, 648, 649]. Overflow [1755]. Overview [174, 1299, 1407, 1501, 1623, 1748, 2058, 2085, 2092, 2327].

Packard [1975]. Packet [819, 1346, 1372, 1443, 1542, 1751, 1927, 2460, 625, 960]. Packets [1959, 2006, 2403, 2451, 2454, 2457, 2484, 1041, 1122]. Pad [2027, 1043]. Padding [70]. paged [673]. Pages [1578, 1674, 1785, 2135, 2153, 2205, 2340]. Paging [1558, 1559, 1635, 1693, 1848]. Paper [150, 1540, 1661, 1700]. papers [468, 624]. paradigm [1206]. Parallel [1308, 1650]. Parallel-printer-like [1308, 1650]. Parameter [2171, 2218]. Parameters [2202, 474]. PARRY [433]. Part [972, 979, 1030, 1103, 1104, 1105, 1411, 1412, 1413, 1414, 1511, 1512, 2032, 2033, 2034, 2035, 2036, 2146, 2148]. Partial [1683]. Parties [1343]. Parts [2145]. Party [1437]. Password [962, 1750, 1916, 1925, 2276, 2431]. Past [1290]. PASTE [2417]. patent [1812]. Patents [1975]. Path [1181, 1425, 1781, 1968, 905, 971]. Pathname [608]. paths [1062, 1175]. Payload [1817, 2016, 2019, 2022, 2025, 2177, 2185, 2237, 2330, 2380, 2381, 2382, 2393, 2416, 2418, 2422]. PCA [1862]. PCMAIL [974, 983, 1046]. PDF [2333]. PDP [190]. PDP-10-IMLAC [190]. peg [2309]. peg-dhcp [2309]. Performance [653, 1313, 1800, 2317, 2480, 1118]. personal [904, 914, 974, 983, 1046]. Perspective [861, 1117, 1596]. PF KEY [2354]. PGP [1978, 2002]. Ph [2365]. Phase [1279, 1320, 1366, 1549, 1752]. philosophy [635]. phone [505]. Photuris [1812]. Physical [1445]. PICS [1698]. Pilot [1792]. Pilots [1374, 1607]. PIM [2104, 2324, 2349]. PIM-SM [2104, 2349]. PINT [2445]. Pioneer [2504]. Pip [1611, 1612]. Pipelining [1841, 2184]. pitfall [718]. PKCS [2300, 2301, 2302, 2424]. PL1 [120]. Places [1280, 1392]. Plain [2469]. Plan [1097, 1420, 1517, 2364, 761, 791, 871, 1005]. Planning [1680]. Plans [149]. plasma [593]. plateau [48]. Plus [1921, 1700]. PNG [2070]. poems [1111]. Point [1124, 1161, 1162, 1210, 1321, 1461, 1462, 1463, 1464, 1537, 1538, 1651, 2105]. Point-To-Point [2105, 1124, 1161, 1162, 1210, 1321, 1461, 1462, 1463, 1464, 1537, 1538, 1651]. POISED [1386, 1630]. Policies [1519, 1776, 1995, 2501, 115]. Policy [1092, 1114, 1115, 1164, 1263, 1468, 1469, 1472, 1520, 1862, 2264, 2267, 892, 1082, 1094]. Policy-Based [1472]. POP [2082, 2182, 2371]. POP3 [1724, 1944, 2436]. Port [2204, 1068]. Portable [1363, 2070]. ports [241, 657, 804]. Possible [48, 202, 619]. Post [908, 927, 1071, 1072, 1215, 1450, 1715, 1926, 1944]. Postel [2504]. Posting [1194]. Postmaster [1638]. Postponement [232]. PostScript [2147, 2333]. Pot [2311]. Potential [1281]. Power [1663]. PPP [1162, 1321, 1322, 1323, 1324, 1366, 1367, 1368, 1538, 1539, 1542, 1560, 1588, 1608, 1609, 1628, 1651, 1652, 1653, 1707, 1752, 1753, 1754, 1829, 1864, 1902, 1949, 1950, 1954, 1955, 1956, 1960, 1961, 1962, 1963, 1964, 1965, 1966, 1976, 1977, 1980, 1981, 2010, 2030, 2084, 2112, 2140, 2271, 2277, 2350, 2351, 2406, 2407, 2420, 2459, 2471, 2492]. Practice [2167]. Practices [1566, 1617, 1808]. Pre [784, 2445, 621]. pre-assigned [621]. Pre-emption [784]. Pre-PINT [2445]. Precision [1579]. Predictor [1965]. Prefixes [1904]. preliminary [34]. presence [912]. present [1486]. Presentation [1268, 1796, 2170, 1075]. Pressure [111, 130]. Pretty [2002]. Prevention [2072]. preventive [439, 626, 631]. Primary [1568, 2064]. prime [952]. Primer [1729, 2138]. Principles [1518, 1519, 1520, 1559, 1693, 1945, 2263]. Print [442]. Printable [1830]. Printer [1636, 1749, 1169, 1308, 1650]. Printers [315]. Printing [1476, 1518, 1519]. priority [570]. Privacy [979, 1030, 1103, 1104, 1105, 1345, 1411, 1412, 1413, 1414, 2002]. Private [1587, 1905]. Problem [1100, 1325, 1525, 1695, 499, 500, 564, 637, 696, 747]. Problems [157, 1201, 953, 954]. Procedure [664, 1040, 1047, 1580, 1821, 1858, 2476, 674, 940]. Procedures [1411, 1518, 1559, 1593, 1693, 2001, 2035, 2265, 2405, 2501, 11, 136, 979, 1030, 1103]. Process [343, 423, 1300, 1592, 1630, 2013, 2014, 2015, 2269, 342, 982, 1386]. Processes [356]. Processing [796, 1612, 2196, 2249, 2259, 520, 945]. processors [640]. Production [1793]. Proffered [123, 165, 247, 143, 151]. Proffering [54]. Profile [1877, 1898, 2293, 2408, 2413, 2446]. PROFORMA [1698]. Program [1058, 1218, 1498, 1813, 2065, 2466, 320, 440, 665]. programmer [222]. Programming [340, 698, 1303]. programs [465]. Project [1792, 2245, 524, 1792]. Prompt [1983]. Proposal [50, 51, 86, 98, 125, 440, 514, 569, 1337, 1355, 2468, 299, 445, 483, 1124]. Proposals [1444, 238]. Proposed [43, 67, 72, 91, 158, 184, 307, 328, 335, 344, 355, 363, 389, 461, 518, 608, 640, 714, 796, 876, 919, 924, 967, 1093, 1178, 1353, 1467, 1525, 1967, 2437, 137, 138, 239, 448, 533, 548, 680, 918, 97]. Protected [2465]. Protection [2372, 2466]. PROTOCOL [1240, 1270, 1770, 1790, 1867, 2187, 2287, 2387, 17, 33, 36, 38, 39, 40, 45, 54, 56, 60, 65, 72, 79, 91, 98, 102, 106, 107, 109, 112, 114, 141, 142, 165, 172, 187, 192, 196, 197, 221, 224, 259, 273, 276, 278, 287, 289, 307, 312, 319, 328, 331, 349, 355, 363, 380, 382, 389, 402, 409, 421, 424, 445, 448, 461, 487, 514, 518, 529, 535, 573, 600, 604, 607, 617, 636, 652, 656, 661, 670, 677, 680, 682, 684, 685, 686, 690, 694, 695, 703, 704, 711, 724, 730, 731, 732, 737, 751, 755, 756, 758, 762, 764, 770, 773, 774, 775, 776, 778, 783, 792, 811, 841, 852, 853, 854, 855, 857, 858, 859, 868, 882, 893, 895, 897, 898, 899, 903, 906, 908, 916, 919, 927, 941, 948, 949, 954, 962, 967, 968, 970, 984, 985, 991, 992, 995, 997, 998, 1012, 1020, 1034, 1035, 1040, 1047, 1048, 1049, 1054, 1057, 1065, 1071, 1072, 1085, 1088, 1109, 1124, 1132, 1141, 1147, 1149]. Protocol [1152, 1155, 1161, 1162, 1166, 1184, 1186, 1193, 1210, 1211, 1215, 1218, 1225, 1228, 1235, 1236, 1239, 1255, 1256, 1278, 1283, 1291, 1295, 1297, 1298, 1299, 1302, 1321, 1322, 1329, 1350, 1351, 1366, 1367, 1377, 1400, 1403, 1419, 1424, 1437, 1438, 1450, 1455, 1461, 1462, 1463, 1464, 1469, 1482, 1490, 1521, 1522, 1528, 1530, 1531, 1532, 1538, 1558, 1571, 1582, 1590, 1600, 1628, 1635, 1644, 1646, 1651, 1698, 1700, 1704, 1707, 1710, 1711, 1712, 1715, 1725, 1742, 1752, 1754, 1759, 1764, 1767, 1785, 1803, 1809, 1814, 1821, 1829, 1848, 1864, 1872, 1876, 1892, 1902, 1907, 1908, 1915, 1921, 1940, 1949, 1950, 1954, 1955, 1956, 1961, 1964, 1965, 1973, 1974, 1977, 1980, 1981, 1987, 1999, 2000, 2017, 2020, 2030, 2061, 2068, 2079, 2080, 2081, 2084, 2093, 2094, 2101, 2104, 2105, 2111, 2116, 2118, 2152, 2154, 2158, 2160, 2161, 2162, 2175, 2190, 2191, 2216, 2233, 2244, 2271, 2278, 2284, 2311, 2313, 2314, 2319, 2320, 2321, 2325, 2328, 2349, 2370, 2376, 2377, 2380, 2395, 2406, 2407, 2433, 2434, 2447, 2450]. Protocol [2472, 2487, 19, 46, 48, 53, 76, 88, 117, 199, 205, 523, 533, 544, 546, 548, 640, 654, 674, 715, 807, 816, 904, 955, 959, 988, 994, 1027, 1073, 1090, 1120, 1130, 1169, 1190, 1216, 1217, 1426, 1763, 1340, 1720, 1746, 2047, 93, 97, 137, 139, 145, 158, 171, 260, 290, 323, 388, 455, 489, 552, 659, 669, 691, 692, 693, 718, 743, 749, 750, 754, 767, 771, 781, 782, 816, 817, 844, 877, 878, 880, 892, 894, 928, 953, 956, 976, 1003, 1018, 1034, 1059, 1064, 1084, 1095, 1098, 1118, 1119, 1153, 1154, 1180, 1194, 1231, 1254, 1257, 1258, 1259, 1322, 1339, 1368, 1375, 1387, 1426, 1432, 1433, 1434, 1435, 1436, 1438, 1439, 1440, 1449, 1466, 1477, 1494, 1537, 1542, 1582, 1645, 1647, 1682, 1719, 1753, 1761, 1762, 1788, 1795, 1815, 1864, 1870, 1872, 1885, 1889, 1890, 1891, 1892, 1893, 1894, 1926, 1932, 1944, 1966, 1998, 2049, 2055, 2112, 2173, 2174, 2192, 2195, 2196, 2223, 2231, 2238, 2239, 2240, 2246, 2249, 2251, 2252, 2259, 2261, 2262]. Protocol [2268, 2358, 2359, 2388, 2399, 2439, 2441, 2486]. Protocols [80, 305, 313, 1324, 1342, 1436, 1448, 1462, 1780, 1823, 2326, 2344, 2464, 163, 545, 575, 628, 779, 830, 870, 881, 891, 914, 925, 929, 932, 934, 951, 981, 1001, 1029, 1092, 1213, 1179]. Prototypical [55]. provide [1036, 1265, 1266]. Provider [1510, 2060, 2257, 2417, 1777, 2247]. Provider-Based [2060]. provides [838]. Providing [984, 916]. Proxies [1906]. PSN [969]. PSTN [2290, 2445]. PSTN/Internet [2445]. Public [1160, 2012, 2446, 2501, 867, 1249, 1285]. Public-Key [2012]. Publishing [1794]. Punycode [2500]. purchase [249].

Q [1133]. QoS [2373]. QoS-based [2373]. Quality [1323, 1976, 2199, 2409, 538]. quench [1006]. Queries [2295]. Query [2246]. Questionnaire [106, 112, 180]. questionnaires [453]. Questions [1167, 1196, 1197, 1315, 1568, 1584, 1852, 1928, 17]. Queue [1972, 2296]. Queuing [1036]. quick [490]. QUOTA [2074]. Quote [855]. Quoted [1446].

Quoted-Readable [1446].

r [2255, 1675, 1826, 2281]. RA [2417]. race [161]. Radio [1303, 1559, 1667, 1693, 1973]. RADIUS [2045, 2046, 2125, 2126]. Rand [113, 399, 78, 225, 227]. Rand/UCLA [227]. Rand/UCSB [225]. randomly [738]. randomly-lose [738]. Randomness [1740]. RAP [1466]. Rapicom [759, 788]. RARP [1918]. rates [227]. Ratio [1705]. RATP [906]. RC2 [2255]. RC5 [2027]. RC5-CBC [2027]. RC5-CBC-Pad [2027]. RC5-CTS [2027]. RCTE [556, 708, 709]. RDBMS [1687]. RDP [1141]. Re [39, 17, 544]. Reachability [1263]. Readable [1741, 575, 1446]. Readings [1453]. Ready [635]. Real [502, 1811, 1876, 2313, 1183]. Real-Time [1876, 502, 1811, 1183]. Reallocation [71]. reassembly [805]. Reassignment [1229]. Rec [2416]. Recall [2014, 2269]. Recommendation [1361, 1471, 1742, 794]. Recommendations [118, 1320, 1726, 1740, 1744, 1793, 1866, 2296, 2489, 1042]. Recommended [1164]. Reconfiguration [138, 166, 194, 431, 83]. Reconnection [421, 661]. Reconsidered [175, 167]. Record [2217, 875]. Records [1347, 1535, 1627, 1629, 1696, 1797]. Recovery [133, 1988, 654, 806]. reduce [19]. Redundancy [2325]. Redundant [2185]. Reference [81, 1171, 490, 819, 861, 990]. Referral [1704, 2154]. Referrals [2180, 2208]. Reflection [1953]. Reflections [57]. Regarding [143]. regime [1031]. Registering [2029]. Registration [1390, 1479, 1580, 1943, 2035, 2265, 2289, 2409, 2410, 1917]. Registries [2348]. REGISTRY [2037, 1776]. Related [1414, 1859, 2099, 2374, 869]. Relational [1687]. Relationship [1384, 2018]. Relative [1798]. Relay [1158, 1284, 1305, 1449, 1480, 1576, 1586, 1594, 1960, 2102, 2414, 2143]. Release [354]. Releases [401]. relevant [146, 182]. reliability [522, 629]. Reliable [898, 906, 925, 928, 1141, 1653, 2344, 203, 230, 623, 2308]. Remember [2455, 2503]. Remembrances [1290]. Remote [88, 105, 278, 302, 320, 355, 363, 402, 471, 808, 1040, 1047, 1155, 1261, 1329, 1362, 1443, 1476, 1503, 1518, 1519, 1747, 1821, 1924, 1972, 2008, 2045, 2061, 2093, 2125, 2175, 2283, 231, 359, 582, 1070, 1746, 553, 574, 716]. Removal [405]. Renumbering [1887, 1903, 2058, 2059]. Repair [2341]. Repeater [1358, 1506, 2095, 2253]. Replay [2072]. Replication [1265, 1266]. Reply [1834, 466, 633]. Report [63, 265, 290, 524, 1009, 1099, 1578, 1626, 1679, 1800, 1849, 1879, 2117, 2303, 2316, 77, 78, 113, 138, 152, 277, 929, 935, 1142, 1182, 1814]. Reporting [1879, 218]. reports [182]. Representation [1332, 1475, 1478, 1548, 1768, 1769, 1776, 1822, 1911, 1947, 2240, 2241, 961, 993, 1004]. Representing [1598, 1826, 1827, 2280, 2281]. Request [81, 129, 504, 613, 689, 790, 815, 889, 990, 1002, 1089, 1101, 1189, 1296, 1789, 1886, 1986, 2086, 2186, 2286, 2301, 2386, 466, 1101]. REQuested [2157]. Requests [815, 989]. require [1247]. Requirement [2106]. Requirements [975, 999, 1112, 1113, 1117, 1265, 1448, 1537, 1639, 1657, 1658, 1662, 1666, 1667, 1668, 1669, 1676, 1706, 1727, 1743, 1802, 2089, 2278, 2359, 2367, 910, 1007, 1115, 1116, 1183, 1200, 1287]. Research [34, 126, 191, 1157, 1207, 1606, 1663, 2001, 1005, 1007, 1142, 1200]. ReSerVation [2192, 2195, 2196, 2338]. Reservations [245]. Resilient [2435]. Resolution [816, 893, 1283, 1319, 1725, 2155, 2156, 2263, 2319, 2377, 2470, 915, 1019]. Resolver [1602]. Resource [61, 62, 458, 513, 1336, 1570, 1620, 1627, 1696, 1718, 1726, 1727, 1728, 1798, 1962, 2003, 2043, 2066, 2098, 2155, 2263, 2275, 2379, 2383, 2400, 440, 477, 519, 585, 697, 715, 877, 2192, 2195, 2196]. Resources [1336, 1504, 346, 498, 538]. Responding [456]. Response [29, 73, 125, 130, 131, 135, 156, 350, 486, 561, 596, 607, 675, 2082, 2182, 2297, 2337, 457, 525, 606]. Responses [548, 2230, 112]. Responsibilities [1163]. Restart [1833]. Restricted [515, 1127]. Results [1866, 566]. Resynchronization [461, 483]. RET [68]. Retransmit [1988]. Retrieval [105, 734, 1177, 1679, 1719, 1843, 2332, 452, 1426]. retrofit [428, 441, 470]. Return [1904, 643]. Returning [2021, 2375]. rev [470]. Revelations [550]. Reverse [893, 2331]. Review [1099, 2181]. Reviewed [197]. Revised [633, 642, 725, 1622, 911, 1028]. Revision [273, 706, 1592, 2013, 773, 1340]. revisited [480]. RFC [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83, 84, 85, 86, 87, 88, 89, 90, 91, 92, 93, 94, 95, 96, 97, 98, 99, 100, 101, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119, 120, 121, 122, 123, 124, 125, 126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136, 137, 138, 139, 140, 141, 142, 143, 144, 145, 146, 147, 148, 149, 150, 151, 152, 153, 154, 155, 156, 157, 158, 159, 160, 161, 162, 163, 164, 165, 166, 167]. RFC [168, 169, 170, 171, 172, 173, 174, 175, 176, 177, 178, 179, 180, 181, 182, 183, 184, 185, 186, 187, 188, 189, 190, 191, 192, 193, 194, 195, 196, 197, 198, 199, 200, 201, 202, 203, 204, 205, 206, 207, 208, 209, 210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 224, 225, 226, 227, 228, 229, 230, 231, 232, 233, 234, 235, 236, 237, 238, 239, 240, 241, 242, 243, 244, 245, 246, 247, 248, 249, 250, 251, 252, 253, 254, 255, 256, 257, 258, 259, 260, 261, 262, 263, 264, 265, 266, 267, 268, 269, 270, 271, 272, 273, 274, 275, 276, 277, 278, 279, 280, 281, 282, 283, 284, 285, 286, 287, 288, 289, 290, 291, 292, 293, 294, 295, 296, 297, 298, 299, 300, 301, 302, 303, 304, 305, 306, 307, 308, 309, 310, 311, 312, 313, 314, 315, 316, 317, 318, 319, 320, 321, 322, 323, 324, 325, 326, 327, 328, 329, 330, 331, 332, 333, 334]. RFC [335, 336, 337, 338, 339, 340, 341, 342, 343, 344, 345, 346, 347, 348, 349, 350, 351, 352, 353, 354, 355, 356, 357, 358, 359, 360, 361, 362, 363, 364, 365, 366, 367, 368, 369, 370, 371, 372, 373, 374, 375, 376, 377, 378, 379, 380, 381, 382, 383, 384, 385, 386, 387, 388, 389, 390, 391, 392, 393, 394, 395, 396, 397, 398, 399, 400, 401, 402, 403, 404, 405, 406, 407, 408, 409, 410, 411, 412, 413, 414, 415, 416, 417, 418, 419, 420, 421, 422, 423, 424, 425, 426, 427, 428, 429, 430, 431, 432, 433, 434, 435, 436, 437, 438, 439, 440, 441, 442, 443, 444, 445, 446, 447, 448, 449, 450, 451, 452, 453, 454, 455, 456, 457, 458, 459, 460, 461, 462, 463, 464, 465, 466, 467, 468, 469, 470, 471, 472, 473, 474, 475, 476, 477, 478, 479, 480, 481, 482, 483, 484, 485, 486, 487, 488, 489, 490, 491, 492, 493, 494, 495, 496, 497, 498, 499, 500, 501]. RFC [502, 503, 504, 505, 506, 507, 508, 509, 510, 511, 512, 513, 514, 515, 516, 517, 518, 519, 520, 521, 522, 523, 524, 525, 526, 527, 528, 529, 530, 531, 532, 533, 534, 535, 536, 537, 538, 539, 540, 541, 542, 543, 544, 545, 546, 547, 548, 549, 550, 551, 552, 553, 554, 555, 556, 557, 558, 559, 560, 561, 562, 563, 564, 565, 566, 567, 568, 569, 570, 571, 572, 573, 574, 575, 576, 577, 578, 579, 580, 581, 582, 583, 584, 585, 586, 587, 588, 589, 590, 591, 592, 593, 594, 595, 596, 597, 598, 599, 600, 601, 602, 603, 604, 605, 606, 607, 608, 609, 610, 611, 612, 613, 614, 615, 616, 617, 618, 619, 620, 621, 622, 623, 624, 625, 626, 627, 628, 629, 630, 631, 632, 633, 634, 635, 636, 637, 638, 639, 640, 641, 642, 643, 644, 645, 646, 647, 648, 649, 650, 651, 652, 653, 654, 655, 656, 657, 658, 659, 660, 661, 662, 663, 664, 665, 666, 667, 668]. RFC [669, 670, 671, 672, 673, 674, 675, 676, 677, 678, 679, 680, 681, 682, 683, 684, 685, 686, 687, 688, 689, 690, 691, 692, 693, 694, 695, 696, 697, 698, 699, 700, 701, 702, 703, 704, 705, 706, 707, 708, 709, 710, 711, 712, 713, 714, 715, 716, 717, 718, 719, 720, 721, 722, 723, 724, 725, 726, 727, 728, 729, 730, 731, 732, 733, 734, 735, 736, 737, 738, 739, 740, 741, 742, 743, 744, 745, 746, 747, 748, 749, 750, 751, 752, 753, 754, 755, 756, 757, 758, 759, 760, 761, 762, 763, 764, 765, 766, 767, 768, 769, 770, 771, 772, 773, 774, 775, 776, 777, 778, 779, 780, 781, 782, 783, 784, 785, 786, 787, 788, 789, 790, 791, 792, 793, 794, 795, 796, 797, 798, 799, 800, 801, 802, 803, 804, 805, 806, 807, 808, 809, 810, 811, 812, 813, 814, 815, 816, 817, 818, 819, 820, 821, 822, 823, 824, 825, 826, 827, 828, 829, 830, 831, 832, 833, 834, 835]. RFC [836, 837, 838, 839, 840, 841, 842, 843, 844, 845, 846, 847, 848, 849, 850, 851, 852, 853, 854, 855, 856, 857, 858, 859, 860, 861, 862, 863, 864, 865, 866, 867, 868, 869, 870, 871, 872, 873, 874, 875, 876, 877, 878, 879, 880, 881, 882, 883, 884, 885, 886, 887, 888, 889, 890, 891, 892, 893, 894, 895, 896, 897, 898, 899, 900, 901, 902, 903, 904, 905, 906, 907, 908, 909, 910, 911, 912, 913, 914, 915, 916, 917, 918, 919, 920, 921, 922, 923, 924, 925, 926, 927, 928, 929, 930, 931, 932, 933, 934, 935, 936, 937, 938, 939, 940, 941, 942, 943, 944, 945, 946, 947, 948, 949, 950, 951, 952, 953, 954, 955, 956, 957, 958, 959, 960, 961, 962, 963, 964, 965, 966, 967, 968, 969, 970, 971, 972, 973, 974, 975, 976, 977, 978, 979, 980, 981, 982, 983, 984, 985, 986, 987, 988, 989, 990, 991, 992, 993, 994, 995, 996, 997, 998, 999, 1000, 1001, 1002]. RFC [1003, 1004, 1005, 1006, 1007, 1008, 1009, 1010, 1011, 1012, 1013, 1014, 1015, 1016, 1017, 1018, 1019, 1020, 1021, 1022, 1023, 1024, 1025, 1026, 1027, 1028, 1029, 1030, 1031, 1032, 1033, 1034, 1035, 1036, 1037, 1038, 1039, 1040, 1041, 1042, 1043, 1044, 1045, 1046, 1047, 1048, 1049, 1050, 1051, 1052, 1053, 1054, 1055, 1056, 1057, 1058, 1059, 1060, 1061, 1062, 1063, 1064, 1065, 1066, 1067, 1068, 1069, 1070, 1071, 1072, 1073, 1074, 1075, 1076, 1077, 1078, 1079, 1080, 1081, 1082, 1083, 1084, 1085, 1086, 1087, 1088, 1089, 1090, 1091, 1092, 1093, 1094, 1095, 1096, 1097, 1098, 1099, 1100, 1101, 1102, 1103, 1104, 1105, 1106, 1107, 1108, 1109, 1110, 1111, 1112, 1113, 1114, 1115, 1116, 1117, 1118, 1119, 1120, 1121, 1122, 1123, 1124, 1125, 1126, 1127, 1128, 1129, 1130, 1131, 1132, 1133, 1134, 1135, 1136, 1137, 1138, 1139, 1140, 1141, 1142, 1143, 1144, 1145, 1146, 1147, 1148, 1149, 1150, 1151, 1152, 1153, 1154, 1155, 1156, 1157, 1158, 1159, 1160, 1161, 1162, 1163, 1164, 1165, 1166, 1167, 1168, 1169]. RFC [1170, 1171, 1172, 1173, 1174, 1175, 1176, 1177, 1178, 1179, 1180, 1181, 1182, 1183, 1184, 1185, 1186, 1187, 1188, 1189, 1190, 1191, 1192, 1193, 1194, 1195, 1196, 1197, 1198, 1199, 1200, 1201, 1202, 1203, 1204, 1205, 1206, 1207, 1208, 1209, 1210, 1211, 1212, 1213, 1214, 1215, 1216, 1217, 1218, 1219, 1220, 1221, 1222, 1223, 1224, 1225, 1226, 1227, 1228, 1229, 1230, 1231, 1232, 1233, 1234, 1235, 1236, 1237, 1238, 1239, 1240, 1241, 1242, 1243, 1244, 1245, 1246, 1247, 1248, 1249, 1250, 1251, 1252, 1253, 1254, 1255, 1256, 1257, 1258, 1259, 1260, 1261, 1262, 1263, 1264, 1265, 1266, 1267, 1268, 1269, 1270, 1271, 1272, 1273, 1274, 1275, 1276, 1277, 1278, 1279, 1280, 1281, 1282, 1283, 1284, 1285, 1286, 1287, 1288, 1289, 1290, 1291, 1292, 1293, 1294, 1295, 1296, 1297, 1298, 1299, 1300, 1301, 1302, 1303, 1304, 1305, 1306, 1307, 1308, 1309, 1310, 1311, 1312, 1313, 1314, 1315, 1316, 1317, 1318, 1319, 1320, 1321, 1322, 1323, 1324, 1325, 1326, 1327, 1328, 1329, 1330, 1331, 1332, 1333, 1334, 1335, 1336]. RFC [1337, 1338, 1339, 1340, 1341, 1342, 1343, 1344, 1345, 1346, 1347, 1348, 1349, 1350, 1351, 1352, 1353, 1354, 1355, 1356, 1357, 1358, 1359, 1360, 1361, 1362, 1363, 1364, 1365, 1366, 1367, 1368, 1369, 1370, 1371, 1372, 1373, 1374, 1375, 1376, 1377, 1378, 1379, 1380, 1381, 1382, 1383, 1384, 1385, 1386, 1387, 1388, 1389, 1390, 1391, 1392, 1393, 1394, 1395, 1396, 1397, 1398, 1399, 1400, 1401, 1402, 1403, 1404, 1405, 1406, 1407, 1408, 1409, 1410, 1411, 1412, 1413, 1414, 1415, 1416, 1417, 1418, 1419, 1420, 1421, 1422, 1423, 1424, 1425, 1426, 1427, 1428, 1429, 1430, 1431, 1432, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1441, 1442, 1443, 1444, 1445, 1446, 1447, 1448, 1449, 1450, 1451, 1452, 1453, 1454, 1455, 1456, 1457, 1458, 1459, 1460, 1461, 1462, 1463, 1464, 1465, 1466, 1467, 1468, 1469, 1470, 1471, 1472, 1473, 1474, 1475, 1476, 1477, 1478, 1479, 1480, 1481, 1482, 1483, 1484, 1485, 1486, 1487, 1488, 1489, 1490, 1491, 1492, 1493, 1494, 1495, 1496, 1497, 1498, 1499, 1500, 1501, 1502, 1503]. RFC [1504, 1505, 1506, 1507, 1508, 1509, 1510, 1511, 1512, 1513, 1514, 1515, 1516, 1517, 1518, 1519, 1520, 1521, 1522, 1523, 1524, 1525, 1526, 1527, 1528, 1529, 1530, 1531, 1532, 1533, 1534, 1535, 1536, 1537, 1538, 1539, 1540, 1541, 1542, 1543, 1544, 1545, 1546, 1547, 1548, 1549, 1550, 1551, 1552, 1553, 1554, 1555, 1556, 1557, 1558, 1559, 1560, 1561, 1562, 1563, 1564, 1565, 1566, 1567, 1568, 1569, 1570, 1571, 1572, 1573, 1574, 1575, 1576, 1577, 1578, 1579, 1580, 1581, 1582, 1583, 1584, 1585, 1586, 1587, 1588, 1589, 1590, 1591, 1592, 1593, 1594, 1595, 1596, 1597, 1598, 1599, 1600, 1601, 1602, 1603, 1604, 1605, 1606, 1607, 1608, 1609, 1610, 1611, 1612, 1613, 1614, 1615, 1616, 1617, 1618, 1619, 1620, 1621, 1622, 1623, 1624, 1625, 1626, 1627, 1628, 1629, 1630, 1631, 1632, 1633, 1634, 1635, 1636, 1637, 1638, 1639, 1640, 1641, 1642, 1643, 1644, 1645, 1646, 1647, 1648, 1649, 1650, 1651, 1652, 1653, 1654, 1655, 1656, 1657, 1658, 1659, 1660, 1661, 1662, 1663, 1664, 1665, 1666, 1667, 1668, 1669, 1670]. RFC [1671, 1672, 1673, 1674, 1675, 1676, 1677, 1678, 1679, 1680, 1681, 1682, 1683, 1684, 1685, 1686, 1687, 1688, 1689, 1690, 1691, 1692, 1693, 1694, 1695, 1696, 1697, 1698, 1699, 1700, 1701, 1702, 1703, 1704, 1705, 1706, 1707, 1708, 1709, 1710, 1711, 1712, 1713, 1714, 1715, 1716, 1717, 1718, 1719, 1720, 1721, 1722, 1723, 1724, 1725, 1726, 1727, 1728, 1729, 1730, 1731, 1732, 1733, 1734, 1735, 1736, 1737, 1738, 1739, 1740, 1741, 1742, 1743, 1744, 1745, 1746, 1747, 1748, 1749, 1750, 1751, 1752, 1753, 1754, 1755, 1756, 1757, 1758, 1759, 1760, 1761, 1762, 1763, 1764, 1765, 1766, 1767, 1768, 1769, 1770, 1771, 1772, 1773, 1774, 1775, 1776, 1777, 1778, 1779, 1780, 1781, 1782, 1783, 1784, 1785, 1786, 1787, 1788, 1789, 1790, 1791, 1792, 1793, 1794, 1795, 1796, 1797, 1798, 1799, 1800, 1801, 1802, 1803, 1804, 1805, 1806, 1807, 1808, 1809, 1810, 1811, 1812, 1813, 1814, 1815, 1816, 1817, 1818, 1819, 1820, 1821, 1822, 1823, 1824, 1825, 1826, 1827, 1828, 1829, 1830, 1831, 1832, 1833, 1834, 1835, 1836, 1837]. RFC [1838, 1839, 1840, 1841, 1842, 1843, 1844, 1845, 1846, 1847, 1848, 1849, 1850, 1851, 1852, 1853, 1854, 1855, 1856, 1857, 1858, 1859, 1860, 1861, 1862, 1863, 1864, 1865, 1866, 1867, 1868, 1869, 1870, 1871, 1872, 1873, 1874, 1875, 1876, 1877, 1878, 1879, 1880, 1881, 1882, 1883, 1884, 1885, 1886, 1887, 1888, 1889, 1890, 1891, 1892, 1893, 1894, 1895, 1896, 1897, 1898, 1899, 1900, 1901, 1902, 1903, 1904, 1905, 1906, 1907, 1908, 1909, 1910, 1911, 1912, 1913, 1914, 1915, 1916, 1917, 1918, 1919, 1920, 1921, 1922, 1923, 1924, 1925, 1926, 1927, 1928, 1929, 1930, 1931, 1932, 1933, 1934, 1935, 1936, 1937, 1938, 1939, 1940, 1941, 1942, 1943, 1944, 1945, 1946, 1947, 1948, 1949, 1950, 1951, 1952, 1953, 1954, 1955, 1956, 1957, 1958, 1959, 1960, 1961, 1962, 1963, 1964, 1965, 1966, 1967, 1968, 1969, 1970, 1971, 1972, 1973, 1974, 1975, 1976, 1977, 1978, 1979, 1980, 1981, 1982, 1983, 1984, 1985, 1986, 1987, 1988, 1989, 1990, 1991, 1992, 1993, 1994, 1995, 1996, 1997, 1998, 1999, 2000, 2001, 2002, 2003, 2004]. RFC [2005, 2006, 2007, 2008, 2009, 2010, 2011, 2012, 2013, 2014, 2015, 2016, 2017, 2018, 2019, 2020, 2021, 2022, 2023, 2024, 2025, 2026, 2027, 2028, 2029, 2030, 2031, 2032, 2033, 2034, 2035, 2036, 2037, 2038, 2039, 2040, 2041, 2042, 2043, 2044, 2045, 2046, 2047, 2048, 2049, 2050, 2051, 2052, 2053, 2054, 2055, 2056, 2057, 2058, 2059, 2060, 2061, 2062, 2063, 2064, 2065, 2066, 2067, 2068, 2069, 2070, 2071, 2072, 2073, 2074, 2075, 2076, 2077, 2078, 2079, 2080, 2081, 2082, 2083, 2084, 2085, 2086, 2087, 2088, 2089, 2090, 2091, 2092, 2093, 2094, 2095, 2096, 2097, 2098, 2099, 2100, 2101, 2102, 2103, 2104, 2105, 2106, 2107, 2108, 2109, 2110, 2111, 2112, 2113, 2114, 2115, 2116, 2117, 2118, 2119, 2120, 2121, 2122, 2123, 2124, 2125, 2126, 2127, 2128, 2129, 2130, 2131, 2132, 2133, 2134, 2135, 2136, 2137, 2138, 2139, 2140, 2141, 2142, 2143, 2144, 2145, 2146, 2147, 2148, 2149, 2150, 2151, 2152, 2153, 2154, 2155, 2156, 2157, 2158, 2159, 2160, 2161, 2162, 2163, 2164, 2165, 2166, 2167, 2168, 2169, 2170, 2171]. RFC [2172, 2173, 2174, 2175, 2176, 2177, 2178, 2179, 2180, 2181, 2182, 2183, 2184, 2185, 2186, 2187, 2188, 2189, 2190, 2191, 2192, 2193, 2194, 2195, 2196, 2197, 2198, 2199, 2200, 2201, 2202, 2203, 2204, 2205, 2206, 2207, 2208, 2209, 2210, 2211, 2212, 2213, 2214, 2215, 2216, 2217, 2218, 2219, 2220, 2221, 2222, 2223, 2224, 2225, 2226, 2227, 2228, 2229, 2230, 2231, 2232, 2233, 2234, 2235, 2236, 2237, 2238, 2239, 2240, 2241, 2242, 2243, 2244, 2245, 2246, 2247, 2248, 2249, 2250, 2251, 2252, 2253, 2254, 2255, 2256, 2257, 2258, 2259, 2260, 2261, 2262, 2263, 2264, 2265, 2266, 2267, 2268, 2269, 2270, 2271, 2272, 2273, 2274, 2275, 2276, 2277, 2278, 2279, 2280, 2281, 2282, 2283, 2284, 2285, 2286, 2287, 2288, 2289, 2290, 2291, 2292, 2293, 2294, 2295, 2296, 2297, 2298, 2299, 2300, 2301, 2302, 2303, 2304, 2305, 2306, 2307, 2308, 2309, 2310, 2311, 2312, 2313, 2314, 2315, 2316, 2317, 2318, 2319, 2320, 2321, 2322, 2323, 2324, 2325, 2326, 2327, 2328, 2329, 2330, 2331, 2332, 2333, 2334, 2335, 2336, 2337, 2338]. RFC [2339, 2340, 2341, 2342, 2343, 2344, 2345, 2346, 2347, 2348, 2349, 2350, 2351, 2352, 2353, 2354, 2355, 2356, 2357, 2358, 2359, 2360, 2361, 2362, 2363, 2364, 2365, 2366, 2367, 2368, 2369, 2370, 2371, 2372, 2373, 2374, 2375, 2376, 2377, 2378, 2379, 2380, 2381, 2382, 2383, 2384, 2385, 2386, 2387, 2388, 2389, 2390, 2391, 2392, 2393, 2394, 2395, 2396, 2397, 2398, 2399, 2400, 2401, 2402, 2403, 2404, 2405, 2406, 2407, 2408, 2409, 2410, 2411, 2412, 2413, 2414, 2415, 2416, 2417, 2418, 2419, 2420, 2421, 2422, 2423, 2424, 2425, 2426, 2427, 2428, 2429, 2430, 2431, 2432, 2433, 2434, 2435, 2436, 2437, 2438, 2439, 2440, 2441, 2442, 2443, 2444, 2445, 2446, 2447, 2448, 2449, 2450, 2451, 2452, 2453, 2454, 2455, 2456, 2457, 2458, 2459, 2460, 2461, 2462, 2463, 2464, 2465, 2466, 2467, 2468, 2469, 2470, 2471, 2472, 2473, 2474, 2475, 2476, 2477, 2478, 2479, 2480, 2481, 2482, 2483, 2484, 2485, 2486, 2487, 2488, 2489, 2490, 2491, 2492, 2493, 2494, 2495, 2496, 2497, 2498, 2499, 2500, 2501, 2502]. RFC-822 [1016, 1484, 1485, 2144]. RFC-822/MIME [2144]. RFC1006 [1846]. RFC1327 [1654]. RFCs [151, 1117, 1786, 2106, 2421, 100, 682]. RFNM [68]. rid [64]. RIG [1785]. Rights [1812, 1975, 1285]. Ring [1221, 1459, 1503, 2457]. RIP [1377, 1378, 1379, 1571, 1572, 1711, 1712, 1713, 1714, 2069, 2078, 2079, 2440]. RIP-2 [2069]. ripe [1776, 1171]. ripe-81 [1776]. RIPEMD128 [2273]. RIPEMD160 [2273]. RIPng [2067, 2068]. RIPv1 [1910]. RITA [2308]. RJE [217, 319, 333, 493, 715]. RJE/RJOR [217]. RJOR [217]. RJS [430, 484]. Rlogin [1248, 1272]. road [1249]. Roadmap [1646, 2398]. Roaming [2181, 2464]. Role [753, 1159]. Roles [2129]. Root [1997, 2107]. round [612]. round-trip [612]. Route [1387, 1466, 1850, 1953, 2426, 2498]. Router [1246, 2059, 2085, 2100, 2268, 2325]. Routers [1706, 1802, 1920, 2324]. routes [804]. Routing [985, 1048, 1065, 1123, 1126, 1132, 1254, 1312, 1337, 1370, 1373, 1455, 1468, 1469, 1472, 1494, 1507, 1509, 1510, 1515, 1658, 1691, 1692, 1701, 1739, 1743, 1776, 1777, 1791, 1792, 1807, 1927, 1942, 1979, 1985, 1995, 1996, 2172, 2176, 2188, 2373, 964, 971, 1082, 1083, 1092, 1094, 1115, 1116, 1185, 1212, 1850, 2267]. RPC [1040, 1047, 1780, 1821, 1823]. RPCSEC GSS [2190]. RPSL [2267]. RR [1173, 2039]. RRs [1338]. RS [1307, 1649]. RS-232-like [1307, 1649]. RSA [2300, 2424]. RSVP [2192, 2193, 2194, 2195, 2196, 2197, 2366, 2367, 2369, 2477]. RTP [1876, 1877, 2016, 2019, 2022, 2025, 2177, 2185, 2237, 2330, 2416, 2418, 2422, 2491]. RTSP [2313]. Rule [2437]. Rules [1486, 2196]. Running [1576]. RVSA [2283]. RVSA/1.0 [2283]. RWhois [2154, 1704].

s [2409, 1750, 2298, 2299]. S. [49]. S/KEY [1750]. S/MIME [2298, 2299]. Safe [1632, 2139, 2297]. sampling [444]. SASL [2209, 2232, 2431]. Satellite [341, 2475, 352, 819]. SATNET [821]. Scalable [1936, 2247]. Scaling [1471]. Scenario [622, 385]. Scenarios [254]. Schedule [296, 1357, 223, 234, 428, 439, 441, 470, 586, 626, 631, 678, 871, 880, 887, 911]. Scheduled [401, 434]. Scheduling [615, 2432, 2433]. Schema [1264, 1794, 2205, 2243]. Scheme [1231, 1245, 1942, 2179, 2211, 2371, 436, 443, 922, 926, 930, 994, 1208, 2355, 2384]. School [1568]. Schools [1928]. Science [174]. scientific [1007]. Scoped [2352]. SCSP [2321, 2322, 2430]. SDH [1585, 1609, 2158, 2162]. SDLC [1737]. SDP [2314]. SDTP [1950]. Search [1548, 1947, 2241, 1426]. Searching [1280, 1392]. Second [157, 253, 391, 405, 588, 589, 1099]. Secondary [1568, 2169]. Section [2421]. Secure [1271, 2124, 2474]. Security [1098, 1234, 1342, 1436, 1445, 1447, 1462, 1497, 1498, 1499, 1525, 1626, 1665, 1740, 1814, 1815, 1817, 1835, 1836, 1845, 1897, 2002, 2052, 2065, 2071, 2166, 2183, 2209, 2215, 2251, 2261, 2303, 2310, 2337, 2388, 2393, 2394, 2395, 2398, 2466, 2467, 2488, 1028]. segment [869]. Selection [1636, 2014, 2169, 2269, 2283, 1917]. Selective [2005]. Send [1149, 1302, 1418, 769]. send-location [769]. Sender [1430, 1760]. Sender-Initiated [1430]. Sender-Initiated/Unsolicited [1430]. September [207, 395, 572, 692]. Sequence [1935]. Sequences [1688, 523]. sequencing [619]. Serial [1950, 1969, 2491, 1045, 1134]. Server [106, 109, 112, 413, 526, 801, 943, 1601, 1843, 1850, 1864, 2042, 2136, 2216, 2321, 110, 432, 460, 472, 582, 728, 746, 921, 986]. Server-FTP [526]. server-server [432, 472]. Servers [693, 1779, 1997, 2026, 2169, 659, 669, 691, 692]. Service [90, 138, 166, 194, 231, 278, 416, 431, 471, 734, 785, 984, 1155, 1199, 1263, 1339, 1390, 1415, 1416, 1417, 1445, 1497, 1498, 1499, 1536, 1586, 1594, 1641, 1642, 1643, 1683, 1717, 1793, 1820, 1824, 1833, 1841, 1856, 1857, 1878, 1900, 1930, 1972, 2021, 2045, 2062, 2065, 2113, 2125, 2152, 2178, 2184, 2198, 2199, 2201, 2202, 2203, 2254, 2294, 2322, 2368, 2430, 2466, 2474, 398, 492, 505, 602, 616, 618, 623, 763, 808, 820, 902, 905, 931, 945, 991, 992, 1036, 1068, 1072, 1192, 1825, 768, 1420, 1500, 1552, 2135, 2205]. Service-Level [1263]. Services [703, 712, 1085, 1179, 1260, 1281, 1292, 1298, 1299, 1320, 1414, 1455, 1555, 1623, 1670, 1674, 1779, 1811, 1836, 2129, 2197, 2200, 2201, 2206, 2235, 2369, 2417, 2461, 2462, 2470, 364, 798, 838, 916, 996, 1075, 1097, 1158, 1183, 1230, 1251, 2039, 1320, 2228]. Serving [359]. Session [2314]. Sessions [2372]. Set [289, 1479, 2117, 2306, 247]. Sets [368, 1335, 1492, 1805, 2171, 2218, 2264]. Settings [390]. SGML [1861]. SHA [2391, 1839, 2189]. Shared [242, 1610]. Sharing [61, 62, 2137, 2378, 15, 144, 146, 243, 285, 519, 585, 697, 715]. shifts [1206].

Short [1453, 2175, 842]. Should [1349, 1517]. Shouldn’t [1617]. Shows [2166]. Side [1967, 821]. SIFT [1430]. SIFT/UFT [1430]. Signaling [1745]. Signalling [2318]. Signals [711]. Signature [2372, 637]. Signatures [2141]. Signed [1835]. Simple [122, 404, 778, 811, 903, 1057, 1088, 1147, 1218, 1337, 1351, 1437, 1528, 1558, 1635, 1759, 1848, 1973, 2012, 2017, 2082, 2182, 2214, 2292, 2465, 1018, 1432, 1433, 1434, 1435, 1436, 1438, 1439, 1440, 1582, 1700, 1889, 1890, 1891, 1892, 1893, 1894, 2209, 2246, 2249, 2251, 2252, 2259, 2261, 2262]. Simple-Minded [122, 404]. Simplified [60]. Simulation [1657, 2402, 2477, 2486]. Single [545, 2257, 625]. SIP [1294]. Site [106, 112, 198, 235, 240, 347, 1234, 2183, 156, 182, 379, 465, 662]. Sites [2257, 482, 505]. Six [1695]. Size [1417, 1643, 1774, 1857, 2108, 2336, 2402, 176, 869, 1063]. SKIP [2343]. SLIP [1045]. Slow [1207, 1988]. SM [2104, 2349]. Small [2130]. Smallberg [837]. SMDS [1199, 1684]. SMFS [217, 394, 425]. SMIv2 [1640, 1647, 1648, 1649, 1650, 1655, 1656, 1684, 1685, 1686, 1687, 1733, 1737, 1738, 1739, 1993, 1998, 1999, 2000, 2008, 2011, 2024, 2038, 2095, 2102, 2114, 2115, 2142, 2193, 2200, 2201, 2219, 2220, 2225, 2312, 2226, 2307]. SMTP [866, 1037, 1080, 1415, 1416, 1417, 1641, 1642, 1643, 1820, 1833, 1834, 1841, 1856, 1857, 1878, 1972, 2021, 2184, 2429, 2474, 2489]. SMTP/MIME [1418]. SNA [1528, 1583, 1655, 1656, 1737, 2030]. SNA/IP [1528]. SNACP [2030]. SNMP [1079, 1088, 1147, 1151, 1177, 1205, 1217, 1218, 1260, 1273, 1288, 1293, 1341, 1342, 1343, 1371, 1372, 1408, 1409, 1410, 1451, 2076, 2248, 2249, 2252, 2258, 2259, 2262]. SNMP-based [1293]. SNMP-DPI [1218]. SNMPv1 [2076]. SNMPv2 [1432, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1493, 1896, 1897, 1998, 1999, 2000, 2076]. SNMPv3 [2250, 2251, 2260, 2261]. Snoop [1751]. SNQP [2246]. SNTP [1351, 1759, 2017]. SOBs [1428]. Society [1780, 2121, 2122]. Socket [167, 175, 344, 427, 497, 2120, 129, 147, 317, 464, 610]. Sockets [204, 393, 2279]. SOCKS [1915, 1916, 1948]. Software [1, 377, 401, 406, 416, 522, 1525, 1795, 2, 9, 11, 434, 520]. Solicitation [1540, 1669, 1903]. Solution [161, 1325, 2089, 2090, 747]. solutions [499]. Some [17, 34, 89, 94, 115, 117, 192, 250, 264, 382, 566, 585, 651, 953, 954, 1008, 1617, 1944, 2195, 2385]. Something [1006]. Sometimes [1482]. SONET [1585, 1595, 1609, 2158, 2162]. SONET/SDH [1585, 1609, 2158, 2162]. Sonnet [1595]. Source [1515, 1739, 1927, 2254, 1006, 1006, 2044]. sources [819]. Space [1325, 1356, 1357, 1456, 1734, 2023]. Spam [2489]. Sparse [2104, 2324, 2349]. Spec [67]. special [888]. Specific [396, 972, 1812, 1975]. Specification [638, 656, 665, 710, 771, 796, 831, 844, 1239, 1287, 1295, 1353, 1405, 1446, 1469, 1634, 1782, 1803, 1809, 1821, 1870, 1927, 1937, 1938, 1974, 2041, 2042, 2070, 2080, 2104, 2109, 2111, 2116, 2168, 2175, 2190, 2192, 2198, 2199, 2203, 2267, 2284, 2298, 2349, 2370, 2432, 2447, 2460, 166, 540, 555, 754, 755, 800, 873, 882, 894, 895, 897, 928, 942, 953, 954, 969, 972, 1025, 1034, 1035, 1040, 1047, 1049, 1084, 1109, 1121, 1211, 1939, 1940, 2450]. Specifications [74, 105, 217, 509, 731, 845, 2221, 2424, 8, 122, 189, 489, 507, 584, 735, 992, 2425]. Specifying [1331, 1511, 2039]. Speed [2491, 1069, 1134, 1142, 1175]. Speedup [587]. SPF [1064]. SPKM [2012]. Spoofing [2254]. SPX [1402]. SQuID [1006, 1008]. SRI [152, 153, 537, 614]. SRI-ARC [537]. SRV [2039]. ST [1180]. ST-II [1180]. ST2 [1809, 1933, 2370]. Stability [2501]. Stac [1961]. Staff [384, 418]. Stage [2137]. Standard [86, 125, 229, 236, 344, 608, 638, 668, 723, 812, 840, 867, 884, 885, 940, 953, 954, 967, 1026, 1032, 1041, 1078, 1122, 1139, 1178, 1446, 1467, 1478, 1537, 1768, 1785, 1822, 2475, 237, 247, 268, 545, 714, 750, 751, 876, 918, 924, 930, 966, 972, 991, 992, 993, 1004, 1093, 1229, 796, 1975, 1431, 1442, 1895]. Standardization [226, 233, 1254, 2316]. Standardizing [554]. Standards [28, 1300, 1350, 1386, 1400, 1490, 1590, 1592, 1600, 1630, 1710, 1786, 2013, 2015, 2026, 2347, 2425, 231, 961, 1042, 1073, 1090, 1120, 1130, 1160, 1190, 1240, 1270, 1770, 1790, 1867, 2187, 2287, 2387, 1530, 1907, 1987]. Standby [2268]. Standing [1407, 1748]. Start [1988, 958, 1165]. start-up [958]. Starting [1972]. Starts [2403]. State [2096, 679]. Stateless [1958, 2449]. Statement [602, 1360, 1507, 1712, 1910, 1971, 1992, 2068, 2195, 2320, 935, 1029]. Statements [1428, 1434, 1862, 1891]. Static [1455]. Station [1739, 414]. Statistics [395, 986, 1394, 1843, 1844, 373, 383, 386, 408, 417, 437, 449, 476, 491, 503, 516, 532, 549, 559, 572, 579, 594, 605, 611]. Status [138, 156, 240, 255, 262, 282, 288, 310, 314, 321, 325, 327, 337, 339, 354, 357, 361, 365, 371, 621, 849, 1164, 1457, 1679, 1878, 1880, 1881, 1910, 78, 152, 153, 218, 235, 252, 261, 283, 293, 301, 348, 362, 409, 461, 483, 590, 596, 642]. STD [1301]. steering [1142]. stockings [595]. Storage [2493, 960]. Store [1454, 939]. Storing [558]. Strategic [1420]. Strategy [803, 1328, 1471, 1509, 352, 763]. Stream [86, 125, 967, 1180, 1306, 1648, 1809]. Stream-Based [967]. Streaming [2313, 2341]. Streams [2019, 2177]. String [1454, 1475, 1478, 1548, 1768, 1769, 1947, 2240, 2241, 1268]. Structure [116, 1055, 1145, 1325, 1432, 1581, 1889, 129, 972]. Structured [757]. Structuring [1607]. STUB [878]. Studies [2402]. Study [96, 1263]. Style [154]. SU-DSL [630]. Sub [2289, 2409, 2410]. Sub-type [2289, 2409, 2410]. Subdomain [1518, 1519, 1520, 1559, 1693]. Subject [222]. subliminal [1087]. subliminal-message [1087]. Submission [2463, 536]. Subnet [1787, 1847, 1865, 1866, 619, 926, 1017, 1209]. subnets [907, 912]. Subnetting [940, 930]. Subnetwork [922, 1060]. Subnetworks [1924]. subprograms [119, 120]. Subsystem [334, 15]. Subtrees [1827, 2280]. such [2097]. Suggested [276, 323, 747, 1526, 1914]. Suggestion [1365]. Suggestions [199, 839, 19]. Suite [1339, 2486]. Summary [798, 837, 1089, 1189, 1289, 1389, 1489, 1589, 1689, 1789, 1886, 1986, 2086, 2186, 2243, 2286, 2386, 2501, 463, 689, 790, 889, 929, 989, 1163, 1182]. Sun [2343, 1780, 2016]. SUPDUP [724, 726, 736, 737, 739]. SUPDUP-Output [739]. Supernetting [1328]. supplement [997]. Supplemental [2359]. supplies [249]. Support [1267, 1472, 1571, 1572, 1670, 1688, 1745, 1783, 1784, 1792, 1933, 1989, 1992, 1993, 2009, 2078, 2089, 2090, 2247, 2318, 863, 982, 1113, 1791, 1828, 1873, 2151]. Supporting [1296]. Suppress [848]. Surrogate [484]. Survey [524, 741, 834, 866, 961, 1244, 1481, 454, 558, 659, 669, 691, 692, 693, 777, 832, 833, 835, 836, 517]. survey/tutorial [777]. surveys [837]. swap [19]. Switch [390, 1424, 1785, 1974, 2160, 2161, 2284]. Switch-Switch [2161]. Switch-to-Switch [1424, 1785]. Switched [1296, 1297, 1924]. switches [960]. Switching [328, 1424, 1785, 2011, 2092, 2093, 2101, 2272, 2327]. Symposium [157]. synch [523]. Synchronization [58, 2321, 947, 1119]. synchronizing [946, 2075]. Syntax [710, 1620, 2049, 2128, 2221, 2239, 2301, 2302, 2356, 2383, 608, 638]. Syntaxes [1478, 1768]. System [74, 122, 213, 216, 222, 266, 281, 294, 326, 338, 345, 354, 369, 404, 411, 563, 698, 702, 985, 1003, 1011, 1084, 1152, 1228, 1287, 1454, 1516, 1581, 1687, 1750, 1814, 1863, 1879, 1880, 1917, 1925, 1952, 2052, 2123, 2124, 2130, 2274, 2276, 180, 190, 291, 299, 447, 469, 585, 799, 820, 887, 911, 963, 964, 974, 983, 1019, 1046, 1116, 1118, 1188, 1034, 2155]. System-Level [2274]. Systems [62, 157, 831, 1029, 1213, 1603, 494, 520, 747, 796, 2092].

T [1634, 2416, 2423]. T.4 [794]. T/TCP [1634]. T3 [1296]. tab [645, 648]. Table [756, 764, 1177, 1344, 1455, 1847, 1865, 2083, 613, 742, 800, 839, 942]. Tables [1654, 1827, 1929, 2280, 208]. tablet [199]. tabstops [644, 647]. TACACS [917, 1482]. Tactical [1667]. Tag [2092, 2289, 2293]. Tagging [2469]. Tags [1756]. Talk [1303]. talks [822, 823, 824, 825, 826, 827, 828, 829, 832, 833, 834, 835, 836]. Tao [1381, 1529, 1708]. Task [1254, 1381, 1428, 1529, 1708, 1975, 1007]. TCP [763, 791, 803, 822, 823, 824, 825, 826, 827, 828, 829, 832, 833, 835, 836, 838, 862, 869, 886, 952, 973, 991, 992, 996, 1015, 1055, 1056, 1062, 1068, 1075, 1076, 1085, 1096, 1100, 1134, 1135, 1136, 1137, 1145, 1146, 1148, 1170, 1175, 1185, 1203, 1253, 1263, 1313, 1327, 1337, 1369, 1460, 1603, 1634, 1683, 1729, 1781, 1782, 1846, 1988, 2005, 2113, 2127, 2134, 2138, 2372, 2385, 2401, 2402, 2403, 2475]. TCP-4 [952]. TCP-on-a-LAN [862]. TCP/IP [1055, 1056, 1075, 1085, 1134, 1137, 1145, 1146, 1148, 1170, 1185, 1203, 1263, 1460, 1729, 2138]. TCP/IP-based [1055, 1056, 1145, 1146, 1148, 1203]. TCP/IPX [1782]. TCP/UDP [991, 992]. TCPMUX [1068]. Tech [2504]. Technical [520, 1281, 1299, 1518, 1559, 1693, 1716, 1743]. Technique [2435]. Techniques [1214]. Technology [1501, 1869, 1971, 819]. TELENET [446]. Telephone [1779]. Teletype [41]. Television [1676]. Telex [1384]. Telnet [97, 313, 1133, 2053, 137, 139, 158, 206, 215, 216, 323, 334, 335, 388, 429, 455, 460, 489, 507, 552, 553, 555, 556, 574, 580, 586, 588, 589, 642, 643, 644, 645, 646, 647, 648, 649, 650, 659, 669, 678, 688, 691, 692, 693, 716, 717, 718, 719, 721, 722, 725, 726, 738, 739, 754, 769, 808, 844, 845, 846, 847, 848, 849, 850, 851, 874, 875, 917, 920, 923, 936, 1031, 1033, 1043, 1063, 1069, 1070, 1081, 1086, 1087, 1106, 1174, 1195, 1362, 1398, 1399, 1401, 1402, 1406, 1561, 1562, 2204]. Template [2203]. Tenex [334, 404, 410, 539, 564, 679, 629, 673, 708]. Tentative [445, 678]. term [1611]. Terminal [215, 372, 772, 110, 593, 874, 920, 936, 1069, 1081, 721, 722, 1033]. terminal-type [1081]. terminals [230]. Terminating [1963]. Terminology [2272, 2419, 1232]. Terms [1171, 1198]. TESS [1814]. Test [2189, 2273, 1249]. test-bed [1249]. Testing [1884, 2385, 2458, 1020]. Text [13, 1332, 1512, 2034, 2311, 2469, 546, 620, 723, 812, 984, 1513, 1553, 1883, 2305]. text/css [2305]. text/enriched [1513, 1553, 1883]. text/graphics [546]. Texts [1546]. Textual [1433, 1890, 2452, 2480, 2494]. TFTP [1340, 773, 896, 1772, 1773, 1774, 1775, 2077, 2334, 2335, 2336]. Thanksgiving [411]. thar [1280, 1392]. their [183, 2356]. them [1280, 1392]. There [1280, 1392]. Things [1290]. Thinwire [904]. Third [56, 66, 88]. Thoughts [57, 533, 712, 1157, 1319, 94, 250, 585, 588, 589]. Three [376, 2034, 2403]. throughout [1391]. Throughput [625]. Ticket [1287]. Ticketing [2338]. Tier [1325]. tiff [2289, 2289, 2293]. Time [28, 142, 728, 858, 948, 1049, 1109, 1118, 1155, 1295, 1351, 1698, 1750, 1759, 1876, 1925, 2017, 2276, 2313, 2431, 15, 502, 615, 675, 1119, 1183, 1811, 1327]. Time-out [142]. TIME-WAIT [1327]. Timekeeping [1579]. Timeline [2222]. Timeout [1774, 2336, 444]. times [612]. timestamp [771]. timetable [4]. Timing [850]. TIP [230, 292, 347, 360, 381, 428, 439, 441, 464, 470, 615, 626, 629, 631, 678]. TIP/Tenex [629]. TIPS [278, 390]. TIPUG [451]. TLA [2437]. TLS [2233, 2474]. TMux [1682]. TN3270 [1566, 1636, 1637, 2342]. TNLS [490]. TNVIP [1908]. To-dos [2433]. Today [1303, 2088]. together [957]. Token [1220, 1221, 1459, 1503, 2457]. Token-Ring [1459]. tolerant [982, 1019]. Toll [2409]. too [834]. Tool [1460, 1137]. Tools [1137, 1460, 1564, 1570, 1679, 1703, 1729, 2138, 2385]. Top [1913, 973, 996, 1075, 1230, 2113]. Topic [87]. topics [869]. TOPS20 [776, 774, 775]. Toshiba [2085, 2116]. TP [1465]. TP/IX [1465]. TP0 [1076]. TPC.INT [1518, 1519, 1520, 1559, 1693]. Traceroute [1383]. Tracing [2028]. Track [2026, 2425]. Trade [2166]. tradition [1163]. Traffic [373, 386, 395, 408, 417, 437, 449, 476, 491, 503, 516, 532, 549, 559, 572, 579, 594, 605, 2050, 2051, 2110, 2338, 2417, 1191, 1224]. Trailer [883]. Training [1994]. Transaction [289, 928, 1035, 2071, 945, 2358, 2359]. Transactions [1369, 1634]. transatlantic [1200]. transcoding [793]. Transfer [114, 133, 141, 171, 172, 259, 260, 264, 276, 289, 304, 305, 322, 349, 380, 409, 424, 448, 535, 600, 607, 617, 755, 762, 770, 774, 775, 776, 778, 811, 903, 906, 949, 967, 1058, 1430, 1774, 1932, 1973, 1982, 2020, 2055, 2191, 2336, 2376, 163, 227, 250, 413, 480, 481, 495, 499, 566, 959, 988]. transfers [653]. Transform [1819, 1838]. Transformation [1632, 2139, 2502, 2031, 2266]. Transition [1251, 1390, 1418, 1661, 1920, 2172, 2323, 761, 763, 791, 1021]. translating [1187]. Translation [1595, 2378]. Translator [1621]. Transmission [23, 553, 665, 670, 703, 716, 751, 783, 846, 954, 967, 1178, 1199, 1321, 1380, 1653, 1820, 1941, 1959, 1999, 2006, 2435, 2451, 2454, 2457, 2484, 502, 757, 777, 816, 867, 884, 885, 938, 1032, 1041, 1045, 1078, 1093, 1122, 1124, 1139, 1161, 574, 2439]. Transmitting [1191, 387]. Transparent [1906, 2282, 1017]. Transponders [1718]. Transport [882, 895, 932, 997, 998, 1291, 1439, 1443, 1528, 1846, 1893, 1950, 2113, 2344, 2356, 2433, 929, 945, 973, 991, 992, 996, 1230, 1682]. Transport-Independent [2433]. Transport/Network [1443]. traps [1205, 2443]. Traversal [2343]. Treasure [1280, 1392]. Tree [1826, 2281]. Trees [2188, 2176]. Trends [2504]. Tribute [2503, 2428]. Triggered [2078, 2079]. trip [612]. Triple [1838, 2407]. Triple-DES [2407]. Trivial [1973, 2156]. Trouble [1287]. Troubleshooting [2308]. Truths [1912]. try [681]. TSO [372, 374, 385]. TSS [413]. TSS/360 [413]. TUBA [1337, 1516, 1551]. TUBA/CLNP [1516]. Tunnel [2094]. Tunneling [1224, 1840, 2331, 2460]. Tutorial [1496, 777, 1170]. Twas [958]. Twelve [1912]. Two [19, 389, 499, 604, 938, 1325, 1512, 1848, 2033, 2328, 525]. Two-Tier [1325]. Two-Way [1848]. Type [289, 704, 1294, 1339, 1397, 1445, 1580, 1585, 1731, 1860, 1879, 2004, 2064, 2066, 2305, 2312, 2412, 2429, 2481, 2483, 874, 920, 1036, 1039, 1081, 1222, 1223, 1513, 1553, 1859, 1882, 1883, 2099, 2207, 2289, 2374, 2409, 2410]. type-of-service [1036]. Types [42, 1274, 1388, 1396, 1427, 1613, 1633, 1640, 1861, 1914, 2029, 2033, 2345, 2363, 2482, 2487, 1091, 1486]. Typographical [124, 132].

U [2306]. U.S. [1098, 1801, 1806, 2133]. UCL [799]. UCLA [49, 174, 227, 384, 418, 484, 2428]. UCLA-CCN [484]. UCSB [74, 78, 105, 113, 122, 216, 225, 345, 404, 430, 431, 471, 479, 519, 538, 563, 567]. UCSB-OLS [519]. UCSB/Rand [78]. UCSD [526]. UCSD-CC [526]. UDP [991, 992, 1230, 1337, 1781, 2134, 2491]. UFT [1430]. UIDPLUS [2346]. Ukrainian [2306]. Un-muddling [495]. UNARP [1855]. Unavailable [411]. UNI [2009, 2318, 2353, 2370, 2404]. Unicast [1874, 2060, 2178, 2361]. Unicode [1631, 1632, 2031, 2139, 2469, 2500, 2501, 2502]. Unified [445, 656, 1312, 1658]. Uniform [1727, 1728, 1798, 2043, 2098, 2155, 2275, 2379, 2003, 2066, 2263, 2383]. Unifying [1620]. UNINETT [1862]. Unique [1429, 1804, 939]. unique-named [939]. Uniqueness [1429]. Unit [2466]. Units [2226, 1505]. Universal [742, 1620]. UNIX [901, 671]. Unknown [300]. Unsolicited [1430]. Unused [1904]. upcoming [602]. Update [425, 592, 1494, 1614, 2124, 1003, 2123]. Update-Based [1494]. Updated [52]. Updates [2123, 613]. updating [1131]. Upload [1854]. Upper [1688]. Upper-Layer [1688]. UPS [1618]. URAs [2003]. Urbana [101, 108]. URI [2383, 2470]. URIs [2066]. URL [1728, 1946, 2004, 2109, 2179, 2211, 2242, 2355, 2371, 2384]. URLs [2356]. URN [2128, 2156, 2470]. Usage [1596, 2214, 212, 231, 269, 412, 623]. Usage-Limiting [2214]. Usages [1481]. USCB [306]. Use [150, 289, 329, 411, 473, 1060, 1185, 1336, 1492, 1551, 1625, 1812, 1828, 2132, 2151, 2197, 2206, 2356, 2390, 2391, 2397, 2487, 74, 204, 515, 976, 985, 1059, 1205, 1228, 1391, 1616, 1791, 1846, 2023, 2106, 2243]. used [1620]. USENET [840, 1026]. User [76, 91, 106, 112, 206, 219, 345, 407, 445, 656, 758, 808, 1167, 1184, 1186, 1278, 1315, 1333, 1382, 1474, 1491, 1514, 1568, 1584, 1677, 1736, 1771, 1810, 1832, 1897, 1970, 2000, 2045, 2125, 2243, 2472, 2488, 110, 231, 456, 582, 604, 614, 809, 917, 1196, 1197, 1200, 1285, 2251, 2261, 2441]. User-based [1897, 2251, 2261]. User-Level [656]. User-User [91]. user/network [604]. User/Server [106, 112]. Username [1916]. Username/Password [1916]. Users [360, 416, 577, 223, 359, 381, 578, 856]. users-2 [381]. Using [302, 372, 374, 571, 1017, 1187, 1298, 1299, 1383, 1454, 1474, 1631, 1654, 1719, 1771, 1799, 2102, 2150, 2234, 2256, 2257, 2275, 2292, 2294, 2307, 2322, 2339, 2381, 2382, 2430, 2480, 110, 254, 541, 622, 896, 1265, 1266, 1640, 1647, 1648, 1649, 1650, 1655, 1656, 1684, 1685, 1686, 1687, 1733, 1737, 1738, 1739, 1818, 1839, 1973, 1993, 1998, 1999, 2000, 2008, 2011, 2024, 2038, 2095, 2114, 2115, 2142, 2155, 2156, 2193, 2200, 2201, 2219, 2220, 2225, 2226, 2312, 2324, 2378, 2475]. USM [2251, 2261]. UTF [1632, 2031, 2139, 2240, 2266, 2502]. UTF-18 [2502]. UTF-7 [1632, 2139]. UTF-8 [2031, 2240, 2266]. UTF-9 [2502]. Utilities [2138]. utility [746]. UUCP [966].

V.4 [2326]. V1.5 [2154]. v12 [1474]. v2 [1421]. v2.0 [2153]. V2ToV1 [2076]. v3 [1554, 2238, 2239, 2240, 1455]. v5 [1475, 1500, 1916]. VACM [2252, 2262]. Value [2171, 2218]. Values [2375]. Variable [1847, 1865, 1962, 1014]. Variance [1858, 1902]. Variant [2283]. Various [1352, 1541, 1624, 1995]. vCard [2413]. Vector [1065]. VEMMI [2109]. Vendor [1385, 1487, 1523, 2119, 2140, 863, 1038, 1074]. VENUS [2178]. Versatile [1035]. Version [221, 509, 664, 927, 996, 1047, 1054, 1071, 1072, 1141, 1166, 1180, 1193, 1211, 1215, 1230, 1231, 1237, 1238, 1242, 1243, 1259, 1295, 1377, 1378, 1379, 1387, 1444, 1450, 1469, 1558, 1573, 1582, 1596, 1635, 1647, 1685, 1698, 1711, 1712, 1713, 1714, 1715, 1744, 1751, 1785, 1802, 1803, 1809, 1821, 1823, 1837, 1848, 1870, 1871, 1872, 1885, 1889, 1890, 1891, 1892, 1893, 1894, 1895, 1915, 1926, 1927, 1940, 1941, 1948, 1957, 1974, 2008, 2010, 2017, 2065, 2070, 2111, 2132, 2158, 2159, 2163, 2165, 2175, 2192, 2195, 2196, 2223, 2233, 2244, 2284, 2298, 2299, 2300, 2301, 2302, 2315, 2354, 2358, 2360, 2370, 2406, 2416, 2424, 2439, 2440, 2441, 2447, 2448, 2450, 2452, 2453, 2459, 695, 1003, 1049, 1109, 1431, 1432, 1433, 1434, 1435, 1436, 1437, 1438, 1439, 1440, 1442, 1873, 1937, 1938, 1939, 1968, 2160, 2161, 2173, 2174, 2176, 2251, 2261, 2408, 1720, 1746, 2047, 1401, 1809, 1951]. Versioning [2278]. Versus [59, 1906]. vertical [647, 648]. Very [258, 540, 584, 2178, 1142]. very-high-speed [1142]. VFIP [968]. via [372, 452, 640, 1614, 2372]. Viability [1659]. victims [957]. Video [1443, 1877, 2016, 2019, 2022, 2025, 2177, 2237, 2418, 2422, 2435]. VIEW [1597, 1664, 1677, 219, 231, 237, 2252, 2262]. View-based [2252, 2262]. Viewpoint [1676]. Views [146]. Vines [1753]. violation [412]. VIQR [1446]. Virtual [372, 772, 814, 1695, 2325, 2327]. Visas [2504]. VISCII [1446]. Vision [1717]. VMTP [1035]. VNS [2327]. Voice [731, 968, 1898, 2408, 2409, 2410]. VPIM [2410]. Vulnerabilities [779].

WAIS [1615]. WAIT [1327]. WAN [1352, 1541, 1543, 1624]. want [2058]. Washington [1200]. WAVE [2348]. Way [1848]. Ways [1736]. Weather [251, 415]. Web [1620, 2026, 2071, 2278]. WEBDAV [2497]. WebNFS [2041, 2042]. Week [411]. weight [2111]. Well [317, 676]. were [595]. where [1165]. which [192, 2254]. White [1540, 1578, 1661, 1674, 1700, 2135, 2205]. Who [822, 823, 824, 825, 826, 827, 828, 829, 832, 833, 834, 835, 836, 838, 1241, 1326]. Whois [1704, 1824, 2154, 802, 944, 1825, 1900, 1901]. Wide [2026, 2278, 1620]. Widely [1525]. Will [2504, 411]. Window [803, 1003, 2401, 2402, 1063, 1096, 1100, 1188]. Wishlist [1287]. Within [1455, 937, 1320, 1990, 1991, 2023, 2076, 2390, 2391]. Word [2171, 2218]. words [2106]. Work [1887, 508]. Working [85, 87, 101, 150, 164, 207, 234, 309, 311, 391, 577, 1386, 1593, 1630, 1669, 1744, 2405, 2428, 578]. Workshop [290, 304, 315, 465, 1009, 1142, 1626, 1849, 2153, 2303, 2340, 222, 322, 498, 2117]. World [1620, 2026, 2278]. World-Wide [1620]. would [2058]. WRITE [1746]. Writers [2347]. Writing [1675, 2421]. Wrong [1280, 1392]. WWW [1620]. WWWSeeker [2496].

X [1003, 1086, 1188]. X.121 [1226]. X.25 [864, 1076, 1080, 1346, 1371, 1372, 1451, 1588, 1603]. X.400 [2143, 977, 1016, 1128, 1138, 1317, 1318, 1320, 1395, 1455, 1484, 1485, 1486, 1492, 1496, 1605, 1606, 1638, 1639, 1675, 1792, 1828, 2143, 2144, 2145, 2147, 2149]. X.400/84 [1486]. X.400/88 [1486]. X.500 [1420, 1552, 1264, 1265, 1266, 1269, 1282, 1298, 1299, 1320, 1477, 1478, 1481, 1557, 1598, 1599, 1607, 1622, 1674, 1788, 1791, 1792, 1793, 1794, 1826, 1827, 1828, 1930, 2066, 2103, 2151, 2234, 2243, 2280, 2281, 2107]. X.500/LDAP [2151]. X.509 [2446]. X3 [1043]. XDR [1004, 1780, 1822]. XML [2363]. XNS [1754]. XNSCP [1754]. XOT [1603]. XRCP [733]. XRSQ [733]. XRSQ/XRCP [733]. XSEN [727].

Z39.50 [1615, 1719, 2043]. Z39.50-1988 [1615]. Zero [13]. ZLIB [1937]. Zone [1982, 1983].

References

RFC0001

[1] Steve Crocker. RFC 1: Host software, April 7, 1969. URL ftp:

      //ftp.internic.net/rfc/rfc1.txt,   ftp://ftp.math.utah.edu/
      pub/rfc/rfc1.txt. Status: UNKNOWN.

RFC0002

[2] B. Duvall. RFC 2: Host software, April 9, 1969. URL ftp:

//ftp.internic.net/rfc/rfc2.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2.txt. Status: UNKNOWN. Not online.

RFC0003

[3] S. D. Crocker. RFC 3: Documentation conventions, April 9, 1969. URL

ftp://ftp.internic.net/rfc/rfc10.txt,   ftp://ftp.internic.
net/rfc/rfc3.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc10.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc3.txt. Obsoleted by RFC0010

[10]. Status: UNKNOWN.

RFC0004

[4] E. B. Shapiro. RFC 4: Network timetable, March 24, 1969. URL

ftp://ftp.internic.net/rfc/rfc4.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc4.txt. Status: UNKNOWN. Not online.

RFC0005

[5] J. Rulifson. RFC 5: Decode encode language (DEL), June 2,

1969. URL ftp://ftp.internic.net/rfc/rfc5.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc5.txt. Status: UNKNOWN.

RFC0006

[6] S. D. Crocker. RFC 6: Conversation with Bob Kahn, April 10,

1969. URL ftp://ftp.internic.net/rfc/rfc6.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc6.txt. Status: UNKNOWN.

RFC0007

[7] G. Deloche. RFC 7: Host-IMP interface, May 1, 1969. URL ftp:

//ftp.internic.net/rfc/rfc7.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc7.txt. Status: UNKNOWN. Not online.

RFC0008

[8] G. Deloche. RFC 8: Functional specifications for the ARPA network,

May 5, 1969. URL ftp://ftp.internic.net/rfc/rfc8.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc8.txt. Status: UNKNOWN. Not

online.

RFC0009

[9] G. Deloche. RFC 9: Host software, May 1, 1969. URL ftp:

//ftp.internic.net/rfc/rfc9.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc9.txt. Status: UNKNOWN. Not online.

RFC0010

[10] S. D. Crocker. RFC 10: Documentation conventions, July 29,

1969.  URL ftp://ftp.internic.net/rfc/rfc10.txt,   ftp://
ftp.internic.net/rfc/rfc16.txt, ftp://ftp.internic.net/rfc/
rfc24.txt,   ftp://ftp.internic.net/rfc/rfc27.txt, ftp://ftp.
internic.net/rfc/rfc3.txt,   ftp://ftp.internic.net/rfc/rfc30.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc10.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc16.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc24.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc27.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc3.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc30.txt. Obsoleted by RFC0016 [16].

Obsoletes RFC0003 [3]. Updated by RFC0024, RFC0027, RFC0030 [24, 27, 30]. Status: UNKNOWN.

RFC0011

[11] G. Deloche. RFC 11: Implementation of the host-host software procedures in GORDO, August 1, 1969. URL ftp://ftp.internic.net/rfc/

rfc11.txt,   ftp://ftp.internic.net/rfc/rfc33.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc11.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc33.txt. Obsoleted by RFC0033 [33]. Status: UN-KNOWN. Not online.

RFC0012

[12] M. Wingfield. RFC 12: IMP-host interface flow diagrams, August 26,

1969. URL ftp://ftp.internic.net/rfc/rfc12.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc12.txt. Status: UNKNOWN. Not online.

RFC0013

[13] V. G. Cerf. RFC 13: Zero text length EOF message, August 20, 1969.

URL ftp://ftp.internic.net/rfc/rfc13.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc13.txt. Status: UNKNOWN.

RFC0014

[14] ???? RFC 14: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc14.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc14.txt.

RFC0015

[15] C. S. Carr. RFC 15: Network subsystem for time sharing hosts, Septem-

ber 25, 1969. URL ftp://ftp.internic.net/rfc/rfc15.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc15.txt. Status: UNKNOWN. Not

online.

RFC0016

[16] S. D. Crocker. RFC 16: M.I.T., September 27, 1969. URL ftp:

//ftp.internic.net/rfc/rfc10.txt,   ftp://ftp.internic.net/
rfc/rfc16.txt, ftp://ftp.internic.net/rfc/rfc24.txt,   ftp:
//ftp.internic.net/rfc/rfc27.txt, ftp://ftp.internic.net/rfc/
rfc30.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc10.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc16.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc24.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc27.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc30.txt. Obsoleted by

RFC0024 [24]. Obsoletes RFC0010 [10]. Updated by RFC0024, RFC0027, RFC0030 [24, 27, 30]. Status: UNKNOWN.

RFC0017

[17] J. E. Kreznar. RFC 17: Some questions re: Host-IMP protocol, August

27, 1969. URL ftp://ftp.internic.net/rfc/rfc17.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc17.txt. Status: UNKNOWN. Not

online.

RFC0018

[18] Vint Cerf. RFC 18: IMP-IMP and HOST-HOST control links, 19xx.

URL ftp://ftp.internic.net/rfc/rfc18.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc18.txt. Status: UNKNOWN.

RFC0019

[19] J. E. Kreznar. RFC 19: Two protocol suggestions to reduce congestion at swap bound nodes, October 7, 1969. URL ftp://ftp.internic.

net/rfc/rfc19.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc19.

txt. Status: UNKNOWN.

RFC0020

[20] V. G. Cerf. RFC 20: ASCII format for network interchange, October 16,

1969. URL ftp://ftp.internic.net/rfc/rfc20.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc20.txt. Status: UNKNOWN. Not online.

RFC0021

[21] V. G. Cerf. RFC 21: Network meeting, October 17, 1969. URL ftp:/

/ftp.internic.net/rfc/rfc21.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc21.txt. Status: UNKNOWN.

RFC0022

[22] V. G. Cerf. RFC 22: Host-host control message formats, October 17,

1969. URL ftp://ftp.internic.net/rfc/rfc22.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc22.txt. Status: UNKNOWN. Not online.

RFC0023

[23] G. Gregg. RFC 23: Transmission of multiple control messages, October

16, 1969. URL ftp://ftp.internic.net/rfc/rfc23.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc23.txt. Status: UNKNOWN.

RFC0024

[24] S. D. Crocker. RFC 24: Documentation conventions, November

21, 1969. URL ftp://ftp.internic.net/rfc/rfc10.txt,   ftp://
ftp.internic.net/rfc/rfc16.txt, ftp://ftp.internic.net/rfc/
rfc24.txt,   ftp://ftp.internic.net/rfc/rfc27.txt, ftp://ftp.
internic.net/rfc/rfc30.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc10.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc16.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc24.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc27.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc30.

txt. Obsoletes RFC0016 [16]. Updates RFC0010, RFC0016 [10, 16]. Updated by RFC0027, RFC0030 [27, 30]. Status: UNKNOWN.

RFC0025

[25] S. D. Crocker. RFC 25: No high link numbers, October 30, 1969.

URL ftp://ftp.internic.net/rfc/rfc25.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc25.txt. Status: UNKNOWN.

RFC0026

[26] ???? RFC 26: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc26.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc26.txt.

RFC0027

[27] S. D. Crocker. RFC 27: Documentation conventions, December

9, 1969. URL ftp://ftp.internic.net/rfc/rfc10.txt,   ftp://
ftp.internic.net/rfc/rfc16.txt, ftp://ftp.internic.net/rfc/
rfc24.txt,   ftp://ftp.internic.net/rfc/rfc27.txt, ftp://ftp.
internic.net/rfc/rfc30.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc10.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc16.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc24.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc27.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc30.

txt. Updates RFC0010, RFC0016, RFC0024 [10, 16, 24]. Updated by RFC0030 [30]. Status: UNKNOWN.

RFC0028

[28] W. K. English. RFC 28: Time standards, January 13, 1970.

URL ftp://ftp.internic.net/rfc/rfc28.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc28.txt. Status: UNKNOWN.

RFC0029

[29] R. E. Kahn. RFC 29: Response to RFC 28, January 19, 1970. URL ftp:

//ftp.internic.net/rfc/rfc28.txt,   ftp://ftp.internic.net/
rfc/rfc29.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc28.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc29.txt. See also RFC0028

[28]. Status: UNKNOWN.

RFC0030

[30] S. D. Crocker. RFC 30: Documentation conventions, February

4, 1970. URL ftp://ftp.internic.net/rfc/rfc10.txt,   ftp://
ftp.internic.net/rfc/rfc16.txt, ftp://ftp.internic.net/rfc/
rfc24.txt,   ftp://ftp.internic.net/rfc/rfc27.txt, ftp://ftp.
internic.net/rfc/rfc30.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc10.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc16.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc24.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc27.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc30.

txt. Updates RFC0010, RFC0016, RFC0024, RFC0027 [10, 16, 24, 27]. Status: UNKNOWN.

RFC0031

[31] D. Bobrow and W. R. Sutherland. RFC 31: Binary message forms in computer, February 1, 1968. URL ftp://ftp.internic.net/rfc/

rfc31.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc31.txt. Sta-

tus: UNKNOWN.

RFC0032

[32] D. Vedder. RFC 32: Connecting M.I.T. computers to the ARPA computer-to-computer communication network, January 31, 1969.

URL ftp://ftp.internic.net/rfc/rfc32.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc32.txt. Status: UNKNOWN. Not online.

RFC0033

[33] S. D. Crocker. RFC 33: New host-host protocol, February 12,

1970.  URL ftp://ftp.internic.net/rfc/rfc11.txt,   ftp://
ftp.internic.net/rfc/rfc33.txt, ftp://ftp.internic.net/rfc/
rfc36.txt,   ftp://ftp.internic.net/rfc/rfc47.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc11.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc33.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc36.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc47.txt. Obsoletes

RFC0011 [11]. Updated by RFC0036, RFC0047 [36, 47]. Status: UN-KNOWN. Not online.

RFC0034

[34] W. K. English. RFC 34: Some brief preliminary notes on the augmentation research center clock, February 26, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc34.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc34.txt. Status: UNKNOWN.

RFC0035

[35] S. D. Crocker. RFC 35: Network meeting, March 3, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc35.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc35.txt. Status: INFORMATIONAL.

RFC0036

[36] S. D. Crocker. RFC 36: Protocol notes, March 16, 1970. URL ftp:

//ftp.internic.net/rfc/rfc33.txt,   ftp://ftp.internic.net/
rfc/rfc36.txt, ftp://ftp.internic.net/rfc/rfc39.txt,   ftp:
//ftp.internic.net/rfc/rfc44.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc33.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc36.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc39.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc44.txt. Updates RFC0033 [33]. Updated

by RFC0039, RFC0044 [39, 44]. Status: UNKNOWN. Not online.

RFC0037

[37] S. D. Crocker. RFC 37: Network meeting epilogue, etc, March 20, 1970.

URL ftp://ftp.internic.net/rfc/rfc37.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc37.txt. Status: UNKNOWN.

RFC0038

[38] S. M. Wolfe. RFC 38: Comments on network protocol from NWG/

RFC #36, March 20, 1970. URL ftp://ftp.internic.net/rfc/rfc38.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc38.txt. Status: UN-

KNOWN.

RFC0039

[39] E. Harslem and J. F. Heafner. RFC 39: Comments on protocol re:

NWG/RFC #36, March 25, 1970. URL ftp://ftp.internic.net/rfc/
rfc36.txt,   ftp://ftp.internic.net/rfc/rfc39.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc36.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc39.txt. Updates RFC0036 [36]. Status: UNKNOWN. Not online.

RFC0040

[40] E. Harslem and J. F. Heafner. RFC 40: More comments on the forthcoming protocol, March 27, 1970. URL ftp://ftp.internic.net/rfc/

rfc40.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc40.txt. Sta-

tus: UNKNOWN.

RFC0041

[41] J. T. Melvin. RFC 41: IMP-IMP teletype communication, March 30,

1970. URL ftp://ftp.internic.net/rfc/rfc41.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc41.txt. Status: UNKNOWN.

RFC0042

[42] E. Ancona. RFC 42: Message data types, March 31, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc42.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc42.txt. Status: UNKNOWN. Not online.

RFC0043

[43] A. G. Nemeth. RFC 43: Proposed meeting, April 8, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc122.txt,   ftp://ftp.internic.net/
rfc/rfc43.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc122.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc43.txt. Updates RFC0122

[122]. Status: UNKNOWN.

RFC0044

[44] A. Shoshani, R. Long, and A. Landsberg. RFC 44: Comments on NWG/

RFC 33 and 36, April 10, 1970. URL ftp://ftp.internic.net/rfc/
rfc36.txt,   ftp://ftp.internic.net/rfc/rfc44.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc36.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc44.txt. Updates RFC0036 [36]. Status: UNKNOWN. Not online.

RFC0045

[45] J. Postel and S. D. Crocker. RFC 45: New protocol is coming, April 14,

1970. URL ftp://ftp.internic.net/rfc/rfc45.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc45.txt. Status: UNKNOWN.

RFC0046

[46] E. W. Meyer, Jr. RFC 46: ARPA network protocol notes, April 17,

1970. URL ftp://ftp.internic.net/rfc/rfc46.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc46.txt. Status: UNKNOWN. Not online.

RFC0047

[47] W. R. Crowther. RFC 47: BBN’s comments on NWG/RFC #33, April

20, 1970. URL ftp://ftp.internic.net/rfc/rfc33.txt,   ftp:/
/ftp.internic.net/rfc/rfc47.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc33.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc47.

txt. Updates RFC0033 [33]. Status: UNKNOWN.

RFC0048

[48] J. Postel and S. D. Crocker. RFC 48: Possible protocol plateau, April 21,

1970. URL ftp://ftp.internic.net/rfc/rfc48.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc48.txt. Status: UNKNOWN. Not online.

RFC0049

[49] E. W. Meyer, Jr. RFC 49: Conversations with S. Crocker (UCLA), April

23, 1970. URL ftp://ftp.internic.net/rfc/rfc49.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc49.txt. Status: UNKNOWN.

RFC0050

[50] E. Harslem and J. Haverty. RFC 50: Comments on the Meyer proposal,

April 30, 1970. URL ftp://ftp.internic.net/rfc/rfc50.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc50.txt. Status: UNKNOWN.

RFC0051

[51] M. Elie. RFC 51: Proposal for a network interchange language, May 4,

1970. URL ftp://ftp.internic.net/rfc/rfc51.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc51.txt. Status: UNKNOWN. Not online.

RFC0052

[52] J. Postel and S. D. Crocker. RFC 52: Updated distribution list, July

1, 1970. URL ftp://ftp.internic.net/rfc/rfc52.txt,   ftp:/
/ftp.internic.net/rfc/rfc69.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc52.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc69.

txt. Updated by RFC0069 [69]. Status: UNKNOWN.

RFC0053

[53] S. D. Crocker. RFC 53: Official protocol mechanism, June 9, 1970.

URL ftp://ftp.internic.net/rfc/rfc53.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc53.txt. Status: UNKNOWN.

RFC0054

[54] S. D. Crocker, J. Postel, J. Newkirk, and M. Kraley. RFC 54: Official protocol proffering, June 18, 1970. URL ftp://ftp.internic.net/rfc/

rfc54.txt,   ftp://ftp.internic.net/rfc/rfc57.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc54.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc57.txt. Updated by RFC0057 [57]. Status: UN-KNOWN.

RFC0055

[55] J. Newkirk, M. Kraley, J. Postel, and S. D. Crocker. RFC 55: Prototypical implementation of the NCP, June 19, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc55.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc55.txt. Status: UNKNOWN. Not online.

RFC0056

[56] E. Belove, D. Black, R. Flegal, and L. G. Farquar. RFC 56: Third level protocol: Logger protocol, June 1, 1970. URL ftp://ftp.internic.

net/rfc/rfc56.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc56.

txt. Status: UNKNOWN.

RFC0057

[57] M. Kraley and J. Newkirk. RFC 57: Thoughts and reflections on

NWG/RFC 54, June 19, 1970. URL ftp://ftp.internic.net/rfc/
rfc54.txt,   ftp://ftp.internic.net/rfc/rfc57.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc54.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc57.txt. Updates RFC0054 [54]. Status: UNKNOWN.

RFC0058

[58] T. P. Skinner. RFC 58: Logical message synchronization, June 26, 1970.

URL ftp://ftp.internic.net/rfc/rfc58.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc58.txt. Status: UNKNOWN.

RFC0059

[59] E. W. Meyer, Jr. RFC 59: Flow control — fixed versus demand alloca-

tion, June 27, 1970. URL ftp://ftp.internic.net/rfc/rfc59.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc59.txt. Status: UNKNOWN.

RFC0060

[60] R. B. Kalin. RFC 60: Simplified NCP protocol, July 15, 1970.

URL ftp://ftp.internic.net/rfc/rfc60.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc60.txt. Status: UNKNOWN.

RFC0061

[61] D. C. Walden. RFC 61: Note on interprocess communication in a resource sharing computer network, July 17, 1970. URL

ftp://ftp.internic.net/rfc/rfc61.txt,   ftp://ftp.internic.
net/rfc/rfc62.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc61.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc62.txt. Obsoleted by

RFC0062 [62]. Status: UNKNOWN. Not online.

RFC0062

[62] D. C. Walden. RFC 62: Systems for interprocess communication in a resource sharing computer network, August 3, 1970. URL ftp:

//ftp.internic.net/rfc/rfc61.txt,   ftp://ftp.internic.net/

rfc/rfc62.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc61.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc62.txt. Obsoletes RFC0061

[61]. Status: UNKNOWN.

RFC0063

[63] V. G. Cerf. RFC 63: Belated network meeting report, July 31, 1970.

URL ftp://ftp.internic.net/rfc/rfc63.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc63.txt. Status: UNKNOWN.

RFC0064

[64] M. Elie. RFC 64: Getting rid of marking, July 1, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc64.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc64.txt. Status: UNKNOWN. Not online.

RFC0065

[65] D. C. Walden. RFC 65: Comments on host/host protocol document #1,

August 29, 1970. URL ftp://ftp.internic.net/rfc/rfc65.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc65.txt. Status: UNKNOWN.

Not online.

RFC0066

[66] S. D. Crocker. RFC 66: NIC — third level ideas and other noise, August

26, 1970. URL ftp://ftp.internic.net/rfc/rfc123.txt,   ftp://
ftp.internic.net/rfc/rfc66.txt, ftp://ftp.internic.net/rfc/
rfc80.txt,   ftp://ftp.internic.net/rfc/rfc93.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc123.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc66.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc80.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc93.txt. Obsoleted by

RFC0123 [123]. Updated by RFC0080, RFC0093 [80, 93]. Status: UN-KNOWN. Not online.

RFC0067

[67] W. R. Crowther. RFC 67: Proposed change to host/IMP spec to eliminate marking, January 1, 1970. URL ftp://ftp.internic.net/rfc/

rfc67.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc67.txt. Sta-

tus: UNKNOWN.

RFC0068

[68] M. Elie. RFC 68: Comments on memory allocation control commands: CEASE, ALL, GVB, RET, and RFNM, August 31, 1970.

URL ftp://ftp.internic.net/rfc/rfc68.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc68.txt. Status: UNKNOWN.

RFC0069

[69] A. K. Bhushan. RFC 69: Distribution list change for MIT, September

22, 1970. URL ftp://ftp.internic.net/rfc/rfc52.txt,   ftp:/
/ftp.internic.net/rfc/rfc69.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc52.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc69.

txt. Updates RFC0052 [52]. Status: UNKNOWN.

RFC0070

[70] S. D. Crocker. RFC 70: Note on padding, October 15, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc228.txt,   ftp://ftp.internic.net/
rfc/rfc70.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc228.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc70.txt. Updated by RFC0228

[228]. Status: UNKNOWN.

RFC0071

[71] T. Schipper. RFC 71: Reallocation in case of input error, September 25,

1970. URL ftp://ftp.internic.net/rfc/rfc71.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc71.txt. Status: UNKNOWN.

RFC0072

[72] R. D. Bressler. RFC 72: Proposed moratorium on changes to network protocol, September 28, 1970. URL ftp://ftp.internic.net/rfc/

rfc72.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc72.txt. Sta-

tus: UNKNOWN.

RFC0073

[73] S. D. Crocker. RFC 73: Response to NWG/RFC 67, September 25,

1970. URL ftp://ftp.internic.net/rfc/rfc73.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc73.txt. Status: UNKNOWN.

RFC0074

[74] J. E. White. RFC 74: Specifications for network use of the UCSB online system, October 16, 1970. URL ftp://ftp.internic.net/rfc/

rfc217.txt,   ftp://ftp.internic.net/rfc/rfc225.txt, ftp://
ftp.internic.net/rfc/rfc74.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc217.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc225.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc74.txt. Updated by RFC0217,

RFC0225 [217, 225]. Status: UNKNOWN. Not online.

RFC0075

[75] S. D. Crocker. RFC 75: Network meeting, October 14, 1970.

URL ftp://ftp.internic.net/rfc/rfc75.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc75.txt. Status: UNKNOWN.

RFC0076

[76] J. Bouknight, J. Madden, and G. R. Grossman. RFC 76: Connection by name: User oriented protocol, October 28, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc76.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc76.txt. Status: UNKNOWN. Not online.

RFC0077

[77] J. Postel. RFC 77: Network meeting report, November 20, 1970.

URL ftp://ftp.internic.net/rfc/rfc77.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc77.txt. Status: UNKNOWN. Not online.

RFC0078

[78] E. Harslem, J. F. Heafner, and J. E. White. RFC 78: NCP status report:

UCSB/Rand, October 1, 1970. URL ftp://ftp.internic.net/rfc/
rfc78.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc78.txt. Sta-

tus: UNKNOWN.

RFC0079

[79] E. W. Meyer, Jr. RFC 79: Logger protocol error, November 16, 1970.

URL ftp://ftp.internic.net/rfc/rfc79.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc79.txt. Status: UNKNOWN. Not online.

RFC0080

[80] E. Harslem and J. F. Heafner. RFC 80: Protocols and data formats, De-

cember 1, 1970. URL ftp://ftp.internic.net/rfc/rfc123.txt,
ftp://ftp.internic.net/rfc/rfc66.txt, ftp://ftp.internic.net/
rfc/rfc80.txt,   ftp://ftp.internic.net/rfc/rfc93.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc123.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc66.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc80.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc93.txt. Obsoleted by

RFC0123 [123]. Updates RFC0066 [66]. Updated by RFC0093 [93]. Status: UNKNOWN. Not online.

RFC0081

[81] J. Bouknight. RFC 81: Request for reference information, December 3,

1970. URL ftp://ftp.internic.net/rfc/rfc81.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc81.txt. Status: UNKNOWN.

RFC0082

[82] E. W. Meyer, Jr. RFC 82: Network meeting notes, December 9, 1970.

URL ftp://ftp.internic.net/rfc/rfc82.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc82.txt. Status: UNKNOWN. Not online.

RFC0083

[83] R. H. Anderson, E. Harslem, and J. F. Heafner. RFC 83: Languagemachine for data reconfiguration, December 18, 1970. URL ftp:/

/ftp.internic.net/rfc/rfc83.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc83.txt. Status: UNKNOWN. Not online.

RFC0084

[84] J. B. North. RFC 84: List of NWG/RFC’s 1–80, December 23, 1970.

URL ftp://ftp.internic.net/rfc/rfc84.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc84.txt. Status: UNKNOWN. Not online.

RFC0085

[85] S. D. Crocker. RFC 85: Network working group meeting, December 28,

1970. URL ftp://ftp.internic.net/rfc/rfc85.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc85.txt. Status: UNKNOWN.

RFC0086

[86] S. D. Crocker. RFC 86: Proposal for a network standard format for a data stream to control graphics display, January 5,

1971.  URL ftp://ftp.internic.net/rfc/rfc125.txt,   ftp:/
/ftp.internic.net/rfc/rfc86.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc125.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc86.

txt. Updated by RFC0125 [125]. Status: UNKNOWN.

RFC0087

[87] A. Vezza. RFC 87: Topic for discussion at the next network working group meeting, January 12, 1971. URL ftp://ftp.internic.net/rfc/

rfc87.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc87.txt. Sta-

tus: UNKNOWN.

RFC0088

[88] R. T. Braden and S. M. Wolfe. RFC 88: NETRJS: A third level protocol for remote job entry, January 13, 1971. URL ftp://ftp.internic.

net/rfc/rfc189.txt,   ftp://ftp.internic.net/rfc/rfc88.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc189.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc88.txt. Obsoleted by RFC0189 [189]. Status: UNKNOWN. Not online.

RFC0089

[89] R. M. Metcalfe. RFC 89: Some historic moments in networking, January

19, 1971. URL ftp://ftp.internic.net/rfc/rfc89.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc89.txt. Status: UNKNOWN. Not

online.

RFC0090

[90] R. T. Braden. RFC 90: CCN as a network service center, January 15,

1971. URL ftp://ftp.internic.net/rfc/rfc90.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc90.txt. Status: UNKNOWN.

RFC0091

[91] G. H. Mealy. RFC 91: Proposed user-user protocol, December 27, 1970.

URL ftp://ftp.internic.net/rfc/rfc91.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc91.txt. Status: UNKNOWN.

RFC0092

[92] ???? RFC 92: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc92.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc92.txt.

RFC0093

[93] A. M. McKenzie. RFC 93: Initial Connection Protocol, January

27, 1971. URL ftp://ftp.internic.net/rfc/rfc66.txt,   ftp://
ftp.internic.net/rfc/rfc80.txt, ftp://ftp.internic.net/rfc/
rfc93.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc66.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc80.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc93.txt. Updates RFC0066, RFC0080 [66, 80]. Status: UNKNOWN.

RFC0094

[94] E. Harslem and J. F. Heafner. RFC 94: Some thoughts on network graph-

ics, February 3, 1971. URL ftp://ftp.internic.net/rfc/rfc94.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc94.txt. Status: UN-

KNOWN. Not online.

RFC0095

[95] S. D. Crocker. RFC 95: Distribution of NWG/RFC’s through

the NIC, February 4, 1971. URL ftp://ftp.internic.net/rfc/
rfc155.txt,   ftp://ftp.internic.net/rfc/rfc95.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc155.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc95.txt. Obsoleted by RFC0155 [155]. Status: UN-KNOWN.

RFC0096

[96] R. W. Watson. RFC 96: Interactive network experiment to study modes of access to the Network Information Center, February 12, 1971.

URL ftp://ftp.internic.net/rfc/rfc96.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc96.txt. Status: INFORMATIONAL.

RFC0097

[97] J. T. Melvin and R. W. Watson. RFC 97: First cut at a Proposed Telnet Protocol, February 15, 1971. URL ftp://ftp.internic.net/rfc/

rfc97.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc97.txt. Sta-

tus: UNKNOWN. Not online.

RFC0098

[98] E. W. Meyer, Jr. and T. P. Skinner. RFC 98: Logger protocol proposal, February 11, 1971. URL ftp://ftp.internic.net/rfc/

rfc123.txt,   ftp://ftp.internic.net/rfc/rfc98.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc123.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc98.txt. Updated by RFC0123 [123]. Status: UN-KNOWN.

RFC0099

[99] P. M. Karp. RFC 99: Network meeting, February 22, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc116.txt,   ftp://ftp.internic.net/
rfc/rfc99.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc116.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc99.txt. Updated by RFC0116

[116]. Status: UNKNOWN.

RFC0100

[100] P. M. Karp. RFC 100: Categorization and guide to NWG/RFCs,

February 26, 1971. URL ftp://ftp.internic.net/rfc/rfc100.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc100.txt.  Status: UN-

KNOWN. Not online.

RFC0101

[101] R. W. Watson. RFC 101: Notes on the Network Working Group meeting, Urbana, Illinois, February 17, 1971, February 23, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc101.txt,   ftp://ftp.internic.net/
rfc/rfc108.txt, ftp://ftp.internic.net/rfc/rfc123.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc101.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc108.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc123.

txt. Updated by RFC0108, RFC0123 [108, 123]. Status: UNKNOWN. Not online.

RFC0102

[102] S. D. Crocker. RFC 102: Output of the host-host protocol glitch cleaning committee, February 22, 1971. URL ftp://ftp.internic.net/rfc/

rfc102.txt,   ftp://ftp.internic.net/rfc/rfc107.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc102.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc107.txt. Updated by RFC0107 [107]. Status: UN-KNOWN. Not online.

RFC0103

[103] R. B. Kalin. RFC 103: Implementation of interrupt keys, February

24, 1971. URL ftp://ftp.internic.net/rfc/rfc103.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc103.txt. Status: UNKNOWN.

RFC0104

[104] J. B. Postel and S. D. Crocker. RFC 104: Link 191, February

25, 1971. URL ftp://ftp.internic.net/rfc/rfc104.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc104.txt. Status: UNKNOWN.

RFC0105

[105] J. E. White. RFC 105: Network specifications for remote job entry and remote job output retrieval at UCSB, March 22, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc105.txt,   ftp://ftp.internic.net/
rfc/rfc217.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc105.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc217.txt. Updated by RFC0217

[217]. Status: UNKNOWN.

RFC0106

[106] T. C. O’Sullivan. RFC 106: User/server site protocol network host ques-

tionnaire, March 3, 1971. URL ftp://ftp.internic.net/rfc/rfc106.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc106.txt.  Status:

UNKNOWN.

RFC0107

[107] R. D. Bressler, S. D. Crocker, W. R. Crowther, G. R. Grossman, R. S. Tomlinson, and J. E. White. RFC 107: Output of the Host-Host Protocol Glitch Cleaning Committee, March 23, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc102.txt,   ftp://ftp.internic.net/
rfc/rfc107.txt, ftp://ftp.internic.net/rfc/rfc111.txt,   ftp:
//ftp.internic.net/rfc/rfc124.txt, ftp://ftp.internic.net/rfc/
rfc132.txt,   ftp://ftp.internic.net/rfc/rfc154.txt, ftp://
ftp.internic.net/rfc/rfc179.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc102.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc107.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc111.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc124.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc132.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc154.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc179.txt. Updates

RFC0102 [102]. Updated by RFC0111, RFC0124, RFC0132, RFC0154, RFC0179 [111, 124, 132, 154, 179]. Status: UNKNOWN.

RFC0108

[108] R. W. Watson. RFC 108: Attendance list at the Urbana NWG meeting, February 17–19, 1971, March 25, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc101.txt,   ftp://ftp.internic.net/
rfc/rfc108.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc101.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc108.txt. Updates RFC0101

[101]. Status: UNKNOWN. Not online.

RFC0109

[109] J. M. Winett. RFC 109: Level III server protocol for the Lincoln Laboratory NIC 360/67 host, March 24, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc109.txt,   ftp://ftp.internic.net/
rfc/rfc393.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc109.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc393.txt. See also RFC0393

[388]. Status: UNKNOWN. Not online.

RFC0110

[110] J. M. Winett. RFC 110: Conventions for using an IBM 2741 terminal as a user console for access to network server hosts, March

25, 1971. URL ftp://ftp.internic.net/rfc/rfc110.txt,   ftp://
ftp.internic.net/rfc/rfc135.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc110.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc135.

txt. Updated by RFC0135 [135]. Status: UNKNOWN. Not online.

RFC0111

[111] S. D. Crocker. RFC 111: Pressure from the chairman, March 31,

1971. URL ftp://ftp.internic.net/rfc/rfc107.txt,   ftp://
ftp.internic.net/rfc/rfc111.txt, ftp://ftp.internic.net/rfc/
rfc130.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc107.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc111.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc130.txt. Updates RFC0107 [107]. Updated by RFC0130 [130]. Status: UNKNOWN.

RFC0112

[112] T. C. O’Sullivan. RFC 112: User/server site protocol: Network host questionnaire responses, April 1, 1971. URL ftp://ftp.

internic.net/rfc/rfc112.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc112.txt. Status: UNKNOWN. Not online.

RFC0113

[113] E. Harslem, J. F. Heafner, and J. E. White. RFC 113: Network activity report: UCSB rand, April 5, 1971. URL ftp://ftp.internic.net/rfc/

rfc113.txt,   ftp://ftp.internic.net/rfc/rfc227.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc113.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc227.txt. Updated by RFC0227 [227]. Status: UN-KNOWN.

RFC0114

[114] A. K. Bhushan. RFC 114: File transfer protocol, April 10,

1971. URL ftp://ftp.internic.net/rfc/rfc114.txt,   ftp://
ftp.internic.net/rfc/rfc141.txt, ftp://ftp.internic.net/rfc/
rfc171.txt,   ftp://ftp.internic.net/rfc/rfc172.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc114.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc141.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc171.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc172.txt. Updated by

RFC0141, RFC0172, RFC0171 [141, 172, 171]. Status: UNKNOWN. Not online.

RFC0115

[115] R. W. Watson and J. B. North. RFC 115: Some network information center policies on handling documents, April 16, 1971. URL ftp://

ftp.internic.net/rfc/rfc115.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc115.txt. Status: UNKNOWN. Not online.

RFC0116

[116] S. D. Crocker. RFC 116: Structure of the may NWG meeting,

April 12, 1971. URL ftp://ftp.internic.net/rfc/rfc116.txt,
 ftp://ftp.internic.net/rfc/rfc131.txt, ftp://ftp.internic.
net/rfc/rfc156.txt,   ftp://ftp.internic.net/rfc/rfc99.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc116.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc131.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc156.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc99.txt.

Updates RFC0099 [99]. Updated by RFC0131, RFC0156 [131, 156]. Status: UNKNOWN.

RFC0117

[117] J. Wong. RFC 117: Some comments on the official protocol, April

7, 1971. URL ftp://ftp.internic.net/rfc/rfc117.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc117.txt. Status: UNKNOWN. Not

online.

RFC0118

[118] R. W. Watson. RFC 118: Recommendations for facility documen-

tation, April 16, 1971. URL ftp://ftp.internic.net/rfc/rfc118.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc118.txt.  Status:

UNKNOWN. Not online.

RFC0119

[119] M. Krilanovich. RFC 119: Network Fortran subprograms, April

21, 1971. URL ftp://ftp.internic.net/rfc/rfc119.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc119.txt. Status: UNKNOWN. Not

online.

RFC0120

[120] M. Krilanovich. RFC 120: Network PL1 subprograms, April 21,

1971. URL ftp://ftp.internic.net/rfc/rfc120.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc120.txt. Status: UNKNOWN.

RFC0121

[121] M. Krilanovich. RFC 121: Network on-line operators, April 21,

1971. URL ftp://ftp.internic.net/rfc/rfc121.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc121.txt. Status: UNKNOWN. Not

online.

RFC0122

[122] J. E. White. RFC 122: Network specifications for UCSB’s simple-minded

file system, April 26, 1971. URL ftp://ftp.internic.net/rfc/
rfc122.txt,   ftp://ftp.internic.net/rfc/rfc217.txt, ftp://
ftp.internic.net/rfc/rfc269.txt,   ftp://ftp.internic.net/rfc/
rfc399.txt, ftp://ftp.internic.net/rfc/rfc43.txt,   ftp://ftp.
internic.net/rfc/rfc431.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc122.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc217.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc269.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc399.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc43.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc431.txt.

Updated by RFC0217, RFC0269, RFC0399, RFC0043, RFC0431 [217, 264, 394, 43, 425]. Status: UNKNOWN. Not online.

RFC0123

[123] S. D. Crocker. RFC 123: Proffered official ICP, April 20, 1971. URL ftp:

//ftp.internic.net/rfc/rfc101.txt,   ftp://ftp.internic.net/
rfc/rfc123.txt, ftp://ftp.internic.net/rfc/rfc127.txt,   ftp:
//ftp.internic.net/rfc/rfc143.txt, ftp://ftp.internic.net/rfc/
rfc148.txt,   ftp://ftp.internic.net/rfc/rfc165.txt, ftp://
ftp.internic.net/rfc/rfc66.txt,   ftp://ftp.internic.net/rfc/
rfc80.txt, ftp://ftp.internic.net/rfc/rfc98.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc101.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc123.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc127.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc143.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc148.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc165.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc66.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc80.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc98.txt. Obsoleted by RFC0165 [165].

Obsoletes RFC0066, RFC0080 [66, 80]. Updates RFC0098, RFC0101 [98, 101]. Updated by RFC0127, RFC0148, RFC0143 [127, 148, 143]. Status: UNKNOWN.

RFC0124

[124] J. T. Melvin. RFC 124: Typographical error in RFC 107, April

19, 1971. URL ftp://ftp.internic.net/rfc/rfc107.txt,   ftp://
ftp.internic.net/rfc/rfc124.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc107.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc124.

txt. Updates RFC0107 [107]. Status: UNKNOWN.

RFC0125

[125] J. McConnell. RFC 125: Response to RFC 86: Proposal for network standard format for a graphics data stream, April 18, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc125.txt,   ftp://ftp.internic.net/
rfc/rfc177.txt, ftp://ftp.internic.net/rfc/rfc86.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc125.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc177.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc86.

txt. Updates RFC0086 [86]. Updated by RFC0177 [177]. Status: UN-KNOWN.

RFC0126

[126] J. McConnell. RFC 126: Graphics facilities at ames research center, April

18, 1971. URL ftp://ftp.internic.net/rfc/rfc126.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc126.txt. Status: UNKNOWN.

RFC0127

[127] J. Postel. RFC 127: Comments on RFC 123, April 20, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc123.txt,   ftp://ftp.internic.net/
rfc/rfc127.txt, ftp://ftp.internic.net/rfc/rfc145.txt,   ftp:
//ftp.internic.net/rfc/rfc151.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc123.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc127.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc145.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc151.txt. Obsoleted by RFC0145

[145]. Updates RFC0123 [123]. Updated by RFC0151 [151]. Status: UN-KNOWN.

RFC0128

[128] J. Postel. RFC 128: Bytes, April 21, 1971. URL ftp://ftp.

internic.net/rfc/rfc128.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc128.txt. Status: UNKNOWN. Not online.

RFC0129

[129] E. Harslem, J. F. Heafner, and E. W. Meyer, Jr. RFC 129: Request for comments on socket name structure, April 22, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc129.txt,   ftp://ftp.internic.net/
rfc/rfc147.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc129.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc147.txt. Updated by RFC0147

[147]. Status: UNKNOWN.

RFC0130

[130] J. F. Heafner. RFC 130: Response to RFC 111: Pressure from the chairman, April 22, 1971. URL ftp://ftp.internic.net/rfc/

rfc111.txt,   ftp://ftp.internic.net/rfc/rfc130.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc111.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc130.txt. Updates RFC0111 [111]. Status: UN-KNOWN. Not online.

RFC0131

[131] E. Harslem and J. F. Heafner. RFC 131: Response to RFC 116: May

NWG meeting, April 22, 1971. URL ftp://ftp.internic.net/rfc/
rfc116.txt,   ftp://ftp.internic.net/rfc/rfc131.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc116.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc131.txt. Updates RFC0116 [116]. Status: UN-KNOWN.

RFC0132

[132] J. E. White. RFC 132: Typographical error in RFC 107, April

28, 1971. URL ftp://ftp.internic.net/rfc/rfc107.txt,   ftp://
ftp.internic.net/rfc/rfc132.txt, ftp://ftp.internic.net/rfc/
rfc154.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc107.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc132.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc154.txt. Obsoleted by RFC0154 [154]. Updates RFC0107 [107]. Status: UNKNOWN.

RFC0133

[133] R. L. Sundberg. RFC 133: File transfer and recovery, April 27,

1971. URL ftp://ftp.internic.net/rfc/rfc133.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc133.txt. Status: UNKNOWN. Not

online.

RFC0134

[134] A. Vezza. RFC 134: Network graphics meeting, April 29, 1971. URL ftp:

//ftp.internic.net/rfc/rfc134.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc134.txt. Status: UNKNOWN.

RFC0135

[135] W. Hathaway. RFC 135: Response to NWG/RFC 110, April 29,

1971. URL ftp://ftp.internic.net/rfc/rfc110.txt,   ftp://
ftp.internic.net/rfc/rfc135.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc110.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc135.

txt. Updates RFC0110 [110]. Status: UNKNOWN. Not online.

RFC0136

[136] R. E. Kahn. RFC 136: Host accounting and administrative proce-

dures, April 29, 1971. URL ftp://ftp.internic.net/rfc/rfc136.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc136.txt.  Status:

UNKNOWN. Not online.

RFC0137

[137] T. C. O’Sullivan. RFC 137: Telnet Protocol — a proposed

document, April 30, 1971. URL ftp://ftp.internic.net/rfc/
rfc137.txt,   ftp://ftp.internic.net/rfc/rfc139.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc137.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc139.txt. Updated by RFC0139 [139]. Status: UN-KNOWN. Not online.

RFC0138

[138] R. H. Anderson, V. G. Cerf, E. Harslem, J. F. Heafner, J. Madden, R. M. Metcalfe, A. Shoshani, J. E. White, and D. C. M. Wood. RFC 138: Status report on proposed data reconfiguration service, April

28, 1971. URL ftp://ftp.internic.net/rfc/rfc138.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc138.txt. Status: UNKNOWN. Not

online.

RFC0139

[139] T. C. O’Sullivan. RFC 139: Discussion of Telnet Protocol, May

7, 1971. URL ftp://ftp.internic.net/rfc/rfc137.txt,   ftp://
ftp.internic.net/rfc/rfc139.txt, ftp://ftp.internic.net/rfc/
rfc158.txt,   ftp://ftp.internic.net/rfc/rfc393.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc137.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc139.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc158.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc393.txt. See also

RFC0393 [388]. Updates RFC0137 [137]. Updated by RFC0158 [158]. Status: UNKNOWN. Not online.

RFC0140

[140] S. D. Crocker. RFC 140: Agenda for the may NWG meeting, May

4, 1971. URL ftp://ftp.internic.net/rfc/rfc140.txt,   ftp://

ftp.internic.net/rfc/rfc149.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc140.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc149.

txt. Updated by RFC0149 [149]. Status: UNKNOWN.

RFC0141

[141] E. Harslem and J. F. Heafner. RFC 141: Comments on RFC 114: A file transfer protocol, April 29, 1971. URL ftp://ftp.internic.net/rfc/

rfc114.txt,   ftp://ftp.internic.net/rfc/rfc141.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc114.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc141.txt. Updates RFC0114 [114]. Status: UN-KNOWN.

RFC0142

[142] C. Kline and J. Wong. RFC 142: Time-out mechanism in the host-host

protocol, May 3, 1971. URL ftp://ftp.internic.net/rfc/rfc142.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc142.txt.  Status:

UNKNOWN.

RFC0143

[143] W. Naylor, J. Wong, C. Kline, and J. Postel. RFC 143: Regarding proffered official ICP, May 3, 1971. URL ftp://ftp.internic.net/rfc/

rfc123.txt,   ftp://ftp.internic.net/rfc/rfc143.txt, ftp://
ftp.internic.net/rfc/rfc145.txt,   ftp://ftp.internic.net/rfc/
rfc165.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc123.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc143.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc145.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc165.txt. Obsoleted by RFC0165 [165]. Updates RFC0123, RFC0145 [123, 145]. Status: UNKNOWN.

RFC0144

[144] A. Shoshani. RFC 144: Data sharing on computer networks, April

30, 1971. URL ftp://ftp.internic.net/rfc/rfc144.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc144.txt. Status: UNKNOWN. Not

online.

RFC0145

[145] J. Postel. RFC 145: Initial Connection Protocol control com-

mands, May 4, 1971. URL ftp://ftp.internic.net/rfc/rfc127.txt,
 ftp://ftp.internic.net/rfc/rfc143.txt, ftp://ftp.internic.
net/rfc/rfc145.txt,   ftp://ftp.internic.net/rfc/rfc165.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc127.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc143.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc145.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc165.txt.

Obsoleted by RFC0165 [165]. Obsoletes RFC0127 [127]. Updated by RFC0143 [143]. Status: UNKNOWN. Not online.

RFC0146

[146] P. M. Karp, D. B. McKay, and D. C. M. Wood. RFC 146: Views on issues relevant to data sharing on computer networks, May

12, 1971. URL ftp://ftp.internic.net/rfc/rfc146.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc146.txt. Status: UNKNOWN.

RFC0147

[147] J. M. Winett. RFC 147: Definition of a socket, May 7, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc129.txt,   ftp://ftp.internic.net/
rfc/rfc147.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc129.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc147.txt. Updates RFC0129

[129]. Status: UNKNOWN.

RFC0148

[148] A. K. Bhushan. RFC 148: Comments on RFC 123, May 7,

1971. URL ftp://ftp.internic.net/rfc/rfc123.txt,   ftp://
ftp.internic.net/rfc/rfc148.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc123.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc148.

txt. Updates RFC0123 [123]. Status: UNKNOWN.

RFC0149

[149] S. D. Crocker. RFC 149: Best laid plans, May 10, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc140.txt,   ftp://ftp.internic.net/
rfc/rfc149.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc140.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc149.txt. Updates RFC0140

[140]. Status: UNKNOWN.

RFC0150

[150] R. B. Kalin. RFC 150: Use of IPC facilities: A working paper, May

5, 1971. URL ftp://ftp.internic.net/rfc/rfc150.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc150.txt. Status: UNKNOWN.

RFC0151

[151] A. Shoshani. RFC 151: Comments on a proffered official ICP:

RFCs 123, 127, May 10, 1971. URL ftp://ftp.internic.net/rfc/
rfc127.txt,   ftp://ftp.internic.net/rfc/rfc151.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc127.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc151.txt. Updates RFC0127 [127]. Status: UN-KNOWN.

RFC0152

[152] M. Wilber. RFC 152: SRI artificial intelligence status report, May

10, 1971. URL ftp://ftp.internic.net/rfc/rfc152.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc152.txt. Status: UNKNOWN.

RFC0153

[153] J. T. Melvin and R. W. Watson. RFC 153: SRI ARC-NIC status, May

15, 1971. URL ftp://ftp.internic.net/rfc/rfc153.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc153.txt. Status: UNKNOWN. Not

online.

RFC0154

[154] S. D. Crocker. RFC 154: Exposition style, May 12, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc107.txt,   ftp://ftp.internic.net/
rfc/rfc132.txt, ftp://ftp.internic.net/rfc/rfc154.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc107.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc132.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc154.

txt. Obsoletes RFC0132 [132]. Updates RFC107, RFC0107 [107, 107]. Status: UNKNOWN.

RFC0155

[155] J. B. North. RFC 155: ARPA network mailing lists, May 1,

1971. URL ftp://ftp.internic.net/rfc/rfc155.txt,   ftp://
ftp.internic.net/rfc/rfc168.txt, ftp://ftp.internic.net/rfc/
rfc95.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc155.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc168.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc95.txt. Obsoleted by RFC0168 [168]. Obsoletes RFC0095 [95]. Status: UNKNOWN.

RFC0156

[156] J. Bouknight. RFC 156: Status of the Illinois site: Response

to RFC 116, April 26, 1971. URL ftp://ftp.internic.net/rfc/
rfc116.txt,   ftp://ftp.internic.net/rfc/rfc156.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc116.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc156.txt. Updates RFC0116 [116]. Status: UN-KNOWN.

RFC0157

[157] V. G. Cerf. RFC 157: Invitation to the second symposium on problems in the optimization of data communications systems, May

12, 1971. URL ftp://ftp.internic.net/rfc/rfc157.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc157.txt. Status: UNKNOWN. Not

online.

RFC0158

[158] T. C. O’Sullivan. RFC 158: Telnet Protocol: A proposed docu-

ment, May 19, 1971. URL ftp://ftp.internic.net/rfc/rfc139.txt,
 ftp://ftp.internic.net/rfc/rfc158.txt, ftp://ftp.internic.
net/rfc/rfc318.txt,   ftp://ftp.internic.net/rfc/rfc393.txt,
ftp://ftp.internic.net/rfc/rfc495.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc139.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc158.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc318.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc393.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc495.txt. See also RFC0393 [388]. Obsoleted by RFC0495 [489]. Updates RFC0139 [139]. Updated by RFC0318 [313]. Status: UNKNOWN. Not online.

RFC0159

[159] ???? RFC 159: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc159.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc159.txt.

RFC0160

[160] Network Information Center. Stanford Research Institute. RFC 160:

RFC brief list, May 18, 1971. URL ftp://ftp.internic.net/rfc/
nic6716.txt,   ftp://ftp.internic.net/rfc/rfc160.txt, ftp:/
/ftp.internic.net/rfc/rfc200.txt,   ftp://ftp.internic.net/
rfc/rfc999.txt,   ftp://ftp.math.utah.edu/pub/rfc/nic6716.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc160.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc200.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc999.txt. Obsoleted by RFC0200, RFC0999 [200, 989]. Updates NIC6716 [?]. Status: UNKNOWN. Not online.

RFC0161

[161] A. Shoshani. RFC 161: Solution to the race condition in the ICP, May

19, 1971. URL ftp://ftp.internic.net/rfc/rfc161.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc161.txt. Status: UNKNOWN. Not

online.

RFC0162

[162] M. Kampe. RFC 162: NETBUGGER3, May 22, 1971. URL ftp://

ftp.internic.net/rfc/rfc162.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc162.txt. Status: UNKNOWN. Not online.

RFC0163

[163] V. G. Cerf. RFC 163: Data transfer protocols, May 19, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc163.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc163.txt. Status: UNKNOWN.

RFC0164

[164] J. F. Heafner. RFC 164: Minutes of network working group meeting, 5/16 through 5/1971, May 25, 1971. URL ftp://ftp.

internic.net/rfc/rfc164.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc164.txt. Status: UNKNOWN. Not online.

RFC0165

[165] J. Postel. RFC 165: Proffered official initial connection protocol,

May 25, 1971. URL ftp://ftp.internic.net/rfc/nic7101.txt,
 ftp://ftp.internic.net/rfc/rfc123.txt, ftp://ftp.internic.
net/rfc/rfc143.txt,   ftp://ftp.internic.net/rfc/rfc145.txt,
ftp://ftp.internic.net/rfc/rfc165.txt,   ftp://ftp.math.utah.
edu/pub/rfc/nic7101.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc123.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc143.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc145.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc165.txt.  Obsoletes RFC0145, RFC0143,

RFC0123 [145, 143, 123]. Updated by NIC7101 [?]. Status: UNKNOWN. Not online.

RFC0166

[166] R. H. Anderson, V. G. Cerf, E. Harslem, J. F. Heafner, J. Madden, R. M. Metcalfe, A. Shoshani, J. E. White, and D. C. M. Wood. RFC 166: Data reconfiguration service: An implementation specification, May

25, 1971. URL ftp://ftp.internic.net/rfc/rfc166.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc166.txt. Status: UNKNOWN. Not

online.

RFC0167

[167] A. K. Bhushan, R. M. Metcalfe, and J. M. Winett. RFC 167: Socket conventions reconsidered, May 24, 1971. URL ftp://ftp.internic.net/

rfc/rfc129.txt,   ftp://ftp.internic.net/rfc/rfc147.txt, ftp:
//ftp.internic.net/rfc/rfc167.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc129.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc147.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc167.txt. See also

RFC0147, RFC0129 [147, 129]. Status: UNKNOWN.

RFC0168

[168] J. B. North. RFC 168: ARPA network mailing lists, May 26,

1971. URL ftp://ftp.internic.net/rfc/rfc155.txt,   ftp://
ftp.internic.net/rfc/rfc168.txt, ftp://ftp.internic.net/rfc/
rfc211.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc155.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc168.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc211.txt. Obsoleted by RFC0211 [211]. Obsoletes RFC0155 [155]. Status: UNKNOWN.

RFC0169

[169] S. D. Crocker. RFC 169: Computer networks, May 27, 1971. URL ftp://

ftp.internic.net/rfc/rfc169.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc169.txt. Status: UNKNOWN. Not online.

RFC0170

[170] Network Information Center. Stanford Research Institute. RFC 170: RFC list by number, June 1, 1971. URL ftp://ftp.internic.net/rfc/

rfc170.txt,   ftp://ftp.internic.net/rfc/rfc200.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc170.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc200.txt. Obsoleted by RFC0200 [200]. Status: UN-KNOWN.

RFC0171

[171] A. Bhushan, R. Braden, W. Crowther, E. Harslem, J. Heafner, A. McKenzie, J. Melvin, B. Sundberg, D. Watson, and J. White. RFC 171: The Data Transfer Protocol, June 1, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc114.txt,   ftp://ftp.internic.net/
rfc/rfc171.txt, ftp://ftp.internic.net/rfc/rfc238.txt,   ftp:
//ftp.internic.net/rfc/rfc264.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc114.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc171.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc238.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc264.txt. Obsoleted by RFC0264

[259]. Updates RFC0114 [114]. Updated by RFC0238 [238]. Not online. Status: UNKNOWN.

RFC0172

[172] A. Bhushan, B. Braden, W. Crowther, E. Harslem, J. Heafner, A. McKenzie, J. Melvin, B. Sundberg, D. Watson, and J. White. RFC 172: The file transfer protocol, June 23, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc114.txt,   ftp://ftp.internic.net/
rfc/rfc172.txt, ftp://ftp.internic.net/rfc/rfc238.txt,   ftp:
//ftp.internic.net/rfc/rfc265.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc114.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc172.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc238.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc265.txt. Obsoleted by RFC0265

[260]. Updates RFC0114 [114]. Updated by RFC0238 [238]. Status: UN-KNOWN.

RFC0173

[173] P. M. Karp and D. B. McKay. RFC 173: Network data management committee meeting announcement, June 4, 1971. URL ftp://ftp.

internic.net/rfc/rfc173.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc173.txt. Status: UNKNOWN.

RFC0174

[174] J. Postel and V. G. Cerf. RFC 174: UCLA — computer science graphics

overview, June 8, 1971. URL ftp://ftp.internic.net/rfc/rfc174.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc174.txt.  Status:

UNKNOWN.

RFC0175

[175] E. Harslem and J. F. Heafner. RFC 175: Comments on “socket conventions reconsidered”, June 11, 1971. URL ftp://ftp.internic.net/

rfc/rfc175.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc175.txt.

Status: UNKNOWN.

RFC0176

[176] A. K. Bhushan, R. Kanodia, R. M. Metcalfe, and J. Postel. RFC 176: Comments on “byte size for connections”, June 14, 1971. URL ftp://

ftp.internic.net/rfc/rfc176.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc176.txt. Status: UNKNOWN.

RFC0177

[177] J. McConnell. RFC 177: Device independent graphical display

description, June 15, 1971. URL ftp://ftp.internic.net/rfc/
rfc125.txt,   ftp://ftp.internic.net/rfc/rfc177.txt, ftp://
ftp.internic.net/rfc/rfc181.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc125.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc177.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc181.txt. Updates

RFC0125 [125]. Updated by RFC0181 [181]. Status: UNKNOWN. Not online.

RFC0178

[178] I. W. Cotton. RFC 178: Network graphic attention handling, June

27, 1971. URL ftp://ftp.internic.net/rfc/rfc178.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc178.txt. Status: UNKNOWN. Not

online.

RFC0179

[179] A. M. McKenzie. RFC 179: Link number assignments, June 22,

1971. URL ftp://ftp.internic.net/rfc/rfc107.txt,   ftp://

ftp.internic.net/rfc/rfc179.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc107.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc179.

txt. Updates RFC0107 [107]. Status: UNKNOWN.

RFC0180

[180] A. M. McKenzie. RFC 180: File system questionnaire, June 25,

1971. URL ftp://ftp.internic.net/rfc/rfc180.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc180.txt. Status: UNKNOWN.

RFC0181

[181] J. McConnell. RFC 181: Modifications to RFC 177, July 21,

1971. URL ftp://ftp.internic.net/rfc/rfc177.txt,   ftp://
ftp.internic.net/rfc/rfc181.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc177.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc181.

txt. Updates RFC0177 [177]. Status: UNKNOWN. Format: TXT=4892 bytes.

RFC0182

[182] J. B. North. RFC 182: Compilation of list of relevant site reports, June

25, 1971. URL ftp://ftp.internic.net/rfc/rfc182.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc182.txt. Status: UNKNOWN. Not

online.

RFC0183

[183] J. M. Winett. RFC 183: EBCDIC codes and their mapping to

ASCII, July 21, 1971. URL ftp://ftp.internic.net/rfc/rfc183.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc183.txt.  Status:

UNKNOWN. Not online.

RFC0184

[184] K. C. Kelley. RFC 184: Proposed graphic display modes, July

6, 1971. URL ftp://ftp.internic.net/rfc/rfc184.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc184.txt. Status: UNKNOWN. Not

online.

RFC0185

[185] J. B. North. RFC 185: NIC distribution of manuals and handbooks, July

7, 1971. URL ftp://ftp.internic.net/rfc/rfc185.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc185.txt. Status: UNKNOWN.

RFC0186

[186] J. C. Michener. RFC 186: Network graphics loader, July 12,

1971. URL ftp://ftp.internic.net/rfc/rfc186.txt,   ftp://

ftp.math.utah.edu/pub/rfc/rfc186.txt. Status: UNKNOWN. Not

online.

RFC0187

[187] D. B. McKay and D. P. Karp. RFC 187: Network/440 protocol

concept, July 1, 1971. URL ftp://ftp.internic.net/rfc/rfc187.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc187.txt.  Status:

UNKNOWN.

RFC0188

[188] P. M. Karp and D. B. McKay. RFC 188: Data management meeting announcement, January 28, 1971. URL ftp://ftp.internic.net/rfc/

rfc188.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc188.txt. Sta-

tus: UNKNOWN. Not online.

RFC0189

[189] R. T. Braden. RFC 189: Interim NETRJS specifications, July 15,

1971. URL ftp://ftp.internic.net/rfc/rfc189.txt,   ftp://
ftp.internic.net/rfc/rfc283.txt, ftp://ftp.internic.net/rfc/
rfc599.txt,   ftp://ftp.internic.net/rfc/rfc88.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc189.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc283.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc599.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc88.txt. Obsoleted by

RFC0599 [592]. Obsoletes RFC0088 [88]. Updated by RFC0283 [278]. Status: UNKNOWN. Not online.

RFC0190

[190] L. P. Deutsch. RFC 190: DEC PDP-10-IMLAC communications sys-

tem, July 13, 1971. URL ftp://ftp.internic.net/rfc/rfc190.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc190.txt.  Status:

UNKNOWN. Not online.

RFC0191

[191] C. H. Irby. RFC 191: Graphics implementation and conceptualization at augmentation research center, July 13, 1971. URL ftp://ftp.

internic.net/rfc/rfc191.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc191.txt. Status: UNKNOWN.

RFC0192

[192] R. W. Watson. RFC 192: Some factors which a network graphics protocol must consider, July 12, 1971. URL ftp://ftp.

internic.net/rfc/rfc192.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc192.txt. Status: UNKNOWN. Not online.

RFC0193

[193] E. Harslem and J. F. Heafner. RFC 193: Network checkout, July

14, 1971. URL ftp://ftp.internic.net/rfc/rfc193.txt,   ftp://
ftp.internic.net/rfc/rfc198.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc193.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc198.

txt. Obsoleted by RFC0198 [198]. Status: UNKNOWN. Updated by RFC0198 [198].

RFC0194

[194] V. Cerf, E. Harslem, J. Heafner, R. Metcalfe, and J. White. RFC 194: Data reconfiguration service — compiler/interpreter implementa-

tion notes, July 14, 1971. URL ftp://ftp.internic.net/rfc/rfc194.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc194.txt. Format:

TXT=32716 bytes. Status: UNKNOWN.

RFC0195

[195] G. H. Mealy. RFC 195: Data computers-data descriptions and access

language, July 16, 1971. URL ftp://ftp.internic.net/rfc/rfc195.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc195.txt.  Status:

UNKNOWN. Not online.

RFC0196

[196] R. W. Watson. RFC 196: Mail box protocol, July 20, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc196.txt,   ftp://ftp.internic.net/
rfc/rfc221.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc196.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc221.txt.  Obsoleted by

RFC0221 [221]. Status: UNKNOWN.

RFC0197

[197] A. Shoshani and E. Harslem. RFC 197: Initial connection protocol —

reviewed, July 14, 1971. URL ftp://ftp.internic.net/rfc/rfc197.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc197.txt.  Status:

UNKNOWN.

RFC0198

[198] J. F. Heafner. RFC 198: Site certification — Lincoln Labs 360/

67, July 20, 1971. URL ftp://ftp.internic.net/rfc/rfc193.txt,
 ftp://ftp.internic.net/rfc/rfc198.txt, ftp://ftp.internic.
net/rfc/rfc214.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc193.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc198.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc214.txt. Obsoleted by RFC0214

[214]. Obsoletes RFC0193 [193]. Updates RFC0193 [193]. Status: UN-KNOWN.

RFC0199

[199] T. Williams. RFC 199: Suggestions for a network data-tablet graphics

protocol, July 15, 1971. URL ftp://ftp.internic.net/rfc/rfc199.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc199.txt.  Status:

UNKNOWN. Not online.

RFC0200

[200] J. B. North. RFC 200: RFC list by number, August 1, 1971. URL ftp://

ftp.internic.net/rfc/nic7724.txt,   ftp://ftp.internic.net/
rfc/rfc160.txt, ftp://ftp.internic.net/rfc/rfc170.txt,   ftp:
//ftp.internic.net/rfc/rfc200.txt,   ftp://ftp.math.utah.edu/
pub/rfc/nic7724.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc160.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc170.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc200.txt. Obsoleted by NIC7724 [?].

Obsoletes RFC0170, RFC0160 [170, 160]. Status: UNKNOWN. Not online.

RFC0201

[201] ???? RFC 201: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc201.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc201.txt.

RFC0202

[202] S. M. Wolfe and J. Postel. RFC 202: Possible deadlock in ICP, July

26, 1971. URL ftp://ftp.internic.net/rfc/rfc202.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc202.txt. Status: UNKNOWN. For-

mat: TXT=2796 bytes.

RFC0203

[203] R. B. Kalin. RFC 203: Achieving reliable communication, August

10, 1971. URL ftp://ftp.internic.net/rfc/rfc203.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc203.txt. Status: UNKNOWN. Not

online.

RFC0204

[204] J. Postel. RFC 204: Sockets in use, August 5, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc204.txt,   ftp://ftp.internic.net/
rfc/rfc234.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc204.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc234.txt. Updated by RFC0234

[234]. Status: UNKNOWN. Format: TXT=1379 bytes.

RFC0205

[205] R. T. Braden. RFC 205: NETCRT — a character display proto-

col, August 6, 1971. URL ftp://ftp.internic.net/rfc/rfc205.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc205.txt.  Status:

UNKNOWN. Not online.

RFC0206

[206] J. E. White. RFC 206: User Telnet — description of an initial implemen-

tation, August 9, 1971. URL ftp://ftp.internic.net/rfc/rfc206.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc206.txt.  Status:

UNKNOWN. Not online.

RFC0207

[207] A. Vezza. RFC 207: September network working group meeting, August

9, 1971. URL ftp://ftp.internic.net/rfc/rfc207.txt,   ftp://
ftp.internic.net/rfc/rfc212.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc207.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc212.

txt. Obsoleted by RFC0212 [212]. Status: UNKNOWN. Format: TXT=3356 bytes.

RFC0208

[208] A. M. McKenzie. RFC 208: Address tables, August 9, 1971. URL ftp://

ftp.internic.net/rfc/rfc208.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc208.txt. Status: UNKNOWN. Not online.

RFC0209

[209] B. Cosell. RFC 209: Host/IMP interface documentation, August

13, 1971. URL ftp://ftp.internic.net/rfc/rfc209.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc209.txt. Status: UNKNOWN. Not

online.

RFC0210

[210] W. Conrad. RFC 210: Improvement of flow control, August 16,

1971. URL ftp://ftp.internic.net/rfc/rfc210.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc210.txt. Status: UNKNOWN. For-

mat: TXT=3329 bytes.

RFC0211

[211] J. B. North. RFC 211: ARPA network mailing lists, August 18,

1971. URL ftp://ftp.internic.net/rfc/rfc168.txt,   ftp://
ftp.internic.net/rfc/rfc211.txt, ftp://ftp.internic.net/rfc/
rfc300.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc168.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc211.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc300.txt. Obsoleted by RFC0300 [295]. Obsoletes RFC0168 [168]. Status: UNKNOWN. Not online.

RFC0212

[212] Information Sciences Institute. University of Southern California. RFC 212: NWG meeting on network usage, August 23, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc207.txt,   ftp://ftp.internic.net/
rfc/rfc212.txt, ftp://ftp.internic.net/rfc/rfc222.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc207.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc212.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc222.

txt. Obsoletes RFC0207 [207]. Updated by RFC0222 [222]. Status: UN-KNOWN.

RFC0213

[213] B. Cosell. RFC 213: IMP system change notification, August

20, 1971. URL ftp://ftp.internic.net/rfc/rfc213.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc213.txt. Status: UNKNOWN.

RFC0214

[214] E. Harslem. RFC 214: Network checkpoint, August 21, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc198.txt,   ftp://ftp.internic.net/
rfc/rfc214.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc198.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc214.txt. Obsoletes RFC0198

[198]. Status: UNKNOWN.

RFC0215

[215] A. M. McKenzie. RFC 215: NCP, ICP, and Telnet: The terminal IMP implementation, August 30, 1971. URL ftp://ftp.

internic.net/rfc/rfc215.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc215.txt. Status: UNKNOWN.

RFC0216

[216] J. E. White. RFC 216: Telnet access to UCSB’s on-line system, Septem-

ber 8, 1971. URL ftp://ftp.internic.net/rfc/rfc216.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc216.txt. Status: UNKNOWN.

Not online.

RFC0217

[217] J. E. White. RFC 217: Specifications changes for OLS, RJE/RJOR,

and SMFS, September 8, 1971. URL ftp://ftp.internic.net/rfc/
rfc105.txt,   ftp://ftp.internic.net/rfc/rfc122.txt, ftp://
ftp.internic.net/rfc/rfc217.txt,   ftp://ftp.internic.net/rfc/
rfc74.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc105.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc122.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc217.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc74.

txt. Updates RFC0074, RFC0105, RFC0122 [74, 105, 122]. Status: UN-KNOWN.

RFC0218

[218] B. Cosell. RFC 218: Changing the IMP status reporting facility, Septem-

ber 8, 1971. URL ftp://ftp.internic.net/rfc/rfc218.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc218.txt. Status: UNKNOWN.

RFC0219

[219] R. Winter. RFC 219: User’s view of the datacomputer, September

3, 1971. URL ftp://ftp.internic.net/rfc/rfc219.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc219.txt. Status: UNKNOWN. Not

online.

RFC0220

[220] ???? RFC 220: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc220.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc220.txt.

RFC0221

[221] R. W. Watson. RFC 221: Mail box protocol: Version 2, Au-

gust 27, 1971. URL ftp://ftp.internic.net/rfc/rfc196.txt,
 ftp://ftp.internic.net/rfc/rfc221.txt, ftp://ftp.internic.
net/rfc/rfc278.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc196.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc221.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc278.txt. Obsoleted by RFC0278

[273]. Obsoletes RFC0196 [196]. Status: UNKNOWN. Not online.

RFC0222

[222] R. M. Metcalfe. RFC 222: Subject: System programmer’s work-

shop, September 13, 1971. URL ftp://ftp.internic.net/rfc/
rfc212.txt,   ftp://ftp.internic.net/rfc/rfc222.txt, ftp://
ftp.internic.net/rfc/rfc234.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc212.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc222.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc234.txt. Updates

RFC0212 [212]. Updated by RFC0234 [234]. Status: UNKNOWN. Not online.

RFC0223

[223] J. T. Melvin and R. W. Watson. RFC 223: Network information center schedule for network users, September 14, 1971. URL ftp://ftp.

internic.net/rfc/rfc223.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc223.txt. Status: UNKNOWN.

RFC0224

[224] A. M. McKenzie. RFC 224: Comments on mailbox protocol, September

14, 1971. URL ftp://ftp.internic.net/rfc/rfc224.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc224.txt. Status: UNKNOWN.

RFC0225

[225] E. Harslem and R. Stoughton. RFC 225: Rand/UCSB network graphics experiment, September 13, 1971. URL ftp://ftp.internic.net/rfc/

rfc225.txt,   ftp://ftp.internic.net/rfc/rfc74.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc225.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc74.txt. Updates RFC0074 [74]. Status: UNKNOWN. Not online.

RFC0226

[226] P. M. Karp. RFC 226: Standardization of host mnemonics, September

20, 1971. URL ftp://ftp.internic.net/rfc/rfc226.txt,   ftp://
ftp.internic.net/rfc/rfc247.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc226.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc247.

txt. Obsoleted by RFC0247 [247]. Status: UNKNOWN.

RFC0227

[227] J. F. Heafner and E. Harslem. RFC 227: Data transfer rates (Rand/

UCLA), September 17, 1971. URL ftp://ftp.internic.net/rfc/
rfc113.txt,   ftp://ftp.internic.net/rfc/rfc227.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc113.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc227.txt. Updates RFC0113 [113]. Status: UN-KNOWN. Not online.

RFC0228

[228] D. C. Walden. RFC 228: Clarification, September 22, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc228.txt,   ftp://ftp.internic.net/
rfc/rfc70.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc228.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc70.txt. Updates RFC0070

[70]. Status: UNKNOWN.

RFC0229

[229] J. Postel. RFC 229: Standard host names, September 22, 1971. URL

ftp://ftp.internic.net/rfc/rfc229.txt,   ftp://ftp.internic.
net/rfc/rfc236.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc229.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc236.txt. Obsoleted

by RFC0236 [236]. Status: UNKNOWN.

RFC0230

[230] T. N. Pyke, Jr. RFC 230: Toward reliable operation of minicomputerbased terminals on a TIP, September 24, 1971. URL ftp://ftp.

internic.net/rfc/rfc230.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc230.txt. Status: UNKNOWN.

RFC0231

[231] J. F. Heafner and E. Harslem. RFC 231: Service center standards for remote usage: A user’s view, September 21, 1971. URL ftp://ftp.

internic.net/rfc/rfc231.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc231.txt. Status: UNKNOWN.

RFC0232

[232] A. Vezza. RFC 232: Postponement of network graphics meeting,

September 23, 1971. URL ftp://ftp.internic.net/rfc/rfc232.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc232.txt.  Status:

UNKNOWN.

RFC0233

[233] A. K. Bhushan and R. M. Metcalfe. RFC 233: Standardization of host call letters, September 28, 1971. URL ftp://ftp.

internic.net/rfc/rfc233.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc233.txt. Status: UNKNOWN.

RFC0234

[234] A. Vezza. RFC 234: Network working group meeting schedule,

October 5, 1971. URL ftp://ftp.internic.net/rfc/rfc204.txt,
 ftp://ftp.internic.net/rfc/rfc222.txt, ftp://ftp.internic.
net/rfc/rfc234.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc204.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc222.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc234.txt. Updates RFC0222, RFC0204

[222, 204]. Status: UNKNOWN. Not online.

RFC0235

[235] E. Westheimer. RFC 235: Site status, September 27, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc235.txt,   ftp://ftp.internic.net/
rfc/rfc240.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc235.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc240.txt.  Obsoleted by

RFC0240 [240]. Status: UNKNOWN. Not online.

RFC0236

[236] J. Postel. RFC 236: Standard host names, September 27, 1971. URL

ftp://ftp.internic.net/rfc/rfc229.txt,   ftp://ftp.internic.

net/rfc/rfc236.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc229.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc236.txt. Obsoletes

RFC0229 [229]. Status: UNKNOWN.

RFC0237

[237] R. W. Watson. RFC 237: NIC view of standard host names, September

29, 1971. URL ftp://ftp.internic.net/rfc/rfc237.txt,   ftp://
ftp.internic.net/rfc/rfc273.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc237.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc273.

txt. Obsoleted by RFC0273 [268]. Status: UNKNOWN. Not online.

RFC0238

[238] R. T. Braden. RFC 238: Comments on DTP and FTP proposals,

September 29, 1971. URL ftp://ftp.internic.net/rfc/rfc171.txt,
 ftp://ftp.internic.net/rfc/rfc172.txt, ftp://ftp.internic.
net/rfc/rfc238.txt,   ftp://ftp.internic.net/rfc/rfc269.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc171.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc172.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc238.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc269.txt.

Updates RFC0171, RFC0172 [171, 172]. Updated by RFC0269 [264]. Status: UNKNOWN.

RFC0239

[239] R. T. Braden. RFC 239: Host mnemonics proposed in RFC 226 (NIC 7625), September 23, 1971. URL ftp://ftp.internic.net/rfc/

rfc226.txt,   ftp://ftp.internic.net/rfc/rfc229.txt, ftp://
ftp.internic.net/rfc/rfc236.txt,   ftp://ftp.internic.net/rfc/
rfc239.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc226.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc229.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc236.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc239.txt. See also RFC0226, RFC0229, RFC0236 [226, 229, 236]. Status: UNKNOWN.

RFC0240

[240] A. M. McKenzie. RFC 240: Site status, September 30, 1971. URL ftp:/

/ftp.internic.net/rfc/rfc235.txt,   ftp://ftp.internic.net/
rfc/rfc240.txt, ftp://ftp.internic.net/rfc/rfc252.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc235.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc240.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc252.

txt. Obsoleted by RFC0252 [252]. Obsoletes RFC0235 [235]. Status: UNKNOWN. Format: TXT=7948 bytes.

RFC0241

[241] A. M. McKenzie. RFC 241: Connecting computers to MLC ports,

September 29, 1971. URL ftp://ftp.internic.net/rfc/rfc241.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc241.txt.  Status:

UNKNOWN.

RFC0242

[242] L. Haibt and A. P. Mullery. RFC 242: Data descriptive language for shared data, July 19, 1971. URL ftp://ftp.internic.net/rfc/

rfc242.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc242.txt. Sta-

tus: UNKNOWN. Format: TXT=18151 bytes.

RFC0243

[243] A. P. Mullery. RFC 243: Network and data sharing bibliography, October

5, 1971. URL ftp://ftp.internic.net/rfc/rfc243.txt,   ftp://
ftp.internic.net/rfc/rfc290.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc243.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc290.

txt. Obsoleted by RFC0290 [285]. Status: UNKNOWN. Not online.

RFC0244

[244] ???? RFC 244: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc244.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc244.txt.

RFC0245

[245] C. Falls. RFC 245: Reservations for network group meeting, October

5, 1971. URL ftp://ftp.internic.net/rfc/rfc245.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc245.txt. Status: UNKNOWN.

RFC0246

[246] A. Vezza. RFC 246: Network graphics meeting, October 5,

1971. URL ftp://ftp.internic.net/rfc/rfc246.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc246.txt. Status: UNKNOWN.

RFC0247

[247] P. M. Karp. RFC 247: Proffered set of standard host names, October

12, 1971. URL ftp://ftp.internic.net/rfc/rfc226.txt,   ftp://
ftp.internic.net/rfc/rfc247.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc226.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc247.

txt. Obsoletes RFC0226 [226]. Status: UNKNOWN.

RFC0248

RFC0249

[249] R. F. Borelli. RFC 249: Coordination of equipment and supplies pur-

chase, October 8, 1971. URL ftp://ftp.internic.net/rfc/rfc249.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc249.txt.  Status:

UNKNOWN.

RFC0250

[250] H. Brodie. RFC 250: Some thoughts on file transfer, October

7, 1971. URL ftp://ftp.internic.net/rfc/rfc250.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc250.txt. Status: UNKNOWN.

RFC0251

[251] D. Stern. RFC 251: Weather data, October 13, 1971. URL ftp://

ftp.internic.net/rfc/rfc251.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc251.txt. Status: UNKNOWN.

RFC0252

[252] E. Westheimer. RFC 252: Network host status, October 8,

1971. URL ftp://ftp.internic.net/rfc/rfc240.txt,   ftp://
ftp.internic.net/rfc/rfc252.txt, ftp://ftp.internic.net/rfc/
rfc255.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc240.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc252.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc255.txt. Obsoleted by RFC0255 [255]. Obsoletes RFC0240 [240]. Status: UNKNOWN.

RFC0253

[253] J. A. Moorer. RFC 253: Second network graphics meeting details,

October 19, 1971. URL ftp://ftp.internic.net/rfc/rfc253.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc253.txt.  Status: UN-

KNOWN.

RFC0254

[254] A. K. Bhushan. RFC 254: Scenarios for using ARPANET comput-

ers, October 29, 1971. URL ftp://ftp.internic.net/rfc/rfc254.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc254.txt.  Status:

UNKNOWN. Not online.

RFC0255

[255] E. Westheimer. RFC 255: Status of network hosts, October 26,

1971. URL ftp://ftp.internic.net/rfc/rfc252.txt,   ftp://
ftp.internic.net/rfc/rfc255.txt, ftp://ftp.internic.net/rfc/
rfc266.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc252.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc255.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc266.txt. Obsoleted by RFC0266 [261]. Obsoletes RFC0252 [252]. Status: UNKNOWN. Not online.

RFC0256

[256] B. Cosell. RFC 256: IMPSYS change notification, November 3,

1971. URL ftp://ftp.internic.net/rfc/rfc256.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc256.txt. Status: UNKNOWN.

RFC0257

[257] ???? RFC 257: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc257.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc257.txt.

RFC0263

[258] A. M. McKenzie. RFC 263: “Very Distant” host interface, December

17, 1971. URL ftp://ftp.internic.net/rfc/rfc263.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc263.txt. Status: UNKNOWN.

RFC0264

[259] A. Bhushan, B. Braden, W. Crowther, E. Harslem, J. Heafner, A. McKenzie, B. Sundberg, D. Watson, and J. White. RFC 264: The data transfer protocol, January 4, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc171.txt,   ftp://ftp.internic.net/
rfc/rfc264.txt, ftp://ftp.internic.net/rfc/rfc354.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc171.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc264.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc354.

txt. Obsoleted by RFC0354 [349]. Obsoletes RFC0171 [171]. Status: UNKNOWN. Not online.

RFC0265

[260] A. Bhushan, R. Braden, W. Crowther, E. Harslem, J. F. Heafner, A. M. McKenzie, J. T. Melvin, R. L. Sundberg, R. W. Watson, and J. E. White. RFC 265: The File Transfer Protocol, De-

cember 17, 1971. URL ftp://ftp.internic.net/rfc/rfc172.txt,
 ftp://ftp.internic.net/rfc/rfc265.txt, ftp://ftp.internic.
net/rfc/rfc294.txt,   ftp://ftp.internic.net/rfc/rfc354.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc172.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc265.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc354.txt.

Obsoleted by RFC0354 [349]. Obsoletes RFC0172 [172]. Updated by RFC0294 [289]. Status: UNKNOWN.

RFC0266

[261] E. Westheimer. RFC 266: Network host status, November 8,

1971. URL ftp://ftp.internic.net/rfc/rfc255.txt,   ftp://
ftp.internic.net/rfc/rfc266.txt, ftp://ftp.internic.net/rfc/
rfc267.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc255.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc266.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc267.txt. Obsoleted by RFC0267 [262]. Obsoletes RFC0255 [255]. Status: UNKNOWN.

RFC0267

[262] E. Westheimer. RFC 267: Network host status, November 22,

1971. URL ftp://ftp.internic.net/rfc/rfc266.txt,   ftp://
ftp.internic.net/rfc/rfc267.txt, ftp://ftp.internic.net/rfc/
rfc287.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc266.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc267.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc287.txt. Obsoleted by RFC0287 [282]. Obsoletes RFC0266 [261]. Status: UNKNOWN. Format: TXT=7862 bytes.

RFC0268

[263] J. Postel. RFC 268: Graphics facilities information, November

24, 1971. URL ftp://ftp.internic.net/rfc/rfc268.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc268.txt. Status: UNKNOWN.

RFC0269

[264] H. Brodie. RFC 269: Some experience with file transfer, De-

cember 6, 1971. URL ftp://ftp.internic.net/rfc/rfc122.txt,
 ftp://ftp.internic.net/rfc/rfc238.txt, ftp://ftp.internic.
net/rfc/rfc269.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc122.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc238.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc269.txt. Updates RFC0122, RFC0238

[122, 238]. Status: UNKNOWN. Format: TXT=5961 bytes.

RFC0270

[265] A. M. McKenzie. RFC 270: Correction to BBN report no. 1822 (NIC NO 7958), January 1, 1972. URL ftp://ftp.internic.net/

rfc/nic7959.txt,   ftp://ftp.internic.net/rfc/rfc270.txt,
ftp://ftp.math.utah.edu/pub/rfc/nic7959.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc270.txt. Updates NIC7959 [?]. Status: UN-KNOWN. Not online.

RFC0271

[266] B. Cosell. RFC 271: IMP system change notifications, January

3, 1972. URL ftp://ftp.internic.net/rfc/rfc271.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc271.txt. Status: UNKNOWN.

RFC0272

[267] ???? RFC 272: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc272.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc272.txt.

RFC0273

[268] R. W. Watson. RFC 273: More on standard host names, October

18, 1971. URL ftp://ftp.internic.net/rfc/rfc237.txt,   ftp://
ftp.internic.net/rfc/rfc273.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc237.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc273.

txt. Obsoletes RFC0237 [237]. Status: UNKNOWN.

RFC0274

[269] E. Forman. RFC 274: Establishing a local guide for network us-

age, November 1, 1971. URL ftp://ftp.internic.net/rfc/rfc274.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc274.txt.  Status:

UNKNOWN.

RFC0275

[270] ???? RFC 275: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc275.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc275.txt.

RFC0276

[271] R. W. Watson. RFC 276: NIC course, November 8, 1971. URL ftp://

ftp.internic.net/rfc/rfc276.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc276.txt. Status: UNKNOWN.

RFC0277

[272] ???? RFC 277: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc277.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc277.txt.

RFC0278

[273] A. K. Bhushan, R. T. Braden, E. Harslem, J. F. Heafner, A. M. McKenzie, J. T. Melvin, R. L. Sundberg, R. W. Watson, and J. E. White. RFC 278: Revision of the mail box protocol, November

17, 1971. URL ftp://ftp.internic.net/rfc/rfc221.txt,   ftp://
ftp.internic.net/rfc/rfc278.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc221.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc278.

txt. Obsoletes RFC0221 [221]. Status: UNKNOWN.

RFC0279

[274] ???? RFC 279: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc279.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc279.txt.

RFC0280

[275] R. W. Watson. RFC 280: A draft of host names, November

17, 1971. URL ftp://ftp.internic.net/rfc/rfc280.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc280.txt. Status: UNKNOWN.

RFC0281

[276] A. M. McKenzie. RFC 281: Suggested addition to file transfer proto-

col, December 8, 1971. URL ftp://ftp.internic.net/rfc/rfc281.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc281.txt.  Status:

UNKNOWN. Not online.

RFC0282

[277] M. A. Padlipsky. RFC 282: Graphics meeting report, December

8, 1971. URL ftp://ftp.internic.net/rfc/rfc282.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc282.txt. Status: UNKNOWN. Not

online.

RFC0283

[278] R. T. Braden. RFC 283: NETRJT: Remote job service protocol

for TIPS, December 20, 1971. URL ftp://ftp.internic.net/rfc/
rfc189.txt,   ftp://ftp.internic.net/rfc/rfc283.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc189.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc283.txt. Updates RFC0189 [189]. Status: UN-KNOWN.

RFC0284

[279] ???? RFC 284: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc284.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc284.txt.

RFC0285

[280] D. Huff. RFC 285: Network graphics, December 15, 1971. URL ftp://

ftp.internic.net/rfc/rfc285.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc285.txt. Status: UNKNOWN. Not online.

RFC0286

[281] E. Forman. RFC 286: Network library information system, December

21, 1971. URL ftp://ftp.internic.net/rfc/rfc286.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc286.txt. Status: UNKNOWN.

RFC0287

[282] E. Westheimer. RFC 287: Status of network hosts, Decem-

ber 22, 1971. URL ftp://ftp.internic.net/rfc/rfc267.txt,
 ftp://ftp.internic.net/rfc/rfc287.txt, ftp://ftp.internic.
net/rfc/rfc288.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc267.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc287.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc288.txt. Obsoleted by RFC0288

[283]. Obsoletes RFC0267 [262]. Status: UNKNOWN.

RFC0288

[283] E. Westheimer. RFC 288: Network host status, January 6,

1972. URL ftp://ftp.internic.net/rfc/rfc287.txt,   ftp://
ftp.internic.net/rfc/rfc288.txt, ftp://ftp.internic.net/rfc/
rfc293.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc287.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc288.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc293.txt. Obsoleted by RFC0293 [288]. Obsoletes RFC0287 [282]. Updated by RFC0293 [288]. Status: UNKNOWN. Not online.

RFC0289

[284] R. W. Watson. RFC 289: What we hope is an official list of host

names, December 21, 1971. URL ftp://ftp.internic.net/rfc/
rfc289.txt,   ftp://ftp.internic.net/rfc/rfc384.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc289.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc384.txt. Obsoleted by RFC0384 [379]. Status: UN-KNOWN.

RFC0290

[285] A. P. Mullery. RFC 290: Computer networks and data sharing: A bibliography, January 11, 1972. URL ftp://ftp.internic.net/rfc/

rfc243.txt,   ftp://ftp.internic.net/rfc/rfc290.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc243.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc290.txt. Obsoletes RFC0243 [243]. Status: UN-KNOWN. Not online.

RFC0291

[286] D. B. McKay. RFC 291: Data management meeting announcement,

January 14, 1972. URL ftp://ftp.internic.net/rfc/rfc291.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc291.txt.  Status: UN-

KNOWN.

RFC0292

[287] J. C. Michener, I. W. Cotton, K. C. Kelley, D. E. Liddle, and E. W. Meyer, Jr. RFC 292: Graphics protocol: Level 0 only, January

12, 1972. URL ftp://ftp.internic.net/rfc/rfc292.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc292.txt. Status: UNKNOWN. Not

online.

RFC0293

[288] E. Westheimer. RFC 293: Network host status, January 18,

1972. URL ftp://ftp.internic.net/rfc/rfc288.txt,   ftp://
ftp.internic.net/rfc/rfc293.txt, ftp://ftp.internic.net/rfc/
rfc298.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc288.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc293.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc298.txt. Obsoleted by RFC0298 [293]. Obsoletes RFC0288 [283]. Updates RFC0288 [283]. Status: UNKNOWN.

RFC0294

[289] A. K. Bhushan. RFC 294: The use of “set data type” transaction in file transfer protocol, January 25, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc265.txt,   ftp://ftp.internic.net/
rfc/rfc294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc265.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc294.txt. Updates RFC0265

[260]. Status: UNKNOWN.

RFC0295

[290] J. Postel. RFC 295: Report of the Protocol Workshop, 12 October

1971, January 2, 1972. URL ftp://ftp.internic.net/rfc/rfc295.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc295.txt.  Status:

UNKNOWN.

RFC0296

[291] D. E. Liddle. RFC 296: DS-1 display system, January 27,

1972. URL ftp://ftp.internic.net/rfc/rfc296.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc296.txt. Status: UNKNOWN. Not

online.

RFC0297

[292] D. C. Walden. RFC 297: TIP message buffers, January 31,

1972. URL ftp://ftp.internic.net/rfc/rfc297.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc297.txt. Status: UNKNOWN.

RFC0298

[293] E. Westheimer. RFC 298: Network host status, February 11,

1972. URL ftp://ftp.internic.net/rfc/rfc293.txt,   ftp://
ftp.internic.net/rfc/rfc298.txt, ftp://ftp.internic.net/rfc/
rfc306.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc293.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc298.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc306.txt. Obsoleted by RFC0306 [301]. Obsoletes RFC0293 [288]. Status: UNKNOWN. Not online.

RFC0299

[294] D. Hopkin. RFC 299: Information management system, February

11, 1972. URL ftp://ftp.internic.net/rfc/rfc299.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc299.txt. Status: UNKNOWN.

RFC0300

[295] J. B. North. RFC 300: ARPA network mailing lists, Jan-

uary 25, 1972. URL ftp://ftp.internic.net/rfc/rfc211.txt,
 ftp://ftp.internic.net/rfc/rfc300.txt, ftp://ftp.internic.
net/rfc/rfc303.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc211.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc300.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc303.txt. Obsoleted by RFC0303

[298]. Obsoletes RFC0211 [211]. Status: UNKNOWN. Not online.

RFC0301

[296] R. Alter. RFC 301: BBN IMP (#5) and NCC schedule March 4,

1971, February 11, 1972. URL ftp://ftp.internic.net/rfc/rfc301.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc301.txt.  Status:

UNKNOWN.

RFC0302

[297] R. F. Bryan. RFC 302: Exercising the ARPANET, February 8,

1972. URL ftp://ftp.internic.net/rfc/rfc302.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc302.txt. Status: UNKNOWN.

RFC0303

[298] Network Information Center. Stanford Research Institute. RFC 303: ARPA network mailing lists, February 23, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc300.txt,   ftp://ftp.internic.net/
rfc/rfc303.txt, ftp://ftp.internic.net/rfc/rfc329.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc300.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc303.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc329.

txt. Obsoleted by RFC0329 [324]. Obsoletes RFC0300 [295]. Status: UNKNOWN. Not online.

RFC0304

[299] D. B. McKay. RFC 304: Data management system proposal for the ARPA network, February 17, 1972. URL ftp://ftp.

internic.net/rfc/rfc304.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc304.txt. Status: UNKNOWN. Not online.

RFC0305

[300] R. Alter. RFC 305: Unknown host numbers, February 23, 1972. URL

ftp://ftp.internic.net/rfc/rfc305.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc305.txt. Status: UNKNOWN.

RFC0306

[301] E. Westheimer. RFC 306: Network host status, February 15,

1972. URL ftp://ftp.internic.net/rfc/rfc298.txt,   ftp://
ftp.internic.net/rfc/rfc306.txt, ftp://ftp.internic.net/rfc/
rfc315.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc298.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc306.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc315.txt. Obsoleted by RFC0315 [310]. Obsoletes RFC0298 [293]. Status: UNKNOWN. Not online.

RFC0307

[302] E. Harslem. RFC 307: Using network remote job entry, February

24, 1972. URL ftp://ftp.internic.net/rfc/rfc307.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc307.txt. Status: UNKNOWN. Not

online.

RFC0308

[303] M. Seriff. RFC 308: ARPANET host availability data, March

13, 1972. URL ftp://ftp.internic.net/rfc/rfc308.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc308.txt. Status: UNKNOWN.

RFC0309

[304] A. K. Bhushan. RFC 309: Data and file transfer workshop announce-

ment, March 17, 1972. URL ftp://ftp.internic.net/rfc/rfc309.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc309.txt.  Status:

UNKNOWN.

RFC0310

[305] A. K. Bhushan. RFC 310: Another look at data and file transfer pro-

tocols, April 3, 1972. URL ftp://ftp.internic.net/rfc/rfc310.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc310.txt.  Status:

UNKNOWN. Not online.

RFC0311

[306] R. F. Bryan. RFC 311: New console attachments to the USCB

host, February 29, 1972. URL ftp://ftp.internic.net/rfc/rfc311.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc311.txt.  Status:

UNKNOWN.

RFC0312

[307] A. M. McKenzie. RFC 312: Proposed change in IMP-to-host pro-

tocol, March 22, 1972. URL ftp://ftp.internic.net/rfc/rfc312.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc312.txt.  Status:

UNKNOWN.

RFC0313

[308] T. C. O’Sullivan. RFC 313: Computer based instruction, March

6, 1972. URL ftp://ftp.internic.net/rfc/rfc313.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc313.txt. Status: UNKNOWN. Not

online.

RFC0314

[309] I. W. Cotton. RFC 314: Network graphics working group meeting, March

14, 1972. URL ftp://ftp.internic.net/rfc/rfc314.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc314.txt. Status: UNKNOWN.

RFC0315

[310] E. Westheimer. RFC 315: Network host status, March 8, 1972. URL ftp:

//ftp.internic.net/rfc/rfc306.txt,   ftp://ftp.internic.net/
rfc/rfc315.txt, ftp://ftp.internic.net/rfc/rfc319.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc306.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc315.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc319.

txt. Obsoleted by RFC0319 [314]. Obsoletes RFC0306 [301]. Status: UNKNOWN. Not online.

RFC0316

[311] D. B. McKay and A. P. Mullery. RFC 316: ARPA network data management working group, February 23, 1972. URL ftp://ftp.

internic.net/rfc/rfc316.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc316.txt. Status: UNKNOWN. Not online.

RFC0317

[312] J. Postel. RFC 317: Official host-host protocol modification: Assigned link numbers, March 20, 1972. URL ftp://ftp.internic.net/rfc/

rfc317.txt,   ftp://ftp.internic.net/rfc/rfc604.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc317.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc604.txt. Obsoleted by RFC0604 [597]. Status: UN-KNOWN. Not online.

RFC0318

[313] J. Postel. RFC 318: Telnet protocols, April 3, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc139.txt,   ftp://ftp.internic.net/

rfc/rfc158.txt, ftp://ftp.internic.net/rfc/rfc318.txt,   ftp:
//ftp.internic.net/rfc/rfc435.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc139.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc158.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc318.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc435.txt. Updates RFC0158 [158].

Updated by RFC0435 [429]. Online: no. Status: UNKNOWN. See also RFC0139 [139], RFC0158 [158].

RFC0319

[314] E. Westheimer. RFC 319: Network host status, March 21,

1972. URL ftp://ftp.internic.net/rfc/rfc315.txt,   ftp://
ftp.internic.net/rfc/rfc319.txt, ftp://ftp.internic.net/rfc/
rfc326.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc315.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc319.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc326.txt. Obsoleted by RFC0326 [321]. Obsoletes RFC0315 [310]. Status: UNKNOWN. Not online.

RFC0320

[315] R. Reddy. RFC 320: Workshop on hard copy line printers, March

27, 1972. URL ftp://ftp.internic.net/rfc/rfc320.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc320.txt. Status: UNKNOWN. Not

online.

RFC0321

[316] P. M. Karp. RFC 321: CBI networking activity at MITRE, March

24, 1972. URL ftp://ftp.internic.net/rfc/rfc321.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc321.txt. Status: UNKNOWN.

RFC0322

[317] V. G. Cerf and J. Postel. RFC 322: Well known socket numbers, March

26, 1972. URL ftp://ftp.internic.net/rfc/rfc322.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc322.txt. Status: UNKNOWN.

RFC0323

[318] V. G. Cerf. RFC 323: Formation of network measurement group

(NMG), March 23, 1972. URL ftp://ftp.internic.net/rfc/
rfc323.txt,   ftp://ftp.internic.net/rfc/rfc388.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc323.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc388.txt. Updated by RFC0388 [383]. Status: UN-KNOWN. Not online.

RFC0324

[319] J. Postel. RFC 324: RJE protocol meeting, April 3, 1972. URL ftp://

ftp.internic.net/rfc/rfc324.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc324.txt. Status: UNKNOWN. Format: TXT=1176 bytes.

RFC0325

[320] G. Hicks. RFC 325: Network remote job entry program — NETRJS,

April 6, 1972. URL ftp://ftp.internic.net/rfc/rfc325.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc325.txt.  Status: UN-

KNOWN. Not online.

RFC0326

[321] E. Westheimer. RFC 326: Network host status, April 3, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc319.txt,   ftp://ftp.internic.net/
rfc/rfc326.txt, ftp://ftp.internic.net/rfc/rfc330.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc319.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc326.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc330.

txt. Obsoleted by RFC0330 [325]. Obsoletes RFC0319 [314]. Updated by RFC0330 [325]. Status: UNKNOWN.

RFC0327

[322] A. K. Bhushan. RFC 327: Data and file transfer workshop notes, April

27, 1972. URL ftp://ftp.internic.net/rfc/rfc327.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc327.txt. Status: UNKNOWN. Not

online.

RFC0328

[323] J. Postel. RFC 328: Suggested Telnet Protocol changes, April

29, 1972. URL ftp://ftp.internic.net/rfc/rfc328.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc328.txt. Status: UNKNOWN.

RFC0329

[324] Network Information Center. Stanford Research Institute. RFC 329: ARPA network mailing lists, May 17, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc303.txt,   ftp://ftp.internic.net/
rfc/rfc329.txt, ftp://ftp.internic.net/rfc/rfc363.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc303.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc329.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc363.

txt. Obsoleted by RFC0363 [358]. Obsoletes RFC0303 [298]. Status: UNKNOWN. Not online.

RFC0330

[325] E. Westheimer. RFC 330: Network host status, April 13, 1972. URL ftp:

//ftp.internic.net/rfc/rfc326.txt,   ftp://ftp.internic.net/
rfc/rfc330.txt, ftp://ftp.internic.net/rfc/rfc332.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc326.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc330.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc332.

txt. Obsoleted by RFC0332 [327]. Obsoletes RFC0326 [321]. Updates RFC0326 [321]. Updated by RFC0332 [327]. Status: UNKNOWN.

RFC0331

[326] J. M. McQuillan. RFC 331: IMP system change notification, April

19, 1972. URL ftp://ftp.internic.net/rfc/rfc331.txt,   ftp://
ftp.internic.net/rfc/rfc343.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc331.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc343.

txt. Obsoleted by RFC0343 [338]. Status: UNKNOWN.

RFC0332

[327] E. Westheimer. RFC 332: Network host status, April 25, 1972. URL ftp:

//ftp.internic.net/rfc/rfc330.txt,   ftp://ftp.internic.net/
rfc/rfc332.txt, ftp://ftp.internic.net/rfc/rfc342.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc330.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc332.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc342.

txt. Obsoleted by RFC0342 [337]. Obsoletes RFC0330 [325]. Updates RFC0330 [325]. Updated by RFC0342 [337]. Status: UNKNOWN.

RFC0333

[328] R. D. Bressler, D. Murphy, and D. C. Walden. RFC 333: Proposed experiment with a message switching protocol, May 15, 1972. URL ftp://

ftp.internic.net/rfc/rfc333.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc333.txt. Status: UNKNOWN. Not online.

RFC0334

[329] A. M. McKenzie. RFC 334: Network use on may 8, May 1,

1972. URL ftp://ftp.internic.net/rfc/rfc334.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc334.txt. Status: UNKNOWN.

RFC0335

[330] R. F. Bryan. RFC 335: New interface — IMP/360, May 1,

1972. URL ftp://ftp.internic.net/rfc/rfc335.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc335.txt. Status: UNKNOWN.

RFC0336

[331] I. W. Cotton. RFC 336: Level 0 graphic input protocol, May

5, 1972. URL ftp://ftp.internic.net/rfc/rfc336.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc336.txt. Status: UNKNOWN.

RFC0337

[332] ???? RFC 337: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc337.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc337.txt.

RFC0338

[333] R. T. Braden. RFC 338: EBCDIC/ASCII mapping for network

RJE, May 17, 1972. URL ftp://ftp.internic.net/rfc/rfc338.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc338.txt.  Status:

UNKNOWN. Not online.

RFC0339

[334] R. Thomas. RFC 339: MLTNET: A “Multi Telnet” subsystem for

tenex, May 5, 1972. URL ftp://ftp.internic.net/rfc/rfc339.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc339.txt.  Status:

UNKNOWN. Format: TXT=7941 bytes.

RFC0340

[335] T. C. O’Sullivan. RFC 340: Proposed Telnet changes, May 15,

1972. URL ftp://ftp.internic.net/rfc/rfc328.txt,   ftp://
ftp.internic.net/rfc/rfc340.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc328.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc340.

txt. See also RFC0328 [323]. Status: UNKNOWN.

RFC0341

[336] ???? RFC 341: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc341.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc341.txt.

RFC0342

[337] E. Westheimer. RFC 342: Network host status, May 15, 1972. URL ftp:

//ftp.internic.net/rfc/rfc332.txt,   ftp://ftp.internic.net/
rfc/rfc342.txt, ftp://ftp.internic.net/rfc/rfc344.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc332.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc342.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc344.

txt. Obsoleted by RFC0344 [339]. Obsoletes RFC0332 [327]. Updates RFC0332 [327]. Updated by RFC0344 [339]. Status: UNKNOWN.

RFC0343

[338] A. M. McKenzie. RFC 343: IMP system change notification,

May 19, 1972. URL ftp://ftp.internic.net/rfc/rfc331.txt,
 ftp://ftp.internic.net/rfc/rfc343.txt, ftp://ftp.internic.
net/rfc/rfc359.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc331.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc343.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc359.txt. Obsoleted by RFC0359

[354]. Obsoletes RFC0331 [326]. Status: UNKNOWN. Not online.

RFC0344

[339] E. Westheimer. RFC 344: Network host status, May 22, 1972. URL ftp:

//ftp.internic.net/rfc/rfc342.txt,   ftp://ftp.internic.net/
rfc/rfc344.txt, ftp://ftp.internic.net/rfc/rfc353.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc342.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc344.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc353.

txt. Obsoleted by RFC0353 [348]. Obsoletes RFC0342 [337]. Updates RFC0342 [337]. Status: UNKNOWN.

RFC0345

[340] K. C. Kelley. RFC 345: Interest in mixed integer programming (MPSX on NIC 360/91 at CCN), May 26, 1972. URL ftp://ftp.

internic.net/rfc/rfc345.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc345.txt. Status: UNKNOWN.

RFC0346

[341] J. Postel. RFC 346: Satellite considerations, May 30, 1972. URL ftp://

ftp.internic.net/rfc/rfc346.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc346.txt. Status: UNKNOWN.

RFC0347

[342] J. Postel. RFC 347: Echo process, May 30, 1972. URL ftp://ftp.

internic.net/rfc/rfc347.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc347.txt. Status: UNKNOWN.

RFC0348

[343] J. Postel. RFC 348: Discard process, May 30, 1972. URL ftp://

ftp.internic.net/rfc/rfc348.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc348.txt. Status: UNKNOWN.

RFC0349

[344] J. Postel. RFC 349: Proposed standard socket numbers, May 30,

1972. URL ftp://ftp.internic.net/rfc/rfc204.txt,   ftp://

ftp.internic.net/rfc/rfc322.txt, ftp://ftp.internic.net/rfc/
rfc349.txt,   ftp://ftp.internic.net/rfc/rfc433.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc204.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc322.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc349.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc433.txt. See also

RFC0322, RFC0204 [317, 204]. Obsoleted by RFC0433 [427]. Status: UNKNOWN. Format: TXT=1663 bytes.

RFC0350

[345] R. Stoughton. RFC 350: User accounts for UCSB on-line system, May

18, 1972. URL ftp://ftp.internic.net/rfc/rfc350.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc350.txt. Status: UNKNOWN.

RFC0351

[346] D. Crocker. RFC 351: Graphics information form for the ARPANET graphics resources notebook, June 5, 1972. URL ftp://ftp.

internic.net/rfc/rfc351.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc351.txt. Status: UNKNOWN. Not online.

RFC0352

[347] D. Crocker. RFC 352: TIP site information form, June 5,

1972. URL ftp://ftp.internic.net/rfc/rfc352.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc352.txt. Status: UNKNOWN.

RFC0353

[348] E. Westheimer. RFC 353: Network host status, June 12, 1972. URL ftp:

//ftp.internic.net/rfc/rfc344.txt,   ftp://ftp.internic.net/
rfc/rfc353.txt, ftp://ftp.internic.net/rfc/rfc362.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc344.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc353.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc362.

txt. Obsoleted by RFC0362 [357]. Obsoletes RFC0344 [339]. Status: UNKNOWN. Not online.

RFC0354

[349] A. K. Bhushan. RFC 354: File transfer protocol, July 8, 1972. URL ftp:

//ftp.internic.net/rfc/rfc264.txt,   ftp://ftp.internic.net/
rfc/rfc265.txt, ftp://ftp.internic.net/rfc/rfc354.txt,   ftp:
//ftp.internic.net/rfc/rfc385.txt, ftp://ftp.internic.net/rfc/
rfc542.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc264.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc265.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc354.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc385.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc542.txt.

Obsoleted by RFC0542 [535]. Obsoletes RFC0264, RFC0265 [259, 260].

Updated by RFC0385 [380]. Status: UNKNOWN. Format: TXT=58074 bytes.

RFC0355

[350] J. Davidson. RFC 355: Response to NWG/RFC 346, June 9,

1972. URL ftp://ftp.internic.net/rfc/rfc355.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc355.txt. Status: UNKNOWN.

RFC0356

[351] R. Alter. RFC 356: ARPA network control center, June 21,

1972. URL ftp://ftp.internic.net/rfc/rfc356.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc356.txt. Status: UNKNOWN.

RFC0357

[352] J. Davidson. RFC 357: Echoing strategy for satellite links, June

26, 1972. URL ftp://ftp.internic.net/rfc/rfc357.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc357.txt. Status: UNKNOWN. Not

online.

RFC0358

[353] ???? RFC 358: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc358.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc358.txt.

RFC0359

[354] D. C. Walden. RFC 359: Status of the release of the new IMP system (2600), June 22, 1972. URL ftp://ftp.internic.net/rfc/

rfc343.txt,   ftp://ftp.internic.net/rfc/rfc359.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc343.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc359.txt. Obsoletes RFC0343 [338]. Status: UN-KNOWN.

RFC0360

[355] C. Holland. RFC 360: Proposed remote job entry protocol, June

24, 1972. URL ftp://ftp.internic.net/rfc/rfc360.txt,   ftp://
ftp.internic.net/rfc/rfc407.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc360.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc407.

txt. Obsoleted by RFC0407 [402]. Status: UNKNOWN. Not online.

RFC0361

[356] R. D. Bressler. RFC 361: Deamon processes on host 106, July

5, 1972. URL ftp://ftp.internic.net/rfc/rfc361.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc361.txt. Status: UNKNOWN.

RFC0362

[357] E. Westheimer. RFC 362: Network host status, June 28, 1972. URL ftp:

//ftp.internic.net/rfc/rfc353.txt,   ftp://ftp.internic.net/
rfc/rfc362.txt, ftp://ftp.internic.net/rfc/rfc366.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc353.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc362.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc366.

txt. Obsoleted by RFC0366 [361]. Obsoletes RFC0353 [348]. Status: UNKNOWN.

RFC0363

[358] Network Information Center. Stanford Research Institute. RFC 363: ARPA network mailing lists, August 8, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc329.txt,   ftp://ftp.internic.net/
rfc/rfc363.txt, ftp://ftp.internic.net/rfc/rfc402.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc329.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc363.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc402.

txt. Obsoleted by RFC0402 [397]. Obsoletes RFC0329 [324]. Status: UNKNOWN. Not online.

RFC0364

[359] M. D. Abrams. RFC 364: Serving remote users on the ARPANET, July

11, 1972. URL ftp://ftp.internic.net/rfc/rfc364.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc364.txt. Status: UNKNOWN. Not

online.

RFC0365

[360] D. C. Walden. RFC 365: Letter to all TIP users, July 11,

1972. URL ftp://ftp.internic.net/rfc/rfc365.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc365.txt. Status: UNKNOWN.

RFC0366

[361] E. Westheimer. RFC 366: Network host status, July 11, 1972. URL ftp:

//ftp.internic.net/rfc/rfc362.txt,   ftp://ftp.internic.net/
rfc/rfc366.txt, ftp://ftp.internic.net/rfc/rfc367.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc362.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc366.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc367.

txt. Obsoleted by RFC0367 [362]. Obsoletes RFC0362 [357]. Status: UNKNOWN.

RFC0367

[362] E. Westheimer. RFC 367: Network host status, July 19, 1972. URL ftp:

//ftp.internic.net/rfc/rfc366.txt,   ftp://ftp.internic.net/
rfc/rfc367.txt, ftp://ftp.internic.net/rfc/rfc370.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc366.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc367.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc370.

txt. Obsoleted by RFC0370 [365]. Obsoletes RFC0366 [361]. Updated by RFC0370 [365]. Status: UNKNOWN. Not online.

RFC0368

[363] R. T. Braden. RFC 368: Comments on “proposed remote job entry

protocol”, July 21, 1972. URL ftp://ftp.internic.net/rfc/rfc368.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc368.txt.  Status:

UNKNOWN.

RFC0369

[364] J. R. Pickens. RFC 369: Evaluation of ARPANET services january–

march, 1972, July 25, 1972. URL ftp://ftp.internic.net/rfc/
rfc369.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc369.txt. Sta-

tus: UNKNOWN. Not online.

RFC0370

[365] E. Westheimer. RFC 370: Network host status, July 31, 1972. URL ftp:

//ftp.internic.net/rfc/rfc367.txt,   ftp://ftp.internic.net/
rfc/rfc370.txt, ftp://ftp.internic.net/rfc/rfc376.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc367.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc370.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc376.

txt. Obsoleted by RFC0376 [371]. Obsoletes RFC0367 [362]. Updates RFC0367 [362]. Updated by RFC0376 [371]. Status: UNKNOWN.

RFC0371

[366] R. E. Kahn. RFC 371: Demonstration at International Computer Communications Conference, July 12, 1972. URL ftp://ftp.

internic.net/rfc/rfc371.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc371.txt. Status: UNKNOWN.

RFC0372

[367] R. W. Watson. RFC 372: Notes on a conversation with Bob Kahn on

the ICCC, July 12, 1972. URL ftp://ftp.internic.net/rfc/rfc372.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc372.txt.  Status:

UNKNOWN.

RFC0373

[368] John McCarthy. RFC 373: Arbitrary character sets, July 14,

1972. URL ftp://ftp.internic.net/rfc/rfc373.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc373.txt. Status: UNKNOWN.

RFC0374

[369] A. M. McKenzie. RFC 374: IMP system announcement, July

19, 1972. URL ftp://ftp.internic.net/rfc/rfc374.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc374.txt. Status: UNKNOWN.

RFC0375

[370] ???? RFC 375: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc375.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc375.txt.

RFC0376

[371] E. Westheimer. RFC 376: Network host status, August 8, 1972. URL

ftp://ftp.internic.net/rfc/rfc370.txt,   ftp://ftp.internic.
net/rfc/rfc376.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc370.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc376.txt. Obsoletes

RFC0370 [365]. Updates RFC0370 [365]. Status: UNKNOWN.

RFC0377

[372] R. T. Braden. RFC 377: Using TSO via ARPA network virtual termi-

nal, August 10, 1972. URL ftp://ftp.internic.net/rfc/rfc377.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc377.txt.  Status:

UNKNOWN. Not online.

RFC0378

[373] A. M. McKenzie. RFC 378: Traffic statistics (july 1972), August

10, 1972. URL ftp://ftp.internic.net/rfc/rfc378.txt,   ftp://
ftp.internic.net/rfc/rfc391.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc378.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc391.

txt. Obsoleted by RFC0391 [386]. Status: UNKNOWN. Not online.

RFC0379

[374] R. T. Braden. RFC 379: Using TSO at CCN, August 11,

1972. URL ftp://ftp.internic.net/rfc/rfc379.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc379.txt. Status: UNKNOWN. For-

mat: TXT=8674 bytes.

RFC0380

[375] ???? RFC 380: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc380.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc380.txt.

RFC0381

[376] J. M. McQuillan. RFC 381: Three aids to improved network

operation, July 26, 1972. URL ftp://ftp.internic.net/rfc/

rfc381.txt,   ftp://ftp.internic.net/rfc/rfc394.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc381.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc394.txt. Updated by RFC0394 [389]. Status: UN-KNOWN. Not online.

RFC0382

[377] L. McDaniel. RFC 382: Mathematical software on the ARPA net-

work, August 3, 1972. URL ftp://ftp.internic.net/rfc/rfc382.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc382.txt.  Status:

UNKNOWN.

RFC0383

[378] ???? RFC 383: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc383.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc383.txt.

RFC0384

[379] J. B. North. RFC 384: Official site idents for organizations in the ARPA network, August 28, 1972. URL ftp://ftp.internic.net/rfc/

rfc289.txt,   ftp://ftp.internic.net/rfc/rfc384.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc289.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc384.txt. Obsoletes RFC0289 [284]. Status: UN-KNOWN. Not online.

RFC0385

[380] A. K. Bhushan. RFC 385: Comments on the file transfer protocol,

August 18, 1972. URL ftp://ftp.internic.net/rfc/rfc354.txt,
 ftp://ftp.internic.net/rfc/rfc385.txt, ftp://ftp.internic.
net/rfc/rfc414.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc354.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc385.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc414.txt. Updates RFC0354 [349].

Updated by RFC0414 [409]. Status: UNKNOWN.

RFC0386

[381] B. Cosell and D. C. Walden. RFC 386: Letter to TIP users-2, August

16, 1972. URL ftp://ftp.internic.net/rfc/rfc386.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc386.txt. Status: UNKNOWN.

RFC0387

[382] K. C. Kelley and J. Meir. RFC 387: Some experiences in implementing network graphics protocol level 0, August 10, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc387.txt,   ftp://ftp.internic.net/
rfc/rfc401.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc387.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc401.txt. Updated by RFC0401

[396]. Status: UNKNOWN. Not online.

RFC0388

[383] V. G. Cerf. RFC 388: NCP statistics, August 23, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc323.txt,   ftp://ftp.internic.net/
rfc/rfc388.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc323.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc388.txt. Updates RFC0323

[318]. Status: UNKNOWN. Not online.

RFC0389

[384] B. Noble. RFC 389: UCLA campus computing network liaison staff for ARPA network, August 30, 1972. URL ftp://ftp.internic.net/rfc/

rfc389.txt,   ftp://ftp.internic.net/rfc/rfc423.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc389.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc423.txt. Obsoleted by RFC0423 [418]. Status: UN-KNOWN.

RFC0390

[385] R. T. Braden. RFC 390: TSO scenario, September 12, 1972. URL ftp://

ftp.internic.net/rfc/rfc390.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc390.txt. Status: UNKNOWN. Format: TXT=6103 bytes.

RFC0391

[386] A. M. McKenzie. RFC 391: Traffic statistics (august 1972), September

15, 1972. URL ftp://ftp.internic.net/rfc/rfc378.txt,   ftp://
ftp.internic.net/rfc/rfc391.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc378.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc391.

txt. Obsoletes RFC0378 [373]. Status: UNKNOWN. Not online.

RFC0392

[387] G. Hicks and B. D. Wessler. RFC 392: Measurement of host costs for transmitting network data, September 20, 1972. URL ftp://ftp.

internic.net/rfc/rfc392.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc392.txt. Status: UNKNOWN. Not online.

RFC0393

[388] J. M. Winett. RFC 393: Comments on Telnet Protocol changes,

October 3, 1972. URL ftp://ftp.internic.net/rfc/rfc109.txt,
 ftp://ftp.internic.net/rfc/rfc139.txt, ftp://ftp.internic.
net/rfc/rfc158.txt,   ftp://ftp.internic.net/rfc/rfc318.txt,
ftp://ftp.internic.net/rfc/rfc328.txt,   ftp://ftp.internic.
net/rfc/rfc393.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc109.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc139.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc158.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc318.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc328.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc393.txt. See also

RFC0109, RFC0139, RFC0158, RFC0318, RFC0328 [109, 139, 158, 313, 323]. Status: UNKNOWN.

RFC0394

[389] J. M. McQuillan. RFC 394: Two proposed changes to the IMP-host protocol, September 27, 1972. URL ftp://ftp.internic.net/rfc/

rfc381.txt,   ftp://ftp.internic.net/rfc/rfc394.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc381.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc394.txt. Updates RFC381, RFC0381 [376, 376]. Status: UNKNOWN.

RFC0395

[390] J. M. McQuillan. RFC 395: Switch settings on IMPs and TIPs, October

3, 1972. URL ftp://ftp.internic.net/rfc/rfc395.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc395.txt. Status: UNKNOWN.

RFC0396

[391] S. Bunch. RFC 396: Network graphics working group meeting — second iteration, November 13, 1972. URL ftp://ftp.internic.net/rfc/

rfc396.txt,   ftp://ftp.internic.net/rfc/rfc474.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc396.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc474.txt. Updated by RFC0474 [468]. Status: UN-KNOWN.

RFC0397

[392] ???? RFC 397: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc397.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc397.txt.

RFC0398

[393] J. R. Pickens and E. Faeh. RFC 398: ICP sockets, September

22, 1972. URL ftp://ftp.internic.net/rfc/rfc398.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc398.txt. Status: UNKNOWN.

RFC0399

[394] M. Krilanovich. RFC 399: SMFS login and logout, Septem-

ber 26, 1972. URL ftp://ftp.internic.net/rfc/rfc122.txt,
 ftp://ftp.internic.net/rfc/rfc399.txt, ftp://ftp.internic.
net/rfc/rfc431.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc122.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc399.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc431.txt. Obsoleted by RFC0431

[425]. Updates RFC0122 [122]. Status: UNKNOWN.

RFC0400

[395] A. M. McKenzie. RFC 400: Traffic statistics (september 1972), October

18, 1972. URL ftp://ftp.internic.net/rfc/rfc400.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc400.txt. Status: UNKNOWN. Not

online.

RFC0401

[396] J. Hansen. RFC 401: Conversion of NGP-0 coordinates to device specific coordinates, October 23, 1972. URL ftp://ftp.internic.net/rfc/

rfc387.txt,   ftp://ftp.internic.net/rfc/rfc401.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc387.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc401.txt. Updates RFC387, RFC0387 [382, 382]. Status: UNKNOWN.

RFC0402

[397] J. B. North. RFC 402: ARPA network mailing lists, October 26,

1972. URL ftp://ftp.internic.net/rfc/rfc363.txt,   ftp://
ftp.internic.net/rfc/rfc402.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc363.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc402.

txt. Obsoletes RFC0363 [358]. Status: UNKNOWN. Not online.

RFC0403

[398] G. Hicks. RFC 403: Desirability of a network 1108 service, January

10, 1973. URL ftp://ftp.internic.net/rfc/rfc403.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc403.txt. Status: UNKNOWN. Not

online.

RFC0404

[399] A. M. McKenzie. RFC 404: Host address changes involving rand

and ISI, October 5, 1972. URL ftp://ftp.internic.net/rfc/
rfc404.txt,   ftp://ftp.internic.net/rfc/rfc405.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc404.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc405.txt. Updated by RFC0405 [400]. Status: UN-KNOWN.

RFC0405

[400] A. M. McKenzie. RFC 405: Correction to RFC 404, October 10,

1972. URL ftp://ftp.internic.net/rfc/rfc404.txt,   ftp://
ftp.internic.net/rfc/rfc405.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc404.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc405.

txt. Updates RFC0404 [399]. Status: UNKNOWN.

RFC0406

[401] J. M. McQuillan. RFC 406: Scheduled IMP software releases, October

10, 1972. URL ftp://ftp.internic.net/rfc/rfc406.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc406.txt. Status: UNKNOWN.

RFC0407

[402] R. D. Bressler, R. Guida, and A. M. McKenzie. RFC 407: Remote job entry protocol, October 16, 1972. URL ftp://ftp.internic.net/rfc/

rfc360.txt,   ftp://ftp.internic.net/rfc/rfc407.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc360.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc407.txt. Obsoletes RFC0360 [355]. Status: HIS-TORIC. Format: TXT=47556 bytes.

RFC0408

[403] A. D. Owen and J. Postel. RFC 408: NETBANK, October 25,

1972. URL ftp://ftp.internic.net/rfc/rfc408.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc408.txt. Status: UNKNOWN.

RFC0409

[404] J. E. White. RFC 409: Tenex interface to UCSB’s simple-minded file sys-

tem, December 8, 1972. URL ftp://ftp.internic.net/rfc/rfc409.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc409.txt.  Status:

UNKNOWN. Not online.

RFC0410

[405] J. M. McQuillan. RFC 410: Removal of the 30-second delay when hosts come up, November 10, 1972. URL ftp://ftp.

internic.net/rfc/rfc410.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc410.txt. Status: UNKNOWN.

RFC0411

[406] M. A. Padlipsky. RFC 411: New MULTICS network software features,

November 14, 1972. URL ftp://ftp.internic.net/rfc/rfc411.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc411.txt.  Status:

UNKNOWN.

RFC0412

[407] G. Hicks. RFC 412: User FTP documentation, November 27,

1972. URL ftp://ftp.internic.net/rfc/rfc412.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc412.txt. Status: UNKNOWN.

RFC0413

[408] A. M. McKenzie. RFC 413: Traffic statistics (october 1972), November

13, 1972. URL ftp://ftp.internic.net/rfc/rfc413.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc413.txt. Status: UNKNOWN. Not

online.

RFC0414

[409] A. K. Bhushan. RFC 414: File transfer protocol (FTP) status and further comments, December 29, 1972. URL ftp://ftp.internic.net/rfc/

rfc385.txt,   ftp://ftp.internic.net/rfc/rfc414.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc385.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc414.txt. Updates RFC0385 [380]. Status: UN-KNOWN. Not online.

RFC0415

[410] H. Murray. RFC 415: Tenex bandwidth, November 29, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc415.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc415.txt. Status: UNKNOWN. Not online.

RFC0416

[411] J. C. Norton. RFC 416: ARC system will be unavailable for use during Thanksgiving week, November 7, 1972. URL ftp://ftp.

internic.net/rfc/rfc416.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc416.txt. Status: UNKNOWN.

RFC0417

[412] J. Postel and C. Kline. RFC 417: Link usage violation, December

6, 1972. URL ftp://ftp.internic.net/rfc/rfc417.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc417.txt. Status: UNKNOWN. Not

online.

RFC0418

[413] W. Hathaway. RFC 418: Server file transfer under TSS/360 at NASA Ames, November 27, 1972. URL ftp://ftp.internic.net/rfc/

rfc418.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc418.txt. Sta-

tus: UNKNOWN. Not online.

RFC0419

[414] A. Vezza. RFC 419: To: Network liaisons and station agents, December

12, 1972. URL ftp://ftp.internic.net/rfc/rfc419.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc419.txt. Status: UNKNOWN. Not

online.

RFC0420

[415] H. Murray. RFC 420: CCA ICCC weather demo, January 4,

1973. URL ftp://ftp.internic.net/rfc/rfc420.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc420.txt. Status: UNKNOWN. Not

online.

RFC0421

[416] A. M. McKenzie. RFC 421: Software consulting service for network users,

November 27, 1972. URL ftp://ftp.internic.net/rfc/rfc421.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc421.txt.  Status:

UNKNOWN.

RFC0422

[417] A. M. McKenzie. RFC 422: Traffic statistics (november 1972), December

11, 1972. URL ftp://ftp.internic.net/rfc/rfc422.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc422.txt. Status: UNKNOWN. Not

online.

RFC0423

[418] B. Noble. RFC 423: UCLA campus computing network liaison staff for

ARPANET, December 12, 1972. URL ftp://ftp.internic.net/rfc/
rfc389.txt,   ftp://ftp.internic.net/rfc/rfc423.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc389.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc423.txt. Obsoletes RFC0389 [384]. Status: UN-KNOWN.

RFC0424

[419] ???? RFC 424: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc424.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc424.txt.

RFC0425

[420] R. D. Bressler. RFC 425: “but my NCP costs $500 a day”, December

19, 1972. URL ftp://ftp.internic.net/rfc/rfc425.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc425.txt. Status: UNKNOWN.

RFC0426

[421] R. Thomas. RFC 426: Reconnection protocol, January 26, 1973. URL

ftp://ftp.internic.net/rfc/rfc426.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc426.txt. Status: UNKNOWN. Not online.

RFC0427

RFC0429

[423] J. Postel. RFC 429: Character generator process, December 12,

1972. URL ftp://ftp.internic.net/rfc/rfc429.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc429.txt. Status: UNKNOWN.

RFC0430

[424] R. T. Braden. RFC 430: Comments on file transfer protocol, February

7, 1973. URL ftp://ftp.internic.net/rfc/rfc430.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc430.txt. Status: UNKNOWN. Not

online.

RFC0431

[425] M. Krilanovich. RFC 431: Update on SMFS login and lo-

gout, December 15, 1972. URL ftp://ftp.internic.net/rfc/
rfc122.txt,   ftp://ftp.internic.net/rfc/rfc399.txt, ftp://
ftp.internic.net/rfc/rfc431.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc122.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc399.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc431.txt. Obsoletes

RFC0399 [394]. Updates RFC0122 [122]. Status: UNKNOWN.

RFC0432

[426] N. Neigus. RFC 432: Network logical map, December 29, 1972. URL

ftp://ftp.internic.net/rfc/rfc432.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc432.txt. Status: UNKNOWN. Not online.

RFC0433

[427] J. Postel. RFC 433: Socket number list, December 22, 1972. URL ftp:/

/ftp.internic.net/rfc/rfc349.txt,   ftp://ftp.internic.net/
rfc/rfc433.txt, ftp://ftp.internic.net/rfc/rfc503.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc349.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc433.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc503.

txt. Obsoleted by RFC0503 [497]. Obsoletes RFC0349 [344]. Status: UNKNOWN. Not online.

RFC0434

[428] A. M. McKenzie. RFC 434: IMP/TIP memory retrofit schedule, January

4, 1973. URL ftp://ftp.internic.net/rfc/rfc434.txt,   ftp://
ftp.internic.net/rfc/rfc447.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc434.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc447.

txt. Obsoleted by RFC0447 [441]. Status: UNKNOWN. Not online.

RFC0435

[429] B. Cosell and D. C. Walden. RFC 435: Telnet issues, January 5,

1973. URL ftp://ftp.internic.net/rfc/rfc318.txt,   ftp://
ftp.internic.net/rfc/rfc435.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc318.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc435.

txt. Updates RFC0318 [313]. Status: UNKNOWN. Not online.

RFC0436

[430] M. Krilanovich. RFC 436: Announcement of RJS at UCSB, January

10, 1973. URL ftp://ftp.internic.net/rfc/rfc436.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc436.txt. Status: UNKNOWN. Not

online.

RFC0437

[431] E. Faeh. RFC 437: Data reconfiguration service at UCSB, June

30, 1973. URL ftp://ftp.internic.net/rfc/rfc437.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc437.txt. Status: UNKNOWN. Not

online.

RFC0438

[432] R. Thomas and R. Clements. RFC 438: FTP server-server interac-

tion, January 15, 1973. URL ftp://ftp.internic.net/rfc/rfc438.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc438.txt.  Status:

UNKNOWN. Not online.

RFC0439

[433] V. G. Cerf. RFC 439: PARRY encounters the DOCTOR, January

21, 1973. URL ftp://ftp.internic.net/rfc/rfc439.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc439.txt. Status: UNKNOWN. Not

online.

RFC0440

[434] D. C. Walden. RFC 440: Scheduled network software maintenance,

January 1, 1973. URL ftp://ftp.internic.net/rfc/rfc440.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc440.txt.  Status: UN-

KNOWN. Not online.

RFC0441

[435] R. D. Bressler and R. Thomas. RFC 441: Inter-entity communication — an experiment, January 19, 1973. URL ftp://ftp.

internic.net/rfc/rfc441.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc441.txt. Status: UNKNOWN. Not online.

RFC0442

[436] V. G. Cerf. RFC 442: Current flow-control scheme for IMPSYS, January

24, 1973. URL ftp://ftp.internic.net/rfc/rfc442.txt,   ftp://
ftp.internic.net/rfc/rfc449.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc442.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc449.

txt. Updated by RFC0449 [443]. Status: UNKNOWN. Not online.

RFC0443

[437] A. M. McKenzie. RFC 443: Traffic statistics (december 1972), January

18, 1973. URL ftp://ftp.internic.net/rfc/rfc443.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc443.txt. Status: UNKNOWN. Not

online.

RFC0444

[438] ???? RFC 444: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc444.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc444.txt.

RFC0445

[439] A. M. McKenzie. RFC 445: IMP/TIP preventive maintenance sched-

ule, January 22, 1973. URL ftp://ftp.internic.net/rfc/rfc445.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc445.txt.  Status:

UNKNOWN. Not online.

RFC0446

[440] L. P. Deutsch. RFC 446: Proposal to consider a network program resource notebook, January 25, 1973. URL ftp://ftp.

internic.net/rfc/rfc446.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc446.txt. Status: UNKNOWN. Not online.

RFC0447

[441] A. M. McKenzie. RFC 447: IMP/TIP memory retrofit schedule,

January 29, 1973. URL ftp://ftp.internic.net/rfc/rfc434.txt,
 ftp://ftp.internic.net/rfc/rfc447.txt, ftp://ftp.internic.
net/rfc/rfc476.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc434.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc447.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc476.txt. Obsoleted by RFC0476

[470]. Obsoletes RFC0434 [428]. Status: UNKNOWN. Not online.

RFC0448

[442] R. T. Braden. RFC 448: Print files in FTP, February 27,

1973. URL ftp://ftp.internic.net/rfc/rfc448.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc448.txt. Status: UNKNOWN. Not

online.

RFC0449

[443] D. C. Walden. RFC 449: Current flow-control scheme for IMPSYS,

January 6, 1973. URL ftp://ftp.internic.net/rfc/rfc442.txt,
ftp://ftp.internic.net/rfc/rfc449.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc442.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc449.

txt. Updates RFC0442 [436]. Status: UNKNOWN. Not online.

RFC0450

[444] M. A. Padlipsky. RFC 450: MULTICS sampling timeout change, Febru-

ary 8, 1973. URL ftp://ftp.internic.net/rfc/rfc450.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc450.txt. Status: UNKNOWN.

Not online.

RFC0451

[445] M. A. Padlipsky. RFC 451: Tentative proposal for a unified user level pro-

tocol, February 22, 1973. URL ftp://ftp.internic.net/rfc/rfc451.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc451.txt.  Status:

UNKNOWN. Not online.

RFC0452

[446] ???? RFC 452: TELENET command at host LL, February 8,

1973. URL ftp://ftp.internic.net/rfc/rfc452.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc452.txt. Status: UNKNOWN. Not

online.

RFC0453

[447] M. D. Kudlick. RFC 453: Meeting announcement to discuss a network mail system, February 7, 1973. URL ftp://ftp.

internic.net/rfc/rfc453.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc453.txt. Status: UNKNOWN. Not online.

RFC0454

[448] A. M. McKenzie. RFC 454: File transfer protocol — meeting announcement and a new proposed document, February 16, 1973. URL ftp://

ftp.internic.net/rfc/rfc454.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc454.txt. Status: UNKNOWN. Not online.

RFC0455

[449] A. M. McKenzie. RFC 455: Traffic statistics (january 1973), February

12, 1973. URL ftp://ftp.internic.net/rfc/rfc455.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc455.txt. Status: UNKNOWN. Not

online.

RFC0456

[450] M. D. Kudlick. RFC 456: Memorandum: Date change of mail meet-

ing, February 13, 1973. URL ftp://ftp.internic.net/rfc/rfc456.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc456.txt.  Status:

UNKNOWN. Not online.

RFC0457

[451] D. C. Walden. RFC 457: TIPUG, February 15, 1973. URL ftp://

ftp.internic.net/rfc/rfc457.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc457.txt. Status: UNKNOWN. Not online.

RFC0458

[452] R. D. Bressler and R. Thomas. RFC 458: Mail retrieval via FTP,

February 20, 1973. URL ftp://ftp.internic.net/rfc/rfc458.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc458.txt.  Status: UN-

KNOWN. Not online.

RFC0459

[453] W. Kantrowitz. RFC 459: Network questionnaires, February 26,

1973. URL ftp://ftp.internic.net/rfc/rfc459.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc459.txt. Status: UNKNOWN. Not

online.

RFC0460

[454] C. Kline. RFC 460: NCP survey, February 13, 1973. URL ftp://

ftp.internic.net/rfc/rfc460.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc460.txt. Status: UNKNOWN. Not online.

RFC0461

[455] A. M. McKenzie. RFC 461: Telnet Protocol meeting announce-

ment, February 14, 1973. URL ftp://ftp.internic.net/rfc/rfc461.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc461.txt.  Status:

UNKNOWN. Not online.

RFC0462

[456] J. Iseli and D. Crocker. RFC 462: Responding to user needs, February

22, 1973. URL ftp://ftp.internic.net/rfc/rfc462.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc462.txt. Status: UNKNOWN. Not

online.

RFC0463

[457] A. K. Bhushan. RFC 463: FTP comments and response to RFC

430, February 21, 1973. URL ftp://ftp.internic.net/rfc/rfc463.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc463.txt.  Status:

UNKNOWN. Not online.

RFC0464

[458] M. D. Kudlick. RFC 464: Resource notebook framework, February

27, 1973. URL ftp://ftp.internic.net/rfc/rfc464.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc464.txt. Status: UNKNOWN. Not

online.

RFC0465

[459] ???? RFC 465: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc465.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc465.txt.

RFC0466

[460] J. M. Winett. RFC 466: Telnet logger/server for host LL-67, February

27, 1973. URL ftp://ftp.internic.net/rfc/rfc466.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc466.txt. Status: UNKNOWN. Not

online.

RFC0467

[461] J. D. Burchfiel and R. S. Tomlinson. RFC 467: Proposed change to host-host protocol: Resynchronization of connection status, February

20, 1973. URL ftp://ftp.internic.net/rfc/rfc467.txt,   ftp://
ftp.internic.net/rfc/rfc492.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc467.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc492.

txt. Updated by RFC0492 [486]. Status: UNKNOWN. Not online.

RFC0468

[462] R. T. Braden. RFC 468: FTP data compression, March 8,

1973. URL ftp://ftp.internic.net/rfc/rfc468.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc468.txt. Status: UNKNOWN. Not

online.

RFC0469

[463] M. D. Kudlick. RFC 469: Network mail meeting summary, March

8, 1973. URL ftp://ftp.internic.net/rfc/rfc469.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc469.txt. Status: UNKNOWN. Not

online.

RFC0470

[464] R. Thomas. RFC 470: Change in socket for TIP news facility, March

13, 1973. URL ftp://ftp.internic.net/rfc/rfc470.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc470.txt. Status: UNKNOWN. Not

online.

RFC0471

[465] R. Thomas. RFC 471: Workshop on multi-site executive programs,

March 13, 1973. URL ftp://ftp.internic.net/rfc/rfc471.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc471.txt.  Status: UN-

KNOWN. Not online.

RFC0472

[466] S. Bunch. RFC 472: Illinois’ reply to maxwell’s request for graphics information (NIC 14925), March 1, 1973. URL ftp://ftp.

internic.net/rfc/rfc472.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc472.txt. Status: UNKNOWN. Not online.

RFC0473

[467] D. C. Walden. RFC 473: MIX and MIXAL?, February 28,

1973. URL ftp://ftp.internic.net/rfc/rfc473.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc473.txt. Status: UNKNOWN. Not

online.

RFC0474

[468] S. Bunch. RFC 474: Announcement of NGWG meeting: Call for

papers, March 1, 1973. URL ftp://ftp.internic.net/rfc/rfc474.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc474.txt. Updated by

RFC0396 [391]. Status: UNKNOWN. Not online.

RFC0475

[469] A. K. Bhushan. RFC 475: FTP and network mail system, March

6, 1973. URL ftp://ftp.internic.net/rfc/rfc475.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc475.txt. Status: UNKNOWN. Not

online.

RFC0476

[470] A. M. McKenzie. RFC 476: IMP/TIP memory retrofit sched-

ule (rev 2.), March 7, 1973. URL ftp://ftp.internic.net/rfc/
rfc447.txt,   ftp://ftp.internic.net/rfc/rfc476.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc447.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc476.txt. Obsoletes RFC0447 [441]. Status: UN-KNOWN. Not online.

RFC0477

[471] M. Krilanovich. RFC 477: Remote job service at UCSB, May

23, 1973. URL ftp://ftp.internic.net/rfc/rfc477.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc477.txt. Status: UNKNOWN. Not

online.

RFC0478

[472] R. D. Bressler and R. Thomas. RFC 478: FTP server-server interaction

— II, March 26, 1973. URL ftp://ftp.internic.net/rfc/rfc478.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc478.txt.  Status:

UNKNOWN. Not online.

RFC0479

[473] J. E. White. RFC 479: Use of FTP by the NIC journal, March

8, 1973. URL ftp://ftp.internic.net/rfc/rfc479.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc479.txt. Status: UNKNOWN. Not

online.

RFC0480

[474] J. E. White. RFC 480: Host-dependent FTP parameters, March

8, 1973. URL ftp://ftp.internic.net/rfc/rfc480.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc480.txt. Status: UNKNOWN. Not

online.

RFC0481

[475] ???? RFC 481: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc481.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc481.txt.

RFC0482

[476] A. M. McKenzie. RFC 482: Traffic statistics (february 1973), March

12, 1973. URL ftp://ftp.internic.net/rfc/rfc482.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc482.txt. Status: UNKNOWN. Not

online.

RFC0483

[477] M. D. Kudlick. RFC 483: Cancellation of the resource notebook framework meeting, March 14, 1973. URL ftp://ftp.

internic.net/rfc/rfc483.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc483.txt. Status: UNKNOWN. Not online.

RFC0484

[478] ???? RFC 484: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc484.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc484.txt.

RFC0485

[479] J. R. Pickens. RFC 485: MIX and MIXAL at UCSB, March

19, 1973. URL ftp://ftp.internic.net/rfc/rfc485.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc485.txt. Status: UNKNOWN. Not

online.

RFC0486

[480] R. D. Bressler. RFC 486: Data transfer revisited, March 20,

1973. URL ftp://ftp.internic.net/rfc/rfc486.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc486.txt. Status: UNKNOWN. Not

online.

RFC0487

[481] R. D. Bressler. RFC 487: Free file transfer, April 6, 1973. URL ftp://

ftp.internic.net/rfc/rfc487.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc487.txt. Status: UNKNOWN. Not online.

RFC0488

[482] M. F. Auerbach. RFC 488: NLS classes at network sites, March

23, 1973. URL ftp://ftp.internic.net/rfc/rfc488.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc488.txt. Status: UNKNOWN. Not

online.

RFC0489

[483] J. Postel. RFC 489: Comment on resynchronization of connection status proposal, March 26, 1973. URL ftp://ftp.internic.net/rfc/

rfc489.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc489.txt. Sta-

tus: UNKNOWN. Not online.

RFC0490

[484] J. R. Pickens. RFC 490: Surrogate RJS for UCLA-CCN, March

6, 1973. URL ftp://ftp.internic.net/rfc/rfc490.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc490.txt. Status: UNKNOWN. Not

online.

RFC0491

[485] M. A. Padlipsky. RFC 491: What is “free”?, April 12, 1973. URL ftp://

ftp.internic.net/rfc/rfc491.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc491.txt. Status: UNKNOWN. Not online.

RFC0492

[486] E. W. Meyer, Jr. RFC 492: Response to RFC 467, April 18,

1973. URL ftp://ftp.internic.net/rfc/rfc467.txt,   ftp://
ftp.internic.net/rfc/rfc492.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc467.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc492.

txt. Updates RFC0467 [461]. Status: UNKNOWN. Not online.

RFC0493

[487] J. C. Michener, I. W. Cotton, K. C. Kelley, D. E. Liddle, and E. W. Meyer, Jr. RFC 493: Graphics protocol, April 26, 1973. URL ftp://

ftp.internic.net/rfc/rfc493.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc493.txt. Status: UNKNOWN. Not online.

RFC0494

[488] D. C. Walden. RFC 494: Availability of MIX and MIXAL in the net-

work, April 20, 1973. URL ftp://ftp.internic.net/rfc/rfc494.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc494.txt.  Status:

UNKNOWN. Not online.

RFC0495

[489] A. M. McKenzie. RFC 495: Telnet Protocol specifications, May

1, 1973. URL ftp://ftp.internic.net/rfc/rfc158.txt,   ftp://
ftp.internic.net/rfc/rfc495.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc158.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc495.

txt. Obsoletes RFC0158 [158]. Status: UNKNOWN. Not online.

RFC0496

[490] M. F. Auerbach. RFC 496: TNLS quick reference card is available, April

5, 1973. URL ftp://ftp.internic.net/rfc/rfc496.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc496.txt. Status: UNKNOWN. Not

online.

RFC0497

[491] A. M. McKenzie. RFC 497: Traffic statistics (march 1973), April

10, 1973. URL ftp://ftp.internic.net/rfc/rfc497.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc497.txt. Status: UNKNOWN. Not

online.

RFC0498

[492] R. T. Braden. RFC 498: On mail service to CCN, April 17,

1973. URL ftp://ftp.internic.net/rfc/rfc498.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc498.txt. Status: UNKNOWN. Not

online.

RFC0499

[493] B. R. Reussow. RFC 499: Harvard’s network RJE, April 1,

1973. URL ftp://ftp.internic.net/rfc/rfc499.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc499.txt. Status: UNKNOWN. Not

online.

RFC0500

[494] A. Shoshani and I. Spiegler. RFC 500: Integration of data management systems on a computer network, April 16, 1973. URL ftp://

ftp.internic.net/rfc/rfc500.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc500.txt. Status: UNKNOWN. Not online.

RFC0501

[495] K. T. Pogran. RFC 501: Un-muddling “free file transfer”, May

11, 1973. URL ftp://ftp.internic.net/rfc/rfc501.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc501.txt. Status: UNKNOWN. Not

online.

RFC0502

[496] ???? RFC 502: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc502.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc502.txt.

RFC0503

[497] N. Neigus and J. Postel. RFC 503: Socket number list, April 12,

1973. URL ftp://ftp.internic.net/rfc/rfc433.txt,   ftp://
ftp.internic.net/rfc/rfc503.txt, ftp://ftp.internic.net/rfc/
rfc739.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc433.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc503.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc739.txt. Obsoleted by RFC0739 [729]. Obsoletes RFC0433 [427]. Status: UNKNOWN. Not online.

RFC0504

[498] R. Thomas. RFC 504: Distributed resources workshop announce-

ment, April 30, 1973. URL ftp://ftp.internic.net/rfc/rfc504.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc504.txt.  Status:

UNKNOWN. Not online.

RFC0505

[499] M. A. Padlipsky. RFC 505: Two solutions to a file transfer access

problem, June 25, 1973. URL ftp://ftp.internic.net/rfc/rfc505.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc505.txt.  Status:

UNKNOWN. Not online.

RFC0506

[500] M. A. Padlipsky. RFC 506: FTP command naming problem, June

26, 1973. URL ftp://ftp.internic.net/rfc/rfc506.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc506.txt. Status: UNKNOWN. Not

online.

RFC0507

[501] ???? RFC 507: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc507.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc507.txt.

RFC0508

[502] L. Pfeifer and J. McAfee. RFC 508: Real-time data transmission on the

ARPANET, May 7, 1973. URL ftp://ftp.internic.net/rfc/rfc508.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc508.txt.  Status:

UNKNOWN. Not online.

RFC0509

[503] A. M. McKenzie. RFC 509: Traffic statistics (april 1973), April

7, 1973. URL ftp://ftp.internic.net/rfc/rfc509.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc509.txt. Status: UNKNOWN. Not

online.

RFC0510

[504] J. E. White. RFC 510: Request for network mailbox addresses, May

30, 1973. URL ftp://ftp.internic.net/rfc/rfc510.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc510.txt. Status: UNKNOWN. Not

online.

RFC0511

[505] J. B. North. RFC 511: Enterprise phone service to NIC from ARPANET

sites, May 23, 1973. URL ftp://ftp.internic.net/rfc/rfc511.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc511.txt.  Status:

UNKNOWN. Not online.

RFC0512

[506] W. Hathaway. RFC 512: More on lost message detection, May

23, 1973. URL ftp://ftp.internic.net/rfc/rfc512.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc512.txt. Status: UNKNOWN. Not

online.

RFC0513

[507] W. Hathaway. RFC 513: Comments on the new Telnet specifica-

tions, May 30, 1973. URL ftp://ftp.internic.net/rfc/rfc513.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc513.txt.  Status:

UNKNOWN. Not online.

RFC0514

[508] W. Kantrowitz. RFC 514: Network make-work, June 5, 1973. URL ftp:/

/ftp.internic.net/rfc/rfc514.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc514.txt. Status: UNKNOWN. Not online.

RFC0515

[509] R. Winter. RFC 515: Specifications for datalanguage: Version 0/9, June

6, 1973. URL ftp://ftp.internic.net/rfc/rfc515.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc515.txt. Status: UNKNOWN. Not

online.

RFC0516

[510] J. Postel. RFC 516: Lost message detection, May 18, 1973. URL ftp://

ftp.internic.net/rfc/rfc516.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc516.txt. Status: UNKNOWN. Not online.

RFC0517

[511] ???? RFC 517: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc517.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc517.txt.

RFC0518

[512] N. Vaughan and E. J. Feinler. RFC 518: ARPANET accounts, June

19, 1973. URL ftp://ftp.internic.net/rfc/rfc518.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc518.txt. Status: UNKNOWN. Not

online.

RFC0519

[513] J. R. Pickens. RFC 519: Resource evaluation, June 1, 1973. URL ftp://

ftp.internic.net/rfc/rfc519.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc519.txt. Status: UNKNOWN. Not online.

RFC0520

[514] J. D. Day. RFC 520: Memo to FTP group: Proposal for file access

protocol, June 25, 1973. URL ftp://ftp.internic.net/rfc/rfc520.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc520.txt.  Status:

UNKNOWN. Not online.

RFC0521

[515] A. M. McKenzie. RFC 521: Restricted use of IMP DDT, May

30, 1973. URL ftp://ftp.internic.net/rfc/rfc521.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc521.txt. Status: UNKNOWN. Not

online.

RFC0522

[516] A. M. McKenzie. RFC 522: Traffic statistics (may 1973), June

5, 1973. URL ftp://ftp.internic.net/rfc/rfc522.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc522.txt. Status: UNKNOWN. Not

online.

RFC0523

[517] A. K. Bhushan. RFC 523: SURVEY is in operation again, June

5, 1973. URL ftp://ftp.internic.net/rfc/rfc523.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc523.txt. Status: UNKNOWN. Not

online.

RFC0524

[518] J. E. White. RFC 524: Proposed mail protocol, June 13, 1973. URL ftp:

//ftp.internic.net/rfc/rfc524.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc524.txt. Status: UNKNOWN. Not online.

RFC0525

[519] W. Parrish and J. R. Pickens. RFC 525: MIT-MATHLAB meets UCSB-OLS — an example of resource sharing, June 1, 1973. URL ftp://

ftp.internic.net/rfc/rfc525.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc525.txt. Status: UNKNOWN. Not online.

RFC0526

[520] W. K. Pratt. RFC 526: Technical meeting: Digital image processing software systems, June 25, 1973. URL ftp://ftp.

internic.net/rfc/rfc526.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc526.txt. Status: UNKNOWN. Not online.

RFC0527

[521] D. L. Covill. RFC 527: ARPAWOCKY, May 1, 1973. URL ftp://

ftp.internic.net/rfc/rfc527.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc527.txt. Status: UNKNOWN. Format: TXT=1857 bytes.

RFC0528

[522] J. M. McQuillan. RFC 528: Software checksumming in the IMP and network reliability, June 20, 1973. URL ftp://ftp.

internic.net/rfc/rfc528.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc528.txt. Status: UNKNOWN. Not online.

RFC0529

[523] A. M. McKenzie, R. Thomas, R. S. Tomlinson, and K. T. Pogran. RFC 529: Note on protocol synch sequences, June 29, 1973. URL ftp://

ftp.internic.net/rfc/rfc529.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc529.txt. Status: UNKNOWN. Not online.

RFC0530

[524] A. K. Bhushan. RFC 530: Report on the survey project, June

22, 1973. URL ftp://ftp.internic.net/rfc/rfc530.txt,   ftp:/

/ftp.math.utah.edu/pub/rfc/rfc530.txt. Status: UNKNOWN. Not

online.

RFC0531

[525] M. A. Padlipsky. RFC 531: Feast or famine? A response to two recent RFC’s about network information, June 26, 1973. URL ftp://ftp.

internic.net/rfc/rfc531.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc531.txt. Status: UNKNOWN. Not online.

RFC0532

[526] R. G. Merryman. RFC 532: UCSD-CC server-FTP facility, July

12, 1973. URL ftp://ftp.internic.net/rfc/rfc532.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc532.txt. Status: UNKNOWN. Not

online.

RFC0533

[527] D. C. Walden. RFC 533: Message-ID numbers, July 17, 1973. URL ftp:/

/ftp.internic.net/rfc/rfc533.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc533.txt. Status: UNKNOWN. Not online.

RFC0534

[528] D. C. Walden. RFC 534: Lost message detection, July 17,

1973. URL ftp://ftp.internic.net/rfc/rfc534.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc534.txt. Status: UNKNOWN. Not

online.

RFC0535

[529] R. Thomas. RFC 535: Comments on file access protocol, July

25, 1973. URL ftp://ftp.internic.net/rfc/rfc535.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc535.txt. Status: UNKNOWN. Not

online.

RFC0536

[530] ???? RFC 536: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc536.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc536.txt.

RFC0537

[531] S. Bunch. RFC 537: Announcement of NGG meeting July 16–17, June

27, 1973. URL ftp://ftp.internic.net/rfc/rfc537.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc537.txt. Status: UNKNOWN. Not

online.

RFC0538

[532] A. M. McKenzie. RFC 538: Traffic statistics (june 1973), July

5, 1973. URL ftp://ftp.internic.net/rfc/rfc538.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc538.txt. Status: UNKNOWN. Not

online.

RFC0539

[533] D. Crocker and J. Postel. RFC 539: Thoughts on the mail protocol proposed in RFC 524, July 7, 1973. URL ftp://ftp.

internic.net/rfc/rfc539.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc539.txt. Status: UNKNOWN. Not online.

RFC0540

[534] ???? RFC 540: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc540.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc540.txt.

RFC0542

[535] N. Neigus. RFC 542: File transfer protocol, July 12, 1973. URL ftp:/

/ftp.internic.net/rfc/rfc354.txt,   ftp://ftp.internic.net/
rfc/rfc454.txt, ftp://ftp.internic.net/rfc/rfc495.txt,   ftp:
//ftp.internic.net/rfc/rfc542.txt, ftp://ftp.internic.net/rfc/
rfc765.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc354.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc454.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc495.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc542.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc765.txt.

Obsoleted by RFC0765 [755]. Obsoletes RFC0354 [349]. Status: UN-KNOWN. Format: TXT=100666 bytes. See also RFC354 [?], RFC454 [?], RFC495 [?].

RFC0543

[536] N. D. Meyer. RFC 543: Network journal submission and delivery, July

24, 1973. URL ftp://ftp.internic.net/rfc/rfc543.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc543.txt. Status: UNKNOWN. Not

online.

RFC0544

[537] N. D. Meyer and K. Kelley. RFC 544: Locating on-line documentation at

SRI-ARC, July 13, 1973. URL ftp://ftp.internic.net/rfc/rfc544.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc544.txt.  Status:

UNKNOWN. Not online.

RFC0545

[538] J. R. Pickens. RFC 545: Of what quality be the UCSB resources eval-

uators?, July 23, 1973. URL ftp://ftp.internic.net/rfc/rfc545.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc545.txt.  Status:

UNKNOWN. Not online.

RFC0546

[539] R. Thomas. RFC 546: Tenex load averages for July 1973, August

10, 1973. URL ftp://ftp.internic.net/rfc/rfc546.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc546.txt. Status: UNKNOWN. Not

online.

RFC0547

[540] D. C. Walden. RFC 547: Change to the very distant host specifica-

tion, August 13, 1973. URL ftp://ftp.internic.net/rfc/rfc547.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc547.txt.  Status:

UNKNOWN. Not online.

RFC0548

[541] D. C. Walden. RFC 548: Hosts using the IMP going down mes-

sage, August 16, 1973. URL ftp://ftp.internic.net/rfc/rfc548.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc548.txt.  Status:

UNKNOWN. Not online.

RFC0549

[542] J. C. Michener. RFC 549: Minutes of Network Graphics Group meeting, 15–17 July 1973, July 17, 1973. URL ftp://ftp.

internic.net/rfc/rfc549.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc549.txt. Status: UNKNOWN. Not online.

RFC0550

[543] L. P. Deutsch. RFC 550: NIC NCP experiment, August 24,

1973. URL ftp://ftp.internic.net/rfc/rfc550.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc550.txt. Status: UNKNOWN. Not

online.

RFC0551

[544] Y. Feinroth and R. Fink. RFC 551: [letter from Feinroth re: NYU, ANL, and LBL entering the net, and FTP protocol], August

27, 1973. URL ftp://ftp.internic.net/rfc/rfc551.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc551.txt.

RFC0552

[545] A. D. Owen. RFC 552: Single access to standard protocols, July

13, 1973. URL ftp://ftp.internic.net/rfc/rfc552.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc552.txt. Status: UNKNOWN. Not

online.

RFC0553

[546] C. H. Irby and K. Victor. RFC 553: Draft design for a text/graphics

protocol, July 14, 1973. URL ftp://ftp.internic.net/rfc/rfc553.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc553.txt.  Status:

UNKNOWN. Not online.

RFC0554

[547] ???? RFC 554: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc554.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc554.txt.

RFC0555

[548] J. E. White. RFC 555: Responses to critiques of the proposed mail

protocol, July 30, 1973. URL ftp://ftp.internic.net/rfc/rfc555.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc555.txt.  Status:

UNKNOWN. Not online.

RFC0556

[549] A. M. McKenzie. RFC 556: Traffic statistics (july 1973), August

13, 1973. URL ftp://ftp.internic.net/rfc/rfc556.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc556.txt. Status: UNKNOWN. Not

online.

RFC0557

[550] B. D. Wessler. RFC 557: Revelations in network host measure-

ments, August 30, 1973. URL ftp://ftp.internic.net/rfc/rfc557.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc557.txt.  Status:

UNKNOWN. Not online.

RFC0558

[551] ???? RFC 558: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc558.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc558.txt.

RFC0559

[552] A. K. Bhushan. RFC 559: Comments on the new Telnet Protocol and its implementation, August 15, 1973. URL ftp://ftp.

internic.net/rfc/rfc559.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc559.txt. Status: UNKNOWN. Format: TXT=10643 bytes.

RFC0560

[553] D. Crocker and J. Postel. RFC 560: Remote Controlled Transmission and Echoing Telnet option, August 18, 1973. URL ftp://ftp.

internic.net/rfc/rfc560.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc560.txt. Status: UNKNOWN. Not online.

RFC0561

[554] A. K. Bhushan, K. T. Pogran, R. S. Tomlinson, and J. E. White. RFC 561: Standardizing network mail headers, September 5, 1973. URL ftp:/

/ftp.internic.net/rfc/rfc561.txt,   ftp://ftp.internic.net/
rfc/rfc680.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc561.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc680.txt. Updated by RFC0680

[670]. Status: UNKNOWN. Format: TXT=6159 bytes.

RFC0562

[555] A. M. McKenzie. RFC 562: Modifications to the Telnet specifica-

tion, August 28, 1973. URL ftp://ftp.internic.net/rfc/rfc562.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc562.txt.  Status:

UNKNOWN. Not online.

RFC0563

[556] J. Davidson. RFC 563: Comments on the RCTE Telnet option, August

28, 1973. URL ftp://ftp.internic.net/rfc/rfc563.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc563.txt. Status: UNKNOWN. Not

online.

RFC0564

[557] ???? RFC 564: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc564.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc564.txt.

RFC0565

[558] D. Cantor. RFC 565: Storing network survey data at the datacom-

puter, August 28, 1973. URL ftp://ftp.internic.net/rfc/rfc565.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc565.txt.  Status:

UNKNOWN. Not online.

RFC0566

[559] A. M. McKenzie. RFC 566: Traffic statistics (august 1973), September

4, 1973. URL ftp://ftp.internic.net/rfc/rfc566.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc566.txt. Status: UNKNOWN. Not

online.

RFC0567

[560] L. P. Deutsch. RFC 567: Cross country network bandwidth, September

6, 1973. URL ftp://ftp.internic.net/rfc/rfc567.txt,   ftp://
ftp.internic.net/rfc/rfc568.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc567.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc568.

txt. Updated by RFC0568 [561]. Status: UNKNOWN. Format: TXT=1529 bytes.

RFC0568

[561] J. M. McQuillan. RFC 568: Response to RFC 567 — cross country network bandwidth, September 18, 1973. URL ftp:/

/ftp.internic.net/rfc/rfc567.txt,   ftp://ftp.internic.net/
rfc/rfc568.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc567.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc568.txt. Updates RFC0567

[560]. Status: UNKNOWN. Not online.

RFC0569

[562] M. A. Padlipsky. RFC 569: NETED: A common editor for the ARPA net-

work, October 15, 1973. URL ftp://ftp.internic.net/rfc/rfc569.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc569.txt.  Status:

HISTORIC. Format: TXT=17717 bytes.

RFC0570

[563] J. R. Pickens. RFC 570: Experimental input mapping between NVT ASCII and UCSB on line system, October 30, 1973. URL ftp://ftp.

internic.net/rfc/rfc570.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc570.txt. Status: UNKNOWN. Not online.

RFC0571

[564] R. T. Braden. RFC 571: Tenex FTP problem, November 15,

1973. URL ftp://ftp.internic.net/rfc/rfc571.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc571.txt. Status: UNKNOWN. Not

online.

RFC0572

[565] ???? RFC 572: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc572.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc572.txt.

RFC0573

[566] A. K. Bhushan. RFC 573: Data and file transfer: Some measurement results, September 14, 1973. URL ftp://ftp.internic.net/rfc/

rfc573.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc573.txt. Sta-

tus: UNKNOWN. Not online.

RFC0574

[567] M. Krilanovich. RFC 574: Announcement of a mail facility at UCSB,

September 26, 1973. URL ftp://ftp.internic.net/rfc/rfc574.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc574.txt.  Status:

UNKNOWN. Not online.

RFC0575

[568] ???? RFC 575: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc575.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc575.txt.

RFC0576

[569] K. Victor. RFC 576: Proposal for modifying linking, September

26, 1973. URL ftp://ftp.internic.net/rfc/rfc576.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc576.txt. Status: UNKNOWN. Not

online.

RFC0577

[570] D. Crocker. RFC 577: Mail priority, October 18, 1973. URL ftp://

ftp.internic.net/rfc/rfc577.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc577.txt. Status: UNKNOWN. Not online.

RFC0578

[571] A. K. Bhushan and N. D. Ryan. RFC 578: Using MIT-Mathlab MAC-SYMA from MIT-DMS muddle, October 29, 1973. URL ftp://ftp.

internic.net/rfc/rfc578.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc578.txt. Status: UNKNOWN. Not online.

RFC0579

[572] A. M. McKenzie. RFC 579: Traffic statistics (september 1973),

November 26, 1973. URL ftp://ftp.internic.net/rfc/rfc579.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc579.txt.  Status:

UNKNOWN. Not online.

RFC0580

[573] J. Postel. RFC 580: Note to protocol designers and imple-

menters, October 25, 1973. URL ftp://ftp.internic.net/rfc/
rfc580.txt,   ftp://ftp.internic.net/rfc/rfc582.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc580.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc582.txt. Updated by RFC0582 [575]. Status: UN-KNOWN. Format: TXT=1451 bytes.

RFC0581

[574] D. Crocker and J. Postel. RFC 581: Corrections to RFC 560: Remote Controlled Transmission and Echoing Telnet option, November

2, 1973. URL ftp://ftp.internic.net/rfc/rfc581.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc581.txt. Status: UNKNOWN. Not

online.

RFC0582

[575] R. Clements. RFC 582: Comments on RFC 580: Machine readable protocols, November 5, 1973. URL ftp://ftp.internic.net/rfc/

rfc580.txt,   ftp://ftp.internic.net/rfc/rfc582.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc580.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc582.txt. Updates RFC0580 [573]. Status: UN-KNOWN. Not online.

RFC0583

[576] ???? RFC 583: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc583.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc583.txt.

RFC0584

[577] J. Iseli, D. Crocker, and N. Neigus. RFC 584: Charter for ARPANET users interest working group, November 6, 1973. URL ftp://ftp.

internic.net/rfc/rfc584.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc584.txt. Status: UNKNOWN. Not online.

RFC0585

[578] D. Crocker, N. Neigus, E. J. Feinler, and J. Iseli. RFC 585: ARPANET users interest working group meeting, November 6, 1973. URL ftp://

ftp.internic.net/rfc/rfc585.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc585.txt. Status: UNKNOWN. Not online.

RFC0586

[579] A. M. McKenzie. RFC 586: Traffic statistics (october 1973), November

8, 1973. URL ftp://ftp.internic.net/rfc/rfc586.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc586.txt. Status: UNKNOWN. Not

online.

RFC0587

[580] J. Postel. RFC 587: Announcing new Telnet options, November

13, 1973. URL ftp://ftp.internic.net/rfc/rfc587.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc587.txt. Status: UNKNOWN. Not

online.

RFC0588

[581] A. V. Stokes. RFC 588: London node is now up, October 29,

1973. URL ftp://ftp.internic.net/rfc/rfc588.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc588.txt. Status: UNKNOWN. Not

online.

RFC0589

[582] R. T. Braden. RFC 589: CCN NETRJS server messages to remote

user, November 26, 1973. URL ftp://ftp.internic.net/rfc/rfc589.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc589.txt.  Status:

UNKNOWN. Not online.

RFC0590

[583] M. A. Padlipsky. RFC 590: MULTICS address change, November

19, 1973. URL ftp://ftp.internic.net/rfc/rfc590.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc590.txt. Status: UNKNOWN. Not

online.

RFC0591

[584] D. C. Walden. RFC 591: Addition to the very distant host specifications,

November 29, 1973. URL ftp://ftp.internic.net/rfc/rfc591.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc591.txt.  Status:

UNKNOWN. Not online.

RFC0592

[585] R. W. Watson. RFC 592: Some thoughts on system design to facilitate resource sharing, November 20, 1973. URL ftp://ftp.

internic.net/rfc/rfc592.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc592.txt. Status: UNKNOWN. Not online.

RFC0593

[586] A. M. McKenzie and J. Postel. RFC 593: Telnet and FTP implementation schedule change, November 29, 1973. URL ftp://ftp.

internic.net/rfc/rfc593.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc593.txt. Status: UNKNOWN. Not online.

RFC0594

[587] J. D. Burchfiel. RFC 594: Speedup of host-IMP interface, December

10, 1973. URL ftp://ftp.internic.net/rfc/rfc594.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc594.txt. Status: UNKNOWN. Not

online.

RFC0595

[588] W. Hathaway. RFC 595: Second thoughts in defense of the Telnet go-ahead, December 12, 1973. URL ftp://ftp.internic.net/rfc/

rfc595.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc595.txt. Sta-

tus: UNKNOWN. Not online.

RFC0596

[589] E. A. Taft. RFC 596: Second thoughts on Telnet go-ahead, December

8, 1973. URL ftp://ftp.internic.net/rfc/rfc596.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc596.txt. Status: UNKNOWN. Not

online.

RFC0597

[590] N. Neigus and E. J. Feinler. RFC 597: Host status, December 12,

1973. URL ftp://ftp.internic.net/rfc/rfc597.txt,   ftp://
ftp.internic.net/rfc/rfc603.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc597.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc603.

txt. Updated by RFC0603 [596]. Status: UNKNOWN. Not online.

RFC0598

[591] Network Information Center. Stanford Research Institute. RFC 598: RFC index — December 5, 1973, December 5, 1973. URL ftp://ftp.

internic.net/rfc/rfc598.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc598.txt. Status: UNKNOWN. Not online.

RFC0599

[592] R. T. Braden. RFC 599: Update on NETRJS, December 13,

1973. URL ftp://ftp.internic.net/rfc/rfc189.txt,   ftp://
ftp.internic.net/rfc/rfc599.txt, ftp://ftp.internic.net/rfc/
rfc740.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc189.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc599.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc740.txt. Obsoleted by RFC0740 [730]. Obsoletes RFC0189 [189]. Status: UNKNOWN. Format: TXT=16590 bytes.

RFC0600

[593] A. Berggreen. RFC 600: Interfacing an Illinois plasma terminal to the ARPANET, November 26, 1973. URL ftp://ftp.

internic.net/rfc/rfc600.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc600.txt. Status: UNKNOWN. Not online.

RFC0601

[594] A. M. McKenzie. RFC 601: Traffic statistics (november 1973), December

14, 1973. URL ftp://ftp.internic.net/rfc/rfc601.txt,   ftp://

ftp.math.utah.edu/pub/rfc/rfc601.txt. Status: UNKNOWN. Not

online.

RFC0602

[595] R. M. Metcalfe. RFC 602: “The stockings were hung by the chimney with care”, December 27, 1973. URL ftp://ftp.internic.net/rfc/

rfc602.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc602.txt. Sta-

tus: UNKNOWN. Not online.

RFC0603

[596] J. D. Burchfiel. RFC 603: Response to RFC 597: Host sta-

tus, December 31, 1973. URL ftp://ftp.internic.net/rfc/
rfc597.txt,   ftp://ftp.internic.net/rfc/rfc603.txt, ftp://
ftp.internic.net/rfc/rfc613.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc597.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc603.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc613.txt. Updates

RFC0597 [590]. Updated by RFC0613 [606]. Status: UNKNOWN. Not online.

RFC0604

[597] J. Postel. RFC 604: Assigned link numbers, December 26,

1973. URL ftp://ftp.internic.net/rfc/rfc317.txt,   ftp://
ftp.internic.net/rfc/rfc604.txt, ftp://ftp.internic.net/rfc/
rfc739.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc317.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc604.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc739.txt. Obsoleted by RFC0739 [729]. Obsoletes RFC0317 [312]. Status: UNKNOWN. Not online.

RFC0605

[598] ???? RFC 605: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc605.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc605.txt.

RFC0606

[599] L. P. Deutsch. RFC 606: Host names on-line, December 29,

1973. URL ftp://ftp.internic.net/rfc/rfc606.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc606.txt. Status: UNKNOWN. Not

online.

RFC0607

[600] M. Krilanovich and G. Gregg. RFC 607: Comments on the file transfer protocol, January 7, 1974. URL ftp://ftp.internic.net/rfc/

rfc607.txt,   ftp://ftp.internic.net/rfc/rfc614.txt, ftp://
ftp.internic.net/rfc/rfc624.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc607.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc614.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc624.txt. Obsoleted

by RFC0624 [617]. Updated by RFC0614 [607]. Status: UNKNOWN. Not online.

RFC0608

[601] M. D. Kudlick. RFC 608: Host names on-line, January 10,

1974. URL ftp://ftp.internic.net/rfc/rfc608.txt,   ftp://
ftp.internic.net/rfc/rfc810.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc608.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc810.

txt. Obsoleted by RFC0810 [800]. Status: UNKNOWN. Not online.

RFC0609

[602] B. Ferguson. RFC 609: Statement of upcoming move of NIC/NLS ser-

vice, January 10, 1974. URL ftp://ftp.internic.net/rfc/rfc609.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc609.txt.  Status:

UNKNOWN. Not online.

RFC0610

[603] R. Winter, J. Hill, and W. Greiff. RFC 610: Further datalanguage design concepts, December 15, 1973. URL ftp://ftp.

internic.net/rfc/rfc610.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc610.txt. Status: UNKNOWN. Not online.

RFC0611

[604] D. C. Walden. RFC 611: Two changes to the IMP/host protocol to improve user/network communications, February 14, 1974. URL ftp://

ftp.internic.net/rfc/rfc611.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc611.txt. Status: UNKNOWN. Not online.

RFC0612

[605] A. M. McKenzie. RFC 612: Traffic statistics (december 1973), January

16, 1974. URL ftp://ftp.internic.net/rfc/rfc612.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc612.txt. Status: UNKNOWN. Not

online.

RFC0613

[606] A. M. McKenzie. RFC 613: Network connectivity: A response to

RFC 603, January 21, 1974. URL ftp://ftp.internic.net/rfc/
rfc603.txt,   ftp://ftp.internic.net/rfc/rfc613.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc603.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc613.txt. Updates RFC0603 [596]. Status: UN-KNOWN. Not online.

RFC0614

[607] K. T. Pogran and N. Neigus. RFC 614: Response to RFC 607: “Comments on the File Transfer Protocol”, January 28,

1974. URL ftp://ftp.internic.net/rfc/rfc607.txt,   ftp://
ftp.internic.net/rfc/rfc614.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc607.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc614.

txt. Updates RFC0607 [600]. Status: UNKNOWN. Not online.

RFC0615

[608] D. Crocker. RFC 615: Proposed network standard data pathname

syntax, March 1, 1974. URL ftp://ftp.internic.net/rfc/rfc615.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc615.txt.  Status:

UNKNOWN. Not online.

RFC0616

[609] D. C. Walden. RFC 616: Latest network maps, February 11,

1973. URL ftp://ftp.internic.net/rfc/rfc616.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc616.txt. Status: UNKNOWN. Not

online.

RFC0617

[610] E. A. Taft. RFC 617: Note on socket number assignment, February

19, 1974. URL ftp://ftp.internic.net/rfc/rfc617.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc617.txt. Status: UNKNOWN. Not

online.

RFC0618

[611] E. A. Taft. RFC 618: Few observations on NCP statistics, February

19, 1974. URL ftp://ftp.internic.net/rfc/rfc618.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc618.txt. Status: UNKNOWN. Not

online.

RFC0619

[612] W. Naylor and H. Opderbeck. RFC 619: Mean round-trip times in

the ARPANET, March 7, 1974. URL ftp://ftp.internic.net/rfc/
rfc619.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc619.txt. Sta-

tus: UNKNOWN. Not online.

RFC0620

[613] B. Ferguson. RFC 620: Request for monitor host table updates, March

1, 1974. URL ftp://ftp.internic.net/rfc/rfc620.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc620.txt. Status: UNKNOWN. Not

online.

RFC0621

[614] M. D. Kudlick. RFC 621: NIC user directories at SRI ARC, March

6, 1974. URL ftp://ftp.internic.net/rfc/rfc621.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc621.txt. Status: UNKNOWN. Not

online.

RFC0622

[615] A. M. McKenzie. RFC 622: Scheduling IMP/TIP down time, March

13, 1974. URL ftp://ftp.internic.net/rfc/rfc622.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc622.txt. Status: UNKNOWN. Not

online.

RFC0623

[616] M. Krilanovich. RFC 623: Comments on on-line host name ser-

vice, February 22, 1974. URL ftp://ftp.internic.net/rfc/rfc623.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc623.txt.  Status:

UNKNOWN. Not online.

RFC0624

[617] M. Krilanovich, G. Gregg, W. Hathaway, and J. E. White. RFC 624: Comments on the file transfer protocol, February 28, 1974. URL ftp:/

/ftp.internic.net/rfc/rfc607.txt,   ftp://ftp.internic.net/
rfc/rfc624.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc607.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc624.txt. Obsoletes RFC0607

[600]. Status: UNKNOWN. Not online.

RFC0625

[618] M. D. Kudlick and E. J. Feinler. RFC 625: On-line hostnames ser-

vice, March 7, 1974. URL ftp://ftp.internic.net/rfc/rfc625.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc625.txt.  Status:

UNKNOWN. Not online.

RFC0626

[619] L. Kleinrock and H. Opderbeck. RFC 626: On a possible lockup condition in IMP subnet due to message sequencing, March 14,

1974. URL ftp://ftp.internic.net/rfc/rfc626.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc626.txt. Status: UNKNOWN. Not

online.

RFC0627

[620] M. D. Kudlick and E. J. Feinler. RFC 627: ASCII text file of host-

names, March 25, 1974. URL ftp://ftp.internic.net/rfc/rfc627.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc627.txt.  Status:

UNKNOWN. Not online.

RFC0628

[621] M. L. Keeney. RFC 628: Status of RFC numbers and a note on pre-assigned journal numbers, March 27, 1974. URL ftp://ftp.

internic.net/rfc/rfc628.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc628.txt. Status: UNKNOWN. Not online.

RFC0629

[622] J. B. North. RFC 629: Scenario for using the network journal, March

27, 1974. URL ftp://ftp.internic.net/rfc/rfc629.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc629.txt. Status: UNKNOWN. Not

online.

RFC0630

[623] J. Sussmann. RFC 630: FTP error code usage for more reliable mail

service, April 10, 1974. URL ftp://ftp.internic.net/rfc/rfc630.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc630.txt.  Status:

UNKNOWN. Not online.

RFC0631

[624] A. Danthine. RFC 631: International meeting on minicomputers and data communication: Call for papers, April 17, 1974. URL ftp://

ftp.internic.net/rfc/rfc631.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc631.txt. Status: UNKNOWN. Not online.

RFC0632

[625] H. Opderbeck. RFC 632: Throughput degradations for single packet

messages, May 20, 1974. URL ftp://ftp.internic.net/rfc/rfc632.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc632.txt.  Status:

UNKNOWN. Not online.

RFC0633

[626] A. M. McKenzie. RFC 633: IMP/TIP preventive maintenance

schedule, March 18, 1974. URL ftp://ftp.internic.net/rfc/
rfc633.txt,   ftp://ftp.internic.net/rfc/rfc638.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc633.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc638.txt. Obsoleted by RFC0638 [631]. Status: UN-KNOWN. Not online.

RFC0634

[627] A. M. McKenzie. RFC 634: Change in network address for hask-

ins lab, April 10, 1974. URL ftp://ftp.internic.net/rfc/rfc634.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc634.txt.  Status:

UNKNOWN. Not online.

RFC0635

[628] V. G. Cerf. RFC 635: Assessment of ARPANET protocols, April

22, 1974. URL ftp://ftp.internic.net/rfc/rfc635.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc635.txt. Status: UNKNOWN. Not

online.

RFC0636

[629] J. D. Burchfiel, B. Cosell, R. S. Tomlinson, and D. C. Walden. RFC 636: TIP/Tenex reliability improvements, June 10, 1974. URL ftp://ftp.

internic.net/rfc/rfc636.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc636.txt. Status: UNKNOWN. Not online.

RFC0637

[630] A. M. McKenzie. RFC 637: Change of network address for SU-DSL, April

23, 1974. URL ftp://ftp.internic.net/rfc/rfc637.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc637.txt. Status: UNKNOWN. Not

online.

RFC0638

[631] A. M. McKenzie. RFC 638: IMP/TIP preventive maintenance

schedule, April 25, 1974. URL ftp://ftp.internic.net/rfc/
rfc633.txt,   ftp://ftp.internic.net/rfc/rfc638.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc633.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc638.txt. Obsoletes RFC0633 [626]. Status: UN-KNOWN. Not online.

RFC0639

[632] ???? RFC 639: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc639.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc639.txt.

RFC0640

[633] J. Postel. RFC 640: Revised FTP reply codes, June 5, 1974. URL ftp://

ftp.internic.net/rfc/rfc640.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc640.txt. Status: UNKNOWN. Not online.

RFC0641

[634] ???? RFC 641: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc641.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc641.txt.

RFC0642

[635] J. D. Burchfiel. RFC 642: Ready line philosophy and implemen-

tation, July 5, 1974. URL ftp://ftp.internic.net/rfc/rfc642.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc642.txt.  Status:

UNKNOWN. Not online.

RFC0643

[636] E. Mader. RFC 643: Network debugging protocol, July 1, 1974. URL

ftp://ftp.internic.net/rfc/rfc643.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc643.txt. Status: UNKNOWN. Not online.

RFC0644

[637] R. Thomas. RFC 644: On the problem of signature authentication for network mail, July 22, 1974. URL ftp://ftp.internic.net/rfc/

rfc644.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc644.txt. Sta-

tus: UNKNOWN. Not online.

RFC0645

[638] D. Crocker. RFC 645: Network standard data specification syntax, June

26, 1974. URL ftp://ftp.internic.net/rfc/rfc645.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc645.txt. Status: UNKNOWN. Not

online.

RFC0646

[639] ???? RFC 646: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc646.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc646.txt.

RFC0647

[640] M. A. Padlipsky. RFC 647: Proposed protocol for connecting host computers to ARPA-like networks via front end processors, November

12, 1974. URL ftp://ftp.internic.net/rfc/rfc647.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc647.txt. Status: UNKNOWN. Not

online.

RFC0648

[641] ???? RFC 648: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc648.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc648.txt.

RFC0651

[642] D. Crocker. RFC 651: Revised Telnet status option, October 25,

1974. URL ftp://ftp.internic.net/rfc/rfc651.txt,   ftp://
ftp.internic.net/rfc/rfc859.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc651.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc859.

txt. Obsoleted by RFC0859 [849]. Status: UNKNOWN. Format: TXT=4360 bytes.

RFC0652

[643] D. Crocker. RFC 652: Telnet output carriage-return disposition option,

October 25, 1974. URL ftp://ftp.internic.net/rfc/rfc652.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc652.txt. Status: HISTORIC.

Format: TXT=7003 bytes.

RFC0653

[644] D. Crocker. RFC 653: Telnet output horizontal tabstops option, Oc-

tober 25, 1974. URL ftp://ftp.internic.net/rfc/rfc653.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc653.txt. Status: HISTORIC.

Format: TXT=4694 bytes.

RFC0654

[645] D. Crocker. RFC 654: Telnet output horizontal tab disposition option,

October 25, 1974. URL ftp://ftp.internic.net/rfc/rfc654.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc654.txt. Status: HISTORIC.

Format: TXT=6158 bytes.

RFC0655

[646] D. Crocker. RFC 655: Telnet output formfeed disposition option, Oc-

tober 25, 1974. URL ftp://ftp.internic.net/rfc/rfc655.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc655.txt. Status: HISTORIC.

Format: TXT=5991 bytes.

RFC0656

[647] D. Crocker. RFC 656: Telnet output vertical tabstops option, October

25, 1974. URL ftp://ftp.internic.net/rfc/rfc656.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc656.txt. Status: HISTORIC. For-

mat: TXT=4858 bytes.

RFC0657

[648] D. Crocker. RFC 657: Telnet output vertical tab disposition option,

October 25, 1974. URL ftp://ftp.internic.net/rfc/rfc657.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc657.txt. Status: HISTORIC.

Format: TXT=5759 bytes.

RFC0658

[649] D. Crocker. RFC 658: Telnet output linefeed disposition, October

25, 1974. URL ftp://ftp.internic.net/rfc/rfc658.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc658.txt. Status: HISTORIC. For-

mat: TXT=6484 bytes.

RFC0659

[650] J. Postel. RFC 659: Announcing additional Telnet options, October

18, 1974. URL ftp://ftp.internic.net/rfc/rfc659.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc659.txt. Status: UNKNOWN. Not

online.

RFC0660

[651] D. C. Walden. RFC 660: Some changes to the IMP and the IMP/host in-

terface, October 23, 1974. URL ftp://ftp.internic.net/rfc/rfc660.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc660.txt.  Status:

UNKNOWN. Format: TXT=4991 bytes.

RFC0661

[652] J. Postel. RFC 661: Protocol information, November 23, 1974. URL ftp:

//ftp.internic.net/rfc/rfc661.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc661.txt. Status: UNKNOWN. Not online.

RFC0662

[653] R. Kanodia. RFC 662: Performance improvement in ARPANET file transfers from multics, November 26, 1974. URL ftp://ftp.

internic.net/rfc/rfc662.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc662.txt. Status: UNKNOWN. Format: TXT=8872 bytes.

RFC0663

[654] R. Kanodia. RFC 663: Lost message detection and recovery proto-

col, November 29, 1974. URL ftp://ftp.internic.net/rfc/rfc663.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc663.txt.  Status:

UNKNOWN.

RFC0664

[655] ???? RFC 664: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc664.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc664.txt.

RFC0666

[656] M. A. Padlipsky. RFC 666: Specification of the unified user-level proto-

col, November 26, 1974. URL ftp://ftp.internic.net/rfc/rfc666.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc666.txt.  Status:

UNKNOWN. Not online.

RFC0667

[657] S. G. Chipman. RFC 667: BBN host ports, December 17,

1974. URL ftp://ftp.internic.net/rfc/rfc667.txt,   ftp://

ftp.math.utah.edu/pub/rfc/rfc667.txt. Status: UNKNOWN. Not

online.

RFC0668

[658] ???? RFC 668: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc668.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc668.txt.

RFC0669

[659] D. W. Dodds. RFC 669: November, 1974, survey of New-Protocol Telnet servers, December 4, 1974. URL ftp://ftp.

internic.net/rfc/rfc669.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc669.txt. Status: UNKNOWN. Not online.

RFC0670

[660] ???? RFC 670: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc670.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc670.txt.

RFC0671

[661] R. Schantz. RFC 671: Note on reconnection protocol, December

6, 1974. URL ftp://ftp.internic.net/rfc/rfc671.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc671.txt. Status: UNKNOWN. Not

online.

RFC0672

[662] R. Schantz. RFC 672: Multi-site data collection facility, December

6, 1974. URL ftp://ftp.internic.net/rfc/rfc672.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc672.txt. Status: UNKNOWN. Not

online.

RFC0673

[663] ???? RFC 673: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc673.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc673.txt.

RFC0674

[664] J. Postel and J. E. White. RFC 674: Procedure call documents: Version

2, December 12, 1974. URL ftp://ftp.internic.net/rfc/rfc674.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc674.txt.  Status:

UNKNOWN. Not online.

RFC0675

[665] V. G. Cerf, Y. K. Dalal, and C. A. Sunshine. RFC 675: Specification of Internet Transmission Control Program, December 1,

1974. URL ftp://ftp.internic.net/rfc/rfc675.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc675.txt. Status: UNKNOWN. Not

online.

RFC0676

[666] ???? RFC 676: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc676.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc676.txt.

RFC0677

[667] P. R. Johnson and R. Thomas. RFC 677: Maintenance of duplicate databases, January 27, 1975. URL ftp://ftp.internic.net/rfc/

rfc677.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc677.txt. Sta-

tus: UNKNOWN. Not online.

RFC0678

[668] J. Postel. RFC 678: Standard file formats, December 19, 1974. URL ftp:

//ftp.internic.net/rfc/rfc678.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc678.txt. Status: UNKNOWN. Not online.

RFC0679

[669] D. W. Dodds. RFC 679: February, 1975, survey of New-Protocol Telnet servers, February 21, 1975. URL ftp://ftp.internic.net/rfc/

rfc679.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc679.txt. Sta-

tus: UNKNOWN. Not online.

RFC0680

[670] T. H. Myer and D. A. Henderson. RFC 680: Message transmission protocol, April 30, 1975. URL ftp://ftp.internic.net/rfc/

rfc561.txt,   ftp://ftp.internic.net/rfc/rfc680.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc561.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc680.txt. Updates RFC0561 [554]. Status: UN-KNOWN. Not online.

RFC0681

[671] S. Holmgren. RFC 681: Network UNIX, March 18, 1975. URL ftp://

ftp.internic.net/rfc/rfc681.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc681.txt. Status: UNKNOWN. Not online.

RFC0682

RFC0683

[673] R. Clements. RFC 683: FTPSRV — Tenex extension for paged

files, April 3, 1975. URL ftp://ftp.internic.net/rfc/rfc683.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc683.txt.  Status:

UNKNOWN. Not online.

RFC0684

[674] R. Schantz. RFC 684: Commentary on procedure calling as a network

protocol, April 15, 1975. URL ftp://ftp.internic.net/rfc/rfc684.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc684.txt.  Status:

UNKNOWN. Not online.

RFC0685

[675] M. Beeler. RFC 685: Response time in cross network debugging, April

16, 1975. URL ftp://ftp.internic.net/rfc/rfc685.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc685.txt. Status: UNKNOWN. Not

online.

RFC0686

[676] B. Harvey. RFC 686: Leaving well enough alone, May 10,

1975. URL ftp://ftp.internic.net/rfc/rfc686.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc686.txt. Status: UNKNOWN. Not

online.

RFC0687

[677] D. C. Walden. RFC 687: IMP/host and host/IMP protocol changes,

June 2, 1975. URL ftp://ftp.internic.net/rfc/rfc687.txt,
 ftp://ftp.internic.net/rfc/rfc690.txt, ftp://ftp.internic.
net/rfc/rfc704.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc687.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc690.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc704.txt. Obsoleted by RFC0704

[694]. Updated by RFC0690 [680]. Status: UNKNOWN. Not online.

RFC0688

[678] D. C. Walden. RFC 688: Tentative schedule for the new Telnet implementation for the TIP, June 4, 1975. URL ftp://ftp.

internic.net/rfc/rfc688.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc688.txt. Status: UNKNOWN. Not online.

RFC0689

[679] R. Clements. RFC 689: Tenex NCP finite state machine for connec-

tions, May 23, 1975. URL ftp://ftp.internic.net/rfc/rfc689.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc689.txt.  Status:

UNKNOWN. Not online.

RFC0690

[680] J. Postel. RFC 690: Comments on the proposed host/IMP pro-

tocol changes, June 6, 1975. URL ftp://ftp.internic.net/rfc/
rfc687.txt,   ftp://ftp.internic.net/rfc/rfc690.txt, ftp://
ftp.internic.net/rfc/rfc692.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc687.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc690.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc692.txt. Updates

RFC0687 [677]. Updated by RFC0692 [682]. Status: UNKNOWN. Not online.

RFC0691

[681] B. Harvey. RFC 691: One more try on the FTP, May 28,

1975. URL ftp://ftp.internic.net/rfc/rfc691.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc691.txt. Status: UNKNOWN. Not

online.

RFC0692

[682] S. M. Wolfe. RFC 692: Comments on IMP/host protocol changes (RFCs

687 and 690), June 20, 1975. URL ftp://ftp.internic.net/rfc/
rfc690.txt,   ftp://ftp.internic.net/rfc/rfc692.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc690.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc692.txt. Updates RFC0690 [680]. Status: UN-KNOWN. Not online.

RFC0693

[683] ???? RFC 693: Not issued, 19xx. URL ftp://ftp.internic.net/rfc/

rfc693.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc693.txt.

RFC0694

[684] J. Postel. RFC 694: Protocol information, June 18, 1975. URL ftp://

ftp.internic.net/rfc/rfc694.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc694.txt. Status: UNKNOWN. Not online.

RFC0695

[685] M. Krilanovich. RFC 695: Official change in host-host protocol, July

5, 1975. URL ftp://ftp.internic.net/rfc/rfc695.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc695.txt. Status: UNKNOWN. Not

online.

RFC0696

[686] V. G. Cerf. RFC 696: Comments on the IMP/host and host/IMP protocol changes, July 13, 1975. URL ftp://ftp.internic.net/rfc/

rfc696.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc696.txt. Sta-

tus: UNKNOWN. Not online.

RFC0697

[687] J. Lieb. RFC 697: CWD command of FTP, July 14, 1975. URL ftp://

ftp.internic.net/rfc/rfc697.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc697.txt. Status: UNKNOWN. Not online.

RFC0698

[688] T. Mock. RFC 698: Telnet extended ASCII option, July 23,

1975.  URL ftp://ftp.internic.net/rfc/rfc698.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc698.txt. Status: PROPOSED

STANDARD. Not online.

RFC0699

[689] J. Postel and J. Vernon. RFC 699: Request for comments summary notes: 600-699, November 1, 1982. URL ftp://ftp.

internic.net/rfc/rfc699.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc699.txt. Status: INFORMATIONAL. Not online.

RFC0700

[690] E. Mader, W. W. Plummer, and R. S. Tomlinson. RFC 700: Protocol experiment, August 1, 1974. URL ftp://ftp.internic.net/rfc/

rfc700.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc700.txt. Sta-

tus: UNKNOWN. Not online.

RFC0701

[691] D. W. Dodds. RFC 701: August, 1974, survey of New-Protocol Telnet

servers, August 1, 1974. URL ftp://ftp.internic.net/rfc/rfc701.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc701.txt.  Status:

UNKNOWN. Not online.

RFC0702

[692] D. W. Dodds. RFC 702: September, 1974, survey of New-Protocol Telnet servers, September 25, 1974. URL ftp://ftp.

internic.net/rfc/rfc702.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc702.txt. Status: UNKNOWN. Not online.

RFC0703

[693] D. W. Dodds. RFC 703: July, 1975, survey of New-Protocol Telnet

servers, July 11, 1975. URL ftp://ftp.internic.net/rfc/rfc703.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc703.txt.  Status:

UNKNOWN. Not online.

RFC0704

[694] P. J. Santos. RFC 704: IMP/host and host/IMP protocol change,

September 15, 1975. URL ftp://ftp.internic.net/rfc/rfc687.txt,

 ftp://ftp.internic.net/rfc/rfc704.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc687.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc704.

txt. Obsoletes RFC0687 [677]. Status: UNKNOWN. Not online.

RFC0705

[695] R. F. Bryan. RFC 705: Front-end protocol B6700 version, November

5, 1975. URL ftp://ftp.internic.net/rfc/rfc705.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc705.txt. Status: UNKNOWN. Not

online.

RFC0706

[696] J. Postel. RFC 706: On the junk mail problem, November 8,

1975. URL ftp://ftp.internic.net/rfc/rfc706.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc706.txt. Status: UNKNOWN. Not

online.

RFC0707

[697] J. E. White. RFC 707: High-level framework for network-based resource

sharing, December 23, 1975. URL ftp://ftp.internic.net/rfc/
rfc707.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc707.txt. Sta-

tus: UNKNOWN. Not online.

RFC0708

[698] J. E. White. RFC 708: Elements of a distributed programming sys-

tem, January 28, 1976. URL ftp://ftp.internic.net/rfc/rfc708.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc708.txt.  Status:

UNKNOWN. Not online.

RFC0709

[699] ???? RFC 709: (Not Issued), November 1993. URL ftp://ftp.

internic.net/rfc/rfc709.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc709.txt. Status: UNKNOWN. Format: TXT=27 bytes.

RFC0710

[700] ???? RFC 710: (Not Issued), November 1993. URL ftp://ftp.

internic.net/rfc/rfc710.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc710.txt. Status: UNKNOWN. Format: TXT=27 bytes.

RFC0711

[701] ???? RFC 711: (Not Issued), November 1993. URL ftp://ftp.

internic.net/rfc/rfc711.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc711.txt. Status: UNKNOWN. Format: TXT=27 bytes.

RFC0712

[702] J. E. Donnelley. RFC 712: Distributed capability computing system (DCCS), February 5, 1976. URL ftp://ftp.internic.net/rfc/

rfc712.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc712.txt. Sta-

tus: UNKNOWN. Not online.

RFC0713

[703] J. Haverty. RFC 713: MSDTP-message services data transmission

protocol, April 6, 1976. URL ftp://ftp.internic.net/rfc/rfc713.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc713.txt.  Status:

UNKNOWN. Not online.

RFC0714

[704] A. M. McKenzie. RFC 714: Host-host protocol for an ARPANET-type

network, April 21, 1976. URL ftp://ftp.internic.net/rfc/rfc714.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc714.txt.  Status:

UNKNOWN. Not online.

RFC0715

[705] ???? RFC 715: (Not Issued), November 1993. URL ftp://ftp.

internic.net/rfc/rfc715.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc715.txt. Status: UNKNOWN. Format: TXT=27 bytes.

RFC0716

[706] D. C. Walden and J. Levin. RFC 716: Interim revision to Appendix F of

BBN 1822, May 24, 1976. URL ftp://ftp.internic.net/rfc/rfc716.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc716.txt.  Status:

UNKNOWN. Not online.

RFC0717

[707] J. Postel. RFC 717: Assigned network numbers, July 1, 1976. URL ftp:/

/ftp.internic.net/rfc/rfc717.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc717.txt. Status: HISTORIC. Not online.

RFC0718

[708] J. Postel. RFC 718: Comments on RCTE from the Tenex implementation experience, June 30, 1976. URL ftp://ftp.internic.net/rfc/

rfc718.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc718.txt. Sta-

tus: UNKNOWN. Not online.

RFC0719

[709] J. Postel. RFC 719: Discussion on RCTE, July 22, 1976. URL ftp://

ftp.internic.net/rfc/rfc719.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc719.txt. Status: UNKNOWN. Not online.

RFC0720

[710] D. Crocker. RFC 720: Address specification syntax for network

mail, August 5, 1976. URL ftp://ftp.internic.net/rfc/rfc720.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc720.txt.  Status:

UNKNOWN. Not online.

RFC0721

[711] L. L. Garlick. RFC 721: Out-of-band control signals in a host-to-host pro-

tocol, September 1, 1976. URL ftp://ftp.internic.net/rfc/rfc721.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc721.txt.  Status:

UNKNOWN. Not online.

RFC0722

[712] J. Haverty. RFC 722: Thoughts on interactions in distributed services,

September 16, 1976. URL ftp://ftp.internic.net/rfc/rfc722.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc722.txt.  Status:

UNKNOWN. Not online.

RFC0723

[713] ???? RFC 723, November 1993. URL ftp://ftp.internic.net/rfc/

rfc723.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc723.txt. Sta-

tus: UNKNOWN. Not online.

RFC0724

[714] D. Crocker, K. T. Pogran, J. Vittal, and D. A. Henderson. RFC 724: Proposed official standard for the format of ARPA network messages, May

12, 1977. URL ftp://ftp.internic.net/rfc/rfc724.txt,   ftp://
ftp.internic.net/rfc/rfc733.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc724.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc733.

txt. Obsoleted by RFC0733 [723]. Status: UNKNOWN. Not online.

RFC0725

[715] J. D. Day and G. R. Grossman. RFC 725: RJE protocol for a resource sharing network, March 1, 1977. URL ftp://ftp.

internic.net/rfc/rfc725.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc725.txt. Status: UNKNOWN. Not online.

RFC0726

[716] J. Postel and D. Crocker. RFC 726: Remote Controlled Transmission and Echoing Telnet option, March 8, 1977. URL ftp://ftp.

internic.net/rfc/rfc726.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc726.txt. Status: PROPOSED STANDARD. Not online.

RFC0727

[717] M. R. Crispin. RFC 727: Telnet logout option, April 27, 1977. URL ftp:

//ftp.internic.net/rfc/rfc727.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc727.txt. Status: PROPOSED STANDARD. Not online.

RFC0728

[718] J. D. Day. RFC 728: Minor pitfall in the Telnet Protocol, April

27, 1977. URL ftp://ftp.internic.net/rfc/rfc728.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc728.txt. Status: UNKNOWN. Not

online.

RFC0729

[719] D. Crocker. RFC 729: Telnet byte macro option, May 13,

1977. URL ftp://ftp.internic.net/rfc/rfc729.txt,   ftp://
ftp.internic.net/rfc/rfc735.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc729.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc735.

txt. Obsoleted by RFC0735 [725]. Status: UNKNOWN. Not online.

RFC0730

[720] J. Postel. RFC 730: Extensible field addressing, May 20, 1977. URL ftp:

//ftp.internic.net/rfc/rfc730.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc730.txt. Status: UNKNOWN.

RFC0731

[721] J. D. Day. RFC 731: Telnet Data Entry Terminal option, June

27, 1977. URL ftp://ftp.internic.net/rfc/rfc731.txt,   ftp://
ftp.internic.net/rfc/rfc732.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc731.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc732.

txt. Obsoleted by RFC0732 [722]. Status: UNKNOWN.

RFC0732

[722] J. D. Day. RFC 732: Telnet Data Entry Terminal option, Septem-

ber 12, 1977. URL ftp://ftp.internic.net/rfc/rfc1043.txt,
 ftp://ftp.internic.net/rfc/rfc731.txt, ftp://ftp.internic.
net/rfc/rfc732.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1043.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc731.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc732.txt. Obsoletes RFC0731 [721].

Updated by RFC1043 [1033]. Status: UNKNOWN.

RFC0733

[723] D. Crocker, J. Vittal, K. T. Pogran, and D. A. Henderson. RFC 733: Standard for the format of ARPA network text messages,

November 21, 1977. URL ftp://ftp.internic.net/rfc/rfc724.txt,
 ftp://ftp.internic.net/rfc/rfc733.txt, ftp://ftp.internic.
net/rfc/rfc822.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc724.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc733.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc822.txt. Obsoleted by RFC0822

[812]. Obsoletes RFC0724 [714]. Status: UNKNOWN.

RFC0734

[724] M. R. Crispin. RFC 734: SUPDUP protocol, October 7, 1977. URL ftp:

//ftp.internic.net/rfc/rfc734.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc734.txt. Status: HISTORIC.

RFC0735

[725] D. Crocker and R. H. Gumpertz. RFC 735: Revised Telnet byte macro option, November 3, 1977. URL ftp://ftp.internic.net/rfc/

rfc729.txt,   ftp://ftp.internic.net/rfc/rfc735.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc729.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc735.txt. Obsoletes RFC0729 [719]. Status: PRO-POSED STANDARD.

RFC0736

[726] M. R. Crispin. RFC 736: Telnet SUPDUP option, October 31,

1977.  URL ftp://ftp.internic.net/rfc/rfc736.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc736.txt. Status: PROPOSED

STANDARD.

RFC0737

[727] K. Harrenstien. RFC 737: FTP extension: XSEN, October 31,

1977. URL ftp://ftp.internic.net/rfc/rfc737.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc737.txt. Status: UNKNOWN.

RFC0738

[728] K. Harrenstien. RFC 738: Time server, October 31, 1977. URL ftp://

ftp.internic.net/rfc/rfc738.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc738.txt. Status: UNKNOWN.

RFC0739

[729] J. Postel. RFC 739: Assigned numbers, November 11, 1977. URL ftp:/

/ftp.internic.net/rfc/rfc503.txt,   ftp://ftp.internic.net/
rfc/rfc604.txt, ftp://ftp.internic.net/rfc/rfc739.txt,   ftp:
//ftp.internic.net/rfc/rfc750.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc503.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc604.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc739.txt,   ftp://

ftp.math.utah.edu/pub/rfc/rfc750.txt. Obsoleted by RFC0750

[740]. Obsoletes RFC0604, RFC0503 [597, 497]. Status: HISTORIC.

RFC0740

[730] R. T. Braden. RFC 740: NETRJS protocol, November 22,

1977. URL ftp://ftp.internic.net/rfc/rfc599.txt,   ftp://
ftp.internic.net/rfc/rfc740.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc599.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc740.

txt. Obsoletes RFC0599 [592]. Status: HISTORIC. Not online.

RFC0741

[731] D. Cohen. RFC 741: Specifications for the network voice protocol (NVP),

November 22, 1977. URL ftp://ftp.internic.net/rfc/rfc741.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc741.txt.  Status:

UNKNOWN. Not online.

RFC0742

[732] K. Harrenstien. RFC 742: NAME/FINGER protocol, December 30,

1977. URL ftp://ftp.internic.net/rfc/rfc1194.txt,   ftp://
ftp.internic.net/rfc/rfc1288.txt,   ftp://ftp.internic.net/
rfc/rfc742.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1194.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1288.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc742.txt. Obsoleted by RFC1288, RFC1194

[1278, 1184]. Status: UNKNOWN. Not online.

RFC0743

[733] K. Harrenstien. RFC 743: FTP extension: XRSQ/XRCP, December

30, 1977. URL ftp://ftp.internic.net/rfc/rfc743.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc743.txt. Status: UNKNOWN. Not

online.

RFC0744

[734] J. Sattley. RFC 744: MARS — a message archiving and retrieval ser-

vice, January 8, 1978. URL ftp://ftp.internic.net/rfc/rfc744.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc744.txt.  Status:

UNKNOWN. Not online.

RFC0745

[735] M. Beeler. RFC 745: JANUS interface specifications, March 30,

1978. URL ftp://ftp.internic.net/rfc/rfc745.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc745.txt. Status: UNKNOWN. Not

online.

RFC0746

[736] R. Stallman. RFC 746: SUPDUP graphics extension, March 17,

1978. URL ftp://ftp.internic.net/rfc/rfc746.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc746.txt. Status: UNKNOWN. Not

online.

RFC0747

[737] M. R. Crispin. RFC 747: Recent extensions to the SUPDUP pro-

tocol, March 21, 1978. URL ftp://ftp.internic.net/rfc/rfc747.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc747.txt.  Status:

UNKNOWN. Not online.

RFC0748

[738] M. R. Crispin. RFC 748: Telnet randomly-lose option, April 1,

1978. URL ftp://ftp.internic.net/rfc/rfc748.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc748.txt. Status: UNKNOWN. Not

online.

RFC0749

[739] B. Greenberg. RFC 749: Telnet SUPDUP-output option, September

18, 1978. URL ftp://ftp.internic.net/rfc/rfc749.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc749.txt. Status: PROPOSED

STANDARD. Not online.

RFC0750

[740] J. Postel. RFC 750: Assigned numbers, September 26, 1978. URL ftp:/

/ftp.internic.net/rfc/rfc739.txt,   ftp://ftp.internic.net/
rfc/rfc750.txt, ftp://ftp.internic.net/rfc/rfc755.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc739.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc750.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc755.

txt. Obsoleted by RFC0755 [745]. Obsoletes RFC0739 [729]. Status: HISTORIC. Not online.

RFC0751

[741] P. D. Lebling. RFC 751: Survey of FTP mail and MLFL, December

10, 1978. URL ftp://ftp.internic.net/rfc/rfc751.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc751.txt. Status: UNKNOWN. Not

online.

RFC0752

[742] M. R. Crispin. RFC 752: Universal host table, January 2,

1979. URL ftp://ftp.internic.net/rfc/rfc752.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc752.txt. Status: UNKNOWN. Not

online.

RFC0753

[743] J. Postel. RFC 753: Internet Message Protocol, March 1, 1979. URL

ftp://ftp.internic.net/rfc/rfc753.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc753.txt. Status: UNKNOWN. Not online.

RFC0754

[744] J. Postel. RFC 754: Out-of-net host addresses for mail, April

6, 1979. URL ftp://ftp.internic.net/rfc/rfc754.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc754.txt. Status: UNKNOWN. Not

online.

RFC0755

[745] J. Postel. RFC 755: Assigned numbers, May 3, 1979. URL ftp:/

/ftp.internic.net/rfc/rfc750.txt,   ftp://ftp.internic.net/
rfc/rfc755.txt, ftp://ftp.internic.net/rfc/rfc758.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc750.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc755.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc758.

txt. Obsoleted by RFC0758 [748]. Obsoletes RFC0750 [740]. Status: HISTORIC. Not online.

RFC0756

[746] J. R. Pickens, E. J. Feinler, and J. E. Mathis. RFC 756: NIC name server — a datagram-based information utility, July 1, 1979. URL ftp://ftp.

internic.net/rfc/rfc756.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc756.txt. Status: UNKNOWN. Not online.

RFC0757

[747] Debra P. Deutsch. RFC 757: Suggested solution to the naming, addressing, and delivery problem for ARPANET message systems, September

10, 1979. URL ftp://ftp.internic.net/rfc/rfc757.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc757.txt. Status: UNKNOWN. Not

online.

RFC0758

[748] J. Postel. RFC 758: Assigned numbers, August 1, 1979. URL ftp:/

/ftp.internic.net/rfc/rfc755.txt,   ftp://ftp.internic.net/
rfc/rfc758.txt, ftp://ftp.internic.net/rfc/rfc762.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc755.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc758.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc762.

txt. Obsoleted by RFC0762 [752]. Obsoletes RFC0755 [745]. Status: HISTORIC. Not online.

RFC0759

[749] J. Postel. RFC 759: Internet Message Protocol, August 1, 1980. URL

ftp://ftp.internic.net/rfc/rfc759.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc759.txt. Status: HISTORIC. Not online.

RFC0760

[750] J. Postel. RFC 760: DoD standard Internet Protocol, Jan-

uary 1, 1980. URL ftp://ftp.internic.net/rfc/ien123.txt,
 ftp://ftp.internic.net/rfc/rfc760.txt, ftp://ftp.internic.
net/rfc/rfc777.txt,   ftp://ftp.internic.net/rfc/rfc791.txt,
 ftp://ftp.math.utah.edu/pub/rfc/ien123.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc760.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc777.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc791.txt.

Obsoleted by RFC0791, RFC0777 [781, 767]. Obsoletes IEN123 [?]. Status: UNKNOWN. Not online.

RFC0761

[751] J. Postel. RFC 761: DoD standard transmission control protocol,

January 1, 1980. URL ftp://ftp.internic.net/rfc/rfc761.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc761.txt.  Status: UN-

KNOWN.

RFC0762

[752] J. Postel. RFC 762: Assigned numbers, January 1, 1980. URL ftp:/

/ftp.internic.net/rfc/rfc758.txt,   ftp://ftp.internic.net/
rfc/rfc762.txt, ftp://ftp.internic.net/rfc/rfc770.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc758.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc762.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc770.

txt. Obsoleted by RFC0770 [760]. Obsoletes RFC0758 [748]. Status: HISTORIC. Not online.

RFC0763

[753] M. D. Abrams. RFC 763: Role mailboxes, May 7, 1980. URL ftp://

ftp.internic.net/rfc/rfc763.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc763.txt. Status: UNKNOWN. Not online.

RFC0764

[754] J. Postel. RFC 764: Telnet Protocol specification, June 1, 1980. URL

ftp://ftp.internic.net/rfc/rfc764.txt,   ftp://ftp.internic.
net/rfc/rfc854.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc764.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc854.txt. Obsoleted

by RFC0854 [844]. Status: UNKNOWN. Not online.

RFC0765

[755] J. Postel. RFC 765: File transfer protocol specification, June 1,

1980. URL ftp://ftp.internic.net/rfc/rfc542.txt,   ftp://
ftp.internic.net/rfc/rfc765.txt, ftp://ftp.internic.net/rfc/
rfc959.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc542.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc765.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc959.txt. Obsoleted by RFC0959 [949]. Obsoletes RFC0542 [535]. Status: UNKNOWN. Not online.

RFC0766

[756] J. Postel. RFC 766: Internet Protocol Handbook: Table of contents, July

1, 1980. URL ftp://ftp.internic.net/rfc/rfc766.txt,   ftp://
ftp.internic.net/rfc/rfc774.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc766.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc774.

txt. Obsoleted by RFC0774 [764]. Status: UNKNOWN. Not online.

RFC0767

[757] J. Postel. RFC 767: Structured format for transmission of multi-media

documents, August 1, 1980. URL ftp://ftp.internic.net/rfc/
rfc767.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc767.txt. Sta-

tus: UNKNOWN. Not online.

RFC0768

[758] J. Postel. RFC 768: User datagram protocol, August 28, 1980. URL ftp:

//ftp.internic.net/rfc/rfc768.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc768.txt. Status: STANDARD. See also STD0006 [?].

RFC0769

[759] J. Postel. RFC 769: Rapicom 450 facsimile file format, September

26, 1980. URL ftp://ftp.internic.net/rfc/rfc769.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc769.txt. Status: UNKNOWN. Not

online.

RFC0770

[760] J. Postel. RFC 770: Assigned numbers, September 1, 1980. URL ftp:/

/ftp.internic.net/rfc/rfc762.txt,   ftp://ftp.internic.net/
rfc/rfc770.txt, ftp://ftp.internic.net/rfc/rfc776.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc762.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc770.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc776.

txt. Obsoleted by RFC0776 [766]. Obsoletes RFC0762 [752]. Status: HISTORIC. Not online.

RFC0771

[761] V. G. Cerf and J. Postel. RFC 771: Mail transition plan, September

1, 1980. URL ftp://ftp.internic.net/rfc/rfc771.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc771.txt. Status: UNKNOWN. Not

online.

RFC0772

[762] S. Sluizer and J. Postel. RFC 772: Mail transfer protocol, September

1, 1980. URL ftp://ftp.internic.net/rfc/rfc772.txt,   ftp://
ftp.internic.net/rfc/rfc780.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc772.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc780.

txt. Obsoleted by RFC0780 [770]. Status: UNKNOWN. Not online.

RFC0773

[763] V. G. Cerf. RFC 773: Comments on NCP/TCP mail service transition strategy, October 1, 1980. URL ftp://ftp.internic.net/rfc/

rfc773.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc773.txt. Sta-

tus: UNKNOWN. Not online.

RFC0774

[764] J. Postel. RFC 774: Internet Protocol Handbook: Table of

contents, October 1, 1980. URL ftp://ftp.internic.net/rfc/
rfc766.txt,   ftp://ftp.internic.net/rfc/rfc774.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc766.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc774.txt. Obsoletes RFC0766 [756]. Status: UN-KNOWN. Not online.

RFC0775

[765] D. Mankins, D. Franklin, and A. D. Owen. RFC 775: Directory oriented FTP commands, December 1, 1980. URL ftp://ftp.

internic.net/rfc/rfc775.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc775.txt. Status: UNKNOWN. Not online.

RFC0776

[766] J. Postel. RFC 776: Assigned numbers, January 1, 1981. URL ftp:/

/ftp.internic.net/rfc/rfc770.txt,   ftp://ftp.internic.net/
rfc/rfc776.txt, ftp://ftp.internic.net/rfc/rfc790.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc770.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc776.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc790.

txt. Obsoleted by RFC0790 [780]. Obsoletes RFC0770 [760]. Status: HISTORIC. Not online.

RFC0777

[767] J. Postel. RFC 777: Internet Control Message Protocol, April 1,

1981. URL ftp://ftp.internic.net/rfc/rfc760.txt,   ftp://
ftp.internic.net/rfc/rfc777.txt, ftp://ftp.internic.net/rfc/
rfc792.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc760.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc777.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc792.txt. Obsoleted by RFC0792 [782]. Obsoletes RFC0760 [750]. Status: UNKNOWN. Not online.

RFC0778

[768] D. L. Mills. RFC 778: DCNET Internet Clock Service, April

18, 1981. URL ftp://ftp.internic.net/rfc/rfc778.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc778.txt. Status: HISTORIC. Not

online.

RFC0779

[769] E. Killian. RFC 779: Telnet send-location option, April 1, 1981. URL

ftp://ftp.internic.net/rfc/rfc779.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc779.txt. Status: PROPOSED STANDARD. Not online.

RFC0780

[770] S. Sluizer and J. Postel. RFC 780: Mail transfer protocol,

May 1, 1981. URL ftp://ftp.internic.net/rfc/rfc772.txt,
 ftp://ftp.internic.net/rfc/rfc780.txt, ftp://ftp.internic.
net/rfc/rfc788.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc772.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc780.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc788.txt. Obsoleted by RFC0788

[778]. Obsoletes RFC0772 [762]. Status: UNKNOWN. Not online.

RFC0781

[771] Z. Su. RFC 781: Specification of the Internet Protocol (IP) timestamp

option, May 1, 1981. URL ftp://ftp.internic.net/rfc/rfc781.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc781.txt.  Status:

UNKNOWN. Not online.

RFC0782

[772] J. Nabielsky and A. P. Skelton. RFC 782: Virtual terminal management

model, January 1, 1981. URL ftp://ftp.internic.net/rfc/rfc782.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc782.txt.  Status:

UNKNOWN. Not online.

RFC0783

[773] K. R. Sollins. RFC 783: TFTP protocol (revision 2), June 1,

1981. URL ftp://ftp.internic.net/rfc/ien133.txt,   ftp://
ftp.internic.net/rfc/rfc783.txt,   ftp://ftp.math.utah.edu/
pub/rfc/ien133.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc783.

txt. Obsoletes IEN133 [?]. Status: UNKNOWN.

RFC0784

[774] S. Sluizer and J. Postel. RFC 784: Mail transfer protocol: ISI TOPS20 implementation, July 1, 1981. URL ftp://ftp.internic.net/

rfc/rfc1350.txt,   ftp://ftp.internic.net/rfc/rfc784.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1350.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc784.txt. Obsoleted by RFC1350 [1340]. Status: UNKNOWN. Not online.

RFC0785

[775] S. Sluizer and J. Postel. RFC 785: Mail transfer protocol: ISI TOPS20 file definitions, July 1, 1981. URL ftp://ftp.

internic.net/rfc/rfc785.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc785.txt. Status: UNKNOWN. Not online.

RFC0786

[776] S. Sluizer and J. Postel. RFC 786: Mail transfer protocol: ISI TOPS20 MTP-NIMAIL interface, July 1, 1981. URL ftp://ftp.

internic.net/rfc/rfc786.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc786.txt. Status: UNKNOWN. Not online.

RFC0787

[777] A. L. Chapin. RFC 787: Connectionless data transmission survey/

tutorial, July 1, 1981. URL ftp://ftp.internic.net/rfc/rfc787.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc787.txt.  Status:

UNKNOWN. Not online.

RFC0788

[778] J. Postel. RFC 788: Simple mail transfer protocol, Novem-

ber 1, 1981.  URL ftp://ftp.internic.net/rfc/rfc780.txt,
 ftp://ftp.internic.net/rfc/rfc788.txt, ftp://ftp.internic.
net/rfc/rfc821.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc780.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc788.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc821.txt. Obsoleted by RFC0821

[811]. Obsoletes RFC0780 [770]. Status: UNKNOWN. Not online.

RFC0789

[779] E. C. Rosen. RFC 789: Vulnerabilities of network control protocols: An

example, July 1, 1981. URL ftp://ftp.internic.net/rfc/rfc789.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc789.txt.  Status:

UNKNOWN. Not online.

RFC0790

[780] J. Postel. RFC 790: Assigned numbers, September 1, 1981. URL ftp:/

/ftp.internic.net/rfc/rfc776.txt,   ftp://ftp.internic.net/
rfc/rfc790.txt, ftp://ftp.internic.net/rfc/rfc820.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc776.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc790.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc820.

txt. Obsoleted by RFC0820 [810]. Obsoletes RFC0776 [766]. Status: HISTORIC.

RFC0791

[781] J. Postel. RFC 791: Internet Protocol, September 1, 1981. URL ftp:/

/ftp.internic.net/rfc/rfc760.txt,   ftp://ftp.internic.net/
rfc/rfc791.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc760.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc791.txt. Obsoletes RFC0760

[750]. See also STD0005 [?]. Status: STANDARD.

RFC0792

[782] J. Postel. RFC 792: Internet Control Message Protocol, Septem-

ber 1, 1981.  URL ftp://ftp.internic.net/rfc/rfc777.txt,
 ftp://ftp.internic.net/rfc/rfc792.txt, ftp://ftp.internic.
net/rfc/rfc950.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc777.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc792.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc950.txt. Obsoletes RFC0777 [767].

Updated by RFC0950 [940]. See also STD0005 [?]. Status: STANDARD.

RFC0793

[783] J. Postel. RFC 793: Transmission control protocol, September

1, 1981. URL ftp://ftp.internic.net/rfc/rfc793.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc793.txt. See also STD0007 [?]. Sta-

tus: STANDARD.

RFC0794

[784] V. G. Cerf. RFC 794: Pre-emption, September 1, 1981. URL ftp:/

/ftp.internic.net/rfc/ien125.txt,   ftp://ftp.internic.net/
rfc/rfc794.txt,   ftp://ftp.math.utah.edu/pub/rfc/ien125.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc794.txt. Updates IEN125

[?]. Status: UNKNOWN. Not online.

RFC0795

[785] J. Postel. RFC 795: Service mappings, September 1, 1981. URL ftp://

ftp.internic.net/rfc/rfc795.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc795.txt. Status: UNKNOWN. Not online.

RFC0796

[786] J. Postel. RFC 796: Address mappings, September 1, 1981. URL ftp:/

/ftp.internic.net/rfc/ien115.txt,   ftp://ftp.internic.net/
rfc/rfc796.txt,   ftp://ftp.math.utah.edu/pub/rfc/ien115.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc796.txt. Obsoletes IEN115

[?]. Status: UNKNOWN. Not online.

RFC0797

[787] A. R. Katz. RFC 797: Format for bitmap files, September 1,

1981. URL ftp://ftp.internic.net/rfc/rfc797.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc797.txt. Status: UNKNOWN. Not

online.

RFC0798

[788] A. R. Katz. RFC 798: Decoding facsimile data from the rapicom

450, September 1, 1981. URL ftp://ftp.internic.net/rfc/rfc798.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc798.txt.  Status:

UNKNOWN. Not online.

RFC0799

[789] D. L. Mills. RFC 799: Internet name domains, September 1,

1981. URL ftp://ftp.internic.net/rfc/rfc799.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc799.txt. Status: UNKNOWN. Not

online.

RFC0800

[790] J. Postel and J. Vernon. RFC 800: Request for comments summary notes: 700-799, November 1, 1982. URL ftp://ftp.

internic.net/rfc/rfc800.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc800.txt. Status: INFORMATIONAL.

RFC0801

[791] J. Postel. RFC 801: NCP/TCP transition plan, November 1,

1981. URL ftp://ftp.internic.net/rfc/rfc801.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc801.txt. Status: UNKNOWN.

RFC0802

[792] A. G. Malis. RFC 802: ARPANET 1822L host access protocol, November

1, 1981. URL ftp://ftp.internic.net/rfc/rfc802.txt,   ftp://
ftp.internic.net/rfc/rfc851.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc802.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc851.

txt. Obsoleted by RFC0851 [841]. Status: UNKNOWN.

RFC0803

[793] A. Agarwal, M. J. O’Connor, and D. L. Mills. RFC 803: Dacom 450/ 500 facsimile data transcoding, November 2, 1981. URL ftp://ftp.

internic.net/rfc/rfc803.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc803.txt. Status: UNKNOWN.

RFC0804

[794] International Telegraph and Telephone Consultative Committee. RFC 804: CCITT draft recommendation T.4, January 1, 1981. URL ftp://

ftp.internic.net/rfc/rfc804.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc804.txt. Status: UNKNOWN.

RFC0805

[795] J. Postel. RFC 805: Computer mail meeting notes, February 8,

1982. URL ftp://ftp.internic.net/rfc/rfc805.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc805.txt. Status: UNKNOWN.

RFC0806

[796] National Bureau of Standards. RFC 806: Proposed Federal Information Processing Standard: Specification for message format for computer based message systems, September 1, 1981. URL ftp:/

/ftp.internic.net/rfc/rfc806.txt,   ftp://ftp.internic.net/
rfc/rfc841.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc806.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc841.txt.  Obsoleted by

RFC0841 [831]. Status: UNKNOWN.

RFC0807

[797] J. Postel. RFC 807: Multimedia mail meeting notes, February

9, 1982. URL ftp://ftp.internic.net/rfc/rfc807.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc807.txt. Status: UNKNOWN.

RFC0808

[798] J. Postel. RFC 808: Summary of computer mail services meeting held at BBN on 10 January 1979, March 1, 1982. URL ftp://ftp.

internic.net/rfc/rfc808.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc808.txt. Status: UNKNOWN.

RFC0809

[799] T. Chang. RFC 809: UCL facsimile system, February 1, 1982. URL ftp:

//ftp.internic.net/rfc/rfc809.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc809.txt. Status: UNKNOWN.

RFC0810

[800] E. J. Feinler, K. Harrenstien, Z. Su, and V. White. RFC 810: DoD Internet host table specification, March 1, 1982. URL ftp:/

/ftp.internic.net/rfc/rfc608.txt,   ftp://ftp.internic.net/
rfc/rfc810.txt, ftp://ftp.internic.net/rfc/rfc952.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc608.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc810.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc952.

txt. Obsoleted by RFC0952 [942]. Obsoletes RFC0608 [601]. Status: UNKNOWN.

RFC0811

[801] K. Harrenstien, V. White, and E. J. Feinler. RFC 811: Host-

names server, March 1, 1982. URL ftp://ftp.internic.net/rfc/
rfc811.txt,   ftp://ftp.internic.net/rfc/rfc953.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc811.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc953.txt. Obsoleted by RFC0953 [943]. Status: UN-KNOWN.

RFC0812

[802] K. Harrenstien and V. White. RFC 812: NICNAME/WHOIS, March

1, 1982. URL ftp://ftp.internic.net/rfc/rfc812.txt,   ftp://
ftp.internic.net/rfc/rfc954.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc812.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc954.

txt. Obsoleted by RFC0954 [944]. Status: UNKNOWN.

RFC0813

[803] D. D. Clark. RFC 813: Window and acknowledgement strat-

egy in TCP, July 1, 1982. URL ftp://ftp.internic.net/rfc/
rfc813.txt,   ftp://ftp.internic.net/rfc/rfc954.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc813.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc954.txt. Status: UNKNOWN.

RFC0814

[804] D. D. Clark. RFC 814: Name, addresses, ports, and routes, July

1, 1982. URL ftp://ftp.internic.net/rfc/rfc814.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc814.txt. Status: UNKNOWN.

RFC0815

[805] D. D. Clark. RFC 815: IP datagram reassembly algorithms, July

1, 1982. URL ftp://ftp.internic.net/rfc/rfc815.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc815.txt. Status: UNKNOWN.

RFC0816

[806] D. D. Clark. RFC 816: Fault isolation and recovery, July 1,

1982. URL ftp://ftp.internic.net/rfc/rfc816.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc816.txt. Status: UNKNOWN.

RFC0817

[807] D. D. Clark. RFC 817: Modularity and efficiency in protocol imple-

mentation, July 1, 1982. URL ftp://ftp.internic.net/rfc/rfc817.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc817.txt.  Status:

UNKNOWN.

RFC0818

[808] J. Postel. RFC 818: Remote user Telnet service, November 1,

1982. URL ftp://ftp.internic.net/rfc/rfc818.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc818.txt. Status: HISTORIC.

RFC0819

[809] Z. Su and J. Postel. RFC 819: Domain naming convention for Internet user applications, August 1, 1982. URL ftp://ftp.

internic.net/rfc/rfc819.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc819.txt. Status: UNKNOWN.

RFC0820

[810] J. Postel. RFC 820: Assigned numbers, August 14, 1982. URL ftp:/

/ftp.internic.net/rfc/rfc790.txt,   ftp://ftp.internic.net/
rfc/rfc820.txt, ftp://ftp.internic.net/rfc/rfc870.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc790.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc820.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc870.

txt. Obsoleted by RFC0870 [860]. Obsoletes RFC0790 [780]. Status: HISTORIC.

RFC0821

[811] J. Postel. RFC 821: Simple mail transfer protocol, August 1,

1982. URL ftp://ftp.internic.net/rfc/rfc788.txt,   ftp://
ftp.internic.net/rfc/rfc821.txt, ftp://ftp.internic.net/rfc/
std10.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc788.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc821.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std10.txt. See also STD0010 [?]. Obsoletes RFC0788 [778]. Status: STANDARD.

RFC0822

[812] D. Crocker. RFC 822: Standard for the format of ARPA Internet text messages, August 13, 1982. URL ftp://ftp.internic.net/rfc/

rfc1123.txt,   ftp://ftp.internic.net/rfc/rfc1138.txt,   ftp:
//ftp.internic.net/rfc/rfc1148.txt,   ftp://ftp.internic.net/
rfc/rfc1327.txt,   ftp://ftp.internic.net/rfc/rfc2156.txt,
ftp://ftp.internic.net/rfc/rfc733.txt, ftp://ftp.internic.net/
rfc/rfc822.txt,   ftp://ftp.internic.net/rfc/std11.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1123.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1138.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1148.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1327.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2156.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc733.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc822.txt,   ftp://ftp.math.utah.edu/pub/rfc/std11.txt.

See also STD0011 [?]. Obsoletes RFC0733 [723]. Updated by RFC1123, RFC1138, RFC1148, RFC1327, RFC2156 [1113, 1128, 1138, 1317, 2143]. Status: STANDARD.

RFC0823

[813] R. M. Hinden and A. Sheltzer. RFC 823: DARPA Internet gateway,

September 1, 1982. URL ftp://ftp.internic.net/rfc/ien109.txt,
 ftp://ftp.internic.net/rfc/ien30.txt, ftp://ftp.internic.net/
rfc/rfc823.txt,   ftp://ftp.math.utah.edu/pub/rfc/ien109.txt,
 ftp://ftp.math.utah.edu/pub/rfc/ien30.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc823.txt. Updates IEN109, IEN30 [?, ?]. Status: HISTORIC.

RFC0824

[814] W. I. MacGregor and D. C. Tappan. RFC 824: CRONUS virtual local network, August 25, 1982. URL ftp://ftp.internic.net/rfc/

rfc824.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc824.txt. Sta-

tus: UNKNOWN.

RFC0825

[815] J. Postel. RFC 825: Request for comments on requests for comments,

November 1, 1982. URL ftp://ftp.internic.net/rfc/rfc1111.
txt,   ftp://ftp.internic.net/rfc/rfc1543.txt,   ftp://ftp.
internic.net/rfc/rfc825.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1111.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1543.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc825.txt. Obsoleted

by RFC1111, RFC1543 [1101, 1533]. Status: UNKNOWN.

RFC0826

[816] D. C. Plummer. RFC 826: Ethernet Address Resolution Protocol: Or converting network protocol addresses to 48.bit Ethernet address for transmission on Ethernet hardware, November 1,

1982. URL ftp://ftp.internic.net/rfc/rfc826.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc826.txt. Status: STANDARD.

RFC0827

[817] E. C. Rosen. RFC 827: Exterior Gateway Protocol (EGP), October

1, 1982. URL ftp://ftp.internic.net/rfc/rfc827.txt,   ftp://
ftp.internic.net/rfc/rfc904.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc827.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc904.

txt. Updated by RFC0904 [894]. Status: UNKNOWN.

RFC0828

[818] K. Owen. RFC 828: Data communications: IFIP’s international “network” of experts, August 1, 1982. URL ftp://ftp.

internic.net/rfc/rfc828.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc828.txt. Status: UNKNOWN.

RFC0829

[819] V. G. Cerf. RFC 829: Packet satellite technology reference sources,

November 1, 1982. URL ftp://ftp.internic.net/rfc/rfc829.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc829.txt.  Status: UN-

KNOWN.

RFC0830

[820] Z. Su. RFC 830: Distributed system for Internet name service, October

1, 1982. URL ftp://ftp.internic.net/rfc/rfc830.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc830.txt. Status: UNKNOWN. Not

online.

RFC0831

[821] R. T. Braden. RFC 831: Backup access to the European side of SAT-

NET, December 1, 1982. URL ftp://ftp.internic.net/rfc/rfc831.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc831.txt.  Status:

UNKNOWN. Not online.

RFC0832

[822] D. Smallberg. RFC 832: Who talks TCP?, December 7, 1982. URL ftp:/

/ftp.internic.net/rfc/rfc832.txt,   ftp://ftp.internic.net/
rfc/rfc833.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc832.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc833.txt.  Obsoleted by

RFC0833 [823]. Status: UNKNOWN. Not online.

RFC0833

[823] D. Smallberg. RFC 833: Who talks TCP?, December 14, 1982. URL ftp:

//ftp.internic.net/rfc/rfc832.txt,   ftp://ftp.internic.net/
rfc/rfc833.txt, ftp://ftp.internic.net/rfc/rfc834.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc832.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc833.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc834.

txt. Obsoleted by RFC0834 [824]. Obsoletes RFC0832 [822]. Status: UNKNOWN. Not online.

RFC0834

[824] D. Smallberg. RFC 834: Who talks TCP?, December 22, 1982. URL ftp:

//ftp.internic.net/rfc/rfc833.txt,   ftp://ftp.internic.net/
rfc/rfc834.txt, ftp://ftp.internic.net/rfc/rfc835.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc833.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc834.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc835.

txt. Obsoleted by RFC0835 [825]. Obsoletes RFC0833 [823]. Status: UNKNOWN. Not online.

RFC0835

[825] D. Smallberg. RFC 835: Who talks TCP?, December 29, 1982. URL ftp:

//ftp.internic.net/rfc/rfc834.txt,   ftp://ftp.internic.net/
rfc/rfc835.txt, ftp://ftp.internic.net/rfc/rfc836.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc834.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc835.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc836.

txt. Obsoleted by RFC0836 [826]. Obsoletes RFC0834 [824]. Status: UNKNOWN. Not online.

RFC0836

[826] D. Smallberg. RFC 836: Who talks TCP?, January 5, 1983. URL ftp:/

/ftp.internic.net/rfc/rfc835.txt,   ftp://ftp.internic.net/
rfc/rfc836.txt, ftp://ftp.internic.net/rfc/rfc837.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc835.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc836.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc837.

txt. Obsoleted by RFC0837 [827]. Obsoletes RFC0835 [825]. Status: UNKNOWN. Not online.

RFC0837

[827] D. Smallberg. RFC 837: Who talks TCP?, January 12, 1983. URL ftp:/

/ftp.internic.net/rfc/rfc836.txt,   ftp://ftp.internic.net/
rfc/rfc837.txt, ftp://ftp.internic.net/rfc/rfc838.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc836.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc837.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc838.

txt. Obsoleted by RFC0838 [828]. Obsoletes RFC0836 [826]. Status: UNKNOWN. Not online.

RFC0838

[828] D. Smallberg. RFC 838: Who talks TCP?, January 20, 1983. URL ftp:/

/ftp.internic.net/rfc/rfc837.txt,   ftp://ftp.internic.net/
rfc/rfc838.txt, ftp://ftp.internic.net/rfc/rfc839.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc837.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc838.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc839.

txt. Obsoleted by RFC0839 [829]. Obsoletes RFC0837 [827]. Status: UNKNOWN. Not online.

RFC0839

[829] D. Smallberg. RFC 839: Who talks TCP?, January 26, 1983. URL ftp:/

/ftp.internic.net/rfc/rfc838.txt,   ftp://ftp.internic.net/
rfc/rfc839.txt, ftp://ftp.internic.net/rfc/rfc842.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc838.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc839.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc842.

txt. Obsoleted by RFC0842 [832]. Obsoletes RFC0838 [828]. Status: UNKNOWN. Not online.

RFC0840

[830] J. Postel. RFC 840: Official protocols, April 13, 1983. URL ftp:/

/ftp.internic.net/rfc/rfc840.txt,   ftp://ftp.internic.net/
rfc/rfc880.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc840.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc880.txt.  Obsoleted by

RFC0880 [870]. Status: HISTORIC. Not online.

RFC0841

[831] National Bureau of Standards. RFC 841: Specification for message format for computer based message systems, January 27,

1983. URL ftp://ftp.internic.net/rfc/rfc806.txt,   ftp://
ftp.internic.net/rfc/rfc841.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc806.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc841.

txt. Obsoletes RFC0806 [796]. Status: UNKNOWN. Not online.

RFC0842

[832] D. Smallberg. RFC 842: Who talks TCP? — survey of 1 Febru-

ary 83, February 3, 1983. URL ftp://ftp.internic.net/rfc/
rfc839.txt,   ftp://ftp.internic.net/rfc/rfc842.txt, ftp://
ftp.internic.net/rfc/rfc843.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc839.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc842.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc843.txt. Obsoleted

by RFC0843 [833]. Obsoletes RFC0839 [829]. Status: UNKNOWN. Not online.

RFC0843

[833] D. Smallberg. RFC 843: Who talks TCP? — survey of 8 February 83,

February 9, 1983. URL ftp://ftp.internic.net/rfc/rfc842.txt,
 ftp://ftp.internic.net/rfc/rfc843.txt, ftp://ftp.internic.
net/rfc/rfc844.txt,   ftp://ftp.internic.net/rfc/rfc845.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc842.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc843.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc844.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc845.txt.

Obsoleted by RFC0845 [835]. Obsoletes RFC0842 [832]. Updated by RFC0844 [834]. Status: UNKNOWN. Not online.

RFC0844

[834] R. Clements. RFC 844: Who talks ICMP, too? — survey of 18 Febru-

ary 1983, February 18, 1983. URL ftp://ftp.internic.net/rfc/
rfc843.txt,   ftp://ftp.internic.net/rfc/rfc844.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc843.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc844.txt. Updates RFC0843 [833]. Status: UN-KNOWN. Not online.

RFC0845

[835] D. Smallberg. RFC 845: Who talks TCP? — survey of 15 Febru-

ary 1983, February 17, 1983. URL ftp://ftp.internic.net/rfc/
rfc843.txt,   ftp://ftp.internic.net/rfc/rfc845.txt, ftp://
ftp.internic.net/rfc/rfc846.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc843.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc845.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc846.txt. Obsoleted

by RFC0846 [836]. Obsoletes RFC0843 [833]. Status: UNKNOWN. Not online.

RFC0846

[836] D. Smallberg. RFC 846: Who talks TCP? — survey of 22 Febru-

ary 1983, February 23, 1983. URL ftp://ftp.internic.net/rfc/
rfc845.txt,   ftp://ftp.internic.net/rfc/rfc846.txt, ftp://
ftp.internic.net/rfc/rfc847.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc845.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc846.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc847.txt. Obsoleted

by RFC0847 [837]. Obsoletes RFC0845 [835]. Status: UNKNOWN. Not online.

RFC0847

[837] D. Smallberg, A. Westine, and J. Postel. RFC 847: Summary of Smallberg surveys, February 1, 1983. URL ftp://ftp.internic.net/rfc/

rfc846.txt,   ftp://ftp.internic.net/rfc/rfc847.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc846.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc847.txt. Obsoletes RFC0846 [836]. Status: UN-KNOWN. Not online.

RFC0848

[838] D. Smallberg. RFC 848: Who provides the “little” TCP services?, March

14, 1983. URL ftp://ftp.internic.net/rfc/rfc848.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc848.txt. Status: UNKNOWN. Not

online.

RFC0849

[839] M. R. Crispin. RFC 849: Suggestions for improved host table distri-

bution, May 1, 1983. URL ftp://ftp.internic.net/rfc/rfc849.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc849.txt.  Status:

UNKNOWN. Not online.

RFC0850

[840] M. R. Horton. RFC 850: Standard for interchange of USENET messages,

June 1, 1983. URL ftp://ftp.internic.net/rfc/rfc1036.txt,
ftp://ftp.internic.net/rfc/rfc850.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1036.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc850.txt. Obsoleted by RFC1036 [1026]. Status: UNKNOWN.

RFC0851

[841] A. G. Malis. RFC 851: ARPANET 1822L host access protocol,

April 18, 1983. URL ftp://ftp.internic.net/rfc/rfc802.txt,
 ftp://ftp.internic.net/rfc/rfc851.txt, ftp://ftp.internic.
net/rfc/rfc878.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc802.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc851.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc878.txt. Obsoleted by RFC0878

[868]. Obsoletes RFC0802 [792]. Status: UNKNOWN.

RFC0852

[842] A. G. Malis. RFC 852: ARPANET short blocking feature, April

1, 1983. URL ftp://ftp.internic.net/rfc/rfc852.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc852.txt. Status: UNKNOWN.

RFC0853

[843] ???? RFC 853, November 1993. URL ftp://ftp.internic.net/rfc/

rfc853.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc853.txt. Sta-

tus: UNKNOWN. Not online.

RFC0854

[844] J. Postel and J. K. Reynolds. RFC 854: Telnet Protocol specifica-

tion, May 1, 1983. URL ftp://ftp.internic.net/rfc/nic18639.txt,
 ftp://ftp.internic.net/rfc/rfc764.txt, ftp://ftp.internic.
net/rfc/rfc854.txt,   ftp://ftp.math.utah.edu/pub/rfc/nic18639.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc764.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc854.txt. Obsoletes RFC0764, NIC18639

[754, ?]. See also STD0008 [?]. Status: STANDARD.

RFC0855

[845] J. Postel and J. K. Reynolds. RFC 855: Telnet option specifications, May

1, 1983. URL ftp://ftp.internic.net/rfc/nic18640.txt,   ftp:/
/ftp.internic.net/rfc/rfc855.txt,   ftp://ftp.math.utah.edu/
pub/rfc/nic18640.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc855.

txt. Obsoletes NIC18640 [?]. Status: STANDARD.

RFC0856

[846] J. Postel and J. K. Reynolds. RFC 856: Telnet binary transmission, May

1, 1983. URL ftp://ftp.internic.net/rfc/nic15389.txt,   ftp:/
/ftp.internic.net/rfc/rfc856.txt,   ftp://ftp.math.utah.edu/
pub/rfc/nic15389.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc856.

txt. Obsoletes NIC15389 [?]. Status: UNKNOWN.

RFC0857

[847] J. Postel and J. K. Reynolds. RFC 857: Telnet echo option, May 1, 1983.

URL ftp://ftp.internic.net/rfc/nic15390.txt,   ftp://ftp.
internic.net/rfc/rfc857.txt,   ftp://ftp.math.utah.edu/pub/
rfc/nic15390.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc857.

txt. Obsoletes NIC15390 [?]. Status: UNKNOWN.

RFC0858

[848] J. Postel and J. K. Reynolds. RFC 858: Telnet Suppress Go Ahead op-

tion, May 1, 1983. URL ftp://ftp.internic.net/rfc/nic15392.txt,
 ftp://ftp.internic.net/rfc/rfc858.txt,   ftp://ftp.math.utah.
edu/pub/rfc/nic15392.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc858.txt. Obsoletes NIC15392 [?]. Status: UNKNOWN.

RFC0859

[849] J. Postel and J. K. Reynolds. RFC 859: Telnet status option, May

1, 1983. URL ftp://ftp.internic.net/rfc/rfc651.txt,   ftp://
ftp.internic.net/rfc/rfc859.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc651.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc859.

txt. Obsoletes RFC0651 [642]. Status: UNKNOWN.

RFC0860

[850] J. Postel and J. K. Reynolds. RFC 860: Telnet timing mark option, May

1, 1983. URL ftp://ftp.internic.net/rfc/nic16238.txt,   ftp:/
/ftp.internic.net/rfc/rfc860.txt,   ftp://ftp.math.utah.edu/
pub/rfc/nic16238.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc860.

txt. Obsoletes NIC16238 [?]. Status: UNKNOWN.

RFC0861

[851] J. Postel and J. K. Reynolds. RFC 861: Telnet extended options: List op-

tion, May 1, 1983. URL ftp://ftp.internic.net/rfc/nic16239.txt,
 ftp://ftp.internic.net/rfc/rfc861.txt,   ftp://ftp.math.utah.
edu/pub/rfc/nic16239.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc861.txt. Obsoletes NIC16239 [?]. Status: UNKNOWN.

RFC0862

[852] J. Postel. RFC 862: Echo protocol, May 1, 1983. URL ftp://ftp.

internic.net/rfc/rfc862.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc862.txt. See also STD0020 [?]. Status: STANDARD.

RFC0863

[853] J. Postel. RFC 863: Discard protocol, May 1, 1983. URL ftp://

ftp.internic.net/rfc/rfc863.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc863.txt. See also STD0021 [?]. Status: STANDARD.

RFC0864

[854] J. Postel. RFC 864: Character generator protocol, May 1, 1983. URL

ftp://ftp.internic.net/rfc/rfc864.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc864.txt. See also STD0022 [?]. Status: UNKNOWN.

RFC0865

[855] J. Postel. RFC 865: Quote of the day protocol, May 1, 1983. URL ftp://

ftp.internic.net/rfc/rfc865.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc865.txt. See also STD0023 [?]. Status: STANDARD.

RFC0866

[856] J. Postel. RFC 866: Active users, May 1, 1983. URL ftp://ftp.

internic.net/rfc/rfc866.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc866.txt. See also STD0024 [?]. Status: STANDARD.

RFC0867

[857] J. Postel. RFC 867: Daytime protocol, May 1, 1983. URL ftp://

ftp.internic.net/rfc/rfc867.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc867.txt. See also STD0025 [?]. Status: STANDARD.

RFC0868

[858] J. Postel and K. Harrenstien. RFC 868: Time protocol, May

1, 1983. URL ftp://ftp.internic.net/rfc/rfc868.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc868.txt. See also STD0026 [?]. Sta-

tus: STANDARD.

RFC0869

[859] R. M. Hinden. RFC 869: Host monitoring protocol, December

1, 1983. URL ftp://ftp.internic.net/rfc/rfc869.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc869.txt. Status: HISTORIC. For-

mat: TXT=94550 bytes.

RFC0870

[860] J. K. Reynolds and J. Postel. RFC 870: Assigned numbers, Oc-

tober 1, 1983. URL ftp://ftp.internic.net/rfc/rfc820.txt,
 ftp://ftp.internic.net/rfc/rfc870.txt, ftp://ftp.internic.
net/rfc/rfc900.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc820.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc870.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc900.txt. Obsoleted by RFC0900

[890]. Obsoletes RFC0820 [810]. Status: HISTORIC.

RFC0871

[861] M. A. Padlipsky. RFC 871: Perspective on the ARPANET reference model, September 1, 1982. URL ftp://ftp.internic.net/rfc/

rfc871.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc871.txt. Sta-

tus: UNKNOWN.

RFC0872

[862] M. A. Padlipsky. RFC 872: TCP-on-a-LAN, September 1, 1982. URL

ftp://ftp.internic.net/rfc/rfc872.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc872.txt. Status: UNKNOWN.

RFC0873

[863] M. A. Padlipsky. RFC 873: Illusion of vendor support, September

1, 1982. URL ftp://ftp.internic.net/rfc/rfc873.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc873.txt. Status: UNKNOWN.

RFC0874

[864] M. A. Padlipsky. RFC 874: Critique of X.25, September 1,

1982. URL ftp://ftp.internic.net/rfc/rfc874.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc874.txt. Status: UNKNOWN.

RFC0875

[865] M. A. Padlipsky. RFC 875: Gateways, architectures, and heffalumps,

September 1, 1982. URL ftp://ftp.internic.net/rfc/rfc875.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc875.txt.  Status: UN-

KNOWN.

RFC0876

[866] D. Smallberg. RFC 876: Survey of SMTP implementations, September

1, 1983. URL ftp://ftp.internic.net/rfc/rfc876.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc876.txt. Status: UNKNOWN.

RFC0877

[867] J. T. Korb. RFC 877: Standard for the transmission of IP datagrams over public data networks, September 1, 1983. URL ftp://ftp.

internic.net/rfc/rfc1356.txt,   ftp://ftp.internic.net/rfc/
rfc877.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1356.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc877.txt. Obsoleted by RFC1356

[1346]. Status: UNKNOWN.

RFC0878

[868] A. G. Malis. RFC 878: ARPANET 1822L host access protocol, December

1, 1983. URL ftp://ftp.internic.net/rfc/rfc851.txt,   ftp://
ftp.internic.net/rfc/rfc878.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc851.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc878.

txt. Obsoletes RFC0851 [841]. Status: UNKNOWN.

RFC0879

[869] J. Postel. RFC 879: TCP maximum segment size and related top-

ics, November 1, 1983. URL ftp://ftp.internic.net/rfc/rfc879.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc879.txt.  Status:

UNKNOWN.

RFC0880

[870] J. K. Reynolds and J. Postel. RFC 880: Official protocols, Oc-

tober 1, 1983. URL ftp://ftp.internic.net/rfc/rfc840.txt,
 ftp://ftp.internic.net/rfc/rfc880.txt, ftp://ftp.internic.
net/rfc/rfc901.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc840.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc880.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc901.txt. Obsoleted by RFC0901

[891]. Obsoletes RFC0840 [830]. Status: HISTORIC.

RFC0881

[871] J. Postel. RFC 881: Domain names plan and schedule, November

1, 1983. URL ftp://ftp.internic.net/rfc/rfc881.txt,   ftp://
ftp.internic.net/rfc/rfc897.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc881.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc897.

txt. Updated by RFC0897 [887]. Status: UNKNOWN.

RFC0882

[872] P. V. Mockapetris. RFC 882: Domain names: Concepts and facili-

ties, November 1, 1983. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.
internic.net/rfc/rfc882.txt, ftp://ftp.internic.net/rfc/rfc973.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc882.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc973.

txt. Obsoleted by RFC1034, RFC1035 [1024, 1025]. Updated by RFC0973 [963]. Status: UNKNOWN.

RFC0883

[873] P. V. Mockapetris. RFC 883: Domain names: Implementation specifica-

tion, November 1, 1983. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.
internic.net/rfc/rfc883.txt, ftp://ftp.internic.net/rfc/rfc973.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc883.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc973.

txt. Obsoleted by RFC1034, RFC1035 [1024, 1025]. Updated by RFC0973 [963]. Status: UNKNOWN.

RFC0884

[874] M. Solomon and E. Wimmers. RFC 884: Telnet terminal type

option, December 1, 1983. URL ftp://ftp.internic.net/rfc/
rfc884.txt,   ftp://ftp.internic.net/rfc/rfc930.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc884.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc930.txt. Obsoleted by RFC0930 [920]. Status: UN-KNOWN.

RFC0885

[875] J. Postel. RFC 885: Telnet end of record option, December

1, 1983. URL ftp://ftp.internic.net/rfc/rfc885.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc885.txt. Status: PROPOSED

STANDARD.

RFC0886

[876] M. T. Rose. RFC 886: Proposed standard for message header mung-

ing, December 15, 1983. URL ftp://ftp.internic.net/rfc/rfc886.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc886.txt.  Status:

UNKNOWN.

RFC0887

[877] M. Accetta. RFC 887: Resource Location Protocol, December

1, 1983. URL ftp://ftp.internic.net/rfc/rfc887.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc887.txt. Status: EXPERIMENTAL.

RFC0888

[878] L. Seamonson and E. C. Rosen. RFC 888: “STUB” Exterior Gateway Protocol, January 1, 1984. URL ftp://ftp.internic.net/rfc/

rfc888.txt,   ftp://ftp.internic.net/rfc/rfc904.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc888.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc904.txt. Updated by RFC0904 [894]. Status: UN-KNOWN.

RFC0889

[879] D. L. Mills. RFC 889: Internet delay experiments, December 1,

1983. URL ftp://ftp.internic.net/rfc/rfc889.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc889.txt. Status: UNKNOWN.

RFC0890

[880] J. Postel. RFC 890: Exterior Gateway Protocol implementation sched-

ule, February 1, 1984. URL ftp://ftp.internic.net/rfc/rfc890.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc890.txt.  Status:

UNKNOWN.

RFC0891

[881] D. L. Mills. RFC 891: DCN local-network protocols, December

1, 1983. URL ftp://ftp.internic.net/rfc/rfc891.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc891.txt. Status: STANDARD.

RFC0892

[882] International Organization for Standardization. RFC 892: ISO transport protocol specification, December 1, 1983. URL ftp:/

/ftp.internic.net/rfc/rfc892.txt,   ftp://ftp.internic.net/
rfc/rfc905.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc892.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc905.txt.  Obsoleted by

RFC0905 [895]. Status: UNKNOWN.

RFC0893

[883] S. Leffler and M. J. Karels. RFC 893: Trailer encapsulations, April

1, 1984. URL ftp://ftp.internic.net/rfc/rfc893.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc893.txt. Status: UNKNOWN.

RFC0894

[884] C. Hornig. RFC 894: Standard for the transmission of IP datagrams over Ethernet networks, April 1, 1984. URL ftp://ftp.

internic.net/rfc/rfc894.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc894.txt. Status: STANDARD.

RFC0895

[885] J. Postel. RFC 895: Standard for the transmission of IP datagrams over experimental Ethernet networks, April 1, 1984. URL ftp://ftp.

internic.net/rfc/rfc895.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc895.txt. Status: STANDARD.

RFC0896

[886] J. Nagle. RFC 896: Congestion control in IP/TCP internetworks,

January 6, 1984. URL ftp://ftp.internic.net/rfc/rfc896.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc896.txt.  Status: UN-

KNOWN.

RFC0897

[887] J. Postel. RFC 897: Domain name system implementation schedule,

February 1, 1984. URL ftp://ftp.internic.net/rfc/rfc881.txt,
 ftp://ftp.internic.net/rfc/rfc897.txt, ftp://ftp.internic.
net/rfc/rfc921.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc881.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc897.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc921.txt. Updates RFC0881 [871].

Updated by RFC0921 [911]. Status: UNKNOWN.

RFC0898

[888] R. M. Hinden, J. Postel, M. Muuss, and J. K. Reynolds. RFC 898: Gateway special interest group meeting notes, April 1, 1984. URL ftp://

ftp.internic.net/rfc/rfc898.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc898.txt. Status: UNKNOWN.

RFC0899

[889] J. Postel and A. Westine. RFC 899: Request for comments summary

notes: 800-899, May 1, 1984. URL ftp://ftp.internic.net/rfc/
rfc899.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc899.txt. Sta-

tus: INFORMATIONAL.

RFC0900

[890] J. K. Reynolds and J. Postel. RFC 900: Assigned numbers,

June 1, 1984. URL ftp://ftp.internic.net/rfc/rfc870.txt,
 ftp://ftp.internic.net/rfc/rfc900.txt, ftp://ftp.internic.
net/rfc/rfc923.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc870.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc900.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc923.txt. Obsoleted by RFC0923

[913]. Obsoletes RFC0870 [860]. Status: HISTORIC.

RFC0901

[891] J. K. Reynolds and J. Postel. RFC 901: Official ARPA-Internet

protocols, June 1, 1984.  URL ftp://ftp.internic.net/rfc/
rfc880.txt,   ftp://ftp.internic.net/rfc/rfc901.txt, ftp://
ftp.internic.net/rfc/rfc924.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc880.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc901.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc924.txt. Obsoleted

by RFC0924 [914]. Obsoletes RFC0880 [870]. Status: UNKNOWN.

RFC0902

[892] J. K. Reynolds and J. Postel. RFC 902: ARPA Internet Protocol

policy, July 1, 1984. URL ftp://ftp.internic.net/rfc/rfc902.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc902.txt.  Status:

UNKNOWN.

RFC0903

[893] R. Finlayson, T. Mann, J. C. Mogul, and M. Theimer. RFC 903: Reverse address resolution protocol, June 1, 1984. URL ftp://ftp.

internic.net/rfc/rfc903.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc903.txt. Status: STANDARD.

RFC0904

[894] D. L. Mills. RFC 904: Exterior Gateway Protocol formal specifica-

tion, April 1, 1984. URL ftp://ftp.internic.net/rfc/rfc827.txt,

 ftp://ftp.internic.net/rfc/rfc888.txt, ftp://ftp.internic.
net/rfc/rfc904.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc827.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc888.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc904.txt. Updates RFC0827, RFC0888

[817, 878]. Status: HISTORIC.

RFC0905

[895] A. M. McKenzie. RFC 905: ISO transport protocol specification

ISO DP 8073, April 1, 1984. URL ftp://ftp.internic.net/rfc/
rfc892.txt,   ftp://ftp.internic.net/rfc/rfc905.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc892.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc905.txt. Obsoletes RFC0892 [882]. Status: UN-KNOWN.

RFC0906

[896] R. Finlayson. RFC 906: Bootstrap loading using TFTP, June

1, 1984. URL ftp://ftp.internic.net/rfc/rfc906.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc906.txt. Status: UNKNOWN.

RFC0907

[897] Bolt Beranek and Newman, Inc. RFC 907: Host access protocol specification, July 1, 1984. URL ftp://ftp.internic.net/rfc/

rfc1221.txt,   ftp://ftp.internic.net/rfc/rfc907.txt, ftp:/
/ftp.internic.net/rfc/std40.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1221.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc907.
txt,   ftp://ftp.math.utah.edu/pub/rfc/std40.txt. Obsoleted by

STD0040 [?]. Updated by RFC1221 [1211]. Status: STANDARD.

RFC0908

[898] D. Velten, R. M. Hinden, and J. Sax. RFC 908: Reliable data protocol, July 1, 1984. URL ftp://ftp.internic.net/

rfc/rfc1151.txt,   ftp://ftp.internic.net/rfc/rfc908.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1151.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc908.txt. Updated by RFC1151 [1141]. Status: EXPERIMENTAL.

RFC0909

[899] C. Welles and W. Milliken. RFC 909: Loader debugger protocol, July

1, 1984. URL ftp://ftp.internic.net/rfc/rfc909.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc909.txt. Status: EXPERIMENTAL.

RFC0910

[900] H. C. Forsdick. RFC 910: Multimedia mail meeting notes, August

1, 1984. URL ftp://ftp.internic.net/rfc/rfc910.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc910.txt. Status: UNKNOWN.

RFC0911

[901] P. Kirton. RFC 911: EGP Gateway under Berkeley UNIX 4.2, August

22, 1984. URL ftp://ftp.internic.net/rfc/rfc911.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc911.txt. Status: UNKNOWN.

RFC0912

[902] M. St. Johns. RFC 912: Authentication service, September 1,

1984. URL ftp://ftp.internic.net/rfc/rfc912.txt,   ftp://
ftp.internic.net/rfc/rfc931.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc912.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc931.

txt. Obsoleted by RFC0931 [921]. Status: UNKNOWN.

RFC0913

[903] M. Lottor. RFC 913: Simple file transfer protocol, September

1, 1984. URL ftp://ftp.internic.net/rfc/rfc913.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc913.txt. Status: HISTORIC.

RFC0914

[904] D. J. Farber, G. Delp, and T. M. Conte. RFC 914: Thinwire protocol for connecting personal computers to the Internet, September

1, 1984. URL ftp://ftp.internic.net/rfc/rfc914.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc914.txt. Status: HISTORIC.

RFC0915

[905] M. A. Elvy and R. Nedved. RFC 915: Network mail path service, Decem-

ber 1, 1984. URL ftp://ftp.internic.net/rfc/rfc915.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc915.txt. Status: UNKNOWN.

RFC0916

[906] G. G. Finn. RFC 916: Reliable asynchronous transfer protocol (RATP),

October 1, 1984. URL ftp://ftp.internic.net/rfc/rfc916.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc916.txt. Status: HISTORIC.

RFC0917

[907] J. C. Mogul. RFC 917: Internet subnets, October 1, 1984. URL ftp://

ftp.internic.net/rfc/rfc917.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc917.txt. Status: UNKNOWN.

RFC0918

[908] J. K. Reynolds. RFC 918: Post office protocol, October 1,

1984. URL ftp://ftp.internic.net/rfc/rfc918.txt,   ftp://
ftp.internic.net/rfc/rfc937.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc918.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc937.

txt. Obsoleted by RFC0937 [927]. Status: UNKNOWN.

RFC0919

[909] J. C. Mogul. RFC 919: Broadcasting Internet datagrams, October

1, 1984. URL ftp://ftp.internic.net/rfc/rfc919.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc919.txt. See also STD0005 [?]. Sta-

tus: STANDARD.

RFC0920

[910] J. Postel and J. K. Reynolds. RFC 920: Domain requirements, October

1, 1984. URL ftp://ftp.internic.net/rfc/rfc920.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc920.txt. Status: UNKNOWN.

RFC0921

[911] J. Postel. RFC 921: Domain name system implementation schedule

— revised, October 1, 1984. URL ftp://ftp.internic.net/rfc/
rfc897.txt,   ftp://ftp.internic.net/rfc/rfc921.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc897.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc921.txt. Updates RFC0897 [887]. Status: UN-KNOWN.

RFC0922

[912] J. C. Mogul. RFC 922: Broadcasting Internet datagrams in the presence of subnets, October 1, 1984. URL ftp://ftp.internic.net/rfc/

rfc922.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc922.txt. See

also STD0005 [?]. Status: STANDARD.

RFC0923

[913] J. K. Reynolds and J. Postel. RFC 923: Assigned numbers, Oc-

tober 1, 1984. URL ftp://ftp.internic.net/rfc/rfc900.txt,
 ftp://ftp.internic.net/rfc/rfc923.txt, ftp://ftp.internic.
net/rfc/rfc943.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc900.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc923.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc943.txt. Obsoleted by RFC0943

[933]. Obsoletes RFC0900 [890]. Status: HISTORIC.

RFC0924

[914] J. K. Reynolds and J. Postel. RFC 924: Official ARPA-Internet protocols for connecting personal computers to the Internet, Oc-

tober 1, 1984. URL ftp://ftp.internic.net/rfc/rfc901.txt,
 ftp://ftp.internic.net/rfc/rfc924.txt, ftp://ftp.internic.
net/rfc/rfc944.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc901.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc924.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc944.txt. Obsoleted by RFC0944

[934]. Obsoletes RFC0901 [891]. Status: UNKNOWN.

RFC0925

[915] J. Postel. RFC 925: Multi-LAN address resolution, October 1,

1984. URL ftp://ftp.internic.net/rfc/rfc925.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc925.txt. Status: UNKNOWN.

RFC0926

[916] International Organization for Standardization. RFC 926: Protocol for providing the connectionless mode network services, December

1, 1984. URL ftp://ftp.internic.net/rfc/rfc926.txt,   ftp://
ftp.internic.net/rfc/rfc994.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc926.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc994.

txt. Obsoleted by RFC0994 [984]. Status: UNKNOWN.

RFC0927

[917] B. A. Anderson. RFC 927: TACACS user identification Telnet option,

December 1, 1984. URL ftp://ftp.internic.net/rfc/rfc927.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc927.txt. Status: PRO-

POSED STANDARD.

RFC0928

[918] M. A. Padlipsky. RFC 928: Introduction to proposed DoD standard

H-FP, December 1, 1984. URL ftp://ftp.internic.net/rfc/rfc928.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc928.txt.  Status:

UNKNOWN.

RFC0929

[919] J. Lilienkamp, R. Mandell, and M. A. Padlipsky. RFC 929: Proposed host-front end protocol, December 1, 1984. URL ftp://ftp.

internic.net/rfc/rfc929.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc929.txt. Status: HISTORIC.

RFC0930

[920] M. Solomon and E. Wimmers. RFC 930: Telnet terminal type option,

January 1, 1985. URL ftp://ftp.internic.net/rfc/rfc1091.txt,
 ftp://ftp.internic.net/rfc/rfc884.txt, ftp://ftp.internic.
net/rfc/rfc930.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1091.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc884.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc930.txt. Obsoleted by RFC1091

[1081]. Obsoletes RFC0884 [874]. Status: UNKNOWN.

RFC0931

[921] M. St. Johns. RFC 931: Authentication server, January 1,

1985. URL ftp://ftp.internic.net/rfc/rfc912.txt,   ftp://
ftp.internic.net/rfc/rfc931.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc912.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc931.

txt. Obsoleted RFC1413 [1403]. Obsoletes RFC0912 [902]. Status: UN-KNOWN.

RFC0932

[922] D. D. Clark. RFC 932: Subnetwork addressing scheme, January

1, 1985. URL ftp://ftp.internic.net/rfc/rfc932.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc932.txt. Status: UNKNOWN.

RFC0933

[923] S. Silverman. RFC 933: Output marking Telnet option, January

1, 1985. URL ftp://ftp.internic.net/rfc/rfc933.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc933.txt. Status: PROPOSED

STANDARD.

RFC0934

[924] M. T. Rose and E. A. Stefferud. RFC 934: Proposed standard for message encapsulation, January 1, 1985. URL ftp://ftp.

internic.net/rfc/rfc934.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc934.txt. Status: UNKNOWN.

RFC0935

[925] J. G. Robinson. RFC 935: Reliable link layer protocols, January

1, 1985. URL ftp://ftp.internic.net/rfc/rfc935.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc935.txt. Status: UNKNOWN.

RFC0936

[926] M. J. Karels. RFC 936: Another Internet subnet addressing scheme,

February 1, 1985. URL ftp://ftp.internic.net/rfc/rfc936.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc936.txt.  Status: UN-

KNOWN.

RFC0937

[927] M. Butler, J. Postel, D. Chase, J. Goldberger, and J. K. Reynolds. RFC 937: Post office protocol: Version 2, February 1, 1985. URL ftp:/

/ftp.internic.net/rfc/rfc918.txt,   ftp://ftp.internic.net/
rfc/rfc937.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc918.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc937.txt. Obsoletes RFC0918

[908]. Status: HISTORIC.

RFC0938

[928] T. Miller. RFC 938: Internet Reliable Transaction Protocol functional and interface specification, February 1, 1985. URL ftp://ftp.

internic.net/rfc/rfc938.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc938.txt. Status: EXPERIMENTAL.

RFC0939

[929] National Research Council. RFC 939: Executive summary of the NRC report on transport protocols for department of defense data net-

works, February 1, 1985. URL ftp://ftp.internic.net/rfc/rfc939.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc939.txt.  Status:

UNKNOWN.

RFC0940

[930] Gateway Algorithms and Data Structures Task Force. RFC 940: Toward an Internet standard scheme for subnetting, April 1, 1985. URL ftp://

ftp.internic.net/rfc/rfc940.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc940.txt. Status: UNKNOWN.

RFC0941

[931] International Organization for Standardization. RFC 941: Addendum to the network service definition covering network layer addressing, April

1, 1985. URL ftp://ftp.internic.net/rfc/rfc941.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc941.txt. Status: UNKNOWN.

RFC0942

[932] National Research Council. RFC 942: Transport protocols for Department of Defense data networks, February 1, 1985. URL ftp://ftp.

internic.net/rfc/rfc942.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc942.txt. Status: UNKNOWN.

RFC0943

[933] J. K. Reynolds and J. Postel. RFC 943: Assigned numbers,

April 1, 1985. URL ftp://ftp.internic.net/rfc/rfc923.txt,
 ftp://ftp.internic.net/rfc/rfc943.txt, ftp://ftp.internic.
net/rfc/rfc960.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc923.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc943.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc960.txt. Obsoleted by RFC0960

[950]. Obsoletes RFC0923 [913]. Status: HISTORIC.

RFC0944

[934] J. K. Reynolds and J. Postel. RFC 944: Official ARPA-Internet

protocols, April 1, 1985. URL ftp://ftp.internic.net/rfc/
rfc924.txt,   ftp://ftp.internic.net/rfc/rfc944.txt, ftp://
ftp.internic.net/rfc/rfc961.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc924.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc944.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc961.txt. Obsoleted

by RFC0961 [951]. Obsoletes RFC0924 [914]. Status: UNKNOWN.

RFC0945

[935] J. Postel. RFC 945: DoD statement on the NRC report, May 1,

1985. URL ftp://ftp.internic.net/rfc/rfc1039.txt,   ftp://
ftp.internic.net/rfc/rfc945.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1039.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc945.

txt. Obsoleted by RFC1039 [1029]. Status: UNKNOWN.

RFC0946

[936] R. Nedved. RFC 946: Telnet terminal location number option, May

1, 1985. URL ftp://ftp.internic.net/rfc/rfc946.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc946.txt. Status: PROPOSED

STANDARD.

RFC0947

[937] K. Lebowitz and D. Mankins. RFC 947: Multi-network broadcasting within the Internet, June 1, 1985. URL ftp://ftp.

internic.net/rfc/rfc947.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc947.txt. Status: UNKNOWN.

RFC0948

[938] I. Winston. RFC 948: Two methods for the transmission of IP datagrams over IEEE 802.3 networks, June 1, 1985. URL ftp://ftp.

internic.net/rfc/rfc1042.txt,   ftp://ftp.internic.net/rfc/
rfc948.txt, ftp://ftp.internic.net/rfc/std43.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1042.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc948.txt,   ftp://ftp.math.utah.edu/pub/rfc/std43.

txt. Obsoleted by RFC1042, STD0043 [1032, ?]. Status: UNKNOWN.

RFC0949

[939] M. A. Padlipsky. RFC 949: FTP unique-named store command, July

1, 1985. URL ftp://ftp.internic.net/rfc/rfc949.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc949.txt. Status: UNKNOWN.

RFC0950

[940] J. C. Mogul and J. Postel. RFC 950: Internet Standard Subnetting Procedure, August 1, 1985. URL ftp://ftp.internic.net/rfc/

rfc792.txt,   ftp://ftp.internic.net/rfc/rfc950.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc792.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc950.txt. Updates RFC0792 [782]. See also STD0005 [?]. Status: STANDARD.

RFC0951

[941] W. J. Croft and J. Gilmore. RFC 951: Bootstrap protocol,

September 1, 1985. URL ftp://ftp.internic.net/rfc/rfc1395.
txt,   ftp://ftp.internic.net/rfc/rfc1497.txt,   ftp://ftp.
internic.net/rfc/rfc1532.txt,   ftp://ftp.internic.net/rfc/
rfc1542.txt,   ftp://ftp.internic.net/rfc/rfc951.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1395.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1497.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1532.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1542.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc951.txt. Updated by RFC1395,

RFC1497, RFC1532, RFC1542 [1385, 1487, 1522, 1532]. Status: UN-KNOWN.

RFC0952

[942] K. Harrenstien, M. K. Stahl, and E. J. Feinler. RFC 952: DoD Internet host table specification, October 1, 1985. URL ftp:/

/ftp.internic.net/rfc/rfc810.txt,   ftp://ftp.internic.net/
rfc/rfc952.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc810.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc952.txt. Obsoletes RFC0810

[800]. Status: UNKNOWN.

RFC0953

[943] K. Harrenstien, M. K. Stahl, and E. J. Feinler. RFC 953: Hostname server, October 1, 1985. URL ftp://ftp.internic.net/rfc/

rfc811.txt,   ftp://ftp.internic.net/rfc/rfc953.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc811.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc953.txt. Obsoletes RFC0811 [801]. Status: HIS-TORIC.

RFC0954

[944] K. Harrenstien, M. K. Stahl, and E. J. Feinler. RFC 954: NIC-NAME/WHOIS, October 1, 1985. URL ftp://ftp.internic.net/rfc/

rfc812.txt,   ftp://ftp.internic.net/rfc/rfc813.txt, ftp://
ftp.internic.net/rfc/rfc954.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc812.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc813.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc954.txt. Obsoletes

RFC0812 [802]. Status: DRAFT STANDARD.

RFC0955

[945] R. T. Braden. RFC 955: Towards a transport service for transaction processing applications, September 1, 1985. URL ftp://ftp.

internic.net/rfc/rfc955.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc955.txt. Status: UNKNOWN.

RFC0956

[946] D. L. Mills. RFC 956: Algorithms for synchronizing network clocks,

September 1, 1985. URL ftp://ftp.internic.net/rfc/rfc956.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc956.txt.  Status: UN-

KNOWN.

RFC0957

[947] D. L. Mills. RFC 957: Experiments in network clock synchroniza-

tion, September 1, 1985. URL ftp://ftp.internic.net/rfc/rfc957.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc957.txt.  Status:

UNKNOWN.

RFC0958

[948] D. L. Mills. RFC 958: Network time protocol (NTP), September 1,

1985. URL ftp://ftp.internic.net/rfc/rfc1059.txt,   ftp://
ftp.internic.net/rfc/rfc1119.txt,   ftp://ftp.internic.net/
rfc/rfc1305.txt,   ftp://ftp.internic.net/rfc/rfc958.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1059.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1119.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1305.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc958.txt.

Obsoleted by RFC1059, RFC1119, RFC1305 [1049, 1109, 1295]. Status: UNKNOWN.

RFC0959

[949] J. Postel and J. K. Reynolds. RFC 959: File transfer protocol, Oc-

tober 1, 1985. URL ftp://ftp.internic.net/rfc/rfc2228.txt,

 ftp://ftp.internic.net/rfc/rfc765.txt, ftp://ftp.internic.
net/rfc/rfc959.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2228.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc765.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc959.txt. Obsoletes RFC0765 [755].

Updated by RFC2228 [2215]. Status: STANDARD.

RFC0960

[950] J. K. Reynolds and J. Postel. RFC 960: Assigned numbers, De-

cember 1, 1985. URL ftp://ftp.internic.net/rfc/rfc943.txt,
 ftp://ftp.internic.net/rfc/rfc960.txt, ftp://ftp.internic.
net/rfc/rfc990.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc943.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc960.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc990.txt. Obsoleted by RFC0990

[980]. Obsoletes RFC0943 [933]. Status: HISTORIC.

RFC0961

[951] J. K. Reynolds and J. Postel. RFC 961: Official ARPA-Internet protocols, December 1, 1985. URL ftp://ftp.internic.net/rfc/

rfc944.txt,   ftp://ftp.internic.net/rfc/rfc961.txt, ftp://
ftp.internic.net/rfc/rfc991.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc944.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc961.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc991.txt. Obsoleted

by RFC0991 [981]. Obsoletes RFC0944 [934]. Status: UNKNOWN.

RFC0962

[952] M. A. Padlipsky. RFC 962: TCP-4 prime, November 1, 1985. URL ftp:/

/ftp.internic.net/rfc/rfc962.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc962.txt. Status: UNKNOWN.

RFC0963

[953] D. P. Sidhu. RFC 963: Some problems with the specification of the Military Standard Internet Protocol, November 1, 1985. URL ftp://

ftp.internic.net/rfc/rfc963.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc963.txt. Status: UNKNOWN.

RFC0964

[954] D. P. Sidhu. RFC 964: Some problems with the specification of the military standard transmission control protocol, November

1, 1985. URL ftp://ftp.internic.net/rfc/rfc964.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc964.txt. Status: UNKNOWN.

RFC0965

[955] L. Aguilar. RFC 965: Format for a graphical communication proto-

col, December 1, 1985. URL ftp://ftp.internic.net/rfc/rfc965.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc965.txt.  Status:

UNKNOWN.

RFC0966

[956] S. E. Deering and D. R. Cheriton. RFC 966: Host groups: A multicast extension to the Internet Protocol, December 1, 1985. URL ftp:/

/ftp.internic.net/rfc/rfc966.txt,   ftp://ftp.internic.net/
rfc/rfc988.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc966.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc988.txt.  Obsoleted by

RFC0988 [978]. Status: UNKNOWN.

RFC0967

[957] M. A. Padlipsky. RFC 967: All victims together, December 1,

1985. URL ftp://ftp.internic.net/rfc/rfc967.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc967.txt. Status: UNKNOWN.

RFC0968

[958] V. G. Cerf. RFC 968: Twas the night before start-up, December

1, 1985. URL ftp://ftp.internic.net/rfc/rfc968.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc968.txt. Status: UNKNOWN.

RFC0969

[959] D. D. Clark, M. L. Lambert, and L. Zhang. RFC 969: NET-BLT: A bulk data transfer protocol, December 1, 1985. URL ftp:/

/ftp.internic.net/rfc/rfc969.txt,   ftp://ftp.internic.net/
rfc/rfc998.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc969.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc998.txt.  Obsoleted by

RFC0998 [988]. Status: UNKNOWN.

RFC0970

[960] J. Nagle. RFC 970: On packet switches with infinite storage, December

1, 1985. URL ftp://ftp.internic.net/rfc/rfc970.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc970.txt. Status: UNKNOWN.

RFC0971

[961] A. L. DeSchon. RFC 971: Survey of data representation standards,

January 1, 1986. URL ftp://ftp.internic.net/rfc/rfc971.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc971.txt.  Status: UN-

KNOWN.

RFC0972

[962] F. J. Wancho. RFC 972: Password generator protocol, January

1, 1986. URL ftp://ftp.internic.net/rfc/rfc972.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc972.txt. Status: UNKNOWN.

RFC0973

[963] P. V. Mockapetris. RFC 973: Domain system changes and observa-

tions, January 1, 1986. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.
internic.net/rfc/rfc882.txt, ftp://ftp.internic.net/rfc/rfc883.
txt,   ftp://ftp.internic.net/rfc/rfc973.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1034.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1035.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc882.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc883.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc973.txt. Obsoleted by RFC1034, RFC1035

[1024, 1025]. Updates RFC0882, RFC0883 [872, 873]. Status: UN-KNOWN.

RFC0974

[964] C. Partridge. RFC 974: Mail routing and the domain system, January

1, 1986. URL ftp://ftp.internic.net/rfc/rfc974.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc974.txt. See also STD0014 [?]. Sta-

tus: STANDARD.

RFC0975

[965] D. L. Mills. RFC 975: Autonomous confederations, February 1,

1986. URL ftp://ftp.internic.net/rfc/rfc975.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc975.txt. Status: UNKNOWN.

RFC0976

[966] M. R. Horton. RFC 976: UUCP mail interchange format standard,

February 1, 1986. URL ftp://ftp.internic.net/rfc/rfc1137.txt,
 ftp://ftp.internic.net/rfc/rfc976.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1137.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc976.txt. Updated by RFC1137 [1127]. Status: UNKNOWN.

RFC0977

[967] B. Kantor and P. Lapsley. RFC 977: Network news transfer protocol: A proposed standard for the stream-based transmission of news, Febru-

ary 1, 1986. URL ftp://ftp.internic.net/rfc/rfc977.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc977.txt. Status: PROPOSED

STANDARD.

RFC0978

[968] J. K. Reynolds, R. Gillman, W. A. Brackenridge, A. Witkowski, and J. Postel. RFC 978: Voice file interchange protocol (VFIP), February

1, 1986. URL ftp://ftp.internic.net/rfc/rfc978.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc978.txt. Status: UNKNOWN.

RFC0979

[969] A. G. Malis. RFC 979: PSN end-to-end functional specification, March

1, 1986. URL ftp://ftp.internic.net/rfc/rfc979.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc979.txt. Status: UNKNOWN.

RFC0980

[970] O. J. Jacobsen and J. Postel. RFC 980: Protocol document order infor-

mation, March 1, 1986. URL ftp://ftp.internic.net/rfc/rfc980.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc980.txt.  Status:

UNKNOWN.

RFC0981

[971] D. L. Mills. RFC 981: Experimental multiple-path routing algo-

rithm, March 1, 1986. URL ftp://ftp.internic.net/rfc/rfc981.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc981.txt.  Status:

UNKNOWN.

RFC0982

[972] H. W. Braun. RFC 982: Guidelines for the specification of the structure of the domain specific part (DSP) of the ISO standard NSAP

address, April 1, 1986. URL ftp://ftp.internic.net/rfc/rfc982.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc982.txt.  Status:

UNKNOWN.

RFC0983

[973] D. E. Cass and M. T. Rose. RFC 983: ISO transport arrives on top of the TCP, April 1, 1986. URL ftp://ftp.internic.net/

rfc/rfc1006.txt,   ftp://ftp.internic.net/rfc/rfc983.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1006.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc983.txt. Obsoleted by RFC1006 [996]. Status: UNKNOWN.

RFC0984

[974] D. D. Clark and M. L. Lambert. RFC 984: PCMAIL: A distributed mail system for personal computers, May 1, 1986. URL ftp:/

/ftp.internic.net/rfc/rfc984.txt,   ftp://ftp.internic.net/
rfc/rfc993.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc984.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc993.txt.  Obsoleted by

RFC0993 [983]. Status: UNKNOWN.

RFC0985

[975] National Science Foundation and Network Technical Advisory Group. RFC 985: Requirements for Internet gateways — draft, May 1,

1986. URL ftp://ftp.internic.net/rfc/rfc1009.txt,   ftp://
ftp.internic.net/rfc/rfc985.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1009.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc985.

txt. Obsoleted by RFC1009 [999]. Status: UNKNOWN.

RFC0986

[976] R. W. Callon and H. W. Braun. RFC 986: Guidelines for the use of Internet-IP addresses in the ISO Connectionless-Mode Network Protocol, June 1, 1986. URL ftp://ftp.internic.net/

rfc/rfc1069.txt,   ftp://ftp.internic.net/rfc/rfc986.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1069.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc986.txt. Obsoleted by RFC1069 [1059]. Status: UNKNOWN.

RFC0987

[977] S. E. Kille. RFC 987: Mapping between X.400 and RFC 822, June

1, 1986. URL ftp://ftp.internic.net/rfc/rfc1026.txt,   ftp://
ftp.internic.net/rfc/rfc1138.txt,   ftp://ftp.internic.net/
rfc/rfc1148.txt,   ftp://ftp.internic.net/rfc/rfc2156.txt,
ftp://ftp.internic.net/rfc/rfc987.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1026.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1138.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1148.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2156.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc987.txt. Obsoleted by RFC2156 [2143]. Updated by RFC1026, RFC1138, RFC1148 [1016, 1128, 1138]. Status: UN-KNOWN.

RFC0988

[978] S. E. Deering. RFC 988: Host extensions for IP multicasting, July

1, 1986. URL ftp://ftp.internic.net/rfc/rfc1054.txt,   ftp://
ftp.internic.net/rfc/rfc1112.txt,   ftp://ftp.internic.net/
rfc/rfc966.txt, ftp://ftp.internic.net/rfc/rfc988.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1054.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1112.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc966.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc988.txt. Ob-

soleted by RFC1054, RFC1112 [1044, 1102]. Obsoletes RFC0966 [956]. Status: UNKNOWN.

RFC0989

[979] J. Linn. RFC 989: Privacy enhancement for Internet electronic mail: Part I: Message encipherment and authentication procedures,

February 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1040.
txt,   ftp://ftp.internic.net/rfc/rfc1113.txt,   ftp://ftp.
internic.net/rfc/rfc989.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1040.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1113.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc989.txt. Obsoleted

by RFC1040, RFC1113 [1030, 1103]. Status: UNKNOWN.

RFC0990

[980] J. K. Reynolds and J. Postel. RFC 990: Assigned numbers, Novem-

ber 1, 1986. URL ftp://ftp.internic.net/rfc/rfc1010.txt,
 ftp://ftp.internic.net/rfc/rfc960.txt, ftp://ftp.internic.
net/rfc/rfc990.txt,   ftp://ftp.internic.net/rfc/rfc997.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1010.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc960.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc990.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc997.txt.

Obsoleted by RFC1010 [1000]. Obsoletes RFC0960 [950]. Updated by RFC0997 [987]. Status: HISTORIC.

RFC0991

[981] J. K. Reynolds and J. Postel. RFC 991: Official ARPA-Internet protocols, November 1, 1986. URL ftp://ftp.internic.net/rfc/

rfc1011.txt,   ftp://ftp.internic.net/rfc/rfc961.txt, ftp:/
/ftp.internic.net/rfc/rfc991.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1011.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc961.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc991.txt. Obsoleted

by RFC1011 [1001]. Obsoletes RFC0961 [951]. Status: UNKNOWN.

RFC0992

[982] K. P. Birman and T. A. Joseph. RFC 992: On communication support for fault tolerant process groups, November 1, 1986. URL ftp://ftp.

internic.net/rfc/rfc992.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc992.txt. Status: UNKNOWN.

RFC0993

[983] D. D. Clark and M. L. Lambert. RFC 993: PCMAIL: A distributed mail system for personal computers, December 1, 1986. URL ftp://ftp.

internic.net/rfc/rfc1056.txt,   ftp://ftp.internic.net/rfc/
rfc984.txt, ftp://ftp.internic.net/rfc/rfc993.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1056.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc984.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc993.

txt. Obsoleted by RFC1056 [1046]. Obsoletes RFC0984 [974]. Status: UNKNOWN.

RFC0994

[984] International Organization for Standardization. RFC 994: Final text of DIS 8473, protocol for providing the connectionless-mode net-

work service, March 1, 1986. URL ftp://ftp.internic.net/rfc/
rfc926.txt,   ftp://ftp.internic.net/rfc/rfc994.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc926.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc994.txt. Obsoletes RFC0926 [916]. Status: UN-KNOWN.

RFC0995

[985] International Organization for Standardization. RFC 995: End system to intermediate system routing exchange protocol for use in conjunction with ISO 8473, April 1, 1986. URL ftp://ftp.

internic.net/rfc/rfc995.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc995.txt. Status: UNKNOWN.

RFC0996

[986] D. L. Mills. RFC 996: Statistics server, February 1, 1987. URL ftp://

ftp.internic.net/rfc/rfc996.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc996.txt. Status: HISTORIC.

RFC0997

[987] J. K. Reynolds and J. Postel. RFC 997: Internet numbers, March 1,

1987. URL ftp://ftp.internic.net/rfc/rfc1020.txt,   ftp://
ftp.internic.net/rfc/rfc1117.txt,   ftp://ftp.internic.net/
rfc/rfc990.txt, ftp://ftp.internic.net/rfc/rfc997.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1020.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1117.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc990.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc997.txt. Ob-

soleted by RFC1020, RFC1117 [1010, 1107]. Updates RFC0990 [980]. Status: UNKNOWN.

RFC0998

[988] D. D. Clark, M. L. Lambert, and L. Zhang. RFC 998: NET-BLT: A bulk data transfer protocol, March 1, 1987. URL ftp:/

/ftp.internic.net/rfc/rfc969.txt,   ftp://ftp.internic.net/
rfc/rfc998.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc969.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc998.txt. Obsoletes RFC0969

[959]. Status: EXPERIMENTAL.

RFC0999

[989] A. Westine and J. Postel. RFC 999: Requests for comments summary notes: 900-999, April 1, 1987. URL ftp://ftp.internic.net/rfc/

rfc1000.txt,   ftp://ftp.internic.net/rfc/rfc160.txt, ftp:/
/ftp.internic.net/rfc/rfc999.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1000.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc160.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc999.txt. Obsoleted

by RFC1000 [990]. Obsoletes RFC0160 [160]. Status: UNKNOWN.

RFC1000

[990] J. K. Reynolds and J. Postel. RFC 1000: Request for comments reference guide, August 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1000.txt,   ftp://ftp.internic.net/rfc/rfc999.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1000.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc999.txt. Obsoletes RFC0999 [989]. Status: UN-KNOWN.

RFC1001

[991] NetBIOS Working Group. RFC 1001: Protocol standard for a net-BIOS service on a TCP/UDP transport: Concepts and methods, March

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1001.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1001.txt. See also STD0019 [?]. Sta-

tus: STANDARD.

RFC1002

[992] NetBIOS Working Group. RFC 1002: Protocol standard for a net-BIOS service on a TCP/UDP transport: Detailed specifications, March

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1002.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1002.txt. Status: STANDARD.

RFC1003

[993] A. R. Katz. RFC 1003: Issues in defining an equations representation standard, March 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1003.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1003.

txt. Status: UNKNOWN.

RFC1004

[994] D. L. Mills. RFC 1004: Distributed-protocol authentication scheme,

April 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1004.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1004.txt. Status: EXPERI-

MENTAL.

RFC1005

[995] A. Khanna and A. G. Malis. RFC 1005: ARPANET AHIP-E host access protocol (enhanced AHIP), May 1, 1987. URL ftp://ftp.internic.

net/rfc/rfc1005.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1005.

txt. Status: UNKNOWN.

RFC1006

[996] M. T. Rose and D. E. Cass. RFC 1006: ISO transport services on top of the TCP: Version 3, May 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1006.txt,   ftp://ftp.internic.net/rfc/rfc983.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1006.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc983.txt. Obsoletes RFC0983 [973]. Updated by RFC 2126 [2113]. Status: STANDARD.

RFC1007

[997] W. McCoy. RFC 1007: Military supplement to the ISO transport pro-

tocol, June 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1007.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1007.txt. Status:

UNKNOWN.

RFC1008

[998] W. McCoy. RFC 1008: Implementation guide for the ISO transport

protocol, June 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1008.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1008.txt. Status:

UNKNOWN.

RFC1009

[999] R. T. Braden and J. Postel. RFC 1009: Requirements for Internet

gateways, June 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1009.
txt,   ftp://ftp.internic.net/rfc/rfc1812.txt,   ftp://ftp.
internic.net/rfc/rfc985.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1009.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1812.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc985.txt. Obsoleted

by RFC1812 [1802]. Obsoletes RFC0985 [975]. Status: HISTORIC.

RFC1010

[1000] J. K. Reynolds and J. Postel. RFC 1010: Assigned numbers, May 1,

1987. URL ftp://ftp.internic.net/rfc/rfc1010.txt,   ftp://
ftp.internic.net/rfc/rfc1060.txt,   ftp://ftp.internic.net/
rfc/rfc990.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1010.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1060.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc990.txt. Obsoleted by RFC1060 [1050]. Obsoletes RFC0990 [980]. Status: HISTORIC.

RFC1011

[1001] J. K. Reynolds and J. Postel. RFC 1011: Official Internet protocols, May

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1011.txt,   ftp://
ftp.internic.net/rfc/rfc991.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1011.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc991.

txt. Obsoletes RFC0991 [981]. Status: UNKNOWN.

RFC1012

[1002] J. K. Reynolds and J. Postel. RFC 1012: Bibliography of request for comments 1 through 999, June 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1012.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1012.

txt. Status: INFORMATIONAL.

RFC1013

[1003] R. W. Scheifler. RFC 1013: X Window System Protocol, version 11: Alpha update April 1987, June 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1013.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1013.

txt. Status: UNKNOWN.

RFC1014

[1004] Sun Microsystems, Inc. RFC 1014: XDR: External Data Representation

standard, June 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1014.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1014.txt. Status:

UNKNOWN.

RFC1015

[1005] B. M. Leiner. RFC 1015: Implementation plan for interagency research

Internet, July 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1015.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1015.txt. Status:

UNKNOWN.

RFC1016

[1006] W. Prue and J. Postel. RFC 1016: Something a host could do with source quench: The Source Quench Introduced Delay (SQuID), July

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1016.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1016.txt. Status: UNKNOWN.

RFC1017

[1007] B. M. Leiner. RFC 1017: Network requirements for scientific research: Internet task force on scientific computing, August 1,

1987. URL ftp://ftp.internic.net/rfc/rfc1017.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1017.txt. Status: UNKNOWN.

RFC1018

[1008] A. M. McKenzie. RFC 1018: Some comments on SQuID, August

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1018.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1018.txt. Status: UNKNOWN.

RFC1019

[1009] D. Arnon. RFC 1019: Report of the workshop on environments for computational mathematics, September 1, 1987. URL ftp://ftp.

internic.net/rfc/rfc1019.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1019.txt. Status: UNKNOWN.

RFC1020

[1010] S. Romano and M. K. Stahl. RFC 1020: Internet numbers,

November 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1020.
txt,   ftp://ftp.internic.net/rfc/rfc1062.txt,   ftp://ftp.
internic.net/rfc/rfc1117.txt,   ftp://ftp.internic.net/rfc/
rfc1166.txt,   ftp://ftp.internic.net/rfc/rfc997.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1020.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1062.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1117.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1166.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc997.txt.  Obsoleted by

RFC1062, RFC1117, RFC1166 [1052, 1107, 1156]. Obsoletes RFC0997 [987]. Status: UNKNOWN.

RFC1021

[1011] C. Partridge and G. Trewitt. RFC 1021: High-level entity management system (HEMS), October 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1021.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1021.

txt. Status: HISTORIC.

RFC1022

[1012] C. Partridge and G. Trewitt. RFC 1022: High-level entity management protocol (HEMP), October 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1022.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1022.

txt. Status: UNKNOWN.

RFC1023

[1013] G. Trewitt and C. Partridge. RFC 1023: HEMS monitoring and control language, October 1, 1987. URL ftp://ftp.internic.net/rfc/

rfc1023.txt,   ftp://ftp.internic.net/rfc/rfc1076.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1023.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1076.txt. Obsoleted by RFC1076 [1066]. Status: UN-KNOWN.

RFC1024

[1014] C. Partridge and G. Trewitt. RFC 1024: HEMS variable defini-

tions, October 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1024.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1024.txt. Status:

UNKNOWN.

RFC1025

[1015] J. Postel. RFC 1025: TCP and IP bake off, September 1,

1987. URL ftp://ftp.internic.net/rfc/rfc1025.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1025.txt. Status: UNKNOWN.

RFC1026

[1016] S. E. Kille. RFC 1026: Addendum to RFC 987: (mapping between X.400 and RFC-822), September 1, 1987. URL ftp://ftp.

internic.net/rfc/rfc1026.txt,   ftp://ftp.internic.net/rfc/
rfc1138.txt,   ftp://ftp.internic.net/rfc/rfc1148.txt,   ftp:
//ftp.internic.net/rfc/rfc1327.txt,   ftp://ftp.internic.net/
rfc/rfc1495.txt,   ftp://ftp.internic.net/rfc/rfc2156.txt,
ftp://ftp.internic.net/rfc/rfc987.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1026.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1138.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1148.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1327.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1495.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2156.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc987.txt.

Obsoleted by RFC1327, RFC1495, RFC2156 [1317, 1485, 2143]. Updates RFC0987 [977]. Updated by RFC1138, RFC1148 [1128, 1138]. Status: UNKNOWN.

RFC1027

[1017] S. Carl-Mitchell and J. S. Quarterman. RFC 1027: Using ARP to implement transparent subnet gateways, October 1, 1987. URL ftp://ftp.

internic.net/rfc/rfc1027.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1027.txt. Status: UNKNOWN.

RFC1028

[1018] J. Davin, J. D. Case, M. Fedor, and M. L. Schoffstall. RFC 1028: Simple Gateway Monitoring Protocol, November 1, 1987. URL ftp://ftp.

internic.net/rfc/rfc1028.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1028.txt. Status: HISTORIC.

RFC1029

[1019] G. Parr. RFC 1029: More fault tolerant approach to address resolution for a multi-LAN system of ethernets, May 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1029.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1029.txt. Status: UNKNOWN.

RFC1030

[1020] M. L. Lambert. RFC 1030: On testing the NETBLT protocol over divers networks, November 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1030.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1030.

txt. Status: UNKNOWN.

RFC1031

[1021] W. D. Lazear. RFC 1031: MILNET name domain transition, November

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1031.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1031.txt. Status: UNKNOWN.

RFC1032

[1022] M. K. Stahl. RFC 1032: Domain administrators guide, November

1, 1987. URL ftp://ftp.internic.net/rfc/rfc1032.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1032.txt. Status: UNKNOWN.

RFC1033

[1023] M. Lottor. RFC 1033: Domain administrators operations guide,

November 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1033.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1033.txt. Status:

UNKNOWN.

RFC1034

[1024] P. V. Mockapetris. RFC 1034: Domain names — concepts and facili-

ties, November 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1065.txt,   ftp://ftp.
internic.net/rfc/rfc1101.txt,   ftp://ftp.internic.net/rfc/
rfc1183.txt,   ftp://ftp.internic.net/rfc/rfc1348.txt,   ftp:
//ftp.internic.net/rfc/rfc1876.txt,   ftp://ftp.internic.net/
rfc/rfc1982.txt,   ftp://ftp.internic.net/rfc/rfc2065.txt,
ftp://ftp.internic.net/rfc/rfc2181.txt,   ftp://ftp.internic.
net/rfc/rfc2308.txt,   ftp://ftp.internic.net/rfc/rfc882.txt,
ftp://ftp.internic.net/rfc/rfc883.txt,   ftp://ftp.internic.
net/rfc/rfc973.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1065.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1101.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1183.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1348.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1876.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1982.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2065.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2181.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2308.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc882.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc883.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc973.txt. Obsoletes RFC0973, RFC0882, RFC0883 [963, 872, 873]. See also STD0013 [?]. Updated by RFC1101, RFC1183, RFC1348, RFC1876, RFC1982, RFC2065, RFC2181, RFC2308 [1091, 1173, 1338, 1863, 1969, 2052, 2168, 2295]. Status: STANDARD.

RFC1035

[1025] P. V. Mockapetris. RFC 1035: Domain names — implementation and specification, November 1, 1987. URL ftp://ftp.

internic.net/rfc/rfc1035.txt,   ftp://ftp.internic.net/rfc/
rfc1101.txt,   ftp://ftp.internic.net/rfc/rfc1183.txt,   ftp:
//ftp.internic.net/rfc/rfc1348.txt,   ftp://ftp.internic.net/
rfc/rfc1876.txt,   ftp://ftp.internic.net/rfc/rfc1982.txt,
ftp://ftp.internic.net/rfc/rfc1995.txt,   ftp://ftp.internic.
net/rfc/rfc1996.txt,   ftp://ftp.internic.net/rfc/rfc2065.txt,
 ftp://ftp.internic.net/rfc/rfc2136.txt,   ftp://ftp.internic.
net/rfc/rfc2137.txt,   ftp://ftp.internic.net/rfc/rfc2181.txt,
 ftp://ftp.internic.net/rfc/rfc2308.txt,   ftp://ftp.internic.
net/rfc/rfc882.txt, ftp://ftp.internic.net/rfc/rfc883.txt,
ftp://ftp.internic.net/rfc/rfc973.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1101.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1183.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1348.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1876.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1982.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1995.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1996.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2065.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2136.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2137.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2181.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2308.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc882.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc883.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc973.txt.

Obsoletes RFC0973, RFC0882, RFC0883 [963, 872, 873]. See also STD0013 [?]. Updated by RFC1101, RFC1183, RFC1348, RFC1876, RFC1982, RFC1995, RFC1996, RFC2065, RFC2181, RFC2136, RFC2137, RFC2308 [1091, 1173, 1338, 1863, 1969, 1982, 1983, 2052, 2168, 2123, 2124, 2295]. Status: STANDARD.

RFC1036

[1026] M. R. Horton and R. Adams. RFC 1036: Standard for interchange of USENET messages, December 1, 1987. URL ftp://ftp.internic.net/

rfc/rfc1036.txt,   ftp://ftp.internic.net/rfc/rfc850.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1036.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc850.txt. Obsoletes RFC0850 [840]. Status: UN-KNOWN.

RFC1037

[1027] B. Greenberg and S. Keene. RFC 1037: NFILE — a file access protocol,

December 1, 1987. URL ftp://ftp.internic.net/rfc/rfc1037.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1037.txt. Status: HIS-

TORIC.

RFC1038

[1028] M. St. Johns. RFC 1038: Draft revised IP security option,

January 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1038.
txt,   ftp://ftp.internic.net/rfc/rfc1108.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1038.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1108.txt. Obsoleted by RFC1108 [1098]. Status: UN-KNOWN.

RFC1039

[1029] D. Latham. RFC 1039: DoD statement on open systems interconnection protocols, January 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1039.txt,   ftp://ftp.internic.net/rfc/rfc945.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1039.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc945.txt. Obsoletes RFC0945 [935]. Status: UN-KNOWN.

RFC1040

[1030] J. Linn. RFC 1040: Privacy enhancement for Internet electronic mail: Part I: Message encipherment and authentication procedures,

January 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1040.
txt,   ftp://ftp.internic.net/rfc/rfc1113.txt,   ftp://ftp.
internic.net/rfc/rfc989.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1040.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1113.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc989.txt. Obsoleted

by RFC1113 [1103]. Obsoletes RFC0989 [979]. Status: UNKNOWN.

RFC1041

[1031] Y. Rekhter. RFC 1041: Telnet 3270 regime option, January 1,

1988. URL ftp://ftp.internic.net/rfc/rfc1041.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1041.txt. Status: PROPOSED

STANDARD.

RFC1042

[1032] J. Postel and J. K. Reynolds. RFC 1042: Standard for the transmission of IP datagrams over IEEE 802 networks, February 1,

1988. URL ftp://ftp.internic.net/rfc/rfc1042.txt,   ftp://
ftp.internic.net/rfc/rfc948.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1042.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc948.

txt. Obsoletes RFC0948 [938]. Status: STANDARD.

RFC1043

[1033] A. Yasuda and T. Thompson. RFC 1043: Telnet Data Entry Terminal option: DODIIS implementation, February 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1043.txt,   ftp://ftp.internic.net/rfc/
rfc732.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1043.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc732.txt. Updates RFC0732

[722]. Status: PROPOSED STANDARD.

RFC1044

[1034] K. Hardwick and J. Lekashman. RFC 1044: Internet Protocol on Network System’s HYPERchannel: Protocol specification, February

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1044.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1044.txt. Status: STANDARD.

RFC1045

[1035] D. R. Cheriton. RFC 1045: VMTP: Versatile message transaction protocol: Protocol specification, February 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1045.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1045.txt. Status: EXPERIMENTAL.

RFC1046

[1036] W. Prue and J. Postel. RFC 1046: Queuing algorithm to provide type-of-service for IP links, February 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1046.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1046.

txt. Status: UNKNOWN.

RFC1047

[1037] C. Partridge. RFC 1047: Duplicate messages and SMTP, February

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1047.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1047.txt. Status: UNKNOWN.

RFC1048

[1038] P. A. Prindeville. RFC 1048: BOOTP vendor information exten-

sions, February 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1048.

txt,   ftp://ftp.internic.net/rfc/rfc1084.txt,   ftp://ftp.
internic.net/rfc/rfc1395.txt,   ftp://ftp.internic.net/rfc/
rfc1497.txt,   ftp://ftp.internic.net/rfc/rfc1533.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1048.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1084.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1395.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1497.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1533.txt. Obsoleted by

RFC1084, RFC1395, RFC1497, RFC1533 [1074, 1385, 1487, 1523]. Status: UNKNOWN.

RFC1049

[1039] M. A. Sirbu. RFC 1049: Content-type header field for Internet messages,

March 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1049.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1049.txt. See also STD0011

[?]. Status: HISTORIC.

RFC1050

[1040] Sun Microsystems, Inc. RFC 1050: RPC: Remote procedure call protocol specification, April 1, 1988. URL ftp://ftp.internic.net/rfc/

rfc1050.txt,   ftp://ftp.internic.net/rfc/rfc1057.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1050.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1057.txt. Obsoleted by RFC1057 [1047]. Status: HIS-TORIC.

RFC1051

[1041] P. A. Prindeville. RFC 1051: Standard for the transmission of IP datagrams and ARP packets over ARCNET networks, March

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1051.txt,   ftp:
//ftp.internic.net/rfc/rfc1201.txt, ftp://ftp.internic.net/
rfc/std46.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1051.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1201.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std46.txt. Obsoleted by RFC1201, STD0046 [1191, ?]. Status: UNKNOWN.

RFC1052

[1042] V. G. Cerf. RFC 1052: IAB recommendations for the development of Internet network management standards, April 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1052.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1052.txt. Status: UNKNOWN.

RFC1053

[1043] S. Levy and T. Jacobson. RFC 1053: Telnet X3 PAD option, April

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1053.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc1053.txt. Status: PROPOSED

STANDARD.

RFC1054

[1044] S. E. Deering. RFC 1054: Host extensions for IP multicast-

ing, May 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1054.
txt,   ftp://ftp.internic.net/rfc/rfc1112.txt,   ftp://ftp.
internic.net/rfc/rfc988.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1054.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1112.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc988.txt. Obsoleted

by RFC1112 [1102]. Obsoletes RFC0988 [978]. Status: UNKNOWN.

RFC1055

[1045] J. L. Romkey. RFC 1055: Nonstandard for transmission of IP datagrams over serial lines: SLIP, June 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1055.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1055.

txt. Status: STANDARD.

RFC1056

[1046] M. L. Lambert. RFC 1056: PCMAIL: A distributed mail system for personal computers, June 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1056.txt,   ftp://ftp.internic.net/rfc/rfc993.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1056.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc993.txt. Obsoletes RFC0993 [983]. Status: IN-FORMATIONAL.

RFC1057

[1047] Sun Microsystems, Inc. RFC 1057: RPC: Remote procedure call protocol specification: Version 2, June 1, 1988. URL ftp://

ftp.internic.net/rfc/rfc1050.txt,   ftp://ftp.internic.net/
rfc/rfc1057.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1050.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1057.txt. Obsoletes

RFC1050 [1040]. Status: INFORMATIONAL.

RFC1058

[1048] C. L. Hedrick. RFC 1058: Routing information protocol, June 1,

1988. URL ftp://ftp.internic.net/rfc/rfc1058.txt,   ftp://
ftp.internic.net/rfc/rfc1388.txt,   ftp://ftp.internic.net/
rfc/rfc1723.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1058.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1388.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1723.txt. Updated by RFC1388,

RFC1723 [1378, 1713]. Status: HISTORIC.

RFC1059

[1049] D. L. Mills. RFC 1059: Network time protocol (version 1) specification and implementation, July 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1059.txt,   ftp://ftp.internic.net/rfc/
rfc1119.txt,   ftp://ftp.internic.net/rfc/rfc1305.txt,   ftp:
//ftp.internic.net/rfc/rfc958.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1059.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1119.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1305.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc958.txt. Obsoleted by RFC1119,

RFC1305 [1109, 1295]. Obsoletes RFC0958 [948]. Status: UNKNOWN.

RFC1060

[1050] J. K. Reynolds and J. Postel. RFC 1060: Assigned numbers, March

1, 1990. URL ftp://ftp.internic.net/rfc/rfc1010.txt,   ftp://
ftp.internic.net/rfc/rfc1060.txt,   ftp://ftp.internic.net/
rfc/rfc1340.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1010.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1060.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1340.txt. Obsoleted by RFC1340

[1330]. Obsoletes RFC1010 [1000]. Status: HISTORIC.

RFC1061

[1051] ???? RFC 1061, November 1993. URL ftp://ftp.internic.net/

rfc/rfc1061.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1061.

txt. Status: UNKNOWN. Not online.

RFC1062

[1052] S. Romano, M. K. Stahl, and M. Recker. RFC 1062: Internet num-

bers, August 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1020.
txt,   ftp://ftp.internic.net/rfc/rfc1062.txt,   ftp://ftp.
internic.net/rfc/rfc1117.txt,   ftp://ftp.internic.net/rfc/
rfc1166.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1020.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1062.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1117.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1166.txt. Obsoleted by RFC1117, RFC1166 [1107, 1156]. Obsoletes RFC1020 [1010]. Status: UNKNOWN.

RFC1063

[1053] J. C. Mogul, C. A. Kent, C. Partridge, and K. McCloghrie. RFC 1063: IP MTU discovery options, July 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1063.txt,   ftp://ftp.internic.net/rfc/
rfc1191.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1063.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1191.txt. Obsoleted by

RFC1191 [1181]. Status: UNKNOWN.

RFC1064

[1054] M. R. Crispin. RFC 1064: Interactive mail access protocol: Ver-

sion 2, July 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1064.
txt,   ftp://ftp.internic.net/rfc/rfc1176.txt,   ftp://ftp.
internic.net/rfc/rfc1203.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1064.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1176.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1203.txt. Obsoleted

by RFC1176, RFC1203 [1166, 1193]. Status: UNKNOWN.

RFC1065

[1055] K. McCloghrie and M. T. Rose. RFC 1065: Structure and identification of management information for TCP/IP-based internets, August

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1034.txt,   ftp://
ftp.internic.net/rfc/rfc1065.txt,   ftp://ftp.internic.net/
rfc/rfc1155.txt, ftp://ftp.internic.net/rfc/std16.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1034.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1065.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1155.txt,   ftp://ftp.math.utah.edu/pub/rfc/std16.txt. Ob-

soleted by RFC1155, STD0016 [1145, ?]. Status: UNKNOWN.

RFC1066

[1056] K. McCloghrie and M. T. Rose. RFC 1066: Management information base for network management of TCP/IP-based inter-

nets, August 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1066.
txt,   ftp://ftp.internic.net/rfc/rfc1156.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1066.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1156.txt. Obsoleted by RFC1156 [1146]. Status: UN-KNOWN.

RFC1067

[1057] J. D. Case, M. Fedor, M. L. Schoffstall, and J. Davin. RFC 1067: Simple network management protocol, August 1, 1988. URL ftp://

ftp.internic.net/rfc/rfc1067.txt,   ftp://ftp.internic.net/
rfc/rfc1098.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1067.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1098.txt. Obsoleted

by RFC1098 [1088]. Status: UNKNOWN.

RFC1068

[1058] A. L. DeSchon and R. T. Braden. RFC 1068: Background file transfer program (BFTP), August 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1068.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1068.

txt. Status: UNKNOWN.

RFC1069

[1059] R. W. Callon and H. W. Braun. RFC 1069: Guidelines for the use of Internet-IP addresses in the ISO Connectionless-Mode Network Protocol, February 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1069.txt,   ftp://ftp.internic.net/rfc/rfc986.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1069.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc986.txt. Obsoletes RFC0986 [976]. Status: UN-KNOWN.

RFC1070

[1060] R. A. Hagens, N. E. Hall, and M. T. Rose. RFC 1070: Use of the Internet as a subnetwork for experimentation with the OSI network

layer, February 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1070.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1070.txt. Status:

UNKNOWN.

RFC1071

[1061] R. T. Braden, D. A. Borman, and C. Partridge. RFC 1071: Computing the Internet checksum, September 1, 1988. URL ftp://

ftp.internic.net/rfc/rfc1071.txt,   ftp://ftp.internic.net/
rfc/rfc1141.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1071.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1141.txt. Updated

by RFC1141 [1131]. Status: UNKNOWN.

RFC1072

[1062] V. Jacobson and R. T. Braden. RFC 1072: TCP extensions for long-delay

paths, October 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1072.
txt,   ftp://ftp.internic.net/rfc/rfc1323.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1072.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1323.txt. Obsoleted by RFC1323 [1313]. Status: UN-KNOWN.

RFC1073

[1063] D. Waitzman. RFC 1073: Telnet window size option, October

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1073.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1073.txt. Status: PROPOSED

STANDARD.

RFC1074

[1064] J. Rekhter. RFC 1074: NSFNET backbone SPF based Interior Gateway Protocol, October 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1074.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1074.

txt. Status: UNKNOWN.

RFC1075

[1065] D. Waitzman, C. Partridge, and S. E. Deering. RFC 1075: Distance vector multicast routing protocol, November 1, 1988. URL ftp://ftp.

internic.net/rfc/rfc1075.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1075.txt. Status: EXPERIMENTAL.

RFC1076

[1066] G. Trewitt and C. Partridge. RFC 1076: HEMS monitoring and control language, November 1, 1988. URL ftp://ftp.internic.net/rfc/

rfc1023.txt,   ftp://ftp.internic.net/rfc/rfc1076.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1023.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1076.txt. Obsoletes RFC1023 [1013]. Status: UN-KNOWN.

RFC1077

[1067] B. M. Leiner. RFC 1077: Critical issues in high bandwidth network-

ing, November 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1077.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1077.txt. Status:

UNKNOWN.

RFC1078

[1068] M. Lottor. RFC 1078: TCP port service multiplexer (TCPMUX),

November 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1078.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1078.txt. Status:

UNKNOWN.

RFC1079

[1069] C. L. Hedrick. RFC 1079: Telnet terminal speed option, December

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1079.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1079.txt. Status: PROPOSED

STANDARD.

RFC1080

[1070] C. L. Hedrick. RFC 1080: Telnet remote flow control option,

November 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1080.
txt,   ftp://ftp.internic.net/rfc/rfc1372.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1080.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1372.txt. Obsoleted by RFC1372 [1362]. Status: UN-KNOWN.

RFC1081

[1071] M. T. Rose. RFC 1081: Post office protocol: Version 3, November

1, 1988. URL ftp://ftp.internic.net/rfc/rfc1081.txt,   ftp://

ftp.internic.net/rfc/rfc1225.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1081.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1225.

txt. Obsoleted by RFC1225 [1215]. Status: UNKNOWN.

RFC1082

[1072] M. T. Rose. RFC 1082: Post office protocol: Version 3: Extended service offerings, November 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1082.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1082.

txt. Status: UNKNOWN.

RFC1083

[1073] Defense Advanced Research Projects Agency and Internet Activities Board. RFC 1083: IAB official protocol standards, De-

cember 1, 1988.  URL ftp://ftp.internic.net/rfc/rfc1083.
txt,   ftp://ftp.internic.net/rfc/rfc1100.txt,   ftp://ftp.
internic.net/rfc/rfc1250.txt,   ftp://ftp.internic.net/rfc/
rfc2200.txt,   ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp:
//ftp.internic.net/rfc/rfc2400.txt, ftp://ftp.internic.net/
rfc/std1.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1083.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1100.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1250.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std1.txt. Obsoleted by RFC1100,

RFC1250, RFC2200, RFC2300, STD0001, RFC2400 [1090, 1240, 2187, 2287, ?, 2387]. Status: HISTORIC.

RFC1084

[1074] J. K. Reynolds. RFC 1084: BOOTP vendor information extensions,

December 1, 1988. URL ftp://ftp.internic.net/rfc/rfc1048.
txt,   ftp://ftp.internic.net/rfc/rfc1084.txt,   ftp://ftp.
internic.net/rfc/rfc1395.txt,   ftp://ftp.internic.net/rfc/
rfc1497.txt,   ftp://ftp.internic.net/rfc/rfc1533.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1048.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1084.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1395.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1497.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1533.txt. Obsoleted by

RFC1395, RFC1497, RFC1533 [1385, 1487, 1523]. Obsoletes RFC1048 [1038]. Status: UNKNOWN.

RFC1085

[1075] M. T. Rose. RFC 1085: ISO presentation services on top of TCP/IP based internets, December 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1085.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1085.

txt. Status: UNKNOWN.

RFC1086

[1076] J. P. Onions and M. T. Rose. RFC 1086: ISO-TP0 bridge between TCP and X.25, December 1, 1988. URL ftp://ftp.internic.net/

rfc/rfc1086.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1086.

txt. Status: UNKNOWN.

RFC1087

[1077] Defense Advanced Research Projects Agency and Internet Activities Board. RFC 1087: Ethics and the Internet, January 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1087.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1087.txt. Status: UNKNOWN.

RFC1088

[1078] L. J. McLaughlin. RFC 1088: Standard for the transmission of IP datagrams over netBIOS networks, February 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1088.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1088.txt. Status: STANDARD.

RFC1089

[1079] M. L. Schoffstall, C. Davin, M. Fedor, and J. D. Case. RFC 1089: SNMP over ethernet, February 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1089.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1089.

txt. Status: UNKNOWN.

RFC1090

[1080] R. Ullmann. RFC 1090: SMTP on X.25, February 1, 1989. URL ftp://

ftp.internic.net/rfc/rfc1090.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1090.txt. Status: UNKNOWN.

RFC1091

[1081] J. VanBokkelen. RFC 1091: Telnet terminal-type option, February 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1091.txt,   ftp://
ftp.internic.net/rfc/rfc930.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1091.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc930.

txt. Obsoletes RFC0930 [920]. Status: PROPOSED STANDARD.

RFC1092

[1082] J. Rekhter. RFC 1092: EGP and policy based routing in the new NSFNET backbone, February 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1092.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1092.

txt. Status: UNKNOWN.

RFC1093

[1083] H. W. Braun. RFC 1093: NSFNET routing architecture, February

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1093.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1093.txt. Status: UNKNOWN.

RFC1094

[1084] Sun Microsystems, Inc. RFC 1094: NFS: Network File System Protocol specification, March 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1094.txt,   ftp://ftp.internic.net/rfc/
rfc1813.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1094.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1813.txt. See also RFC1813

[1803]. Status: INFORMATIONAL.

RFC1095

[1085] U. S. Warrier and L. Besaw. RFC 1095: Common Management Information Services and Protocol over TCP/IP (CMOT), April 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1095.txt,   ftp://
ftp.internic.net/rfc/rfc1189.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1095.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1189.

txt. Obsoleted by RFC1189 [1179]. Status: UNKNOWN.

RFC1096

[1086] G. A. Marcy. RFC 1096: Telnet X display location option, March

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1096.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1096.txt. Status: PROPOSED

STANDARD.

RFC1097

[1087] B. Miller. RFC 1097: Telnet subliminal-message option, April 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1097.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1097.txt. Status: UNKNOWN.

RFC1098

[1088] J. D. Case, M. Fedor, M. L. Schoffstall, and C. Davin. RFC 1098: Simple network management protocol (SNMP), April 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1067.txt,   ftp://
ftp.internic.net/rfc/rfc1098.txt,   ftp://ftp.internic.net/
rfc/rfc1157.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1067.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1098.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1157.txt. Obsoleted by RFC1157

[1147]. Obsoletes RFC1067 [1057]. Status: UNKNOWN.

RFC1099

[1089] J. Reynolds. RFC 1099: Request for comments summary: RFC numbers 1000–1099, December 1991. URL ftp://ftp.internic.net/

rfc/rfc1099.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1099.

txt. Status: INFORMATIONAL.

RFC1100

[1090] Defense Advanced Research Projects Agency and Internet Activities Board. RFC 1100: IAB official protocol standards, April 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1083.txt,   ftp://
ftp.internic.net/rfc/rfc1100.txt,   ftp://ftp.internic.net/
rfc/rfc1130.txt,   ftp://ftp.internic.net/rfc/rfc1250.txt,
ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.internic.
net/rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/rfc2400.txt,
ftp://ftp.internic.net/rfc/std1.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1083.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1100.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1130.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1250.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.
txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt. Obsoleted by

RFC1130, RFC1250, RFC2200, RFC2300, STD0001, RFC2400 [1120, 1240, 2187, 2287, ?, 2387]. Obsoletes RFC1083 [1073]. Status: HIS-TORIC.

RFC1101

[1091] P. V. Mockapetris. RFC 1101: DNS encoding of network names and other

types, April 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.
internic.net/rfc/rfc1101.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1034.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1101.txt. Updates

RFC1034, RFC1035 [1024, 1025]. Status: UNKNOWN.

RFC1102

[1092] D. D. Clark. RFC 1102: Policy routing in Internet protocols, May

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1102.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1102.txt. Status: UNKNOWN.

RFC1103

[1093] D. Katz. RFC 1103: Proposed standard for the transmission of IP datagrams over FDDI networks, June 1, 1989. URL ftp://

ftp.internic.net/rfc/rfc1103.txt,   ftp://ftp.internic.net/

rfc/rfc1188.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1103.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1188.txt. Obsoleted

by RFC1188 [1178]. Status: UNKNOWN.

RFC1104

[1094] H. W. Braun. RFC 1104: Models of policy based routing, June 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1104.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1104.txt. Status: UNKNOWN.

RFC1105

[1095] K. Lougheed and Y. Rekhter. RFC 1105: Border Gateway Protocol

(BGP), June 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1105.
txt,   ftp://ftp.internic.net/rfc/rfc1163.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1105.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1163.txt. Obsoleted by RFC1163 [1153]. Status: EXPER-IMENTAL.

RFC1106

[1096] R. Fox. RFC 1106: TCP big window and NAK options, June 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1106.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1106.txt. Status: UNKNOWN.

RFC1107

[1097] K. R. Sollins. RFC 1107: Plan for Internet directory services, July

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1107.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1107.txt. Status: INFORMA-

TIONAL.

RFC1108

[1098] S. Kent. RFC 1108: U.S. Department of Defense security options for the Internet Protocol, November 1991. URL ftp://

ftp.internic.net/rfc/rfc1038.txt,   ftp://ftp.internic.net/
rfc/rfc1108.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1038.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1108.txt. Obsoletes

RFC1038 [1028]. Status: HISTORIC.

RFC1109

[1099] V. G. Cerf. RFC 1109: Report of the second ad hoc network management review group, August 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1109.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1109.

txt. Status: UNKNOWN.

RFC1110

[1100] A. M. McKenzie. RFC 1110: Problem with the TCP big window op-

tion, August 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1110.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1110.txt. Status:

UNKNOWN.

RFC1111

[1101] J. Postel. RFC 1111: Request for comments on Request for Comments: Instructions to RFC authors, August 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1111.txt,   ftp://ftp.internic.net/rfc/
rfc1543.txt,   ftp://ftp.internic.net/rfc/rfc2223.txt,   ftp:
//ftp.internic.net/rfc/rfc825.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1111.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1543.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2223.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc825.txt. Obsoleted by RFC1543,

RFC2223 [1533, 2210]. Obsoletes RFC0825 [815]. Status: INFORMA-TIONAL.

RFC1112

[1102] S. E. Deering. RFC 1112: Host extensions for IP multicasting, August

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1054.txt,   ftp://
ftp.internic.net/rfc/rfc1112.txt,   ftp://ftp.internic.net/
rfc/rfc2236.txt,   ftp://ftp.internic.net/rfc/rfc988.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1054.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1112.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2236.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc988.txt.

Obsoletes RFC0988, RFC1054 [978, 1044]. See also STD0005 [?]. Updated by RFC2236 [2223]. Status: STANDARD.

RFC1113

[1103] J. Linn. RFC 1113: Privacy enhancement for Internet electronic mail: Part I — message encipherment and authentication procedures, August

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1040.txt,   ftp://
ftp.internic.net/rfc/rfc1113.txt,   ftp://ftp.internic.net/
rfc/rfc1421.txt,   ftp://ftp.internic.net/rfc/rfc989.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1040.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1113.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1421.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc989.txt.

Obsoleted by RFC1421 [1411]. Obsoletes RFC0989, RFC1040 [979, 1030]. Status: HISTORIC.

RFC1114

[1104] S. T. Kent and J. Linn. RFC 1114: Privacy enhancement for Internet electronic mail: Part II — certificate-based key manage-

ment, August 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1114.
txt,   ftp://ftp.internic.net/rfc/rfc1422.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1114.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1422.txt. Obsoleted by RFC1422 [1412]. Status: HIS-TORIC.

RFC1115

[1105] J. Linn. RFC 1115: Privacy enhancement for Internet electronic mail: Part III — algorithms, modes, and identifiers, August 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1115.txt,   ftp://
ftp.internic.net/rfc/rfc1423.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1115.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1423.

txt. Obsoleted by RFC1423 [1413]. Status: HISTORIC.

RFC1116

[1106] D. A. Borman. RFC 1116: Telnet Linemode option, August 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1116.txt,   ftp://
ftp.internic.net/rfc/rfc1184.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1116.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1184.

txt. Obsoleted by RFC1184 [1174]. Status: PROPOSED STANDARD.

RFC1117

[1107] S. Romano, M. K. Stahl, and M. Recker. RFC 1117: Internet num-

bers, August 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1020.
txt,   ftp://ftp.internic.net/rfc/rfc1062.txt,   ftp://ftp.
internic.net/rfc/rfc1117.txt,   ftp://ftp.internic.net/rfc/
rfc1166.txt,   ftp://ftp.internic.net/rfc/rfc997.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1020.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1062.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1117.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1166.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc997.txt.  Obsoleted by

RFC1166 [1156]. Obsoletes RFC1062, RFC1020, RFC0997 [1052, 1010, 987]. Status: INFORMATIONAL.

RFC1118

[1108] E. Krol. RFC 1118: Hitchhikers guide to the Internet, September

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1118.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1118.txt. Status: INFORMA-

TIONAL.

RFC1119

[1109] D. L. Mills. RFC 1119: Network time protocol (version 2) specification and implementation, September 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1059.txt,   ftp://ftp.internic.net/rfc/
rfc1119.txt,   ftp://ftp.internic.net/rfc/rfc1305.txt,   ftp:
//ftp.internic.net/rfc/rfc958.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1059.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1119.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1305.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc958.txt. Obsoleted by RFC1305

[1295]. Obsoletes RFC0958, RFC1059 [948, 1049]. See also STD0012 [?]. Status: STANDARD.

RFC1120

[1110] V. Cerf. RFC 1120: Internet Activities Board, September 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1120.txt,   ftp://
ftp.internic.net/rfc/rfc1160.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1120.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1160.

txt. Obsoleted by RFC1160 [1150]. Status: INFORMATIONAL.

RFC1121

[1111] J. Postel, L. Kleinrock, V. G. Cerf, and B. Boehm. RFC 1121: Act one — the poems, September 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1121.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1121.

txt. Status: INFORMATIONAL.

RFC1122

[1112] R. T. Braden. RFC 1122: Requirements for Internet hosts — communication layers, October 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1122.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1122.

txt. See also STD0003 [?]. Status: STANDARD.

RFC1123

[1113] R. T. Braden. RFC 1123: Requirements for Internet hosts — application and support, October 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1123.txt,   ftp://ftp.internic.net/rfc/
rfc2181.txt,   ftp://ftp.internic.net/rfc/rfc822.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1123.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2181.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc822.txt. See also STD0003 [?]. Updates RFC0822 [812]. Updated by RFC2181 [2168]. Status: STANDARD.

RFC1124

[1114] B. M. Leiner. RFC 1124: Policy issues in interconnecting networks,

September 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1124.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1124.txt. Status:

UNKNOWN.

RFC1125

[1115] D. Estrin. RFC 1125: Policy requirements for inter administrative domain routing, November 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1125.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1125.

txt. Status: UNKNOWN.

RFC1126

[1116] M. Little. RFC 1126: Goals and functional requirements for interautonomous system routing, October 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1126.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1126.txt. Status: UNKNOWN.

RFC1127

[1117] R. T. Braden. RFC 1127: Perspective on the host requirements RFCs,

October 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1127.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1127.txt. Status: INFOR-

MATIONAL.

RFC1128

[1118] D. L. Mills. RFC 1128: Measured performance of the Network Time Protocol in the Internet system, October 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1128.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1128.txt. Status: UNKNOWN.

RFC1129

[1119] D. L. Mills. RFC 1129: Internet time synchronization: The Network Time Protocol, October 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1129.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1129.

txt. Status: UNKNOWN.

RFC1130

[1120] Defense Advanced Research Projects Agency and Internet Activities Board. RFC 1130: IAB official protocol standards, October 1,

1989. URL ftp://ftp.internic.net/rfc/rfc1100.txt,   ftp://
ftp.internic.net/rfc/rfc1130.txt,   ftp://ftp.internic.net/
rfc/rfc1140.txt,   ftp://ftp.internic.net/rfc/rfc1250.txt,

ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.internic.
net/rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/rfc2400.txt,
ftp://ftp.internic.net/rfc/std1.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1100.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1130.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1140.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1250.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.
txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt. Obsoleted by

RFC1140, RFC1250, RFC2200, RFC2300, STD0001, RFC2400 [1130, 1240, 2187, 2287, ?, 2387]. Obsoletes RFC1100 [1090]. Status: HIS-TORIC.

RFC1131

[1121] J. Moy. RFC 1131: OSPF specification, October 1, 1989. URL ftp://

ftp.internic.net/rfc/rfc1131.txt,   ftp://ftp.internic.net/
rfc/rfc1247.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1131.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1247.txt. Obsoleted

by RFC1247 [1237]. Status: PROPOSED STANDARD.

RFC1132

[1122] L. J. McLaughlin. RFC 1132: Standard for the transmission of 802.2 packets over IPX networks, November 1, 1989. URL ftp://ftp.

internic.net/rfc/rfc1132.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1132.txt. Status: STANDARD.

RFC1133

[1123] J. Y. Yu and H. W. Braun. RFC 1133: Routing between the NSFNET and the DDN, November 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1133.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1133.

txt. Status: INFORMATIONAL.

RFC1134

[1124] D. Perkins. RFC 1134: Point-to-point protocol: A proposal for multi-protocol transmission of datagrams over point-to-point links,

November 1, 1989. URL ftp://ftp.internic.net/rfc/rfc1134.
txt,   ftp://ftp.internic.net/rfc/rfc1171.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1134.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1171.txt. Obsoleted by RFC1171 [1161]. Status: PRO-POSED STANDARD.

RFC1135

[1125] J. K. Reynolds. RFC 1135: Helminthiasis of the Internet, December

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1135.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc1135.txt. Status: INFORMA-

TIONAL.

RFC1136

[1126] S. Hares and D. Katz. RFC 1136: Administrative domains and routing domains: A model for routing in the Internet, December

1, 1989. URL ftp://ftp.internic.net/rfc/rfc1136.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1136.txt. Status: INFORMA-

TIONAL.

RFC1137

[1127] S. Kille. RFC 1137: Mapping between full RFC 822 and RFC 822 with restricted encoding, December 1, 1989. URL ftp://ftp.internic.net/

rfc/rfc1137.txt,   ftp://ftp.internic.net/rfc/rfc976.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1137.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc976.txt. Updates RFC0976 [966]. Status: HIS-TORIC.

RFC1138

[1128] S. E. Kille. RFC 1138: Mapping between X.400(1988) /ISO 10021 and RFC 822, December 1, 1989. URL ftp://ftp.internic.net/rfc/

rfc1026.txt,   ftp://ftp.internic.net/rfc/rfc1138.txt,   ftp:
//ftp.internic.net/rfc/rfc1148.txt,   ftp://ftp.internic.net/
rfc/rfc1327.txt,   ftp://ftp.internic.net/rfc/rfc1495.txt,
ftp://ftp.internic.net/rfc/rfc2156.txt,   ftp://ftp.internic.
net/rfc/rfc822.txt, ftp://ftp.internic.net/rfc/rfc987.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1026.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1138.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1148.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1327.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1495.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2156.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc822.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc987.

txt. Obsoleted by RFC1327, RFC1495, RFC2156 [1317, 1485, 2143]. Updates RFC0822, RFC0987, RFC1026 [812, 977, 1016]. Updated by RFC1148 [1138]. Status: EXPERIMENTAL.

RFC1139

[1129] R. A. Hagens. RFC 1139: Echo function for ISO 8473, January 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1139.txt,   ftp://
ftp.internic.net/rfc/rfc1574.txt,   ftp://ftp.internic.net/
rfc/rfc1575.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1139.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1574.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1575.txt. Obsoleted by RFC1574,

RFC1575 [1564, 1565]. Status: PROPOSED STANDARD.

RFC1140

[1130] Defense Advanced Research Projects Agency and Internet Activities Board. RFC 1140: IAB official protocol standards, May 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1130.txt,   ftp://
ftp.internic.net/rfc/rfc1140.txt,   ftp://ftp.internic.net/
rfc/rfc1200.txt,   ftp://ftp.internic.net/rfc/rfc1250.txt,
ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.internic.
net/rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/rfc2400.txt,
ftp://ftp.internic.net/rfc/std1.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1130.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1140.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1200.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1250.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.
txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt. Obsoleted by

RFC1200, RFC1250, RFC2200, RFC2300, STD0001, RFC2400 [1190, 1240, 2187, 2287, ?, 2387]. Obsoletes RFC1130 [1120]. Status: HIS-TORIC.

RFC1141

[1131] T. Mallory and A. Kullberg. RFC 1141: Incremental updating of the Internet checksum, January 1, 1990. URL ftp://ftp.

internic.net/rfc/rfc1071.txt,   ftp://ftp.internic.net/rfc/
rfc1141.txt,   ftp://ftp.internic.net/rfc/rfc1624.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1071.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1141.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1624.txt. Updates RFC1071 [1061]. Updated by RFC1624 [1614]. Status: INFORMATIONAL.

RFC1142

[1132] D. Oran. RFC 1142: OSI IS-IS intra-domain routing protocol, February

1, 1990. URL ftp://ftp.internic.net/rfc/rfc1142.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1142.txt. Status: INFORMA-

TIONAL.

RFC1143

[1133] D. J. Bernstein. RFC 1143: The Q method of implementing TELNET option negotiation, February 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1143.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1143.

txt. Status: EXPERIMENTAL.

RFC1144

[1134] V. Jacobson. RFC 1144: Compressing TCP/IP headers for lowspeed serial links, February 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1144.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1144.

txt. Status: PROPOSED STANDARD.

RFC1145

[1135] J. Zweig and C. Partridge. RFC 1145: TCP alternate checksum op-

tions, February 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1145.
txt,   ftp://ftp.internic.net/rfc/rfc1146.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1145.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1146.txt. Obsoleted by RFC1146 [1136]. Status: EXPER-IMENTAL.

RFC1146

[1136] J. Zweig and C. Partridge. RFC 1146: TCP alternate checksum op-

tions, March 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1145.
txt,   ftp://ftp.internic.net/rfc/rfc1146.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1145.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1146.txt. Obsoletes RFC1145 [1135]. Status: EXPERI-MENTAL.

RFC1147

[1137] R. H. Stine. RFC 1147: FYI on a network management tool catalog: Tools for monitoring and debugging TCP/IP internets and interconnected devices, April 1, 1990. URL ftp://ftp.internic.net/rfc/

fyi2.txt,   ftp://ftp.internic.net/rfc/rfc1147.txt,   ftp:/
/ftp.internic.net/rfc/rfc1470.txt,   ftp://ftp.math.utah.edu/
pub/rfc/fyi2.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1147.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1470.txt. See also

FYI0002 [?]. Obsoleted by RFC1470 [1460]. Updated by FYI0002 [?]. Status: INFORMATIONAL.

RFC1148

[1138] S. E. Kille. RFC 1148: Mapping between X.400(1988) /ISO 10021

and RFC 822, March 1, 1990. URL ftp://ftp.internic.net/rfc/
rfc1026.txt,   ftp://ftp.internic.net/rfc/rfc1138.txt,   ftp:
//ftp.internic.net/rfc/rfc1148.txt,   ftp://ftp.internic.net/
rfc/rfc1327.txt,   ftp://ftp.internic.net/rfc/rfc1495.txt,
ftp://ftp.internic.net/rfc/rfc2156.txt,   ftp://ftp.internic.
net/rfc/rfc822.txt, ftp://ftp.internic.net/rfc/rfc987.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1026.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1138.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1148.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1327.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1495.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2156.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc822.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc987.

txt. Obsoleted by RFC1327, RFC1495, RFC2156 [1317, 1485, 2143]. Updates RFC0822, RFC0987, RFC1026, RFC1138 [812, 977, 1016, 1128]. Status: EXPERIMENTAL.

RFC1149

[1139] D. Waitzman. RFC 1149: Standard for the transmission of IP datagrams on avian carriers, April 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1149.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1149.

txt. Status: EXPERIMENTAL.

RFC1150

[1140] G. S. Malkin and J. K. Reynolds. RFC 1150: F.Y.I. on F.Y.I.: Introduction to the F.Y.I. notes, March 1, 1990. URL ftp://ftp.internic.

net/rfc/fyi1.txt,   ftp://ftp.internic.net/rfc/rfc1150.txt,
 ftp://ftp.math.utah.edu/pub/rfc/fyi1.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc1150.txt. See also FYI0001 [?]. Status: INFOR-MATIONAL.

RFC1151

[1141] C. Partridge and R. M. Hinden. RFC 1151: Version 2 of the reliable data protocol (RDP), April 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1151.txt,   ftp://ftp.internic.net/rfc/rfc908.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1151.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc908.txt. Updates RFC0908 [898]. Status: EX-PERIMENTAL.

RFC1152

[1142] C. Partridge. RFC 1152: Workshop report: Internet research steering group workshop on very-high-speed networks, April 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1152.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1152.txt. Status: INFORMA-

TIONAL.

RFC1153

[1143] F. J. Wancho. RFC 1153: Digest message format, April 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1153.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1153.txt. Status: EXPERIMEN-

TAL.

RFC1154

[1144] D. Robinson and R. Ullmann. RFC 1154: Encoding header field for Internet messages, April 1, 1990. URL ftp://ftp.

internic.net/rfc/rfc1154.txt,   ftp://ftp.internic.net/rfc/
rfc1505.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1154.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc1505.txt. Obsoleted by

RFC1505 [1495]. Status: EXPERIMENTAL.

RFC1155

[1145] M. T. Rose and K. McCloghrie. RFC 1155: Structure and identification of management information for TCP/IP-based internets, May

1, 1990. URL ftp://ftp.internic.net/rfc/rfc1065.txt,   ftp:
//ftp.internic.net/rfc/rfc1155.txt, ftp://ftp.internic.net/
rfc/std16.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1065.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1155.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std16.txt. See also STD0016 [?]. Obsoletes RFC1065 [1055]. Status: STANDARD.

RFC1156

[1146] K. McCloghrie and M. T. Rose. RFC 1156: Management information base for network management of TCP/IP-based inter-

nets, May 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1066.
txt,   ftp://ftp.internic.net/rfc/rfc1156.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1066.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1156.txt. Obsoletes RFC1066 [1056]. Status: HISTORIC.

RFC1157

[1147] J. D. Case, M. Fedor, M. L. Schoffstall, and C. Davin. RFC 1157: Simple network management protocol (SNMP), May 1, 1990. URL ftp://ftp.

internic.net/rfc/rfc1098.txt,   ftp://ftp.internic.net/rfc/
rfc1157.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1098.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1157.txt. See also STD0015

[?]. Obsoletes RFC1098 [1088]. Status: STANDARD.

RFC1158

[1148] M. T. Rose. RFC 1158: Management information base for network management of TCP/IP-based internets: MIB-II, May 1,

1990.  URL ftp://ftp.internic.net/rfc/rfc1158.txt,   ftp:
//ftp.internic.net/rfc/rfc1213.txt, ftp://ftp.internic.net/
rfc/std17.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1158.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1213.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std17.txt. Obsoleted by RFC1213, STD0017 [1203, ?]. Status: DRAFT STANDARD.

RFC1159

[1149] R. Nelson. RFC 1159: Message send protocol, June 1, 1990. URL ftp://

ftp.internic.net/rfc/rfc1159.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1159.txt. Status: EXPERIMENTAL.

RFC1160

[1150] V. Cerf. RFC 1160: Internet Activities Board, May 1, 1990. URL ftp://

ftp.internic.net/rfc/rfc1120.txt,   ftp://ftp.internic.net/
rfc/rfc1160.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1120.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1160.txt. Obsoletes

RFC1120 [1110]. Status: INFORMATIONAL.

RFC1161

[1151] M. T. Rose. RFC 1161: SNMP over OSI, June 1, 1990. URL ftp://

ftp.internic.net/rfc/rfc1161.txt,   ftp://ftp.internic.net/
rfc/rfc1418.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1161.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1418.txt. Obsoleted

by RFC1418 [1408]. Status: EXPERIMENTAL.

RFC1162

[1152] G. Satz. RFC 1162: Connectionless network protocol (ISO 8473) and end system to intermediate system (ISO 9542) management informa-

tion base, June 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1162.
txt,   ftp://ftp.internic.net/rfc/rfc1238.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1162.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1238.txt. Obsoleted by RFC1238 [1228]. Status: EXPER-IMENTAL.

RFC1163

[1153] K. Lougheed and Y. Rekhter. RFC 1163: Border Gateway Protocol

(BGP), June 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1105.
txt,   ftp://ftp.internic.net/rfc/rfc1163.txt,   ftp://ftp.
internic.net/rfc/rfc1267.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1105.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1163.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1267.txt. Obsoleted

by RFC1267 [1257]. Obsoletes RFC1105 [1095]. Status: HISTORIC.

RFC1164

[1154] J. C. Honig, D. Katz, M. Mathis, Y. Rekhter, and J. Y. Yu. RFC 1164: Application of the Border Gateway Protocol in the Inter-

net, June 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1164.
txt,   ftp://ftp.internic.net/rfc/rfc1268.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1164.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1268.txt. Obsoleted by RFC1268 [1258]. Status: HIS-TORIC.

RFC1165

[1155] J. Crowcroft and J. P. Onions. RFC 1165: Network time protocol (NTP) over the OSI remote operations service, June 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1165.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1165.txt. Status: EXPERIMEN-

TAL.

RFC1166

[1156] S. Kirkpatrick, M. K. Stahl, and M. Recker. RFC 1166: Internet

numbers, July 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1020.
txt,   ftp://ftp.internic.net/rfc/rfc1062.txt,   ftp://ftp.
internic.net/rfc/rfc1117.txt,   ftp://ftp.internic.net/rfc/
rfc1166.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1020.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1062.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1117.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1166.txt. Obsoletes RFC1117, RFC1062, RFC1020 [1107, 1052, 1010]. Status: INFORMATIONAL.

RFC1167

[1157] V. G. Cerf. RFC 1167: Thoughts on the National Research and Education Network, July 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1167.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1167.

txt. Status: INFORMATIONAL.

RFC1168

[1158] A. Westine, A. L. DeSchon, J. Postel, and C. E. Ward. RFC 1168: Intermail and commercial mail relay services, July 1, 1990. URL ftp://ftp.

internic.net/rfc/rfc1168.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1168.txt. Status: INFORMATIONAL.

RFC1169

[1159] V. G. Cerf and K. L. Mills. RFC 1169: Explaining the role of GOSIP,

August 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1169.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1169.txt. Status: INFOR-

MATIONAL.

RFC1170

[1160] R. B. Fougner. RFC 1170: Public key standards and licenses, Jan-

uary 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1170.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1170.txt. Status: INFOR-

MATIONAL.

RFC1171

[1161] D. Perkins. RFC 1171: Point-to-point protocol for the transmission of multi-protocol datagrams over point-to-point links, July 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1134.txt,   ftp://
ftp.internic.net/rfc/rfc1171.txt,   ftp://ftp.internic.net/
rfc/rfc1331.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1134.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1171.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1331.txt. Obsoleted by RFC1331

[1321]. Obsoletes RFC1134 [1124]. Status: DRAFT STANDARD.

RFC1172

[1162] D. Perkins and R. Hobby. RFC 1172: Point-to-point protocol (PPP) initial configuration options, July 1, 1990. URL ftp://

ftp.internic.net/rfc/rfc1172.txt,   ftp://ftp.internic.net/
rfc/rfc1331.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1172.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1331.txt. Obsoleted

by RFC1331 [1321]. Status: PROPOSED STANDARD.

RFC1173

[1163] J. Van Bokkelen. RFC 1173: Responsibilities of host and network managers: A summary of the “oral tradition” of the Internet, Au-

gust 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1173.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1173.txt. Status: INFOR-

MATIONAL.

RFC1174

[1164] V. G. Cerf. RFC 1174: IAB recommended policy on distributing Internet identifier assignment and IAB recommended policy change to Internet “connected” status, August 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1174.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1174.

txt. Status: INFORMATIONAL.

RFC1175

[1165] K. L. Bowers, T. L. LaQuey, J. K. Reynolds, K. Roubicek, M. K. Stahl, and A. Yuan. RFC 1175: FYI on where to start: A bibliography of internetworking information, August 1, 1990. URL ftp://ftp.internic.

net/rfc/fyi3.txt,   ftp://ftp.internic.net/rfc/rfc1175.txt,
 ftp://ftp.math.utah.edu/pub/rfc/fyi3.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc1175.txt. See also FYI0003 [?]. Status: INFOR-MATIONAL.

RFC1176

[1166] M. R. Crispin. RFC 1176: Interactive mail access protocol: Version

2, August 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1064.
txt,   ftp://ftp.internic.net/rfc/rfc1176.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1064.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1176.txt. Obsoletes RFC1064 [1054]. Status: EXPERI-MENTAL.

RFC1177

[1167] G. S. Malkin, A. N. Marine, and J. K. Reynolds. RFC 1177: FYI on questions and answers: Answers to commonly asked “new Internet user”

questions, August 1, 1990. URL ftp://ftp.internic.net/rfc/fyi4.
txt,   ftp://ftp.internic.net/rfc/rfc1177.txt,   ftp://ftp.
internic.net/rfc/rfc1206.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi4.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1177.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1206.txt. See also FYI0004

[?]. Obsoleted by RFC1206, FYI0004 [1196, ?]. Status: INFORMA-TIONAL.

RFC1178

[1168] D. Libes. RFC 1178: Choosing a name for your computer, August 1,

1990. URL ftp://ftp.internic.net/rfc/fyi5.txt,   ftp://ftp.
internic.net/rfc/rfc1178.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi5.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1178.txt.

See also FYI0005 [?]. Status: INFORMATIONAL.

RFC1179

[1169] L. McLaughlin, III. RFC 1179: Line printer daemon protocol, Au-

gust 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1179.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1179.txt. Status: INFOR-

MATIONAL.

RFC1180

[1170] T. J. Socolofsky and C. J. Kale. RFC 1180: TCP/IP tutorial, Jan-

uary 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1180.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1180.txt. Status: INFOR-

MATIONAL.

RFC1181

[1171] R. Blokzijl. RFC 1181: RIPE terms of reference, September 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1181.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1181.txt. Status: INFORMA-

TIONAL.

RFC1182

[1172] ???? RFC 1182, 19xx. URL ftp://ftp.internic.net/rfc/rfc1182.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1182.txt. Status:

UNKNOWN. Not online.

RFC1183

[1173] C. F. Everhart, L. A. Mamakos, R. Ullmann, and P. V. Mockapetris. RFC 1183: New DNS RR definitions, October 1, 1990. URL ftp://ftp.

internic.net/rfc/rfc1034.txt,   ftp://ftp.internic.net/rfc/
rfc1035.txt,   ftp://ftp.internic.net/rfc/rfc1183.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1034.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1183.txt. Updates RFC1034, RFC1035 [1024, 1025]. Status: EX-PERIMENTAL.

RFC1184

[1174] D. A. Borman. RFC 1184: Telnet Linemode option, October 1,

1990. URL ftp://ftp.internic.net/rfc/rfc1116.txt,   ftp://
ftp.internic.net/rfc/rfc1184.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1116.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1184.

txt. Obsoletes RFC1116 [1106]. Status: DRAFT STANDARD.

RFC1185

[1175] V. Jacobson, R. T. Braden, and L. Zhang. RFC 1185: TCP extension for high-speed paths, October 1, 1990. URL ftp://ftp.

internic.net/rfc/rfc1185.txt,   ftp://ftp.internic.net/rfc/
rfc1323.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1185.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1323.txt. Obsoleted by

RFC1323 [1313]. Status: EXPERIMENTAL.

RFC1186

[1176] R. L. Rivest. RFC 1186: MD4 message digest algorithm, October

1, 1990. URL ftp://ftp.internic.net/rfc/rfc1186.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1186.txt. Status: INFORMA-

TIONAL.

RFC1187

[1177] M. T. Rose, K. McCloghrie, and J. R. Davin. RFC 1187: Bulk table retrieval with the SNMP, October 1, 1990. URL ftp://ftp.internic.

net/rfc/rfc1187.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1187.

txt. Status: EXPERIMENTAL.

RFC1188

[1178] D. Katz. RFC 1188: Proposed standard for the transmission of IP datagrams over FDDI networks, October 1, 1990. URL ftp://

ftp.internic.net/rfc/rfc1103.txt,   ftp://ftp.internic.net/
rfc/rfc1188.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1103.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1188.txt. Obsoletes

RFC1103 [1093]. Status: DRAFT STANDARD.

RFC1189

[1179] U. S. Warrier, L. Besaw, L. LaBarre, and B. D. Handspicker. RFC 1189: Common Management Information Services and Protocols for the Internet (CMOT and CMIP), October 1, 1990. URL ftp://

ftp.internic.net/rfc/rfc1095.txt,   ftp://ftp.internic.net/
rfc/rfc1189.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1095.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1189.txt. Obsoletes

RFC1095 [1085]. Status: HISTORIC.

RFC1190

[1180] C. Topolcic. RFC 1190: Experimental Internet Stream Protocol: Version 2 (ST-II), October 1, 1990. URL ftp://ftp.internic.net/

rfc/ien119.txt,   ftp://ftp.internic.net/rfc/rfc1190.txt,
ftp://ftp.internic.net/rfc/rfc1819.txt,   ftp://ftp.math.utah.
edu/pub/rfc/ien119.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1190.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1819.txt. Obsoleted

by RFC1819 [1809]. Obsoletes IEN119 [?]. Status: EXPERIMENTAL.

RFC1191

[1181] J. C. Mogul and S. E. Deering. RFC 1191: Path MTU discovery,

November 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1063.
txt,   ftp://ftp.internic.net/rfc/rfc1191.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1063.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1191.txt. Obsoletes RFC1063 [1053]. Status: DRAFT STANDARD.

RFC1192

[1182] B. Kahin. RFC 1192: Commercialization of the Internet summary report,

November 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1192.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1192.txt. Status: INFOR-

MATIONAL.

RFC1193

[1183] D. Ferrari. RFC 1193: Client requirements for real-time communication services, November 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1193.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1193.

txt. Status: INFORMATIONAL.

RFC1194

[1184] D. P. Zimmerman. RFC 1194: Finger user information protocol,

November 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1194.
txt,   ftp://ftp.internic.net/rfc/rfc1196.txt,   ftp://ftp.
internic.net/rfc/rfc1288.txt,   ftp://ftp.internic.net/rfc/
rfc742.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1194.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1196.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1288.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc742.txt. Obsoleted by RFC1288, RFC1196 [1278, 1186]. Obsoletes RFC0742 [732]. Status: DRAFT STANDARD.

RFC1195

[1185] R. W. Callon. RFC 1195: Use of OSI IS-IS for routing in TCP/IP and dual environments, December 1, 1990. URL ftp://ftp.internic.net/

rfc/rfc1195.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1195.

txt. Status: PROPOSED STANDARD.

RFC1196

[1186] D. P. Zimmerman. RFC 1196: Finger user information protocol,

December 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1194.
txt,   ftp://ftp.internic.net/rfc/rfc1196.txt,   ftp://ftp.
internic.net/rfc/rfc1288.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1194.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1196.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1288.txt. Obsoleted

by RFC1288 [1278]. Obsoletes RFC1194, RFC0742 [1184, 732]. Status: DRAFT STANDARD.

RFC1197

[1187] M. Sherman. RFC 1197: Using ODA for translating multimedia informa-

tion, December 1, 1990. URL ftp://ftp.internic.net/rfc/rfc1197.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1197.txt. Status:

INFORMATIONAL.

RFC1198

[1188] R. W. Scheifler. RFC 1198: FYI on the X window system, January 1,

1991. URL ftp://ftp.internic.net/rfc/fyi6.txt,   ftp://ftp.
internic.net/rfc/rfc1198.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi6.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1198.txt.

See also FYI0006 [?]. Status: INFORMATIONAL.

RFC1199

[1189] J. Reynolds. RFC 1199: Request for comments summary notes: 1100–

1199, December 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1199.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1199.txt. Status:

INFORMATIONAL.

RFC1200

[1190] Defense Advanced Research Projects Agency and Internet Activities Board. RFC 1200: IAB official protocol standards, April 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1140.txt,   ftp://
ftp.internic.net/rfc/rfc1200.txt,   ftp://ftp.internic.net/
rfc/rfc1250.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1140.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1250.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. Obsoleted by RFC1250, RFC2200, RFC2300, STD0001, RFC2400 [1240, 2187, 2287, ?, 2387]. Obsoletes RFC1140 [1130]. Status: HISTORIC.

RFC1201

[1191] D. Provan. RFC 1201: Transmitting IP traffic over ARCNET networks,

February 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1051.
txt,   ftp://ftp.internic.net/rfc/rfc1201.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1051.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1201.txt. Obsoletes RFC1051 [1041]. Status: STAN-DARD.

RFC1202

[1192] M. T. Rose. RFC 1202: Directory assistance service, February

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1202.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1202.txt. Status: INFORMA-

TIONAL.

RFC1203

[1193] J. Rice. RFC 1203: Interactive mail access protocol: Version 3,

February 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1064.
txt,   ftp://ftp.internic.net/rfc/rfc1203.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1064.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1203.txt. Obsoletes RFC1064 [1054]. Status: HISTORIC.

RFC1204

[1194] S. Yeh and D. Lee. RFC 1204: Message Posting Protocol (MPP),

February 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1204.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1204.txt. Status: EXPER-

IMENTAL.

RFC1205

[1195] P. Chmielewski. RFC 1205: 5250 Telnet interface, February 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1205.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1205.txt. Status: INFORMA-

TIONAL.

RFC1206

[1196] G. S. Malkin and A. N. Marine. RFC 1206: FYI on questions and answers: Answers to commonly asked “new Internet user” ques-

tions, February 1, 1991. URL ftp://ftp.internic.net/rfc/fyi4.
txt,   ftp://ftp.internic.net/rfc/rfc1177.txt,   ftp://ftp.
internic.net/rfc/rfc1206.txt,   ftp://ftp.internic.net/rfc/
rfc1325.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi4.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1177.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1206.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1325.txt. See also FYI0004 [?]. Obsoleted by FYI0004, RFC1325 [?, 1315]. Obsoletes RFC1177 [1167]. Status: INFORMATIONAL.

RFC1207

[1197] G. S. Malkin, A. N. Marine, and J. K. Reynolds. RFC 1207: FYI on questions and answers: Answers to commonly asked “experienced Internet user” questions, February 1, 1991. URL ftp://ftp.internic.net/rfc/

fyi7.txt,   ftp://ftp.internic.net/rfc/rfc1207.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/fyi7.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1207.txt. See also FYI0007 [?]. Status: INFORMA-TIONAL.

RFC1208

[1198] O. J. Jacobsen and D. C. Lynch. RFC 1208: Glossary of networking

terms, March 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1208.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1208.txt. Status:

INFORMATIONAL.

RFC1209

[1199] D. M. Piscitello and J. Lawrence. RFC 1209: Transmission of IP datagrams over the SMDS service, March 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1209.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1209.txt. See also STD0052 [?]. Status: STANDARD.

RFC1210

[1200] V. G. Cerf, P. T. Kirstein, and B. Randell. RFC 1210: Network and infrastructure user requirements for transatlantic research collaboration: Brussels, July 16–18, and Washington, July 24–25, 1990,

March 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1210.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1210.txt. Status: INFOR-

MATIONAL.

RFC1211

[1201] A. Westine and J. Postel. RFC 1211: Problems with the maintenance of large mailing lists, March 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1211.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1211.

txt. Status: INFORMATIONAL.

RFC1212

[1202] M. T. Rose and K. McCloghrie. RFC 1212: Concise MIB definitions,

March 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1212.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1212.txt. See also STD0016

[?]. Status: STANDARD.

RFC1213

[1203] K. McCloghrie and M. T. Rose. RFC 1213: Management information base for network management of TCP/IP-based internets:MIB-

II, March 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1158.
txt,   ftp://ftp.internic.net/rfc/rfc1213.txt,   ftp://ftp.
internic.net/rfc/rfc2011.txt,   ftp://ftp.internic.net/rfc/
rfc2012.txt,   ftp://ftp.internic.net/rfc/rfc2013.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1158.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1213.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2011.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2012.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2013.txt. Obsoletes RFC1158

[1148]. See also STD0017 [?]. Updated by RFC2011, RFC2012, RFC2013 [1998, 1999, 2000]. Status: STANDARD.

RFC1214

[1204] L. LaBarre. RFC 1214: OSI Internet management: Management Information Base, April 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1214.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1214.

txt. Status: HISTORIC.

RFC1215

[1205] M. T. Rose. RFC 1215: Convention for defining traps for use with the

SNMP, March 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1215.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1215.txt. Status:

INFORMATIONAL.

RFC1216

[1206] P. Richard and P. Kynikos. RFC 1216: Gigabit network economics and paradigm shifts, April 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1216.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1216.

txt. Status: INFORMATIONAL.

RFC1217

[1207] V. G. Cerf. RFC 1217: Memo from the consortium for slow commotion research (CSCR), April 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1217.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1217.

txt. Status: INFORMATIONAL.

RFC1218

[1208] North American Directory Forum. RFC 1218: Naming scheme for

c=US, April 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1218.
txt,   ftp://ftp.internic.net/rfc/rfc1255.txt,   ftp://ftp.
internic.net/rfc/rfc1417.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1218.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1255.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1417.txt. Obsoleted

by RFC1255, RFC1417 [1245, 1407]. Status: INFORMATIONAL.

RFC1219

[1209] P. F. Tsuchiya. RFC 1219: On the assignment of subnet numbers,

April 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1219.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1219.txt. Status: INFOR-

MATIONAL.

RFC1220

[1210] F. Baker. RFC 1220: Point-to-point protocol extensions for bridg-

ing, April 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1220.
txt,   ftp://ftp.internic.net/rfc/rfc1638.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1220.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1638.txt. Obsoleted by RFC1638 [1628]. Status: PRO-POSED STANDARD.

RFC1221

[1211] W. Edmond. RFC 1221: Host access protocol (HAP) specification: Version 2, April 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1221.txt,   ftp://ftp.internic.net/rfc/rfc907.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc1221.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc907.txt. Updates RFC0907 [897]. Status: IN-FORMATIONAL.

RFC1222

[1212] H. W. Braun and Y. Rekhter. RFC 1222: Advancing the NSFNET routing architecture, May 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1222.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1222.

txt. Status: INFORMATIONAL.

RFC1223

[1213] J. M. Halpern. RFC 1223: OSI CLNS and LLC1 protocols on network systems HYPERchannel, May 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1223.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1223.

txt. Status: INFORMATIONAL.

RFC1224

[1214] L. Steinberg. RFC 1224: Techniques for managing asynchronously gener-

ated alerts, May 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1224.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1224.txt. Status:

EXPERIMENTAL.

RFC1225

[1215] M. T. Rose. RFC 1225: Post office protocol: Version 3, May 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1081.txt,   ftp://
ftp.internic.net/rfc/rfc1225.txt,   ftp://ftp.internic.net/
rfc/rfc1460.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1081.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1225.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1460.txt. Obsoleted by RFC1460

[1450]. Obsoletes RFC1081 [1071]. Status: DRAFT STANDARD.

RFC1226

[1216] B. Kantor. RFC 1226: Internet protocol encapsulation of AX.25 frames,

May 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1226.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1226.txt. Status: EXPERI-

MENTAL.

RFC1227

[1217] M. T. Rose. RFC 1227: SNMP MUX protocol and MIB, May 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1227.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1227.txt. Status: HISTORIC.

RFC1228

[1218] G. Carpenter and B. Wijnen. RFC 1228: SNMP-DPI: Simple network management protocol distributed program interface, May 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1228.txt,   ftp://
ftp.internic.net/rfc/rfc1592.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1228.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1592.

txt. Obsoleted by RFC1592 [1582]. Status: EXPERIMENTAL.

RFC1229

[1219] K. McCloghrie. RFC 1229: Extensions to the generic-interface MIB, May

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1229.txt,   ftp://
ftp.internic.net/rfc/rfc1239.txt,   ftp://ftp.internic.net/
rfc/rfc1573.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1229.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1239.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1573.txt. Obsoleted by RFC1573

[1563]. Updated by RFC1239 [1229]. Status: PROPOSED STANDARD.

RFC1230

[1220] K. McCloghrie and R. Fox. RFC 1230: IEEE 802.4 Token Bus

MIB, May 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1230.
txt,   ftp://ftp.internic.net/rfc/rfc1239.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1230.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1239.txt. Updated by RFC1239 [1229]. Status: HIS-TORIC.

RFC1231

[1221] K. McCloghrie, R. Fox, and E. Decker. RFC 1231: IEEE 802.5 Token

Ring MIB, May 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1231.
txt,   ftp://ftp.internic.net/rfc/rfc1239.txt,   ftp://ftp.
internic.net/rfc/rfc1743.txt,   ftp://ftp.internic.net/rfc/
rfc1748.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1231.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1239.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1743.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1748.txt. Obsoleted by RFC1743, RFC1748 [1733, 1738]. Updated by RFC1239 [1229]. Status: PROPOSED STANDARD.

RFC1232

[1222] F. Baker and C. P. Kolb. RFC 1232: Definitions of managed objects for the DS1 interface type, May 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1232.txt,   ftp://ftp.internic.net/rfc/
rfc1239.txt,   ftp://ftp.internic.net/rfc/rfc1406.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1232.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1239.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1406.txt. Obsoleted by RFC1406 [1396]. Updated by RFC1239 [1229]. Status: PROPOSED STANDARD.

RFC1233

[1223] T. A. Cox and K. Tesink. RFC 1233: Definitions of managed objects for the DS3 interface type, May 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1233.txt,   ftp://ftp.internic.net/rfc/
rfc1239.txt,   ftp://ftp.internic.net/rfc/rfc1407.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1233.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1239.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1407.txt. Obsoleted by RFC1407 [1397]. Updated by RFC1239 [1229]. Status: PROPOSED STANDARD.

RFC1234

[1224] D. Provan. RFC 1234: Tunneling IPX traffic through IP networks, June

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1234.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1234.txt. Status: PROPOSED

STANDARD.

RFC1235

[1225] J. Ioannidis and G. Q. Maguire, Jr. RFC 1235: Coherent file distribution protocol, June 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1235.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1235.

txt. Status: EXPERIMENTAL.

RFC1236

[1226] L. F. Morales, Jr. and P. R. Hasse. RFC 1236: IP to X.121 address mapping for DDN, June 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1236.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1236.

txt. Status: INFORMATIONAL.

RFC1237

[1227] R. Collela, E. P. Gardner, and R. W. Callon. RFC 1237: Guidelines for OSI NSAP allocation in the Internet, July 1, 1991. URL ftp://

ftp.internic.net/rfc/rfc1237.txt,   ftp://ftp.internic.net/
rfc/rfc1629.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1237.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1629.txt. Obsoleted

by RFC1629 [1619]. Status: PROPOSED STANDARD.

RFC1238

[1228] G. Satz. RFC 1238: CLNS MIB for use with connectionless network protocol (ISO 8473) and end system to intermediate system (ISO

9542), June 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1162.
txt,   ftp://ftp.internic.net/rfc/rfc1238.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1162.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1238.txt. Obsoletes RFC1162 [1152]. Status: EXPERI-MENTAL.

RFC1239

[1229] J. K. Reynolds. RFC 1239: Reassignment of experimental MIBs to

standard MIBs, June 1, 1991. URL ftp://ftp.internic.net/rfc/
rfc1229.txt,   ftp://ftp.internic.net/rfc/rfc1230.txt,   ftp:
//ftp.internic.net/rfc/rfc1231.txt,   ftp://ftp.internic.net/
rfc/rfc1232.txt,   ftp://ftp.internic.net/rfc/rfc1233.txt,
ftp://ftp.internic.net/rfc/rfc1239.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1229.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1230.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1231.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1232.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1233.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1239.txt. Updates RFC1229, RFC1230, RFC1231, RFC1232, RFC1233 [1219, 1220, 1221, 1222, 1223]. Status: PROPOSED STAN-DARD.

RFC1240

[1230] C. Shue, W. Haggerty, and K. Dobbins. RFC 1240: OSI connectionless transport services on top of UDP: Version 1, June 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1240.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1240.txt. Status: PROPOSED

STANDARD.

RFC1241

[1231] R. A. Woodburn and D. L. Mills. RFC 1241: Scheme for an Internet Encapsulation Protocol: Version 1, July 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1241.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1241.txt. Status: EXPERIMENTAL.

RFC1242

[1232] S. Bradner. RFC 1242: Benchmarking terminology for network interconnection devices, July 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1242.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1242.

txt. Status: INFORMATIONAL.

RFC1243

[1233] S. Waldbusser. RFC 1243: AppleTalk Management Information

Base, July 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1243.
txt,   ftp://ftp.internic.net/rfc/rfc1742.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1243.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1742.txt. Obsoleted by RFC1742 [1732]. Status: PRO-POSED STANDARD.

RFC1244

[1234] J. P. Holbrook and J. K. Reynolds. RFC 1244: Site security

handbook, July 1, 1991. URL ftp://ftp.internic.net/rfc/fyi8.
txt,   ftp://ftp.internic.net/rfc/rfc1244.txt,   ftp://ftp.
internic.net/rfc/rfc2196.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi8.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1244.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2196.txt. See also FYI0008

[?]. Obsoleted by RFC2196, FYI0008 [2183, ?]. Status: INFORMA-TIONAL.

RFC1245

[1235] J. Moy. RFC 1245: OSPF protocol analysis, July 1, 1991. URL ftp://

ftp.internic.net/rfc/rfc1245.txt,   ftp://ftp.internic.net/
rfc/rfc1246.txt,   ftp://ftp.internic.net/rfc/rfc1247.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1245.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1246.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1247.txt. See also RFC1247, RFC1246 [1237, 1236]. Status: INFORMATIONAL.

RFC1246

[1236] J. Moy. RFC 1246: Experience with the OSPF protocol, July 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1245.txt,   ftp://
ftp.internic.net/rfc/rfc1246.txt,   ftp://ftp.internic.net/
rfc/rfc1247.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1245.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1246.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1247.txt.  See also RFC1247,

RFC1245 [1237, 1235]. Status: INFORMATIONAL.

RFC1247

[1237] J. Moy. RFC 1247: OSPF version 2, July 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1131.txt,   ftp://ftp.internic.net/rfc/
rfc1245.txt,   ftp://ftp.internic.net/rfc/rfc1246.txt,   ftp:
//ftp.internic.net/rfc/rfc1247.txt,   ftp://ftp.internic.net/
rfc/rfc1583.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1131.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1245.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1246.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1247.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1583.txt. See also RFC1246, RFC1245 [1236, 1235]. Obsoleted by RFC1583 [1573]. Obsoletes RFC1131 [1121]. Status: DRAFT STAN-DARD.

RFC1248

[1238] F. Baker and R. Coltun. RFC 1248: OSPF version 2 management information base, July 1, 1991. URL ftp://ftp.internic.net/rfc/

rfc1248.txt,   ftp://ftp.internic.net/rfc/rfc1252.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1248.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1252.txt. Obsoleted by RFC1252 [1242]. Status: PROPOSED STANDARD.

RFC1249

[1239] T. Howes, M. Smith, and B. Beecher. RFC 1249: DIXIE protocol specifi-

cation, August 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1202.
txt,   ftp://ftp.internic.net/rfc/rfc1249.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1202.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1249.txt. See also RFC1202 [1192]. Status: INFORMA-TIONAL.

RFC1250

[1240] J. Postel. RFC 1250: IAB OFFICIAL PROTOCOL STANDARDS,

August 1, 1991.  URL ftp://ftp.internic.net/rfc/rfc1011.
txt,   ftp://ftp.internic.net/rfc/rfc1060.txt,   ftp://ftp.
internic.net/rfc/rfc1083.txt,   ftp://ftp.internic.net/rfc/
rfc1100.txt,   ftp://ftp.internic.net/rfc/rfc1130.txt,   ftp:
//ftp.internic.net/rfc/rfc1140.txt,   ftp://ftp.internic.net/
rfc/rfc1160.txt,   ftp://ftp.internic.net/rfc/rfc1200.txt,
ftp://ftp.internic.net/rfc/rfc1250.txt,   ftp://ftp.internic.
net/rfc/rfc2200.txt,   ftp://ftp.internic.net/rfc/rfc2300.txt,
 ftp://ftp.internic.net/rfc/rfc2400.txt, ftp://ftp.internic.
net/rfc/std1.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1011.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1060.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1083.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1100.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1130.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1140.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1160.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1250.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. See also RFC1060, RFC1011, RFC1160 [1050, 1001, 1150]. Obsoleted by RFC2200, RFC2300, STD0001, RFC2400 [2187, 2287, ?, 2387]. Obsoletes RFC1200, RFC1100, RFC1083, RFC1130, RFC1140 [1190, 1090, 1073, 1120, 1130]. Status: HISTORIC.

RFC1251

[1241] G. Malkin. RFC 1251: Who’s who in the Internet: Biographies of IAB, IESG and IRSG members, August 1, 1991. URL ftp://ftp.internic.

net/rfc/fyi9.txt,   ftp://ftp.internic.net/rfc/rfc1251.txt,
 ftp://ftp.math.utah.edu/pub/rfc/fyi9.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc1251.txt. See also FYI0009 [?]. Status: INFOR-MATIONAL.

RFC1252

[1242] F. Baker and R. Coltun. RFC 1252: OSPF version 2 management information base, August 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1245.txt,   ftp://ftp.internic.net/rfc/
rfc1247.txt,   ftp://ftp.internic.net/rfc/rfc1248.txt,   ftp:
//ftp.internic.net/rfc/rfc1252.txt,   ftp://ftp.internic.net/
rfc/rfc1253.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1245.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1247.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1248.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1252.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1253.txt. See also RFC1247, RFC1245 [1237, 1235]. Obsoleted by RFC1253 [1243]. Obsoletes RFC1248 [1238]. Status: PROPOSED STANDARD.

RFC1253

[1243] F. Baker and R. Coltun. RFC 1253: OSPF version 2 management information base, August 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1245.txt,   ftp://ftp.internic.net/rfc/
rfc1246.txt,   ftp://ftp.internic.net/rfc/rfc1247.txt,   ftp:
//ftp.internic.net/rfc/rfc1252.txt,   ftp://ftp.internic.net/
rfc/rfc1253.txt,   ftp://ftp.internic.net/rfc/rfc1850.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1245.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1246.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1247.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1252.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1253.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1850.txt.  See also RFC1247,

RFC1245, RFC1246 [1237, 1235, 1236]. Obsoleted by RFC1850 [1837]. Obsoletes RFC1252 [1242]. Status: PROPOSED STANDARD.

RFC1254

[1244] A. Mankin and K. Ramakrishnan. RFC 1254: Gateway congestion con-

trol survey, July 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1254.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1254.txt. Status:

INFORMATIONAL.

RFC1255

[1245] The North American Directory Forum. RFC 1255: A naming scheme for c=US, September 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1218.txt,   ftp://ftp.internic.net/rfc/
rfc1255.txt,   ftp://ftp.internic.net/rfc/rfc1417.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1218.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1255.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1417.txt. Obsoleted by RFC1417 [1407]. Obsoletes RFC1218 [1208]. Status: INFORMATIONAL.

RFC1256

[1246] S. Deering. RFC 1256: ICMP router discovery messages, September

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1256.txt,   ftp://
ftp.internic.net/rfc/rfc792.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1256.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc792.

txt. See also RFC0792 [782]. Status: PROPOSED STANDARD.

RFC1257

[1247] C. Partridge. RFC 1257: Isochronous applications do not require jittercontrolled networks, September 1, 1991. URL ftp://ftp.internic.

net/rfc/rfc1257.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1257.

txt. Status: INFORMATIONAL.

RFC1258

[1248] B. Kantor. RFC 1258: BSD rlogin, September 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1258.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1258.txt. Status: INFORMATIONAL.

RFC1259

[1249] M. Kapor. RFC 1259: Building the open road: The NREN as test-bed for the national public network, September 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1259.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1259.txt. Status: INFORMATIONAL.

RFC1260

[1250] ???? RFC 1260, 19xx. URL ftp://ftp.internic.net/rfc/rfc1260.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1260.txt. Status:

UNKNOWN. Not online.

RFC1261

[1251] S. Williamson and L. Nobile. RFC 1261: Transition of NIC services,

September 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1261.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc1261.txt. Status: INFOR-

MATIONAL.

RFC1262

[1252] V. G. Cerf. RFC 1262: Guidelines for Internet measurement activities,

October 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1262.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1262.txt. Status: INFOR-

MATIONAL.

RFC1263

[1253] S. O’Malley and L. L. Peterson. RFC 1263: TCP extensions considered harmful, October 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1263.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1263.

txt. Status: INFORMATIONAL.

RFC1264

[1254] R. M. Hinden. RFC 1264: Internet Engineering Task Force Internet Routing Protocol standardization criteria, October 1, 1991. URL ftp://

ftp.internic.net/rfc/rfc1264.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1264.txt. Status: INFORMATIONAL.

RFC1265

[1255] Y. Rekhter. RFC 1265: BGP protocol analysis, October 1,

1991. URL ftp://ftp.internic.net/rfc/rfc1265.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1265.txt. Status: INFORMA-

TIONAL.

RFC1266

[1256] Y. Rekhter. RFC 1266: Experience with the BGP protocol, Oc-

tober 1, 1991. URL ftp://ftp.internic.net/rfc/rfc1266.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1266.txt. Status: INFOR-

MATIONAL.

RFC1267

[1257] K. Lougheed and Y. Rekhter. RFC 1267: Border Gateway Protocol

3 (BGP-3), October 1, 1991. URL ftp://ftp.internic.net/rfc/
rfc1163.txt,   ftp://ftp.internic.net/rfc/rfc1267.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1163.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1267.txt. Obsoletes RFC1163 [1153]. Status: HIS-TORIC.

RFC1268

[1258] Y. Rekhter and P. Gross. RFC 1268: Application of the Border Gateway Protocol in the Internet, October 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1164.txt,   ftp://ftp.internic.net/rfc/
rfc1268.txt,   ftp://ftp.internic.net/rfc/rfc1655.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1164.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1268.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1655.txt. Obsoleted by RFC1655 [1645]. Obsoletes RFC1164 [1154]. Status: HISTORIC.

RFC1269

[1259] S. Willis and J. W. Burruss. RFC 1269: Definitions of managed objects for the Border Gateway Protocol: Version 3, October

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1269.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1269.txt. Status: PROPOSED

STANDARD.

RFC1270

[1260] F. Kastenholz. RFC 1270: SNMP communications services, October

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1270.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1270.txt. Status: INFORMA-

TIONAL.

RFC1271

[1261] S. Waldbusser. RFC 1271: Remote network monitoring Management Information Base, November 1, 1991. URL ftp://ftp.

internic.net/rfc/rfc1271.txt,   ftp://ftp.internic.net/rfc/
rfc1757.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1271.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1757.txt. Obsoleted by

RFC1757 [1747]. Status: PROPOSED STANDARD.

RFC1272

[1262] C. Mills, D. Hirsh, and G. R. Ruth. RFC 1272: Internet accounting: Background, November 1, 1991. URL ftp://ftp.internic.net/

rfc/rfc1272.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1272.

txt. Status: INFORMATIONAL.

RFC1273

[1263] M. F. Schwartz. RFC 1273: Measurement study of changes in service-level reachability in the global TCP/IP Internet: Goals, experimental design, implementation, and policy considerations, November

1, 1991. URL ftp://ftp.internic.net/rfc/rfc1273.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1273.txt. Status: INFORMA-

TIONAL.

RFC1274

[1264] P. Barker and S. Kille. RFC 1274: The COSINE and Internet X.500 schema, November 1991. URL ftp://ftp.internic.net/

rfc/rfc1274.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1274.

txt. Status: PROPOSED STANDARD.

RFC1275

[1265] S. E. Hardcastle-Kille. RFC 1275: Replication requirements to provide an Internet Directory using X.500, November 1991. URL ftp://ftp.

internic.net/rfc/rfc1275.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1275.txt. Status: INFORMATIONAL.

RFC1276

[1266] S. E. Hardcastle-Kille. RFC 1276: Replication and distributed operations extensions to provide an Internet Directory using X.500, Novem-

ber 1991. URL ftp://ftp.internic.net/rfc/rfc1276.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1276.txt. Status: PROPOSED

STANDARD.

RFC1277

[1267] S. E. Hardcastle-Kille. RFC 1277: Encoding network addresses to support operation over non-OSI lower layers, November 1991. URL ftp://

ftp.internic.net/rfc/rfc1277.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1277.txt. Status: PROPOSED STANDARD.

RFC1278

[1268] S. E. Hardcastle-Kille. RFC 1278: A string encoding of presentation ad-

dress, November 1991. URL ftp://ftp.internic.net/rfc/rfc1278.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1278.txt. Status:

INFORMATIONAL.

RFC1279

[1269] S. E. Hardcastle-Kille. RFC 1279: X.500 and domains, November

1991. URL ftp://ftp.internic.net/rfc/rfc1279.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1279.txt. Status: EXPERIMEN-

TAL.

RFC1280

[1270] J. Postel. RFC 1280: IAB OFFICIAL PROTOCOL STAN-

DARDS, March 1992. URL ftp://ftp.internic.net/rfc/rfc1280.
txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.
internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/
rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1280.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std1.txt. Obsoleted by RFC2200, ,

RFC2300, STD0001, RFC2400 [2187, 2287, ?, 2387]. Status: HISTORIC.

RFC1281

[1271] R. Pethia, S. Crocker, and B. Fraser. RFC 1281: Guidelines for the secure operation of the Internet, November 1991. URL ftp://ftp.internic.

net/rfc/rfc1281.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1281.

txt. Status: INFORMATIONAL.

RFC1282

[1272] B. Kantor. RFC 1282: BSD Rlogin, December 1991. URL ftp://ftp.

internic.net/rfc/rfc1282.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1282.txt. Status: INFORMATIONAL.

RFC1283

[1273] M. Rose. RFC 1283: SNMP over OSI, December 1991. URL ftp://

ftp.internic.net/rfc/rfc1283.txt,   ftp://ftp.internic.net/
rfc/rfc1418.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1283.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1418.txt. Obsoleted

by RFC1418 [1408]. Status: EXPERIMENTAL.

RFC1284

[1274] J. Cook. RFC 1284: Definitions of managed objects for the Ethernet-like interface types, December 1991. URL ftp://ftp.

internic.net/rfc/rfc1284.txt,   ftp://ftp.internic.net/rfc/
rfc1398.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1284.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1398.txt. Obsoleted by

RFC1398 [1388]. Status: PROPOSED STANDARD.

RFC1285

[1275] J. Case. RFC 1285: FDDI management information base, January

1992. URL ftp://ftp.internic.net/rfc/rfc1285.txt,   ftp://
ftp.internic.net/rfc/rfc1512.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1285.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1512.

txt. Updated by RFC1512 [1502]. Status: PROPOSED STANDARD.

RFC1286

[1276] E. Decker, P. Langille, A. Rijsinghani, and K. McCloghrie. RFC 1286: Definitions of managed objects for bridges, December 1991. URL ftp://

ftp.internic.net/rfc/rfc1286.txt,   ftp://ftp.internic.net/
rfc/rfc1493.txt,   ftp://ftp.internic.net/rfc/rfc1525.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc1286.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1493.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1525.txt. Obsoleted by RFC1493, RFC1525 [1483, 1515]. Status: PROPOSED STANDARD.

RFC1287

[1277] D. Clark, L. Chapin, V. Cerf, R. Braden, and R. Hobby. RFC 1287: Towards the future Internet architecture, December 1991. URL ftp://ftp.

internic.net/rfc/rfc1287.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1287.txt. Status: INFORMATIONAL.

RFC1288

[1278] D. Zimmerman. RFC 1288: The finger user information proto-

col, December 1991. URL ftp://ftp.internic.net/rfc/rfc1194.
txt,   ftp://ftp.internic.net/rfc/rfc1196.txt,   ftp://ftp.
internic.net/rfc/rfc1288.txt,   ftp://ftp.internic.net/rfc/
rfc742.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1194.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1196.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1288.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc742.txt. Obsoletes RFC1196, RFC1194, RFC0742 [1186, 1184, 732]. Status: DRAFT STANDARD.

RFC1289

[1279] J. Saperia. RFC 1289: DECnet Phase IV MIB extensions, December

1991. URL ftp://ftp.internic.net/rfc/rfc1289.txt,   ftp://
ftp.internic.net/rfc/rfc1559.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1289.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1559.

txt. Obsoleted by RFC1559 [1549]. Status: PROPOSED STANDARD.

RFC1290

[1280] J. Martin. RFC 1290: There’s gold in them thar networks! or searching for treasure in all the wrong places, December 1991. URL ftp:

//ftp.internic.net/rfc/fyi10.txt,   ftp://ftp.internic.net/
rfc/rfc1290.txt,   ftp://ftp.internic.net/rfc/rfc1402.txt,
ftp://ftp.math.utah.edu/pub/rfc/fyi10.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1290.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1402.txt. See also FYI0010 [?]. Obsoleted by RFC1402, FYI0010 [1392, ?]. Status: INFORMATIONAL.

RFC1291

[1281] V. Aggarwal. RFC 1291: Mid-level networks potential technical services,

December 1991. URL ftp://ftp.internic.net/rfc/rfc1291.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1291.txt. Status: INFOR-

MATIONAL.

RFC1292

[1282] R. Lang and R. Wright. RFC 1292: A catalog of available X.500 imple-

mentations, January 1992. URL ftp://ftp.internic.net/rfc/fyi11.
txt,   ftp://ftp.internic.net/rfc/rfc1292.txt,   ftp://ftp.
internic.net/rfc/rfc1632.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi11.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1292.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1632.txt. Obsoleted by

RFC1632, FYI0011 [1622, ?]. Status: INFORMATIONAL.

RFC1293

[1283] T. Bradley and C. Brown. RFC 1293: Inverse address resolution pro-

tocol, January 1992. URL ftp://ftp.internic.net/rfc/rfc1293.
txt,   ftp://ftp.internic.net/rfc/rfc2390.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1293.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2390.txt. Obsoleted by RFC2390 [2377]. Status: PRO-POSED STANDARD.

RFC1294

[1284] T. Bradley, C. Brown, and A. Malis. RFC 1294: Multiprotocol interconnect over frame relay, January 1992. URL ftp://ftp.

internic.net/rfc/rfc1294.txt,   ftp://ftp.internic.net/rfc/
rfc1490.txt,   ftp://ftp.internic.net/rfc/rfc2427.txt, ftp:
//ftp.internic.net/rfc/std55.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1490.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2427.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std55.txt. Obsoleted by RFC1490,

RFC2427, STD0055 [1480, 2414, ?]. Status: PROPOSED STANDARD.

RFC1295

[1285] The North American Directory Forum. RFC 1295: User Bill of Rights for entries and listings in the Public Directory, January 1992. URL ftp://

ftp.internic.net/rfc/rfc1295.txt,   ftp://ftp.internic.net/
rfc/rfc1417.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1295.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1417.txt. Obsoleted

by RFC1417 [1407]. Status: INFORMATIONAL.

RFC1296

[1286] M. Lottor. RFC 1296: Internet growth (1981–1991), January

1992. URL ftp://ftp.internic.net/rfc/rfc1296.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1296.txt. Status: INFORMA-

TIONAL.

RFC1297

[1287] D. Johnson. RFC 1297: NOC internal integrated trouble ticket system functional specification wishlist (“NOC TT REQUIREMENTS”),

January 1992. URL ftp://ftp.internic.net/rfc/rfc1297.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1297.txt. Status: INFOR-

MATIONAL.

RFC1298

[1288] R. Wormley and S. Bostock. RFC 1298: SNMP over IPX, February

1992. URL ftp://ftp.internic.net/rfc/rfc1298.txt,   ftp://
ftp.internic.net/rfc/rfc1420.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1298.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1420.

txt. Obsoleted by RFC1420 [1410]. Status: INFORMATIONAL.

RFC1299

[1289] M. Kennedy. RFC 1299: Summary of 1200–1299, January 1997. URL

ftp://ftp.internic.net/rfc/rfc1299.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1299.txt. Status: INFORMATIONAL.

RFC1300

[1290] S. Greenfield. RFC 1300: Remembrances of things past, Febru-

ary 1992. URL ftp://ftp.internic.net/rfc/rfc1300.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1300.txt. Status: INFORMA-

TIONAL.

RFC1301

[1291] S. Armstrong, A. Freier, and K. Marzullo. RFC 1301: Multicast transport protocol, February 1992. URL ftp://ftp.internic.net/

rfc/rfc1301.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1301.

txt. Status: INFORMATIONAL.

RFC1302

[1292] D. Sitzler, P. Smith, and A. Marine. RFC 1302: Building a network information services infrastructure, February 1992. URL ftp:

//ftp.internic.net/rfc/fyi12.txt,   ftp://ftp.internic.net/
rfc/rfc1302.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi12.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1302.txt. See also FYI0012

[?]. Status: INFORMATIONAL.

RFC1303

[1293] K. McCloghrie and M. Rose. RFC 1303: A convention for describing SNMP-based agents, February 1992. URL ftp://ftp.

internic.net/rfc/rfc1155.txt,   ftp://ftp.internic.net/rfc/

rfc1157.txt,   ftp://ftp.internic.net/rfc/rfc1212.txt,   ftp:
//ftp.internic.net/rfc/rfc1213.txt,   ftp://ftp.internic.net/
rfc/rfc1303.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1155.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1157.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1212.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1213.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1303.txt. See also RFC1155, RFC1212, RFC1213, RFC1157 [1145, 1202, 1203, 1147]. Status: INFORMATIONAL.

RFC1304

[1294] T. Cox and K. Tesinks. RFC 1304: Definitions of managed objects for the SIP interface type, February 1992. URL ftp://ftp.

internic.net/rfc/rfc1304.txt,   ftp://ftp.internic.net/rfc/
rfc1694.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1304.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1694.txt. Obsoleted by

RFC1694 [1684]. Status: PROPOSED STANDARD.

RFC1305

[1295] David L. Mills. RFC 1305: Network time protocol (version 3) specification, implementation, March 1992. URL ftp://ftp.

internic.net/rfc/rfc1059.txt,   ftp://ftp.internic.net/rfc/
rfc1119.txt,   ftp://ftp.internic.net/rfc/rfc1305.txt,   ftp:
//ftp.internic.net/rfc/rfc958.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1059.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1119.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1305.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc958.txt.  Obsoletes RFC0958,

RFC1059, RFC1119 [948, 1049, 1109]. Status: DRAFT STANDARD.

RFC1306

[1296] A. Nicholson and J. Young. RFC 1306: Experiences supporting byrequest circuit-switched T3 networks, March 1992. URL ftp://ftp.

internic.net/rfc/rfc1306.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1306.txt. Status: INFORMATIONAL.

RFC1307

[1297] J. Young and A. Nicholson. RFC 1307: Dynamically switched link control

protocol, March 1992. URL ftp://ftp.internic.net/rfc/rfc1307.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1307.txt. Status:

EXPERIMENTAL.

RFC1308

[1298] C. Weider and J. Reynolds. RFC 1308: Executive introduction to directory services using the X.500 protocol, March 1992. URL ftp://ftp.

internic.net/rfc/rfc1308.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1308.txt. Status: INFORMATIONAL.

RFC1309

[1299] C. Weider, J. Reynolds, and S. Heker. RFC 1309: Technical overview of directory services using the X.500 protocol, March

1992. URL ftp://ftp.internic.net/rfc/rfc1309.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1309.txt. Status: INFORMA-

TIONAL.

RFC1310

[1300] Lyman Chapin. RFC 1310: The Internet standards process, March

1992. URL ftp://ftp.internic.net/rfc/rfc1310.txt,   ftp://
ftp.internic.net/rfc/rfc1602.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1310.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1602.

txt. Obsoleted by RFC1602 [1592]. Status: INFORMATIONAL.

RFC1311

[1301] J. Postel. RFC 1311: Introduction to the STD notes, March

1992. URL ftp://ftp.internic.net/rfc/rfc1311.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1311.txt. Status: INFORMA-

TIONAL.

RFC1312

[1302] R. Nelson and G. Arnold. RFC 1312: Message send protocol 2, April

1992. URL ftp://ftp.internic.net/rfc/rfc1312.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1312.txt. Status: EXPERIMEN-

TAL.

RFC1313

[1303] C. Partridge. RFC 1313: Today’s programming for KRFC AM 1313 Internet talk radio, April 1992. URL ftp://ftp.internic.net/

rfc/rfc1313.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1313.

txt. Status: INFORMATIONAL.

RFC1314

[1304] A. Katz and D. Cohen. RFC 1314: A file format for the exchange of images in the Internet, April 1992. URL ftp://ftp.internic.net/

rfc/rfc1314.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1314.

txt. Status: PROPOSED STANDARD.

RFC1315

[1305] C. Brown, F. Baker, and C. Carvalho. RFC 1315: Management information base for frame relay DTEs, April 1992. URL ftp://

ftp.internic.net/rfc/rfc1315.txt,   ftp://ftp.internic.net/
rfc/rfc2115.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1315.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2115.txt. Obsoleted

by RFC2115 [2102]. Status: PROPOSED STANDARD.

RFC1316

[1306] B. Stewart. RFC 1316: Definitions of managed objects for character stream devices, April 1992. URL ftp://ftp.internic.net/rfc/

rfc1316.txt,   ftp://ftp.internic.net/rfc/rfc1658.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1316.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1658.txt. Obsoleted by RFC1658 [1648]. Status: PROPOSED STANDARD.

RFC1317

[1307] B. Stewart. RFC 1317: Definitions of managed objects for RS-232-like hardware devices, April 1992. URL ftp://ftp.internic.net/rfc/

rfc1317.txt,   ftp://ftp.internic.net/rfc/rfc1659.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1317.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1659.txt. Obsoleted by RFC1659 [1649]. Status: PROPOSED STANDARD.

RFC1318

[1308] B. Stewart. RFC 1318: Definitions of managed objects for parallelprinter-like hardware devices, April 1992. URL ftp://ftp.internic.

net/rfc/rfc1318.txt,   ftp://ftp.internic.net/rfc/rfc1660.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1318.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc1660.txt. Obsoleted by RFC1660 [1650]. Status: PROPOSED STANDARD.

RFC1319

[1309] B. Kaliski. RFC 1319: The MD2 message-digest algorithm, April

1992. URL ftp://ftp.internic.net/rfc/rfc1319.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1319.txt. Status: INFORMA-

TIONAL.

RFC1320

[1310] R. Rivest. RFC 1320: The MD4 message-digest algorithm, April

1992. URL ftp://ftp.internic.net/rfc/rfc1320.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1320.txt. Status: INFORMA-

TIONAL.

RFC1321

[1311] R. Rivest. RFC 1321: The MD5 message-digest algorithm, April

1992. URL ftp://ftp.internic.net/rfc/rfc1321.txt,   ftp:/

/ftp.math.utah.edu/pub/rfc/rfc1321.txt. Status: INFORMA-

TIONAL.

RFC1322

[1312] D. Estrin, Y. Rekhter, and S. Hotz. RFC 1322: A unified approach to inter-domain routing, May 1992. URL ftp://ftp.internic.net/

rfc/rfc1322.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1322.

txt. Status: INFORMATIONAL.

RFC1323

[1313] V. Jacobson, R. Braden, and D. Borman. RFC 1323: TCP extensions for high performance, May 1992. URL ftp://ftp.

internic.net/rfc/rfc1072.txt,   ftp://ftp.internic.net/rfc/
rfc1185.txt,   ftp://ftp.internic.net/rfc/rfc1323.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1072.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1185.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1323.txt. Obsoletes RFC1072, RFC1185 [1062, 1175]. Status: PRO-POSED STANDARD.

RFC1324

[1314] D. Reed. RFC 1324: A discussion on computer network conferenc-

ing, May 1992. URL ftp://ftp.internic.net/rfc/rfc1324.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1324.txt. Status: INFOR-

MATIONAL.

RFC1325

[1315] G. Malkin and A. Marine. RFC 1325: FYI on questions and answers: Answers to commonly asked “New Internet User” questions, May

1992. URL ftp://ftp.internic.net/rfc/fyi4.txt,   ftp://ftp.
internic.net/rfc/rfc1206.txt,   ftp://ftp.internic.net/rfc/
rfc1325.txt,   ftp://ftp.internic.net/rfc/rfc1594.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/fyi4.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1206.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1325.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1594.txt.

See also FYI0004 [?]. Obsoleted by RFC1594, FYI0004 [1584, ?]. Obsoletes RFC1206 [1196]. Status: INFORMATIONAL.

RFC1326

[1316] P. Tsuchiya. RFC 1326: Mutual encapsulation considered danger-

ous, May 1992. URL ftp://ftp.internic.net/rfc/rfc1326.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1326.txt. Status: INFOR-

MATIONAL.

RFC1327

[1317] S. Hardcastle-Kille. RFC 1327: Mapping between X.400(1988) /ISO

10021 and RFC 822, May 1992. URL ftp://ftp.internic.net/rfc/
rfc1026.txt,   ftp://ftp.internic.net/rfc/rfc1138.txt,   ftp:
//ftp.internic.net/rfc/rfc1148.txt,   ftp://ftp.internic.net/
rfc/rfc1327.txt,   ftp://ftp.internic.net/rfc/rfc1495.txt,
ftp://ftp.internic.net/rfc/rfc2156.txt,   ftp://ftp.internic.
net/rfc/rfc822.txt, ftp://ftp.internic.net/rfc/rfc987.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1026.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1138.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1148.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1327.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1495.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2156.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc822.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc987.

txt. Obsoleted by RFC1495, RFC2156 [1485, 2143]. Obsoletes RFC987, RFC1026, RFC1138, RFC1148 [977, 1016, 1128, 1138]. Updates RFC0822, RFC0822 [812, 812]. Status: PROPOSED STANDARD.

RFC1328

[1318] S. Hardcastle-Kille. RFC 1328: X.400 1988 to 1984 downgrading,

May 1992. URL ftp://ftp.internic.net/rfc/rfc1328.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1328.txt. Status: PROPOSED

STANDARD.

RFC1329

[1319] P. Kuehn. RFC 1329: Thoughts on address resolution for dual MAC FDDI networks, May 1992. URL ftp://ftp.internic.net/

rfc/rfc1329.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1329.

txt. Status: INFORMATIONAL.

RFC1330

[1320] ESCC X.500/X.400 Task Force and ESnet Site Coordinating Committee (ESCC). RFC 1330: Recommendations for the phase I deployment of OSI directory services (X.500) and OSI Message Handling Services (X.400) within the ESNET community, May

1992. URL ftp://ftp.internic.net/rfc/rfc1330.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1330.txt. Status: INFORMA-

TIONAL.

RFC1331

[1321] W. Simpson. RFC 1331: The point-to-point protocol (PPP) for the transmission of multi-protocol datagrams over point-to-point

links, May 1992. URL ftp://ftp.internic.net/rfc/rfc1171.
txt,   ftp://ftp.internic.net/rfc/rfc1172.txt,   ftp://ftp.
internic.net/rfc/rfc1331.txt,   ftp://ftp.internic.net/rfc/
rfc1548.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1171.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1172.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1331.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1548.txt. Obsoleted by RFC1548 [1538]. Obsoletes RFC1171, RFC1172 [1161, 1162]. Status: PROPOSED STANDARD.

RFC1332

[1322] G. McGregor. RFC 1332: The PPP Internet Protocol Control Proto-

col (IPCP), May 1992. URL ftp://ftp.internic.net/rfc/rfc1332.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1332.txt. Obsoletes

RFC1172 [1162]. Status: PROPOSED STANDARD.

RFC1333

[1323] W. Simpson. RFC 1333: PPP link quality monitoring, May

1992. URL ftp://ftp.internic.net/rfc/rfc1333.txt,   ftp://
ftp.internic.net/rfc/rfc1989.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1333.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1989.

txt. Obsoleted by RFC1989 [1976]. Status: PROPOSED STANDARD.

RFC1334

[1324] B. Lloyd and W. Simpson. RFC 1334: PPP authentication proto-

cols, October 1992. URL ftp://ftp.internic.net/rfc/rfc1334.
txt,   ftp://ftp.internic.net/rfc/rfc1994.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1334.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1994.txt. Obsoleted by RFC1994 [1981]. Status: PRO-POSED STANDARD.

RFC1335

[1325] Z. Wang and J. Crowcroft. RFC 1335: A two-tier address structure for the Internet: A solution to the problem of address space exhaus-

tion, May 1992. URL ftp://ftp.internic.net/rfc/rfc1335.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1335.txt. Status: INFOR-

MATIONAL.

RFC1336

[1326] G. Malkin. RFC 1336: Who’s who in the Internet: Biographies of IAB, IESG and IRSG members, May 1992. URL ftp://ftp.internic.net/

rfc/rfc1336.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1336.

txt. Status: INFORMATIONAL.

RFC1337

[1327] R. Braden. RFC 1337: TIME-WAIT assassination hazards in TCP,

May 1992. URL ftp://ftp.internic.net/rfc/rfc1337.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1337.txt. Status: INFORMA-

TIONAL.

RFC1338

[1328] V. Fuller, T. Li, J. Yu, and K. Varadhan. RFC 1338: Supernetting: an address assignment and aggregation strategy, June 1992. URL ftp://

ftp.internic.net/rfc/rfc1338.txt,   ftp://ftp.internic.net/
rfc/rfc1519.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1338.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1519.txt. Obsoleted

by RFC1519 [1509]. Status: INFORMATIONAL.

RFC1339

[1329] S. Dorner and P. Resnick. RFC 1339: Remote mail checking proto-

col, June 1992. URL ftp://ftp.internic.net/rfc/rfc1339.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1339.txt. Status: EXPERI-

MENTAL.

RFC1340

[1330] J. Reynolds and J. Postel. RFC 1340: ASSIGNED NUMBERS, July

1992. URL ftp://ftp.internic.net/rfc/rfc1060.txt,   ftp://
ftp.internic.net/rfc/rfc1340.txt,   ftp://ftp.internic.net/
rfc/rfc1700.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1060.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1340.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1700.txt. Obsoleted by RFC1700

[1690]. Obsoletes RFC1060 [1050]. Status: HISTORIC.

RFC1341

[1331] N. Borenstein and N. Freed. RFC 1341: MIME (Multipurpose Internet Mail Extensions): Mechanisms for specifying and describing the format of Internet message bodies, June 1992. URL ftp://

ftp.internic.net/rfc/rfc1341.txt,   ftp://ftp.internic.net/
rfc/rfc1521.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1341.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1521.txt. Obsoleted

by RFC1521 [1511]. Status: PROPOSED STANDARD.

RFC1342

[1332] K. Moore. RFC 1342: Representation of non-ASCII text in Internet mes-

sage headers, June 1992. URL ftp://ftp.internic.net/rfc/rfc1342.
txt,   ftp://ftp.internic.net/rfc/rfc1522.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1342.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1522.txt. Obsoleted by RFC1522 [1512]. Status: INFOR-MATIONAL.

RFC1343

[1333] N. Borenstein. RFC 1343: A user agent configuration mechanism for multimedia mail format information, June 1992. URL ftp://ftp.

internic.net/rfc/rfc1343.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1343.txt. Status: INFORMATIONAL.

RFC1344

[1334] N. Borenstein. RFC 1344: Implications of MIME for Internet mail gate-

ways, June 1992. URL ftp://ftp.internic.net/rfc/rfc1344.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1344.txt. Status: INFOR-

MATIONAL.

RFC1345

[1335] K. Simonsen. RFC 1345: Character mnemonics and character sets,

June 1992. URL ftp://ftp.internic.net/rfc/rfc1345.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1345.txt. Status: INFORMA-

TIONAL.

RFC1346

[1336] P. Jones. RFC 1346: Resource allocation, control, and accounting for the use of network resources, June 1992. URL ftp://ftp.internic.net/

rfc/rfc1346.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1346.

txt. Status: INFORMATIONAL.

RFC1347

[1337] R. Callon. RFC 1347: TCP and UDP with bigger addresses (TUBA), A simple proposal for Internet addressing and routing,

June 1992. URL ftp://ftp.internic.net/rfc/rfc1347.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1347.txt. Status: INFORMA-

TIONAL.

RFC1348

[1338] B. Manning. RFC 1348: DNS NSAP RRs, July 1992. URL ftp://ftp.

internic.net/rfc/rfc1034.txt,   ftp://ftp.internic.net/rfc/
rfc1035.txt,   ftp://ftp.internic.net/rfc/rfc1348.txt,   ftp:
//ftp.internic.net/rfc/rfc1637.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1034.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1348.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1637.txt. Obsoleted by RFC1637

[1627]. Updates RFC1034, RFC1035 [1024, 1025]. Updated by RFC1637 [1627]. Status: EXPERIMENTAL.

RFC1349

[1339] P. Almquist. RFC 1349: Type of service in the Internet Protocol suite,

July 1992. URL ftp://ftp.internic.net/rfc/rfc1349.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1349.txt. Status: PROPOSED

STANDARD.

RFC1350

[1340] K. Sollins. RFC 1350: THE TFTP PROTOCOL (REVISION 2), July

1992. URL ftp://ftp.internic.net/rfc/rfc1350.txt,   ftp://
ftp.internic.net/rfc/rfc1782.txt,   ftp://ftp.internic.net/
rfc/rfc1783.txt,   ftp://ftp.internic.net/rfc/rfc1784.txt,
ftp://ftp.internic.net/rfc/rfc1785.txt,   ftp://ftp.internic.
net/rfc/rfc2347.txt,   ftp://ftp.internic.net/rfc/rfc2348.txt,
 ftp://ftp.internic.net/rfc/rfc2349.txt,   ftp://ftp.internic.
net/rfc/rfc784.txt, ftp://ftp.internic.net/rfc/std33.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1350.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1782.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1783.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1784.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1785.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2347.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2348.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2349.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc784.txt,
ftp://ftp.math.utah.edu/pub/rfc/std33.txt. See also STD0033 [?].

Obsoletes RFC0784 [774]. Updates RFC1350, RFC1783 [1340, 1773]. Updated by RFC1782, RFC1783, RFC1784, RFC1350, RFC1785, RFC2347, RFC2348, RFC2349 [1772, 1773, 1774, 1340, 1775, 2334, 2335, 2336]. Status: STANDARD.

RFC1351

[1341] J. Davin, J. Galvin, and K. McCloghrie. RFC 1351: SNMP administra-

tive model, July 1992. URL ftp://ftp.internic.net/rfc/rfc1351.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1351.txt. Status:

PROPOSED STANDARD.

RFC1352

[1342] J. Galvin, K. McCloghrie, and J. Davin. RFC 1352: SNMP security pro-

tocols, July 1992. URL ftp://ftp.internic.net/rfc/rfc1352.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1352.txt. Status: PRO-

POSED STANDARD.

RFC1353

[1343] K. McCloghrie, J. Davin, and J. Galvin. RFC 1353: Definitions of managed objects for administration of SNMP parties, July

1992. URL ftp://ftp.internic.net/rfc/rfc1353.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1353.txt. Status: PROPOSED

STANDARD.

RFC1354

[1344] F. Baker. RFC 1354: IP forwarding table MIB, July 1992. URL ftp://

ftp.internic.net/rfc/rfc1354.txt,   ftp://ftp.internic.net/
rfc/rfc2096.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1354.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2096.txt. Obsoleted

by RFC2096 [2083]. Status: PROPOSED STANDARD.

RFC1355

[1345] J. Curran and A. Marine. RFC 1355: Privacy and accuracy issues in network information center databases, August 1992. URL ftp:

//ftp.internic.net/rfc/fyi15.txt,   ftp://ftp.internic.net/
rfc/rfc1355.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi15.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1355.txt. See also FYI0015

[?]. Status: INFORMATIONAL.

RFC1356

[1346] A. Malis, D. Robinson, and R. Ullmann. RFC 1356: Multiprotocol interconnect on X.25 and ISDN in the packet mode, August

1992. URL ftp://ftp.internic.net/rfc/rfc1356.txt,   ftp://
ftp.internic.net/rfc/rfc877.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1356.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc877.

txt. Obsoletes RFC0877 [867]. Status: DRAFT STANDARD.

RFC1357

[1347] D. Cohen. RFC 1357: A format for E-mailing bibliographic

records, July 1992. URL ftp://ftp.internic.net/rfc/rfc1357.
txt,   ftp://ftp.internic.net/rfc/rfc1807.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1357.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1807.txt. Obsoleted by RFC1807 [1797]. Status: INFOR-MATIONAL.

RFC1358

[1348] L. Chapin. RFC 1358: Charter of the Internet Architecture Board

(IAB), August 1992. URL ftp://ftp.internic.net/rfc/rfc1358.
txt,   ftp://ftp.internic.net/rfc/rfc1601.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1358.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1601.txt. Obsoleted by RFC1601 [1591]. Status: INFOR-MATIONAL.

RFC1359

[1349] ACM SIGUCCS. RFC 1359: Connecting to the Internet — what connecting institutions should anticipate, August 1992. URL ftp:

//ftp.internic.net/rfc/fyi16.txt,   ftp://ftp.internic.net/
rfc/rfc1359.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi16.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1359.txt. See also FYI0016

[?]. Status: INFORMATIONAL.

RFC1360

[1350] J. Postel. RFC 1360: IAB official protocol standards, September

1992. URL ftp://ftp.internic.net/rfc/rfc1280.txt,   ftp://
ftp.internic.net/rfc/rfc1360.txt,   ftp://ftp.internic.net/
rfc/rfc1410.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1280.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1360.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1410.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. See also RFC1280 [1270]. Obsoleted by RFC1410, RFC2200, RFC2300, STD0001, RFC2400 [1400, 2187, 2287, ?, 2387]. Status: HISTORIC.

RFC1361

[1351] D. Mills. RFC 1361: Simple network time protocol (SNTP), August

1992. URL ftp://ftp.internic.net/rfc/rfc1305.txt,   ftp://
ftp.internic.net/rfc/rfc1361.txt,   ftp://ftp.internic.net/
rfc/rfc1769.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1305.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1361.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1769.txt. See also RFC1305 [1295].

Obsoleted by RFC1769 [1759]. Status: INFORMATIONAL.

RFC1362

[1352] M. Allen. RFC 1362: Novell IPX over various WAN media (IPXWAN),

September 1992.  URL ftp://ftp.internic.net/rfc/rfc1362.
txt,   ftp://ftp.internic.net/rfc/rfc1634.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1362.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1634.txt. Obsoleted by RFC1634 [1624]. Status: INFOR-MATIONAL.

RFC1363

[1353] C. Partridge. RFC 1363: A proposed flow specification, Septem-

ber 1992. URL ftp://ftp.internic.net/rfc/rfc1363.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1363.txt. Status: INFORMA-

TIONAL.

RFC1364

[1354] K. Varadhan. RFC 1364: BGP OSPF interaction, September

1992. URL ftp://ftp.internic.net/rfc/rfc1247.txt,   ftp://
ftp.internic.net/rfc/rfc1267.txt,   ftp://ftp.internic.net/
rfc/rfc1364.txt,   ftp://ftp.internic.net/rfc/rfc1403.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1247.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1267.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1364.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1403.

txt. See also RFC1247, RFC1267 [1237, 1257]. Obsoleted by RFC1403 [1393]. Status: PROPOSED STANDARD.

RFC1365

[1355] K. Siyan. RFC 1365: An IP address extension proposal, Septem-

ber 1992. URL ftp://ftp.internic.net/rfc/rfc1365.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1365.txt. Status: INFORMA-

TIONAL.

RFC1366

[1356] E. Gerich. RFC 1366: Guidelines for management of IP address

space, October 1992. URL ftp://ftp.internic.net/rfc/rfc1366.
txt,   ftp://ftp.internic.net/rfc/rfc1466.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1366.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1466.txt. Obsoleted by RFC1466 [1456]. Status: INFOR-MATIONAL.

RFC1367

[1357] C. Topolcic. RFC 1367: Schedule for IP address space management guidelines, October 1992. URL ftp://ftp.internic.net/rfc/

rfc1367.txt,   ftp://ftp.internic.net/rfc/rfc1467.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1367.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1467.txt. Obsoleted by RFC1467 [1457]. Status: IN-FORMATIONAL.

RFC1368

[1358] D. McMaster and K. McCloghrie. RFC 1368: Definitions of managed objects for IEEE 802.3 repeater devices, October 1992. URL ftp://

ftp.internic.net/rfc/rfc1368.txt,   ftp://ftp.internic.net/

rfc/rfc1516.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1368.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1516.txt. Obsoleted

by RFC1516 [1506]. Status: PROPOSED STANDARD.

RFC1369

[1359] F. Kastenholz. RFC 1369: Implementation notes and experience for the Internet Ethernet MIB, October 1992. URL ftp://ftp.internic.net/

rfc/rfc1369.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1369.

txt. Status: PROPOSED STANDARD.

RFC1370

[1360] Internet Architecture Board and Lyman Chapin. RFC 1370: Applicability statement for OSPF, October 1992. URL ftp://ftp.internic.net/

rfc/rfc1370.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1370.

txt. Status: PROPOSED STANDARD.

RFC1371

[1361] P. Gross. RFC 1371: Choosing a common IGP for the IP Internet: (the IESG’s recommendation to the IAB), October 1992. URL ftp://ftp.

internic.net/rfc/rfc1371.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1371.txt. Status: INFORMATIONAL.

RFC1372

[1362] C. Hedrick and D. Borman. RFC 1372: Telnet Remote Flow Control

Option, October 1992. URL ftp://ftp.internic.net/rfc/rfc1080.
txt,   ftp://ftp.internic.net/rfc/rfc1372.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1080.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1372.txt. Obsoletes RFC1080 [1070]. Status: PROPOSED STANDARD.

RFC1373

[1363] T. Tignor. RFC 1373: Portable DUAs, October 1992. URL ftp://ftp.

internic.net/rfc/rfc1373.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1373.txt. Status: INFORMATIONAL.

RFC1374

[1364] J. Renwick and A. Nicholson. RFC 1374: IP and ARP on HIPPI, Octo-

ber 1992. URL ftp://ftp.internic.net/rfc/rfc1374.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1374.txt. Status: PROPOSED

STANDARD.

RFC1375

[1365] P. Robinson. RFC 1375: Suggestion for new classes of IP addresses,

October 1992. URL ftp://ftp.internic.net/rfc/rfc1375.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc1375.txt. Status: INFOR-

MATIONAL.

RFC1376

[1366] S. Senum. RFC 1376: The PPP DECnet phase IV control proto-

col (DNCP), November 1992. URL ftp://ftp.internic.net/rfc/
rfc1376.txt,   ftp://ftp.internic.net/rfc/rfc1762.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1376.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1762.txt. Obsoleted by RFC1762 [1752]. Status: PROPOSED STANDARD.

RFC1377

[1367] D. Katz. RFC 1377: The PPP OSI network layer control protocol (OSINLCP), November 1992. URL ftp://ftp.internic.net/

rfc/rfc1377.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1377.

txt. Status: PROPOSED STANDARD.

RFC1378

[1368] B. Parker. RFC 1378: The PPP AppleTalk Control Protocol (ATCP),

November 1992. URL ftp://ftp.internic.net/rfc/rfc1378.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1378.txt. Status: PRO-

POSED STANDARD.

RFC1379

[1369] R. Braden. RFC 1379: Extending TCP for transactions — concepts,

November 1992. URL ftp://ftp.internic.net/rfc/rfc1379.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1379.txt. Status: INFOR-

MATIONAL.

RFC1380

[1370] P. Gross and P. Almquist. RFC 1380: IESG deliberations on routing and addressing, November 1992. URL ftp://ftp.internic.net/

rfc/rfc1380.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1380.

txt. Status: INFORMATIONAL.

RFC1381

[1371] D. Throop and F. Baker. RFC 1381: SNMP MIB extension for X.25

LAPB, November 1992. URL ftp://ftp.internic.net/rfc/rfc1381.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1381.txt. Status:

PROPOSED STANDARD.

RFC1382

[1372] D. Throop. RFC 1382: SNMP MIB extension for the X.25 packet layer,

November 1992. URL ftp://ftp.internic.net/rfc/rfc1382.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc1382.txt. Status: PRO-

POSED STANDARD.

RFC1383

[1373] C. Huitema. RFC 1383: An experiment in DNS based IP routing, Decem-

ber 1992. URL ftp://ftp.internic.net/rfc/rfc1383.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1383.txt. Status: EXPERIMEN-

TAL.

RFC1384

[1374] P. Barker and S. E. Hardcastle-Kille. RFC 1384: Naming guidelines for directory pilots, February 1993. URL ftp://ftp.internic.net/rfc/

rfc1384.txt,   ftp://ftp.internic.net/rfc/rfc1617.txt, ftp:
//ftp.internic.net/rfc/rtr11.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1384.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1617.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rtr11.txt. Obsoleted by

RFC1617, RTR0011 [1607, ?]. Status: INFORMATIONAL.

RFC1385

[1375] Z. Wang. RFC 1385: EIP: The Extended Internet Protocol: A framework for maintaining backward compatibility, November

1992. URL ftp://ftp.internic.net/rfc/rfc1385.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1385.txt. Status: INFORMA-

TIONAL.

RFC1386

[1376] A. Cooper and J. Postel. RFC 1386: The US domain, December

1992. URL ftp://ftp.internic.net/rfc/rfc1386.txt,   ftp://
ftp.internic.net/rfc/rfc1480.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1386.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1480.

txt. Obsoleted by RFC1480 [1470]. Status: INFORMATIONAL.

RFC1387

[1377] G. Malkin. RFC 1387: RIP version 2 protocol analysis, January

1993. URL ftp://ftp.internic.net/rfc/rfc1387.txt,   ftp://
ftp.internic.net/rfc/rfc1721.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1387.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1721.

txt. Obsoleted by RFC1721 [1711]. Status: INFORMATIONAL.

RFC1388

[1378] G. Malkin. RFC 1388: RIP version 2 carrying additional informa-

tion, January 1993. URL ftp://ftp.internic.net/rfc/rfc1058.
txt,   ftp://ftp.internic.net/rfc/rfc1388.txt,   ftp://ftp.

internic.net/rfc/rfc1723.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1058.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1388.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1723.txt. Obsoleted

by RFC1723, RFC2453 [1713, 2440]. Updates RFC1058 [1048]. Updated by RFC2453 [2440], STD0056 [?]. Status: PROPOSED STANDARD.

RFC1389

[1379] G. Malkin and F. Baker. RFC 1389: RIP version 2 MIB exten-

sion, January 1993. URL ftp://ftp.internic.net/rfc/rfc1389.
txt,   ftp://ftp.internic.net/rfc/rfc1724.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1389.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1724.txt. Obsoleted by RFC1724 [1714]. Status: PRO-POSED STANDARD.

RFC1390

[1380] D. Katz. RFC 1390: Transmission of IP and ARP over FDDI networks,

January 1993. URL ftp://ftp.internic.net/rfc/rfc1390.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1390.txt. Status: STAN-

DARD.

RFC1391

[1381] G. Malkin. RFC 1391: The tao of IETF: A guide for new attendees of the Internet Engineering Task Force, January 1993. URL ftp:

//ftp.internic.net/rfc/fyi17.txt,   ftp://ftp.internic.net/
rfc/rfc1391.txt,   ftp://ftp.internic.net/rfc/rfc1539.txt,
ftp://ftp.math.utah.edu/pub/rfc/fyi17.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1391.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1539.txt. Obsoleted by RFC1539, FYI0017 [1529, ?]. See also FYI0017 [?]. Status: INFORMATIONAL.

RFC1392

[1382] G. Malkin and T. LaQuey Parker. RFC 1392: Internet users’

glossary, January 1993. URL ftp://ftp.internic.net/rfc/fyi18.
txt,   ftp://ftp.internic.net/rfc/rfc1392.txt,   ftp://ftp.
internic.net/rfc/rfc1983.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi18.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1392.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1983.txt. See also FYI0018

[?]. Obsoleted by RFC1983 [1970], FYI0018 [?]. Status: INFORMA-TIONAL.

RFC1393

[1383] G. Malkin. RFC 1393: Traceroute using an IP option, January

1993. URL ftp://ftp.internic.net/rfc/rfc1393.txt,   ftp://

ftp.math.utah.edu/pub/rfc/rfc1393.txt. Status: EXPERIMEN-

TAL.

RFC1394

[1384] P. Robinson. RFC 1394: Relationship of Telex answerback codes to Internet domains, January 1993. URL ftp://ftp.internic.net/

rfc/rfc1394.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1394.

txt. Status: INFORMATIONAL.

RFC1395

[1385] J. Reynolds. RFC 1395: BOOTP vendor information exten-

sions, January 1993. URL ftp://ftp.internic.net/rfc/rfc1048.
txt,   ftp://ftp.internic.net/rfc/rfc1084.txt,   ftp://ftp.
internic.net/rfc/rfc1395.txt,   ftp://ftp.internic.net/rfc/
rfc1497.txt,   ftp://ftp.internic.net/rfc/rfc1533.txt,   ftp:
//ftp.internic.net/rfc/rfc951.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1048.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1084.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1395.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1497.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1533.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc951.txt. Obsoleted by RFC1497, RFC1533 [1487, 1523]. Obsoletes RFC1084, RFC1048 [1074, 1038]. Updates RFC951, RFC0951 [941, 941]. Status: DRAFT STANDARD.

RFC1396

[1386] S. Crocker. RFC 1396: The Process for Organization of Internet Standards Working Group (POISED), January 1993. URL ftp://ftp.

internic.net/rfc/rfc1396.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1396.txt. Status: INFORMATIONAL.

RFC1397

[1387] D. Haskin. RFC 1397: Default route advertisement in BGP2 and BGP3 version of the Border Gateway Protocol, January 1993. URL ftp://ftp.

internic.net/rfc/rfc1397.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1397.txt. Status: PROPOSED STANDARD.

RFC1398

[1388] F. Kastenholz. RFC 1398: Definitions of managed objects for the Ethernet-like interface types, January 1993. URL ftp://ftp.

internic.net/rfc/rfc1284.txt,   ftp://ftp.internic.net/rfc/
rfc1398.txt,   ftp://ftp.internic.net/rfc/rfc1623.txt,   ftp:
//ftp.internic.net/rfc/rfc1643.txt, ftp://ftp.internic.net/
rfc/std50.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1284.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1398.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc1623.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1643.txt,   ftp://ftp.math.utah.edu/pub/rfc/std50.txt.

Obsoleted by RFC1643, STD0050, RFC1623 [1633, ?, 1613]. Obsoletes RFC1284 [1274]. Status: DRAFT STANDARD.

RFC1399

[1389] J. Elliott. RFC 1399: Summary of 1300–1399, January 1997. URL ftp:/

/ftp.internic.net/rfc/rfc1399.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1399.txt. Status: INFORMATIONAL.

RFC1400

[1390] S. Williamson. RFC 1400: Transition and modernization of the Internet registration service, March 1993. URL ftp://ftp.internic.net/

rfc/rfc1400.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1400.

txt. Status: INFORMATIONAL.

RFC1401

[1391] Internet Architecture Board. RFC 1401: Correspondence between the IAB and DISA on the use of DNS throughout the Internet, Jan-

uary 1993. URL ftp://ftp.internic.net/rfc/rfc1401.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1401.txt. Status: INFORMA-

TIONAL.

RFC1402

[1392] J. Martin. RFC 1402: There’s gold in them thar networks! or searching for treasure in all the wrong places, January 1993. URL ftp:

//ftp.internic.net/rfc/fyi10.txt,   ftp://ftp.internic.net/
rfc/rfc1290.txt,   ftp://ftp.internic.net/rfc/rfc1402.txt,
ftp://ftp.math.utah.edu/pub/rfc/fyi10.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1290.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1402.txt. See also FYI0010 [?]. Obsoletes RFC1290 [1280]. Status: INFORMATIONAL.

RFC1403

[1393] K. Varadhan. RFC 1403: BGP OSPF interaction, January 1993. URL

ftp://ftp.internic.net/rfc/rfc1364.txt,   ftp://ftp.internic.
net/rfc/rfc1403.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1364.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1403.txt. Obsoletes

RFC1364 [1354]. Status: PROPOSED STANDARD.

RFC1404

[1394] B. Stockman. RFC 1404: A model for common operational statis-

tics, January 1993. URL ftp://ftp.internic.net/rfc/rfc1404.

txt,   ftp://ftp.internic.net/rfc/rfc1857.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1404.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1857.txt. Obsoleted by RFC1857 [1844]. Status: INFOR-MATIONAL.

RFC1405

[1395] C. Allocchio. RFC 1405: Mapping between X.400(1984/1988) and Mail-11 (DECnet mail), January 1993. URL ftp://ftp.

internic.net/rfc/rfc1405.txt,   ftp://ftp.internic.net/rfc/
rfc2162.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1405.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2162.txt. Obsoleted by

RFC2162 [2149]. Status: EXPERIMENTAL.

RFC1406

[1396] F. Baker and J. Watts. RFC 1406: Definitions of managed objects for the DS1 and E1 interface types, January 1993. URL ftp://

ftp.internic.net/rfc/rfc1232.txt,   ftp://ftp.internic.net/
rfc/rfc1406.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1232.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1406.txt. Obsoletes

RFC1232 [1222]. Status: PROPOSED STANDARD.

RFC1407

[1397] Tracy A. Cox and Kaj Tesink. RFC 1407: Definitions of managed objects for the DS3/E3 interface type, January 1993. URL ftp://

ftp.internic.net/rfc/rfc1233.txt,   ftp://ftp.internic.net/
rfc/rfc1407.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1233.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1407.txt. Obsoletes

RFC1233 [1223]. Status: PROPOSED STANDARD.

RFC1408

[1398] D. Borman. RFC 1408: Telnet Environment Option, January

1993. URL ftp://ftp.internic.net/rfc/rfc1408.txt,   ftp://
ftp.internic.net/rfc/rfc1571.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1408.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1571.

txt. Updated by RFC1571 [1561]. Status: HISTORIC.

RFC1409

[1399] D. Borman. RFC 1409: Telnet Authentication Option, January

1993. URL ftp://ftp.internic.net/rfc/rfc1409.txt,   ftp://
ftp.internic.net/rfc/rfc1416.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1409.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1416.

txt. Obsoleted by RFC1416 [1406]. Status: EXPERIMENTAL.

RFC1410

[1400] J. Postel. RFC 1410: IAB official protocol standards, March

1993. URL ftp://ftp.internic.net/rfc/rfc1360.txt,   ftp://
ftp.internic.net/rfc/rfc1410.txt,   ftp://ftp.internic.net/
rfc/rfc1500.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1360.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1410.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1500.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. Obsoleted by RFC1500, RFC2200, RFC2300, STD0001, RFC2400 [1490, 2187, 2287, ?, 2387]. Obsoletes RFC1360 [1350]. Status: HISTORIC.

RFC1411

[1401] D. Borman. RFC 1411: Telnet Authentication: Kerberos Version 4, Jan-

uary 1993. URL ftp://ftp.internic.net/rfc/rfc1411.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1411.txt. Status: EXPERIMEN-

TAL.

RFC1412

[1402] K. Alagappan. RFC 1412: Telnet Authentication: SPX, January

1993. URL ftp://ftp.internic.net/rfc/rfc1412.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1412.txt. Status: EXPERIMEN-

TAL.

RFC1413

[1403] M. St. Johns. RFC 1413: Identification protocol, January 1993. URL

ftp://ftp.internic.net/rfc/rfc1413.txt,   ftp://ftp.internic.
net/rfc/rfc931.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1413.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc931.txt. Obsoletes

RFC0931 [921]. Status: PROPOSED STANDARD.

RFC1414

[1404] M. St. Johns and M. Rose. RFC 1414: Identification MIB, Jan-

uary 1993. URL ftp://ftp.internic.net/rfc/rfc1414.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1414.txt. Status: PROPOSED

STANDARD.

RFC1415

[1405] J. Mindel and R. Slaski. RFC 1415: FTP-FTAM gateway specifica-

tion, January 1993. URL ftp://ftp.internic.net/rfc/rfc1415.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1415.txt. Status: PRO-

POSED STANDARD.

RFC1416

[1406] D. Borman. RFC 1416: Telnet Authentication Option, February

1993. URL ftp://ftp.internic.net/rfc/rfc1409.txt,   ftp://
ftp.internic.net/rfc/rfc1416.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1409.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1416.

txt. Obsoletes RFC1409 [1399]. Status: EXPERIMENTAL.

RFC1417

[1407] The North American Directory Forum. RFC 1417: NADF standing documents: A brief overview, February 1993. URL ftp://ftp.

internic.net/rfc/rfc1218.txt,   ftp://ftp.internic.net/rfc/
rfc1255.txt,   ftp://ftp.internic.net/rfc/rfc1295.txt,   ftp:
//ftp.internic.net/rfc/rfc1417.txt,   ftp://ftp.internic.net/
rfc/rfc1758.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1218.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1255.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1295.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1417.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1758.txt. Obsoleted by RFC1758 [1748]. Obsoletes RFC1295, RFC1255, RFC1218 [1285, 1245, 1208]. Status: INFORMATIONAL.

RFC1418

[1408] M. Rose. RFC 1418: SNMP over OSI, March 1993. URL ftp://ftp.

internic.net/rfc/rfc1161.txt,   ftp://ftp.internic.net/rfc/
rfc1283.txt,   ftp://ftp.internic.net/rfc/rfc1418.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1161.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1283.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1418.txt. Obsoletes RFC1161, RFC1283 [1151, 1273]. Status: PRO-POSED STANDARD.

RFC1419

[1409] G. Minshall and M. Ritter. RFC 1419: SNMP over AppleTalk, March

1993. URL ftp://ftp.internic.net/rfc/rfc1419.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1419.txt. Status: PROPOSED

STANDARD.

RFC1420

[1410] S. Bostock. RFC 1420: SNMP over IPX, March 1993. URL ftp://

ftp.internic.net/rfc/rfc1298.txt,   ftp://ftp.internic.net/
rfc/rfc1420.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1298.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1420.txt. Obsoletes

RFC1298 [1288]. Status: PROPOSED STANDARD.

RFC1421

[1411] J. Linn. RFC 1421: Privacy enhancement for Internet electronic mail: Part I: Message encryption and authentication proce-

dures, February 1993. URL ftp://ftp.internic.net/rfc/rfc1113.
txt,   ftp://ftp.internic.net/rfc/rfc1421.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1113.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1421.txt. Obsoletes RFC1113 [1103]. Status: PROPOSED STANDARD.

RFC1422

[1412] S. Kent. RFC 1422: Privacy enhancement for Internet electronic mail: Part II: Certificate-based key management, February 1993. URL ftp://

ftp.internic.net/rfc/rfc1114.txt,   ftp://ftp.internic.net/
rfc/rfc1422.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1114.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1422.txt. Obsoletes

RFC1114 [1104]. Status: PROPOSED STANDARD.

RFC1423

[1413] D. Balenson. RFC 1423: Privacy enhancement for Internet electronic mail: Part III: Algorithms, modes, and identifiers, February

1993. URL ftp://ftp.internic.net/rfc/rfc1115.txt,   ftp://
ftp.internic.net/rfc/rfc1423.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1115.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1423.

txt. Obsoletes RFC1115 [1105]. Status: PROPOSED STANDARD.

RFC1424

[1414] B. Kaliski. RFC 1424: Privacy enhancement for Internet electronic mail: Part IV: Key certification and related services, Febru-

ary 1993. URL ftp://ftp.internic.net/rfc/rfc1424.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1424.txt. Status: PROPOSED

STANDARD.

RFC1425

[1415] J. Klensin, WG Chair, N. Freed, M. Rose, E. Stefferud, and D. Crocker. RFC 1425: SMTP service extensions, February 1993. URL ftp://

ftp.internic.net/rfc/rfc1425.txt,   ftp://ftp.internic.net/
rfc/rfc1651.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1425.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1651.txt. Obsoleted

by RFC1651 [1641]. Status: PROPOSED STANDARD.

RFC1426

[1416] J. Klensin, WG Chair, N. Freed, M. Rose, E. Stefferud, and D. Crocker. RFC 1426: SMTP service extension for 8bit-MIMEtransport, February

1993. URL ftp://ftp.internic.net/rfc/rfc1426.txt,   ftp://
ftp.internic.net/rfc/rfc1652.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1426.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1652.

txt. Obsoleted by RFC1652 [1642]. Status: PROPOSED STANDARD.

RFC1427

[1417] J. Klensin, WG Chair, N. Freed, and K. Moore. RFC 1427: SMTP service extension for message size declaration, February 1993. URL ftp://

ftp.internic.net/rfc/rfc1427.txt,   ftp://ftp.internic.net/
rfc/rfc1653.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1427.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1653.txt. Obsoleted

by RFC1653 [1643]. Status: PROPOSED STANDARD.

RFC1428

[1418] G. Vaudreuil. RFC 1428: Transition of Internet mail from just-send-8 to 8bit-SMTP/MIME, February 1993. URL ftp://ftp.internic.net/

rfc/rfc1428.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1428.

txt. Status: INFORMATIONAL.

RFC1429

[1419] E. Thomas. RFC 1429: Listserv distribute protocol, February

1993. URL ftp://ftp.internic.net/rfc/rfc1429.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1429.txt. Status: INFORMA-

TIONAL.

RFC1430

[1420] S. Hardcastle-Kille, E. Huizer, V. Cerf, R. Hobby, and S. Kent. RFC 1430: A strategic plan for deploying an Internet X.500 Directory Service,

February 1993. URL ftp://ftp.internic.net/rfc/rfc1430.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1430.txt. Status: INFOR-

MATIONAL.

RFC1431

[1421] P. Barker. RFC 1431: DUA metrics (OSI-DS 33 (v2)), Febru-

ary 1993. URL ftp://ftp.internic.net/rfc/rfc1431.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc1431.txt. Status: INFORMA-

TIONAL.

RFC1432

[1422] J. Quarterman. RFC 1432: Recent Internet books, March 1993. URL

ftp://ftp.internic.net/rfc/rfc1432.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1432.txt. Status: INFORMATIONAL.

RFC1433

[1423] J. Garrett, J. Hagan, and J. Wong. RFC 1433: Directed ARP, March

1993. URL ftp://ftp.internic.net/rfc/rfc1433.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1433.txt. Status: EXPERIMEN-

TAL.

RFC1434

[1424] R. Dixon and D. Kushi. RFC 1434: Data link switching: Switch-to-switch protocol, March 1993. URL ftp://ftp.internic.net/rfc/

rfc1434.txt,   ftp://ftp.internic.net/rfc/rfc1795.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1434.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1795.txt. Obsoleted by RFC1795 [1785]. Status: IN-FORMATIONAL.

RFC1435

[1425] S. Knowles. RFC 1435: IESG advice from experience with path MTU

discovery, March 1993. URL ftp://ftp.internic.net/rfc/rfc1435.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1435.txt. Status:

INFORMATIONAL.

RFC1436

[1426] F. Anklesaria, M. McCahill, P. Lindner, D. Johnson, D. Torrey, and B. Albert. RFC 1436: The Internet Gopher Protocol (a distributed document search and retrieval protocol), March

1993. URL ftp://ftp.internic.net/rfc/rfc1436.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1436.txt. Status: INFORMA-

TIONAL.

RFC1437

[1427] N. Borenstein and M. Linimon. RFC 1437: The extension of MIME content-types to a new medium, April 1993. URL ftp://ftp.internic.

net/rfc/rfc1437.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1437.

txt. Status: INFORMATIONAL.

RFC1438

[1428] A. Lyman Chapin and C. Huitema. RFC 1438: Internet Engineering Task Force statements of boredom (SOBs), April 1993. URL ftp://ftp.

internic.net/rfc/rfc1438.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1438.txt. Status: INFORMATIONAL.

RFC1439

[1429] C. Finseth. RFC 1439: The uniqueness of unique identifiers, March

1993. URL ftp://ftp.internic.net/rfc/rfc1439.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1439.txt. Status: INFORMA-

TIONAL.

RFC1440

[1430] R. Troth. RFC 1440: SIFT/UFT: Sender-initiated/unsolicited file trans-

fer, July 1993. URL ftp://ftp.internic.net/rfc/rfc1440.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1440.txt. Status: EXPERI-

MENTAL.

RFC1441

[1431] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1441: Introduction to version 2 of the Internet-standard Network Management

Framework, April 1993. URL ftp://ftp.internic.net/rfc/rfc1441.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1441.txt. Status:

PROPOSED STANDARD.

RFC1442

[1432] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1442: Structure of management information for version 2 of the Simple Network Management Protocol (SNMPv2), April 1993. URL ftp://

ftp.internic.net/rfc/rfc1442.txt,   ftp://ftp.internic.net/
rfc/rfc1902.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1442.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1902.txt. Obsoleted

by RFC1902 [1889]. Status: PROPOSED STANDARD.

RFC1443

[1433] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1443: Textual conventions for version 2 of the Simple Network Management Protocol (SNMPv2), April 1993. URL ftp://ftp.

internic.net/rfc/rfc1443.txt,   ftp://ftp.internic.net/rfc/
rfc1903.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1443.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1903.txt. Obsoleted by

RFC1903 [1890]. Status: PROPOSED STANDARD.

RFC1444

[1434] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1444: Conformance statements for version 2 of the Simple Network Management Protocol (SNMPv2), April 1993. URL ftp://

ftp.internic.net/rfc/rfc1444.txt,   ftp://ftp.internic.net/
rfc/rfc1904.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1444.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1904.txt. Obsoleted

by RFC1904 [1891]. Status: PROPOSED STANDARD.

RFC1445

[1435] J. Galvin and K. McCloghrie. RFC 1445: Administrative model for version 2 of the Simple Network Management Protocol (SNMPv2), April

1993. URL ftp://ftp.internic.net/rfc/rfc1445.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1445.txt. Status: HISTORIC.

RFC1446

[1436] J. Galvin and K. McCloghrie. RFC 1446: Security protocols for version 2 of the Simple Network Management Protocol (SNMPv2), April

1993. URL ftp://ftp.internic.net/rfc/rfc1446.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1446.txt. Status: HISTORIC.

RFC1447

[1437] K. McCloghrie and J. Galvin. RFC 1447: Party MIB for version 2 of the simple network management protocol (SNMPv2), April

1993. URL ftp://ftp.internic.net/rfc/rfc1447.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1447.txt. Status: HISTORIC.

RFC1448

[1438] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1448: Protocol operations for version 2 of the Simple Network Management Protocol (SNMPv2), April 1993. URL ftp://ftp.

internic.net/rfc/rfc1448.txt,   ftp://ftp.internic.net/rfc/
rfc1905.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1448.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1905.txt. Obsoleted by

RFC1905 [1892]. Status: PROPOSED STANDARD.

RFC1449

[1439] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1449: Transport mappings for version 2 of the Simple Network Management Protocol (SNMPv2), April 1993. URL ftp://ftp.

internic.net/rfc/rfc1449.txt,   ftp://ftp.internic.net/rfc/
rfc1906.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1449.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1906.txt. Obsoleted by

RFC1906 [1893]. Status: PROPOSED STANDARD.

RFC1450

[1440] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1450: Management information base for version 2 of the Simple Network Management Protocol (SNMPv2), April 1993. URL ftp://ftp.

internic.net/rfc/rfc1450.txt,   ftp://ftp.internic.net/rfc/
rfc1907.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1450.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1907.txt. Obsoleted by

RFC1907 [1894]. Status: PROPOSED STANDARD.

RFC1451

[1441] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1451: Manager-to-manager management information base, April

1993. URL ftp://ftp.internic.net/rfc/rfc1451.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1451.txt. Status: HISTORIC.

RFC1452

[1442] J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1452: Coexistence between version 1 and version 2 of the Internet-standard Network Management Framework, April 1993. URL ftp://ftp.

internic.net/rfc/rfc1452.txt,   ftp://ftp.internic.net/rfc/
rfc1908.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1452.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1908.txt. Obsoleted by

RFC1908 [1895]. Status: PROPOSED STANDARD.

RFC1453

[1443] W. Chimiak. RFC 1453: A comment on packet video remote conferencing and the transport/network layers, April 1993. URL ftp://ftp.

internic.net/rfc/rfc1453.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1453.txt. Status: INFORMATIONAL.

RFC1454

[1444] T. Dixon. RFC 1454: Comparison of proposals for next version of

IP, May 1993. URL ftp://ftp.internic.net/rfc/rfc1454.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1454.txt. Status: INFOR-

MATIONAL.

RFC1455

[1445] D. Eastlake, III. RFC 1455: Physical link security type of service,

May 1993. URL ftp://ftp.internic.net/rfc/rfc1455.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1455.txt. Obsoleted by RCF2474

[?]. Status: EXPERIMENTAL.

RFC1456

[1446] Vietnamese Standardization Working Group. RFC 1456: Conventions for encoding the Vietnamese language. VISCII: VIetnamese Standard Code for Information Interchange. VIQR: VIetnamese Quoted-Readable speci-

fication, May 1993. URL ftp://ftp.internic.net/rfc/rfc1456.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1456.txt. Status: INFOR-

MATIONAL.

RFC1457

[1447] R. Housley. RFC 1457: Security label framework for the Internet,

May 1993. URL ftp://ftp.internic.net/rfc/rfc1457.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1457.txt. Status: INFORMA-

TIONAL.

RFC1458

[1448] R. Braudes and S. Zabele. RFC 1458: Requirements for multicast pro-

tocols, May 1993. URL ftp://ftp.internic.net/rfc/rfc1458.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1458.txt. Status: INFOR-

MATIONAL.

RFC1459

[1449] J. Oikarinen and D. Reed. RFC 1459: Internet Relay Chat Protocol,

May 1993. URL ftp://ftp.internic.net/rfc/rfc1459.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1459.txt. Status: EXPERIMEN-

TAL.

RFC1460

[1450] M. Rose. RFC 1460: Post office protocol — version 3, June

1993. URL ftp://ftp.internic.net/rfc/rfc1225.txt,   ftp://
ftp.internic.net/rfc/rfc1460.txt,   ftp://ftp.internic.net/
rfc/rfc1725.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1225.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1460.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1725.txt. Obsoleted by RFC1725

[1715]. Obsoletes RFC1225 [1215]. Status: DRAFT STANDARD.

RFC1461

[1451] D. Throop. RFC 1461: SNMP MIB extension for multiprotocol interconnect over X.25, May 1993. URL ftp://ftp.internic.net/

rfc/rfc1461.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1461.

txt. Status: PROPOSED STANDARD.

RFC1462

[1452] E. Krol and E. Hoffman. RFC 1462: FYI on “What is the In-

ternet?”, May 1993. URL ftp://ftp.internic.net/rfc/fyi20.

txt,   ftp://ftp.internic.net/rfc/rfc1462.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi20.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1462.txt. See also FYI0020 [?]. Status: INFORMA-TIONAL.

RFC1463

[1453] E. Hoffman and L. Jackson. RFC 1463: FYI on introducing the Internet — A short bibliography of introductory internetworking readings for the

network novice, May 1993. URL ftp://ftp.internic.net/rfc/fyi19.
txt,   ftp://ftp.internic.net/rfc/rfc1463.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi19.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1463.txt. See also FYI0019 [?]. Status: INFORMA-TIONAL.

RFC1464

[1454] R. Rosenbaum. RFC 1464: Using the domain name system to store arbitrary string attributes, May 1993. URL ftp://ftp.internic.net/

rfc/rfc1464.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1464.

txt. Status: EXPERIMENTAL.

RFC1465

[1455] D. Eppenberger. RFC 1465: Routing coordination for X.400 MHS services within a multi protocol /multi network environment table format V3 for static routing, May 1993. URL ftp://ftp.internic.net/

rfc/rfc1465.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1465.

txt. Status: EXPERIMENTAL.

RFC1466

[1456] E. Gerich. RFC 1466: Guidelines for management of IP address

space, May 1993. URL ftp://ftp.internic.net/rfc/rfc1366.
txt,   ftp://ftp.internic.net/rfc/rfc1466.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1366.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1466.txt. Obsoletes RFC1366 [1356]. Status: INFORMA-TIONAL.

RFC1467

[1457] C. Topolcic. RFC 1467: Status of CIDR deployment in the In-

ternet, August 1993. URL ftp://ftp.internic.net/rfc/rfc1367.
txt,   ftp://ftp.internic.net/rfc/rfc1467.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1367.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1467.txt. Obsoletes RFC1367 [1357]. Status: INFORMA-TIONAL.

RFC1468

[1458] J. Murai, M. Crispin, and E. van der Poel. RFC 1468: Japanese character encoding for Internet messages, June 1993. URL ftp://ftp.internic.

net/rfc/rfc1468.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1468.

txt. Status: INFORMATIONAL.

RFC1469

[1459] T. Pusateri. RFC 1469: IP multicast over token-ring local area networks,

June 1993. URL ftp://ftp.internic.net/rfc/rfc1469.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1469.txt. Status: PROPOSED

STANDARD.

RFC1470

[1460] R. Enger and J. Reynolds. RFC 1470: FYI on a network management tool catalog: Tools for monitoring and debugging TCP/IP internets and interconnected devices, June 1993. URL ftp://ftp.internic.

net/rfc/fyi2.txt,   ftp://ftp.internic.net/rfc/rfc1147.txt,
 ftp://ftp.internic.net/rfc/rfc1470.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi2.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1147.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1470.txt. See also

FYI0002 [?]. Obsoletes RFC1147 [1137]. Status: INFORMATIONAL.

RFC1471

[1461] F. Kastenholz. RFC 1471: The definitions of managed objects for the link control protocol of the point-to-point protocol, June

1993. URL ftp://ftp.internic.net/rfc/rfc1471.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1471.txt. Status: PROPOSED

STANDARD.

RFC1472

[1462] F. Kastenholz. RFC 1472: The definitions of managed objects for the security protocols of the point-to-point protocol, June 1993. URL ftp://

ftp.internic.net/rfc/rfc1472.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1472.txt. Status: PROPOSED STANDARD.

RFC1473

[1463] F. Kastenholz. RFC 1473: The definitions of managed objects for the IP network control protocol of the point-to-point protocol,

June 1993. URL ftp://ftp.internic.net/rfc/rfc1473.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1473.txt. Status: PROPOSED

STANDARD.

RFC1474

[1464] F. Kastenholz. RFC 1474: The definitions of managed objects for the bridge network control protocol of the point-to-point protocol,

June 1993. URL ftp://ftp.internic.net/rfc/rfc1474.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1474.txt. Status: PROPOSED

STANDARD.

RFC1475

[1465] R. Ullmann. RFC 1475: TP/IX: The next Internet, June 1993. URL ftp:

//ftp.internic.net/rfc/rfc1475.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1475.txt. Status: EXPERIMENTAL.

RFC1476

[1466] R. Ullmann. RFC 1476: RAP: Internet Route Access Protocol, June

1993. URL ftp://ftp.internic.net/rfc/rfc1476.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1476.txt. Status: EXPERIMEN-

TAL.

RFC1477

[1467] M. Steenstrup. RFC 1477: IDPR as a proposed standard, July

1993. URL ftp://ftp.internic.net/rfc/rfc1477.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1477.txt. Status: PROPOSED

STANDARD.

RFC1478

[1468] M. Steenstrup. RFC 1478: An architecture for inter-domain policy

routing, July 1993. URL ftp://ftp.internic.net/rfc/rfc1478.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1478.txt. Status: PRO-

POSED STANDARD.

RFC1479

[1469] M. Steenstrup. RFC 1479: Inter-domain policy routing protocol specification: Version 1, July 1993. URL ftp://ftp.internic.net/

rfc/rfc1479.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1479.

txt. Status: PROPOSED STANDARD.

RFC1480

[1470] A. Cooper and J. Postel. RFC 1480: The US domain, June

1993. URL ftp://ftp.internic.net/rfc/rfc1386.txt,   ftp://
ftp.internic.net/rfc/rfc1480.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1386.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1480.

txt. Obsoletes RFC1386 [1376]. Status: INFORMATIONAL.

RFC1481

[1471] C. Huitema. RFC 1481: IAB recommendation for an intermediate strategy to address the issue of scaling, July 1993. URL ftp://ftp.

internic.net/rfc/rfc1481.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1481.txt. Status: INFORMATIONAL.

RFC1482

[1472] Mark Knopper and Steven J. Richardson. RFC 1482: Aggregation support in the NSFNET policy-based routing database, July

1993. URL ftp://ftp.internic.net/rfc/rfc1482.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1482.txt. Status: INFORMA-

TIONAL.

RFC1483

[1473] Juha Heinanen. RFC 1483: Multiprotocol encapsulation over ATM adaptation layer 5, July 1993. URL ftp://ftp.internic.net/

rfc/rfc1483.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1483.

txt. Status: PROPOSED STANDARD.

RFC1484

[1474] S. Hardcastle-Kille. RFC 1484: Using the OSI directory to achieve user friendly naming (OSI-DS 24 (v12)), July 1993. URL ftp://

ftp.internic.net/rfc/rfc1484.txt,   ftp://ftp.internic.net/
rfc/rfc1781.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1484.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1781.txt. Obsoleted

by RFC1781 [1771]. Updated by RFC1781 [1771]. Status: EXPERIMEN-TAL.

RFC1485

[1475] S. Hardcastle-Kille. RFC 1485: A string representation of distinguished names (OSI-DS 23 (v5)), July 1993. URL ftp://ftp.

internic.net/rfc/rfc1485.txt,   ftp://ftp.internic.net/rfc/
rfc1779.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1485.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1779.txt. Obsoleted by

RFC1779 [1769]. Status: PROPOSED STANDARD.

RFC1486

[1476] M. Rose and C. Malamud. RFC 1486: An experiment in remote

printing, August 1993. URL ftp://ftp.internic.net/rfc/rfc1486.
txt,   ftp://ftp.internic.net/rfc/rfc1528.txt,   ftp://ftp.
internic.net/rfc/rfc1529.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1486.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1528.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1529.txt. Obsoleted

by RFC1528, RFC1529 [1518, 1519]. Status: EXPERIMENTAL.

RFC1487

[1477] W. Yeong, T. Howes, and S. Kille. RFC 1487: X.500 Lightweight Directory Access Protocol, July 1993. URL ftp://ftp.internic.net/rfc/

rfc1487.txt,   ftp://ftp.internic.net/rfc/rfc1777.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1487.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1777.txt. Obsoleted by RFC1777 [1767]. Status: PROPOSED STANDARD.

RFC1488

[1478] T. Howes, S. Kille, W. Yeong, and C. Robbins. RFC 1488: The X.500 string representation of standard attribute syntaxes, July

1993. URL ftp://ftp.internic.net/rfc/rfc1488.txt,   ftp://
ftp.internic.net/rfc/rfc1778.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1488.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1778.

txt. Obsoleted by RFC1778 [1768]. Status: PROPOSED STANDARD.

RFC1489

[1479] A. Chernov. RFC 1489: Registration of a Cyrillic character set,

July 1993. URL ftp://ftp.internic.net/rfc/rfc1489.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1489.txt. Status: INFORMA-

TIONAL.

RFC1490

[1480] T. Bradley, C. Brown, and A. Malis. RFC 1490: Multiprotocol interconnect over frame relay, July 1993. URL ftp://ftp.

internic.net/rfc/rfc1294.txt,   ftp://ftp.internic.net/rfc/
rfc1490.txt,   ftp://ftp.internic.net/rfc/rfc2427.txt, ftp:
//ftp.internic.net/rfc/std55.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1490.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2427.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std55.txt. Obsoleted by RFC2427,

STD0055 [2414, ?]. Obsoletes RFC1294 [1284]. Status: DRAFT STAN-DARD.

RFC1491

[1481] C. Weider and R. Wright. RFC 1491: A survey of advanced us-

ages of X.500, July 1993. URL ftp://ftp.internic.net/rfc/fyi21.
txt,   ftp://ftp.internic.net/rfc/rfc1491.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi21.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1491.txt. See also FYI0021 [?]. Status: INFORMA-TIONAL.

RFC1492

[1482] C. Finseth. RFC 1492: An access control protocol, sometimes called

TACACS, July 1993. URL ftp://ftp.internic.net/rfc/rfc1492.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1492.txt. Status:

INFORMATIONAL.

RFC1493

[1483] E. Decker, P. Langille, A. Rijsinghani, and K. McCloghrie. RFC 1493: Definitions of managed objects for bridges, July 1993. URL ftp://

ftp.internic.net/rfc/rfc1286.txt,   ftp://ftp.internic.net/
rfc/rfc1493.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1286.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1493.txt. Obsoletes

RFC1286 [1276]. Status: DRAFT STANDARD.

RFC1494

[1484] H. Alvestrand and S. Thompson. RFC 1494: Equivalences between 1988 X.400 and RFC-822 message bodies, August 1993. URL ftp://ftp.

internic.net/rfc/rfc1494.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1494.txt. Status: PROPOSED STANDARD.

RFC1495

[1485] H. Alvestrand, S. Kille, R. Miles, M. Rose, and S. Thompson. RFC 1495: Mapping between X.400 and RFC-822 message bod-

ies, August 1993. URL ftp://ftp.internic.net/rfc/rfc1026.
txt,   ftp://ftp.internic.net/rfc/rfc1138.txt,   ftp://ftp.
internic.net/rfc/rfc1148.txt,   ftp://ftp.internic.net/rfc/
rfc1327.txt,   ftp://ftp.internic.net/rfc/rfc1495.txt,   ftp:
//ftp.internic.net/rfc/rfc2156.txt,   ftp://ftp.internic.net/
rfc/rfc987.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1026.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1138.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1148.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1327.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1495.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2156.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc987.txt. Obsoleted by RFC2156

[2143]. Obsoletes RFC987, RFC1026, RFC1138, RFC1148, RFC1327 [977, 1016, 1128, 1138, 1317]. Status: PROPOSED STANDARD.

RFC1496

[1486] H. Alvestrand, J. Romaguera, and K. Jordan. RFC 1496: Rules for downgrading messages from X.400/88 to X.400/84 when MIME content-types are present in the messages, August 1993. URL ftp://ftp.internic.

net/rfc/rfc1496.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1496.

txt. Status: PROPOSED STANDARD.

RFC1497

[1487] J. Reynolds. RFC 1497: BOOTP vendor information exten-

sions, August 1993. URL ftp://ftp.internic.net/rfc/rfc1048.
txt,   ftp://ftp.internic.net/rfc/rfc1084.txt,   ftp://ftp.
internic.net/rfc/rfc1395.txt,   ftp://ftp.internic.net/rfc/
rfc1497.txt,   ftp://ftp.internic.net/rfc/rfc1533.txt,   ftp:
//ftp.internic.net/rfc/rfc951.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1048.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1084.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1395.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1497.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1533.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc951.txt. Obsoleted by RFC1533 [1523]. Obsoletes RFC1395, RFC1084, RFC1048 [1385, 1074, 1038]. Updates RFC951, RFC0951 [941, 941]. Status: DRAFT STANDARD.

RFC1498

[1488] J. Saltzer. RFC 1498: On the naming and binding of network destina-

tions, August 1993. URL ftp://ftp.internic.net/rfc/rfc1498.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1498.txt. Status: INFOR-

MATIONAL.

RFC1499

[1489] J. Elliott. RFC 1499: Summary of 1400–1499, January 1997. URL ftp:/

/ftp.internic.net/rfc/rfc1499.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1499.txt. Status: INFORMATIONAL.

RFC1500

[1490] J. Postel. RFC 1500: Internet official protocol standards, August

1993. URL ftp://ftp.internic.net/rfc/rfc1410.txt,   ftp://
ftp.internic.net/rfc/rfc1500.txt,   ftp://ftp.internic.net/
rfc/rfc1540.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1410.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1500.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1540.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. Obsoleted by RFC1540, RFC2200, RFC2300, STD0001, RFC2400 [1530, 2187, 2287, ?, 2387]. Obsoletes RFC1410 [1400]. Status: HISTORIC.

RFC1501

[1491] E. Brunsen. RFC 1501: OS/2 user group, August 1993. URL ftp://ftp.

internic.net/rfc/rfc1501.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1501.txt. Status: INFORMATIONAL.

RFC1502

[1492] H. Alvestrand. RFC 1502: X.400 use of extended character sets, Au-

gust 1993. URL ftp://ftp.internic.net/rfc/rfc1502.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1502.txt. Status: PROPOSED

STANDARD.

RFC1503

[1493] K. McCloghrie and M. Rose. RFC 1503: Algorithms for automating administration in SNMPv2 managers, August 1993. URL ftp://ftp.

internic.net/rfc/rfc1503.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1503.txt. Status: INFORMATIONAL.

RFC1504

[1494] A. Oppenheimer. RFC 1504: AppleTalk Update-Based Routing Protocol: Enhanced AppleTalk routing, August 1993. URL ftp://ftp.

internic.net/rfc/rfc1504.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1504.txt. Status: INFORMATIONAL.

RFC1505

[1495] A. Costanzo, D. Robinson, and R. Ullmann. RFC 1505: Encoding header field for Internet messages, August 1993. URL ftp://

ftp.internic.net/rfc/rfc1154.txt,   ftp://ftp.internic.net/
rfc/rfc1505.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1154.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1505.txt. Obsoletes

RFC1154 [1144]. Status: EXPERIMENTAL.

RFC1506

[1496] J. Houttuin. RFC 1506: A tutorial on gatewaying between X.400 and Internet mail, September 1993. URL ftp://ftp.internic.net/

rfc/rfc1506.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1506.

txt. Status: INFORMATIONAL.

RFC1507

[1497] C. Kaufman. RFC 1507: DASS — distributed authentication security service, September 1993. URL ftp://ftp.internic.net/

rfc/rfc1507.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1507.

txt. Status: PROPOSED STANDARD.

RFC1508

[1498] J. Linn. RFC 1508: Generic security service application program inter-

face, September 1993. URL ftp://ftp.internic.net/rfc/rfc1508.
txt,   ftp://ftp.internic.net/rfc/rfc2078.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1508.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2078.txt. Obsoleted by RFC2078 [2065]. Status: PRO-POSED STANDARD.

RFC1509

[1499] J. Wray. RFC 1509: Generic security service API: C-bindings, Septem-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1509.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1509.txt. Status: PROPOSED

STANDARD.

RFC1510

[1500] J. Kohl and C. Neuman. RFC 1510: The Kerberos Network Authentication Service (V5), September 1993. URL ftp://ftp.internic.net/

rfc/rfc1510.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1510.

txt. Status: PROPOSED STANDARD.

RFC1511

[1501] J. Linn. RFC 1511: Common authentication technology overview,

September 1993. URL ftp://ftp.internic.net/rfc/rfc1511.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1511.txt. Status: INFOR-

MATIONAL.

RFC1512

[1502] J. Case and A. Rijsinghani. RFC 1512: FDDI management information

base, September 1993. URL ftp://ftp.internic.net/rfc/rfc1285.
txt,   ftp://ftp.internic.net/rfc/rfc1512.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1285.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1512.txt. Updates RFC1285 [1275]. Status: PROPOSED STANDARD.

RFC1513

[1503] S. Waldbusser. RFC 1513: Token ring extensions to the remote network monitoring MIB, September 1993. URL ftp://ftp.internic.net/

rfc/rfc1513.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1513.

txt. Updates RFC1271 [1261]. Status: PROPOSED STANDARD.

RFC1514

[1504] P. Grillo and S. Waldbusser. RFC 1514: Host resources MIB, Septem-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1514.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc1514.txt. Status: PROPOSED

STANDARD.

RFC1515

[1505] D. McMaster, K. McCloghrie, and S. Roberts. RFC 1515: Definitions of managed objects for IEEE 802.3 Medium Attachment Units (MAUs),

September 1993. URL ftp://ftp.internic.net/rfc/rfc1515.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1515.txt. Status: PRO-

POSED STANDARD.

RFC1516

[1506] D. McMaster and K. McCloghrie. RFC 1516: Definitions of managed objects for IEEE 802.3 repeater devices, September 1993. URL ftp://ftp.

internic.net/rfc/rfc1368.txt,   ftp://ftp.internic.net/rfc/
rfc1516.txt,   ftp://ftp.internic.net/rfc/rfc2108.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1368.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1516.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2108.txt. Obsoleted by RFC2108 [2095]. Obsoletes RFC1368 [1358]. Status: DRAFT STANDARD.

RFC1517

[1507] Internet Engineering Steering Group and R. Hinden. RFC 1517: Applicability statement for the implementation of classless inter-domain routing (CIDR), September 1993. URL ftp://ftp.internic.net/

rfc/rfc1517.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1517.

txt. Status: PROPOSED STANDARD.

RFC1518

[1508] Y. Rekhter and T. Li. RFC 1518: An architecture for IP address allocation with CIDR, September 1993. URL ftp://ftp.internic.net/

rfc/rfc1518.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1518.

txt. Status: PROPOSED STANDARD.

RFC1519

[1509] V. Fuller, T. Li, J. Yu, and K. Varadhan. RFC 1519: Classless interdomain routing (CIDR): an address assignment and aggregation strat-

egy, September 1993. URL ftp://ftp.internic.net/rfc/rfc1338.
txt,   ftp://ftp.internic.net/rfc/rfc1519.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1338.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1519.txt. Obsoletes RFC1338 [1328]. Status: PROPOSED STANDARD.

RFC1520

[1510] Y. Rekhter and C. Topolcic. RFC 1520: Exchanging routing information across provider boundaries in the CIDR environment, Septem-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1520.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1520.txt. Status: INFORMA-

TIONAL.

RFC1521

[1511] N. Borenstein and N. Freed. RFC 1521: MIME (Multipurpose Internet Mail Extensions) part one: Mechanisms for specifying and describing the format of Internet message bodies, September

1993. URL ftp://ftp.internic.net/rfc/bcp13.txt,   ftp://ftp.
internic.net/rfc/rfc1341.txt,   ftp://ftp.internic.net/rfc/
rfc1521.txt,   ftp://ftp.internic.net/rfc/rfc1590.txt,   ftp:
//ftp.internic.net/rfc/rfc2045.txt,   ftp://ftp.internic.net/
rfc/rfc2046.txt,   ftp://ftp.internic.net/rfc/rfc2047.txt,
ftp://ftp.internic.net/rfc/rfc2048.txt,   ftp://ftp.internic.
net/rfc/rfc2049.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp13.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1341.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1521.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1590.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2045.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2046.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2047.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2048.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2049.txt. Obsoleted by RFC2045, RFC2046, RFC2047, RFC2048, RFC2049, BCP0013 [2032, 2033, 2034, 2035, 2036, ?]. Obsoletes RFC1341 [1331]. Updated by RFC1590 [1580]. Status: DRAFT STANDARD.

RFC1522

[1512] K. Moore. RFC 1522: MIME (Multipurpose Internet Mail Extensions) part two: Message header extensions for non-ASCII text, September

1993. URL ftp://ftp.internic.net/rfc/bcp13.txt,   ftp://ftp.
internic.net/rfc/rfc1342.txt,   ftp://ftp.internic.net/rfc/
rfc1522.txt,   ftp://ftp.internic.net/rfc/rfc2045.txt,   ftp:
//ftp.internic.net/rfc/rfc2046.txt,   ftp://ftp.internic.net/
rfc/rfc2047.txt,   ftp://ftp.internic.net/rfc/rfc2048.txt,
ftp://ftp.internic.net/rfc/rfc2049.txt,   ftp://ftp.math.utah.
edu/pub/rfc/bcp13.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1342.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1522.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2045.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2046.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2047.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2048.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2049.txt. Obsoleted by

RFC2045, RFC2046, RFC2047, RFC2048, RFC2049, BCP0013 [2032, 2033, 2034, 2035, 2036, ?]. Obsoletes RFC1342 [1332]. Status: DRAFT STANDARD.

RFC1523

[1513] N. Borenstein. RFC 1523: The text/enriched MIME content-

type, September 1993. URL ftp://ftp.internic.net/rfc/rfc1523.
txt,   ftp://ftp.internic.net/rfc/rfc1563.txt,   ftp://ftp.
internic.net/rfc/rfc1896.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1523.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1563.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1896.txt. Obsoleted

by RFC1563, RFC1896 [1553, 1883]. Status: INFORMATIONAL.

RFC1524

[1514] N. Borenstein. RFC 1524: A user agent configuration mechanism for multimedia mail format information, September 1993. URL ftp://ftp.

internic.net/rfc/rfc1524.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1524.txt. Status: INFORMATIONAL.

RFC1525

[1515] E. Decker, K. McCloghrie, P. Langille, and A. Rijsinghani. RFC 1525: Definitions of managed objects for source routing bridges,

September 1993.  URL ftp://ftp.internic.net/rfc/rfc1286.
txt,   ftp://ftp.internic.net/rfc/rfc1525.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1286.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1525.txt. Obsoletes RFC1286 [1276]. Status: PROPOSED STANDARD.

RFC1526

[1516] D. Piscitello. RFC 1526: Assignment of system identifiers for TUBA/CLNP hosts, September 1993. URL ftp://ftp.internic.net/

rfc/rfc1526.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1526.

txt. Status: INFORMATIONAL.

RFC1527

[1517] G. Cook. RFC 1527: What should we plan given the dilemma of the net-

work?, September 1993. URL ftp://ftp.internic.net/rfc/rfc1527.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1527.txt. Status:

INFORMATIONAL.

RFC1528

[1518] M. Rose and C. Malamud. RFC 1528: Principles of operation for the TPC.INT subdomain: Remote printing — technical proce-

dures, October 1993. URL ftp://ftp.internic.net/rfc/rfc1486.
txt,   ftp://ftp.internic.net/rfc/rfc1528.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1486.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1528.txt. Obsoletes RFC1486 [1476]. Status: EXPERI-MENTAL.

RFC1529

[1519] C. Malamud and M. Rose. RFC 1529: Principles of operation for the TPC.INT subdomain: Remote printing — administrative poli-

cies, October 1993. URL ftp://ftp.internic.net/rfc/rfc1486.
txt,   ftp://ftp.internic.net/rfc/rfc1529.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1486.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1529.txt. Obsoletes RFC1486 [1476]. Status: INFORMA-TIONAL.

RFC1530

[1520] C. Malamud and M. Rose. RFC 1530: Principles of operation for the TPC.INT subdomain: General principles and policy, Octo-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1530.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1530.txt. Status: INFORMA-

TIONAL.

RFC1531

[1521] R. Droms. RFC 1531: Dynamic host configuration protocol, Octo-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1531.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1531.txt. Obsoleted by RFC1541

[1531], RFC2131 [2118]. Status: PROPOSED STANDARD.

RFC1532

[1522] W. Wimer. RFC 1532: Clarifications and extensions for the bootstrap

protocol, October 1993. URL ftp://ftp.internic.net/rfc/rfc1532.
txt,   ftp://ftp.internic.net/rfc/rfc1542.txt,   ftp://ftp.
internic.net/rfc/rfc951.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1532.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1542.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc951.txt. Obsoleted

by RFC1542 [1532]. Updates RFC0951 [941]. Status: PROPOSED STANDARD.

RFC1533

[1523] S. Alexander and R. Droms. RFC 1533: DHCP options and BOOTP vendor extensions, October 1993. URL ftp://ftp.internic.net/rfc/

rfc1048.txt,   ftp://ftp.internic.net/rfc/rfc1084.txt,   ftp:
//ftp.internic.net/rfc/rfc1395.txt,   ftp://ftp.internic.net/
rfc/rfc1497.txt,   ftp://ftp.internic.net/rfc/rfc1533.txt,
ftp://ftp.internic.net/rfc/rfc2132.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1048.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1084.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1395.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc1497.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1533.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2132.txt. Obsoleted by RFC2132 [2119]. Obsoletes RFC1497, RFC1395, RFC1084, RFC1048 [1487, 1385, 1074, 1038]. Status: PRO-POSED STANDARD.

RFC1534

[1524] R. Droms. RFC 1534: Interoperation between DHCP and BOOTP, Octo-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1534.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1534.txt. Status: DRAFT STAN-

DARD.

RFC1535

[1525] E. Gavron. RFC 1535: A security problem and proposed correction with widely deployed DNS software, October 1993. URL ftp://ftp.

internic.net/rfc/rfc1535.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1535.txt. Status: INFORMATIONAL.

RFC1536

[1526] A. Kumar, J. Postel, C. Neuman, P. Danzig, and S. Miller. RFC 1536: Common DNS implementation errors and suggested fixes, Octo-

ber 1993. URL ftp://ftp.internic.net/rfc/rfc1536.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1536.txt. Status: INFORMA-

TIONAL.

RFC1537

[1527] P. Beertema. RFC 1537: Common DNS data file configuration er-

rors, October 1993. URL ftp://ftp.internic.net/rfc/rfc1537.
txt,   ftp://ftp.internic.net/rfc/rfc1912.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1537.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1912.txt. Obsoleted by RFC1912 [1899]. Status: INFOR-MATIONAL.

RFC1538

[1528] W. Behl, B. Sterling, and W. Teskey. RFC 1538: Advanced SNA/ IP: A simple SNA transport protocol, October 1993. URL ftp://ftp.

internic.net/rfc/rfc1538.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1538.txt. Status: INFORMATIONAL.

RFC1539

[1529] G. Malkin. RFC 1539: The tao of IETF — A guide for new attendees of the Internet Engineering Task Force, October 1993. URL ftp:

//ftp.internic.net/rfc/fyi17.txt,   ftp://ftp.internic.net/
rfc/rfc1391.txt,   ftp://ftp.internic.net/rfc/rfc1539.txt,

ftp://ftp.internic.net/rfc/rfc1718.txt,   ftp://ftp.math.utah.
edu/pub/rfc/fyi17.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1391.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1539.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1718.txt. Obsoleted by FYI0017,

RFC1718 [?, 1708]. Obsoletes RFC1391 [1381]. Status: INFORMA-TIONAL.

RFC1540

[1530] J. Postel. RFC 1540: Internet Official Protocol Standards, October

1993. URL ftp://ftp.internic.net/rfc/rfc1500.txt,   ftp://
ftp.internic.net/rfc/rfc1540.txt,   ftp://ftp.internic.net/
rfc/rfc1600.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1500.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1540.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1600.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. See also STD0001 [?]. Obsoleted by RFC1600, RFC2200, RFC2300, STD0001, RFC2400 [1590, 2187, 2287, ?, 2387]. Obsoletes RFC1500 [1490]. Status: HISTORIC.

RFC1541

[1531] R. Droms. RFC 1541: Dynamic host configuration protocol, October

1993. URL ftp://ftp.internic.net/rfc/rfc1531.txt,   ftp://
ftp.internic.net/rfc/rfc1541.txt,   ftp://ftp.internic.net/
rfc/rfc2131.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1531.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1541.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2131.txt. Obsoleted by RFC2131

[2118]. Obsoletes RFC1531 [1521]. Status: PROPOSED STANDARD.

RFC1542

[1532] W. Wimer. RFC 1542: Clarifications and extensions for the bootstrap

protocol, October 1993. URL ftp://ftp.internic.net/rfc/rfc1532.
txt,   ftp://ftp.internic.net/rfc/rfc1542.txt,   ftp://ftp.
internic.net/rfc/rfc951.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1532.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1542.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc951.txt. Obsoletes

RFC1532 [1522]. Updates RFC0951 [941]. Status: DRAFT STANDARD.

RFC1543

[1533] J. Postel. RFC 1543: Instructions to RFC authors, October

1993. URL ftp://ftp.internic.net/rfc/rfc1111.txt,   ftp://

ftp.internic.net/rfc/rfc1543.txt,   ftp://ftp.internic.net/
rfc/rfc2223.txt,   ftp://ftp.internic.net/rfc/rfc825.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1111.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1543.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2223.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc825.txt.

Obsoleted by RFC2223 [2210]. Obsoletes RFC1111, RFC0825 [1101, 815]. Status: INFORMATIONAL.

RFC1544

[1534] M. Rose. RFC 1544: The content-MD5 header field, November

1993. URL ftp://ftp.internic.net/rfc/rfc1544.txt,   ftp://
ftp.internic.net/rfc/rfc1864.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1544.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1864.

txt. Obsoleted by RFC1864 [1851]. Status: PROPOSED STANDARD.

RFC1545

[1535] D. Piscitello. RFC 1545: FTP operation over big address records (FOO-

BAR), November 1993. URL ftp://ftp.internic.net/rfc/rfc1545.
txt,   ftp://ftp.internic.net/rfc/rfc1639.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1545.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1639.txt. Obsoleted by RFC1639 [1629]. Status: EXPER-IMENTAL.

RFC1546

[1536] C. Partridge, T. Mendez, and W. Milliken. RFC 1546: Host anycasting

service, November 1993. URL ftp://ftp.internic.net/rfc/rfc1546.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1546.txt. Status:

INFORMATIONAL.

RFC1547

[1537] D. Perkins. RFC 1547: Requirements for an Internet Standard Pointto-Point Protocol, December 1993. URL ftp://ftp.internic.net/

rfc/rfc1547.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1547.

txt. Status: INFORMATIONAL.

RFC1548

[1538] W. Simpson. RFC 1548: The point-to-point protocol (PPP), December

1993. URL ftp://ftp.internic.net/rfc/rfc1331.txt,   ftp://
ftp.internic.net/rfc/rfc1548.txt,   ftp://ftp.internic.net/
rfc/rfc1570.txt,   ftp://ftp.internic.net/rfc/rfc1661.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1331.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1548.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1570.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1661.

txt. Obsoleted by RFC1661 [1651]. Obsoletes RFC1331 [1321]. Updated by RFC1570 [1560]. Status: DRAFT STANDARD.

RFC1549

[1539] W. Simpson. RFC 1549: PPP in HDLC framing, December

1993.  URL ftp://ftp.internic.net/rfc/rfc1549.txt,   ftp:
//ftp.internic.net/rfc/rfc1662.txt, ftp://ftp.internic.net/
rfc/std51.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1549.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1662.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std51.txt. Obsoleted by RFC1662, STD0051 [1652, ?]. Status: DRAFT STANDARD.

RFC1550

[1540] S. Bradner and A. Mankin. RFC 1550: IP: Next generation (IPng) white paper solicitation, December 1993. URL ftp://ftp.internic.net/

rfc/rfc1550.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1550.

txt. Status: INFORMATIONAL.

RFC1551

[1541] M. Allen. RFC 1551: Novell IPX over various WAN media (IPXWAN),

December 1993. URL ftp://ftp.internic.net/rfc/rfc1551.txt,
ftp://ftp.internic.net/rfc/rfc1634.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1551.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1634.txt. Obsoleted by RFC1634 [1624]. Status: INFORMA-TIONAL.

RFC1552

[1542] W. Simpson. RFC 1552: The PPP Internetwork Packet Exchange Control Protocol (IPXCP), December 1993. URL ftp://ftp.internic.

net/rfc/rfc1552.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1552.

txt. Status: PROPOSED STANDARD.

RFC1553

[1543] S. Mathur and M. Lewis. RFC 1553: Compressing IPX headers over WAN media (CIPX), December 1993. URL ftp://ftp.internic.net/

rfc/rfc1553.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1553.

txt. Status: PROPOSED STANDARD.

RFC1554

[1544] M. Ohta and K. Handa. RFC 1554: ISO-2022-JP-2: Multilingual extension of ISO-2022-JP, December 1993. URL ftp://ftp.internic.net/

rfc/rfc1554.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1554.

txt. Status: INFORMATIONAL.

RFC1555

[1545] H. Nussbacher and Y. Bourvine. RFC 1555: Hebrew character encoding for Internet messages, December 1993. URL ftp://ftp.internic.net/

rfc/rfc1555.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1555.

txt. Status: INFORMATIONAL.

RFC1556

[1546] H. Nussbacher. RFC 1556: Handling of bi-directional texts in MIME,

December 1993. URL ftp://ftp.internic.net/rfc/rfc1556.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1556.txt. Status: INFOR-

MATIONAL.

RFC1557

[1547] U. Choi, K. Chon, and H. Park. RFC 1557: Korean character encoding for Internet messages, December 1993. URL ftp://ftp.internic.net/

rfc/rfc1557.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1557.

txt. Status: INFORMATIONAL.

RFC1558

[1548] T. Howes. RFC 1558: A string representation of LDAP search fil-

ters, December 1993. URL ftp://ftp.internic.net/rfc/rfc1558.
txt,   ftp://ftp.internic.net/rfc/rfc1960.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1558.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1960.txt. Obsoleted by RFC1960 [1947]. Status: INFOR-MATIONAL.

RFC1559

[1549] J. Saperia. RFC 1559: DECnet Phase IV MIB extensions, December

1993. URL ftp://ftp.internic.net/rfc/rfc1289.txt,   ftp://
ftp.internic.net/rfc/rfc1559.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1289.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1559.

txt. Obsoletes RFC1289 [1279]. Status: DRAFT STANDARD.

RFC1560

[1550] B. Leiner and Y. Rekhter. RFC 1560: The MultiProtocol Internet,

December 1993. URL ftp://ftp.internic.net/rfc/rfc1560.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1560.txt. Status: INFOR-

MATIONAL.

RFC1561

[1551] D. Piscitello. RFC 1561: Use of ISO CLNP in TUBA environments,

December 1993. URL ftp://ftp.internic.net/rfc/rfc1561.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1561.txt. Status: EXPERI-

MENTAL.

RFC1562

[1552] G. Michaelson and M. Prior. RFC 1562: Naming guidelines for the AARNet X.500 Directory Service, December 1993. URL ftp://ftp.

internic.net/rfc/rfc1562.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1562.txt. Status: INFORMATIONAL.

RFC1563

[1553] N. Borenstein. RFC 1563: The text/enriched MIME content-

type, January 1994. URL ftp://ftp.internic.net/rfc/rfc1523.
txt,   ftp://ftp.internic.net/rfc/rfc1563.txt,   ftp://ftp.
internic.net/rfc/rfc1896.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1523.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1563.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1896.txt. Obsoleted

by RFC1896 [1883]. Obsoletes RFC1523 [1513]. Status: INFORMA-TIONAL.

RFC1564

[1554] P. Barker and R. Hedberg. RFC 1564: DSA metrics (OSI-DS 34 (v3)),

January 1994. URL ftp://ftp.internic.net/rfc/rfc1564.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1564.txt. Status: INFOR-

MATIONAL.

RFC1565

[1555] S. Kille and N. Freed. RFC 1565: Network services monitoring

MIB, January 1994. URL ftp://ftp.internic.net/rfc/rfc1565.
txt,   ftp://ftp.internic.net/rfc/rfc2248.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1565.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2248.txt. Obsoleted by RFC2248 [2235]. Status: PRO-POSED STANDARD.

RFC1566

[1556] S. Kille and N. Freed. RFC 1566: Mail monitoring MIB, January

1994. URL ftp://ftp.internic.net/rfc/rfc1566.txt,   ftp://
ftp.internic.net/rfc/rfc2249.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1566.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2249.

txt. Obsoleted by RFC2249 [2236]. Status: PROPOSED STANDARD.

RFC1567

[1557] G. Mansfield and S. Kille. RFC 1567: X.500 Directory Monitoring MIB,

January 1994. URL ftp://ftp.internic.net/rfc/rfc1567.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1567.txt. Status: PRO-

POSED STANDARD.

RFC1568

[1558] A. Gwinn. RFC 1568: Simple network paging protocol — version

1(b), January 1994. URL ftp://ftp.internic.net/rfc/rfc1568.
txt,   ftp://ftp.internic.net/rfc/rfc1645.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1568.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1645.txt. Obsoleted by RFC1645 [1635]. Status: INFOR-MATIONAL.

RFC1569

[1559] M. Rose. RFC 1569: Principles of operation for the TPC.INT subdomain: Radio paging — technical procedures, January 1994. URL ftp://

ftp.internic.net/rfc/rfc1569.txt,   ftp://ftp.internic.net/
rfc/rfc1703.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1569.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1703.txt. Obsoleted

by RFC1703 [1693]. Status: INFORMATIONAL.

RFC1570

[1560] W. Simpson. RFC 1570: PPP LCP extensions, January 1994. URL ftp:

//ftp.internic.net/rfc/rfc1548.txt,   ftp://ftp.internic.net/
rfc/rfc1570.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1548.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1570.txt. Updates

RFC1548 [1538]. Status: PROPOSED STANDARD.

RFC1571

[1561] D. Borman. RFC 1571: Telnet Environment Option interoperability

issues, January 1994. URL ftp://ftp.internic.net/rfc/rfc1408.
txt,   ftp://ftp.internic.net/rfc/rfc1571.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1408.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1571.txt. Updates RFC1408 [1398]. Status: INFORMA-TIONAL.

RFC1572

[1562] S. Alexander. RFC 1572: Telnet Environment Option, January

1994. URL ftp://ftp.internic.net/rfc/rfc1572.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1572.txt. Status: PROPOSED

STANDARD.

RFC1573

[1563] K. McCloghrie and F. Kastenholz. RFC 1573: Evolution of the interfaces group of MIB-II, January 1994. URL ftp://ftp.

internic.net/rfc/rfc1229.txt,   ftp://ftp.internic.net/rfc/
rfc1573.txt,   ftp://ftp.internic.net/rfc/rfc2233.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1229.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1573.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2233.txt. Obsoleted by RFC1573, RFC2233 [1563, 2220]. Obsoletes RFC1229, RFC1573 [1219, 1563]. Status: PROPOSED STANDARD.

RFC1574

[1564] S. Hares and C. Wittbrodt. RFC 1574: Essential tools for the OSI In-

ternet, February 1994. URL ftp://ftp.internic.net/rfc/rfc1139.
txt,   ftp://ftp.internic.net/rfc/rfc1574.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1139.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1574.txt. Obsoletes RFC1139 [1129]. Status: INFORMA-TIONAL.

RFC1575

[1565] S. Hares and C. Wittbrodt. RFC 1575: An echo function for CLNP (ISO

8473), February 1994. URL ftp://ftp.internic.net/rfc/rfc1139.
txt,   ftp://ftp.internic.net/rfc/rfc1575.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1139.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1575.txt. Obsoletes RFC1139 [1129]. Status: DRAFT STANDARD.

RFC1576

[1566] J. Penner. RFC 1576: TN3270 current practices, January 1994. URL

ftp://ftp.internic.net/rfc/rfc1576.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1576.txt. Status: INFORMATIONAL.

RFC1577

[1567] M. Laubach. RFC 1577: Classical IP and ARP over ATM, January

1994. URL ftp://ftp.internic.net/rfc/rfc1577.txt,   ftp://
ftp.internic.net/rfc/rfc2225.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1577.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2225.

txt. Obsoleted by RFC2225 [2212]. Status: PROPOSED STANDARD.

RFC1578

[1568] J. Sellers. RFC 1578: FYI on questions and answers — answers to commonly asked “Primary and Secondary School Internet User” ques-

tions, February 1994. URL ftp://ftp.internic.net/rfc/fyi22.
txt,   ftp://ftp.internic.net/rfc/rfc1578.txt,   ftp://ftp.
internic.net/rfc/rfc1941.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi22.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1578.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1941.txt. See also FYI0022

[?]. Obsoleted by FYI0022, RFC1941 [?, 1928]. Status: INFORMA-TIONAL.

RFC1579

[1569] S. Bellovin. RFC 1579: Firewall-friendly FTP, February 1994. URL ftp:

//ftp.internic.net/rfc/rfc1579.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1579.txt. Status: INFORMATIONAL.

RFC1580

[1570] EARN Staff. RFC 1580: Guide to network resource tools, March

1994. URL ftp://ftp.internic.net/rfc/fyi23.txt,   ftp://ftp.
internic.net/rfc/rfc1580.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi23.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1580.txt.

See also FYI0023 [?]. Status: INFORMATIONAL.

RFC1581

[1571] G. Meyer. RFC 1581: Protocol analysis for extensions to RIP to support demand circuits, February 1994. URL ftp://ftp.internic.net/

rfc/rfc1581.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1581.

txt. Status: INFORMATIONAL.

RFC1582

[1572] G. Meyer. RFC 1582: Extensions to RIP to support demand circuits,

February 1994. URL ftp://ftp.internic.net/rfc/rfc1582.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1582.txt. Status: PRO-

POSED STANDARD.

RFC1583

[1573] J. Moy. RFC 1583: OSPF version 2, March 1994. URL ftp://ftp.

internic.net/rfc/rfc1247.txt,   ftp://ftp.internic.net/rfc/
rfc1583.txt,   ftp://ftp.internic.net/rfc/rfc2178.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1247.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1583.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2178.txt. Obsoleted by RFC2178 [2165]. Obsoletes RFC1247 [1237]. Status: DRAFT STANDARD.

RFC1584

[1574] J. Moy. RFC 1584: Multicast extensions to OSPF, March 1994. URL

ftp://ftp.internic.net/rfc/rfc1584.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1584.txt. Status: PROPOSED STANDARD.

RFC1585

[1575] J. Moy. RFC 1585: MOSPF: Analysis and experience, March

1994. URL ftp://ftp.internic.net/rfc/rfc1585.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1585.txt. Status: INFORMA-

TIONAL.

RFC1586

[1576] O. deSouza and M. Rodrigues. RFC 1586: Guidelines for running OSPF over frame relay networks, March 1994. URL ftp://ftp.internic.net/

rfc/rfc1586.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1586.

txt. Status: INFORMATIONAL.

RFC1587

[1577] R. Coltun and V. Fuller. RFC 1587: The OSPF NSSA option, March

1994. URL ftp://ftp.internic.net/rfc/rfc1587.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1587.txt. Status: PROPOSED

STANDARD.

RFC1588

[1578] J. Postel and C. Anderson. RFC 1588: White pages meeting report,

February 1994. URL ftp://ftp.internic.net/rfc/rfc1588.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1588.txt. Status: INFOR-

MATIONAL.

RFC1589

[1579] D. Mills. RFC 1589: A kernel model for precision timekeeping,

March 1994. URL ftp://ftp.internic.net/rfc/rfc1589.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1589.txt. Status: INFOR-

MATIONAL.

RFC1590

[1580] J. Postel. RFC 1590: Media type registration procedure, March

1994. URL ftp://ftp.internic.net/rfc/bcp13.txt,   ftp://ftp.
internic.net/rfc/rfc1521.txt,   ftp://ftp.internic.net/rfc/
rfc1590.txt,   ftp://ftp.internic.net/rfc/rfc2045.txt,   ftp:
//ftp.internic.net/rfc/rfc2046.txt,   ftp://ftp.internic.net/
rfc/rfc2047.txt,   ftp://ftp.internic.net/rfc/rfc2048.txt,
ftp://ftp.internic.net/rfc/rfc2049.txt,   ftp://ftp.math.utah.
edu/pub/rfc/bcp13.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1521.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1590.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2045.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2046.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2047.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2048.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2049.txt. Obsoleted by

RFC2045, RFC2046, RFC2047, RFC2048, RFC2049, BCP0013 [2032, 2033, 2034, 2035, 2036, ?]. Updates RFC1521 [1511]. Status: INFOR-MATIONAL.

RFC1591

[1581] J. Postel. RFC 1591: Domain name system structure and delega-

tion, March 1994. URL ftp://ftp.internic.net/rfc/rfc1591.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1591.txt. Status: INFOR-

MATIONAL.

RFC1592

[1582] B. Wijnen, G. Carpenter, K. Curran, A. Sehgal, and G. Waters. RFC 1592: Simple Network Management Protocol distributed protocol interface version 2.0, March 1994. URL ftp://

ftp.internic.net/rfc/rfc1228.txt,   ftp://ftp.internic.net/
rfc/rfc1592.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1228.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1592.txt. Obsoletes

RFC1228 [1218]. Status: EXPERIMENTAL.

RFC1593

[1583] W. McKenzie and J. Cheng. RFC 1593: SNA APPN node MIB,

March 1994. URL ftp://ftp.internic.net/rfc/rfc1593.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1593.txt. Status: INFOR-

MATIONAL.

RFC1594

[1584] A. Marine, J. Reynolds, and G. Malkin. RFC 1594: FYI on questions and answers — answers to commonly asked “New Internet User”

questions, March 1994. URL ftp://ftp.internic.net/rfc/fyi4.
txt,   ftp://ftp.internic.net/rfc/rfc1325.txt,   ftp://ftp.
internic.net/rfc/rfc1594.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi4.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1325.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1594.txt. See also FYI0004

[?]. Obsoletes RFC1325 [1315]. Status: INFORMATIONAL.

RFC1595

[1585] T. Brown and K. Tesink. RFC 1595: Definitions of managed objects for the SONET/SDH interface type, March 1994. URL ftp://ftp.

internic.net/rfc/rfc1595.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1595.txt. Status: PROPOSED STANDARD.

RFC1596

[1586] T. Brown. RFC 1596: Definitions of managed objects for frame relay

service, March 1994. URL ftp://ftp.internic.net/rfc/rfc1596.
txt,   ftp://ftp.internic.net/rfc/rfc1604.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1596.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1604.txt. Obsoleted by RFC1604 [1594]. Status: PRO-POSED STANDARD.

RFC1597

[1587] Y. Rekhter, B. Moskowitz, D. Karrenberg, and G. de Groot. RFC 1597: Address allocation for private internets, March 1994. URL ftp://ftp.

internic.net/rfc/bcp5.txt,   ftp://ftp.internic.net/rfc/rfc1597.
txt,   ftp://ftp.internic.net/rfc/rfc1918.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp5.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1597.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1918.

txt. Obsoleted by BCP0005, RFC1918 [?, 1905]. Status: INFORMA-TIONAL.

RFC1598

[1588] W. Simpson. RFC 1598: PPP in X.25, March 1994. URL ftp://ftp.

internic.net/rfc/rfc1598.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1598.txt. Status: PROPOSED STANDARD.

RFC1599

[1589] M. Kennedy. RFC 1599: Summary of 1500–1599, January 1997. URL

ftp://ftp.internic.net/rfc/rfc1599.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1599.txt. Status: INFORMATIONAL.

RFC1600

[1590] J. Postel. RFC 1600: Internet official protocol standards, March

1994. URL ftp://ftp.internic.net/rfc/rfc1540.txt,   ftp://
ftp.internic.net/rfc/rfc1600.txt,   ftp://ftp.internic.net/
rfc/rfc1610.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1540.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1600.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1610.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. Obsoleted by RFC1610, RFC2200, RFC2300, STD0001, RFC2400 [1600, 2187, 2287, ?, 2387]. Obsoletes RFC1540 [1530]. Status: HISTORIC.

RFC1601

[1591] C. Huitema. RFC 1601: Charter of the Internet Architecture Board

(IAB), March 1994. URL ftp://ftp.internic.net/rfc/rfc1358.
txt,   ftp://ftp.internic.net/rfc/rfc1601.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1358.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1601.txt. Obsoletes RFC1358 [1348]. Status: INFORMA-TIONAL.

RFC1602

[1592] Internet Architecture Board and Internet Engineering Steering Group. RFC 1602: The Internet standards process — revision 2, March

1994. URL ftp://ftp.internic.net/rfc/bcp2.txt,   ftp://ftp.
internic.net/rfc/bcp9.txt, ftp://ftp.internic.net/rfc/rfc1310.
txt,   ftp://ftp.internic.net/rfc/rfc1602.txt,   ftp://ftp.
internic.net/rfc/rfc1871.txt,   ftp://ftp.internic.net/rfc/
rfc2026.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp2.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/bcp9.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1310.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1602.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1871.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2026.txt. Obsoleted by

RFC2026, BCP0009 [2013, ?]. Obsoletes RFC1310 [1300]. Updated by RFC1871, BCP0002 [1858, ?]. Status: INFORMATIONAL.

RFC1603

[1593] E. Huizer and D. Crocker. RFC 1603: IETF working group guidelines and procedures, March 1994. URL ftp://ftp.internic.net/

rfc/bcp2.txt,   ftp://ftp.internic.net/rfc/bcp25.txt, ftp:/
/ftp.internic.net/rfc/rfc1603.txt,   ftp://ftp.internic.net/
rfc/rfc1871.txt,   ftp://ftp.internic.net/rfc/rfc2418.txt,
ftp://ftp.math.utah.edu/pub/rfc/bcp2.txt,   ftp://ftp.math.
utah.edu/pub/rfc/bcp25.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1603.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1871.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2418.txt. Obsoleted by

RFC2418, BCP0025 [2405, ?]. Updated by RFC1871, BCP0002 [1858, ?]. Status: INFORMATIONAL.

RFC1604

[1594] T. Brown. RFC 1604: Definitions of managed objects for frame relay

service, March 1994. URL ftp://ftp.internic.net/rfc/rfc1596.
txt,   ftp://ftp.internic.net/rfc/rfc1604.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1596.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1604.txt. Obsoletes RFC1596 [1586]. Status: PROPOSED STANDARD.

RFC1605

[1595] W. Shakespeare. RFC 1605: SONET to sonnet translation, April

1994. URL ftp://ftp.internic.net/rfc/rfc1605.txt,   ftp:/

/ftp.math.utah.edu/pub/rfc/rfc1605.txt. Status: INFORMA-

TIONAL.

RFC1606

[1596] J. Onions. RFC 1606: A historical perspective on the usage of IP ver-

sion 9, April 1994. URL ftp://ftp.internic.net/rfc/rfc1606.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1606.txt. Status: INFOR-

MATIONAL.

RFC1607

[1597] V. Cerf. RFC 1607: A VIEW FROM THE 21ST CENTURY, April

1994. URL ftp://ftp.internic.net/rfc/rfc1607.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1607.txt. Status: INFORMA-

TIONAL.

RFC1608

[1598] T. Johannsen, G. Mansfield, M. Kosters, and S. Sataluri. RFC 1608: Representing IP information in the X.500 directory, March

1994. URL ftp://ftp.internic.net/rfc/rfc1608.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1608.txt. Status: EXPERIMEN-

TAL.

RFC1609

[1599] G. Mansfield, T. Johannsen, and M. Knopper. RFC 1609: Charting networks in the X.500 directory, March 1994. URL ftp://ftp.internic.

net/rfc/rfc1609.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1609.

txt. Status: EXPERIMENTAL.

RFC1610

[1600] J. Postel. RFC 1610: Internet official protocol standards, July

1994. URL ftp://ftp.internic.net/rfc/rfc1600.txt,   ftp://
ftp.internic.net/rfc/rfc1610.txt,   ftp://ftp.internic.net/
rfc/rfc1720.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1600.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1610.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1720.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.

edu/pub/rfc/std1.txt. See also STD0001 [?]. Obsoleted by RFC1720, RFC2200, RFC2300, STD0001, RFC2400 [1710, 2187, 2287, ?, 2387]. Obsoletes RFC1600 [1590]. Status: HISTORIC.

RFC1611

[1601] R. Austein and J. Saperia. RFC 1611: DNS server MIB extensions,

May 1994. URL ftp://ftp.internic.net/rfc/rfc1611.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1611.txt. Status: PROPOSED

STANDARD.

RFC1612

[1602] R. Austein and J. Saperia. RFC 1612: DNS resolver MIB extensions,

May 1994. URL ftp://ftp.internic.net/rfc/rfc1612.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1612.txt. Status: PROPOSED

STANDARD.

RFC1613

[1603] J. Forster, G. Satz, G. Glick, and R. Day. RFC 1613: cisco Systems X.25 over TCP (XOT), May 1994. URL ftp://ftp.internic.net/

rfc/rfc1613.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1613.

txt. Status: INFORMATIONAL.

RFC1614

[1604] C. Adie. RFC 1614: Network access to multimedia information,

May 1994. URL ftp://ftp.internic.net/rfc/rfc1614.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1614.txt. Status: INFORMA-

TIONAL.

RFC1615

[1605] J. Houttuin and J. Craigie. RFC 1615: Migrating from X.400(84) to

X.400(88), May 1994. URL ftp://ftp.internic.net/rfc/rfc1615.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1615.txt. Status:

INFORMATIONAL.

RFC1616

[1606] RARE WG-MSG Task Force 88, E. Huizer, and J. Romaguera. RFC 1616: X.400(1988) for the academic and research community in Eu-

rope, May 1994. URL ftp://ftp.internic.net/rfc/rfc1616.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1616.txt. Status: INFOR-

MATIONAL.

RFC1617

[1607] P. Barker, S. Kille, and T. Lenggenhager. RFC 1617: Naming and structuring guidelines for X.500 directory pilots, May 1994. URL ftp://

ftp.internic.net/rfc/rfc1384.txt,   ftp://ftp.internic.net/
rfc/rfc1617.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1384.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1617.txt. Obsoletes

RFC1384 [1374]. Status: INFORMATIONAL.

RFC1618

[1608] W. Simpson. RFC 1618: PPP over ISDN, May 1994. URL ftp://ftp.

internic.net/rfc/rfc1618.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1618.txt. Status: PROPOSED STANDARD.

RFC1619

[1609] W. Simpson. RFC 1619: PPP over SONET/SDH, May 1994. URL ftp:/

/ftp.internic.net/rfc/rfc1619.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1619.txt. Status: PROPOSED STANDARD.

RFC1620

[1610] B. Braden, J. Postel, and Y. Rekhter. RFC 1620: Internet architecture extensions for shared media, May 1994. URL ftp://ftp.internic.net/

rfc/rfc1620.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1620.

txt. Status: INFORMATIONAL.

RFC1621

[1611] P. Francis. RFC 1621: Pip near-term architecture, May 1994. URL ftp:/

/ftp.internic.net/rfc/rfc1621.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1621.txt. Status: INFORMATIONAL.

RFC1622

[1612] P. Francis. RFC 1622: Pip header processing, May 1994. URL ftp://

ftp.internic.net/rfc/rfc1622.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1622.txt. Status: INFORMATIONAL.

RFC1623

[1613] F. Kastenholz. RFC 1623: Definitions of managed objects for the Ethernet-like interface types, May 1994. URL ftp://ftp.

internic.net/rfc/rfc1398.txt,   ftp://ftp.internic.net/rfc/
rfc1623.txt,   ftp://ftp.internic.net/rfc/rfc1643.txt, ftp:
//ftp.internic.net/rfc/std50.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1398.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1623.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1643.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std50.txt. See also STD0050 [?]. Ob-

soleted by RFC1643, STD0050 [1633, ?]. Obsoletes RFC1398 [1388]. Status: STANDARD.

RFC1624

[1614] A. Rijsinghani. RFC 1624: Computation of the Internet checksum via incremental update, May 1994. URL ftp://ftp.internic.net/rfc/

rfc1141.txt,   ftp://ftp.internic.net/rfc/rfc1624.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc1141.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1624.txt. Updates RFC1141 [1131]. Status: INFOR-MATIONAL.

RFC1625

[1615] M. St. Pierre, J. Fullton, K. Gamiel, J. Goldman, B. Kahle, J. Kunze, H. Morris, and F. Schiettecatte. RFC 1625: WAIS over Z39.50-

1988, June 1994. URL ftp://ftp.internic.net/rfc/rfc1625.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1625.txt. Status: INFOR-

MATIONAL.

RFC1626

[1616] R. Atkinson. RFC 1626: Default IP MTU for use over ATM

AAL5, May 1994. URL ftp://ftp.internic.net/rfc/rfc1626.
txt,   ftp://ftp.internic.net/rfc/rfc2225.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1626.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2225.txt. Obsoleted by RFC2225 [2212]. Status: PRO-POSED STANDARD.

RFC1627

[1617] E. Lear, E. Fair, D. Crocker, and T. Kessler. RFC 1627: Network 10 considered harmful (some practices shouldn’t be codified), June

1994. URL ftp://ftp.internic.net/rfc/bcp5.txt,   ftp://ftp.
internic.net/rfc/rfc1627.txt,   ftp://ftp.internic.net/rfc/
rfc1918.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp5.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1627.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1918.txt. Obsoleted by BCP0005, RFC1918 [?, 1905]. Status: INFORMATIONAL.

RFC1628

[1618] J. Case. RFC 1628: UPS management information base, May

1994. URL ftp://ftp.internic.net/rfc/rfc1628.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1628.txt. Status: PROPOSED

STANDARD.

RFC1629

[1619] R. Colella, R. Callon, E. Gardner, and Y. Rekhter. RFC 1629: Guidelines for OSI NSAP allocation in the Internet, May 1994. URL ftp://

ftp.internic.net/rfc/rfc1237.txt,   ftp://ftp.internic.net/
rfc/rfc1629.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1237.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1629.txt. Obsoletes

RFC1237 [1227]. Status: DRAFT STANDARD.

RFC1630

[1620] T. Berners-Lee. RFC 1630: Universal Resource Identifiers in WWW: A unifying syntax for the expression of names and addresses of objects on the network as used in the World-Wide Web, June

1994. URL ftp://ftp.internic.net/rfc/rfc1630.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1630.txt. Status: INFORMA-

TIONAL.

RFC1631

[1621] K. Egevang and P. Francis. RFC 1631: The IP network address translator

(NAT), May 1994. URL ftp://ftp.internic.net/rfc/rfc1631.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1631.txt. Status: INFOR-

MATIONAL.

RFC1632

[1622] A. Getchell and S. Sataluri. RFC 1632: A revised catalog of available X.500 implementations, May 1994. URL ftp://ftp.internic.net/

rfc/fyi11.txt,   ftp://ftp.internic.net/rfc/rfc1292.txt,   ftp:
//ftp.internic.net/rfc/rfc1632.txt,   ftp://ftp.internic.net/
rfc/rfc2116.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi11.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1292.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1632.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2116.txt. See also FYI0011 [?]. Obsoleted by RFC2116, FYI0011 [2103, ?]. Obsoletes RFC1292 [1282]. Status: INFORMA-TIONAL.

RFC1633

[1623] R. Braden, D. Clark, and S. Shenker. RFC 1633: Integrated services in the Internet architecture: an overview, June 1994. URL ftp://ftp.

internic.net/rfc/rfc1633.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1633.txt. Status: INFORMATIONAL.

RFC1634

[1624] M. Allen. RFC 1634: Novell IPX over various WAN media

(IPXWAN), May 1994. URL ftp://ftp.internic.net/rfc/rfc1362.
txt,   ftp://ftp.internic.net/rfc/rfc1551.txt,   ftp://ftp.
internic.net/rfc/rfc1634.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1362.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1551.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1634.txt. Obsoletes

RFC1551, RFC1362 [1541, 1352]. Status: INFORMATIONAL.

RFC1635

[1625] P. Deutsch, A. Emtage, and A. Marine. RFC 1635: How to use anony-

mous FTP, May 1994. URL ftp://ftp.internic.net/rfc/fyi24.
txt,   ftp://ftp.internic.net/rfc/rfc1635.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi24.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1635.txt. See also FYI0024 [?]. Status: INFORMA-TIONAL.

RFC1636

[1626] R. Braden, D. Clark, S. Crocker, and C. Huitema. RFC 1636: Report of IAB Workshop on Security in the Internet Architecture — February 8–10,

1994, June 1994. URL ftp://ftp.internic.net/rfc/rfc1636.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1636.txt. Status: INFOR-

MATIONAL.

RFC1637

[1627] B. Manning and R. Colella. RFC 1637: DNS NSAP resource

records, June 1994. URL ftp://ftp.internic.net/rfc/rfc1348.
txt,   ftp://ftp.internic.net/rfc/rfc1637.txt,   ftp://ftp.
internic.net/rfc/rfc1706.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1348.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1637.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1706.txt. Obsoleted

by RFC1706 [1696]. Obsoletes RFC1348 [1338]. Updates RFC1348 [1338]. Status: EXPERIMENTAL.

RFC1638

[1628] F. Baker and R. Bowen. RFC 1638: PPP bridging control proto-

col (BCP), June 1994. URL ftp://ftp.internic.net/rfc/rfc1220.
txt,   ftp://ftp.internic.net/rfc/rfc1638.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1220.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1638.txt. Obsoletes RFC1220 [1210]. Status: PROPOSED STANDARD.

RFC1639

[1629] D. Piscitello. RFC 1639: FTP operation over big address records

(FOOBAR), June 1994. URL ftp://ftp.internic.net/rfc/rfc1545.
txt,   ftp://ftp.internic.net/rfc/rfc1639.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1545.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1639.txt. Obsoletes RFC1545 [1535]. Status: EXPERI-MENTAL.

RFC1640

[1630] S. Crocker. RFC 1640: The process for organization of Internet Standards Working Group (POISED), June 1994. URL ftp://ftp.

internic.net/rfc/rfc1640.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1640.txt. Status: INFORMATIONAL.

RFC1641

[1631] D. Goldsmith and M. Davis. RFC 1641: Using Unicode with MIME,

July 1994. URL ftp://ftp.internic.net/rfc/rfc1641.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1641.txt. Status: EXPERIMEN-

TAL.

RFC1642

[1632] D. Goldsmith and M. Davis. RFC 1642: UTF-7 — A mail-safe transformation format of Unicode, July 1994. URL ftp://ftp.

internic.net/rfc/rfc1642.txt,   ftp://ftp.internic.net/rfc/
rfc2152.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1642.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2152.txt. Obsoleted by

RFC2152 [2139]. Status: EXPERIMENTAL.

RFC1643

[1633] F. Kastenholz. RFC 1643: Definitions of managed objects for the Ethernet-like interface types, July 1994. URL ftp://ftp.

internic.net/rfc/rfc1398.txt,   ftp://ftp.internic.net/rfc/
rfc1623.txt,   ftp://ftp.internic.net/rfc/rfc1643.txt, ftp:
//ftp.internic.net/rfc/std50.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1398.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1623.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1643.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std50.txt. See also STD0050 [?]. Ob-

soletes RFC1623, RFC1398 [1613, 1388]. Status: STANDARD.

RFC1644

[1634] R. Braden. RFC 1644: T/TCP — TCP extensions for transactions functional specification, July 1994. URL ftp://ftp.internic.net/

rfc/rfc1644.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1644.

txt. Status: EXPERIMENTAL.

RFC1645

[1635] A. Gwinn. RFC 1645: Simple network paging protocol — ver-

sion 2, July 1994. URL ftp://ftp.internic.net/rfc/rfc1568.
txt,   ftp://ftp.internic.net/rfc/rfc1645.txt,   ftp://ftp.
internic.net/rfc/rfc1861.txt,   ftp://ftp.internic.net/rfc/
rfc1863.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1568.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1645.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1861.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1863.txt. Obsoleted by RFC1861, RFC1863 [1848, 1850]. Obsoletes RFC1568 [1558]. Status: INFORMATIONAL.

RFC1646

[1636] C. Graves, T. Butts, and M. Angel. RFC 1646: TN3270 extensions for LUname and printer selection, July 1994. URL ftp://ftp.internic.

net/rfc/rfc1646.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1646.

txt. Status: INFORMATIONAL.

RFC1647

[1637] B. Kelly. RFC 1647: TN3270 enhancements, July 1994. URL ftp://

ftp.internic.net/rfc/rfc1647.txt,   ftp://ftp.internic.net/
rfc/rfc2355.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1647.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2355.txt. Obsoleted

by RFC2355 [2342]. Status: PROPOSED STANDARD.

RFC1648

[1638] A. Cargille. RFC 1648: Postmaster convention for X.400 operations,

July 1994. URL ftp://ftp.internic.net/rfc/rfc1648.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1648.txt. Status: PROPOSED

STANDARD.

RFC1649

[1639] R. Hagens and A. Hansen. RFC 1649: Operational requirements for X.400 management domains in the GO-MHS community, July

1994. URL ftp://ftp.internic.net/rfc/rfc1649.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1649.txt. Status: INFORMA-

TIONAL.

RFC1650

[1640] F. Kastenholz. RFC 1650: Definitions of managed objects for the Ethernet-like interface types using SMIv2, August 1994. URL ftp://

ftp.internic.net/rfc/rfc1650.txt,   ftp://ftp.internic.net/
rfc/rfc2357.txt,   ftp://ftp.internic.net/rfc/rfc2358.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1650.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2357.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2358.txt. Obsoleted by RFC2357, RFC2358 [2344, 2345]. Status: PROPOSED STANDARD.

RFC1651

[1641] J. Klensin, N. Freed, M. Rose, E. Stefferud, and D. Crocker. RFC 1651: SMTP service extensions, July 1994. URL ftp://ftp.

internic.net/rfc/rfc1425.txt,   ftp://ftp.internic.net/rfc/
rfc1651.txt,   ftp://ftp.internic.net/rfc/rfc1869.txt, ftp:
//ftp.internic.net/rfc/std10.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1425.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1651.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1869.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std10.txt. Obsoleted by RFC1869,

STD0010 [1856, ?]. Obsoletes RFC1425 [1415]. Status: DRAFT STAN-DARD.

RFC1652

[1642] J. Klensin, N. Freed, M. Rose, E. Stefferud, and D. Crocker. RFC 1652: SMTP service extension for 8bit-MIMEtransport, July

1994. URL ftp://ftp.internic.net/rfc/rfc1426.txt,   ftp://
ftp.internic.net/rfc/rfc1652.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1426.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1652.

txt. Obsoletes RFC1426 [1416]. Status: DRAFT STANDARD.

RFC1653

[1643] J. Klensin, N. Freed, and K. Moore. RFC 1653: SMTP service extension for message size declaration, July 1994. URL ftp://ftp.

internic.net/rfc/rfc1427.txt,   ftp://ftp.internic.net/rfc/
rfc1653.txt,   ftp://ftp.internic.net/rfc/rfc1870.txt, ftp:
//ftp.internic.net/rfc/std11.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1427.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1653.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1870.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std11.txt. See also STD0011 [?]. Ob-

soleted by RFC1870, STD0011 [1857, ?]. Obsoletes RFC1427 [1417]. Status: DRAFT STANDARD.

RFC1654

[1644] Y. Rekhter and T. Lis. RFC 1654: A Border Gateway Protocol 4

(BGP-4), July 1994. URL ftp://ftp.internic.net/rfc/rfc1654.
txt,   ftp://ftp.internic.net/rfc/rfc1771.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1654.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1771.txt. Obsoleted by RFC1771 [1761]. Status: PRO-POSED STANDARD.

RFC1655

[1645] Y. Rekhter and P. Grosss. RFC 1655: Application of the Border Gateway Protocol in the Internet, July 1994. URL ftp://ftp.

internic.net/rfc/rfc1268.txt,   ftp://ftp.internic.net/rfc/
rfc1655.txt,   ftp://ftp.internic.net/rfc/rfc1772.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1268.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1655.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1772.txt. Obsoleted by RFC1772 [1762]. Obsoletes RFC1268 [1258]. Status: PROPOSED STANDARD.

RFC1656

[1646] P. Traina. RFC 1656: BGP-4 protocol document roadmap and implementation experience, July 1994. URL ftp://ftp.internic.net/rfc/

rfc1656.txt,   ftp://ftp.internic.net/rfc/rfc1773.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1656.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1773.txt. Obsoleted by RFC1773 [1763]. Status: PROPOSED STANDARD.

RFC1657

[1647] S. Willis, J. Burruss, and J. Chu. RFC 1657: Definitions of managed objects for the fourth version of the Border Gateway Protocol (BGP-4) using SMIv2, July 1994. URL ftp://ftp.internic.net/

rfc/rfc1657.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1657.

txt. Status: DRAFT STANDARD.

RFC1658

[1648] B. Stewart. RFC 1658: Definitions of managed objects for character stream devices using SMIv2, July 1994. URL ftp://ftp.

internic.net/rfc/rfc1316.txt,   ftp://ftp.internic.net/rfc/
rfc1658.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1316.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1658.txt. Obsoletes RFC1316

[1306]. Status: DRAFT STANDARD.

RFC1659

[1649] B. Stewart. RFC 1659: Definitions of managed objects for RS-232-like hardware devices using SMIv2, July 1994. URL ftp://

ftp.internic.net/rfc/rfc1317.txt,   ftp://ftp.internic.net/
rfc/rfc1659.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1317.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1659.txt. Obsoletes

RFC1317 [1307]. Status: DRAFT STANDARD.

RFC1660

[1650] B. Stewart. RFC 1660: Definitions of managed objects for parallelprinter-like hardware devices using SMIv2, July 1994. URL ftp://

ftp.internic.net/rfc/rfc1318.txt,   ftp://ftp.internic.net/
rfc/rfc1660.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1318.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1660.txt. Obsoletes

RFC1318 [1308]. Status: DRAFT STANDARD.

RFC1661

[1651] W. Simpson. RFC 1661: The point-to-point protocol (PPP), July

1994. URL ftp://ftp.internic.net/rfc/rfc1548.txt,   ftp://

ftp.internic.net/rfc/rfc1661.txt,   ftp://ftp.internic.net/
rfc/rfc2153.txt, ftp://ftp.internic.net/rfc/std51.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1548.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1661.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2153.txt,   ftp://ftp.math.utah.edu/pub/rfc/std51.txt. See

also STD0051 [?]. Obsoletes RFC1548 [1538]. Updated by RFC2153 [2140]. Status: STANDARD.

RFC1662

[1652] W. Simpson. RFC 1662: PPP in HDLC-like framing, July

1994.  URL ftp://ftp.internic.net/rfc/rfc1549.txt,   ftp:
//ftp.internic.net/rfc/rfc1662.txt, ftp://ftp.internic.net/
rfc/std51.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1549.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1662.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std51.txt. See also STD0051 [?]. Obsoletes RFC1549 [1539]. Status: STANDARD.

RFC1663

[1653] D. Rand. RFC 1663: PPP reliable transmission, July 1994. URL ftp://

ftp.internic.net/rfc/rfc1663.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1663.txt. Status: PROPOSED STANDARD.

RFC1664

[1654] C. Allocchio, A. Bonito, B. Cole, S. Giordano, and R. Hagens. RFC 1664: Using the Internet DNS to distribute RFC1327 mail address mapping

tables, August 1994. URL ftp://ftp.internic.net/rfc/rfc1664.
txt,   ftp://ftp.internic.net/rfc/rfc2163.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1664.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2163.txt. Obsoleted by RFC2163 [2150]. Status: EXPER-IMENTAL.

RFC1665

[1655] Z. Kielczewski, D. Kostick, and K. Shihs. RFC 1665: Definitions of managed objects for SNA NAUs using SMIv2, July 1994. URL ftp://

ftp.internic.net/rfc/rfc1665.txt,   ftp://ftp.internic.net/
rfc/rfc1666.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1665.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1666.txt. Obsoleted

by RFC1666 [1656]. Status: PROPOSED STANDARD.

RFC1666

[1656] Z. Kielczewski, D. Kostick, and K. Shih. RFC 1666: Definitions of managed objects for SNA NAUs using SMIv2, August 1994. URL ftp://

ftp.internic.net/rfc/rfc1665.txt,   ftp://ftp.internic.net/

rfc/rfc1666.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1665.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1666.txt. Obsoletes

RFC1665 [1655]. Status: PROPOSED STANDARD.

RFC1667

[1657] S. Symington, D. Wood, and M. Pullen. RFC 1667: Modeling and simulation requirements for IPng, August 1994. URL ftp://ftp.internic.

net/rfc/rfc1667.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1667.

txt. Status: INFORMATIONAL.

RFC1668

[1658] D. Estrin, T. Li, and Y. Rekhter. RFC 1668: Unified routing requirements for IPng, August 1994. URL ftp://ftp.internic.net/

rfc/rfc1668.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1668.

txt. Status: INFORMATIONAL.

RFC1669

[1659] J. Curran. RFC 1669: Market viability as a IPng criteria, Au-

gust 1994. URL ftp://ftp.internic.net/rfc/rfc1669.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1669.txt. Status: INFORMA-

TIONAL.

RFC1670

[1660] D. Heagerty. RFC 1670: Input to IPng engineering considerations,

August 1994. URL ftp://ftp.internic.net/rfc/rfc1670.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1670.txt. Status: INFOR-

MATIONAL.

RFC1671

[1661] B. Carpenter. RFC 1671: IPng white paper on transition and other considerations, August 1994. URL ftp://ftp.internic.net/

rfc/rfc1671.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1671.

txt. Status: INFORMATIONAL.

RFC1672

[1662] N. Brownlee. RFC 1672: Accounting requirements for IPng, Au-

gust 1994. URL ftp://ftp.internic.net/rfc/rfc1672.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1672.txt. Status: INFORMA-

TIONAL.

RFC1673

[1663] R. Skelton. RFC 1673: Electric power research institute comments on

IPng, August 1994. URL ftp://ftp.internic.net/rfc/rfc1673.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc1673.txt. Status: INFOR-

MATIONAL.

RFC1674

[1664] M. Taylor. RFC 1674: A cellular industry view of IPng, Au-

gust 1994. URL ftp://ftp.internic.net/rfc/rfc1674.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1674.txt. Status: INFORMA-

TIONAL.

RFC1675

[1665] S. Bellovin. RFC 1675: Security concerns for IPng, August

1994. URL ftp://ftp.internic.net/rfc/rfc1675.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1675.txt. Status: INFORMA-

TIONAL.

RFC1676

[1666] A. Ghiselli, D. Salomoni, and C. Vistoli. RFC 1676: INFN requirements for an IPng, August 1994. URL ftp://ftp.internic.net/

rfc/rfc1676.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1676.

txt. Status: INFORMATIONAL.

RFC1677

[1667] B. Adamson. RFC 1677: Tactical radio frequency communication requirements for IPng, August 1994. URL ftp://ftp.internic.net/

rfc/rfc1677.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1677.

txt. Status: INFORMATIONAL.

RFC1678

[1668] E. Britton and J. Tavs. RFC 1678: IPng requirements of large corporate

networks, August 1994. URL ftp://ftp.internic.net/rfc/rfc1678.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1678.txt. Status:

INFORMATIONAL.

RFC1679

[1669] D. Green, P. Irey, D. Marlow, and K. O’Donoghue. RFC 1679: HPN working group input to the IPng requirements solicitation, Au-

gust 1994. URL ftp://ftp.internic.net/rfc/rfc1679.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1679.txt. Status: INFORMA-

TIONAL.

RFC1680

[1670] C. Brazdziunas. RFC 1680: IPng support for ATM services, Au-

gust 1994. URL ftp://ftp.internic.net/rfc/rfc1680.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1680.txt. Status: INFORMA-

TIONAL.

RFC1681

[1671] S. Bellovin. RFC 1681: On many addresses per host, August

1994. URL ftp://ftp.internic.net/rfc/rfc1681.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1681.txt. Status: INFORMA-

TIONAL.

RFC1682

[1672] J. Bound. RFC 1682: IPng BSD host implementation analysis, Au-

gust 1994. URL ftp://ftp.internic.net/rfc/rfc1682.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1682.txt. Status: INFORMA-

TIONAL.

RFC1683

[1673] R. Clark, M. Ammar, and K. Calvert. RFC 1683: Multiprotocol interoperability in IPng, August 1994. URL ftp://ftp.internic.net/

rfc/rfc1683.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1683.

txt. Status: INFORMATIONAL.

RFC1684

[1674] P. Jurg. RFC 1684: Introduction to white pages services based on X.500,

August 1994. URL ftp://ftp.internic.net/rfc/rfc1684.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1684.txt. Status: INFOR-

MATIONAL.

RFC1685

[1675] H. Alvestrand. RFC 1685: Writing X.400 O/R names, August

1994. URL ftp://ftp.internic.net/rfc/rfc1685.txt,   ftp:/
/ftp.internic.net/rfc/rtr12.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1685.txt,   ftp://ftp.math.utah.edu/pub/rfc/rtr12.

txt. See also RTR0012 [?]. Status: INFORMATIONAL.

RFC1686

[1676] M. Vecchi. RFC 1686: IPng requirements: A cable television industry

viewpoint, August 1994. URL ftp://ftp.internic.net/rfc/rfc1686.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1686.txt. Status:

INFORMATIONAL.

RFC1687

[1677] E. Fleischman. RFC 1687: A large corporate user’s view of IPng,

August 1994. URL ftp://ftp.internic.net/rfc/rfc1687.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1687.txt. Status: INFOR-

MATIONAL.

RFC1688

[1678] W. Simpson. RFC 1688: IPng mobility considerations, August

1994. URL ftp://ftp.internic.net/rfc/rfc1688.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1688.txt. Status: INFORMA-

TIONAL.

RFC1689

[1679] J. Foster. RFC 1689: A status report on networked information retrieval: Tools and groups, August 1994. URL ftp://ftp.internic.net/rfc/

fyi25.txt,   ftp://ftp.internic.net/rfc/rfc1689.txt, ftp://
ftp.internic.net/rfc/rtr13.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi25.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1689.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rtr13.txt. See also FYI0025,

RTR0013 [?, ?]. Status: INFORMATIONAL.

RFC1690

[1680] G. Huston. RFC 1690: Introducing the Internet Engineering and Planning Group (IEPG), August 1994. URL ftp://ftp.internic.net/

rfc/rfc1690.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1690.

txt. Status: INFORMATIONAL.

RFC1691

[1681] W. Turner. RFC 1691: The document architecture for the Cornell Digital

Library, August 1994. URL ftp://ftp.internic.net/rfc/rfc1691.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1691.txt. Status:

INFORMATIONAL.

RFC1692

[1682] P. Cameron, D. Crocker, D. Cohen, and J. Postel. RFC 1692: Transport Multiplexing Protocol (TMux), August 1994. URL ftp://ftp.

internic.net/rfc/rfc1692.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1692.txt. Status: PROPOSED STANDARD.

RFC1693

[1683] T. Connolly, P. Amer, and P. Conrad. RFC 1693: An extension to TCP: Partial order service, November 1994. URL ftp://ftp.internic.net/

rfc/rfc1693.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1693.

txt. Status: EXPERIMENTAL.

RFC1694

[1684] T. Brown and K. Tesinks. RFC 1694: Definitions of managed objects for SMDS interfaces using SMIv2, August 1994. URL ftp://

ftp.internic.net/rfc/rfc1304.txt,   ftp://ftp.internic.net/

rfc/rfc1694.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1304.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1694.txt. Obsoletes

RFC1304 [1294]. Status: DRAFT STANDARD.

RFC1695

[1685] M. Ahmed and K. Tesink. RFC 1695: Definitions of managed objects for ATM management version 8.0 using SMIv2, August

1994. URL ftp://ftp.internic.net/rfc/rfc1695.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1695.txt. Status: PROPOSED

STANDARD.

RFC1696

[1686] J. Barnes, L. Brown, R. Royston, and S. Waldbusser. RFC 1696: Modem management information base (MIB) using SMIv2, August

1994. URL ftp://ftp.internic.net/rfc/rfc1696.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1696.txt. Status: PROPOSED

STANDARD.

RFC1697

[1687] D. Brower, B. Purvy (RDBMSMIB Working Group Chair), A. Daniel, M. Sinykin, and J. Smith. RFC 1697: Relational database management system (RDBMS) management information base (MIB) using SMIv2,

August 1994. URL ftp://ftp.internic.net/rfc/rfc1697.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1697.txt. Status: PRO-

POSED STANDARD.

RFC1698

[1688] P. Furniss. RFC 1698: Octet sequences for upper-layer OSI to support basic communications applications, October 1994. URL ftp://ftp.

internic.net/rfc/rfc1698.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1698.txt. Status: INFORMATIONAL.

RFC1699

[1689] J. Elliott. RFC 1699: Summary of 1600–1699, January 1997. URL ftp:/

/ftp.internic.net/rfc/rfc1699.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1699.txt. Status: INFORMATIONAL.

RFC1700

[1690] J. Reynolds and J. Postel. RFC 1700: ASSIGNED NUMBERS, Octo-

ber 1994. URL ftp://ftp.internic.net/rfc/rfc1340.txt,   ftp:
//ftp.internic.net/rfc/rfc1700.txt, ftp://ftp.internic.net/
rfc/std2.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1340.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1700.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std2.txt. See also STD0002 [?]. Obsoletes RFC1340 [1330]. Status: STANDARD.

RFC1701

[1691] S. Hanks, T. Li, D. Farinacci, and P. Traina. RFC 1701: Generic routing encapsulation (GRE), October 1994. URL ftp://ftp.internic.net/

rfc/rfc1701.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1701.

txt. Status: INFORMATIONAL.

RFC1702

[1692] S. Hanks, T. Li, D. Farinacci, and P. Traina. RFC 1702: Generic routing encapsulation over IPv4 networks, October 1994. URL ftp://ftp.

internic.net/rfc/rfc1702.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1702.txt. Status: INFORMATIONAL.

RFC1703

[1693] M. Rose. RFC 1703: Principles of operation for the TPC.INT subdomain: Radio paging — technical procedures, October 1994. URL ftp://

ftp.internic.net/rfc/rfc1569.txt,   ftp://ftp.internic.net/
rfc/rfc1703.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1569.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1703.txt. Obsoletes

RFC1569 [1559]. Status: INFORMATIONAL.

RFC1704

[1694] N. Haller and R. Atkinson. RFC 1704: On Internet authentication,

October 1994. URL ftp://ftp.internic.net/rfc/rfc1704.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1704.txt. Status: INFOR-

MATIONAL.

RFC1705

[1695] R. Carlson and D. Ficarella. RFC 1705: Six virtual inches to the left: The problem with IPng, October 1994. URL ftp://ftp.internic.net/

rfc/rfc1705.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1705.

txt. Status: INFORMATIONAL.

RFC1706

[1696] B. Manning and R. Colella. RFC 1706: DNS NSAP resource

records, October 1994. URL ftp://ftp.internic.net/rfc/rfc1637.
txt,   ftp://ftp.internic.net/rfc/rfc1706.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1637.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1706.txt. Obsoletes RFC1637 [1627]. Status: INFORMA-TIONAL.

RFC1707

[1697] M. McGovern and R. Ullmann. RFC 1707: CATNIP: Common architecture for the Internet, October 1994. URL ftp://ftp.internic.net/

rfc/rfc1707.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1707.

txt. Status: INFORMATIONAL.

RFC1708

[1698] D. Gowin. RFC 1708: NTP PICS PROFORMA — for the network time protocol version 3, October 1994. URL ftp://ftp.internic.net/

rfc/rfc1708.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1708.

txt. Status: INFORMATIONAL.

RFC1709

[1699] J. Gargano and D. Wasley. RFC 1709: K-12 Internetworking guide-

lines, November 1994. URL ftp://ftp.internic.net/rfc/fyi26.
txt,   ftp://ftp.internic.net/rfc/rfc1709.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi26.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1709.txt. See also FYI0026 [?]. Status: INFORMA-TIONAL.

RFC1710

[1700] R. Hinden. RFC 1710: Simple Internet Protocol Plus white paper,

October 1994. URL ftp://ftp.internic.net/rfc/rfc1710.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1710.txt. Status: INFOR-

MATIONAL.

RFC1711

[1701] J. Houttuin. RFC 1711: Classifications in E-mail routing, Octo-

ber 1994. URL ftp://ftp.internic.net/rfc/rfc1711.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1711.txt. Status: INFORMA-

TIONAL.

RFC1712

[1702] C. Farrell, M. Schulze, S. Pleitner, and D. Baldoni. RFC 1712: DNS encoding of geographical location, November 1994. URL ftp://ftp.

internic.net/rfc/rfc1712.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1712.txt. Status: EXPERIMENTAL.

RFC1713

[1703] A. Romao. RFC 1713: Tools for DNS debugging, November 1994. URL

ftp://ftp.internic.net/rfc/fyi27.txt,   ftp://ftp.internic.
net/rfc/rfc1713.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi27.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1713.txt. See also

FYI0027 [?]. Status: INFORMATIONAL.

RFC1714

[1704] S. Williamson and M. Kosters. RFC 1714: Referral whois protocol (RWhois), November 1994. URL ftp://ftp.internic.net/rfc/

rfc1714.txt,   ftp://ftp.internic.net/rfc/rfc2167.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1714.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2167.txt. Obsoleted by RFC2167 [2154]. Status: IN-FORMATIONAL.

RFC1715

[1705] C. Huitema. RFC 1715: The H ratio for address assignment efficiency,

November 1994. URL ftp://ftp.internic.net/rfc/rfc1715.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1715.txt. Status: INFOR-

MATIONAL.

RFC1716

[1706] P. Almquist and F. Kastenholz. RFC 1716: Towards requirements for IP routers, November 1994. URL ftp://ftp.internic.net/rfc/

rfc1716.txt,   ftp://ftp.internic.net/rfc/rfc1812.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1716.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1812.txt. Obsoleted by RFC1812 [1802]. Status: IN-FORMATIONAL.

RFC1717

[1707] K. Sklower, B. Lloyd, G. McGregor, and D. Carr. RFC 1717: The PPP multilink protocol (MP), November 1994. URL ftp://

ftp.internic.net/rfc/rfc1717.txt,   ftp://ftp.internic.net/
rfc/rfc1990.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1717.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1990.txt. Obsoleted

by RFC1990 [1977]. Status: PROPOSED STANDARD.

RFC1718

[1708] The IETF Secretariat and G. Malkin. RFC 1718: The tao of IETF — A guide for new attendees of the Internet Engineering Task Force, Novem-

ber 1994. URL ftp://ftp.internic.net/rfc/fyi17.txt,   ftp://
ftp.internic.net/rfc/rfc1539.txt,   ftp://ftp.internic.net/
rfc/rfc1718.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi17.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1539.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1718.txt. See also FYI0017 [?]. Obsoletes

RFC1539 [1529]. Status: INFORMATIONAL.

RFC1719

[1709] P. Gross. RFC 1719: A direction for IPng, December 1994. URL ftp://

ftp.internic.net/rfc/rfc1719.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1719.txt. Status: INFORMATIONAL.

RFC1720

[1710] J. Postel. RFC 1720: Internet official protocol standards, November

1994. URL ftp://ftp.internic.net/rfc/rfc1610.txt,   ftp://
ftp.internic.net/rfc/rfc1720.txt,   ftp://ftp.internic.net/
rfc/rfc1780.txt,   ftp://ftp.internic.net/rfc/rfc1880.txt,
ftp://ftp.internic.net/rfc/rfc1920.txt,   ftp://ftp.internic.
net/rfc/rfc2200.txt,   ftp://ftp.internic.net/rfc/rfc2300.txt,
 ftp://ftp.internic.net/rfc/rfc2400.txt, ftp://ftp.internic.
net/rfc/std1.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1610.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1720.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1780.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1880.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1920.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.edu/pub/

rfc/std1.txt. See also STD0001 [?]. Obsoleted by RFC1780, STD0001, RFC1880, RFC1920, RFC2200, RFC2300, RFC2400 [1770, ?, 1867, 1907, 2187, 2287, 2387]. Obsoletes RFC1610 [1600]. Status: HISTORIC.

RFC1721

[1711] G. Malkin. RFC 1721: RIP version 2 protocol analysis, November

1994. URL ftp://ftp.internic.net/rfc/rfc1387.txt,   ftp://
ftp.internic.net/rfc/rfc1721.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1387.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1721.

txt. Obsoletes RFC1387 [1377]. Status: INFORMATIONAL.

RFC1722

[1712] G. Malkin. RFC 1722: RIP version 2 protocol applicability statement,

November 1994. URL ftp://ftp.internic.net/rfc/rfc1722.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1722.txt. Status: DRAFT

STANDARD.

RFC1723

[1713] G. Malkin. RFC 1723: RIP version 2 — carrying additional informa-

tion, November 1994. URL ftp://ftp.internic.net/rfc/rfc1058.
txt,   ftp://ftp.internic.net/rfc/rfc1388.txt,   ftp://ftp.
internic.net/rfc/rfc1723.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1058.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1388.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1723.txt. Obsoletes

RFC1388 [1378]. Obsoleted by RFC2453 [2440]. Updates RFC1058 [1048]. Updated by RFC2453 [2440], STD0056 [?]. Status: DRAFT STANDARD.

RFC1724

[1714] G. Malkin and F. Baker. RFC 1724: RIP version 2 MIB exten-

sion, November 1994. URL ftp://ftp.internic.net/rfc/rfc1389.
txt,   ftp://ftp.internic.net/rfc/rfc1724.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1389.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1724.txt. Obsoletes RFC1389 [1379]. Status: DRAFT STANDARD.

RFC1725

[1715] J. Myers and M. Rose. RFC 1725: Post office protocol — version

3, November 1994. URL ftp://ftp.internic.net/rfc/rfc1460.
txt,   ftp://ftp.internic.net/rfc/rfc1725.txt,   ftp://ftp.
internic.net/rfc/rfc1939.txt, ftp://ftp.internic.net/rfc/std53.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1460.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1725.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1939.txt,   ftp://ftp.math.utah.edu/pub/rfc/

std53.txt. Obsoleted by RFC1939, STD0053 [1926, ?]. Obsoletes RFC1460 [1450]. Status: DRAFT STANDARD.

RFC1726

[1716] C. Partridge and F. Kastenholz. RFC 1726: Technical criteria for choosing IP the next generation (IPng), December 1994. URL ftp://ftp.

internic.net/rfc/rfc1726.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1726.txt. Status: INFORMATIONAL.

RFC1727

[1717] C. Weider and P. Deutsch. RFC 1727: A vision of an integrated Internet information service, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1727.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1727.

txt. Status: INFORMATIONAL.

RFC1728

[1718] C. Weider. RFC 1728: Resource transponders, December 1994. URL

ftp://ftp.internic.net/rfc/rfc1728.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1728.txt. Status: INFORMATIONAL.

RFC1729

[1719] C. Lynch. RFC 1729: Using the Z39.50 Information Retrieval Protocol in the Internet environment, December 1994. URL ftp://ftp.internic.

net/rfc/rfc1729.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1729.

txt. Status: INFORMATIONAL.

RFC1730

[1720] M. Crispin. RFC 1730: INTERNET MESSAGE ACCESS PROTOCOL

— VERSION 4, December 1994. URL ftp://ftp.internic.net/rfc/
rfc1730.txt,   ftp://ftp.internic.net/rfc/rfc2060.txt,   ftp:
//ftp.internic.net/rfc/rfc2061.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1730.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2060.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2061.txt. Obsoleted

by RFC2060, RFC2061 [2047, 2048]. Status: PROPOSED STANDARD.

RFC1731

[1721] J. Myers. RFC 1731: IMAP4 authentication mechanisms, Decem-

ber 1994. URL ftp://ftp.internic.net/rfc/rfc1731.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1731.txt. Status: PROPOSED

STANDARD.

RFC1732

[1722] M. Crispin. RFC 1732: IMAP4 COMPATIBILITY WITH IMAP2 AND IMAP2BIS, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1732.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1732.

txt. Status: INFORMATIONAL.

RFC1733

[1723] M. Crispin. RFC 1733: DISTRIBUTED ELECTRONIC MAIL MOD-ELS IN IMAP4, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1733.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1733.

txt. Status: INFORMATIONAL.

RFC1734

[1724] J. Myers. RFC 1734: POP3 AUTHentication command, Decem-

ber 1994. URL ftp://ftp.internic.net/rfc/rfc1734.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1734.txt. Status: PROPOSED

STANDARD.

RFC1735

[1725] J. Heinanen and R. Govindan. RFC 1735: NBMA address resolution protocol (NARP), December 1994. URL ftp://ftp.internic.net/

rfc/rfc1735.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1735.

txt. Status: EXPERIMENTAL.

RFC1736

[1726] J. Kunze. RFC 1736: Functional recommendations for Internet Resource Locators, February 1995. URL ftp://ftp.internic.net/

rfc/rfc1736.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1736.

txt. Status: INFORMATIONAL.

RFC1737

[1727] K. Sollins and L. Masinter. RFC 1737: Functional requirements for uniform resource names, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1737.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1737.

txt. Status: INFORMATIONAL.

RFC1738

[1728] T. Berners-Lee, L. Masinter, and M. McCahill. RFC 1738: Uniform resource locators (URL), December 1994. URL ftp://ftp.

internic.net/rfc/rfc1738.txt,   ftp://ftp.internic.net/rfc/
rfc1808.txt,   ftp://ftp.internic.net/rfc/rfc2368.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1738.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1808.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2368.txt. Updated by RFC1808, RFC2368 [1798, 2355]. Status: PROPOSED STANDARD.

RFC1739

[1729] G. Kessler and S. Shepard. RFC 1739: A primer on Internet and TCP/

IP tools, December 1994. URL ftp://ftp.internic.net/rfc/fyi30.
txt,   ftp://ftp.internic.net/rfc/rfc1739.txt,   ftp://ftp.
internic.net/rfc/rfc2151.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi30.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1739.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2151.txt. Obsoleted by

RFC2151, FYI0030 [2138, ?]. Status: INFORMATIONAL.

RFC1740

[1730] P. Faltstrom, D. Crocker, and E. Fair. RFC 1740: MIME encapsulation of Macintosh files — MacMIME, December 1994. URL ftp://ftp.

internic.net/rfc/rfc1740.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1740.txt. Status: PROPOSED STANDARD.

RFC1741

[1731] P. Faltstrom, D. Crocker, and E. Fair. RFC 1741: MIME content type for BinHex encoded files, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1741.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1741.

txt. Status: INFORMATIONAL.

RFC1742

[1732] S. Waldbusser and K. Frisa. RFC 1742: AppleTalk Management Information Base II, January 1995. URL ftp://ftp.internic.net/rfc/

rfc1243.txt,   ftp://ftp.internic.net/rfc/rfc1742.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1243.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1742.txt. Obsoletes RFC1243 [1233]. Status: PRO-POSED STANDARD.

RFC1743

[1733] K. McCloghrie and E. Decker. RFC 1743: IEEE 802.5 MIB using

SMIv2, December 1994. URL ftp://ftp.internic.net/rfc/rfc1231.
txt,   ftp://ftp.internic.net/rfc/rfc1743.txt,   ftp://ftp.
internic.net/rfc/rfc1748.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1231.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1743.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1748.txt. Obsoleted

by RFC1748 [1738]. Obsoletes RFC1231 [1221]. Status: DRAFT STAN-DARD.

RFC1744

[1734] G. Huston. RFC 1744: Observations on the management of the Internet address space, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1744.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1744.

txt. Status: INFORMATIONAL.

RFC1745

[1735] K. Varadhan, S. Hares, and Y. Rekhter. RFC 1745: BGP4/IDRP for IP-OSPF interaction, December 1994. URL ftp://ftp.internic.net/

rfc/rfc1745.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1745.

txt. Status: PROPOSED STANDARD.

RFC1746

[1736] B. Manning and D. Perkins. RFC 1746: Ways to define user expectations,

December 1994. URL ftp://ftp.internic.net/rfc/rfc1746.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1746.txt. Status: INFOR-

MATIONAL.

RFC1747

[1737] J. Hilgeman, Chair, S. Nix, A. Bartky, and W. Clark. RFC 1747: Definitions of managed objects for SNA data link control (SDLC) using SMIv2,

January 1995. URL ftp://ftp.internic.net/rfc/rfc1747.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1747.txt. Status: PRO-

POSED STANDARD.

RFC1748

[1738] K. McCloghrie and E. Decker. RFC 1748: IEEE 802.5 MIB using

SMIv2, December 1994. URL ftp://ftp.internic.net/rfc/rfc1231.
txt,   ftp://ftp.internic.net/rfc/rfc1743.txt,   ftp://ftp.
internic.net/rfc/rfc1748.txt,   ftp://ftp.internic.net/rfc/
rfc1749.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1231.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1743.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1748.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1749.txt. Obsoletes RFC1743, RFC1231 [1733, 1221]. Updated by RFC1749 [1739]. Status: DRAFT STANDARD.

RFC1749

[1739] K. McCloghrie, F. Baker, and E. Decker. RFC 1749: IEEE 802.5 station source routing MIB using SMIv2, December 1994. URL ftp://

ftp.internic.net/rfc/rfc1748.txt,   ftp://ftp.internic.net/
rfc/rfc1749.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1748.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1749.txt. Updates

RFC1748 [1738]. Status: PROPOSED STANDARD.

RFC1750

[1740] D. Eastlake, 3rd, S. Crocker, and J. Schiller. RFC 1750: Randomness recommendations for security, December 1994. URL ftp://ftp.internic.

net/rfc/rfc1750.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1750.

txt. Status: INFORMATIONAL.

RFC1751

[1741] D. McDonald. RFC 1751: A convention for human-readable 128-bit keys,

December 1994. URL ftp://ftp.internic.net/rfc/rfc1751.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1751.txt. Status: INFOR-

MATIONAL.

RFC1752

[1742] S. Bradner and A. Mankin. RFC 1752: The recommendation for the IP next generation protocol, January 1995. URL ftp://ftp.internic.

net/rfc/rfc1752.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1752.

txt. Status: PROPOSED STANDARD.

RFC1753

[1743] N. Chiappa. RFC 1753: IPng technical requirements of the nimrod routing and addressing architecture, December 1994. URL ftp://ftp.

internic.net/rfc/rfc1753.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1753.txt. Status: INFORMATIONAL.

RFC1754

[1744] M. Laubach. RFC 1754: IP over ATM working group’s recommendations for the ATM forum’s multiprotocol BOF version 1, Jan-

uary 1995. URL ftp://ftp.internic.net/rfc/rfc1754.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1754.txt. Status: INFORMA-

TIONAL.

RFC1755

[1745] M. Perez, F. Liaw, D. Grossman, A. Mankin, E. Hoffman, and A. Malis. RFC 1755: ATM signaling support for IP over ATM, Febru-

ary 1995. URL ftp://ftp.internic.net/rfc/rfc1755.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1755.txt. Status: PROPOSED

STANDARD.

RFC1756

[1746] T. Rinne. RFC 1756: REMOTE WRITE PROTOCOL — VERSION

1.0, January 1995. URL ftp://ftp.internic.net/rfc/rfc1756.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1756.txt. Status: EXPER-

IMENTAL.

RFC1757

[1747] S. Waldbusser. RFC 1757: Remote network monitoring management information base, February 1995. URL ftp://ftp.internic.net/rfc/

rfc1271.txt,   ftp://ftp.internic.net/rfc/rfc1757.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1271.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1757.txt. Obsoletes RFC1271 [1261]. Status: DRAFT STANDARD.

RFC1758

[1748] The North American Directory Forum. RFC 1758: NADF standing documents: A brief overview, February 1995. URL ftp://

ftp.internic.net/rfc/rfc1417.txt,   ftp://ftp.internic.net/
rfc/rfc1758.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1417.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1758.txt. Obsoletes

RFC1417 [1407]. Status: INFORMATIONAL.

RFC1759

[1749] R. Smith, F. Wright, T. Hastings, S. Zilles, and J. Gyllenskog. RFC 1759: Printer MIB, March 1995. URL ftp://ftp.internic.net/

rfc/rfc1759.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1759.

txt. Status: PROPOSED STANDARD.

RFC1760

[1750] N. Haller. RFC 1760: The S/KEY one-time password system, Febru-

ary 1995. URL ftp://ftp.internic.net/rfc/rfc1760.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1760.txt. Status: INFORMA-

TIONAL.

RFC1761

[1751] B. Callaghan and R. Gilligan. RFC 1761: Snoop version 2 packet capture file format, February 1995. URL ftp://ftp.internic.net/

rfc/rfc1761.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1761.

txt. Status: INFORMATIONAL.

RFC1762

[1752] S. Senum. RFC 1762: The PPP DECnet phase IV control protocol

(DNCP), March 1995. URL ftp://ftp.internic.net/rfc/rfc1376.
txt,   ftp://ftp.internic.net/rfc/rfc1762.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1376.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1762.txt. Obsoletes RFC1376 [1366]. Status: DRAFT STANDARD.

RFC1763

[1753] S. Senum. RFC 1763: The PPP Banyan Vines Control Protocol (BVCP),

March 1995. URL ftp://ftp.internic.net/rfc/rfc1763.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1763.txt. Status: PRO-

POSED STANDARD.

RFC1764

[1754] S. Senum. RFC 1764: The PPP XNS IDP control protocol (XN-

SCP), March 1995. URL ftp://ftp.internic.net/rfc/rfc1764.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1764.txt. Status: PRO-

POSED STANDARD.

RFC1765

[1755] J. Moy. RFC 1765: OSPF database overflow, March 1995. URL ftp://

ftp.internic.net/rfc/rfc1765.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1765.txt. Status: EXPERIMENTAL.

RFC1766

[1756] H. Alvestrand. RFC 1766: Tags for the identification of languages, March

1995. URL ftp://ftp.internic.net/rfc/rfc1766.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1766.txt. Status: PROPOSED

STANDARD.

RFC1767

[1757] D. Crocker. RFC 1767: MIME encapsulation of EDI objects, March

1995. URL ftp://ftp.internic.net/rfc/rfc1767.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1767.txt. Status: PROPOSED

STANDARD.

RFC1768

[1758] D. Marlow. RFC 1768: Host group extensions for CLNP multicast-

ing, March 1995. URL ftp://ftp.internic.net/rfc/rfc1768.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1768.txt. Status: EXPER-

IMENTAL.

RFC1769

[1759] D. Mills. RFC 1769: Simple network time protocol (SNTP), March

1995. URL ftp://ftp.internic.net/rfc/rfc1361.txt,   ftp://
ftp.internic.net/rfc/rfc1769.txt,   ftp://ftp.internic.net/
rfc/rfc2030.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1361.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1769.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2030.txt. Obsoleted by RFC2030

[2017]. Obsoletes RFC1361 [1351]. Status: INFORMATIONAL.

RFC1770

[1760] C. Graff. RFC 1770: IPv4 option for sender directed multi-destination

delivery, March 1995. URL ftp://ftp.internic.net/rfc/rfc1770.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1770.txt. Status:

INFORMATIONAL.

RFC1771

[1761] Y. Rekhter and T. Li. RFC 1771: A Border Gateway Protocol 4

(BGP-4), March 1995. URL ftp://ftp.internic.net/rfc/rfc1654.
txt,   ftp://ftp.internic.net/rfc/rfc1771.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1654.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1771.txt. Obsoletes RFC1654 [1644]. Status: DRAFT STANDARD.

RFC1772

[1762] Y. Rekhter and P. Gross. RFC 1772: Application of the Border Gateway Protocol in the Internet, March 1995. URL ftp://

ftp.internic.net/rfc/rfc1655.txt,   ftp://ftp.internic.net/
rfc/rfc1772.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1655.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1772.txt. Obsoletes

RFC1655 [1645]. Status: DRAFT STANDARD.

RFC1773

[1763] P. Traina. RFC 1773: Experience with the BGP-4 protocol, March

1995. URL ftp://ftp.internic.net/rfc/rfc1656.txt,   ftp://
ftp.internic.net/rfc/rfc1773.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1656.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1773.

txt. Obsoletes RFC1656 [1646]. Status: INFORMATIONAL.

RFC1774

[1764] P. Traina. RFC 1774: BGP-4 protocol analysis, March 1995. URL ftp:/

/ftp.internic.net/rfc/rfc1774.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1774.txt. Status: INFORMATIONAL.

RFC1775

[1765] D. Crocker. RFC 1775: To be “on” the Internet, March 1995. URL ftp:/

/ftp.internic.net/rfc/rfc1775.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1775.txt. Status: INFORMATIONAL.

RFC1776

[1766] S. Crocker. RFC 1776: The address is the message, March

1995. URL ftp://ftp.internic.net/rfc/rfc1776.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1776.txt. Status: INFORMA-

TIONAL.

RFC1777

[1767] W. Yeong, T. Howes, and S. Kille. RFC 1777: Lightweight directory access protocol, March 1995. URL ftp://ftp.internic.net/rfc/

rfc1487.txt,   ftp://ftp.internic.net/rfc/rfc1777.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1487.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1777.txt. Obsoletes RFC1487 [1477]. Status: DRAFT STANDARD.

RFC1778

[1768] T. Howes, S. Kille, W. Yeong, and C. Robbins. RFC 1778: The string representation of standard attribute syntaxes, March 1995. URL ftp://

ftp.internic.net/rfc/rfc1488.txt,   ftp://ftp.internic.net/
rfc/rfc1778.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1488.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1778.txt. Obsoletes

RFC1488 [1478]. Status: DRAFT STANDARD.

RFC1779

[1769] S. Kille. RFC 1779: A string representation of distinguished

names, March 1995. URL ftp://ftp.internic.net/rfc/rfc1485.

txt,   ftp://ftp.internic.net/rfc/rfc1779.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1485.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1779.txt. Obsoletes RFC1485 [1475]. Status: DRAFT STANDARD.

RFC1780

[1770] J. Postel. RFC 1780: INTERNET OFFICIAL PROTOCOL STAN-

DARDS, March 1995. URL ftp://ftp.internic.net/rfc/rfc1720.
txt,   ftp://ftp.internic.net/rfc/rfc1780.txt,   ftp://ftp.
internic.net/rfc/rfc1800.txt,   ftp://ftp.internic.net/rfc/
rfc1880.txt,   ftp://ftp.internic.net/rfc/rfc1920.txt,   ftp:
//ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.internic.net/
rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/rfc2400.txt, ftp:
//ftp.internic.net/rfc/std1.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1720.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1780.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1800.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1880.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1920.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2200.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.

utah.edu/pub/rfc/std1.txt. See also STD0001 [?]. Obsoleted by RFC1800, STD0001, RFC1880, RFC1920, RFC2200, RFC2300, RFC2400 [1790, ?, 1867, 1907, 2187, 2287, 2387]. Obsoletes RFC1720 [1710]. Status: HISTORIC.

RFC1781

[1771] S. Kille. RFC 1781: Using the OSI directory to achieve user friendly

naming, March 1995. URL ftp://ftp.internic.net/rfc/rfc1484.
txt,   ftp://ftp.internic.net/rfc/rfc1781.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1484.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1781.txt. Obsoletes RFC1484 [1474]. Updates RFC1484 [1474]. Status: PROPOSED STANDARD.

RFC1782

[1772] G. Malkin and A. Harkin. RFC 1782: TFTP option extension, March

1995. URL ftp://ftp.internic.net/rfc/rfc1350.txt,   ftp://
ftp.internic.net/rfc/rfc1782.txt,   ftp://ftp.internic.net/
rfc/rfc2347.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1350.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1782.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2347.txt. Obsoleted by RFC2347

[2334]. Updates RFC1350 [1340]. Status: PROPOSED STANDARD.

RFC1783

[1773] G. Malkin and A. Harkin. RFC 1783: TFTP blocksize option, March

1995. URL ftp://ftp.internic.net/rfc/rfc1350.txt,   ftp://
ftp.internic.net/rfc/rfc1783.txt,   ftp://ftp.internic.net/
rfc/rfc2348.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1350.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1783.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2348.txt. Obsoleted by RFC2348

[2335]. Updates RFC1350 [1340]. Updated by RFC1350 [1340]. Status: PROPOSED STANDARD.

RFC1784

[1774] G. Malkin and A. Harkin. RFC 1784: TFTP timeout interval and transfer size options, March 1995. URL ftp://ftp.

internic.net/rfc/rfc1350.txt,   ftp://ftp.internic.net/rfc/
rfc1784.txt,   ftp://ftp.internic.net/rfc/rfc2349.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1350.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1784.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2349.txt. Obsoleted by RFC2349 [2336]. Updates RFC1350 [1340]. Status: PROPOSED STANDARD.

RFC1785

[1775] G. Malkin and A. Harkin. RFC 1785: TFTP option negotiation

analysis, March 1995. URL ftp://ftp.internic.net/rfc/rfc1350.
txt,   ftp://ftp.internic.net/rfc/rfc1785.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1350.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1785.txt. Updates RFC1350 [1340]. Status: INFORMA-TIONAL.

RFC1786

[1776] T. Bates, E. Gerich, L. Joncheray, J-M. Jouanigot, D. Karrenberg, M. Terpstra, and J. Yu. RFC 1786: Representation of IP routing policies in a routing registry (ripe-81++), March 1995. URL ftp://ftp.

internic.net/rfc/rfc1786.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1786.txt. Status: INFORMATIONAL.

RFC1787

[1777] Y. Rekhter. RFC 1787: Routing in a multi-provider Internet, April

1995. URL ftp://ftp.internic.net/rfc/rfc1787.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1787.txt. Status: INFORMA-

TIONAL.

RFC1788

[1778] W. Simpson. RFC 1788: ICMP domain name messages, April

1995. URL ftp://ftp.internic.net/rfc/rfc1788.txt,   ftp://

ftp.math.utah.edu/pub/rfc/rfc1788.txt. Status: EXPERIMEN-

TAL.

RFC1789

[1779] C. Yang. RFC 1789: INETPhone: Telephone services and servers on In-

ternet, April 1995. URL ftp://ftp.internic.net/rfc/rfc1789.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1789.txt. Status: INFOR-

MATIONAL.

RFC1790

[1780] V. Cerf. RFC 1790: An agreement between the Internet Society and Sun Microsystems, Inc. in the matter of ONC RPC and XDR proto-

cols, April 1995. URL ftp://ftp.internic.net/rfc/rfc1790.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1790.txt. Status: INFOR-

MATIONAL.

RFC1791

[1781] T. Sung. RFC 1791: TCP and UDP over IPX networks with fixed path

MTU, April 1995. URL ftp://ftp.internic.net/rfc/rfc1791.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1791.txt. Status: EXPER-

IMENTAL.

RFC1792

[1782] T. Sung. RFC 1792: TCP/IPX connection MIB specification, April

1995. URL ftp://ftp.internic.net/rfc/rfc1792.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1792.txt. Status: EXPERIMEN-

TAL.

RFC1793

[1783] J. Moy. RFC 1793: Extending OSPF to support demand cir-

cuits, April 1995. URL ftp://ftp.internic.net/rfc/rfc1793.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1793.txt. Status: PRO-

POSED STANDARD.

RFC1794

[1784] T. Brisco. RFC 1794: DNS support for load balancing, April

1995. URL ftp://ftp.internic.net/rfc/rfc1794.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1794.txt. Status: INFORMA-

TIONAL.

RFC1795

[1785] L. Wells, Chair and A. Bartky, Editor. RFC 1795: Data link switching: Switch-to-switch protocol AIW DLSw RIG: DLSw closed pages, DLSw standard version 1.0, April 1995. URL ftp://

ftp.internic.net/rfc/rfc1434.txt,   ftp://ftp.internic.net/
rfc/rfc1795.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1434.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1795.txt. Obsoletes

RFC1434 [1424]. Status: INFORMATIONAL.

RFC1796

[1786] C. Huitema, J. Postel, and S. Crocker. RFC 1796: Not all RFCs are stan-

dards, April 1995. URL ftp://ftp.internic.net/rfc/rfc1796.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1796.txt. Status: INFOR-

MATIONAL.

RFC1797

[1787] Internet Assigned Numbers Authority (IANA). RFC 1797: Class A subnet experiment, April 1995. URL ftp://ftp.internic.net/

rfc/rfc1797.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1797.

txt. Status: EXPERIMENTAL.

RFC1798

[1788] A. Young. RFC 1798: Connection-less lightweight X.500 Directory Access Protocol, June 1995. URL ftp://ftp.internic.net/

rfc/rfc1798.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1798.

txt. Status: PROPOSED STANDARD.

RFC1799

[1789] M. Kennedy. RFC 1799: Request for comments summary RFC numbers 1700–1799, January 1997. URL ftp://ftp.internic.net/

rfc/rfc1799.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1799.

txt. Status: INFORMATIONAL.

RFC1800

[1790] J. Postel. RFC 1800: INTERNET OFFICIAL PROTOCOL STAN-

DARDS, July 1995. URL ftp://ftp.internic.net/rfc/rfc1780.
txt,   ftp://ftp.internic.net/rfc/rfc1800.txt,   ftp://ftp.
internic.net/rfc/rfc1880.txt,   ftp://ftp.internic.net/rfc/
rfc1920.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp:
//ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.net/
rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1780.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1800.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1880.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1920.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2400.txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt.

See also STD0001 [?]. Obsoleted by RFC1880, STD0001, RFC1920, RFC2200, RFC2300, RFC2400 [1867, ?, 1907, 2187, 2287, 2387]. Obsoletes RFC1780 [1770]. Status: HISTORIC.

RFC1801

[1791] S. Kille. RFC 1801: MHS use of the X.500 directory to support MHS

routing, June 1995. URL ftp://ftp.internic.net/rfc/rfc1801.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1801.txt. Status: EXPER-

IMENTAL.

RFC1802

[1792] H. Alvestrand, K. Jordan, S. Langlois, and J. Romaguera. RFC 1802: Introducing Project Long Bud: Internet pilot project for the deployment of X.500 directory information in support of X.400 rout-

ing, June 1995. URL ftp://ftp.internic.net/rfc/rfc1802.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1802.txt. Status: INFOR-

MATIONAL.

RFC1803

[1793] R. Wright, A. Getchell, T. Howes, S. Sataluri, P. Yee, and W. Yeong. RFC 1803: Recommendations for an X.500 production directory ser-

vice, June 1995. URL ftp://ftp.internic.net/rfc/rfc1803.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1803.txt. Status: INFOR-

MATIONAL.

RFC1804

[1794] G. Mansfield, P. Rajeev, S. Raghavan, and T. Howes. RFC 1804: Schema publishing in X.500 directory, June 1995. URL ftp://ftp.internic.

net/rfc/rfc1804.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1804.

txt. Status: EXPERIMENTAL.

RFC1805

[1795] A. Rubin. RFC 1805: Location-Independent Data/Software Integrity

Protocol, June 1995. URL ftp://ftp.internic.net/rfc/rfc1805.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1805.txt. Status:

INFORMATIONAL.

RFC1806

[1796] R. Troost and S. Dorner. RFC 1806: Communicating presentation information in Internet messages: The content-disposition

header, June 1995. URL ftp://ftp.internic.net/rfc/rfc1806.
txt,   ftp://ftp.internic.net/rfc/rfc2183.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1806.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2183.txt. Updated by RFC2183 [2170]. Status: EXPERI-MENTAL.

RFC1807

[1797] R. Lasher and D. Cohen. RFC 1807: A format for bibliographic

records, June 1995. URL ftp://ftp.internic.net/rfc/rfc1357.
txt,   ftp://ftp.internic.net/rfc/rfc1807.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1357.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1807.txt. Obsoletes RFC1357 [1347]. Status: INFORMA-TIONAL.

RFC1808

[1798] R. Fielding. RFC 1808: Relative uniform resource locators, June

1995. URL ftp://ftp.internic.net/rfc/rfc1738.txt,   ftp://
ftp.internic.net/rfc/rfc1808.txt,   ftp://ftp.internic.net/
rfc/rfc2368.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1738.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1808.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2368.txt. Updates RFC1738 [1728].

Updated by RFC2368 [2355]. Status: PROPOSED STANDARD.

RFC1809

[1799] C. Partridge. RFC 1809: Using the flow label field in IPv6,

June 1995. URL ftp://ftp.internic.net/rfc/rfc1809.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1809.txt. Status: INFORMA-

TIONAL.

RFC1810

[1800] J. Touch. RFC 1810: Report on MD5 performance, June 1995. URL ftp:

//ftp.internic.net/rfc/rfc1810.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1810.txt. Status: INFORMATIONAL.

RFC1811

[1801] Federal Networking Council. RFC 1811: U.S. Government Internet do-

main names, June 1995. URL ftp://ftp.internic.net/rfc/rfc1811.
txt,   ftp://ftp.internic.net/rfc/rfc1816.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1811.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1816.txt. Obsoleted by RFC1816 [1806]. Status: INFOR-MATIONAL.

RFC1812

[1802] F. Baker. RFC 1812: Requirements for IP version 4 routers, June

1995. URL ftp://ftp.internic.net/rfc/rfc1009.txt,   ftp://
ftp.internic.net/rfc/rfc1716.txt,   ftp://ftp.internic.net/
rfc/rfc1812.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1009.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1716.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1812.txt. Obsoletes RFC1716,

RFC1009 [1706, 999]. Status: PROPOSED STANDARD.

RFC1813

[1803] B. Callaghan, B. Pawlowski, and P. Staubach. RFC 1813: NFS version 3 protocol specification, June 1995. URL ftp://ftp.

internic.net/rfc/rfc1094.txt,   ftp://ftp.internic.net/rfc/
rfc1813.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1094.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1813.txt. See also RFC1094

[1084]. Status: INFORMATIONAL.

RFC1814

[1804] E. Gerich. RFC 1814: Unique addresses are good, June 1995. URL ftp:/

/ftp.internic.net/rfc/rfc1814.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1814.txt. Status: INFORMATIONAL.

RFC1815

[1805] M. Ohta. RFC 1815: Character sets ISO-10646 and ISO-10646-J-

1, July 1995. URL ftp://ftp.internic.net/rfc/rfc1815.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1815.txt. Status: INFOR-

MATIONAL.

RFC1816

[1806] Federal Networking Council. RFC 1816: U.S. Government Internet

domain names, August 1995. URL ftp://ftp.internic.net/rfc/
rfc1811.txt,   ftp://ftp.internic.net/rfc/rfc1816.txt,   ftp:
//ftp.internic.net/rfc/rfc2146.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1811.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1816.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2146.txt. Obsoleted

by RFC2146 [2133]. Obsoletes RFC1811 [1801]. Status: INFORMA-TIONAL.

RFC1817

[1807] Y. Rekhter. RFC 1817: CIDR and classful routing, August

1995. URL ftp://ftp.internic.net/rfc/rfc1817.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1817.txt. Status: INFORMA-

TIONAL.

RFC1818

[1808] J. Postel, T. Li, and Y. Rekhter. RFC 1818: Best current

practices, August 1995. URL ftp://ftp.internic.net/rfc/bcp1.
txt,   ftp://ftp.internic.net/rfc/rfc1818.txt,   ftp://ftp.

math.utah.edu/pub/rfc/bcp1.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1818.txt. See also BCP0001 [?]. Status: BEST CURRENT PRACTICE.

RFC1819

[1809] L. Delgrossi and L. Bergers. RFC 1819: Internet Stream Protocol Version 2 (ST2) protocol specification — version ST2+, August

1995. URL ftp://ftp.internic.net/rfc/ien119.txt,   ftp://
ftp.internic.net/rfc/rfc1190.txt,   ftp://ftp.internic.net/
rfc/rfc1819.txt,   ftp://ftp.math.utah.edu/pub/rfc/ien119.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1190.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc1819.txt. Obsoletes RFC1190, IEN119 [1180, ?]. Status: EXPERIMENTAL.

RFC1820

[1810] E. Huizer. RFC 1820: Multimedia E-mail (MIME) user agent check-

list, August 1995. URL ftp://ftp.internic.net/rfc/rfc1820.
txt,   ftp://ftp.internic.net/rfc/rfc1844.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1820.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1844.txt. Obsoleted by RFC1844 [1832]. Status: INFOR-MATIONAL.

RFC1821

[1811] M. Borden, E. Crawley, B. Davie, and S. Batsell. RFC 1821: Integration of real-time services in an IP-ATM network architecture, Au-

gust 1995. URL ftp://ftp.internic.net/rfc/rfc1821.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1821.txt. Status: INFORMA-

TIONAL.

RFC1822

[1812] J. Lowe. RFC 1822: A grant of rights to use a specific IBM patent with

Photuris, August 1995. URL ftp://ftp.internic.net/rfc/rfc1822.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1822.txt. Status:

INFORMATIONAL.

RFC1823

[1813] T. Howes and M. Smith. RFC 1823: The LDAP application program

interface, August 1995. URL ftp://ftp.internic.net/rfc/rfc1823.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1823.txt. Status:

INFORMATIONAL.

RFC1824

[1814] H. Danisch. RFC 1824: The Exponential Security System TESS: An identity-based cryptographic protocol for authenticated key-exchange (E.I.S.S.-Report 1995/4), August 1995. URL ftp://ftp.internic.

net/rfc/rfc1824.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1824.

txt. Status: INFORMATIONAL.

RFC1825

[1815] R. Atkinson. RFC 1825: Security architecture for the Internet Pro-

tocol, August 1995. URL ftp://ftp.internic.net/rfc/rfc1825.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1825.txt. Obsoleted

by RFC2401 [2388]. Status: PROPOSED STANDARD.

RFC1826

[1816] R. Atkinson. RFC 1826: IP authentication header, August 1995. URL

ftp://ftp.internic.net/rfc/rfc1826.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1826.txt. Obsoleted by RFC2402 [2389]. Status: PROPOSED STANDARD.

RFC1827

[1817] R. Atkinson. RFC 1827: IP encapsulating security payload (ESP), Au-

gust 1995. URL ftp://ftp.internic.net/rfc/rfc1827.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1827.txt. Obsoleted by RFC2406

[2393]. Status: PROPOSED STANDARD.

RFC1828

[1818] P. Metzger and W. Simpson. RFC 1828: IP authentication using keyed

MD5, August 1995. URL ftp://ftp.internic.net/rfc/rfc1828.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1828.txt. Status: PRO-

POSED STANDARD.

RFC1829

[1819] P. Karn, P. Metzger, and W. Simpson. RFC 1829: The ESP DES-CBC

transform, August 1995. URL ftp://ftp.internic.net/rfc/rfc1829.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1829.txt. Status:

PROPOSED STANDARD.

RFC1830

[1820] G. Vaudreuil. RFC 1830: SMTP service extensions for transmission of large and binary MIME messages, August 1995. URL ftp://ftp.

internic.net/rfc/rfc1830.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1830.txt. Status: EXPERIMENTAL.

RFC1831

[1821] R. Srinivasan. RFC 1831: RPC: Remote procedure call protocol specification version 2, August 1995. URL ftp://ftp.internic.net/

rfc/rfc1831.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1831.

txt. Status: PROPOSED STANDARD.

RFC1832

[1822] R. Srinivasan. RFC 1832: XDR: External data representation stan-

dard, August 1995. URL ftp://ftp.internic.net/rfc/rfc1832.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1832.txt. Status: DRAFT

STANDARD.

RFC1833

[1823] R. Srinivasan. RFC 1833: Binding protocols for ONC RPC version 2, Au-

gust 1995. URL ftp://ftp.internic.net/rfc/rfc1833.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1833.txt. Status: PROPOSED

STANDARD.

RFC1834

[1824] J. Gargano and K. Weiss. RFC 1834: Whois and network information lookup service, whois++, August 1995. URL ftp://ftp.internic.

net/rfc/rfc1834.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1834.

txt. Status: INFORMATIONAL.

RFC1835

[1825] P. Deutsch, R. Schoultz, P. Faltstrom, and C. Weider. RFC 1835: Architecture of the WHOIS++ service, August 1995. URL ftp://ftp.

internic.net/rfc/rfc1835.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1835.txt. Status: PROPOSED STANDARD.

RFC1836

[1826] S. Kille. RFC 1836: Representing the O/R address hierarchy in the X.500 directory information tree, August 1995. URL ftp://

ftp.internic.net/rfc/rfc1836.txt,   ftp://ftp.internic.net/
rfc/rfc2294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1836.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2294.txt. Obsoleted

by RFC2294 [2281]. Status: EXPERIMENTAL.

RFC1837

[1827] S. Kille. RFC 1837: Representing tables and subtrees in the X.500 di-

rectory, August 1995. URL ftp://ftp.internic.net/rfc/rfc1837.
txt,   ftp://ftp.internic.net/rfc/rfc2293.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1837.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2293.txt. Obsoleted by RFC2293 [2280]. Status: EXPER-IMENTAL.

RFC1838

[1828] S. Kille. RFC 1838: Use of the X.500 directory to support mapping between X.400 and RFC 822 addresses, August 1995. URL ftp://

ftp.internic.net/rfc/rfc1838.txt,   ftp://ftp.internic.net/
rfc/rfc2164.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1838.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2164.txt. Obsoleted

by RFC2164 [2151]. Status: EXPERIMENTAL.

RFC1841

[1829] J. Chapman, D. Coli, A. Harvey, B. Jensen, and K. Rowett. RFC 1841: PPP network control protocol for LAN extension, Septem-

ber 1995. URL ftp://ftp.internic.net/rfc/rfc1841.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1841.txt. Status: INFORMA-

TIONAL.

RFC1842

[1830] Y. Wei, Y. Zhang, J. Li, J. Ding, and Y. Jiang. RFC 1842: ASCII printable characters-based Chinese character encoding for Internet messages,

August 1995. URL ftp://ftp.internic.net/rfc/rfc1842.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1842.txt. Status: INFOR-

MATIONAL.

RFC1843

[1831] F. Lee. RFC 1843: HZ — A data format for exchanging files of arbitrarily mixed Chinese and ASCII characters, August 1995. URL ftp://ftp.

internic.net/rfc/rfc1843.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1843.txt. Status: INFORMATIONAL.

RFC1844

[1832] E. Huizer. RFC 1844: Multimedia E-mail (MIME) user agent check-

list, August 1995. URL ftp://ftp.internic.net/rfc/rfc1820.
txt,   ftp://ftp.internic.net/rfc/rfc1844.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1820.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1844.txt. Obsoletes RFC1820 [1810]. Status: INFORMA-TIONAL.

RFC1845

[1833] D. Crocker, N. Freed, and A. Cargille. RFC 1845: SMTP service extension for checkpoint/restart, September 1995. URL ftp://ftp.

internic.net/rfc/rfc1845.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1845.txt. Status: EXPERIMENTAL.

RFC1846

[1834] A. Durand and F. Dupont. RFC 1846: SMTP 521 reply code, Septem-

ber 1995. URL ftp://ftp.internic.net/rfc/rfc1846.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1846.txt. Status: EXPERIMEN-

TAL.

RFC1847

[1835] J. Galvin, S. Murphy, S. Crocker, and N. Freed. RFC 1847: Security multiparts for MIME: Multipart/signed and multipart/encrypted, Octo-

ber 1995. URL ftp://ftp.internic.net/rfc/rfc1847.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1847.txt. Status: PROPOSED

STANDARD.

RFC1848

[1836] S. Crocker, N. Freed, J. Galvin, and S. Murphy. RFC 1848: MIME object security services, October 1995. URL ftp://ftp.internic.net/

rfc/rfc1848.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1848.

txt. Status: PROPOSED STANDARD.

RFC1850

[1837] F. Baker and R. Coltun. RFC 1850: OSPF version 2 management information base, November 1995. URL ftp://ftp.

internic.net/rfc/rfc1253.txt,   ftp://ftp.internic.net/rfc/
rfc1850.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1253.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1850.txt. Obsoletes RFC1253

[1243]. Status: DRAFT STANDARD.

RFC1851

[1838] P. Karn, P. Metzger, and W. Simpson. RFC 1851: The ESP triple DES transform, September 1995. URL ftp://ftp.internic.net/

rfc/rfc1851.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1851.

txt. Status: EXPERIMENTAL.

RFC1852

[1839] P. Metzger and W. Simpson. RFC 1852: IP authentication using keyed

SHA, September 1995. URL ftp://ftp.internic.net/rfc/rfc1852.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1852.txt. Status:

EXPERIMENTAL.

RFC1853

[1840] W. Simpson. RFC 1853: IP in IP tunneling, October 1995. URL ftp://

ftp.internic.net/rfc/rfc1853.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1853.txt. Status: INFORMATIONAL.

RFC1854

[1841] N. Freed. RFC 1854: SMTP service extension for command pipelin-

ing, October 1995. URL ftp://ftp.internic.net/rfc/rfc1854.
txt,   ftp://ftp.internic.net/rfc/rfc2197.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1854.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2197.txt. Obsoleted by RFC2197 [2184]. Status: PRO-POSED STANDARD.

RFC1855

[1842] S. Hambridge. RFC 1855: Netiquette guidelines, October 1995. URL

ftp://ftp.internic.net/rfc/fyi28.txt,   ftp://ftp.internic.
net/rfc/rfc1855.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi28.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1855.txt. See also

FYI0028 [?]. Status: INFORMATIONAL.

RFC1856

[1843] H. Clark. RFC 1856: The opstat client-server model for statistics re-

trieval, October 1995. URL ftp://ftp.internic.net/rfc/rfc1856.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1856.txt. Status:

INFORMATIONAL.

RFC1857

[1844] M. Lambert. RFC 1857: A model for common operational statis-

tics, October 1995. URL ftp://ftp.internic.net/rfc/rfc1404.
txt,   ftp://ftp.internic.net/rfc/rfc1857.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1404.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1857.txt. Obsoletes RFC1404 [1394]. Status: INFORMA-TIONAL.

RFC1858

[1845] G. Ziemba, D. Reed, and P. Traina. RFC 1858: Security considerations for IP fragment filtering, October 1995. URL ftp://ftp.internic.net/

rfc/rfc1858.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1858.

txt. Status: INFORMATIONAL.

RFC1859

[1846] Y. Pouffary. RFC 1859: ISO transport class 2 non-use of explicit flow control over TCP RFC1006 extension, October 1995. URL ftp://ftp.

internic.net/rfc/rfc1859.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1859.txt. Status: INFORMATIONAL.

RFC1860

[1847] T. Pummill and B. Manning. RFC 1860: Variable length subnet table

for IPv4, October 1995. URL ftp://ftp.internic.net/rfc/rfc1860.

txt,   ftp://ftp.internic.net/rfc/rfc1878.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1860.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1878.txt. Obsoleted by RFC1878 [1865]. Status: INFOR-MATIONAL.

RFC1861

[1848] A. Gwinn. RFC 1861: Simple network paging protocol — version 3 — two-way enhanced, October 1995. URL ftp://ftp.

internic.net/rfc/rfc1645.txt,   ftp://ftp.internic.net/rfc/
rfc1861.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1645.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1861.txt. Obsoletes RFC1645

[1635]. Status: INFORMATIONAL.

RFC1862

[1849] M. McCahill, J. Romkey, M. Schwartz, K. Sollins, T. Verschuren, and C. Weider. RFC 1862: Report of the IAB Workshop on Internet Information Infrastructure, October 12–14, 1994, Novem-

ber 1995. URL ftp://ftp.internic.net/rfc/rfc1862.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1862.txt. Status: INFORMA-

TIONAL.

RFC1863

[1850] D. Haskin. RFC 1863: A BGP/IDRP route server alternative to a full mesh routing, October 1995. URL ftp://ftp.

internic.net/rfc/rfc1645.txt,   ftp://ftp.internic.net/rfc/
rfc1863.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1645.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1863.txt. Obsoletes RFC1645

[1635]. Status: EXPERIMENTAL.

RFC1864

[1851] J. Myers and M. Rose. RFC 1864: The content-MD5 header

field, October 1995. URL ftp://ftp.internic.net/rfc/rfc1544.
txt,   ftp://ftp.internic.net/rfc/rfc1864.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1544.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1864.txt. Obsoletes RFC1544 [1534]. Status: DRAFT STANDARD.

RFC1865

[1852] W. Houser, J. Griffin, and C. Hage. RFC 1865: EDI meets the Internet: Frequently asked questions about Electronic Data Interchange (EDI) on the Internet, January 1996. URL ftp://ftp.internic.net/

rfc/rfc1865.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1865.

txt. Status: INFORMATIONAL.

RFC1866

[1853] T. Berners-Lee and D. Connolly. RFC 1866: Hypertext Markup Language — 2.0, November 1995. URL ftp://ftp.internic.net/

rfc/rfc1866.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1866.

txt. Status: PROPOSED STANDARD.

RFC1867

[1854] E. Nebel and L. Masinter. RFC 1867: Form-based file upload in HTML,

November 1995. URL ftp://ftp.internic.net/rfc/rfc1867.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1867.txt. Status: EXPERI-

MENTAL.

RFC1868

[1855] G. Malkin. RFC 1868: ARP extension — UNARP, November

1995. URL ftp://ftp.internic.net/rfc/rfc1868.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1868.txt. Status: EXPERIMEN-

TAL.

RFC1869

[1856] J. Klensin, N. Freed, M. Rose, E. Stefferud, and D. Crocker. RFC 1869: SMTP service extensions, November 1995. URL ftp://ftp.

internic.net/rfc/rfc1651.txt,   ftp://ftp.internic.net/rfc/
rfc1869.txt, ftp://ftp.internic.net/rfc/std10.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1651.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1869.txt,   ftp://ftp.math.utah.edu/pub/rfc/

std10.txt. See also STD0010 [?]. Obsoletes RFC1651 [1641]. Status: STANDARD.

RFC1870

[1857] J. Klensin, N. Freed, and K. Moore. RFC 1870: SMTP service extension for message size declaration, November 1995. URL ftp://ftp.

internic.net/rfc/rfc1653.txt,   ftp://ftp.internic.net/rfc/
rfc1870.txt, ftp://ftp.internic.net/rfc/std10.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1653.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1870.txt,   ftp://ftp.math.utah.edu/pub/rfc/

std10.txt. See also STD0010 [?]. Obsoletes RFC1653 [1643]. Status: STANDARD.

RFC1871

[1858] J. Postel. RFC 1871: Addendum to RFC 1602 — variance pro-

cedure, November 1995. URL ftp://ftp.internic.net/rfc/bcp2.
txt,   ftp://ftp.internic.net/rfc/rfc1602.txt,   ftp://ftp.
internic.net/rfc/rfc1603.txt,   ftp://ftp.internic.net/rfc/

rfc1871.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp2.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1602.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1603.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1871.txt. See also BCP0002 [?]. Updates RFC1602, RFC1603 [1592, 1593]. Status: BEST CURRENT PRACTICE.

RFC1872

[1859] E. Levinson. RFC 1872: The MIME multipart/related content-

type, December 1995. URL ftp://ftp.internic.net/rfc/rfc1872.
txt,   ftp://ftp.internic.net/rfc/rfc2112.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1872.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2112.txt. Obsoleted by RFC2112 [2099]. Status: EXPER-IMENTAL.

RFC1873

[1860] E. Levinson. RFC 1873: Message/external-body content-ID access type,

December 1995. URL ftp://ftp.internic.net/rfc/rfc1873.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1873.txt. Status: EXPERI-

MENTAL.

RFC1874

[1861] E. Levinson. RFC 1874: SGML media types, December 1995. URL ftp:/

/ftp.internic.net/rfc/rfc1874.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1874.txt. Status: EXPERIMENTAL.

RFC1875

[1862] N. Berge. RFC 1875: UNINETT PCA policy statements, Decem-

ber 1995. URL ftp://ftp.internic.net/rfc/rfc1875.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1875.txt. Status: INFORMA-

TIONAL.

RFC1876

[1863] C. Davis, P. Vixie, T. Goodwin, and I. Dickinson. RFC 1876: A means for expressing location information in the domain name system, January

1996. URL ftp://ftp.internic.net/rfc/rfc1034.txt,   ftp://
ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.internic.net/
rfc/rfc1876.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1876.txt.  Updates RFC1034,

RFC1035 [1024, 1025]. Status: EXPERIMENTAL.

RFC1877

[1864] S. Cobb. RFC 1877: PPP Internet Protocol Control Protocol extensions for name server addresses, December 1995. URL ftp://ftp.internic.

net/rfc/rfc1877.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1877.

txt. Status: INFORMATIONAL.

RFC1878

[1865] T. Pummill and B. Manning. RFC 1878: Variable length subnet table for

IPv4, December 1995. URL ftp://ftp.internic.net/rfc/rfc1860.
txt,   ftp://ftp.internic.net/rfc/rfc1878.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1860.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1878.txt. Obsoletes RFC1860 [1847]. Status: INFORMA-TIONAL.

RFC1879

[1866] B. Manning. RFC 1879: Class A subnet experiment results and recommendations, January 1996. URL ftp://ftp.internic.net/

rfc/rfc1879.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1879.

txt. Status: INFORMATIONAL.

RFC1880

[1867] J. Postel. RFC 1880: INTERNET OFFICIAL PROTOCOL STAN-

DARDS, November 1995.  URL ftp://ftp.internic.net/rfc/
rfc1720.txt,   ftp://ftp.internic.net/rfc/rfc1780.txt,   ftp:
//ftp.internic.net/rfc/rfc1800.txt,   ftp://ftp.internic.net/
rfc/rfc1880.txt,   ftp://ftp.internic.net/rfc/rfc1920.txt,
ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.internic.
net/rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/rfc2400.txt,
ftp://ftp.internic.net/rfc/std1.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1720.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1780.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1800.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1880.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1920.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2300.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std1.txt. See also STD0001 [?]. Obso-

leted by RFC1880, RFC1920, RFC2200, RFC2300, STD0001, RFC2400 [1867, 1907, 2187, 2287, ?, 2387]. Obsoletes RFC1780, RFC1720, RFC1800, RFC1880 [1770, 1710, 1790, 1867]. Status: HISTORIC.

RFC1881

[1868] IAB and IESG. RFC 1881: IPv6 address allocation management,

December 1995. URL ftp://ftp.internic.net/rfc/rfc1881.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1881.txt. Status: INFOR-

MATIONAL.

RFC1882

[1869] B. Hancock. RFC 1882: The 12-days of technology before christmas,

December 1995. URL ftp://ftp.internic.net/rfc/rfc1882.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1882.txt. Status: INFOR-

MATIONAL.

RFC1883

[1870] S. Deering and R. Hinden. RFC 1883: Internet Protocol, version 6 (IPv6) specification, December 1995. URL ftp://ftp.internic.net/

rfc/rfc1883.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1883.

txt. Obsoleted by RFC2460 [2447]. Status: PROPOSED STANDARD.

RFC1884

[1871] R. Hinden and S. Deerings. RFC 1884: IP version 6 addressing architec-

ture, December 1995. URL ftp://ftp.internic.net/rfc/rfc1884.
txt,   ftp://ftp.internic.net/rfc/rfc2373.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1884.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2373.txt. Obsoleted by RFC2373 [2360]. Status: PRO-POSED STANDARD.

RFC1885

[1872] A. Conta and S. Deering. RFC 1885: Internet Control Message Protocol (ICMPv6) for the Internet Protocol Version 6 (IPv6), Decem-

ber 1995. URL ftp://ftp.internic.net/rfc/rfc1885.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1885.txt. Obsoleted by RFC2463

[2360]. Status: PROPOSED STANDARD.

RFC1886

[1873] S. Thomson and C. Huitema. RFC 1886: DNS extensions to support IP version 6, December 1995. URL ftp://ftp.internic.net/

rfc/rfc1886.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1886.

txt. Status: PROPOSED STANDARD.

RFC1887

[1874] Y. Rekhter and T. Lis. RFC 1887: An architecture for IPv6 unicast address allocation, December 1995. URL ftp://ftp.internic.net/

rfc/rfc1887.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1887.

txt. Status: INFORMATIONAL.

RFC1888

[1875] J. Bound, B. Carpenter, D. Harrington, J. Houldsworth, and A. Lloyd. RFC 1888: OSI NSAPs and IPv6, August 1996. URL ftp://ftp.

internic.net/rfc/rfc1888.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1888.txt. Status: EXPERIMENTAL.

RFC1889

[1876] Audio-Video Transport Working Group, H. Schulzrinne, S. Casner, R. Frederick, and V. Jacobson. RFC 1889: RTP: A transport protocol for real-time applications, January 1996. URL ftp://ftp.internic.net/

rfc/rfc1889.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1889.

txt. Status: PROPOSED STANDARD.

RFC1890

[1877] Audio-Video Transport Working Group and H. Schulzrinne. RFC 1890: RTP profile for audio and video conferences with minimal control, Jan-

uary 1996. URL ftp://ftp.internic.net/rfc/rfc1890.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1890.txt. Status: PROPOSED

STANDARD.

RFC1891

[1878] K. Moore. RFC 1891: SMTP service extension for delivery status notifi-

cations, January 1996. URL ftp://ftp.internic.net/rfc/rfc1891.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1891.txt. Status:

PROPOSED STANDARD.

RFC1892

[1879] G. Vaudreuil. RFC 1892: The multipart/report content type for the reporting of mail system administrative messages, January

1996. URL ftp://ftp.internic.net/rfc/rfc1892.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1892.txt. Status: PROPOSED

STANDARD.

RFC1893

[1880] G. Vaudreuil. RFC 1893: Enhanced mail system status codes, Jan-

uary 1996. URL ftp://ftp.internic.net/rfc/rfc1893.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1893.txt. Status: PROPOSED

STANDARD.

RFC1894

[1881] K. Moore and G. Vaudreuil. RFC 1894: An extensible message format for delivery status notifications, January 1996. URL ftp://ftp.internic.

net/rfc/rfc1894.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1894.

txt. Status: PROPOSED STANDARD.

RFC1895

[1882] E. Levinson. RFC 1895: The application/CALS-1840 content-type,

February 1996. URL ftp://ftp.internic.net/rfc/rfc1895.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1895.txt. Status: INFOR-

MATIONAL.

RFC1896

[1883] P. Resnick and A. Walker. RFC 1896: The text/enriched MIME content-

type, February 1996. URL ftp://ftp.internic.net/rfc/rfc1523.
txt,   ftp://ftp.internic.net/rfc/rfc1563.txt,   ftp://ftp.
internic.net/rfc/rfc1896.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1523.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1563.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1896.txt. Obsoletes

RFC1523, RFC1563 [1513, 1553]. Status: INFORMATIONAL.

RFC1897

[1884] R. Hinden and J. Postel. RFC 1897: IPv6 testing address allocation, Jan-

uary 1996. URL ftp://ftp.internic.net/rfc/rfc1897.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1897.txt. Obsoleted by RFC2471

[2458]. Status: EXPERIMENTAL.

RFC1898

[1885] D. Eastlake, 3rd, B. Boesch, S. Crocker, and M. Yesil. RFC 1898: Cyber-Cash Credit Card Protocol version 0.8, February 1996. URL ftp://ftp.

internic.net/rfc/rfc1898.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1898.txt. Status: INFORMATIONAL.

RFC1899

[1886] J. Elliott. RFC 1899: Request for comments summary RFC numbers 1800–1899, January 1997. URL ftp://ftp.internic.net/

rfc/rfc1899.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1899.

txt. Status: INFORMATIONAL.

RFC1900

[1887] B. Carpenter and Y. Rekhter. RFC 1900: Renumbering needs work,

February 1996. URL ftp://ftp.internic.net/rfc/rfc1900.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1900.txt. Status: INFOR-

MATIONAL.

RFC1901

[1888] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1901: Introduction to community-based SNMPv2, Jan-

uary 1996. URL ftp://ftp.internic.net/rfc/rfc1901.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1901.txt. Status: EXPERIMEN-

TAL.

RFC1902

[1889] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1902: Structure of management information for version 2 of the Simple Network Management Protocol (SN-

MPv2), January 1996. URL ftp://ftp.internic.net/rfc/rfc1442.
txt,   ftp://ftp.internic.net/rfc/rfc1902.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1442.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1902.txt. Obsoletes RFC1442 [1432]. Status: DRAFT STANDARD.

RFC1903

[1890] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1903: Textual conventions for version 2 of the Simple Network Management Protocol (SNMPv2), January 1996. URL ftp://

ftp.internic.net/rfc/rfc1443.txt,   ftp://ftp.internic.net/
rfc/rfc1903.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1443.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1903.txt. Obsoletes

RFC1443 [1433]. Status: DRAFT STANDARD.

RFC1904

[1891] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1904: Conformance statements for version 2 of the Simple Network Management Protocol (SNMPv2), January 1996. URL ftp://

ftp.internic.net/rfc/rfc1444.txt,   ftp://ftp.internic.net/
rfc/rfc1904.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1444.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1904.txt. Obsoletes

RFC1444 [1434]. Status: DRAFT STANDARD.

RFC1905

[1892] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1905: Protocol operations for version 2 of the Simple Network Management Protocol (SNMPv2), January 1996. URL ftp://

ftp.internic.net/rfc/rfc1448.txt,   ftp://ftp.internic.net/
rfc/rfc1905.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1448.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1905.txt. Obsoletes

RFC1448 [1438]. Status: DRAFT STANDARD.

RFC1906

[1893] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1906: Transport mappings for version 2 of the Simple Network Management Protocol (SNMPv2), January 1996. URL ftp://

ftp.internic.net/rfc/rfc1449.txt,   ftp://ftp.internic.net/
rfc/rfc1906.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1449.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1906.txt. Obsoletes

RFC1449 [1439]. Status: DRAFT STANDARD.

RFC1907

[1894] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1907: Management information base for version 2 of the Simple Network Management Protocol (SNMPv2), January

1996. URL ftp://ftp.internic.net/rfc/rfc1450.txt,   ftp://
ftp.internic.net/rfc/rfc1907.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1450.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1907.

txt. Obsoletes RFC1450 [1440]. Status: DRAFT STANDARD.

RFC1908

[1895] SNMPv2 Working Group, J. Case, K. McCloghrie, M. Rose, and S. Waldbusser. RFC 1908: Coexistence between version 1 and version 2 of the Internet-standard Network Management Frame-

work, January 1996. URL ftp://ftp.internic.net/rfc/rfc1452.
txt,   ftp://ftp.internic.net/rfc/rfc1908.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1452.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1908.txt. Obsoletes RFC1452 [1442]. Status: DRAFT STANDARD.

RFC1909

[1896] K. McCloghrie. RFC 1909: An administrative infrastructure for SN-

MPv2, February 1996. URL ftp://ftp.internic.net/rfc/rfc1909.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1909.txt. Status:

EXPERIMENTAL.

RFC1910

[1897] G. Waters. RFC 1910: User-based security model for SNMPv2, Febru-

ary 1996. URL ftp://ftp.internic.net/rfc/rfc1910.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1910.txt. Status: EXPERIMEN-

TAL.

RFC1911

[1898] G. Vaudreuil. RFC 1911: Voice profile for Internet mail, February

1996. URL ftp://ftp.internic.net/rfc/rfc1911.txt,   ftp://
ftp.internic.net/rfc/rfc2421.txt,   ftp://ftp.internic.net/
rfc/rfc2422.txt,   ftp://ftp.internic.net/rfc/rfc2423.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1911.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2421.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2422.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2423.

txt. Obsoleted by RFC2421, RFC2422, RFC2423 [2408, 2409, 2410]. Status: EXPERIMENTAL.

RFC1912

[1899] D. Barr. RFC 1912: Common DNS operational and configuration er-

rors, February 1996. URL ftp://ftp.internic.net/rfc/rfc1537.
txt,   ftp://ftp.internic.net/rfc/rfc1912.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1537.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1912.txt. Obsoletes RFC1537 [1527]. Status: INFORMA-TIONAL.

RFC1913

[1900] C. Weider, J. Fullton, and S. Spero. RFC 1913: Architecture of the WHOIS++ index service, February 1996. URL ftp://ftp.internic.

net/rfc/rfc1913.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1913.

txt. Status: PROPOSED STANDARD.

RFC1914

[1901] P. Faltstrom, R. Schoultz, and C. Weider. RFC 1914: How to interact with a Whois++ mesh, February 1996. URL ftp://ftp.internic.net/

rfc/rfc1914.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1914.

txt. Status: PROPOSED STANDARD.

RFC1915

[1902] F. Kastenholz. RFC 1915: Variance for the PPP connection control protocol and the PPP encryption control protocol, February

1996. URL ftp://ftp.internic.net/rfc/bcp3.txt,   ftp://ftp.
internic.net/rfc/rfc1915.txt,   ftp://ftp.math.utah.edu/pub/
rfc/bcp3.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1915.txt.

See also BCP0003 [?]. Status: BEST CURRENT PRACTICE.

RFC1916

[1903] H. Berkowitz, P. Ferguson, W. Leland, and P. Nesser. RFC 1916: Enterprise renumbering: Experience and information solicitation, Febru-

ary 1996. URL ftp://ftp.internic.net/rfc/rfc1916.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1916.txt. Status: INFORMA-

TIONAL.

RFC1917

[1904] P. Nesser, II. RFC 1917: An appeal to the Internet community to return unused IP networks (prefixes) to the IANA, February

1996. URL ftp://ftp.internic.net/rfc/bcp4.txt,   ftp://ftp.
internic.net/rfc/rfc1917.txt,   ftp://ftp.math.utah.edu/pub/
rfc/bcp4.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1917.txt.

See also BCP0004 [?]. Status: BEST CURRENT PRACTICE.

RFC1918

[1905] Y. Rekhter, B. Moskowitz, D. Karrenberg, G. J. de Groot, and E. Lear. RFC 1918: Address allocation for private internets, February

1996. URL ftp://ftp.internic.net/rfc/bcp5.txt,   ftp://ftp.
internic.net/rfc/rfc1597.txt,   ftp://ftp.internic.net/rfc/
rfc1627.txt,   ftp://ftp.internic.net/rfc/rfc1918.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/bcp5.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1597.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1627.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1918.txt.

See also BCP0005 [?]. Obsoletes RFC1627, RFC1597 [1617, 1587]. Status: BEST CURRENT PRACTICE.

RFC1919

[1906] M. Chatel. RFC 1919: Classical versus transparent IP proxies,

March 1996. URL ftp://ftp.internic.net/rfc/rfc1919.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1919.txt. Status: INFOR-

MATIONAL.

RFC1920

[1907] J. Postel. RFC 1920: Internet Official Protocol Standards, March

1996. URL ftp://ftp.internic.net/rfc/rfc1720.txt,   ftp://
ftp.internic.net/rfc/rfc1780.txt,   ftp://ftp.internic.net/
rfc/rfc1800.txt,   ftp://ftp.internic.net/rfc/rfc1880.txt,
ftp://ftp.internic.net/rfc/rfc1920.txt,   ftp://ftp.internic.
net/rfc/rfc2000.txt,   ftp://ftp.internic.net/rfc/rfc2109.txt,
 ftp://ftp.internic.net/rfc/rfc2200.txt,   ftp://ftp.internic.
net/rfc/rfc2300.txt,   ftp://ftp.internic.net/rfc/rfc2400.txt,
ftp://ftp.internic.net/rfc/std1.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1720.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1780.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1800.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1880.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1920.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2000.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2109.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.edu/pub/rfc/

std1.txt. See also STD0001 [?]. Obsoleted by RFC2000, RFC2200, RFC2300, RFC2109, STD0001, RFC2400 [1987, 2187, 2287, 2096, ?, 2387]. Obsoletes RFC1780, RFC1720, RFC1800, RFC1880 [1770, 1710, 1790, 1867]. Status: HISTORIC.

RFC1921

[1908] J. Dujonc. RFC 1921: TNVIP protocol, March 1996. URL ftp://ftp.

internic.net/rfc/rfc1921.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1921.txt. Status: INFORMATIONAL. Not online.

RFC1922

[1909] H. F. Zhu, D. Y. Hu, Z. G. Wang, T. C. Kao, W. C. H. Chang, and M. Crispin. RFC 1922: Chinese character encoding for Internet mes-

sages, March 1996. URL ftp://ftp.internic.net/rfc/rfc1922.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1922.txt. Status: INFOR-

MATIONAL.

RFC1923

[1910] J. Halpern and S. Bradner. RFC 1923: RIPv1 applicability statement for historic status, March 1996. URL ftp://ftp.internic.net/

rfc/rfc1923.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1923.

txt. Status: INFORMATIONAL.

RFC1924

[1911] R. Elz. RFC 1924: A compact representation of IPv6 addresses,

April 1996. URL ftp://ftp.internic.net/rfc/rfc1924.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1924.txt. Status: INFOR-

MATIONAL.

RFC1925

[1912] R. Callon. RFC 1925: The twelve networking truths, April

1996. URL ftp://ftp.internic.net/rfc/rfc1925.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1925.txt. Status: INFORMA-

TIONAL.

RFC1926

[1913] J. Eriksson. RFC 1926: An experimental encapsulation of IP datagrams on top of ATM, April 1996. URL ftp://ftp.internic.net/

rfc/rfc1926.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1926.

txt. Status: INFORMATIONAL.

RFC1927

[1914] C. Rogers. RFC 1927: Suggested additional MIME types for associating

documents, April 1996. URL ftp://ftp.internic.net/rfc/rfc1927.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1927.txt. Status:

INFORMATIONAL.

RFC1928

[1915] M. Leech, M. Ganis, Y. Lee, R. Kuris, D. Koblas, and L. Jones. RFC 1928: SOCKS protocol version 5, April 1996. URL ftp://ftp.

internic.net/rfc/rfc1928.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1928.txt. Status: PROPOSED STANDARD.

RFC1929

[1916] M. Leech. RFC 1929: Username/password authentication for SOCKS

V5, April 1996. URL ftp://ftp.internic.net/rfc/rfc1929.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1929.txt. Status: PRO-

POSED STANDARD.

RFC1930

[1917] J. Hawkinson and T. Bates. RFC 1930: Guidelines for creation, selection, and registration of an autonomous system (AS), March

1996. URL ftp://ftp.internic.net/rfc/bcp6.txt,   ftp://ftp.
internic.net/rfc/rfc1930.txt,   ftp://ftp.math.utah.edu/pub/
rfc/bcp6.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1930.txt.

See also BCP0006 [?]. Status: BEST CURRENT PRACTICE.

RFC1931

[1918] D. Brownell. RFC 1931: Dynamic RARP extensions for automatic network address acquisition, April 1996. URL ftp://ftp.internic.net/

rfc/rfc1931.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1931.

txt. Status: INFORMATIONAL.

RFC1932

[1919] R. Cole, D. Shur, and C. Villamizar. RFC 1932: IP over ATM: A framework document, April 1996. URL ftp://ftp.internic.net/

rfc/rfc1932.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1932.

txt. Status: INFORMATIONAL.

RFC1933

[1920] R. Gilligan and E. Nordmark. RFC 1933: Transition mechanisms for IPv6 hosts and routers, April 1996. URL ftp://ftp.internic.net/

rfc/rfc1933.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1933.

txt. Status: PROPOSED STANDARD.

RFC1934

[1921] K. Smith. RFC 1934: Ascend’s multilink protocol plus (MP+),

April 1996. URL ftp://ftp.internic.net/rfc/rfc1934.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1934.txt. Status: INFOR-

MATIONAL.

RFC1935

[1922] J. Quarterman and S. Carl-Mitchell. RFC 1935: What is the Internet,

anyway?, April 1996. URL ftp://ftp.internic.net/rfc/rfc1935.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1935.txt. Status:

INFORMATIONAL.

RFC1936

[1923] J. Touch and B. Parham. RFC 1936: Implementing the Internet checksum in hardware, April 1996. URL ftp://ftp.internic.net/

rfc/rfc1936.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1936.

txt. Status: INFORMATIONAL.

RFC1937

[1924] Y. Rekhter and D. Kandlur. RFC 1937: “Local/Remote” forwarding decision in switched data link subnetworks, May 1996. URL ftp://ftp.

internic.net/rfc/rfc1937.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1937.txt. Status: INFORMATIONAL.

RFC1938

[1925] N. Haller and C. Metz. RFC 1938: A one-time password sys-

tem, May 1996.  URL ftp://ftp.internic.net/rfc/rfc1938.
txt,   ftp://ftp.internic.net/rfc/rfc2289.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1938.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2289.txt. Obsoleted by RFC2289 [2276]. Status: PRO-POSED STANDARD.

RFC1939

[1926] J. Myers and M. Rose. RFC 1939: Post Office Protocol — version 3, May

1996. URL ftp://ftp.internic.net/rfc/rfc1725.txt,   ftp://
ftp.internic.net/rfc/rfc1939.txt,   ftp://ftp.internic.net/
rfc/rfc1957.txt, ftp://ftp.internic.net/rfc/std53.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1725.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1939.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1957.txt,   ftp://ftp.math.utah.edu/pub/rfc/std53.txt. See

also STD0053 [?]. Obsoletes RFC1725 [1715]. Updated by RFC1957, RFC2449 [1944, 2436]. Status: STANDARD.

RFC1940

[1927] D. Estrin, T. Li, Y. Rekhter, K. Varadhan, and D. Zappala. RFC 1940: Source demand routing: Packet format and forwarding specification (ver-

sion 1), May 1996. URL ftp://ftp.internic.net/rfc/rfc1940.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1940.txt. Status: INFOR-

MATIONAL.

RFC1941

[1928] J. Sellers and J. Robichaux. RFC 1941: Frequently asked questions

for schools, May 1996. URL ftp://ftp.internic.net/rfc/fyi22.
txt,   ftp://ftp.internic.net/rfc/rfc1578.txt,   ftp://ftp.
internic.net/rfc/rfc1941.txt,   ftp://ftp.math.utah.edu/pub/

rfc/fyi22.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1578.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1941.txt. See also FYI0022

[?]. Obsoletes RFC1578 [1568]. Status: INFORMATIONAL.

RFC1942

[1929] D. Raggett. RFC 1942: HTML tables, May 1996. URL ftp://ftp.

internic.net/rfc/rfc1942.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1942.txt. Status: EXPERIMENTAL.

RFC1943

[1930] B. Jennings. RFC 1943: Building an X.500 directory service in the

US, May 1996. URL ftp://ftp.internic.net/rfc/rfc1943.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1943.txt. Status: INFOR-

MATIONAL.

RFC1944

[1931] S. Bradner and J. McQuaid. RFC 1944: Benchmarking methodology for network interconnect devices, May 1996. URL ftp://ftp.internic.

net/rfc/rfc1944.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1944.

txt. Status: INFORMATIONAL.

RFC1945

[1932] T. Berners-Lee, R. Fielding, and H. Frystyk. RFC 1945: Hypertext Transfer Protocol — HTTP/1.0, May 1996. URL ftp://ftp.internic.

net/rfc/rfc1945.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1945.

txt. Status: INFORMATIONAL.

RFC1946

[1933] S. Jackowski. RFC 1946: Native ATM support for ST2+, May

1996. URL ftp://ftp.internic.net/rfc/rfc1946.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1946.txt. Status: INFORMA-

TIONAL.

RFC1947

[1934] D. Spinellis. RFC 1947: Greek character encoding for electronic mail

messages, May 1996. URL ftp://ftp.internic.net/rfc/rfc1947.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1947.txt. Status:

INFORMATIONAL.

RFC1948

[1935] S. Bellovin. RFC 1948: Defending against sequence number attacks,

May 1996. URL ftp://ftp.internic.net/rfc/rfc1948.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1948.txt. Status: INFORMA-

TIONAL.

RFC1949

[1936] A. Ballardie. RFC 1949: Scalable multicast key distribution, May

1996. URL ftp://ftp.internic.net/rfc/rfc1949.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc1949.txt. Status: EXPERIMEN-

TAL.

RFC1950

[1937] L. P. Deutsch and J-L. Gailly. RFC 1950: ZLIB compressed data format specification version 3.3, May 1996. URL ftp://ftp.internic.net/

rfc/rfc1950.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1950.

txt. Status: INFORMATIONAL.

RFC1951

[1938] L. P. Deutsch. RFC 1951: DEFLATE compressed data format specification version 1.3, May 1996. URL ftp://ftp.internic.net/

rfc/rfc1951.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1951.

txt. Status: INFORMATIONAL.

RFC1952

[1939] L. P. Deutsch. RFC 1952: GZIP file format specification version

4.3, May 1996. URL ftp://ftp.internic.net/rfc/rfc1952.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1952.txt. Status: INFOR-

MATIONAL.

RFC1953

[1940] P. Newman, W. Edwards, R. Hinden, E. Hoffman, F. Ching Liaw, T. Lyon, and G. Minshall. RFC 1953: Ipsilon Flow Management Protocol Specification for IPv4 version 1.0, May 1996. URL ftp://ftp.

internic.net/rfc/rfc1953.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1953.txt. Status: INFORMATIONAL.

RFC1954

[1941] P. Newman, W. Edwards, R. Hinden, E. Hoffman, F. Ching Liaw, T. Lyon, and G. Minshall. RFC 1954: Transmission of flow labelled IPv4 on ATM data links ipsilon version 1.0, May

1996. URL ftp://ftp.internic.net/rfc/rfc1954.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1954.txt. Status: INFORMA-

TIONAL.

RFC1955

[1942] R. Hinden. RFC 1955: New scheme for Internet routing and addressing (ENCAPS) for IPNG, June 1996. URL ftp://ftp.internic.net/

rfc/rfc1955.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1955.

txt. Status: INFORMATIONAL.

RFC1956

[1943] D. Engebretson and R. Plzak. RFC 1956: Registration in the MIL do-

main, June 1996. URL ftp://ftp.internic.net/rfc/rfc1956.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1956.txt. Status: INFOR-

MATIONAL.

RFC1957

[1944] R. Nelson. RFC 1957: Some observations on implementations of the Post Office Protocol (POP3), June 1996. URL ftp://

ftp.internic.net/rfc/rfc1939.txt,   ftp://ftp.internic.net/
rfc/rfc1957.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1939.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1957.txt. Updates

RFC1939 [1926]. Status: INFORMATIONAL.

RFC1958

[1945] B. Carpenter. RFC 1958: Architectural principles of the Internet,

June 1996. URL ftp://ftp.internic.net/rfc/rfc1958.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1958.txt. Status: INFORMA-

TIONAL.

RFC1959

[1946] T. Howes and M. Smith. RFC 1959: An LDAP URL format,

June 1996. URL ftp://ftp.internic.net/rfc/rfc1959.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1959.txt. Status: PROPOSED

STANDARD.

RFC1960

[1947] T. Howes. RFC 1960: A string representation of LDAP search

filters, June 1996. URL ftp://ftp.internic.net/rfc/rfc1558.
txt,   ftp://ftp.internic.net/rfc/rfc1960.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1558.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1960.txt. Obsoletes RFC1558 [1548]. Status: PROPOSED STANDARD.

RFC1961

[1948] P. McMahon. RFC 1961: GSS-API authentication method for SOCKS

version 5, June 1996. URL ftp://ftp.internic.net/rfc/rfc1961.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1961.txt. Status:

PROPOSED STANDARD.

RFC1962

[1949] D. Rand. RFC 1962: The PPP compression control protocol

(CCP), June 1996. URL ftp://ftp.internic.net/rfc/rfc1962.

txt,   ftp://ftp.internic.net/rfc/rfc2153.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1962.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2153.txt. Updated by RFC2153 [2140]. Status: PRO-POSED STANDARD.

RFC1963

[1950] K. Schneider and S. Venters. RFC 1963: PPP serial data transport protocol (SDTP), August 1996. URL ftp://ftp.internic.net/

rfc/rfc1963.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1963.

txt. Status: INFORMATIONAL.

RFC1964

[1951] J. Linn. RFC 1964: The Kerberos Version 5 GSS-API mechanism,

June 1996. URL ftp://ftp.internic.net/rfc/rfc1964.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1964.txt. Status: PROPOSED

STANDARD.

RFC1965

[1952] P. Traina. RFC 1965: Autonomous system confederations for BGP,

June 1996. URL ftp://ftp.internic.net/rfc/rfc1965.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1965.txt. Status: EXPERIMEN-

TAL.

RFC1966

[1953] T. Bates and R. Chandrasekeran. RFC 1966: BGP route reflection an alternative to full mesh IBGP, June 1996. URL ftp://ftp.internic.

net/rfc/rfc1966.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1966.

txt. Status: EXPERIMENTAL.

RFC1967

[1954] K. Schneider and R. Friend. RFC 1967: PPP LZS-DCP compression protocol (LZS-DCP), August 1996. URL ftp://ftp.internic.net/

rfc/rfc1967.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1967.

txt. Status: INFORMATIONAL.

RFC1968

[1955] G. Meyer. RFC 1968: The PPP encryption control protocol (ECP),

June 1996. URL ftp://ftp.internic.net/rfc/rfc1968.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1968.txt. Status: PROPOSED

STANDARD.

RFC1969

[1956] K. Sklower and G. Meyer. RFC 1969: The PPP DES encryption proto-

col (DESE), June 1996. URL ftp://ftp.internic.net/rfc/rfc1969.

txt,   ftp://ftp.internic.net/rfc/rfc2419.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1969.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2419.txt. Obsoleted by RFC2419 [2406]. Status: INFOR-MATIONAL.

RFC1970

[1957] T. Narten, E. Nordmark, and W. Simpson. RFC 1970: Neighbor discovery for IP version 6 (IPv6), August 1996. URL ftp://ftp.internic.

net/rfc/rfc1970.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1970.

txt. Obsoleted by RFC2461 [2448]. Status: PROPOSED STANDARD.

RFC1971

[1958] S. Thomson and T. Narten. RFC 1971: IPv6 stateless address autocon-

figuration, August 1996. URL ftp://ftp.internic.net/rfc/rfc1971.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1971.txt. Obsoleted

by RFC2462 [2449]. Status: PROPOSED STANDARD.

RFC1972

[1959] M. Crawford. RFC 1972: A method for the transmission of IPv6 packets over Ethernet networks, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1972.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1972.

txt. Obsoleted by RFC2464 [2451]. Status: PROPOSED STANDARD.

RFC1973

[1960] W. Simpson. RFC 1973: PPP in frame relay, June 1996. URL ftp://

ftp.internic.net/rfc/rfc1973.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1973.txt. Status: PROPOSED STANDARD.

RFC1974

[1961] R. Friend and W. Simpson. RFC 1974: PPP stac LZS compression proto-

col, August 1996. URL ftp://ftp.internic.net/rfc/rfc1974.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1974.txt. Status: INFOR-

MATIONAL.

RFC1975

[1962] D. Schremp, J. Black, and J. Weiss. RFC 1975: PPP magnalink variable resource compression, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1975.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1975.

txt. Status: INFORMATIONAL.

RFC1976

[1963] K. Schneider and S. Venters. RFC 1976: PPP for data compression in data circuit-terminating equipment (DCE), August 1996. URL ftp://

ftp.internic.net/rfc/rfc1976.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1976.txt. Status: INFORMATIONAL.

RFC1977

[1964] V. Schryver. RFC 1977: PPP BSD compression protocol, Au-

gust 1996. URL ftp://ftp.internic.net/rfc/rfc1977.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1977.txt. Status: INFORMA-

TIONAL.

RFC1978

[1965] D. Rand. RFC 1978: PPP predictor compression protocol, Au-

gust 1996. URL ftp://ftp.internic.net/rfc/rfc1978.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1978.txt. Status: INFORMA-

TIONAL.

RFC1979

[1966] J. Woods. RFC 1979: PPP Deflate Protocol, August 1996. URL ftp://

ftp.internic.net/rfc/rfc1979.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc1979.txt. Status: INFORMATIONAL.

RFC1980

[1967] J. Seidman. RFC 1980: A proposed extension to HTML: Client-side im-

age maps, August 1996. URL ftp://ftp.internic.net/rfc/rfc1980.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1980.txt. Status:

INFORMATIONAL.

RFC1981

[1968] J. McCann, S. Deering, and J. Mogul. RFC 1981: Path MTU discovery for IP version 6, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1981.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1981.

txt. Status: PROPOSED STANDARD.

RFC1982

[1969] R. Elz and R. Bush. RFC 1982: Serial number arithmetic, August

1996. URL ftp://ftp.internic.net/rfc/rfc1034.txt,   ftp://
ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.internic.net/
rfc/rfc1982.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1982.txt.  Updates RFC1034,

RFC1035 [1024, 1025]. Status: PROPOSED STANDARD.

RFC1983

[1970] G. Malkin. RFC 1983: Internet users’ glossary, August 1996. URL ftp:

//ftp.internic.net/rfc/fyi18.txt,   ftp://ftp.internic.net/

rfc/rfc1392.txt,   ftp://ftp.internic.net/rfc/rfc1983.txt,
ftp://ftp.math.utah.edu/pub/rfc/fyi18.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1392.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1983.txt. See also FYI0018 [?]. Obsoletes RFC1392 [1382]. Status: INFORMATIONAL.

RFC1984

[1971] IAB and IESG. RFC 1984: IAB and IESG statement on cryptographic technology and the Internet, August 1996. URL ftp://ftp.internic.

net/rfc/rfc1984.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1984.

txt. Status: INFORMATIONAL.

RFC1985

[1972] J. De Winter. RFC 1985: SMTP service extension for remote message queue starting, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1985.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1985.

txt. Status: PROPOSED STANDARD.

RFC1986

[1973] W. Polites, W. Wollman, D. Woo, and R. Langan. RFC 1986: Experiments with a simple file transfer protocol for radio links using enhanced trivial file transfer protocol (ETFTP), August 1996. URL ftp://ftp.

internic.net/rfc/rfc1986.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1986.txt. Status: EXPERIMENTAL.

RFC1987

[1974] P. Newman, W. Edwards, R. Hinden, E. Hoffman, F. Ching Liaw, T. Lyon, and G. Minshall. RFC 1987: Ipsilon’s general switch management protocol specification version 1.1, August 1996. URL ftp://

ftp.internic.net/rfc/rfc1987.txt,   ftp://ftp.internic.net/
rfc/rfc2297.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1987.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2297.txt. Updated

by RFC2297 [2284]. Status: INFORMATIONAL.

RFC1988

[1975] G. McAnally, D. Gilbert, and J. Flick. RFC 1988: Conditional grant of rights to specific Hewlett–Packard patents in conjunction with the Internet Engineering Task Force’s Internet-Standard Network Management Framework, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1988.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1988.

txt. Status: INFORMATIONAL.

RFC1989

[1976] W. Simpson. RFC 1989: PPP link quality monitoring, August

1996. URL ftp://ftp.internic.net/rfc/rfc1333.txt,   ftp://
ftp.internic.net/rfc/rfc1989.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1333.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1989.

txt. Obsoletes RFC1333 [1323]. Status: DRAFT STANDARD.

RFC1990

[1977] K. Sklower, B. Lloyd, G. McGregor, D. Carr, and T. Coradetti. RFC 1990: The PPP multilink protocol (MP), August 1996. URL ftp://

ftp.internic.net/rfc/rfc1717.txt,   ftp://ftp.internic.net/
rfc/rfc1990.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1717.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1990.txt. Obsoletes

RFC1717 [1707]. Status: DRAFT STANDARD.

RFC1991

[1978] D. Atkins, W. Stallings, and P. Zimmermann. RFC 1991: PGP message exchange formats, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1991.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1991.

txt. Status: INFORMATIONAL.

RFC1992

[1979] I. Castineyra, N. Chiappa, and M. Steenstrup. RFC 1992: The nimrod routing architecture, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1992.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1992.

txt. Status: INFORMATIONAL.

RFC1993

[1980] A. Barbir, D. Carr, and W. Simpson. RFC 1993: PPP Gandalf FZA compression protocol, August 1996. URL ftp://ftp.internic.net/

rfc/rfc1993.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1993.

txt. Status: INFORMATIONAL.

RFC1994

[1981] W. Simpson. RFC 1994: PPP challenge handshake authentication protocol (CHAP), August 1996. URL ftp://ftp.internic.net/rfc/

rfc1334.txt,   ftp://ftp.internic.net/rfc/rfc1994.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1334.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc1994.txt. Obsoletes RFC1334 [1324]. Status: DRAFT STANDARD.

RFC1995

[1982] M. Ohta. RFC 1995: Incremental zone transfer in DNS, August

1996. URL ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://

ftp.internic.net/rfc/rfc1995.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1995.

txt. Updates RFC1035 [1025]. Status: PROPOSED STANDARD.

RFC1996

[1983] P. Vixie. RFC 1996: A mechanism for prompt notification of zone changes (DNS NOTIFY), August 1996. URL ftp://ftp.

internic.net/rfc/rfc1035.txt,   ftp://ftp.internic.net/rfc/
rfc1996.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1996.txt. Updates RFC1035

[1025]. Status: PROPOSED STANDARD.

RFC1997

[1984] R. Chandra, P. Traina, and T. Li. RFC 1997: BGP communities at-

tribute, August 1996. URL ftp://ftp.internic.net/rfc/rfc1997.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1997.txt. Status:

PROPOSED STANDARD.

RFC1998

[1985] E. Chen and T. Bates. RFC 1998: An application of the BGP community attribute in multi-home routing, August 1996. URL ftp://ftp.

internic.net/rfc/rfc1998.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc1998.txt. Status: INFORMATIONAL.

RFC1999

[1986] J. Elliott. RFC 1999: Request for comments summary RFC numbers 1900–1999, January 1997. URL ftp://ftp.internic.net/

rfc/rfc1999.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1999.

txt. Status: INFORMATIONAL.

RFC2000

[1987] J. Postel. RFC 2000: Internet Official Protocol Standards, February

1997. URL ftp://ftp.internic.net/rfc/rfc1920.txt,   ftp://
ftp.internic.net/rfc/rfc2000.txt,   ftp://ftp.internic.net/
rfc/rfc2200.txt,   ftp://ftp.internic.net/rfc/rfc2300.txt,
ftp://ftp.internic.net/rfc/rfc2400.txt, ftp://ftp.internic.
net/rfc/std1.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1920.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2000.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2400.txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt. See

also STD0001 [?]. Obsoleted by RFC2200, STD0001, RFC2300, RFC2400 [2187, ?, 2287, 2387]. Obsoletes RFC1920, RFC1880, RFC1800, RFC1780, RFC1720, RFC1610, RFC1600, RFC1540, RFC1500, RFC1410, RFC1360, RFC1280, RFC1250, RFC1200, RFC1140, RFC1130, RFC1100, RFC1083 [1907, 1867, 1790, 1770, 1710, 1600, 1590, 1530, 1490, 1400, 1350, 1270, 1240, 1190, 1130, 1120, 1090, 1073]. Status: HISTORIC.

RFC2001

[1988] W. Stevens. RFC 2001: TCP slow start, congestion avoidance, fast retransmit, and fast recovery algorithms, January 1997. URL ftp://ftp.

internic.net/rfc/rfc2001.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2001.txt. Status: PROPOSED STANDARD.

RFC2002

[1989] C. Perkins. RFC 2002: IP mobility support, October 1996. URL ftp://

ftp.internic.net/rfc/rfc2002.txt,   ftp://ftp.internic.net/
rfc/rfc2290.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2002.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2290.txt. Updated

by RFC2290 [2277]. Status: PROPOSED STANDARD.

RFC2003

[1990] C. Perkins. RFC 2003: IP encapsulation within IP, October

1996. URL ftp://ftp.internic.net/rfc/rfc2003.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2003.txt. Status: PROPOSED

STANDARD.

RFC2004

[1991] C. Perkins. RFC 2004: Minimal encapsulation within IP, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2004.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2004.txt. Status: PROPOSED

STANDARD.

RFC2005

[1992] J. Solomon. RFC 2005: Applicability statement for IP mobility support,

October 1996. URL ftp://ftp.internic.net/rfc/rfc2005.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2005.txt. Status: PRO-

POSED STANDARD.

RFC2006

[1993] D. Cong, M. Hamlen, and C. Perkins. RFC 2006: The definitions of managed objects for IP mobility support using SMIv2, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2006.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2006.txt. Status: PROPOSED

STANDARD.

RFC2007

[1994] J. Foster, M. Isaacs, and M. Prior. RFC 2007: Catalogue of network training materials, October 1996. URL ftp://ftp.internic.net/rfc/

fyi29.txt,   ftp://ftp.internic.net/rfc/rfc2007.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/fyi29.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2007.txt. See also FYI0029 [?]. Status: INFORMA-TIONAL.

RFC2008

[1995] Y. Rekhter and T. Li. RFC 2008: Implications of various address allocation policies for Internet routing, October 1996. URL ftp://ftp.

internic.net/rfc/bcp7.txt,   ftp://ftp.internic.net/rfc/rfc2008.
txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp7.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2008.txt. See also BCP0007 [?]. Status:

BEST CURRENT PRACTICE.

RFC2009

[1996] T. Imielinski and J. Navas. RFC 2009: GPS-based addressing and rout-

ing, November 1996. URL ftp://ftp.internic.net/rfc/rfc2009.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2009.txt. Status:

EXPERIMENTAL.

RFC2010

[1997] B. Manning and P. Vixie. RFC 2010: Operational criteria for root name

servers, October 1996. URL ftp://ftp.internic.net/rfc/rfc2010.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2010.txt. Status:

INFORMATIONAL.

RFC2011

[1998] K. McCloghrie. RFC 2011: SNMPv2 management information base for the Internet Protocol using SMIv2, November 1996. URL ftp://

ftp.internic.net/rfc/rfc1213.txt,   ftp://ftp.internic.net/
rfc/rfc2011.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1213.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2011.txt. Updates

RFC1213 [1203]. Status: PROPOSED STANDARD.

RFC2012

[1999] K. McCloghrie. RFC 2012: SNMPv2 management information base for the transmission control protocol using SMIv2, November

1996. URL ftp://ftp.internic.net/rfc/rfc1213.txt,   ftp://
ftp.internic.net/rfc/rfc2012.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1213.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2012.

txt. Updates RFC1213 [1203]. Status: PROPOSED STANDARD.

RFC2013

[2000] K. McCloghrie. RFC 2013: SNMPv2 management information base for the user datagram protocol using SMIv2, November 1996. URL ftp://

ftp.internic.net/rfc/rfc1213.txt,   ftp://ftp.internic.net/
rfc/rfc2013.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1213.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2013.txt. Updates

RFC1213 [1203]. Status: PROPOSED STANDARD.

RFC2014

[2001] A. Weinrib and J. Postel. RFC 2014: IRTF research group guidelines and procedures, October 1996. URL ftp://ftp.internic.net/rfc/

bcp8.txt,   ftp://ftp.internic.net/rfc/rfc2014.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/bcp8.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2014.txt. See also BCP0008 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2015

[2002] M. Elkins. RFC 2015: MIME security with pretty good privacy (PGP),

October 1996. URL ftp://ftp.internic.net/rfc/rfc2015.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2015.txt. Status: PRO-

POSED STANDARD.

RFC2016

[2003] L. Daigle, P. Deutsch, B. Heelan, C. Alpaugh, and M. Maclachlan. RFC 2016: Uniform Resource Agents (URAs), October

1996. URL ftp://ftp.internic.net/rfc/rfc2016.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2016.txt. Status: EXPERIMEN-

TAL.

RFC2017

[2004] N. Freed, K. Moore, and A. Cargille. RFC 2017: Definition of the URL MIME external-body access-type, October 1996. URL ftp://ftp.

internic.net/rfc/rfc2017.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2017.txt. Status: PROPOSED STANDARD.

RFC2018

[2005] M. Mathis, J. Mahdavi, S. Floyd, and A. Romanow. RFC 2018: TCP selective acknowledgment options, October 1996. URL ftp://ftp.

internic.net/rfc/rfc2018.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2018.txt. Status: PROPOSED STANDARD.

RFC2019

[2006] M. Crawford. RFC 2019: Transmission of IPv6 packets over FDDI, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2019.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc2019.txt. Obsoleted by RFC2467

[2454]. Obsoletes RFC2184 [2171]. Status: PROPOSED STANDARD.

RFC2020

[2007] J. Flick. RFC 2020: Definitions of managed objects for IEEE 802.12 in-

terfaces, October 1996. URL ftp://ftp.internic.net/rfc/rfc2020.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2020.txt. Status:

PROPOSED STANDARD.

RFC2021

[2008] S. Waldbusser. RFC 2021: Remote network monitoring management information base version 2 using SMIv2, January 1997. URL ftp://ftp.

internic.net/rfc/rfc2021.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2021.txt. Status: PROPOSED STANDARD.

RFC2022

[2009] G. Armitage. RFC 2022: Support for multicast over UNI 3.0/3.1 based ATM networks, November 1996. URL ftp://ftp.internic.net/

rfc/rfc2022.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2022.

txt. Status: PROPOSED STANDARD.

RFC2023

[2010] D. Haskin and E. Allen. RFC 2023: IP version 6 over PPP, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2023.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2023.txt. Status: PROPOSED

STANDARD.

RFC2024

[2011] D. Chen, P. Gayek, and S. Nix. RFC 2024: Definitions of managed objects for data link switching using SMIv2, October

1996. URL ftp://ftp.internic.net/rfc/rfc2024.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2024.txt. Status: PROPOSED

STANDARD.

RFC2025

[2012] C. Adams. RFC 2025: The simple public-key GSS-API mechanism

(SPKM), October 1996. URL ftp://ftp.internic.net/rfc/rfc2025.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2025.txt. Status:

PROPOSED STANDARD.

RFC2026

[2013] S. Bradner. RFC 2026: The Internet standards process — revision 3, Oc-

tober 1996. URL ftp://ftp.internic.net/rfc/bcp9.txt,   ftp://
ftp.internic.net/rfc/rfc1602.txt,   ftp://ftp.internic.net/

rfc/rfc2026.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp9.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1602.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2026.txt. See also BCP0009 [?]. Obsoletes

RFC1602 [1592]. Status: BEST CURRENT PRACTICE.

RFC2027

[2014] J. Galvin. RFC 2027: IAB and IESG selection, confirmation, and recall process: Operation of the nominating and recall committees, October

1996. URL ftp://ftp.internic.net/rfc/bcp10.txt,   ftp://ftp.
internic.net/rfc/rfc2027.txt,   ftp://ftp.internic.net/rfc/
rfc2282.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp10.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2027.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc2282.txt. Obsoleted by BCP0010 [?], RFC2282 [2269]. See also BCP0010 [?]. Status: BEST CURRENT PRACTICE.

RFC2028

[2015] R. Hovey and S. Bradner. RFC 2028: The organizations involved in the IETF standards process, October 1996. URL ftp:

//ftp.internic.net/rfc/bcp11.txt,   ftp://ftp.internic.net/
rfc/rfc2028.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp11.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2028.txt. See also BCP0011

[?]. Status: BEST CURRENT PRACTICE.

RFC2029

[2016] M. Speer and D. Hoffman. RFC 2029: RTP payload format of Sun’s CellB video encoding, October 1996. URL ftp://ftp.internic.net/

rfc/rfc2029.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2029.

txt. Status: PROPOSED STANDARD.

RFC2030

[2017] D. Mills. RFC 2030: Simple network time protocol (SNTP) version 4 for IPv4, IPv6 and OSI, October 1996. URL ftp://

ftp.internic.net/rfc/rfc1769.txt,   ftp://ftp.internic.net/
rfc/rfc2030.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1769.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2030.txt. Obsoletes

RFC1769 [1759]. Status: INFORMATIONAL.

RFC2031

[2018] E. Huizer. RFC 2031: IETF-ISOC relationship, October 1996. URL ftp:

//ftp.internic.net/rfc/rfc2031.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2031.txt. Status: INFORMATIONAL.

RFC2032

[2019] T. Turletti and C. Huitema. RFC 2032: RTP payload format for H.261 video streams, October 1996. URL ftp://ftp.internic.net/

rfc/rfc2032.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2032.

txt. Status: PROPOSED STANDARD.

RFC2033

[2020] J. Myers. RFC 2033: Local mail transfer protocol, October

1996. URL ftp://ftp.internic.net/rfc/rfc2033.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2033.txt. Status: INFORMA-

TIONAL.

RFC2034

[2021] N. Freed. RFC 2034: SMTP service extension for returning enhanced er-

ror codes, October 1996. URL ftp://ftp.internic.net/rfc/rfc2034.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2034.txt. Status:

PROPOSED STANDARD.

RFC2035

[2022] L. Berc, W. Fenner, R. Frederick, and S. McCanne. RFC 2035: RTP payload format for JPEG-compressed video, October 1996. URL ftp://

ftp.internic.net/rfc/rfc2035.txt,   ftp://ftp.internic.net/
rfc/rfc2435.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2035.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2435.txt. Obsoleted

by RFC2435 [2422]. Status: PROPOSED STANDARD.

RFC2036

[2023] G. Huston. RFC 2036: Observations on the use of components of the Class A address space within the Internet, October

1996. URL ftp://ftp.internic.net/rfc/rfc2036.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2036.txt. Status: INFORMA-

TIONAL.

RFC2037

[2024] K. McCloghrie and A. Bierman. RFC 2037: Entity MIB using SMIv2,

October 1996. URL ftp://ftp.internic.net/rfc/rfc2037.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2037.txt. Status: PRO-

POSED STANDARD.

RFC2038

[2025] D. Hoffman, G. Fernando, and V. Goyal. RFC 2038: RTP payload format for MPEG1/MPEG2 video, October 1996. URL ftp://

ftp.internic.net/rfc/rfc2038.txt,   ftp://ftp.internic.net/

rfc/rfc2250.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2038.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2250.txt. Obsoleted

by RFC2250 [2237]. Status: PROPOSED STANDARD.

RFC2039

[2026] C. Kalbfleisch. RFC 2039: Applicability of standards track MIBs to management of World Wide Web servers, November 1996. URL ftp://ftp.

internic.net/rfc/rfc2039.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2039.txt. Status: INFORMATIONAL.

RFC2040

[2027] R. Baldwin and R. Rivest. RFC 2040: The RC5, RC5-CBC, RC5-CBC-pad, and RC5-CTS algorithms, October 1996. URL ftp://ftp.

internic.net/rfc/rfc2040.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2040.txt. Status: INFORMATIONAL.

RFC2041

[2028] B. Noble, G. Nguyen, M. Satyanarayanan, and R. Katz. RFC 2041: Mobile network tracing, December 1996. URL ftp://ftp.internic.net/

rfc/rfc2041.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2041.

txt. Status: INFORMATIONAL.

RFC2042

[2029] B. Manning. RFC 2042: Registering new BGP attribute types, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2042.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2042.txt. Status: INFORMA-

TIONAL.

RFC2043

[2030] A. Fuqua. RFC 2043: The PPP SNA control protocol (SNACP), Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2043.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2043.txt. Status: PROPOSED

STANDARD.

RFC2044

[2031] F. Yergeau. RFC 2044: UTF-8, a transformation format of Unicode

and ISO 10646, October 1996. URL ftp://ftp.internic.net/rfc/
rfc2044.txt,   ftp://ftp.internic.net/rfc/rfc2279.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2044.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2279.txt. Obsoleted by RFC2279 [2266]. Status: IN-FORMATIONAL.

RFC2045

[2032] N. Freed and N. Borenstein. RFC 2045: Multipurpose Internet Mail Extensions (MIME) part one: Format of Internet message bod-

ies, November 1996. URL ftp://ftp.internic.net/rfc/rfc1521.
txt,   ftp://ftp.internic.net/rfc/rfc1522.txt,   ftp://ftp.
internic.net/rfc/rfc1590.txt,   ftp://ftp.internic.net/rfc/
rfc2045.txt,   ftp://ftp.internic.net/rfc/rfc2184.txt,   ftp:
//ftp.internic.net/rfc/rfc2231.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1521.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1522.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1590.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2045.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2184.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2231.txt. Obsoletes RFC1521, RFC1522, RFC1590 [1511, 1512, 1580]. Updated by RFC2184, RFC2231 [2171, 2218]. Status: DRAFT STANDARD.

RFC2046

[2033] N. Freed and N. Borenstein. RFC 2046: Multipurpose Internet Mail Extensions (MIME) part two: Media types, November

1996. URL ftp://ftp.internic.net/rfc/rfc1521.txt,   ftp://
ftp.internic.net/rfc/rfc1522.txt,   ftp://ftp.internic.net/
rfc/rfc1590.txt,   ftp://ftp.internic.net/rfc/rfc2046.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1521.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1522.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1590.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2046.

txt. Obsoletes RFC1521, RFC1522, RFC1590 [1511, 1512, 1580]. Status: DRAFT STANDARD.

RFC2047

[2034] K. Moore. RFC 2047: MIME (Multipurpose Internet Mail Extensions) part three: Message header extensions for non-ASCII

text, November 1996. URL ftp://ftp.internic.net/rfc/rfc1521.
txt,   ftp://ftp.internic.net/rfc/rfc1522.txt,   ftp://ftp.
internic.net/rfc/rfc1590.txt,   ftp://ftp.internic.net/rfc/
rfc2047.txt,   ftp://ftp.internic.net/rfc/rfc2184.txt,   ftp:
//ftp.internic.net/rfc/rfc2231.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1521.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1522.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1590.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2047.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2184.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2231.txt. Obsoletes RFC1521, RFC1522, RFC1590 [1511, 1512, 1580]. Updated by RFC2184, RFC2231 [2171, 2218]. Status: DRAFT STANDARD.

RFC2048

[2035] N. Freed, J. Klensin, and J. Postel. RFC 2048: Multipurpose Internet Mail Extensions (MIME) part four: Registration proce-

dures, November 1996. URL ftp://ftp.internic.net/rfc/bcp13.
txt,   ftp://ftp.internic.net/rfc/rfc1521.txt,   ftp://ftp.
internic.net/rfc/rfc1522.txt,   ftp://ftp.internic.net/rfc/
rfc1590.txt,   ftp://ftp.internic.net/rfc/rfc2048.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/bcp13.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1521.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1522.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1590.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2048.txt. See also BCP0013

[?]. Obsoletes RFC1521, RFC1522, RFC1590 [1511, 1512, 1580]. Status: BEST CURRENT PRACTICE.

RFC2049

[2036] N. Freed and N. Borenstein. RFC 2049: Multipurpose Internet Mail Extension (MIME) part five: Conformance criteria and exam-

ples, November 1996. URL ftp://ftp.internic.net/rfc/rfc1521.
txt,   ftp://ftp.internic.net/rfc/rfc1522.txt,   ftp://ftp.
internic.net/rfc/rfc1590.txt,   ftp://ftp.internic.net/rfc/
rfc2049.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1521.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1522.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1590.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2049.txt. Obsoletes RFC1521, RFC1522, RFC1590 [1511, 1512, 1580]. Status: DRAFT STANDARD.

RFC2050

[2037] K. Hubbard, M. Kosters, D. Conrad, D. Karrenberg, and J. Postel. RFC 2050: INTERNET REGISTRY IP ALLOCATION GUIDE-

LINES, November 1996. URL ftp://ftp.internic.net/rfc/bcp12.
txt,   ftp://ftp.internic.net/rfc/rfc2050.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp12.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2050.txt. See also BCP0012 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2051

[2038] M. Allen, B. Clouston, Z. Kielczewski, W. Kwan, and B. Moore. RFC 2051: Definitions of managed objects for APPC using SMIv2, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2051.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2051.txt. Status: PROPOSED

STANDARD.

RFC2052

[2039] A. Gulbrandsen and P. Vixie. RFC 2052: A DNS RR for specifying the location of services (DNS SRV), October 1996. URL ftp://ftp.

internic.net/rfc/rfc2052.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2052.txt. Status: EXPERIMENTAL.

RFC2053

[2040] E. Der-Danieliantz. RFC 2053: The AM (armenia) domain, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2053.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2053.txt. Status: INFORMA-

TIONAL.

RFC2054

[2041] B. Callaghan. RFC 2054: WebNFS client specification, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2054.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2054.txt. Status: INFORMA-

TIONAL.

RFC2055

[2042] B. Callaghan. RFC 2055: WebNFS server specification, Octo-

ber 1996. URL ftp://ftp.internic.net/rfc/rfc2055.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2055.txt. Status: INFORMA-

TIONAL.

RFC2056

[2043] R. Denenberg, J. Kunze, and D. Lynch. RFC 2056: Uniform resource locators for Z39.50, November 1996. URL ftp://ftp.internic.net/

rfc/rfc2056.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2056.

txt. Status: PROPOSED STANDARD.

RFC2057

[2044] S. Bradner. RFC 2057: Source Directed Access Control on the Internet,

November 1996. URL ftp://ftp.internic.net/rfc/rfc2057.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2057.txt. Status: INFOR-

MATIONAL.

RFC2058

[2045] C. Rigney, A. Rubens, W. Simpson, and S. Willens. RFC 2058: Remote authentication dial in user service (RADIUS), January

1997. URL ftp://ftp.internic.net/rfc/rfc2058.txt,   ftp://
ftp.internic.net/rfc/rfc2138.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2058.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2138.

txt. Obsoleted by RFC2138 [2125]. Status: PROPOSED STANDARD.

RFC2059

[2046] C. Rigney. RFC 2059: RADIUS accounting, January 1997. URL ftp://

ftp.internic.net/rfc/rfc2059.txt,   ftp://ftp.internic.net/
rfc/rfc2139.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2059.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2139.txt. Obsoleted

by RFC2139 [2126]. Status: INFORMATIONAL.

RFC2060

[2047] M. Crispin. RFC 2060: INTERNET MESSAGE ACCESS PRO-TOCOL — VERSION 4rev1, December 1996. URL ftp://ftp.

internic.net/rfc/rfc1730.txt,   ftp://ftp.internic.net/rfc/
rfc2060.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1730.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2060.txt. Obsoletes RFC1730

[1720]. Status: PROPOSED STANDARD.

RFC2061

[2048] M. Crispin. RFC 2061: IMAP4 COMPATIBILITY WITH IMAP2BIS,

December 1996. URL ftp://ftp.internic.net/rfc/rfc1730.txt,
ftp://ftp.internic.net/rfc/rfc2061.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1730.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2061.txt. Obsoletes RFC1730 [1720]. Status: INFORMATIONAL.

RFC2062

[2049] M. Crispin. RFC 2062: Internet Message Access Protocol — obsolete

syntax, December 1996. URL ftp://ftp.internic.net/rfc/rfc2062.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2062.txt. Status:

PROPOSED STANDARD.

RFC2063

[2050] N. Brownlee, C. Mills, and G. Ruth. RFC 2063: Traffic flow measurement: Architecture, January 1997. URL ftp://ftp.internic.net/

rfc/rfc2063.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2063.

txt. Status: EXPERIMENTAL.

RFC2064

[2051] N. Brownlee. RFC 2064: Traffic flow measurement: Meter MIB, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2064.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2064.txt. Status: EXPERIMEN-

TAL.

RFC2065

[2052] D. Eastlake, 3rd and C. Kaufman. RFC 2065: Domain name system security extensions, January 1997. URL ftp://ftp.internic.net/rfc/

rfc1034.txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp:
//ftp.internic.net/rfc/rfc2065.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1034.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2065.txt. Updates

RFC1034, RFC1035 [1024, 1025]. Status: PROPOSED STANDARD.

RFC2066

[2053] R. Gellens. RFC 2066: TELNET CHARSET option, January

1997. URL ftp://ftp.internic.net/rfc/rfc2066.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2066.txt. Status: EXPERIMEN-

TAL.

RFC2067

[2054] J. Renwick. RFC 2067: IP over HIPPI, January 1997. URL ftp://ftp.

internic.net/rfc/rfc2067.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2067.txt. Status: DRAFT STANDARD.

RFC2068

[2055] R. Fielding, J. Gettys, J. Mogul, H. Frystyk, and T. Berners-Lee. RFC 2068: Hypertext Transfer Protocol — HTTP/1.1, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2068.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2068.txt. Status: PROPOSED

STANDARD.

RFC2069

[2056] J. Franks, P. Hallam-Baker, J. Hostetler, P. Leach, A. Luotonen, E. Sink, and L. Stewart. RFC 2069: An extension to HTTP: Digest access authen-

tication, January 1997. URL ftp://ftp.internic.net/rfc/rfc2069.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2069.txt. Status:

PROPOSED STANDARD.

RFC2070

[2057] F. Yergeau, G. Nicol, G. Adams, and M. Duerst. RFC 2070: Internationalization of the hypertext markup language, January

1997. URL ftp://ftp.internic.net/rfc/rfc2070.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2070.txt. Status: PROPOSED

STANDARD.

RFC2071

[2058] P. Ferguson and H. Berkowitz. RFC 2071: Network renumbering overview: Why would I want it and what is it anyway?, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2071.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2071.txt. Status: INFORMA-

TIONAL.

RFC2072

[2059] H. Berkowitz. RFC 2072: Router renumbering guide, January

1997. URL ftp://ftp.internic.net/rfc/rfc2072.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2072.txt. Status: INFORMA-

TIONAL.

RFC2073

[2060] Y. Rekhter, P. Lothberg, R. Hinden, S. Deering, and J. Postel. RFC 2073: An IPv6 provider-based unicast address for-

mat, January 1997. URL ftp://ftp.internic.net/rfc/rfc2073.
txt,   ftp://ftp.internic.net/rfc/rfc2374.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2073.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2374.txt. Obsoleted by RFC2374 [2361]. Status: PRO-POSED STANDARD.

RFC2074

[2061] A. Bierman and R. Iddon. RFC 2074: Remote network monitoring MIB protocol identifiers, January 1997. URL ftp://ftp.internic.net/

rfc/rfc2074.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2074.

txt. Status: PROPOSED STANDARD.

RFC2075

[2062] C. Partridge. RFC 2075: IP echo host service, January 1997. URL ftp:/

/ftp.internic.net/rfc/rfc2075.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2075.txt. Status: EXPERIMENTAL.

RFC2076

[2063] J. Palme. RFC 2076: Common Internet message headers, Febru-

ary 1997. URL ftp://ftp.internic.net/rfc/rfc2076.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2076.txt. Status: INFORMA-

TIONAL.

RFC2077

[2064] S. Nelson, C. Parks, and Mitra. RFC 2077: The model primary content type for Multipurpose Internet Mail Extensions, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2077.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2077.txt. Status: PROPOSED

STANDARD.

RFC2078

[2065] J. Linn. RFC 2078: Generic security service application program interface, version 2, January 1997. URL ftp://ftp.internic.net/rfc/

rfc1508.txt,   ftp://ftp.internic.net/rfc/rfc2078.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc1508.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2078.txt. Obsoletes RFC1508 [1498]. Status: PRO-POSED STANDARD.

RFC2079

[2066] M. Smith. RFC 2079: Definition of an X.500 attribute type and an object class to hold Uniform Resource Identifiers (URIs), Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2079.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2079.txt. Status: PROPOSED

STANDARD.

RFC2080

[2067] G. Malkin and R. Minnear. RFC 2080: RIPng for IPv6, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2080.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2080.txt. Status: PROPOSED

STANDARD.

RFC2081

[2068] G. Malkin. RFC 2081: RIPng protocol applicability statement, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2081.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2081.txt. Status: INFORMA-

TIONAL.

RFC2082

[2069] F. Baker and R. Atkinson. RFC 2082: RIP-2 MD5 authentication, Jan-

uary 1997. URL ftp://ftp.internic.net/rfc/rfc2082.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2082.txt. Status: PROPOSED

STANDARD.

RFC2083

[2070] T. Boutell. RFC 2083: PNG (portable network graphics) specification version 1.0, January 1997. URL ftp://ftp.internic.net/

rfc/rfc2083.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2083.

txt. Status: INFORMATIONAL.

RFC2084

[2071] G. Bossert, S. Cooper, and W. Drummond. RFC 2084: Considerations for Web transaction security, January 1997. URL ftp://ftp.internic.

net/rfc/rfc2084.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2084.

txt. Status: INFORMATIONAL.

RFC2085

[2072] M. Oehler and R. Glenn. RFC 2085: HMAC-MD5 IP authentication with replay prevention, February 1997. URL ftp://ftp.internic.net/

rfc/rfc2085.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2085.

txt. Status: PROPOSED STANDARD.

RFC2086

[2073] J. Myers. RFC 2086: IMAP4 ACL extension, January 1997. URL ftp://

ftp.internic.net/rfc/rfc2086.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2086.txt. Status: PROPOSED STANDARD.

RFC2087

[2074] J. Myers. RFC 2087: IMAP4 QUOTA extension, January 1997. URL

ftp://ftp.internic.net/rfc/rfc2087.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2087.txt. Status: PROPOSED STANDARD.

RFC2088

[2075] J. Myers. RFC 2088: IMAP4 non-synchronizing literals, January

1997. URL ftp://ftp.internic.net/rfc/rfc2088.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2088.txt. Status: PROPOSED

STANDARD.

RFC2089

[2076] B. Wijnen and D. Levi. RFC 2089: V2ToV1 mapping SN-MPv2 onto SNMPv1 within a bi-lingual SNMP agent, January

1997. URL ftp://ftp.internic.net/rfc/rfc2089.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2089.txt. Status: INFORMA-

TIONAL.

RFC2090

[2077] A. Emberson. RFC 2090: TFTP multicast option, February

1997. URL ftp://ftp.internic.net/rfc/rfc2090.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2090.txt. Status: EXPERIMEN-

TAL.

RFC2091

[2078] G. Meyer and S. Sherry. RFC 2091: Triggered extensions to RIP to support demand circuits, January 1997. URL ftp://ftp.internic.net/

rfc/rfc2091.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2091.

txt. Status: PROPOSED STANDARD.

RFC2092

[2079] S. Sherry and G. Meyer. RFC 2092: Protocol analysis for triggered

RIP, January 1997. URL ftp://ftp.internic.net/rfc/rfc2092.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2092.txt. Status: INFOR-

MATIONAL.

RFC2093

[2080] H. Harney and C. Muckenhirn. RFC 2093: Group key management protocol (GKMP) specification, July 1997. URL ftp://ftp.internic.

net/rfc/rfc2093.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2093.

txt. Status: EXPERIMENTAL.

RFC2094

[2081] H. Harney and C. Muckenhirn. RFC 2094: Group key management protocol (GKMP) architecture, July 1997. URL ftp://ftp.internic.net/

rfc/rfc2094.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2094.

txt. Status: EXPERIMENTAL.

RFC2095

[2082] J. Klensin, R. Catoe, and P. Krumviede. RFC 2095: IMAP/ POP AUTHorize extension for simple challenge/response, January

1997. URL ftp://ftp.internic.net/rfc/rfc2095.txt,   ftp://
ftp.internic.net/rfc/rfc2195.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2095.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2195.

txt. Obsoleted by RFC2195 [2182]. Status: PROPOSED STANDARD.

RFC2096

[2083] F. Baker. RFC 2096: IP forwarding table MIB, January 1997. URL ftp:

//ftp.internic.net/rfc/rfc1354.txt,   ftp://ftp.internic.net/
rfc/rfc2096.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1354.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2096.txt. Obsoletes

RFC1354 [1344]. Status: PROPOSED STANDARD.

RFC2097

[2084] G. Pall. RFC 2097: The PPP netBIOS frames control protocol (NBFCP),

January 1997. URL ftp://ftp.internic.net/rfc/rfc2097.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2097.txt. Status: PRO-

POSED STANDARD.

RFC2098

[2085] Y. Katsube, K. Nagami, and H. Esaki. RFC 2098: Toshiba’s router architecture extensions for ATM: Overview, February 1997. URL ftp://ftp.

internic.net/rfc/rfc2098.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2098.txt. Status: INFORMATIONAL.

RFC2099

[2086] J. Elliott. RFC 2099: Request for comments summary RFC numbers

2000–2099, March 1997. URL ftp://ftp.internic.net/rfc/rfc2099.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2099.txt. Status:

INFORMATIONAL.

RFC2100

[2087] J. Ashworth. RFC 2100: The naming of hosts, April 1997. URL ftp://

ftp.internic.net/rfc/rfc2100.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2100.txt. Status: INFORMATIONAL.

RFC2101

[2088] B. Carpenter, J. Crowcroft, and Y. Rekhter. RFC 2101: IPv4 address behaviour today, February 1997. URL ftp://ftp.internic.net/

rfc/rfc2101.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2101.

txt. Status: INFORMATIONAL.

RFC2102

[2089] R. Ramanathan. RFC 2102: Multicast support for Nimrod: Requirements and solution approaches, February 1997. URL ftp://ftp.

internic.net/rfc/rfc2102.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2102.txt. Status: INFORMATIONAL.

RFC2103

[2090] R. Ramanathan. RFC 2103: Mobility support for Nimrod: Challenges and solution approaches, February 1997. URL ftp://ftp.internic.

net/rfc/rfc2103.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2103.

txt. Status: INFORMATIONAL.

RFC2104

[2091] H. Krawczyk, M. Bellare, and R. Canetti. RFC 2104: HMAC: Keyedhashing for message authentication, February 1997. URL ftp://ftp.

internic.net/rfc/rfc2104.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2104.txt. Status: INFORMATIONAL.

RFC2105

[2092] Y. Rekhter, B. Davie, D. Katz, E. Rosen, and G. Swallow. RFC 2105: Cisco Systems’ tag switching architecture overview, Febru-

ary 1997. URL ftp://ftp.internic.net/rfc/rfc2105.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2105.txt. Status: INFORMA-

TIONAL.

RFC2106

[2093] S. Chiang, J. Lee, and H. Yasuda. RFC 2106: Data link switching remote access protocol, February 1997. URL ftp://ftp.

internic.net/rfc/rfc2106.txt,   ftp://ftp.internic.net/rfc/
rfc2114.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2106.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc2114.txt. Obsoleted by

RFC2114 [2101]. Status: INFORMATIONAL.

RFC2107

[2094] K. Hamzeh. RFC 2107: Ascend tunnel management protocol — ATMP,

February 1997. URL ftp://ftp.internic.net/rfc/rfc2107.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2107.txt. Status: INFOR-

MATIONAL.

RFC2108

[2095] K. de Graaf, D. Romascanu, D. McMaster, and K. McCloghrie. RFC 2108: Definitions of managed objects for IEEE 802.3 repeater devices using SMIv2, February 1997. URL ftp://ftp.

internic.net/rfc/rfc1516.txt,   ftp://ftp.internic.net/rfc/
rfc2108.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1516.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2108.txt. Obsoletes RFC1516

[1506]. Status: PROPOSED STANDARD.

RFC2109

[2096] D. Kristol and L. Montulli. RFC 2109: HTTP state management mech-

anism, February 1997. URL ftp://ftp.internic.net/rfc/rfc1920.
txt,   ftp://ftp.internic.net/rfc/rfc2109.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1920.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2109.txt. Obsoletes RFC1920 [1907]. Status: PROPOSED STANDARD.

RFC2110

[2097] J. Palme and A. Hopmann. RFC 2110: MIME E-mail encapsulation of aggregate documents, such as HTML (MHTML), March

1997. URL ftp://ftp.internic.net/rfc/rfc2110.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2110.txt. Status: PROPOSED

STANDARD.

RFC2111

[2098] E. Levinson. RFC 2111: Content-ID and message-ID uniform resource

locators, February 1997. URL ftp://ftp.internic.net/rfc/rfc2111.
txt,   ftp://ftp.internic.net/rfc/rfc2392.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2111.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2392.txt. Obsoleted by RFC2392 [2379]. Status: PRO-POSED STANDARD.

RFC2112

[2099] E. Levinson. RFC 2112: The MIME multipart/related content-

type, February 1997. URL ftp://ftp.internic.net/rfc/rfc1872.

txt,   ftp://ftp.internic.net/rfc/rfc2112.txt,   ftp://ftp.
internic.net/rfc/rfc2387.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1872.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2112.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2387.txt. Obsoleted

by RFC2387 [2374]. Obsoletes RFC1872 [1859]. Status: PROPOSED STANDARD.

RFC2113

[2100] D. Katz. RFC 2113: IP router alert option, February 1997. URL ftp://

ftp.internic.net/rfc/rfc2113.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2113.txt. Status: PROPOSED STANDARD.

RFC2114

[2101] S. Chiang, J. Lee, and H. Yasuda. RFC 2114: Data link switching client access protocol, February 1997. URL ftp://

ftp.internic.net/rfc/rfc2106.txt,   ftp://ftp.internic.net/
rfc/rfc2114.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2106.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2114.txt. Obsoletes

RFC2106 [2093]. Status: INFORMATIONAL.

RFC2115

[2102] C. Brown and F. Baker. RFC 2115: Management information base for frame relay DTEs using SMIv2, September 1997. URL ftp://

ftp.internic.net/rfc/rfc1315.txt,   ftp://ftp.internic.net/
rfc/rfc2115.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1315.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2115.txt. Obsoletes

RFC1315 [1305]. Status: DRAFT STANDARD.

RFC2116

[2103] C. Apple and K. Rossen. RFC 2116: X.500 implementations

catalog-96, April 1997. URL ftp://ftp.internic.net/rfc/fyi11.
txt,   ftp://ftp.internic.net/rfc/rfc1632.txt,   ftp://ftp.
internic.net/rfc/rfc2116.txt,   ftp://ftp.math.utah.edu/pub/
rfc/fyi11.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1632.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2116.txt. See also FYI0011

[?]. Obsoletes RFC1632 [1622]. Status: INFORMATIONAL.

RFC2117

[2104] D. Estrin, D. Farinacci, A. Helmy, D. Thaler, S. Deering, M. Handley, V. Jacobson, C. Liu, P. Sharma, and L. Wei. RFC 2117: Protocol independent multicast-sparse mode (PIM-SM): Protocol speci-

fication, June 1997. URL ftp://ftp.internic.net/rfc/rfc2117.
txt,   ftp://ftp.internic.net/rfc/rfc2362.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc2117.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2362.txt. Obsoleted by RFC2362 [2349]. Status: EXPER-IMENTAL.

RFC2118

[2105] G. Pall. RFC 2118: Microsoft point-to-point compression (MPPC) pro-

tocol, March 1997. URL ftp://ftp.internic.net/rfc/rfc2118.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2118.txt. Status: INFOR-

MATIONAL.

RFC2119

[2106] S. Bradner. RFC 2119: Key words for use in RFCs to indicate require-

ment levels, March 1997. URL ftp://ftp.internic.net/rfc/bcp14.
txt,   ftp://ftp.internic.net/rfc/rfc2119.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp14.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2119.txt. See also BCP0014 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2120

[2107] D. Chadwick. RFC 2120: Managing the X.500 root naming con-

text, March 1997. URL ftp://ftp.internic.net/rfc/rfc2120.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2120.txt. Status: EXPER-

IMENTAL.

RFC2121

[2108] G. Armitage. RFC 2121: Issues affecting MARS cluster size,

March 1997. URL ftp://ftp.internic.net/rfc/rfc2121.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2121.txt. Status: INFOR-

MATIONAL.

RFC2122

[2109] D. Mavrakis, H. Layec, and K. Kartmann. RFC 2122: VEMMI URL specification, March 1997. URL ftp://ftp.internic.net/

rfc/rfc2122.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2122.

txt. Status: PROPOSED STANDARD.

RFC2123

[2110] N. Brownlee. RFC 2123: Traffic flow measurement: Experiences with

NeTraMet, March 1997. URL ftp://ftp.internic.net/rfc/rfc2123.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2123.txt. Status:

INFORMATIONAL.

RFC2124

[2111] P. Amsden, J. Amweg, P. Calato, S. Bensley, and G. Lyons. RFC 2124: Cabletron’s Light-weight Flow Admission Protocol Specification Version

1.0, March 1997. URL ftp://ftp.internic.net/rfc/rfc2124.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2124.txt. Status: INFOR-

MATIONAL.

RFC2125

[2112] C. Richards and K. Smith. RFC 2125: The PPP Bandwidth Allocation Protocol (BAP) / the PPP Bandwidth Allocation Control Protocol

(BACP), March 1997. URL ftp://ftp.internic.net/rfc/rfc2125.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2125.txt. Status:

PROPOSED STANDARD.

RFC2126

[2113] Y. Pouffary and A. Young. RFC 2126: ISO transport service on top of TCP (ITOT), March 1997. URL ftp://ftp.internic.net/

rfc/rfc2126.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2126.

txt. Status: PROPOSED STANDARD. Updates RFC1006 [996].

RFC2127

[2114] G. Roeck. RFC 2127: ISDN management information base using SMIv2,

March 1997. URL ftp://ftp.internic.net/rfc/rfc2127.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2127.txt. Status: PRO-

POSED STANDARD.

RFC2128

[2115] G. Roeck. RFC 2128: Dial control management information base using

SMIv2, March 1997. URL ftp://ftp.internic.net/rfc/rfc2128.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2128.txt. Status:

PROPOSED STANDARD.

RFC2129

[2116] K. Nagami, Y. Katsube, Y. Shobatake, A. Mogi, S. Matsuzawa, T. Jinmei, and H. Esaki. RFC 2129: Toshiba’s flow attribute notification protocol (FANP) specification, April 1997. URL ftp://ftp.internic.net/

rfc/rfc2129.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2129.

txt. Status: INFORMATIONAL.

RFC2130

[2117] C. Weider, C. Preston, K. Simonsen, H. Alvestrand, R. Atkinson, M. Crispin, and P. Svanberg. RFC 2130: The report of the IAB Character Set Workshop held 29 February — 1 March, 1996, April

1997. URL ftp://ftp.internic.net/rfc/rfc2130.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2130.txt. Status: INFORMA-

TIONAL.

RFC2131

[2118] R. Droms. RFC 2131: Dynamic host configuration protocol, March

1997. URL ftp://ftp.internic.net/rfc/rfc1541.txt,   ftp://
ftp.internic.net/rfc/rfc2131.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1541.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2131.

txt. Obsoletes RFC1541 [1531]. Status: DRAFT STANDARD.

RFC2132

[2119] S. Alexander and R. Droms. RFC 2132: DHCP options and BOOTP vendor extensions, March 1997. URL ftp://ftp.internic.net/rfc/

rfc1533.txt,   ftp://ftp.internic.net/rfc/rfc2132.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1533.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2132.txt. Obsoletes RFC1533 [1523]. Status: DRAFT STANDARD.

RFC2133

[2120] R. Gilligan, S. Thomson, J. Bound, and W. Stevens. RFC 2133: Basic socket interface extensions for IPv6, April 1997. URL ftp://ftp.

internic.net/rfc/rfc2133.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2133.txt. Status: INFORMATIONAL.

RFC2134

[2121] ISOC Board of Trustees. RFC 2134: Articles of incorporation of Internet

Society, April 1997. URL ftp://ftp.internic.net/rfc/rfc2134.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2134.txt. Status: INFOR-

MATIONAL.

RFC2135

[2122] ISOC Board of Trustees. RFC 2135: Internet Society by-laws, April

1997. URL ftp://ftp.internic.net/rfc/rfc2135.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2135.txt. Status: INFORMA-

TIONAL.

RFC2136

[2123] P. Vixie, Editor, S. Thomson, Y. Rekhter, and J. Bound. RFC 2136: Dynamic updates in the domain name system (DNS UP-

DATE), April 1997. URL ftp://ftp.internic.net/rfc/rfc1035.
txt,   ftp://ftp.internic.net/rfc/rfc2136.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2136.txt. Updates RFC1035 [1025]. Status: PROPOSED STANDARD.

RFC2137

[2124] D. Eastlake. RFC 2137: Secure domain name system dynamic

update, April 1997. URL ftp://ftp.internic.net/rfc/rfc1035.
txt,   ftp://ftp.internic.net/rfc/rfc2137.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2137.txt. Updates RFC1035 [1025]. Status: PROPOSED STANDARD.

RFC2138

[2125] C. Rigney, A. Rubens, W. Simpson, and S. Willens. RFC 2138: Remote authentication dial in user service (RADIUS), April 1997. URL ftp://

ftp.internic.net/rfc/rfc2058.txt,   ftp://ftp.internic.net/
rfc/rfc2138.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2058.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2138.txt. Obsoletes

RFC2058 [2045]. Status: PROPOSED STANDARD.

RFC2139

[2126] C. Rigney. RFC 2139: RADIUS accounting, April 1997. URL ftp://

ftp.internic.net/rfc/rfc2059.txt,   ftp://ftp.internic.net/
rfc/rfc2139.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2059.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2139.txt. Obsoletes

RFC2059 [2046]. Status: INFORMATIONAL.

RFC2140

[2127] J. Touch. RFC 2140: TCP control block interdependence, April

1997. URL ftp://ftp.internic.net/rfc/rfc2140.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2140.txt. Status: INFORMA-

TIONAL.

RFC2141

[2128] R. Moats. RFC 2141: URN syntax, May 1997. URL ftp://ftp.

internic.net/rfc/rfc2141.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2141.txt. Status: PROPOSED STANDARD.

RFC2142

[2129] D. Crocker. RFC 2142: Mailbox names for common services, roles and

functions, May 1997. URL ftp://ftp.internic.net/rfc/rfc2142.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2142.txt. Status:

PROPOSED STANDARD.

RFC2143

[2130] B. Elliston. RFC 2143: Encapsulating IP with the small computer system

interface, May 1997. URL ftp://ftp.internic.net/rfc/rfc2143.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2143.txt. Status:

EXPERIMENTAL.

RFC2144

[2131] C. Adams. RFC 2144: The CAST-128 encryption algorithm, May

21, 1997. URL ftp://ftp.internic.net/rfc/rfc2144.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2144.txt. Status: INFORMA-

TIONAL.

RFC2145

[2132] J. C. Mogul, R. Fielding, J. Gettys, and H. Frystyk. RFC 2145: Use and interpretation of HTTP version numbers, May

1997. URL ftp://ftp.internic.net/rfc/rfc2145.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2145.txt. Status: INFORMA-

TIONAL.

RFC2146

[2133] Federal Networking Council. RFC 2146: U.S. Government Internet do-

main names, May 1997. URL ftp://ftp.internic.net/rfc/rfc1816.
txt,   ftp://ftp.internic.net/rfc/rfc2146.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1816.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2146.txt. Obsoletes RFC1816 [1806]. Status: INFORMA-TIONAL.

RFC2147

[2134] D. Borman. RFC 2147: TCP and UDP over IPv6 jumbograms,

May 1997. URL ftp://ftp.internic.net/rfc/rfc2147.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2147.txt. Status: PROPOSED

STANDARD.

RFC2148

[2135] H. Alvestrand and P. Jurg. RFC 2148: Deployment of the Internet White Pages Service, September 1997. URL ftp://ftp.internic.net/rfc/

bcp15.txt,   ftp://ftp.internic.net/rfc/rfc2148.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/bcp15.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2148.txt. See also BCP0015 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2149

[2136] R. Talpade and M. Ammar. RFC 2149: Multicast server architectures for MARS-based ATM multicasting, May 1997. URL ftp://ftp.internic.

net/rfc/rfc2149.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2149.

txt. Status: INFORMATIONAL.

RFC2150

[2137] J. Max and W. Stickle. RFC 2150: Humanities and arts: Sharing center stage on the Internet, October 1997. URL ftp:

//ftp.internic.net/rfc/fyi31.txt,   ftp://ftp.internic.net/
rfc/rfc2150.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi31.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2150.txt. See also FYI0031

[?]. Status: INFORMATIONAL.

RFC2151

[2138] G. Kessler and S. Shepard. RFC 2151: A primer on Internet and TCP/ IP tools and utilities, June 1997. URL ftp://ftp.internic.net/rfc/

fyi30.txt,   ftp://ftp.internic.net/rfc/rfc1739.txt,   ftp:
//ftp.internic.net/rfc/rfc2151.txt,   ftp://ftp.math.utah.edu/
pub/rfc/fyi30.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1739.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2151.txt. See also

FYI0030 [?]. Obsoletes RFC1739 [1729]. Status: INFORMATIONAL.

RFC2152

[2139] D. Goldsmith and M. Davis. RFC 2152: UTF-7: A mailsafe transformation format of Unicode, May 1997. URL ftp://

ftp.internic.net/rfc/rfc1642.txt,   ftp://ftp.internic.net/
rfc/rfc2152.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1642.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2152.txt. Obsoletes

RFC1642 [1632]. Status: INFORMATIONAL.

RFC2153

[2140] W. Simpson. RFC 2153: PPP vendor extensions, May 1997. URL ftp://

ftp.internic.net/rfc/rfc1661.txt,   ftp://ftp.internic.net/
rfc/rfc1962.txt,   ftp://ftp.internic.net/rfc/rfc2153.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1661.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1962.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2153.txt. Updates RFC1661, RFC1962 [1651, 1949]. Status: INFORMATIONAL.

RFC2154

[2141] S. Murphy, M. Badger, and B. Wellington. RFC 2154: OSPF with digital

signatures, June 1997. URL ftp://ftp.internic.net/rfc/rfc2154.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2154.txt. Status:

EXPERIMENTAL.

RFC2155

[2142] B. Clouston and B. Moore. RFC 2155: Definitions of managed objects for APPN using SMIv2, June 1997. URL ftp://ftp.internic.net/

rfc/rfc2155.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2155.

txt. Obsoleted by RFC2455 [2442]. Status: PROPOSED STANDARD.

RFC2156

[2143] S. Kille. RFC 2156: MIXER (Mime Internet X.400 Enhanced Relay): Mapping between X.400 and RFC 822/MIME, January

1998. URL ftp://ftp.internic.net/rfc/rfc1026.txt,   ftp://
ftp.internic.net/rfc/rfc1138.txt,   ftp://ftp.internic.net/
rfc/rfc1148.txt,   ftp://ftp.internic.net/rfc/rfc1327.txt,
ftp://ftp.internic.net/rfc/rfc1495.txt,   ftp://ftp.internic.
net/rfc/rfc2156.txt,   ftp://ftp.internic.net/rfc/rfc822.txt,
ftp://ftp.internic.net/rfc/rfc987.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1026.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1138.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1148.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1327.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1495.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2156.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc822.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc987.txt. Obsoletes RFC0987,

RFC1026, RFC1138, RFC1148, RFC1327, RFC1495 [977, 1016, 1128, 1138, 1317, 1485]. Updates RFC0822 [812]. Status: PROPOSED STAN-DARD.

RFC2157

[2144] H. Alvestrand. RFC 2157: Mapping between X.400 and RFC-822/ MIME message bodies, January 1998. URL ftp://ftp.internic.net/

rfc/rfc2157.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2157.

txt. Status: PROPOSED STANDARD.

RFC2158

[2145] H. Alvestrand. RFC 2158: X.400 image body parts, January

1998. URL ftp://ftp.internic.net/rfc/rfc2158.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2158.txt. Status: PROPOSED

STANDARD.

RFC2159

[2146] H. Alvestrand. RFC 2159: A MIME body part for FAX, Jan-

uary 1998. URL ftp://ftp.internic.net/rfc/rfc2159.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2159.txt. Status: PROPOSED

STANDARD.

RFC2160

[2147] H. Alvestrand. RFC 2160: Carrying PostScript in X.400 and MIME, Jan-

uary 1998. URL ftp://ftp.internic.net/rfc/rfc2160.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc2160.txt. Status: PROPOSED

STANDARD.

RFC2161

[2148] H. Alvestrand. RFC 2161: A MIME body part for ODA, January

1998. URL ftp://ftp.internic.net/rfc/rfc2161.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2161.txt. Status: EXPERIMEN-

TAL.

RFC2162

[2149] C. Allocchio. RFC 2162: MaXIM-11 — mapping between X.400 /Internet mail and Mail-11 mail, January 1998. URL ftp://

ftp.internic.net/rfc/rfc1405.txt,   ftp://ftp.internic.net/
rfc/rfc2162.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1405.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2162.txt. Obsoletes

RFC1405 [1395]. Status: EXPERIMENTAL.

RFC2163

[2150] C. Allocchio. RFC 2163: Using the Internet DNS to distribute MIXER Conformant Global Address Mapping (MCGAM), January

1998. URL ftp://ftp.internic.net/rfc/rfc1664.txt,   ftp://
ftp.internic.net/rfc/rfc2163.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1664.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2163.

txt. Obsoletes RFC1664 [1654]. Status: PROPOSED STANDARD.

RFC2164

[2151] S. Kille. RFC 2164: Use of an X.500/LDAP directory to support MIXER address mapping, January 1998. URL ftp://ftp.

internic.net/rfc/rfc1838.txt,   ftp://ftp.internic.net/rfc/
rfc2164.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1838.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2164.txt. Obsoletes RFC1838

[1828]. Status: PROPOSED STANDARD.

RFC2165

[2152] J. Veizades, E. Guttman, C. Perkins, and S. Kaplan. RFC 2165: Service location protocol, June 1997. URL ftp://ftp.internic.net/

rfc/rfc2165.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2165.

txt. Status: PROPOSED STANDARD.

RFC2166

[2153] D. Bryant and P. Brittain. RFC 2166: APPN implementer’s workshop closed pages document DLSw v2.0 enhancements, June

1997. URL ftp://ftp.internic.net/rfc/rfc2166.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2166.txt. Status: INFORMA-

TIONAL.

RFC2167

[2154] S. Williamson, M. Kosters, D. Blacka, J. Singh, and K. Zeilstra. RFC 2167: Referral Whois (RWhois) Protocol V1.5, June 1997. URL ftp://

ftp.internic.net/rfc/rfc1714.txt,   ftp://ftp.internic.net/
rfc/rfc2167.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1714.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2167.txt. Obsoletes

RFC1714 [1704]. Status: INFORMATIONAL.

RFC2168

[2155] R. Daniel and M. Mealling. RFC 2168: Resolution of uniform resource identifiers using the Domain Name System, June

1997. URL ftp://ftp.internic.net/rfc/rfc2168.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2168.txt. Status: EXPERIMEN-

TAL.

RFC2169

[2156] R. Daniel. RFC 2169: A trivial convention for using HTTP in URN res-

olution, June 1997. URL ftp://ftp.internic.net/rfc/rfc2169.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2169.txt. Status: EXPER-

IMENTAL.

RFC2170

[2157] W. Almesberger, J. Le Boudec, and P. Oechslin. RFC 2170: Application REQuested IP over ATM (AREQUIPA), June 1997. URL ftp://ftp.

internic.net/rfc/rfc2170.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2170.txt. Status: INFORMATIONAL.

RFC2171

[2158] K. Murakami and M. Maruyama. RFC 2171: MAPOS — multiple access protocol over SONET/SDH version 1, June 1997. URL ftp://ftp.

internic.net/rfc/rfc2171.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2171.txt. Status: INFORMATIONAL.

RFC2172

[2159] M. Maruyama and K. Murakami. RFC 2172: MAPOS version 1 assigned

numbers, June 1997. URL ftp://ftp.internic.net/rfc/rfc2172.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2172.txt. Status:

INFORMATIONAL.

RFC2173

[2160] K. Murakami and M. Maruyama. RFC 2173: A MAPOS version 1 extension — node switch protocol, June 1997. URL ftp://ftp.internic.

net/rfc/rfc2173.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2173.

txt. Status: INFORMATIONAL.

RFC2174

[2161] K. Murakami and M. Maruyama. RFC 2174: A MAPOS version 1 extension — switch-switch protocol, June 1997. URL ftp://ftp.internic.

net/rfc/rfc2174.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2174.

txt. Status: INFORMATIONAL.

RFC2175

[2162] K. Murakami and M. Maruyama. RFC 2175: MAPOS 16 — multiple access protocol over SONET/SDH with 16 bit addressing,

June 1997. URL ftp://ftp.internic.net/rfc/rfc2175.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2175.txt. Status: INFORMA-

TIONAL.

RFC2176

[2163] K. Murakami and M. Maruyama. RFC 2176: IPv4 over MAPOS ver-

sion 1, June 1997. URL ftp://ftp.internic.net/rfc/rfc2176.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2176.txt. Status: INFOR-

MATIONAL.

RFC2177

[2164] B. Leiba. RFC 2177: IMAP4 IDLE command, June 1997. URL ftp://

ftp.internic.net/rfc/rfc2177.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2177.txt. Status: PROPOSED STANDARD.

RFC2178

[2165] J. Moy. RFC 2178: OSPF version 2, July 1997. URL ftp://ftp.

internic.net/rfc/rfc1583.txt,   ftp://ftp.internic.net/rfc/
rfc2178.txt,   ftp://ftp.internic.net/rfc/rfc2328.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1583.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2178.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc2328.txt. Obsoleted by RFC2328 [2315]. Obsoletes RFC1583 [1573]. Status: DRAFT STANDARD.

RFC2179

[2166] A. Gwinn. RFC 2179: Network security for trade shows, July

1997. URL ftp://ftp.internic.net/rfc/rfc2179.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2179.txt. Status: INFORMA-

TIONAL.

RFC2180

[2167] M. Gahrns. RFC 2180: IMAP4 multi-accessed mailbox practice,

July 1997. URL ftp://ftp.internic.net/rfc/rfc2180.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2180.txt. Status: INFORMA-

TIONAL.

RFC2181

[2168] R. Elz and R. Bush. RFC 2181: Clarifications to the DNS spec-

ification, July 1997. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.
internic.net/rfc/rfc1123.txt,   ftp://ftp.internic.net/rfc/
rfc2181.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1034.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1035.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1123.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2181.txt. Updates RFC1034, RFC1035, RFC1123 [1024, 1025, 1113]. Status: PROPOSED STANDARD.

RFC2182

[2169] R. Elz, R. Bush, S. Bradner, and M. Patton. RFC 2182: Selection and operation of secondary DNS servers, July 1997. URL ftp:

//ftp.internic.net/rfc/bcp16.txt,   ftp://ftp.internic.net/
rfc/rfc2182.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp16.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2182.txt. See also BCP0016

[?]. Status: BEST CURRENT PRACTICE.

RFC2183

[2170] R. Troost, S. Dorner, and K. Moore. RFC 2183: Communicating presentation information in Internet messages: The content-disposition header

field, August 1997. URL ftp://ftp.internic.net/rfc/rfc1806.
txt,   ftp://ftp.internic.net/rfc/rfc2183.txt,   ftp://ftp.
internic.net/rfc/rfc2184.txt,   ftp://ftp.internic.net/rfc/
rfc2231.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1806.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2183.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2184.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2231.txt. Updates RFC1806 [1796]. Updated by RFC2184, RFC2231 [2171, 2218]. Status: PROPOSED STANDARD.

RFC2184

[2171] N. Freed and K. Moore. RFC 2184: MIME parameter value and encoded word extensions: Character sets, languages, and continua-

tions, August 1997. URL ftp://ftp.internic.net/rfc/rfc2045.
txt,   ftp://ftp.internic.net/rfc/rfc2047.txt,   ftp://ftp.

internic.net/rfc/rfc2183.txt,   ftp://ftp.internic.net/rfc/
rfc2184.txt,   ftp://ftp.internic.net/rfc/rfc2231.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2045.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2047.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2183.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2184.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2231.txt. Obsoleted by

RFC2184, RFC2231 [2171, 2218]. Updates RFC2045, RFC2047, RFC2183 [2032, 2034, 2170]. Status: PROPOSED STANDARD.

RFC2185

[2172] R. Callon and D. Haskin. RFC 2185: Routing aspects of IPv6 transition,

September 1997. URL ftp://ftp.internic.net/rfc/rfc2185.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2185.txt. Status: INFOR-

MATIONAL.

RFC2186

[2173] D. Wessels and K. Claffy. RFC 2186: Internet Cache Protocol (ICP), ver-

sion 2, September 1997. URL ftp://ftp.internic.net/rfc/rfc2186.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2186.txt. Status:

INFORMATIONAL.

RFC2187

[2174] D. Wessels and K. Claffy. RFC 2187: Application of Internet Cache Protocol (ICP), version 2, September 1997. URL ftp://ftp.internic.

net/rfc/rfc2187.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2187.

txt. Status: INFORMATIONAL.

RFC2188

[2175] M. Banan, M. Taylor, and J. Cheng. RFC 2188: AT&T/Neda’s Efficient Short Remote Operations (ESRO) protocol specification version 1.2,

September 1997. URL ftp://ftp.internic.net/rfc/rfc2188.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2188.txt. Status: INFOR-

MATIONAL.

RFC2189

[2176] A. Ballardie. RFC 2189: Core Based Trees (CBT version 2) multicast routing, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2189.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2189.

txt. Status: EXPERIMENTAL.

RFC2190

[2177] C. Zhu. RFC 2190: RTP payload format for H.263 video streams,

September 1997. URL ftp://ftp.internic.net/rfc/rfc2190.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc2190.txt. Status: PRO-

POSED STANDARD.

RFC2191

[2178] G. Armitage. RFC 2191: VENUS — very extensive non-unicast service,

September 1997. URL ftp://ftp.internic.net/rfc/rfc2191.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2191.txt. Status: INFOR-

MATIONAL.

RFC2192

[2179] C. Newman. RFC 2192: IMAP URL scheme, September 1997. URL ftp:

//ftp.internic.net/rfc/rfc2192.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2192.txt. Status: PROPOSED STANDARD.

RFC2193

[2180] M. Gahrns. RFC 2193: IMAP4 mailbox referrals, September

1997. URL ftp://ftp.internic.net/rfc/rfc2193.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2193.txt. Status: PROPOSED

STANDARD.

RFC2194

[2181] B. Aboba, J. Lu, J. Alsop, J. Ding, and W. Wang. RFC 2194: Review of roaming implementations, September 1997. URL ftp://ftp.internic.

net/rfc/rfc2194.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2194.

txt. Status: INFORMATIONAL.

RFC2195

[2182] J. Klensin, R. Catoe, and P. Krumviede. RFC 2195: IMAP/ POP AUTHorize extension for simple challenge/response, September

1997. URL ftp://ftp.internic.net/rfc/rfc2095.txt,   ftp://
ftp.internic.net/rfc/rfc2195.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2095.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2195.

txt. Obsoletes RFC2095 [2082]. Status: PROPOSED STANDARD.

RFC2196

[2183] B. Fraser. RFC 2196: Site security handbook, September 1997. URL

ftp://ftp.internic.net/rfc/fyi8.txt,   ftp://ftp.internic.net/
rfc/rfc1244.txt,   ftp://ftp.internic.net/rfc/rfc2196.txt,
ftp://ftp.math.utah.edu/pub/rfc/fyi8.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1244.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2196.txt. See also FYI0008 [?]. Obsoletes RFC1244 [1234]. Status: INFORMATIONAL.

RFC2197

[2184] N. Freed. RFC 2197: SMTP service extension for command pipelin-

ing, September 1997. URL ftp://ftp.internic.net/rfc/rfc1854.
txt,   ftp://ftp.internic.net/rfc/rfc2197.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1854.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2197.txt. Obsoletes RFC1854 [1841]. Status: DRAFT STANDARD.

RFC2198

[2185] C. Perkins, I. Kouvelas, O. Hodson, V. Hardman, M. Handley, J. C. Bolot, A. Vega-Garcia, and S. Fosse-Parisis. RFC 2198: RTP payload for redundant audio data, September 1997. URL ftp://ftp.internic.

net/rfc/rfc2198.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2198.

txt. Status: PROPOSED STANDARD.

RFC2199

[2186] A. Ramos. RFC 2199: Request for comments summary RFC numbers 2100–2199, January 1998. URL ftp://ftp.internic.net/

rfc/rfc2199.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2199.

txt. Status: INFORMATIONAL.

RFC2200

[2187] J. Postel. RFC 2200: INTERNET OFFICIAL PROTOCOL STAN-

DARDS, June 1997. URL ftp://ftp.internic.net/rfc/rfc1083.
txt,   ftp://ftp.internic.net/rfc/rfc1100.txt,   ftp://ftp.
internic.net/rfc/rfc1130.txt,   ftp://ftp.internic.net/rfc/
rfc1140.txt,   ftp://ftp.internic.net/rfc/rfc1200.txt,   ftp:
//ftp.internic.net/rfc/rfc1250.txt,   ftp://ftp.internic.net/
rfc/rfc1280.txt,   ftp://ftp.internic.net/rfc/rfc1360.txt,
ftp://ftp.internic.net/rfc/rfc1410.txt,   ftp://ftp.internic.
net/rfc/rfc1500.txt,   ftp://ftp.internic.net/rfc/rfc1540.txt,
 ftp://ftp.internic.net/rfc/rfc1600.txt,   ftp://ftp.internic.
net/rfc/rfc1610.txt,   ftp://ftp.internic.net/rfc/rfc1720.txt,
 ftp://ftp.internic.net/rfc/rfc1780.txt,   ftp://ftp.internic.
net/rfc/rfc1800.txt,   ftp://ftp.internic.net/rfc/rfc1880.txt,
 ftp://ftp.internic.net/rfc/rfc1920.txt,   ftp://ftp.internic.
net/rfc/rfc2000.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
 ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1083.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1100.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1130.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1140.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1200.txt,   ftp:/

/ftp.math.utah.edu/pub/rfc/rfc1250.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1280.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1360.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1410.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1500.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1540.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1600.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1610.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1720.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1780.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1800.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1880.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1920.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2000.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.
txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt.  See also

STD0001 [?]. Obsoleted by RFC2300, RFC2400, STD0001 [2287, 2387, ?]. Obsoletes RFC2000, RFC1920, RFC1880, RFC1800, RFC1780, RFC1720, RFC1610, RFC1600, RFC1540, RFC1500, RFC1410, RFC1360, RFC1280, RFC1250, RFC1200, RFC1140, RFC1130, RFC1100, RFC1083 [1987, 1907, 1867, 1790, 1770, 1710, 1600, 1590, 1530, 1490, 1400, 1350, 1270, 1240, 1190, 1130, 1120, 1090, 1073]. Status: STANDARD.

RFC2201

[2188] A. Ballardie. RFC 2201: Core based trees (CBT) multicast routing architecture, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2201.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2201.

txt. Status: EXPERIMENTAL.

RFC2202

[2189] P. Cheng and R. Glenn. RFC 2202: Test cases for HMAC-MD5 and HMAC-SHA-1, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2202.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2202.

txt. Status: INFORMATIONAL.

RFC2203

[2190] M. Eisler, A. Chiu, and L. Ling. RFC 2203: RPCSEC GSS protocol specification, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2203.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2203.

txt. Status: PROPOSED STANDARD.

RFC2204

[2191] D. Nash. RFC 2204: ODETTE file transfer protocol, Septem-

ber 1997. URL ftp://ftp.internic.net/rfc/rfc2204.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc2204.txt. Status: INFORMA-

TIONAL.

RFC2205

[2192] R. Braden, Ed., L. Zhang, S. Berson, S. Herzog, and S. Jamin. RFC 2205: Resource ReSerVation Protocol (RSVP) — version 1 functional specifi-

cation, September 1997. URL ftp://ftp.internic.net/rfc/rfc2205.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2205.txt. Status:

PROPOSED STANDARD.

RFC2206

[2193] F. Baker, J. Krawczyk, and A. Sastry. RFC 2206: RSVP management information base using SMIv2, September 1997. URL ftp://ftp.

internic.net/rfc/rfc2206.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2206.txt. Status: PROPOSED STANDARD.

RFC2207

[2194] L. Berger and T. O’Malley. RFC 2207: RSVP extensions for IPSEC data

flows, September 1997. URL ftp://ftp.internic.net/rfc/rfc2207.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2207.txt. Status:

PROPOSED STANDARD.

RFC2208

[2195] A. Mankin, Ed., F. Baker, B. Braden, S. Bradner, M. O‘Dell, A. Romanow, A. Weinrib, and L. Zhang. RFC 2208: Resource ReSerVation Protocol (RSVP) — version 1 applicability statement some guidelines on deployment, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2208.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2208.

txt. Status: INFORMATIONAL.

RFC2209

[2196] R. Braden and L. Zhang. RFC 2209: Resource ReSerVation Protocol (RSVP) — version 1 message processing rules, September

1997. URL ftp://ftp.internic.net/rfc/rfc2209.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2209.txt. Status: INFORMA-

TIONAL.

RFC2210

[2197] J. Wroclawski. RFC 2210: The use of RSVP with IETF integrated ser-

vices, September 1997. URL ftp://ftp.internic.net/rfc/rfc2210.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2210.txt. Status:

PROPOSED STANDARD.

RFC2211

[2198] J. Wroclawski. RFC 2211: Specification of the controlled-load network element service, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2211.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2211.

txt. Status: PROPOSED STANDARD.

RFC2212

[2199] S. Shenker, C. Partridge, and R. Guerin. RFC 2212: Specification of guaranteed quality of service, September 1997. URL ftp://ftp.

internic.net/rfc/rfc2212.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2212.txt. Status: PROPOSED STANDARD.

RFC2213

[2200] F. Baker, J. Krawczyk, and A. Sastry. RFC 2213: Integrated services management information base using SMIv2, September

1997. URL ftp://ftp.internic.net/rfc/rfc2213.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2213.txt. Status: PROPOSED

STANDARD.

RFC2214

[2201] F. Baker, J. Krawczyk, and A. Sastry. RFC 2214: Integrated services management information base guaranteed service extensions using SMIv2, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2214.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2214.

txt. Status: PROPOSED STANDARD.

RFC2215

[2202] S. Shenker and J. Wroclawski. RFC 2215: General characterization parameters for integrated service network elements, Septem-

ber 1997. URL ftp://ftp.internic.net/rfc/rfc2215.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2215.txt. Status: PROPOSED

STANDARD.

RFC2216

[2203] S. Shenker and J. Wroclawski. RFC 2216: Network element service specification template, September 1997. URL ftp://ftp.internic.net/

rfc/rfc2216.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2216.

txt. Status: INFORMATIONAL.

RFC2217

[2204] G. Clark. RFC 2217: Telnet com port control option, October

1997. URL ftp://ftp.internic.net/rfc/rfc2217.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2217.txt. Status: EXPERIMEN-

TAL.

RFC2218

[2205] T. Genovese and B. Jennings. RFC 2218: A common schema for the Internet White Pages Service, October 1997. URL ftp://ftp.internic.

net/rfc/rfc2218.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2218.

txt. Status: PROPOSED STANDARD.

RFC2219

[2206] M. Hamilton and R. Wright. RFC 2219: Use of DNS aliases for network

services, October 1997. URL ftp://ftp.internic.net/rfc/bcp17.
txt,   ftp://ftp.internic.net/rfc/rfc2219.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp17.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2219.txt. See also BCP0017 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2220

[2207] R. Guenther. RFC 2220: The application/MARC content-type, Octo-

ber 1997. URL ftp://ftp.internic.net/rfc/rfc2220.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2220.txt. Status: INFORMA-

TIONAL.

RFC2221

[2208] M. Gahrns. RFC 2221: IMAP4 login referrals, October 1997. URL ftp:/

/ftp.internic.net/rfc/rfc2221.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2221.txt. Status: PROPOSED STANDARD.

RFC2222

[2209] J. Myers. RFC 2222: Simple Authentication and Security Layer (SASL),

October 1997. URL ftp://ftp.internic.net/rfc/rfc2222.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2222.txt. Status: PRO-

POSED STANDARD. Updated by RFC2444 [2431].

RFC2223

[2210] J. Postel and J. Reynolds. RFC 2223: Instructions to RFC au-

thors, October 1997. URL ftp://ftp.internic.net/rfc/rfc1111.
txt,   ftp://ftp.internic.net/rfc/rfc1543.txt,   ftp://ftp.
internic.net/rfc/rfc2223.txt,   ftp://ftp.internic.net/rfc/
rfc825.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1111.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1543.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2223.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc825.txt. Obsoletes RFC1543, RFC1111, RFC825 [1533, 1101, 815]. Status: INFORMATIONAL.

RFC2224

[2211] B. Callaghan. RFC 2224: NFS URL scheme, October 1997. URL ftp://

ftp.internic.net/rfc/rfc2224.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2224.txt. Status: INFORMATIONAL.

RFC2225

[2212] M. Laubach and J. Halpern. RFC 2225: Classical IP and ARP

over ATM, April 1998. URL ftp://ftp.internic.net/rfc/rfc1577.
txt,   ftp://ftp.internic.net/rfc/rfc1626.txt,   ftp://ftp.
internic.net/rfc/rfc2225.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1577.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1626.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2225.txt. Obsoletes

RFC1626, RFC1577 [1616, 1567]. Status: PROPOSED STANDARD.

RFC2226

[2213] T. Smith and G. Armitage. RFC 2226: IP broadcast over ATM networks,

October 1997. URL ftp://ftp.internic.net/rfc/rfc2226.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2226.txt. Status: PRO-

POSED STANDARD.

RFC2227

[2214] J. Mogul and P. Leach. RFC 2227: Simple hit-metering and usagelimiting for HTTP, October 1997. URL ftp://ftp.internic.net/

rfc/rfc2227.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2227.

txt. Status: PROPOSED STANDARD.

RFC2228

[2215] M. Horowitz and S. Lunt. RFC 2228: FTP security extensions, October

1997. URL ftp://ftp.internic.net/rfc/rfc2228.txt,   ftp://
ftp.internic.net/rfc/rfc959.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2228.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc959.

txt. Updates RFC0959 [949]. Status: PROPOSED STANDARD.

RFC2229

[2216] R. Faith and B. Martin. RFC 2229: A dictionary server protocol,

October 1997. URL ftp://ftp.internic.net/rfc/rfc2229.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2229.txt. Status: INFOR-

MATIONAL.

RFC2230

[2217] R. Atkinson. RFC 2230: Key exchange delegation record for the DNS,

October 1997. URL ftp://ftp.internic.net/rfc/rfc2230.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc2230.txt. Status: INFOR-

MATIONAL.

RFC2231

[2218] N. Freed and K. Moore. RFC 2231: MIME parameter value and encoded word extensions: Character sets, languages, and continua-

tions, November 1997. URL ftp://ftp.internic.net/rfc/rfc2045.
txt,   ftp://ftp.internic.net/rfc/rfc2047.txt,   ftp://ftp.
internic.net/rfc/rfc2183.txt,   ftp://ftp.internic.net/rfc/
rfc2184.txt,   ftp://ftp.internic.net/rfc/rfc2231.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2045.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2047.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc2183.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2184.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2231.txt. Obsoletes RFC2184

[2171]. Updates RFC2045, RFC2047, RFC2183 [2032, 2034, 2170]. Status: PROPOSED STANDARD.

RFC2232

[2219] B. Clouston and B. Moore. RFC 2232: Definitions of managed objects for DLUR using SMIv2, November 1997. URL ftp://ftp.internic.net/

rfc/rfc2232.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2232.

txt. Status: PROPOSED STANDARD.

RFC2233

[2220] K. McCloghrie and F. Kastenholz. RFC 2233: The interfaces group MIB

using SMIv2, November 1997. URL ftp://ftp.internic.net/rfc/
rfc1573.txt,   ftp://ftp.internic.net/rfc/rfc2233.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1573.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2233.txt. Obsoletes RFC1573 [1563]. Status: PRO-POSED STANDARD.

RFC2234

[2221] D. Crocker, Ed. and P. Overell. RFC 2234: Augmented BNF for syntax specifications: ABNF, November 1997. URL ftp://ftp.internic.net/

rfc/rfc2234.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2234.

txt. Status: PROPOSED STANDARD.

RFC2235

[2222] R. Zakon. RFC 2235: Hobbes’ Internet timeline, November 1997. URL

ftp://ftp.internic.net/rfc/fyi32.txt,   ftp://ftp.internic.
net/rfc/rfc2235.txt,   ftp://ftp.math.utah.edu/pub/rfc/fyi32.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2235.txt. See also

FYI0032 [?]. Status: INFORMATIONAL.

RFC2236

[2223] W. Fenner. RFC 2236: Internet Group Management Protocol, ver-

sion 2, November 1997. URL ftp://ftp.internic.net/rfc/rfc1112.
txt,   ftp://ftp.internic.net/rfc/rfc2236.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1112.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2236.txt. Updates RFC1112 [1102]. Status: PROPOSED STANDARD.

RFC2237

[2224] K. Tamaru. RFC 2237: Japanese character encoding for Internet mes-

sages, November 1997. URL ftp://ftp.internic.net/rfc/rfc2237.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2237.txt. Status:

INFORMATIONAL.

RFC2238

[2225] B. Clouston and B. Moore. RFC 2238: Definitions of managed objects for HPR using SMIv2, November 1997. URL ftp://ftp.internic.net/

rfc/rfc2238.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2238.

txt. Status: PROPOSED STANDARD.

RFC2239

[2226] R. Braden, L. Zhang, S. Berson, S. Herzog, and S. Jamin. RFC 2239: Definitions of managed objects for IEEE 802.3 medium attachment units (MAUs) using SMIv2, November 1997. URL ftp://ftp.internic.net/

rfc/rfc2239.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2239.

txt. Status: PROPOSED STANDARD.

RFC2240

[2227] O. Vaughan. RFC 2240: A legal basis for domain name alloca-

tion, November 1997. URL ftp://ftp.internic.net/rfc/rfc2240.
txt,   ftp://ftp.internic.net/rfc/rfc2352.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2240.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2352.txt. Obsoleted by RFC2352 [2339]. Status: INFOR-MATIONAL.

RFC2241

[2228] D. Provan. RFC 2241: DHCP options for Novell Directory Services,

November 1997. URL ftp://ftp.internic.net/rfc/rfc2241.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2241.txt. Status: PRO-

POSED STANDARD.

RFC2242

[2229] R. Droms and K. Fong. RFC 2242: NetWare/IP domain name and infor-

mation, November 1997. URL ftp://ftp.internic.net/rfc/rfc2242.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2242.txt. Status:

PROPOSED STANDARD.

RFC2243

[2230] C. Metz. RFC 2243: OTP extended responses, November 1997. URL

ftp://ftp.internic.net/rfc/rfc2243.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2243.txt. Status: PROPOSED STANDARD.

RFC2244

[2231] C. Newman and J. G. Myers. RFC 2244: ACAP — Application Configuration Access Protocol, November 1997. URL ftp://ftp.internic.

net/rfc/rfc2244.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2244.

txt. Status: PROPOSED STANDARD.

RFC2245

[2232] C. Newman. RFC 2245: Anonymous SASL mechanism, Novem-

ber 1997. URL ftp://ftp.internic.net/rfc/rfc2245.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2245.txt. Status: PROPOSED

STANDARD.

RFC2246

[2233] T. Dierks and C. Allen. RFC 2246: The TLS protocol version 1, Jan-

uary 1999. URL ftp://ftp.internic.net/rfc/rfc2246.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2246.txt. Status: PROPOSED

STANDARD.

RFC2247

[2234] S. Kille, M. Wahl, A. Grimstad, R. Huber, and S. Sataluri. RFC 2247: Using domains in LDAP/X.500 distinguished names, Jan-

uary 1998. URL ftp://ftp.internic.net/rfc/rfc2247.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2247.txt. Status: PROPOSED

STANDARD.

RFC2248

[2235] N. Freed and S. Kille. RFC 2248: Network services monitoring

MIB, January 1998. URL ftp://ftp.internic.net/rfc/rfc1565.
txt,   ftp://ftp.internic.net/rfc/rfc2248.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1565.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2248.txt. Obsoletes RFC1565 [1555]. Status: PROPOSED STANDARD.

RFC2249

[2236] N. Freed and S. Kille. RFC 2249: Mail monitoring MIB, January

1998. URL ftp://ftp.internic.net/rfc/rfc1566.txt,   ftp://

ftp.internic.net/rfc/rfc2249.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1566.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2249.

txt. Obsoletes RFC1566 [1556]. Status: PROPOSED STANDARD.

RFC2250

[2237] D. Hoffman, G. Fernando, V. Goyal, and M. Civanlar. RFC 2250: RTP payload format for MPEG1/MPEG2 video, January 1998. URL ftp://

ftp.internic.net/rfc/rfc2038.txt,   ftp://ftp.internic.net/
rfc/rfc2250.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2038.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2250.txt. Obsoletes

RFC2038 [2025]. Status: PROPOSED STANDARD.

RFC2251

[2238] M. Wahl, T. Howes, and S. Kille. RFC 2251: Lightweight Directory Access Protocol (v3), December 1997. URL ftp://ftp.internic.net/

rfc/rfc2251.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2251.

txt. Status: PROPOSED STANDARD.

RFC2252

[2239] M. Wahl, A. Coulbeck, T. Howes, and S. Kille. RFC 2252: Lightweight Directory Access Protocol (v3): Attribute syntax definitions, Decem-

ber 1997. URL ftp://ftp.internic.net/rfc/rfc2252.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2252.txt. Status: PROPOSED

STANDARD.

RFC2253

[2240] M. Wahl, S. Kille, and T. Howes. RFC 2253: Lightweight Directory Access Protocol (v3): UTF-8 string representation of distinguished names,

December 1997. URL ftp://ftp.internic.net/rfc/rfc2253.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2253.txt. Status: PRO-

POSED STANDARD.

RFC2254

[2241] T. Howes. RFC 2254: The string representation of LDAP search filters,

December 1997. URL ftp://ftp.internic.net/rfc/rfc2254.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2254.txt. Status: PRO-

POSED STANDARD.

RFC2255

[2242] T. Howes and M. Smith. RFC 2255: The LDAP URL format, Decem-

ber 1997. URL ftp://ftp.internic.net/rfc/rfc2255.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2255.txt. Status: PROPOSED

STANDARD.

RFC2256

[2243] M. Wahl. RFC 2256: A summary of the X.500(96) user schema for use with LDAPv3, December 1997. URL ftp://ftp.internic.net/

rfc/rfc2256.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2256.

txt. Status: PROPOSED STANDARD.

RFC2257

[2244] M. Daniele, B. Wijnen, and D. Francisco. RFC 2257: Agent Extensibility (AgentX) Protocol Version 1, January 1998. URL ftp://ftp.internic.

net/rfc/rfc2257.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2257.

txt. Status: PROPOSED STANDARD.

RFC2258

[2245] J. Ordille. RFC 2258: Internet nomenclator project, January

1998. URL ftp://ftp.internic.net/rfc/rfc2258.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2258.txt. Status: INFORMA-

TIONAL.

RFC2259

[2246] J. Elliott and J. Ordille. RFC 2259: Simple Nomenclator Query Protocol

(SNQP), January 1998. URL ftp://ftp.internic.net/rfc/rfc2259.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2259.txt. Status:

INFORMATIONAL.

RFC2260

[2247] T. Bates and Y. Rekhter. RFC 2260: Scalable support for multi-homed multi-provider connectivity, January 1998. URL ftp://ftp.internic.

net/rfc/rfc2260.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2260.

txt. Status: INFORMATIONAL.

RFC2261

[2248] D. Harrington, R. Presuhn, and B. Wijnen. RFC 2261: An architecture for describing SNMP management frameworks, January

1998. URL ftp://ftp.internic.net/rfc/rfc2261.txt,   ftp://
ftp.internic.net/rfc/rfc2271.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2261.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2271.

txt. Obsoleted by RFC2271 [2258]. Status: PROPOSED STANDARD.

RFC2262

[2249] J. Case, D. Harrington, R. Presuhn, and B. Wijnen. RFC 2262: Message processing and dispatching for the Simple Network Management Protocol (SNMP), January 1998. URL ftp://ftp.

internic.net/rfc/rfc2262.txt,   ftp://ftp.internic.net/rfc/
rfc2272.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2262.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc2272.txt. Obsoleted by

RFC2272 [2259]. Status: PROPOSED STANDARD.

RFC2263

[2250] D. Levi, P. Meyer, and B. Stewart. RFC 2263: SNMPv3 applica-

tions, January 1998. URL ftp://ftp.internic.net/rfc/rfc2263.
txt,   ftp://ftp.internic.net/rfc/rfc2273.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2263.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2273.txt. Obsoleted by RFC2273 [2260]. Status: PRO-POSED STANDARD.

RFC2264

[2251] U. Blumenthal and B. Wijnen. RFC 2264: User-based Security Model (USM) for version 3 of the Simple Network Management Protocol (SNMPv3), January 1998. URL ftp://ftp.

internic.net/rfc/rfc2264.txt,   ftp://ftp.internic.net/rfc/
rfc2274.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2264.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2274.txt. Obsoleted by

RFC2274 [2261]. Status: PROPOSED STANDARD.

RFC2265

[2252] B. Wijnen, R. Presuhn, and K. McCloghrie. RFC 2265: View-based Access Control Model (VACM) for the Simple Network Management Protocol (SNMP), January 1998. URL ftp://ftp.internic.net/rfc/

rfc2265.txt,   ftp://ftp.internic.net/rfc/rfc2275.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2265.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2275.txt. Obsoleted by RFC2275 [2262]. Status: PROPOSED STANDARD.

RFC2266

[2253] J. Flick. RFC 2266: Definitions of managed objects for IEEE 802.12 repeater devices, January 1998. URL ftp://ftp.internic.net/

rfc/rfc2266.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2266.

txt. Status: PROPOSED STANDARD.

RFC2267

[2254] P. Ferguson and D. Senie. RFC 2267: Network ingress filtering: Defeating denial of service attacks which employ IP source address spoof-

ing, January 1998. URL ftp://ftp.internic.net/rfc/rfc2267.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2267.txt. Status: INFOR-

MATIONAL.

RFC2268

[2255] R. Rivest. RFC 2268: A description of the RC2(r) encryption algorithm,

January 1998. URL ftp://ftp.internic.net/rfc/rfc2268.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2268.txt. Status: INFOR-

MATIONAL.

RFC2269

[2256] G. Armitage. RFC 2269: Using the MARS model in non-ATM NBMA

networks, January 1998. URL ftp://ftp.internic.net/rfc/rfc2269.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2269.txt. Status:

INFORMATIONAL.

RFC2270

[2257] J. Stewart, T. Bates, R. Chandra, and E. Chen. RFC 2270: Using a dedicated AS for sites homed to a single provider, Jan-

uary 1998. URL ftp://ftp.internic.net/rfc/rfc2270.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2270.txt. Status: INFORMA-

TIONAL.

RFC2271

[2258] D. Harrington, R. Presuhn, and B. Wijnen. RFC 2271: An architecture for describing SNMP management frameworks, January

1998. URL ftp://ftp.internic.net/rfc/rfc2261.txt,   ftp://
ftp.internic.net/rfc/rfc2271.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2261.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2271.

txt. Obsoletes RFC2261 [2248]. Status: PROPOSED STANDARD.

RFC2272

[2259] J. Case, D. Harrington, R. Presuhn, and B. Wijnen. RFC 2272: Message processing and dispatching for the Simple Network Management Protocol (SNMP), January 1998. URL ftp://

ftp.internic.net/rfc/rfc2262.txt,   ftp://ftp.internic.net/
rfc/rfc2272.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2262.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2272.txt. Obsoletes

RFC2262 [2249]. Status: PROPOSED STANDARD.

RFC2273

[2260] D. Levi, P. Meyer, and B. Stewart. RFC 2273: SNMPv3 applica-

tions, January 1998. URL ftp://ftp.internic.net/rfc/rfc2263.
txt,   ftp://ftp.internic.net/rfc/rfc2273.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2263.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2273.txt. Obsoletes RFC2263 [2250]. Status: PROPOSED STANDARD.

RFC2274

[2261] U. Blumenthal and B. Wijnen. RFC 2274: User-based Security Model (USM) for version 3 of the Simple Network Management Protocol (SNMPv3), January 1998. URL ftp://ftp.

internic.net/rfc/rfc2264.txt,   ftp://ftp.internic.net/rfc/
rfc2274.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2264.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2274.txt. Obsoletes RFC2264

[2251]. Status: PROPOSED STANDARD.

RFC2275

[2262] B. Wijnen, R. Presuhn, and K. McCloghrie. RFC 2275: View-based Access Control Model (VACM) for the Simple Network Management Protocol (SNMP), January 1998. URL ftp://ftp.internic.net/rfc/

rfc2265.txt,   ftp://ftp.internic.net/rfc/rfc2275.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2265.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2275.txt. Obsoletes RFC2265 [2252]. Status: PRO-POSED STANDARD.

RFC2276

[2263] K. Sollins. RFC 2276: Architectural principles of Uniform Resource Name resolution, January 1998. URL ftp://ftp.internic.net/

rfc/rfc2276.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2276.

txt. Status: INFORMATIONAL.

RFC2277

[2264] H. Alvestrand. RFC 2277: IETF policy on character sets and lan-

guages, January 1998. URL ftp://ftp.internic.net/rfc/bcp18.
txt,   ftp://ftp.internic.net/rfc/rfc2277.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp18.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2277.txt. See also BCP0018 [?]. Status: Best Current Practice.

RFC2278

[2265] N. Freed and J. Postel. RFC 2278: IANA charset registration pro-

cedures, January 1998. URL ftp://ftp.internic.net/rfc/bcp19.
txt,   ftp://ftp.internic.net/rfc/rfc2278.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp19.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2278.txt. See also BCP0019 [?]. Status: Best Current Practice.

RFC2279

[2266] F. Yergeau. RFC 2279: UTF-8, a transformation format of ISO

10646, January 1998. URL ftp://ftp.internic.net/rfc/rfc2044.

txt,   ftp://ftp.internic.net/rfc/rfc2279.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2044.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2279.txt. Obsoletes RFC2044 [2031]. Status: PROPOSED STANDARD.

RFC2280

[2267] C. Alaettinoglu, T. Bates, E. Gerich, D. Karrenberg, D. Meyer, M. Terpstra, and C. Villamizar. RFC 2280: Routing Policy Specification Language (RPSL), January 1998. URL ftp://ftp.internic.net/

rfc/rfc2280.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2280.

txt. Status: PROPOSED STANDARD.

RFC2281

[2268] T. Li, B. Cole, P. Morton, and D. Li. RFC 2281: Cisco Hot Standby Router Protocol (HSRP), March 1998. URL ftp://ftp.internic.net/

rfc/rfc2281.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2281.

txt. Status: INFORMATIONAL.

RFC2282

[2269] J. Galvin. RFC 2282: IAB and IESG selection, confirmation, and recall process: Operation of the nominating and recall committees, February

1998. URL ftp://ftp.internic.net/rfc/bcp10.txt,   ftp://ftp.
internic.net/rfc/rfc2027.txt,   ftp://ftp.internic.net/rfc/
rfc2282.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp10.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2027.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2282.txt. See also BCP0010 [?]. Obsoletes

RFC2027 [2014]. Status: BEST CURRENT PRACTICE.

RFC2283

[2270] T. Bates, R. Chandra, D. Katz, and Y. Rekhter. RFC 2283: Multiprotocol extensions for BGP-4, February 1998. URL ftp://ftp.internic.

net/rfc/rfc2283.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2283.

txt. Status: PROPOSED STANDARD.

RFC2284

[2271] L. Blunk and J. Vollbrecht. RFC 2284: PPP extensible authentication protocol (EAP), March 1998. URL ftp://ftp.internic.net/

rfc/rfc2284.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2284.

txt. Status: PROPOSED STANDARD.

RFC2285

[2272] R. Mandeville. RFC 2285: Benchmarking terminology for LAN switching

devices, February 1998. URL ftp://ftp.internic.net/rfc/rfc2285.

txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2285.txt. Status:

INFORMATIONAL.

RFC2286

[2273] J. Kapp. RFC 2286: Test cases for HMAC-RIPEMD160 and HMAC-RIPEMD128, February 1998. URL ftp://ftp.internic.net/

rfc/rfc2286.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2286.

txt. Status: INFORMATIONAL.

RFC2287

[2274] C. Krupczak and J. Saperia. RFC 2287: Definitions of system-level managed objects for applications, February 1998. URL ftp://ftp.

internic.net/rfc/rfc2287.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2287.txt. Status: PROPOSED STANDARD.

RFC2288

[2275] C. Lynch, C. Preston, and R. Daniel. RFC 2288: Using existing bibliographic identifiers as uniform resource names, Febru-

ary 1998. URL ftp://ftp.internic.net/rfc/rfc2288.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2288.txt. Status: INFORMA-

TIONAL.

RFC2289

[2276] N. Haller, C. Metz, P. Nesser, and M. Straw. RFC 2289: A one-time password system, February 1998. URL ftp://ftp.internic.net/rfc/

rfc1938.txt,   ftp://ftp.internic.net/rfc/rfc2289.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1938.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2289.txt. Obsoletes RFC1938 [1925]. Status: DRAFT STANDARD.

RFC2290

[2277] J. Solomon and S. Glass. RFC 2290: Mobile-IPv4 configuration option for PPP IPCP, February 1998. URL ftp://ftp.

internic.net/rfc/rfc2002.txt,   ftp://ftp.internic.net/rfc/
rfc2290.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2002.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2290.txt. Updates RFC2002

[1989]. Status: PROPOSED STANDARD.

RFC2291

[2278] J. Slein, F. Vitali, E. Whitehead, and D. Durand. RFC 2291: Requirements for a distributed authoring and versioning protocol for the World Wide Web, February 1998. URL ftp://ftp.internic.net/

rfc/rfc2291.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2291.

txt. Status: INFORMATIONAL.

RFC2292

[2279] W. Stevens and M. Thomas. RFC 2292: Advanced sockets API for IPv6,

February 1998. URL ftp://ftp.internic.net/rfc/rfc2292.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2292.txt. Status: INFOR-

MATIONAL.

RFC2293

[2280] S. Kille. RFC 2293: Representing tables and subtrees in the X.500 di-

rectory, March 1998. URL ftp://ftp.internic.net/rfc/rfc1837.
txt,   ftp://ftp.internic.net/rfc/rfc2293.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1837.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2293.txt. Obsoletes RFC1837 [1827]. Status: PROPOSED STANDARD.

RFC2294

[2281] S. Kille. RFC 2294: Representing the O/R address hierarchy in the X.500 directory information tree, March 1998. URL ftp://

ftp.internic.net/rfc/rfc1836.txt,   ftp://ftp.internic.net/
rfc/rfc2294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1836.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2294.txt. Obsoletes

RFC1836 [1826]. Status: PROPOSED STANDARD.

RFC2295

[2282] K. Holtman and A. Mutz. RFC 2295: Transparent content negotiation

in HTTP, March 1998. URL ftp://ftp.internic.net/rfc/rfc2295.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2295.txt. Status:

EXPERIMENTAL.

RFC2296

[2283] K. Holtman and A. Mutz. RFC 2296: HTTP remote variant selection algorithm — RVSA/1.0, March 1998. URL ftp://ftp.internic.net/

rfc/rfc2296.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2296.

txt. Status: EXPERIMENTAL.

RFC2297

[2284] P. Newman, W. Edwards, R. Hinden, E. Hoffman, F. Ching Liaw, T. Lyon, and G. Minshall. RFC 2297: Ipsilon’s general switch management protocol specification version 2.0, March 1998. URL ftp://

ftp.internic.net/rfc/rfc1987.txt,   ftp://ftp.internic.net/
rfc/rfc2297.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1987.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2297.txt. Updates

RFC1987 [1974]. Status: INFORMATIONAL.

RFC2298

[2285] R. Fajman. RFC 2298: An extensible message format for message disposition notifications, March 1998. URL ftp://ftp.internic.net/

rfc/rfc2298.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2298.

txt. Status: PROPOSED STANDARD.

RFC2299

[2286] A. Ramos. RFC 2299: Request for comments summary RFC numbers 2200–2299, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2299.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2299.

txt. Status: INFORMATIONAL.

RFC2300

[2287] J. Postel. RFC 2300: INTERNET OFFICIAL PROTOCOL STAN-

DARDS, May 1998. URL ftp://ftp.internic.net/rfc/rfc1083.
txt,   ftp://ftp.internic.net/rfc/rfc1100.txt,   ftp://ftp.
internic.net/rfc/rfc1130.txt,   ftp://ftp.internic.net/rfc/
rfc1140.txt,   ftp://ftp.internic.net/rfc/rfc1200.txt,   ftp:
//ftp.internic.net/rfc/rfc1250.txt,   ftp://ftp.internic.net/
rfc/rfc1280.txt,   ftp://ftp.internic.net/rfc/rfc1360.txt,
ftp://ftp.internic.net/rfc/rfc1410.txt,   ftp://ftp.internic.
net/rfc/rfc1500.txt,   ftp://ftp.internic.net/rfc/rfc1540.txt,
 ftp://ftp.internic.net/rfc/rfc1600.txt,   ftp://ftp.internic.
net/rfc/rfc1610.txt,   ftp://ftp.internic.net/rfc/rfc1720.txt,
 ftp://ftp.internic.net/rfc/rfc1780.txt,   ftp://ftp.internic.
net/rfc/rfc1800.txt,   ftp://ftp.internic.net/rfc/rfc1880.txt,
 ftp://ftp.internic.net/rfc/rfc1920.txt,   ftp://ftp.internic.
net/rfc/rfc2000.txt,   ftp://ftp.internic.net/rfc/rfc2200.txt,
 ftp://ftp.internic.net/rfc/rfc2300.txt,   ftp://ftp.internic.
net/rfc/rfc2400.txt, ftp://ftp.internic.net/rfc/std1.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc1083.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1100.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1130.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1140.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1200.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1250.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1280.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1360.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1410.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1500.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1540.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1600.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1610.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1720.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1780.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1800.txt,   ftp://ftp.math.utah.edu/pub/rfc/

rfc1880.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1920.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2000.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2200.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc2300.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2400.
txt,   ftp://ftp.math.utah.edu/pub/rfc/std1.txt.  See also

STD0001 [?]. Obsoleted by RFC2400, STD0001 [2387, ?]. Obsoletes RFC2200, RFC2000, RFC1920, RFC1880, RFC1800, RFC1780, RFC1720, RFC1610, RFC1600, RFC1540, RFC1500, RFC1410, RFC1360, RFC1280, RFC1250, RFC1200, RFC1140, RFC1130, RFC1100, RFC1083 [2187, 1987, 1907, 1867, 1790, 1770, 1710, 1600, 1590, 1530, 1490, 1400, 1350, 1270, 1240, 1190, 1130, 1120, 1090, 1073]. Status: STANDARDS TRACK.

RFC2301

[2288] L. McIntyre, S. Zilles, R. Buckley, D. Venable, G. Parsons, and J. Rafferty. RFC 2301: File format for Internet Fax, March

1998. URL ftp://ftp.internic.net/rfc/rfc2301.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2301.txt. Status: PROPOSED

STANDARD.

RFC2302

[2289] G. Parsons, J. Rafferty, and S. Zilles. RFC 2302: Tag image file format (TIFF) — image/tiff MIME sub-type registration, March

1998. URL ftp://ftp.internic.net/rfc/rfc2302.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2302.txt. Status: PROPOSED

STANDARD.

RFC2303

[2290] C. Allocchio. RFC 2303: Minimal PSTN address format in Internet

mail, March 1998. URL ftp://ftp.internic.net/rfc/rfc2303.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2303.txt. Status: PRO-

POSED STANDARD.

RFC2304

[2291] C. Allocchio. RFC 2304: Minimal FAX address format in Internet

mail, March 1998. URL ftp://ftp.internic.net/rfc/rfc2304.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2304.txt. Status: PRO-

POSED STANDARD.

RFC2305

[2292] K. Toyoda, H. Ohno, J. Murai, and D. Wing. RFC 2305: A simple mode of facsimile using Internet mail, March 1998. URL ftp://ftp.

internic.net/rfc/rfc2305.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2305.txt. Status: PROPOSED STANDARD.

RFC2306

[2293] G. Parsons and J. Rafferty. RFC 2306: Tag image file format (TIFF) — F profile for facsimile, March 1998. URL ftp://ftp.internic.net/

rfc/rfc2306.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2306.

txt. Status: INFORMATIONAL.

RFC2307

[2294] L. Howard. RFC 2307: An approach for using LDAP as a network information service, March 1998. URL ftp://ftp.internic.net/

rfc/rfc2307.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2307.

txt. Status: EXPERIMENTAL.

RFC2308

[2295] M. Andrews. RFC 2308: Negative caching of DNS queries (DNS

NCACHE), March 1998. URL ftp://ftp.internic.net/rfc/rfc1034.
txt,   ftp://ftp.internic.net/rfc/rfc1035.txt,   ftp://ftp.
internic.net/rfc/rfc2308.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1034.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1035.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2308.txt. Updates

RFC1034, RFC1035 [1024, 1025]. Status: PROPOSED STANDARD.

RFC2309

[2296] B. Braden, D. Clark, J. Crowcroft, B. Davie, S. Deering, D. Estrin, S. Floyd, V. Jacobson, G. Minshall, C. Partridge, L. Peterson, K. Ramakrishnan, S. Shenker, J. Wroclawski, and L. Zhang. RFC 2309: Recommendations on queue management and congestion avoidance in the In-

ternet, April 1998. URL ftp://ftp.internic.net/rfc/rfc2309.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2309.txt. Status: INFOR-

MATIONAL.

RFC2310

[2297] K. Holtman. RFC 2310: The safe response header field, April

1998. URL ftp://ftp.internic.net/rfc/rfc2310.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2310.txt. Status: EXPERIMEN-

TAL.

RFC2311

[2298] S. Dusse, P. Hoffman, B. Ramsdell, L. Lundblade, and L. Repka. RFC 2311: S/MIME version 2 message specification, March 1998. URL ftp:/

/ftp.internic.net/rfc/rfc2311.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2311.txt. Status: INFORMATIONAL.

RFC2312

[2299] S. Dusse, P. Hoffman, B. Ramsdell, and J. Weinstein. RFC 2312: S/ MIME version 2 certificate handling, March 1998. URL ftp://ftp.

internic.net/rfc/rfc2312.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2312.txt. Status: INFORMATIONAL.

RFC2313

[2300] B. Kaliski. RFC 2313: PKCS #1: RSA encryption version

1.5, March 1998. URL ftp://ftp.internic.net/rfc/rfc2313.
txt,   ftp://ftp.internic.net/rfc/rfc2437.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2313.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2437.txt. Obsoleted by RFC2437 [2424]. Status: INFOR-MATIONAL.

RFC2314

[2301] B. Kaliski. RFC 2314: PKCS #10: Certification request syntax version

1.5, March 1998. URL ftp://ftp.internic.net/rfc/rfc2314.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2314.txt. Status: INFOR-

MATIONAL.

RFC2315

[2302] B. Kaliski. RFC 2315: PKCS #7: Cryptographic message syntax version

1.5, March 1998. URL ftp://ftp.internic.net/rfc/rfc2315.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2315.txt. Status: INFOR-

MATIONAL.

RFC2316

[2303] S. Bellovin. RFC 2316: Report of the IAB security architecture work-

shop, April 1998. URL ftp://ftp.internic.net/rfc/rfc2316.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2316.txt. Status: INFOR-

MATIONAL.

RFC2317

[2304] ARPA delegation. RFC 2317: Classless IN-ADDR, March 1998. URL

ftp://ftp.internic.net/rfc/bcp20.txt,   ftp://ftp.internic.
net/rfc/rfc2317.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp20.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2317.txt. See also

BCP0020 [?]. Status: Best Current Practice.

RFC2318

[2305] H. Lie, B. Bos, and C. Lilley. RFC 2318: The text/css media

type, March 1998. URL ftp://ftp.internic.net/rfc/rfc2318.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc2318.txt. Status: INFOR-

MATIONAL.

RFC2319

[2306] KOI8-U Working Group. RFC 2319: Ukrainian character set KOI8-

U, April 1998. URL ftp://ftp.internic.net/rfc/rfc2319.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2319.txt. Status: INFOR-

MATIONAL.

RFC2320

[2307] M. Greene, J. Luciani, K. White, and T. Kuo. RFC 2320: Definitions of managed objects for classical IP and ARP over ATM using SMIv2 (IPOA-MIB), April 1998. URL ftp://ftp.internic.net/

rfc/rfc2320.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2320.

txt. Status: PROPOSED STANDARD.

RFC2321

[2308] A. Bressen. RFC 2321: RITA — the Reliable Internetwork Troubleshoot-

ing Agent, March 1998. URL ftp://ftp.internic.net/rfc/rfc2321.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2321.txt. Status:

INFORMATIONAL.

RFC2322

[2309] K. van den Hout, A. Koopal, and R. van Mook. RFC 2322: Management of IP numbers by peg-dhcp, March 1998. URL ftp://ftp.internic.

net/rfc/rfc2322.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2322.

txt. Status: INFORMATIONAL.

RFC2323

[2310] A. Ramos. RFC 2323: IETF identification and security guidelines,

March 1998. URL ftp://ftp.internic.net/rfc/rfc2323.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2323.txt. Status: INFOR-

MATIONAL.

RFC2324

[2311] L. Masinter. RFC 2324: Hyper text coffee pot control protocol (HTCPCP/1.0), March 1998. URL ftp://ftp.internic.net/

rfc/rfc2324.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2324.

txt. Status: INFORMATIONAL.

RFC2325

[2312] M. Slavitch. RFC 2325: Definitions of managed objects for drip-type heated beverage hardware devices using SMIv2, March

1998. URL ftp://ftp.internic.net/rfc/rfc2325.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2325.txt. Status: INFORMA-

TIONAL.

RFC2326

[2313] H. Schulzrinne, A. Rao, and R. Lanphier. RFC 2326: Real time streaming protocol (RTSP), April 1998. URL ftp://ftp.internic.net/

rfc/rfc2326.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2326.

txt. Status: PROPOSED STANDARD.

RFC2327

[2314] M. Handley and V. Jacobson. RFC 2327: SDP: Session description pro-

tocol, April 1998. URL ftp://ftp.internic.net/rfc/rfc2327.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2327.txt. Status: PRO-

POSED STANDARD.

RFC2328

[2315] J. Moy. RFC 2328: OSPF version 2, April 1998. URL ftp://ftp.

internic.net/rfc/rfc2178.txt,   ftp://ftp.internic.net/rfc/
rfc2328.txt, ftp://ftp.internic.net/rfc/std54.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2178.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2328.txt,   ftp://ftp.math.utah.edu/pub/rfc/

std54.txt. See also STD0054 [?]. Obsoletes RFC2178 [2165]. Status: STANDARD.

RFC2329

[2316] J. Moy. RFC 2329: OSPF standardization report, April 1998. URL ftp:

//ftp.internic.net/rfc/rfc2329.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2329.txt. Status: INFORMATIONAL.

RFC2330

[2317] V. Paxson, G. Almes, J. Mahdavi, and M. Mathis. RFC 2330: Framework for IP performance metrics, May 1998. URL ftp://ftp.internic.net/

rfc/rfc2330.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2330.

txt. Status: INFORMATIONAL.

RFC2331

[2318] Update. RFC 2331: ATM signalling support for IP over ATM — UNI sig-

nalling 4.0, April 1998. URL ftp://ftp.internic.net/rfc/rfc2331.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2331.txt. Status:

PROPOSED STANDARD.

RFC2332

[2319] J. Luciani, D. Katz, D. Piscitello, B. Cole, and N. Doraswamy. RFC 2332: NBMA next hop resolution protocol (NHRP), April

1998. URL ftp://ftp.internic.net/rfc/rfc2332.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2332.txt. Status: PROPOSED

STANDARD.

RFC2333

[2320] D. Cansever. RFC 2333: NHRP protocol applicability statement, April

1998. URL ftp://ftp.internic.net/rfc/rfc2333.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2333.txt. Status: PROPOSED

STANDARD.

RFC2334

[2321] J. Luciani, G. Armitage, J. Halpern, and N. Doraswamy. RFC 2334: Server cache synchronization protocol (SCSP), April 1998. URL ftp://

ftp.internic.net/rfc/rfc2334.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2334.txt. Status: PROPOSED STANDARD.

RFC2335

[2322] J. Luciani. RFC 2335: A distributed NHRP service using SCSP, April

1998. URL ftp://ftp.internic.net/rfc/rfc2335.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2335.txt. Status: PROPOSED

STANDARD.

RFC2336

[2323] J. Luciani. RFC 2336: Classical IP to NHRP transition, July

1998. URL ftp://ftp.internic.net/rfc/rfc2336.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2336.txt. Status: INFORMA-

TIONAL.

RFC2337

[2324] D. Farinacci, D. Meyer, and Y. Rekhter. RFC 2337: Intra-LIS IP multicast among routers over ATM using sparse mode PIM, April

1998. URL ftp://ftp.internic.net/rfc/rfc2337.txt,   ftp://
ftp.math.utah.edu/pub/rfc/rfc2337.txt. Status: EXPERIMEN-

TAL.

RFC2338

[2325] S. Knight, D. Weaver, D. Whipple, R. Hinden, D. Mitzel, P. Hunt, P. Higginson, M. Shand, and A. Lindem. RFC 2338: Virtual router redundancy protocol, April 1998. URL ftp://ftp.internic.net/

rfc/rfc2338.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2338.

txt. Status: PROPOSED STANDARD.

RFC2339

[2326] The Internet Society and Sun Microsystems. RFC 2339: An agreement between the Internet Society, the IETF, and Sun Microsystems, Inc. in the matter of NFS V.4 protocols, May 1998. URL ftp://ftp.internic.

net/rfc/rfc2339.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2339.

txt. Status: INFORMATIONAL.

RFC2340

[2327] B. Jamoussi, D. Jamieson, D. Williston, and S. Gabe. RFC 2340: Nortel’s Virtual Network Switching (VNS) overview, May

1998. URL ftp://ftp.internic.net/rfc/rfc2340.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2340.txt. Status: INFORMA-

TIONAL.

RFC2341

[2328] A. Valencia, M. Littlewood, and T. Kolar. RFC 2341: Cisco Layer Two Forwarding (protocol) “L2F”, May 1998. URL ftp://ftp.internic.

net/rfc/rfc2341.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2341.

txt. Status: HISTORIC.

RFC2342

[2329] M. Gahrns and C. Newman. RFC 2342: IMAP4 namespace, May

1998. URL ftp://ftp.internic.net/rfc/rfc2342.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2342.txt. Status: PROPOSED

STANDARD.

RFC2343

[2330] M. Civanlar, G. Cash, and B. Haskell. RFC 2343: RTP payload format for bundled MPEG, May 1998. URL ftp://ftp.internic.net/

rfc/rfc2343.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2343.

txt. Status: EXPERIMENTAL.

RFC2344

[2331] G. Montenegro. RFC 2344: Reverse tunneling for mobile IP, May

1998. URL ftp://ftp.internic.net/rfc/rfc2344.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2344.txt. Status: PROPOSED

STANDARD.

RFC2345

[2332] J. Klensin, T. Wolf, and G. Oglesby. RFC 2345: Domain names and company name retrieval, May 1998. URL ftp://ftp.internic.net/

rfc/rfc2345.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2345.

txt. Status: EXPERIMENTAL.

RFC2346

[2333] J. Palme. RFC 2346: Making PostScript and PDF international,

May 1998. URL ftp://ftp.internic.net/rfc/rfc2346.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc2346.txt. Status: INFORMA-

TIONAL.

RFC2347

[2334] G. Malkin and A. Harkin. RFC 2347: TFTP option extension, May

1998. URL ftp://ftp.internic.net/rfc/rfc1350.txt,   ftp://
ftp.internic.net/rfc/rfc1782.txt,   ftp://ftp.internic.net/
rfc/rfc2347.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1350.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1782.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2347.txt.  Obsoletes RFC1782

[1772]. Updates RFC1350 [1340]. Status: DRAFT STANDARD.

RFC2348

[2335] G. Malkin and A. Harkin. RFC 2348: TFTP blocksize option, May

1998. URL ftp://ftp.internic.net/rfc/rfc1350.txt,   ftp://
ftp.internic.net/rfc/rfc1783.txt,   ftp://ftp.internic.net/
rfc/rfc2348.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1350.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1783.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2348.txt.  Obsoletes RFC1783

[1773]. Updates RFC1350 [1340]. Status: DRAFT STANDARD.

RFC2349

[2336] G. Malkin and A. Harkin. RFC 2349: TFTP timeout interval and transfer size options, May 1998. URL ftp://ftp.internic.net/rfc/

rfc1350.txt,   ftp://ftp.internic.net/rfc/rfc1784.txt,   ftp:
//ftp.internic.net/rfc/rfc2349.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1350.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1784.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2349.txt. Obsoletes

RFC1784 [1774]. Updates RFC1350 [1340]. Status: DRAFT STAN-DARD.

RFC2350

[2337] N. Brownlee and E. Guttman. RFC 2350: Expectations for computer security incident response, June 1998. URL ftp://ftp.internic.net/

rfc/bcp21.txt,   ftp://ftp.internic.net/rfc/rfc2350.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/bcp21.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2350.txt. See also BCP0021 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2351

[2338] A. Rober. RFC 2351: Mapping of airline reservation, ticketing, and messaging traffic over IP, May 1998. URL ftp://ftp.internic.net/

rfc/rfc2351.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2351.

txt. Status: INFORMATIONAL.

RFC2352

[2339] O. Vaughan. RFC 2352: A convention for using legal names as do-

main names, May 1998. URL ftp://ftp.internic.net/rfc/rfc2240.
txt,   ftp://ftp.internic.net/rfc/rfc2352.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2240.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2352.txt. Obsoletes RFC2240 [2227]. Status: INFORMA-TIONAL.

RFC2353

[2340] G. Dudley. RFC 2353: APPN/HPR in IP networks APPN implementers’ workshop closed pages document, May 1998. URL ftp://ftp.

internic.net/rfc/rfc2353.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2353.txt. Status: INFORMATIONAL.

RFC2354

[2341] C. Perkins and O. Hodson. RFC 2354: Options for repair of streaming

media, June 1998. URL ftp://ftp.internic.net/rfc/rfc2354.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2354.txt. Status: INFOR-

MATIONAL.

RFC2355

[2342] B. Kelly. RFC 2355: TN3270 enhancements, June 1998. URL ftp://

ftp.internic.net/rfc/rfc1647.txt,   ftp://ftp.internic.net/
rfc/rfc2355.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1647.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2355.txt. Obsoletes

RFC1647 [1637]. Status: DRAFT STANDARD.

RFC2356

[2343] G. Montenegro and V. Gupta. RFC 2356: Sun’s SKIP firewall traversal for mobile IP, June 1998. URL ftp://ftp.internic.net/

rfc/rfc2356.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2356.

txt. Status: INFORMATIONAL.

RFC2357

[2344] A. Mankin, A. Romanow, S. Bradner, and V. Paxson. RFC 2357: IETF criteria for evaluating reliable multicast transport and application

protocols, June 1998. URL ftp://ftp.internic.net/rfc/rfc1650.
txt,   ftp://ftp.internic.net/rfc/rfc2357.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1650.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2357.txt. Obsoletes RFC1650 [1640]. Status: INFORMA-TIONAL.

RFC2358

[2345] J. Flick and J. Johnson. RFC 2358: Definitions of managed objects for the Ethernet-like interface types, June 1998. URL ftp://

ftp.internic.net/rfc/rfc1650.txt,   ftp://ftp.internic.net/
rfc/rfc2358.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1650.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2358.txt. Obsoletes

RFC1650 [1640]. Status: PROPOSED STANDARD.

RFC2359

[2346] J. Myers. RFC 2359: IMAP4 UIDPLUS extension, June 1998. URL ftp:

//ftp.internic.net/rfc/rfc2359.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2359.txt. Status: PROPOSED STANDARD.

RFC2360

[2347] G. Scott. RFC 2360: Guide for Internet standards writers, June

1998. URL ftp://ftp.internic.net/rfc/bcp22.txt,   ftp://ftp.
internic.net/rfc/rfc2360.txt,   ftp://ftp.math.utah.edu/pub/
rfc/bcp22.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2360.txt.

See also BCP0022 [?]. Status: BEST CURRENT PRACTICE.

RFC2361

[2348] E. Fleischman. RFC 2361: WAVE and AVI codec registries, June

1998. URL ftp://ftp.internic.net/rfc/rfc2361.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2361.txt. Status: INFORMA-

TIONAL.

RFC2362

[2349] D. Estrin, D. Farinacci, A. Helmy, D. Thaler, S. Deering, M. Handley, V. Jacobson, C. Liu, P. Sharma, and L. Wei. RFC 2362: Protocol independent multicast-sparse mode (PIM-SM): Protocol speci-

fication, June 1998. URL ftp://ftp.internic.net/rfc/rfc2117.
txt,   ftp://ftp.internic.net/rfc/rfc2362.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2117.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2362.txt. Obsoletes RFC2117 [2104]. Status: EXPERI-MENTAL.

RFC2363

[2350] G. Gross, M. Kaycee, A. Li, A. Malis, and J. Stephens. RFC 2363: PPP

over FUNI, July 1998. URL ftp://ftp.internic.net/rfc/rfc2363.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2363.txt. Status:

PROPOSED STANDARD.

RFC2364

[2351] G. Gross, M. Kaycee, A. Li, A. Malis, and J. Stephens. RFC 2364: PPP

over AAL5, July 1998. URL ftp://ftp.internic.net/rfc/rfc2364.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2364.txt. Status:

PROPOSED STANDARD.

RFC2365

[2352] D. Meyer. RFC 2365: Administratively scoped IP multicast, July

1998. URL ftp://ftp.internic.net/rfc/bcp23.txt,   ftp://ftp.
internic.net/rfc/rfc2365.txt,   ftp://ftp.math.utah.edu/pub/
rfc/bcp23.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2365.txt.

See also BCP0023 [?]. Status: BEST CURRENT PRACTICE.

RFC2366

[2353] C. Chung and M. Greene. RFC 2366: Definitions of managed objects for multicast over UNI 3.0/3.1 based ATM networks, July 1998. URL ftp://

ftp.internic.net/rfc/rfc2366.txt,   ftp://ftp.internic.net/
rfc/rfc2417.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2366.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2417.txt. Obsoleted

by RFC2417 [2404]. Status: PROPOSED STANDARD.

RFC2367

[2354] D. McDonald, C. Metz, and B. Phan. RFC 2367: PF KEY key management API, version 2, July 1998. URL ftp://ftp.internic.net/

rfc/rfc2367.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2367.

txt. Status: INFORMATIONAL.

RFC2368

[2355] P. Hoffman, L. Masinter, and J. Zawinski. RFC 2368: The mailto

URL scheme, July 1998. URL ftp://ftp.internic.net/rfc/rfc1738.
txt,   ftp://ftp.internic.net/rfc/rfc1808.txt,   ftp://ftp.
internic.net/rfc/rfc2368.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1738.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1808.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2368.txt. Updates

RFC1738, RFC1808 [1728, 1798]. Status: PROPOSED STANDARD.

RFC2369

[2356] G. Neufeld and J. Baer. RFC 2369: The use of URLs as meta-syntax for core mail list commands and their transport through message header

fields, July 1998. URL ftp://ftp.internic.net/rfc/rfc2369.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2369.txt. Status: PRO-

POSED STANDARD.

RFC2370

[2357] R. Coltun. RFC 2370: The OSPF opaque LSA option, July

1998. URL ftp://ftp.internic.net/rfc/rfc2328.txt,   ftp://
ftp.internic.net/rfc/rfc2370.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2328.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2370.

txt. See also RFC2328 [2315]. Status: PROPOSED STANDARD.

RFC2371

[2358] J. Lyon, K. Evans, and J. Klein. RFC 2371: Transaction Internet Protocol version 3.0, July 1998. URL ftp://ftp.internic.net/

rfc/rfc2371.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2371.

txt. Status: PROPOSED STANDARD.

RFC2372

[2359] K. Evans, J. Klein, and J. Lyon. RFC 2372: Transaction Internet Protocol — requirements and supplemental information, July

1998. URL ftp://ftp.internic.net/rfc/rfc2372.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2372.txt. Status: INFORMA-

TIONAL.

RFC2373

[2360] R. Hinden and S. Deering. RFC 2373: IP version 6 addressing ar-

chitecture, July 1998. URL ftp://ftp.internic.net/rfc/rfc1884.
txt,   ftp://ftp.internic.net/rfc/rfc2373.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1884.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2373.txt. Obsoletes RFC1884 [1871]. Status: PROPOSED STANDARD.

RFC2374

[2361] R. Hinden, M. O’Dell, and S. Deering. RFC 2374: An IPv6 aggregatable global unicast address format, July 1998. URL ftp://

ftp.internic.net/rfc/rfc2073.txt,   ftp://ftp.internic.net/
rfc/rfc2374.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2073.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2374.txt. Obsoletes

RFC2073 [2060]. Status: PROPOSED STANDARD.

RFC2375

[2362] R. Hinden and S. Deering. RFC 2375: IPv6 multicast address assign-

ments, July 1998. URL ftp://ftp.internic.net/rfc/rfc2375.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2375.txt. Status: INFOR-

MATIONAL.

RFC2376

[2363] E. Whitehead and M. Murata. RFC 2376: XML media types,

July 1998. URL ftp://ftp.internic.net/rfc/rfc2376.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2376.txt. Status: INFORMA-

TIONAL.

RFC2377

[2364] A. Grimstad, R. Huber, S. Sataluri, and M. Wahl. RFC 2377: Naming plan for Internet directory-enabled applications, Septem-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc2377.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2377.txt. Status: INFORMA-

TIONAL.

RFC2378

[2365] R. Hedberg and P. Pomes. RFC 2378: The CCSO nameserver (ph) archi-

tecture, August 1998. URL ftp://ftp.internic.net/rfc/rfc2378.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2378.txt. Status:

INFORMATIONAL.

RFC2379

[2366] L. Berger. RFC 2379: RSVP over ATM implementation guide-

lines, August 1998. URL ftp://ftp.internic.net/rfc/bcp24.
txt,   ftp://ftp.internic.net/rfc/rfc2379.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp24.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2379.txt. See also BCP0024 [?]. Status: Best Current Practice.

RFC2380

[2367] L. Berger. RFC 2380: RSVP over ATM implementation requirements,

August 1998. URL ftp://ftp.internic.net/rfc/rfc2380.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2380.txt. Status: PRO-

POSED STANDARD.

RFC2381

[2368] M. Garrett and M. Borden. RFC 2381: Interoperation of controlled-load service and guaranteed service with ATM, August 1998. URL ftp://

ftp.internic.net/rfc/rfc2381.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2381.txt. Status: PROPOSED STANDARD.

RFC2382

[2369] E. Crawley, L. Berger, S. Berson, F. Baker, M. Borden, and J. Krawczyk. RFC 2382: A framework for integrated services and RSVP over ATM,

August 1998. URL ftp://ftp.internic.net/rfc/rfc2382.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc2382.txt. Status: INFOR-

MATIONAL.

RFC2383

[2370] M. Suzuki. RFC 2383: ST2+ over ATM protocol specification — UNI 3.1

version, August 1998. URL ftp://ftp.internic.net/rfc/rfc2383.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2383.txt. Status:

INFORMATIONAL.

RFC2384

[2371] R. Gellens. RFC 2384: POP URL scheme, August 1998. URL ftp://

ftp.internic.net/rfc/rfc2384.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2384.txt. Status: PROPOSED STANDARD.

RFC2385

[2372] A. Heffernan. RFC 2385: Protection of BGP sessions via the TCP MD5 signature option, August 1998. URL ftp://ftp.internic.net/

rfc/rfc2385.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2385.

txt. Status: PROPOSED STANDARD.

RFC2386

[2373] E. Crawley, R. Nair, B. Rajagopalan, and H. Sandick. RFC 2386: A framework for QoS-based routing in the Internet, Au-

gust 1998. URL ftp://ftp.internic.net/rfc/rfc2386.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2386.txt. Status: INFORMA-

TIONAL.

RFC2387

[2374] E. Levinson. RFC 2387: The MIME multipart/related content-

type, August 1998. URL ftp://ftp.internic.net/rfc/rfc2112.
txt,   ftp://ftp.internic.net/rfc/rfc2387.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2112.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2387.txt. Obsoletes RFC2112 [2099]. Status: PROPOSED STANDARD.

RFC2388

[2375] L. Masinter. RFC 2388: Returning values from forms: multipart/form-

data, August 1998. URL ftp://ftp.internic.net/rfc/rfc2388.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2388.txt. Status: PRO-

POSED STANDARD.

RFC2389

[2376] P. Hethmon and R. Elz. RFC 2389: Feature negotiation mechanism for the file transfer protocol, August 1998. URL ftp://ftp.internic.net/

rfc/rfc2389.txt,   ftp://ftp.internic.net/rfc/rfc959.txt,

ftp://ftp.math.utah.edu/pub/rfc/rfc2389.txt,   ftp://ftp.math.

utah.edu/pub/rfc/rfc959.txt. See also RFC0959 [949]. Status: PRO-POSED STANDARD.

RFC2390

[2377] T. Bradley, C. Brown, and A. Malis. RFC 2390: Inverse address resolution protocol, August 1998. URL ftp://ftp.internic.net/rfc/

rfc1293.txt,   ftp://ftp.internic.net/rfc/rfc2390.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1293.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2390.txt. Obsoletes RFC1293 [1283]. Status: DRAFT STANDARD.

RFC2391

[2378] P. Srisuresh and D. Gan. RFC 2391: Load sharing using IP network address translation (LSNAT), August 1998. URL ftp://ftp.internic.

net/rfc/rfc2391.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2391.

txt. Status: INFORMATIONAL.

RFC2392

[2379] E. Levinson. RFC 2392: Content-ID and message-ID uniform resource

locators, August 1998. URL ftp://ftp.internic.net/rfc/rfc2111.
txt,   ftp://ftp.internic.net/rfc/rfc2392.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2111.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2392.txt. Obsoletes RFC2111 [2098]. Status: PROPOSED STANDARD.

RFC2393

[2380] A. Shacham, R. Monsour, R. Pereira, and M. Thomas. RFC 2393: IP Payload Compression Protocol (IPComp), December 1998. URL ftp://

ftp.internic.net/rfc/rfc2393.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2393.txt. Status: PROPOSED STANDARD.

RFC2394

[2381] R. Pereira. RFC 2394: IP payload compression using DEFLATE,

December 1998. URL ftp://ftp.internic.net/rfc/rfc2394.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2394.txt. Status: INFOR-

MATIONAL.

RFC2395

[2382] R. Friend and R. Monsour. RFC 2395: IP payload compression using

LZS, December 1998. URL ftp://ftp.internic.net/rfc/rfc2395.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2395.txt. Status:

INFORMATIONAL.

RFC2396

[2383] T. Berners-Lee, R. Fielding, and L. Masinter. RFC 2396: Uniform Resource Identifiers (URI): Generic syntax, August 1998. URL ftp://ftp.

internic.net/rfc/rfc2396.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2396.txt. Status: DRAFT STANDARD.

RFC2397

[2384] L. Masinter. RFC 2397: The “data” URL scheme, August

1998. URL ftp://ftp.internic.net/rfc/rfc2397.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2397.txt. Status: PROPOSED

STANDARD.

RFC2398

[2385] S. Parker and C. Schmechel. RFC 2398: Some testing tools for TCP im-

plementors, August 1998. URL ftp://ftp.internic.net/rfc/fyi33.
txt,   ftp://ftp.internic.net/rfc/rfc2398.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi33.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2398.txt. See also FYI0033 [?]. Status: INFORMA-TIONAL.

RFC2399

[2386] A. Ramos. RFC 2399: Request for comments summary RFC numbers 2300–2399, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2399.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2399.

txt. Status: INFORMATIONAL.

RFC2400

[2387] J. Postel and J. Reynolds. RFC 2400: INTERNET OFFICIAL PROTOCOL STANDARDS, September 1998. URL ftp://ftp.

internic.net/rfc/rfc1083.txt,   ftp://ftp.internic.net/rfc/
rfc1100.txt,   ftp://ftp.internic.net/rfc/rfc1130.txt,   ftp:
//ftp.internic.net/rfc/rfc1140.txt,   ftp://ftp.internic.net/
rfc/rfc1200.txt,   ftp://ftp.internic.net/rfc/rfc1250.txt,
ftp://ftp.internic.net/rfc/rfc1280.txt,   ftp://ftp.internic.
net/rfc/rfc1360.txt,   ftp://ftp.internic.net/rfc/rfc1410.txt,
 ftp://ftp.internic.net/rfc/rfc1500.txt,   ftp://ftp.internic.
net/rfc/rfc1540.txt,   ftp://ftp.internic.net/rfc/rfc1600.txt,
 ftp://ftp.internic.net/rfc/rfc1610.txt,   ftp://ftp.internic.
net/rfc/rfc1720.txt,   ftp://ftp.internic.net/rfc/rfc1780.txt,
 ftp://ftp.internic.net/rfc/rfc1800.txt,   ftp://ftp.internic.
net/rfc/rfc1880.txt,   ftp://ftp.internic.net/rfc/rfc1920.txt,
 ftp://ftp.internic.net/rfc/rfc2000.txt,   ftp://ftp.internic.
net/rfc/rfc2200.txt,   ftp://ftp.internic.net/rfc/rfc2300.txt,

 ftp://ftp.internic.net/rfc/rfc2400.txt, ftp://ftp.internic.
net/rfc/std1.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1083.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1100.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1130.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1140.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1200.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1250.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1280.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1360.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1410.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1500.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1540.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc1600.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc1610.txt,   ftp://ftp.math.utah.edu/pub/rfc/
rfc1720.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1780.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1800.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc1880.txt,   ftp://ftp.math.utah.edu/pub/
rfc/rfc1920.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2000.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2200.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2300.txt,   ftp://ftp.math.utah.
edu/pub/rfc/rfc2400.txt,   ftp://ftp.math.utah.edu/pub/rfc/

std1.txt. See also STD0001 [?]. Obsoletes RFC2300, RFC2200, RFC2000, RFC1920, RFC1880, RFC1800, RFC1780, RFC1720, RFC1610, RFC1600, RFC1540, RFC1500, RFC1410, RFC1360, RFC1280, RFC1250, RFC1200, RFC1140, RFC1130, RFC1100, RFC1083 [2287, 2187, 1987, 1907, 1867, 1790, 1770, 1710, 1600, 1590, 1530, 1490, 1400, 1350, 1270, 1240, 1190, 1130, 1120, 1090, 1073]. Status: STANDARD.

RFC2401

[2388] S. Kent and R. Atkinson. RFC 2401: Security architecture for the Internet Protocol, November 1998. URL ftp://ftp.

internic.net/rfc/rfc1825.txt,   ftp://ftp.internic.net/rfc/
rfc2401.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1825.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2401.txt. Obsoletes RFC1825

[1815]. Status: PROPOSED STANDARD.

RFC2402

[2389] S. Kent and R. Atkinson. RFC 2402: IP authentication header, Novem-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc1826.txt,   ftp:/
/ftp.internic.net/rfc/rfc2402.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1826.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2402.

txt. Obsoletes RFC1826 [1816]. Status: PROPOSED STANDARD.

RFC2403

[2390] C. Madson and R. Glenn. RFC 2403: The use of HMAC-MD5-96 within ESP and AH, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2403.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2403.

txt. Status: PROPOSED STANDARD.

RFC2404

[2391] C. Madson and R. Glenn. RFC 2404: The use of HMAC-SHA-1-96 within ESP and AH, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2404.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2404.

txt. Status: PROPOSED STANDARD.

RFC2405

[2392] C. Madson and N. Doraswamy. RFC 2405: The ESP DES-CBC cipher algorithm with explicit IV, November 1998. URL ftp://ftp.internic.

net/rfc/rfc2405.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2405.

txt. Status: PROPOSED STANDARD.

RFC2406

[2393] S. Kent and R. Atkinson. RFC 2406: IP Encapsulating Security Payload

(ESP), November 1998. URL ftp://ftp.internic.net/rfc/rfc1827.
txt,   ftp://ftp.internic.net/rfc/rfc2406.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1827.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2406.txt. Obsoletes RFC1827 [1817]. Status: PROPOSED STANDARD.

RFC2407

[2394] D. Piper. RFC 2407: The Internet IP security domain of interpretation for ISAKMP, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2407.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2407.

txt. Status: PROPOSED STANDARD.

RFC2408

[2395] D. Maughan, M. Schertler, M. Schneider, and J. Turner. RFC 2408: Internet Security Association and Key Management Protocol (ISAKMP),

November 1998. URL ftp://ftp.internic.net/rfc/rfc2408.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2408.txt. Status: PRO-

POSED STANDARD.

RFC2409

[2396] D. Harkins and D. Carrel. RFC 2409: The Internet Key Exchange (IKE),

November 1998. URL ftp://ftp.internic.net/rfc/rfc2409.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2409.txt. Status: PRO-

POSED STANDARD.

RFC2410

[2397] R. Glenn and S. Kent. RFC 2410: The NULL encryption algorithm and its use with IPsec, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2410.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2410.

txt. Status: PROPOSED STANDARD.

RFC2411

[2398] R. Thayer, N. Doraswamy, and R. Glenn. RFC 2411: IP security document roadmap, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2411.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2411.

txt. Status: INFORMATIONAL.

RFC2412

[2399] H. Orman. RFC 2412: The OAKLEY Key Determination Protocol,

November 1998. URL ftp://ftp.internic.net/rfc/rfc2412.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2412.txt. Status: INFOR-

MATIONAL.

RFC2413

[2400] S. Weibel, J. Kunze, C. Lagoze, and M. Wolf. RFC 2413: Dublin core metadata for resource discovery, September 1998. URL ftp://ftp.

internic.net/rfc/rfc2413.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2413.txt. Status: INFORMATIONAL.

RFC2414

[2401] M. Allman, S. Floyd, and C. Partridge. RFC 2414: Increasing TCP’s initial window, September 1998. URL ftp://ftp.internic.net/

rfc/rfc2414.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2414.

txt. Status: EXPERIMENTAL.

RFC2415

[2402] K. Poduri and K. Nichols. RFC 2415: Simulation studies of increased initial TCP window size, September 1998. URL ftp://ftp.internic.net/

rfc/rfc2415.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2415.

txt. Status: INFORMATIONAL.

RFC2416

[2403] T. Shepard and C. Partridge. RFC 2416: When TCP starts up with four packets into only three buffers, September 1998. URL ftp://ftp.

internic.net/rfc/rfc2416.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2416.txt. Status: INFORMATIONAL.

RFC2417

[2404] C. Chung and M. Greene. RFC 2417: Definitions of managed objects for multicast over UNI 3.0/3. 1 based atm networks,

September 1998.  URL ftp://ftp.internic.net/rfc/rfc2366.
txt,   ftp://ftp.internic.net/rfc/rfc2417.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2366.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2417.txt. Obsoletes RFC2366 [2353]. Status: PROPOSED STANDARD.

RFC2418

[2405] S. Bradner. RFC 2418: IETF working group guidelines and proce-

dures, September 1998. URL ftp://ftp.internic.net/rfc/bcp25.
txt,   ftp://ftp.internic.net/rfc/rfc1603.txt,   ftp://ftp.
internic.net/rfc/rfc2418.txt,   ftp://ftp.math.utah.edu/pub/
rfc/bcp25.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1603.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2418.txt. See also BCP0025

[?]. Obsoletes RFC1603 [1593]. Status: BEST CURRENT PRACTICE.

RFC2419

[2406] K. Sklower and G. Meyer. RFC 2419: The PPP DES encryption protocol, version 2 (DESE-bis), September 1998. URL ftp://

ftp.internic.net/rfc/rfc1969.txt,   ftp://ftp.internic.net/
rfc/rfc2419.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1969.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2419.txt. Obsoletes

RFC1969 [1956]. Status: PROPOSED STANDARD.

RFC2420

[2407] H. Kummert. RFC 2420: The PPP triple-DES encryption protocol (3DESE), September 1998. URL ftp://ftp.internic.net/

rfc/rfc2420.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2420.

txt. Status: PROPOSED STANDARD.

RFC2421

[2408] G. Vaudreuil and G. Parsons. RFC 2421: Voice profile for Internet mail — version 2, September 1998. URL ftp://ftp.

internic.net/rfc/rfc1911.txt,   ftp://ftp.internic.net/rfc/
rfc2421.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1911.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2421.txt. Obsoletes RFC1911

[1898]. Status: PROPOSED STANDARD.

RFC2422

[2409] G. Vaudreuil and G. Parsons. RFC 2422: Toll quality voice — 32 kbit/s ADPCM MIME sub-type registration, September 1998. URL ftp://

ftp.internic.net/rfc/rfc1911.txt,   ftp://ftp.internic.net/
rfc/rfc2422.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1911.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2422.txt. Obsoletes

RFC1911 [1898]. Status: PROPOSED STANDARD.

RFC2423

[2410] G. Vaudreuil and G. Parsons. RFC 2423: VPIM voice message MIME sub-type registration, September 1998. URL ftp://

ftp.internic.net/rfc/rfc1911.txt,   ftp://ftp.internic.net/
rfc/rfc2423.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1911.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2423.txt. Obsoletes

RFC1911 [1898]. Status: PROPOSED STANDARD.

RFC2424

[2411] G. Vaudreuil and G. Parsons. RFC 2424: Content duration MIME header definition, September 1998. URL ftp://ftp.internic.net/

rfc/rfc2424.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2424.

txt. Status: PROPOSED STANDARD.

RFC2425

[2412] T. Howes, M. Smith, and F. Dawson. RFC 2425: A MIME content-type for directory information, September 1998. URL ftp://ftp.internic.

net/rfc/rfc2425.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2425.

txt. Status: PROPOSED STANDARD.

RFC2426

[2413] F. Dawson and T. Howes. RFC 2426: vCard MIME directory profile,

September 1998. URL ftp://ftp.internic.net/rfc/rfc2426.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2426.txt. Status: PRO-

POSED STANDARD.

RFC2427

[2414] C. Brown and A. Malis. RFC 2427: Multiprotocol interconnect over frame relay, September 1998. URL ftp://ftp.internic.net/rfc/

rfc1294.txt,   ftp://ftp.internic.net/rfc/rfc1490.txt,   ftp:
//ftp.internic.net/rfc/rfc2427.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1294.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1490.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2427.txt. Obsoletes

RFC1490, RFC1294 [1480, 1284]. See also STD0055 [?]. Status: STAN-DARD.

RFC2428

[2415] M. Allman, S. Ostermann, and C. Metz. RFC 2428: FTP extensions for IPv6 and NATs, September 1998. URL ftp://ftp.internic.net/

rfc/rfc2428.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2428.

txt. Status: PROPOSED STANDARD.

RFC2429

[2416] C. Bormann, L. Cline, G. Deisher, T. Gardos, C. Maciocco, D. Newell, J. Ott, G. Sullivan, S. Wenger, and C. Zhu. RFC 2429: RTP payload format for the 1998 version of ITU-T rec, Octo-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc2429.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2429.txt. Status: PROPOSED

STANDARD.

RFC2430

[2417] T. Li and Y. Rekhter. RFC 2430: RA provider architecture for differentiated services and traffic engineering (PASTE), Octo-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc2430.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2430.txt. Status: INFORMA-

TIONAL.

RFC2431

[2418] D. Tynan. RFC 2431: RTP payload format for BT.656 video encod-

ing, October 1998. URL ftp://ftp.internic.net/rfc/rfc2431.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2431.txt. Status: PRO-

POSED STANDARD.

RFC2432

[2419] K. Dubray. RFC 2432: Terminology for IP multicast benchmarking,

October 1998. URL ftp://ftp.internic.net/rfc/rfc2432.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2432.txt. Status: INFOR-

MATIONAL.

RFC2433

[2420] G. Zorn and S. Cobb. RFC 2433: Microsoft PPP CHAP extensions,

October 1998. URL ftp://ftp.internic.net/rfc/rfc2433.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2433.txt. Status: INFOR-

MATIONAL.

RFC2434

[2421] T. Narten and H. Alvestrand. RFC 2434: Guidelines for writing an IANA considerations section in RFCs, October 1998. URL ftp:

//ftp.internic.net/rfc/bcp26.txt,   ftp://ftp.internic.net/
rfc/rfc2434.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp26.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2434.txt. See also BCP0026

[?]. Status: BEST CURRENT PRACTICE.

RFC2435

[2422] L. Berc, W. Fenner, R. Frederick, S. McCanne, and P. Stewart. RFC 2435: RTP payload format for JPEG-compressed video, October

1998. URL ftp://ftp.internic.net/rfc/rfc2035.txt,   ftp://
ftp.internic.net/rfc/rfc2435.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2035.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2435.

txt. Obsoletes RFC2035 [2022]. Status: PROPOSED STANDARD.

RFC2436

[2423] R. Brett, S. Bradner, and G. Parsons. RFC 2436: Collaboration between ISOC/IETF and ITU-T, October 1998. URL ftp://ftp.internic.

net/rfc/rfc2436.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2436.

txt. Status: INFORMATIONAL.

RFC2437

[2424] B. Kaliski and J. Staddon. RFC 2437: PKCS #1: RSA cryptography specifications version 2.0, October 1998. URL ftp://

ftp.internic.net/rfc/rfc2313.txt,   ftp://ftp.internic.net/
rfc/rfc2437.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2313.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2437.txt. Obsoletes

RFC2313 [2300]. Status: INFORMATIONAL.

RFC2438

[2425] M. O’Dell, H. Alvestrand, B. Wijnen, and S. Bradner. RFC 2438: Advancement of MIB specifications on the IETF standards

track, October 1998. URL ftp://ftp.internic.net/rfc/bcp27.
txt,   ftp://ftp.internic.net/rfc/rfc2438.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp27.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2438.txt. See also BCP0027 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2439

[2426] C. Villamizar, R. Chandra, and R. Govindan. RFC 2439: BGP route flap damping, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2439.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2439.

txt. Status: PROPOSED STANDARD.

RFC2440

[2427] J. Callas, L. Donnerhacke, H. Finney, and R. Thayer. RFC 2440: OpenPGP message format, November 1998. URL ftp://ftp.internic.

net/rfc/rfc2440.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2440.

txt. Status: PROPOSED STANDARD.

RFC2441

[2428] D. Cohen. RFC 2441: Working with Jon, tribute delivered at UCLA, October 30, 1998, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2441.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2441.

txt. Status: INFORMATIONAL. See also RFC2441 [2428, 2455, 2503, 2504].

RFC2442

[2429] N. Freed, D. Newman, J. Belissen, and M. Hoy. RFC 2442: The batch SMTP media type, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2442.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2442.

txt. Status: INFORMATIONAL.

RFC2443

[2430] J. Luciani and A. Gallo. RFC 2443: A distributed MARS service using

SCSP, November 1998. URL ftp://ftp.internic.net/rfc/rfc2443.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2443.txt. Status:

EXPERIMENTAL.

RFC2444

[2431] C. Newman. RFC 2444: The one-time-password SASL mecha-

nism, October 1998. URL ftp://ftp.internic.net/rfc/rfc2222.
txt,   ftp://ftp.internic.net/rfc/rfc2444.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2222.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2444.txt. Updates RFC2222 [2209]. Status: PROPOSED STANDARD.

RFC2445

[2432] F. Dawson and D. Stenerson. RFC 2445: Internet calendaring and scheduling core object specification (iCalendar), November

1998. URL ftp://ftp.internic.net/rfc/rfc2445.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2445.txt. Status: PROPOSED

STANDARD.

RFC2446

[2433] S. Silverberg, S. Mansour, F. Dawson, and R. Hopson. RFC 2446: iCalendar Transport-Independent Interoperability Protocol (iTIP) scheduling events, BusyTime, to-dos and journal entries, Novem-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc2446.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2446.txt. Status: PROPOSED

STANDARD.

RFC2447

[2434] F. Dawson, S. Mansour, and S. Silverberg. RFC 2447: iCalendar Message-Based Interoperability Protocol (iMIP), November

1998. URL ftp://ftp.internic.net/rfc/rfc2447.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2447.txt. Status: PROPOSED

STANDARD.

RFC2448

[2435] M. Civanlar, G. Cash, and B. Haskell. RFC 2448: AT&T’s error resilient video transmission technique, November 1998. URL ftp://ftp.

internic.net/rfc/rfc2448.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2448.txt. Status: INFORMATIONAL.

RFC2449

[2436] R. Gellens, C. Newman, and L. Lundblade. RFC 2449: POP3 extension mechanism, November 1998. URL ftp://ftp.internic.net/rfc/

rfc1939.txt,   ftp://ftp.internic.net/rfc/rfc2449.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1939.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2449.txt. Updates RFC1939 [1926]. Status: PRO-POSED STANDARD.

RFC2450

[2437] R. Hinden. RFC 2450: Proposed TLA and NLA assignment rule,

December 1998. URL ftp://ftp.internic.net/rfc/rfc2450.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2450.txt. Status: INFOR-

MATIONAL.

RFC2451

[2438] R. Pereira and R. Adams. RFC 2451: The ESP CBC-Mode cipher algo-

rithms, November 1998. URL ftp://ftp.internic.net/rfc/rfc2451.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2451.txt. Status:

PROPOSED STANDARD.

RFC2452

[2439] M. Daniele. RFC 2452: IP Version 6 Management Information Base for the Transmission Control Protocol, December 1998. URL ftp://ftp.

internic.net/rfc/rfc2452.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2452.txt. Status: PROPOSED STANDARD.

RFC2453

[2440] G. Malkin. RFC 2453: RIP version 2, November 1998. URL ftp://ftp.

internic.net/rfc/rfc1388.txt,   ftp://ftp.internic.net/rfc/
rfc1723.txt,   ftp://ftp.internic.net/rfc/rfc2453.txt, ftp:

//ftp.internic.net/rfc/std56.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1388.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1723.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2453.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/std56.txt. See also STD0056 [?]. Ob-

soletes RFC1388, RFC1723 [1378, 1713]. Status: STANDARD.

RFC2454

[2441] M. Daniele. RFC 2454: IP Version 6 Management Information Base for the User Datagram Protocol, December 1998. URL ftp://ftp.

internic.net/rfc/rfc2454.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2454.txt. Status: PROPOSED STANDARD.

RFC2455

[2442] B. Clouston and B. Moore. RFC 2455: Definitions of managed objects for

APPN, November 1998. URL ftp://ftp.internic.net/rfc/rfc2155.
txt,   ftp://ftp.internic.net/rfc/rfc2455.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2155.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2455.txt. Obsoletes RFC2155 [2142]. Status: PROPOSED STANDARD.

RFC2456

[2443] B. Clouston and B. Moore. RFC 2456: Definitions of managed objects for APPN TRAPS, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2456.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2456.

txt. Status: PROPOSED STANDARD.

RFC2457

[2444] B. Clouston and B. Moore. RFC 2457: Definitions of managed objects for extended border node, November 1998. URL ftp://ftp.internic.net/

rfc/rfc2457.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2457.

txt. Status: PROPOSED STANDARD.

RFC2458

[2445] H. Lu, M. Krishnaswamy, L. Conroy, S. Bellovin, F. Burg, A. DeSimone, K. Tewani, P. Davidson, H. Schulzrinne, and K. Vishwanathan. RFC 2458: Toward the PSTN/Internet inter-networking — pre-PINT implementations, November 1998. URL ftp://ftp.internic.net/rfc/

rfc2458.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2458.txt.

RFC2459

[2446] R. Housley, W. Ford, W. Polk, and D. Solo. RFC 2459: Internet X.509 public key infrastructure certificate and CRL profile, Jan-

uary 1999. URL ftp://ftp.internic.net/rfc/rfc2459.txt,   ftp:

//ftp.math.utah.edu/pub/rfc/rfc2459.txt. Status: PROPOSED

STANDARD.

RFC2460

[2447] S. Deering and R. Hinden. RFC 2460: Internet Protocol, Version 6 (IPv6) specification, December 1998. URL ftp://ftp.

internic.net/rfc/rfc1883.txt,   ftp://ftp.internic.net/rfc/
rfc2460.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1883.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2460.txt. Obsoletes RFC1883

[1870]. Status: DRAFT STANDARD.

RFC2461

[2448] T. Narten, E. Nordmark, and W. Simpson. RFC 2461: Neighbor discovery for IP Version 6 (IPv6), December 1998. URL ftp://

ftp.internic.net/rfc/rfc1970.txt,   ftp://ftp.internic.net/
rfc/rfc2461.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1970.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2461.txt. Obsoletes

RFC1970 [1957]. Status: DRAFT STANDARD.

RFC2462

[2449] S. Thomson and T. Narten. RFC 2462: IPv6 stateless address autoconfiguration, December 1998. URL ftp://ftp.internic.net/rfc/

rfc1971.txt,   ftp://ftp.internic.net/rfc/rfc2462.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1971.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2462.txt. Obsoletes RFC1971 [1958]. Status: DRAFT STANDARD.

RFC2463

[2450] A. Conta and S. Deering. RFC 2463: Internet Control Message Protocol (ICMPv6) for the Internet Protocol Version 6 (IPv6) Specifica-

tion, December 1998. URL ftp://ftp.internic.net/rfc/rfc1885.
txt,   ftp://ftp.internic.net/rfc/rfc2463.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc1885.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2463.txt. Obsoletes RFC1885 [1872]. Status: DRAFT STANDARD.

RFC2464

[2451] M. Crawford. RFC 2464: Transmission of IPv6 packets over Ethernet networks, December 1998. URL ftp://ftp.internic.net/rfc/

rfc1972.txt,   ftp://ftp.internic.net/rfc/rfc2464.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc1972.txt,   ftp://ftp.math.utah.

edu/pub/rfc/rfc2464.txt. Obsoletes RFC1972 [1959]. Status: PRO-POSED STANDARD.

RFC2465

[2452] D. Haskin and S. Onishi. RFC 2465: Management Information Base for IP Version 6: Textual conventions and general group, Decem-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc2465.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2465.txt. Status: PROPOSED

STANDARD.

RFC2466

[2453] D. Haskin and S. Onishi. RFC 2466: Management Information Base for IP Version 6: ICMPv6 group, December 1998. URL ftp://ftp.

internic.net/rfc/rfc2466.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2466.txt. Status: PROPOSED STANDARD.

RFC2467

[2454] M. Crawford. RFC 2467: Transmission of IPv6 packets over FDDI net-

works, December 1998. URL ftp://ftp.internic.net/rfc/rfc2019.
txt,   ftp://ftp.internic.net/rfc/rfc2467.txt,   ftp://ftp.
math.utah.edu/pub/rfc/rfc2019.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2467.txt. Obsoletes RFC2019 [2006]. Status: PROPOSED STANDARD.

RFC2468

[2455] V. Cerf. RFC 2468: I remember IANA, October 17, 1998. URL ftp://

ftp.internic.net/rfc/rfc2468.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2468.txt. Status: INFORMATIONAL. This obituary tribute to Jon Postel was published in [2503]; see also [2428, 2504].

RFC2469

[2456] T. Narten and C. Burton. RFC 2469: A caution on the canonical ordering of link-layer addresses, December 1998. URL ftp://ftp.internic.net/

rfc/rfc2469.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2469.

txt. Status: INFORMATIONAL.

RFC2470

[2457] M. Crawford, T. Narten, and S. Thomas. RFC 2470: Transmission of IPv6 packets over token ring networks, December 1998. URL ftp://ftp.

internic.net/rfc/rfc2470.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2470.txt. Status: PROPOSED STANDARD.

RFC2471

[2458] R. Hinden, R. Fink, and J. Postel. RFC 2471: IPv6 testing address allo-

cation, December 1998. URL ftp://ftp.internic.net/rfc/rfc1897.
txt,   ftp://ftp.internic.net/rfc/rfc2471.txt,   ftp://ftp.

math.utah.edu/pub/rfc/rfc1897.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2471.txt. Obsoletes RFC1897 [1884]. Status: EXPERI-MENTAL.

RFC2472

[2459] D. Haskin and E. Allen. RFC 2472: IP version 6 over PPP, December

1998. URL ftp://ftp.internic.net/rfc/rfc2023.txt,   ftp://
ftp.internic.net/rfc/rfc2472.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc2023.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2472.

txt. Obsoletes RFC2023 [2010]. Status: PROPOSED STANDARD.

RFC2473

[2460] A. Conta and S. Deering. RFC 2473: Generic packet tunneling in IPv6 specification, December 1998. URL ftp://ftp.internic.net/

rfc/rfc2473.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2473.

txt. Status: PROPOSED STANDARD.

RFC2474

[2461] K. Nichols, S. Blake, F. Baker, and D. Black. RFC 2474: Definition of the Differentiated Services Field (DS Field) in the IPv4 and IPv6 headers,

December 1998. URL ftp://ftp.internic.net/rfc/rfc2474.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2474.txt.

RFC2475

[2462] S. Blake, D. Black, M. Carlson, E. Davies, Z. Wang, and W. Weiss. RFC 2475: An architecture for differentiated services, December

1998. URL ftp://ftp.internic.net/rfc/rfc2475.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2475.txt. Status: PROPOSED

STANDARD.

RFC2476

[2463] R. Gellens and J. Klensin. RFC 2476: Message submission, Decem-

ber 1998. URL ftp://ftp.internic.net/rfc/rfc2476.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2476.txt. Status: PROPOSED

STANDARD.

RFC2477

[2464] B. Aboba and G. Zorn. RFC 2477: Criteria for evaluating roaming pro-

tocols, December 1998. URL ftp://ftp.internic.net/rfc/rfc2477.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2477.txt. Status:

INFORMATIONAL.

RFC2478

[2465] E. Baize and D. Pinkas. RFC 2478: The simple and protected GSS-API negotiation mechanism, December 1998. URL ftp://ftp.internic.

net/rfc/rfc2478.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2478.

txt. Status: PROPOSED STANDARD.

RFC2479

[2466] C. Adams. RFC 2479: Independent Data Unit Protection Generic Security Service Application Program Interface (IDUP-GSS-API), De-

cember 1998. URL ftp://ftp.internic.net/rfc/rfc2479.txt,
ftp://ftp.math.utah.edu/pub/rfc/rfc2479.txt. Status: INFOR-

MATIONAL.

RFC2480

[2467] N. Freed. RFC 2480: Gateways and MIME security multiparts, Jan-

uary 1999. URL ftp://ftp.internic.net/rfc/rfc2480.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2480.txt. Status: PROPOSED

STANDARD.

RFC2481

[2468] K. Ramakrishnan and S. Floyd. RFC 2481: A proposal to add Explicit Congestion Notification (ECN) to IP, January 1999. URL ftp://ftp.

internic.net/rfc/rfc1570.txt,   ftp://ftp.internic.net/rfc/
rfc1994.txt,   ftp://ftp.internic.net/rfc/rfc2284.txt,   ftp:
//ftp.internic.net/rfc/rfc2481.txt,   ftp://ftp.math.utah.edu/
pub/rfc/rfc1570.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1994.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2284.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc2481.txt.  Updates RFC2284,

RFC1994, RFC1570 [2271, 1981, 1560]. Status: EXPERIMENTAL.

RFC2482

[2469] K. Whistler and G. Adams. RFC 2482: Language tagging in Unicode plain text, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2482.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2482.

txt. Status: INFORMATIONAL.

RFC2483

[2470] M. Mealling and R. Daniel, Jr. RFC 2483: URI resolution services necessary for URN resolution, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2483.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2483.

txt. Status: EXPERIMENTAL.

RFC2484

[2471] G. Zorn. RFC 2484: PPP LCP internationalization configuration op-

tion, January 1999. URL ftp://ftp.internic.net/rfc/rfc1570.
txt,   ftp://ftp.internic.net/rfc/rfc1994.txt,   ftp://ftp.
internic.net/rfc/rfc2284.txt,   ftp://ftp.internic.net/rfc/
rfc2484.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc1570.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc1994.txt,   ftp://ftp.math.
utah.edu/pub/rfc/rfc2284.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2484.txt. Updates RFC2284, RFC1994, RFC1570 [2271, 1981, 1560]. Status: PROPOSED STANDARD.

RFC2485

[2472] S. Drach. RFC 2485: DHCP option for the Open Group’s user authentication protocol, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2485.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2485.

txt. Status: PROPOSED STANDARD.

RFC2486

[2473] B. Aboba and M. Beadles. RFC 2486: The network access identifier, Jan-

uary 1999. URL ftp://ftp.internic.net/rfc/rfc2486.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2486.txt. Status: PROPOSED

STANDARD.

RFC2487

[2474] P. Hoffman. RFC 2487: SMTP service extension for secure SMTP over

TLS, January 1999. URL ftp://ftp.internic.net/rfc/rfc2487.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2487.txt. Status:

PROPOSED STANDARD.

RFC2488

[2475] M. Allman, D. Glover, and L. Sanchez. RFC 2488: Enhancing TCP over satellite channels using standard mechanisms, January 1999. URL ftp:

//ftp.internic.net/rfc/bcp28.txt,   ftp://ftp.internic.net/
rfc/rfc2488.txt,   ftp://ftp.math.utah.edu/pub/rfc/bcp28.txt,
 ftp://ftp.math.utah.edu/pub/rfc/rfc2488.txt. See also BCP0028

[?]. Status: Best Current Practice.

RFC2489

[2476] R. Droms. RFC 2489: Procedure for defining new DHCP op-

tions, January 1999. URL ftp://ftp.internic.net/rfc/bcp29.
txt,   ftp://ftp.internic.net/rfc/rfc2489.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp29.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2489.txt. See also BCP0029 [?]. Status: Best Current Practice.

RFC2490

[2477] M. Pullen, R. Malghan, L. Lavu, G. Duan, J. Ma, and H. Nah. RFC 2490: A simulation model for IP multicast with RSVP, Jan-

uary 1999. URL ftp://ftp.internic.net/rfc/rfc2490.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2490.txt. Status: INFORMA-

TIONAL.

RFC2491

[2478] G. Armitage, P. Schulter, M. Jork, and G. Harter. RFC 2491: IPv6 over non-broadcast multiple access (NBMA) networks, Jan-

uary 1999. URL ftp://ftp.internic.net/rfc/rfc2491.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2491.txt. Status: PROPOSED

STANDARD.

RFC2492

[2479] G. Armitage, P. Schulter, and M. Jork. RFC 2492: IPv6 over ATM net-

works, January 1999. URL ftp://ftp.internic.net/rfc/rfc2492.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2492.txt. Status:

PROPOSED STANDARD.

RFC2493

[2480] RFC 2493: Textual conventions for MIB modules using performance history based on 15 minute intervals, January 1999. URL ftp://ftp.

internic.net/rfc/rfc2493.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2493.txt. Status: PROPOSED STANDARD.

RFC2494

[2481] RFC 2494: Definitions of managed objects for the DS0 and DS0 bundle interface type, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2494.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2494.

txt. Status: PROPOSED STANDARD.

RFC2495

[2482] RFC 2495: Definitions of managed objects for the DS1, E1, DS2 and E2 interface types, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2495.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2495.

txt. Status: PROPOSED STANDARD.

RFC2496

[2483] RFC 2496: Definitions of managed objects for the DS3/E3 interface type,

January 1999. URL ftp://ftp.internic.net/rfc/rfc2496.txt,

 ftp://ftp.math.utah.edu/pub/rfc/rfc2496.txt. Status: PRO-

POSED STANDARD.

RFC2497

[2484] I. Souvatzis. RFC 2497: Transmission of IPv6 packets over ARCnet net-

works, January 1999. URL ftp://ftp.internic.net/rfc/rfc2497.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2497.txt. Status:

PROPOSED STANDARD.

RFC2498

[2485] J. Mahdavi and V. Paxson. RFC 2498: IPPM metrics for measuring connectivity, January 1999. URL ftp://ftp.internic.net/

rfc/rfc2498.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2498.

txt. Status: EXPERIMENTAL.

RFC2502

[2486] M. Pullen, M. Myjak, and C. Bouwens. RFC 2502: Limitations of Internet Protocol suite for distributed simulation the large multicast environ-

ment, February 1999. URL ftp://ftp.internic.net/rfc/rfc2502.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2502.txt. Status:

INFORMATIONAL.

RFC2503

[2487] R. Moulton and M. Needleman. RFC 2503: MIME types for use with the ISO ILL protocol, February 1999. URL ftp://ftp.internic.net/

rfc/rfc2503.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2503.

txt. Status: INFORMATIONAL.

RFC2504

[2488] E. Guttman, L. Leong, and G. Malkin. RFC 2504: Users’ security

handbook, February 1999. URL ftp://ftp.internic.net/rfc/fyi34.
txt,   ftp://ftp.internic.net/rfc/rfc2504.txt,   ftp://ftp.
math.utah.edu/pub/rfc/fyi34.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2504.txt. See also FYI0034 [?]. Status: INFORMA-TIONAL.

RFC2505

[2489] G. Lindberg. RFC 2505: Anti-spam recommendations for SMTP

MTAs, February 1999. URL ftp://ftp.internic.net/rfc/bcp30.
txt,   ftp://ftp.internic.net/rfc/rfc2505.txt,   ftp://ftp.
math.utah.edu/pub/rfc/bcp30.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc2505.txt. See also BCP0030 [?]. Status: BEST CUR-RENT PRACTICE.

RFC2507

[2490] M. Degermark, B. Nordgren, and S. Pink. RFC 2507: IP header com-

pression, February 1999. URL ftp://ftp.internic.net/rfc/rfc2507.
txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2507.txt. Status:

PROPOSED STANDARD.

RFC2508

[2491] S. Casner and V. Jacobson. RFC 2508: Compressing IP/UDP/RTP headers for low-speed serial links, February 1999. URL ftp://ftp.

internic.net/rfc/rfc2508.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2508.txt. Status: PROPOSED STANDARD.

RFC2509

[2492] M. Engan, S. Casner, and C. Bormann. RFC 2509: IP header compression over PPP, February 1999. URL ftp://ftp.internic.net/

rfc/rfc2509.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2509.

txt. Status: PROPOSED STANDARD.

RFC2513

[2493] K. McCloghrie, J. Heinanen, W. Greene, and A. Prasad. RFC 2513: Managed objects for controlling the collection and storage of accounting information for connection-oriented networks, Febru-

ary 1999. URL ftp://ftp.internic.net/rfc/rfc2513.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2513.txt. Status: PROPOSED

STANDARD.

RFC2514

[2494] M. Noto, E. Spiegel, and K. Tesink. RFC 2514: Definitions of textual conventions and OBJECT-IDENTITIES for ATM management, Febru-

ary 1999. URL ftp://ftp.internic.net/rfc/rfc2514.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2514.txt. Status: PROPOSED

STANDARD.

RFC2515

[2495] K. Tesink and Ed. RFC 2515: Definitions of managed objects for ATM management, February 1999. URL ftp://ftp.internic.net/

rfc/rfc2515.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2515.

txt. Status: PROPOSED STANDARD.

RFC2517

[2496] R. Moats and R. Huber. RFC 2517: Building directories from DNS: Experiences from WWWSeeker, February 1999. URL ftp://ftp.

internic.net/rfc/rfc2517.txt,   ftp://ftp.math.utah.edu/pub/

rfc/rfc2517.txt. Status: INFORMATIONAL.

RFC2518

[2497] Y. Goland, E. Whitehead, A. Faizi, S. Carter, and D. Jensen. RFC 2518: HTTP extensions for distributed authoring — WEBDAV, Febru-

ary 1999. URL ftp://ftp.internic.net/rfc/rfc2518.txt,   ftp:
//ftp.math.utah.edu/pub/rfc/rfc2518.txt. Status: PROPOSED

STANDARD.

RFC2519

[2498] E. Chen and J. Stewart. RFC 2519: A framework for inter-domain route aggregation, February 1999. URL ftp://ftp.internic.net/

rfc/rfc2519.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2519.

txt. Status: INFORMATIONAL.

RFC2520

[2499] J. Luciani, H. Suzuki, N. Doraswamy, and D. Horton. RFC 2520: NHRP with mobile NHCs, February 1999. URL ftp://ftp.internic.net/

rfc/rfc2520.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc2520.

txt. Status: EXPERIMENTAL.

RFC3492

[2500] A. Costello. RFC 3492: Punycode: A bootstring encoding of Unicode for Internationalized Domain Names in Applications (IDNA), March

2003. URL ftp://ftp.internic.net/rfc/rfc3492.txt,   ftp:/
/ftp.math.utah.edu/pub/rfc/rfc3492.txt. Status: PROPOSED

STANDARD.

RFC3718

[2501] R. McGowan. RFC 3718: A summary of Unicode Consortium procedures, policies, stability, and public access, February 2004. URL ftp://

ftp.internic.net/rfc/rfc3718.txt,   ftp://ftp.math.utah.edu/

pub/rfc/rfc3718.txt. Status: INFORMATIONAL.

RFC4042

[2502] M. Crispin. RFC 4042: UTF-9 and UTF-18 efficient transformation formats of Unicode, April 2005. URL ftp://ftp.internic.net/

rfc/rfc4042.txt,   ftp://ftp.math.utah.edu/pub/rfc/rfc4042.

txt. Status: INFORMATIONAL.

Cerf:1998:TRI

[2503] Vinton G. Cerf. Tribute: I remember IANA. Communications of the ACM, 41(12):27–28, December 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://www.acm.org:80/

pubs/citations/journals/cacm/1998-41-12/p27-cerf/.

Garber:1998:ITU

[2504] Lee Garber. Industry trends: US will grant more high-tech visas; Internet pioneer Jon Postel dies. Computer, 31(12):15–16, December 1998. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL

http://dlib.computer.org/co/books/co1998/pdf/rz015.pdf.